A Bibliography of Publications in IEEE Micro

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 April 2024 Version 2.127

Title word cross-reference

#1 [Kah93i].

$1 [Ano17-58, Ano17-59]. 12 [MAT^{+}18]. 16 [ABG^{+}16]. 2 [DTH^{+}95]. 21/2 [Ste00a]. 28 [KBN16]. 3 [ASX19, Alt14e, Ano96o, AOYS95, BWMS19, CMAS11, DDG^{+}19, DFG^{+}13, Joh19c, LXB07, LX10, MKT^{+}13, MAS^{+}07, PMM15, PZB^{+}19, SYW^{+}14, SCSR93, VPV12, WLF^{+}08, ZSS^{+}19]. 60 [TKI^{+}14]. < [BMM15]. > [BMM15]. ^{2} [WHCK18]. ^{3} [KBW95]. ^{II} [BAH^{+}05]. Δ [MKG^{+}20]. k [Eng00j]. μ [AT93, Dia95c, KHS^{+}23, SDF^{+}23, TS95]. N [YW94]. x [And82a].

* [CCD^{+}82].

-3 [LKGL24]. -Core [MAT^{+}18]. -Cubes [YW94]. -D [ASX19, BWMS19, DDG^{+}19, Joh19c, PZB^{+}19, ZSS^{+}19]. -nm [ABG^{+}16, KBN16, TKI^{+}14].

0.18-Micron [HBd^{+}99]. 0.9-micron [Ano02d]. 000-fps [KII09]. 000-Processor [BSP^{+}17]. 024-Core [JJK^{+}11]. 036-TOp [SDF^{+}23]. 036-TOp/s/W [SDF^{+}23].

1 [Ano98s, BH15, Bre10, DtEt22, PFC^{+}02a, SDF^{+}23, Ste02a, Ste14a]. 1-GHz [Ano98s]. 1-terabits [MIM^{+}97]. 10 [KHS^{+}23, Loc03]. 10-Gigabit [Gad07, HcF04]. 10.7 [KHS^{+}23]. 10.7- [KHS^{+}23]. 100 [Kir84a, Pat84, PSW91, YSMH91, ZACM14]. 100-Mops [PSW91]. 1000 [ES84]. 11- [Lyl04]. 11/780 [Abr83]. 115 [JBF94]. 11FO4 [ASD^{+}05]. 12 [DTB01, Dur96, SS05]. 12-DSP [Dur96].

12.2 [SDF^{+}23]. 12.2-mW [SDF^{+}23]. 1284 [Dia94b]. 1284-1994 [Dia94b]. 13 [KW02]. 1394 [SB00]. 1394-1955 [Dia96d]. 14 [WD03]. 15 [FD04]. 15-Billion-Dollar [Gre07d]. 157-nm [Ano01h]. 16 [DD05, GDLT86, HM93, Sho85]. 16-Bit [SZH82, De 83, NN81a, NN81b, mDTG81]. 16-Core [FJL^{+}13, YMA^{+}13]. 16-Kbit [HM93]. 16-nm [FME18]. 16-Way [AK00]. 160 [RT92]. 17 [SS06]. 18 [Ano87d, KS07]. 18-GHz [Ano87d]. 19 [AM08]. 196 [CES^{+}11]. 1970s [Pre21]. 1984 [Jef84]. 1990s [Smi96b]. 1994 [Dia94b]. 1A [XLW^{+}12].

2 [Ano88c, Ano97-28, IKN^{+}99, JAS^{+}22, KSI^{+}96, Lee96, MS03, PFC^{+}02b, RMC04, Ste14b, SSB20]. 2-Petaflop [SB23]. 2.0 [Ano01c, Mat93b, VBC^{+}21]. 2.0-GHz [Ano01c]. 2.5 [Ano03c]. 2.72- [SDF^{+}23]. 2/ [LKGL24]. 20 [Ano88c]. 200 [IKNS88, NG87]. 200-MHz [NG87]. 2000 [Ano99-33, KY91, Mat98d, Mat00e]. 2008 [ET09]. 2011 [FV12, HGPT12]. 2012 [Bel13, FL13, Tor12]. 2013 [Goo14, Mar14, Sco14]. 2014 [Ano14r, Gre15c, KT14, Mud15]. 2015 [Ano14a, Ste16]. 2016 [Ano15b, JQ17, Mar17, Wei17]. 2017 [Ano16a, Ano16b, Ano17y, Bro17]. 2018 [Ano17m, Ano17b, Dwa18, Gon18]. 2019 [Ano19-29, Hil19]. 2020 [Ano20v, Ano20w, Ano20-35, Ano20-36, Ano20-37, Ano20-57, Jim21, Kul20, Kur20a]. 2023 [Ano23a, Ano23b]. 2059 [Ang90]. 21 [AW10]. 2100 [Roe86]. 21164 [ERPR95]. 21264 [Kes99]. 21364 [MBL^{+}02, WPM03]. 21st [Ano22-72, Ano22-73, Ano23-71, Ano23-72, Ano23-73, Ano23-74, Ano23-75, IJ98, Sak99b, Sak00d, Ano14-34, Ano22-71, Ano23-70, Emm07c]. 21st-Century [IJ98, Sak00d]. 22 [RE11]. 22-nm [SDF^{+}23]. 23 [BB12]. 24 [KZ13]. 25 [NN14]. 250 [HYM^{+}90]. 250-MIPS [HYM^{+}90]. 256-Bit [MMG^{+}99]. 256-Kbyte [ASD^{+}05]. 25nm [Ano03b]. 25th [Ano96p]. 26 [GH20, NS15]. 27 [GV21]. 28-nm [CCA^{+}19]. 286 [SKO89]. 286-Based [NC86]. 29 [Eec18d, Lev23]. 2nd [Del91b, Luu90b, Pat90].

3 [Ano03d, EBC22, HRK^{+}24, HWG^{+}09, IPL^{+}23, LKGL24]. 3-D [HRK^{+}24, IPL^{+}23]. 3.0 [Ano96g, Mat93a, SC24]. 3.06 [Ano03b]. 30 [KR19a, LS24, SB23]. 30-Teraflops [SB23]. 30-Teraflops/W [SB23]. 30-Year [Dia96a]. 300 [JBF94, KS90]. 300-MHz [JBF94]. 300-mm [Ano02c, HOHCV99]. 3171 [BSC^{+}90]. 32 [CHH^{+}98, KS90, Kur21a, RY21, RDJ^{+}13]. 32-Bit [BY07, Bor85a, CBLR86, GmDT83, Isa83, Kir83a, MKOK88, Mye83c, Mye84b, NG87, Smo88b, YSMH91, Bor85b, KS90]. 32-nm [RDJ^{+}13]. 32-Way [KAO05]. 33 [Joh22c, SS22]. 34 [DA23, Joh23d]. 360 [AB06]. 370 [Har21]. 376 [PK88]. 386 [Ano88c]. 386-20 [Ano88c]. 386-Monopoly [Sla91a]. 390 [SAC^{+}99]. 3D [Ano95b]. 3DNow [OFW99]. 3rd [Pea95]. 3T1D [LCWB08].

4 [Ano99x, Ano03b, Ano03d, DP97, Das21, GDES08, KSM99, PDT98, Pow94, Sha22, Spr02b]. 4-Bit [HYM^{+}90]. 4-Gbps [DP97, GDES08, PDT98]. 4.1 [Mat93c]. 40-nm [Man09]. 400 [DRM^{+}23]. 400-G [DRM^{+}23]. 4096-Core [ZSB21]. 46 [BCM^{+}14]. 488 [NS81]. 49 [Fan96].

5 [Ano98z, BHM^{+}00, HVS^{+}07]. 5-GHz [HVS^{+}07]. 5.1 [Mat93b]. 5.5 [Mat97c, Mat98e]. 500 [UAN^{+}93]. 5000 [RCC07]. 50th [Ano97l]. 511-Core [DXT^{+}18]. 520 [RHH^{+}03]. 520-MHz [RHH^{+}03]. 533 [Ano96k, Ano97-31]. 533-MHz [Ano96k, Ano97-31]. 56 [Ano97c]. 56-Kbps [Ano97c]. 5G [CHAF22].

6 [Mat93d]. 6.0 [Das21, MBJ08, SC24]. 6.1 [Mat97d]. 6000 [OB91]. 601 [PVYU94]. 604 [SDC94]. 60X [AAWC94]. 6300 [Han85, Mye85a]. 64 [Ano97w, Ano03d, BCC^{+}00, HMR^{+}00, KKL^{+}00, SCV01]. 64-Bit [AT93, BHM^{+}00, HL99, KM89, BBTV15]. 64-Core [DFG^{+}13, XCZ^{+}21]. 64.0 [Das21]. 64K [Mye83b]. 6800 [MM05]. 68HC05 [Ano97u]. 6M [RMC04]. 6T [LCWB08]. 6th [DKyL^{+}17, Kah91a]. 6th-Generation [DKyL^{+}17].

7-nm [XCZ^{+}21]. 780 [Abr83]. 796 [OL85].

80 [Ano88c]. 802.11b [Ano02c]. 802.16 [Ano02e]. 80386 [EAA85]. 8085-Based [CJ85]. 8086 [HF81]. 8088-based [Sho85]. 80960 [Rya88]. 82 [Mye82a]. 82460GX [DGMM00]. 855 [JC84]. 870 [BCC^{+}02]. 88000 [Mel89]. 88000-RISC [Mel89].

90-nm [Ano03c]. 9000 [SGC94]. 91 [Mye91b]. 920 [XCZ^{+}21]. 95 [Ano96t, Mat95d, Mat97d]. 97 [San97a]. 98 [Mat98d, Sca98]. 982S [SGC94]. 9th [Ste84a].

= [Ano87a].

A-Board [Alb09]. A-Changin [Mat08b]. A/V [GDES08]. A100 [CGG^{+}21]. AAI [Ste08a]. Abolishing [Hau88c]. Abstraction [NRS^{+}08]. Abstractions [BMM15, KAK^{+}22, MRJ^{+}15, Sol19]. Abuse [HCPS03, Kir01, Ste01e]. Abusing [MSS15]. AC [GA86]. Academia [Eec17b]. Academic [Gre14c, Yi23c, Gre97d]. Accelerate [ADJK20, FSK^{+}22, HSR18, TONH96]. Accelerated [BCF^{+}14, KMK^{+}19, KBN16, LP21, ML05]. Accelerates [DDHS00]. Accelerating [ABA^{+}21, ABC^{+}20, Bea20, Bha18, DtEt22, ESG^{+}05, GSLK11, HLS^{+}21, HKS16, KZS^{+}22, KRD^{+}20, KLM^{+}15, LPKP22, Lee95, Lee96, LLT^{+}08, LJM^{+}23, MSP^{+}19, OGLG^{+}22, OG24, PCC^{+}15, SFG^{+}22, SJFM19, SFP^{+}23, SMQP10, TSA^{+}22, Wal97]. Acceleration [AMFFM^{+}16, BFZ^{+}22, CKG^{+}09, DWLN20, EWW^{+}19, Eec18c, ESCB13, EP19, GDN^{+}17, GHY^{+}17, HLS^{+}21, KCXmWH17, KS18, KKL^{+}22, KAK^{+}22, MKG^{+}20, MAJ^{+}18, NBS^{+}18, NM24, RSC^{+}22, SAC^{+}21, SB23, SD21, WLF^{+}08, YZW^{+}23, Gre06a]. Accelerator [BGRKR88, BDV^{+}08, CDS^{+}15, CG95, DXT^{+}18, FM91, GTLY22, GCE^{+}21, GRD22, HKS16, HGS^{+}17, mHP18, Joh22b, JJK^{+}11, KSA^{+}19, KYG19, KJL^{+}10, KIR19, KSB21, MAK19, MKM15, MMG^{+}99, OYS^{+}11, PmWH08, PGW^{+}20, PZK^{+}18, QT21, SSY97, SDF^{+}23, SRWB15, SCH^{+}23, SsSMB24, VCS^{+}19, WFW^{+}21, WPO^{+}07, WWZ^{+}08, WLKN22, WHJ^{+}23, YRC^{+}22, ZRB^{+}22, Pri90]. Accelerators [ACA^{+}20, AW22, CES17, DSG^{+}22, GMC18, JLWL20, Joh19d, KJT^{+}11, KSK18, LHMH91, LS22, NM22, NGSW17, OYK^{+}17, PHB15, Ipe19]. Access [Ano02e, Ano14p, HKS16, KMK01, LH12, LTQZ07, SZZ01, WSZS05, ZZ02, Gre01e, Gre05f, Ste96b, Ste96f]. Access-Execute [HKS16]. Access-Interleaving [LTQZ07]. Access-Mode [ZZ02]. Accessed [YYK^{+}20]. Accesses [NAJE22]. Accessible [BMK^{+}21a, BMK^{+}21b]. ACCL [DWF^{+}21]. Accolade [Ste92d]. Accolades [Ste92d]. accommodate [SLM^{+}97]. Accounting [EE10, JGC^{+}11]. Accuracy [BTK^{+}23, KHS^{+}23, TSMS23]. Accurate [BdS98, Hin88, RPE10]. Achieve [Ano17f, Ano17e, Ano17c, Ano17d, LHN95]. Achieved [EM84]. Achieving [LLL^{+}16, MBK^{+}92, SIL^{+}15, SRA^{+}04]. ACM [Ano97l, Ano16c, Ano17g]. acquisition [Jae82a, Jae82b, Jae82c, Jae83, Tau84, Tau87]. Acrobat [Ano99x]. Acronym [War92d, Rob98b]. Act [Ano99a, Gre23d, Ste84b, Pit91, Ste91e, Ste07e].

Action [KZS^{+}22, Noy85]. Activation [CBJ10]. Activations [RAG19]. Active [Ano97s, CMAS11, DRB^{+}12, GGJ^{+}96, GD01, LDF^{+}13, Mye84a, Rob97a, WOM01, ZLTW13, ZHPR17, VBB95]. Activities [Kah91c, STL92]. Activity [Eng00h, RGH^{+}10]. ACTORS [BBE^{+}11]. Acts [Gre19b, Ste06b]. Acyclic [ED18]. Ad [Ano18a, Ano18c, Ano18l, Ano18k, Ano18u, Ano18-29, Ano18z, Ano18-27, Ano18-30, Ano18-28, Ano18-33, Ano18-32]. Adams [Far88b]. Adaptation [ZZ05]. Adapter [Edd02]. Adapting [Bos03b, Hal91]. Adaptive [DRM^{+}23, FAWR^{+}11, HL06, KJT^{+}11, KMPS06, QJP^{+}08, RCC12, RSE01, SMR20, TS91]. Adapts [CR95a]. ADAS [CPS^{+}18]. Add [FBHN04, Ste89b, Ste92c]. Add-on [Ste89b]. Add-Ons [Ste92c]. Adding [ENSD03]. Additional [Mye84b]. Additively [WHJ^{+}23]. Address [Bha17, Bha18, CD97a, CD97b, OG01, PHB15, RLS11, SDG^{+}21, WFA^{+}10, YKG18, Dv87, Mat95d]. Address-Correlated [WFA^{+}10]. Addressable [GGB^{+}15, MC92, PCW15, Rob92]. Addressing [AW03, Her93, SKM23]. Adds [Ano98g]. Adjusting [Gre18a]. Adjustments [Gre20d]. Administration [Gre21b, Gre17d]. Administrative [AKJF22, Ano18t]. Administrative-Distance-Based [AKJF22]. Adobe [Ano98z, Ano99x, Mat97c]. adopt [Gre99c]. Adopted [ST21]. Adoption [Ano98u]. ADSP [Roe86]. ADSP-2100 [Roe86]. Advance [Gre16b, Ste94f]. Advanced [BGRKR88, DQCL24, DG87, DG88, DG89, HOHCV99, Her93, KKL^{+}00, KM05, Mar21, Mis93, SYKM11, SF18, AHO^{+}90, BKM^{+}82, BT84, FMT91, Shl93, VS87, Ano97d, PJ91]. advanced-architecture [BKM^{+}82]. Advances [Ano17l, INKM05, IDI^{+}21, KOI95, Nic84, Ste98f, Ste08d, Ste08e, Web21, ZSS^{+}19, Mat01a]. Advancing [Ano00g, Eng00a, LNLG20, Sak99c, Far84]. Advantage [Ste02d]. Advantages [MKRC97]. Adversarial [NMF^{+}23]. Advert [Ano09a]. Advertisement [Ano13a, Ano13d, Ano13b, Ano13c, Ano13e, Ano13g, Ano13i, Ano14e, Ano14f, Ano14g, Ano14h, Ano14t, Ano14n, Ano14p, Ano14q, Ano14u, Ano14-27, Ano14-29, Ano14-30, Ano14-32, Ano14-33, Ano14-34, Ano14-38, Ano14-39, Ano15j, Ano15c, Ano15f, Ano15l, Ano15s, Ano15t, Ano15u, Ano15v, Ano15b, Ano15-31, Ano15-29, Ano15-30, Ano15-32, Ano15-40, Ano15-34, Ano15-39, Ano15-35, Ano15-36, Ano16f, Ano16d, Ano16e, Ano16i, Ano16j, Ano16p, Ano16q, Ano16s, Ano16v, Ano16t, Ano16w, Ano16y, Ano16-37, Ano16-38, Ano16-41, Ano16-48, Ano17f, Ano17k, Ano17l, Ano17n, Ano17q, Ano17u, Ano17v, Ano17-27, Ano17y, Ano17-28, Ano17-39, Ano17-44, Ano17-45, Ano17-48, Ano17-55, Ano14-28, Ano16h, Ano16u, Ano16-36, Ano16-40, Ano16-47, Ano17e]. Advice [Hil19, Mat07a]. Aérgia [DMMD11]. AFAN [GTF97]. Affair [Kir90b]. Affect [BGK97]. Affects [Ano00b, Eng00l]. Affiliate [Ano13d]. Affordable [CSC^{+}22, Bos05b]. After [Gre24a, McG82, Mil88a, Gre05a]. Aftermath [Gre19a]. Afterword [DM88b]. again [Ano00g, Gre06b]. Against [Hau88b, Kar88a, NMF^{+}23, OFKS23, PZL06, Ste85e, Ste86a, Ste88c, YFDV19, SWL11, Ste89f]. Age [DPY18, Gre09d, Mas21, Gre15c, Gre99b]. Aged [Ste21]. Agenda [Cha85b, Smi17]. Aggregation [AS91a]. Aggressive [Gre00b, KMPS06]. Agile [BCJ^{+}20, BC20, DTS20, GHS17, Joh20a, LWC^{+}16, PGW^{+}20, TGC^{+}20, WDK^{+}20, XYT^{+}23]. Aging [WKK^{+}14, YBNS15]. Agnostic [ECK^{+}22]. Agrees [Ste15b]. Agricultural [BCN^{+}22]. Agriculture [AKJF22, CSC^{+}22, CHAF22, FJB^{+}22, Joh22d, KGT22, LCN^{+}22, MK22, SMM^{+}22].

Ahead [Alb08, Ano98t, FH05, Gre10a, War90b, ZZ05, Yu96, Mat96b]. AI [AKJF22, Ano22a, FOP^{+}19, Gre19d, Gre23a, MD20, OLT^{+}23, PSL^{+}23, RCK^{+}21, VCS^{+}19]. AI-Enabled [AKJF22]. Aid [AJ83, Eng00m, LPL86, LMC^{+}83, Mye82d, HC83a, LDCS09]. AIDA [GTLY22]. Aided [De 94, Yao85]. aids [Mat97a, BM85]. aimed [Ano03d]. Air [Ano02e]. Aladdin [SRWB15]. Alamos [Ano98q]. Alder [RYR^{+}22]. Algebra [HGS^{+}17, Ipe19]. Algebraic [Ber81, Dun82]. Algol [AF82]. Algorithm [Den83, LHC^{+}12, LYP^{+}18, LHN95, RGF96, YZW^{+}23, RGF95]. Algorithm-Driven [LYP^{+}18]. Algorithmic [SS16]. Algorithms [AML05, Ano17l, KSA^{+}19, KMG^{+}03, KAK96, SG01b, SGP02, SZH82, TLYL04, Lou91, MIM^{+}97, TONH96]. Aliasing [SO02]. Aliasing-Free [SO02]. Alice [Ste14a, Ste14b]. Alignment [KYGW17]. Alive [Pre21, Rob00a]. All-Digital [SDF^{+}23]. All-Electronic [Ano97e]. All-Level [HNR10]. All-Programmable [FME18]. Alliance [Ano96m, Ano96s, Ano97t, Wil97]. Allocation [Gre09f, PPBS03]. Allreduce [LPKP22]. allure [Ano96n]. almost [DBDF97, Ste94f]. Alone [HABHW^{+}18]. along [Gre14d]. Alpha [Ano03e, ERPR95, Kes99, McL93, MBL^{+}02, WPM03]. Alpha-21264 [Kes99]. Alphorn [AGH^{+}91]. Also [Ano94c, Mat95d]. Alternative [ARS03, MK10, MSWP03, TC15]. alternatives [VS87]. AltiVec [DDHS00]. Always [BBC^{+}15, CDBY23, ZRB^{+}22]. Always-On [ZRB^{+}22, BBC^{+}15]. AM [KBW95]. AM29000 [Joh87, Man92]. Am9511 [FL84]. Amalgams [Mac93]. Amazon [Ste01a]. AMBA [Fly97]. ambitious [Ano97o]. AMD [Ano99g, Ano03c, Ano03d, BCF^{+}14, BFS12, BCD^{+}11, Chr96, CCS21, CH07, CKD^{+}10, DRM^{+}23, EBC22, HSNJ21, KMAC03, Mat21c, OFW99, OS08, Sla90a, SSB20]. AMD-K5 [Chr96]. America [Wet86]. AMORE [Dv87]. amplifiers [Jae82c]. Ampos [HS85]. amusement [NF81]. Anaheim [Far88a]. Analog [BTHS92, GGJ^{+}96, HGS^{+}17, IKK96, Kra96, LCS92, Lan96, MHW94, MY95, MCC^{+}94, Mur06, OW01, RJHK89, TCF96, YCD^{+}19, ZUNN18, ZRB^{+}22, ACRV96, CT95, Jae82a, Jae82b, Jae82c, Jae83, LC91, LKM92, RS90, SK97, VJ89, VTVM94, VVRV95]. Analog-Digital [LCS92, RS90]. Analog-to-digital [Jae82b]. Analysis [AMR^{+}06, ABC^{+}20, Ano15-35, Bos06c, CDBY23, CG95, CK11, EI87, EPZ02, HO99b, KTK13, KB91, LKGL24, MR85, MCN^{+}18, MKOK88, SG01a, SWM87, SB00, Ste90d, SMAS16, TATC09, WAA^{+}21, Yi21a, Yi22a, Yi22b, Yi22c, Yi22d, Yi23a, Yi23b, Yi24b, Yi24a, BM95, Dur96, GP90, Jae83, Mat98b, Mel87, RLG94]. Analytics [AAC^{+}23, Ano14-30, Ano16-40, Ano16-39, HLS^{+}21, OG24, OYK^{+}17, PAC^{+}14, SMT^{+}14, WBKR14]. Analyzability [UCS^{+}10]. Analyzer [HANR13, HK82]. Analyzing [CG95, JWS^{+}19, KAK96, SG00, Yi23c]. Anatomy [THT^{+}04, Gre05e]. ANDF [Dia94a]. Android [KLM^{+}15, QLLG15]. Andy [Gre06a]. Annals [Ano20d, Ano21j, Ano22j, Ano22k, Ano23m, Ano23n, Ano23o, Ano24e, Ano24f, Ano19f, Ano19d, Ano19e]. anniversary [Ano96p, Gre96f]. Announces [Ano99g]. Annoyances [Mat03b]. Annual [Ano96a, Ano97a, Ano98a, Ano99e, Ano00a, Ano01b, Ano02a, Ano03a, Ano04a, Ano05, Ano06, Ano07, Ano08, Ano09b, Eec16a, Ste90g, Ste90h]. anomalies [KWGG95]. Anomaly [PMS23]. ANSI [Rob97b, Ste99b]. Answers [Ste85a]. ante [Ano03b]. Anti [Ste96a]. Anti-knockoff [Ste96a]. ANTIC [You21]. anticipatory [HC83a]. Antikythera [Mor84]. Antitrust [Gre19b, Gre22f, Ste07a, Ste07c, Ste13, FS05, Ste05d, Ste06b]. Any [Ste08d, Ste08e, Ste94f]. Anyone [Wil95b]. AnySP [WSM^{+}10]. Anytime [WSM^{+}10]. Anyway [WSM^{+}10]. Anywhere [WSM^{+}10]. AOL [Ste97a]. Apache [Gre13d]. Apollo [Mat21a]. Appeals [Ste07c, Ste07e]. Apple [Gre22c, LS98b, Ste12, Ste17c, Ste17a, Ste17b, Ste18]. Appliances [SHTE08]. Application [Ano02c, Ano17l, CR95a, FMN^{+}13, GHSV^{+}11, HANR13, JL87, KLM^{+}15, Koe86, MBA^{+}09, NPC06, Vei04, Bos04e, PW96]. Application-Level [NPC06]. Application-Specific [JL87, Koe86, Vei04, Ano02c, Bos04e]. Applications [Ano00b, Ano00o, Ano10a, AAP^{+}10, BYM^{+}07, BBC^{+}15, BTK^{+}23, BSP^{+}17, CGS10, CDY^{+}18, DLR02, Del91a, ERM08, FBC87, FJB^{+}22, FSH^{+}01, GGC^{+}11, GR92, HSP^{+}01, HHNK09, IBM05, JUP^{+}22, KMN^{+}04, KIM^{+}09, LBD^{+}99, LLT^{+}08, LCP^{+}11, MLL^{+}15, MSP^{+}19, MAM^{+}06, Nic88, NL02, PNDG04, PSL^{+}23, PY87, QLLG15, Rea86, RKA^{+}20, Sak00b, SG00, SC91, SF18, SAC^{+}21, SKA^{+}14a, SWM^{+}20, SsSMB24, UCS^{+}10, VPV12, WLY^{+}21b, YRC^{+}22, vBK98, Ano03b, Cat88, CDGO97, DBDF97, Dia95d, Dia00, Eng00l, FN94, HS92, IKK96, Kah91e, MKRC97, PK88, Rob91, WCH94, Wv92, Yea96, Ano20g, Ano20e, Ano20f, Ano21k, Ano22l, Ano22m]. Applying [CMR97, DP97, HC83a, KSM99, STK88]. Appreciation [Mor84]. Approach [ASK^{+}15, AHK^{+}14, ASD^{+}05, BBE^{+}11, BBSG11, CL04, CML^{+}23, DMWS13, EPM^{+}20, EEKS07, Hil87, JGM^{+}20, KTI^{+}15, KDK^{+}89, KSK18, KCS^{+}20, LWC^{+}16, Mİ09, MM23, MW19, NL02, OHLR94, PFC^{+}02a, SPRK04, SRWB15, Sha22, SNM^{+}13, SMT^{+}14, TES^{+}18, TLM19, vBK98, Hur97, JKN96, Laz89, dG95]. Approaches [DG87, DG88, DGT89, DG89, Hig85, TM17, TM94b, TM94a, Ano95a, TCF96]. Appropriate [Ste89c, Ste89d, Ste89e, Ste89a, Ste90e]. approval [Wal97]. Approximate [AKAK^{+}18, AKK15, Eec18a, ESCB13, JS18a, KYG19, LZX^{+}18, LYP^{+}18, MLL^{+}18, MRJ^{+}15, PPBS03, PPP01, SJB09, THC18, ZUNN18]. Approximation [CYH^{+}18, DSG^{+}22, SMR18, SMR20]. Approximations [TM82, AB83]. apps [Ano96n]. APU [BFS12]. Aquabolt [KKL^{+}22]. Aquabolt-XL [KKL^{+}22]. AR/VR [YRC^{+}22]. Arbitration [Tau84]. Arc [Gre08c, Gre08d]. Arcane [Emm07a]. Archetypes [Gre22a]. ArchExplorer [DGR^{+}10]. Architecting [CLL^{+}20, DWF^{+}21, EEKS07, Gur09, MLM^{+}20, MBJ08]. Architects [Eec23, Mat09c]. Architectural [AW03, BB17, CGO00, CS18, FHP00, GmDT83, GEH^{+}23, HBE^{+}10, IO16, MWE^{+}03, NMHS15, PCDL10, SABR05, TA16, ZQL^{+}04, mDTG81]. Architecture [AAC^{+}23, AS91a, ABZ08, AC05, AFH16, Alb10a, AA93, Als90, AB06, AML05, Ano19g, AH96, BDSC21, BL23, BDH^{+}16, BHM^{+}00, Bro17, BG02, BML^{+}21, CM04, CB04, CGS10, CS15, CWL^{+}14, CLM08, CS08, CFRM04, CEM^{+}95, Cle00a, Cle00b, CAV^{+}14, CH07, CL87, DPY18, DOH94, DS94, DMG00, DKM^{+}92, DVWW05, DRM^{+}98, Dwa19, Eec15e, Eec15f, Eec16b, EAA85, ET09, EKMW02, FL13, FV12, FSK^{+}22, FG00, FCY^{+}20, GM21, GFL^{+}17, GE86, GKS^{+}05, Gon99, Gon06, Gon18, GHSV^{+}11, GA21, GR92, GHF^{+}06, GT22, Han96, HHNK09, HY98, HAWC^{+}11, HMR^{+}00, HF84, Hun87, Hyd00, IHCE07, IST^{+}11, Jag97, JQ17, JSY^{+}16, Jim21, Joh22f, Joh23f, JJK^{+}11, KYGW17, KND02, KMN^{+}04, KT14, Kim20, KBH^{+}04, KKL^{+}00, KIS^{+}00, LL03, LWB09, Lie23, LLW^{+}07, LNOM08, LWML16, MLL^{+}15, MBSP02, MS16]. Architecture [Mar21, Mar14, Mas21, May12, McL93, MCN^{+}18, Mey04, MS87, MCC^{+}94, Mud10, MCM^{+}16, MBL^{+}02, NMU^{+}15, NY22, OFW99, OS08, PPO^{+}04, PKP15, PW96, PFC^{+}02b, PSS^{+}91, RCJ^{+}10, RLV85, RNA^{+}12, RKA^{+}20, STKS17, Sak87b, SK01, SYW^{+}14, SNL^{+}03, Sch91a, SML04, SCS^{+}09, SY06, SLL^{+}18, Tab84, TM14, Tor12, TCC^{+}00, Tua99, Uss91, WZL20, War91c, War91d, WA11, WNW^{+}16, WOM01, Wen18, WGH^{+}07, WKP11, XLX^{+}23, Yeh07, Yi21a, Yi21c, Yi22a, Yi22b, Yi22c, Yi22d, Yi22f, Yi22e, Yi23a, Yi23b, Yi23c, Yi24b, Yi24a, ZES13, ZCW^{+}14, ZSB21, ZZ05, Ano03f, BKM^{+}82, Bos04d, Cat88, Chr96, FN86, Fur88, GDLT86, HF81, HMAF90, KY91, Kai88, KWM89, Kli81b, KWGG95, Lou91, OB91, Pri90, Rya88, SMHB91, SSH88, Sak99a, SPT^{+}92, TO96, VTVM94, BDH03, Dia94a, IG15, LE18, RMBK81, SRU^{+}23]. Architecture [VAD^{+}21]. Architectures [ASK^{+}15, AKAK^{+}18, ACA^{+}20, Ano17h, Ano17l, BNV^{+}15, BM19, CR95a, CF90, CEP^{+}17, DXT^{+}18, DG87, DG88, DG89, DPT^{+}21, Emm08b, FSBA12, Gre17c, Gro02, HFFA10, HGPT12, mHP18, IDI^{+}21, JUP^{+}22, Joh19e, Joh22a, KJL^{+}10, KBK03, KNB14, KC09, MLS^{+}16, MLL^{+}18, MD20, MRSV11, PmWH08, PW89, Rag84, RAA^{+}21, RYR^{+}22, RD90, Rüc02, SAR10, SSLV15, ST19, Sha23b, SKL^{+}92, SSF^{+}14, Sla90c, Sla91b, SMQP10, TS06, TLW^{+}10, TPV89, VCK^{+}13, VDC17, WG97, YRC^{+}22, ZIM^{+}07, DGW^{+}94, HDMT94, IKK96, Laz89, OFG88, Sak00a, Wv92, Luu90a]. Archival [BLC^{+}17]. Area [BF02, BCF^{+}95, CDY^{+}18, HSW98, Hor95, SK01]. Arena [Joh20c]. aren’t [Gre95d]. Argument [Ste09a]. Arguments [Mae87]. Argus [MBS08, SJFM19]. Arithmetic [CCG^{+}84, LJM^{+}23, Mur89, SK88, FL84]. arm [SM85, BBTV15, Har21, Jag97, RNLY23, SBB^{+}17, XCZ^{+}21, Goo21, Gri21, GBW^{+}23, PSB^{+}20, SSR21]. Arm-Based [SSR21]. ARM7100 [MKRC97].

ARM7TDMI [SCG95, Seg97]. ARM996HS [BY07]. Array [ABG^{+}16, BSP^{+}17, BDV^{+}08, But07, JLWL20, KCHA21a, KCHA21b, KHS^{+}23, Kra96, MBK^{+}92, OSS^{+}24, WHJ^{+}23, YNS^{+}14, DGW^{+}94]. Array-Based [Kra96, WHJ^{+}23]. Arrays [AB14, AHKY19, CSL^{+}06, GU98, OYS^{+}11, SFP^{+}23, Sti11, Lan87, MM96]. arrhythmias [CJFP95]. Arrived [Hau88a]. Arrives [Ano96i]. Arriving [Mye83c]. Art [Car98, Hal93, Hin88]. Article [Del93a, Ste96a]. Articles [Ano95a, Ano98d, Ano98e, Ano99b, Ano99c, Ano19k, Ano20s, Ano20t, Ano20-53, Ano21t, Ano21s, Ano21r, Ano23u, Ano23v, Hoo90d]. Artificial [AK24, Ano23a, Ano23b, BG02, FC22, GHRS89, HRK^{+}24, Joh22a, MK22, NHY^{+}22, Rüc02, SMM^{+}22]. Artificial-Intelligence-Based [HRK^{+}24]. Artificial-Intelligence-Enhanced [NHY^{+}22]. Artists [Alt12e]. Arts [Ste08e]. AsAP [BYM^{+}07]. ASIC [AO97, FBGB96, KGMT17, Man09, PKR92, Rob91, RS90, ZBH^{+}00]. Ask [Ste07a]. Asks [Ste08a]. AsmDB [NAA^{+}20]. AsP [Lea88]. aspect [Bos06e]. aspects [Ste89d]. Aspirations [MCF^{+}85]. Assembler [Smi86a, Smi86b, HP81, SL84a]. assemblers [Sko83]. Assembly [Bal84b, Bal84c, SHS85, Kah93d]. Asserting [SHKS19]. Assessing [CLMY96, KAK96, PP82]. Assessment [JLG19]. Assignment [Kah90a]. Assistance [SGL93, IKK96]. Assisted [CHAF22, CDBY23, KTC18, Mur06, SKC^{+}23]. associated [Gre97e]. Association [WHA89]. Associative [FM91, GTLY22, Gro92a, GR92, Gro92b, STS^{+}92, HS92, HM93, KBW95, SPT^{+}92]. Associative-Processors [Gro92a]. Associativity [YKG18, ZZY97]. Asymmetric [MMB^{+}08, SMQP10]. Asymmetry [Gre08e]. Asynchronous [BMG^{+}21, Lin04, SKLY97, XWZ09]. AT&T [FGG^{+}88, Gre00d, HSW^{+}89, Mye85a]. At-Memory [SB23]. Atari [You21]. Athlon [Ano99g, Ano03d]. ATLAS [KPV^{+}99]. ATM [KPV^{+}99, Vic93, VBB95]. Atmospheric [GFL^{+}17]. Atom [BvdGM^{+}15, STT^{+}15, LDCS09]. Atom-Aid [LDCS09]. Atom-Switch [STT^{+}15]. Atomic [Ano92a]. Atomicity [LTQZ07, LDCS09, NRS^{+}08]. Atomristors [Aki18]. Attached [RCBL00, Mon97]. Attack [CDBY23, Gre19a, Ano95b]. Attacking [Mat04a, RNLY23]. Attacks [KPN^{+}20, LWML16, NMF^{+}23, PZL06, SYG^{+}20, WPH^{+}23, YFDV19]. Attaining [CMAS11]. Attestation [ZL16]. Attribute [AAC^{+}16]. Attribute-Based [AAC^{+}16]. Audio [Sav99b]. Auditory [LWK94]. Augmented [KKP^{+}14, SJO01, SsSMB24]. Augmented/Virtual [SsSMB24]. August [Ano95a, Buc85]. Austin [Far87]. Authenticating [RCBL00]. Authentication [RNLY23, ZG96]. Author [Ano97a, Ano98b, Ano98c, Ano98a, Ano00a, Ano01b, Ste98a, Ano96a]. Authority [Rob99c]. Autocuer [Mye83a]. Automata [ATS^{+}22, PVS17]. automate [CMR97, TCF96]. Automated [PBFC21, PRE11, SS16, WHCK18, Kah93d]. Automatic [DGR^{+}10, EPM^{+}20, HIP^{+}22, Joh22b, LPC12, ONS^{+}23, RCA07, SL84a, TLM19]. Automatically [AAW^{+}96]. Automating [CWS^{+}12, KJP^{+}13]. Automation [Bor99a]. AutoML [CLL^{+}20]. Automobile [SV03]. Automotive [AKT^{+}18, Eec18b, Fre02, Koo02, KTC18, LC18, MKAC18, SF18, Van21, vBK98, HDMT94, ZP93]. Autonomous [Gre18a, IEB^{+}14, KTI^{+}15, KSLY17, LCN^{+}22, SNM^{+}22, SSK23, WHP^{+}13, IKK96]. Autovectorization [AS22]. AV [SANK98]. Availability [ERM08, Qua00, JRHM86]. Available [KSR^{+}99, Ond96]. Avena [BCN^{+}22].

Avenues [INKM05]. AVIO [LTQZ07]. Avoiding [Lei98, Mac98]. Award [Ano15f, Ano15-40, Ano16f, Ano16s, Ano16t, Ano16r, Ano17w, Ano17y, Ano17x, Ano17b, Ano18c, Ano18l, Ano19l, Ano19m, Del93a, KT14, LE18, Mar17, MBTS16, Sco14, Ano01d, Ano14o, Ano14a, Ano15b, Ano16c, Ano16a, Ano16b, Ano17g, Ano17-27, Ano17-58, Ano17-59, Ano18a, Ano18k, Ano18-29, Ano21v, Ano22a, Ano22-40, Ano23-83, Ano23-84, Bel12, Bel13, Bro17, Dwa18, Gon18, Goo14, Hil19, KT14, Mar14, Mud15, Ste16, Tor12, Wei17]. awarded [Ano99q, Ano99u]. Awards [Ano17-45, Ano17-58, Ano17-59, Ano19n, Ano21-60, Ano23t, Ano23-58, Bel13, Eng00j, Ano17k, Ano17j, Ano21-58, Ano21-59, Ano24v, MB15]. Aware [ACG03, AS05, Alt12d, BMS16, BAM03, BBS^{+}00, CWL^{+}14, CHSL17, DK14, HAWC^{+}11, JGC^{+}11, KKL^{+}09, KKP^{+}14, MNU^{+}15, MM23, MM09, Red13, RLS11, SPKJ06, SSH^{+}03, SKC^{+}23, TSS18, TSA^{+}22, TLM19, TK21, WB12, ZLBI06]. Awareness [MT05]. AXDIMM [KKL^{+}22, KZS^{+}22]. Axilog [MRJ^{+}15]. AXP [McL93]. Azure [Sti19]. AzusA [AK00].

B [Ano16a, Bel12, Bel13, Mye85b]. Babbage [Ano21v, Ano18a]. Babel [War92d]. BabelFish [SDG^{+}21]. Back [CCS21, DDG^{+}19, Joh21c, Mat04b, Mat07c, Ano19-39, Bos06b, Ste93c, Ste04d]. Back-End-of-Line [DDG^{+}19]. Backbone [Ano99n, Gre03d, Ano99p]. Background [Ste86c]. Backlog [Mat95a]. Backplane [All81, Smo88b, War90a]. Backpressure [KPV^{+}99]. backward [Mat96f]. Bad [Ste88e, Rob00e]. Baking [Gre14a]. Balance [TGF88]. Balanced [ACKM05, BPUH06]. Balancing [AKJF22, KSV^{+}21]. ball [LGJ95]. Band [Ano99a]. Bandwagon [Ano97-32]. Bandwagons [Gre03b]. Bandwidth [BPUH06, BGK97, Das21, OMMB13, PPBS03, SGK^{+}04, TIT^{+}13, WAA^{+}20, Gal97]. Bank [Gre23b, Ste14a, Ste14b]. Banking [Gre99b, LLSS05]. Barriers [DGM^{+}11]. Baseband [BDV^{+}08, FMN^{+}13]. Based [AKJF22, ANC05, AAC^{+}16, ACA^{+}20, Ano16-41, ABC99, AF84, BFZ^{+}22, BEL^{+}23, BLC^{+}17, CL04, Cas95, CPH90, CS08, CJ85, CL87, DMP91, EI87, FMV85, FSH^{+}01, GDN^{+}17, GBW^{+}23, HRK^{+}24, HK82, Har12, HMS^{+}86, HL06, IPL^{+}23, IEB^{+}14, JWS^{+}19, JS18b, Joh84, KG05, KKD^{+}07, KGDW^{+}13, KSE^{+}22, KPHP04, Kra96, KGT22, LNLG20, LS22, LHL09, LSZ82, LMC^{+}83, MR85, MSP^{+}19, MKM15, MM23, MSY^{+}22, MSB^{+}17, Mor88, MAJ^{+}18, MS83, Mye83a, NJZL^{+}17, NMF^{+}23, NC86, NDR^{+}22, NL02, PMM15, PZL06, PMS23, PC01, RBB21, SVA^{+}22, SML04, SAC^{+}21, SFP^{+}23, SWM^{+}20, SSR21, SS16, Tal93, TCF96, WM85, WK13, WLD15, WZL20, WMSH09, WOM01, WHJ^{+}23, XCZ^{+}21, XPZ^{+}19, XLX^{+}23, YKH^{+}19, YZW^{+}23, ZLBI06, ZMVH^{+}83c, ZVHL85, Ano03b, Hsi91, KKT^{+}91, LLLL09, ME95, MST^{+}85, NF81, Sak99a, Sho85, SM85, SU95]. based [ZMVH^{+}83a, ZMVH^{+}83b, GK97, Mel87, RMFG85]. basic [Jag97, KHW85, KHF86, KW83, SB84]. BASIC-DINT [KHW85]. Basics [Spr02a, War89a]. Batch [HOHCV99, MM09]. Battles [Ano97v]. Bazaar [Mat99a]. BB [Ste96f]. Be [Ano15u, Ano16v, Ano16w, Ano16u, Ano17z, Joh20b, Joh22b, Ste83d, Ste86a, Ste92b, Ste13, Mat95d, Mat06d, Sak99a, Sak00e, Ste83c, Ste96f, Ste98b]. beam [LGJ95]. beam-and-ball [LGJ95]. Beards [Del94b]. Bearings [YW88]. Beats [SRJ^{+}91]. Becomes [Ano96r, Raj94]. Becoming [Gre05f]. Beer [Gre18b]. Before [Gre02e, Joh21a]. Beginnings [Bos03d, Sak01c]. Begun [Eng00f]. Behavior [Ano98j, Yi21a, Yi22a, Yi22b, Yi22c, Yi22d, Yi23a, Yi23b, Yi24b, Yi24a, Dan89]. Behavioral [Ano15-35, Gre15a]. Behaviors [RNN^{+}16]. Behind [Gre08c, Gre08d, Gre15a, RCK^{+}21, Sti19]. Being [Mat10a, Ste97b]. Bélády [SJL23]. Bell [Mye84d]. Benchmark [Ano97-28, Ano00h, CBLR86, GHPS93, JC08a, MRC^{+}20, PCLGO09, Pri89, TLYL04, AAW^{+}96, Ano01g, Ano03b, Eng00l]. Benchmarking [SsSMB24, Hin88]. Benchmarks [Far86, JC08b]. Bending [Ano97n]. benefits [Ano00g, Eng00j]. bent [Eng00g]. Berkeley [CFK^{+}10, Pri93a]. Berne [Hau88c, Ste88e]. Bespoke [CDY^{+}18]. Best [Ano89, Del93a, Han87, RY21, SMM^{+}22, TM82, CH94, Emm06e, Ano17-30, Ano19-32]. bets [Wea97a]. Better [AML05, Ano16-33, Ano16-34, Ano17-43, Ano17-40, Ano17-44, Ano17-41, Ano17-42, FSR^{+}05]. Between [SKM23, Yi23a, Yi23b, Das17]. beware [Ste97c]. Beyond [Alt11a, BY17, Eec18b, Lee24a, LCP^{+}11, Loc03]. Bias [KOKA23]. Bidirectional [IGH^{+}99]. Big [Alt11a, Alt14a, Ano14-30, Ano16y, Ano16-36, Ano16-37, Ano16-38, Ano19i, Ano21m, Ano22q, Ano22r, Ano22s, FG14, Gre12a, Gus85, HRSS11, KS11, Mat13a, WBKR14]. Big-Endian [Gus85]. Bigger [CDBY23]. Bill [Ano03d, Gre99c, Gre08c, Gre08d]. Billing [JGC^{+}11]. Billion [Gre07d, LLL^{+}16, RYK18]. billions [Kra96]. Billionth [Ano97u]. Bilski [Ste09a]. Binary [CHH^{+}98, Mae87, MST^{+}85, PBFC21, PO04, ZMVH^{+}83c, ZMVH^{+}83a, ZMVH^{+}83b, ZVHL85]. Binary-Decision-Based [ZMVH^{+}83c, ZVHL85, MST^{+}85, ZMVH^{+}83a, ZMVH^{+}83b]. Bioengineering [Del91a]. Bioimplantable [JC08b]. Biological [Kur20b, GGJ^{+}96]. Biology [Bha20, VN96, Gre97a]. Biology-lnspired [VN96]. Biomechanical [Ano00b, Eng00l]. Birth [Fag21, Goo21]. Bit [AT93, BFZ^{+}22, BY07, BHM^{+}00, Bor85a, CBLR86, EWW^{+}19, GmDT83, HIP^{+}22, HYM^{+}90, HL99, Isa83, Kir83a, KM89, LSZ82, MMG^{+}99, MKOK88, Mye83c, Mye84b, NG87, Sla89, Smo88b, SZH82, TS06, YSMH91, ZLTW13, BBTV15, Bor85b, De 83, NN81a, NN81b, mDTG81, KS90]. Bit-Error-Rate [ZLTW13]. Bit-Serial [EWW^{+}19]. Bit-Split [TS06]. Bitcoin [BH15]. Bitflip [TSMS23]. Bits [Ano00f, Ano01e, Ano01h, Ano02b, Ano02c, Ano02d, Ano02e, Ano03e, Mar21]. bitwise [OZT^{+}22]. Black [OFKS23, Hin88]. Black-Box [OFKS23]. Blackford [RCC07]. BlackParrot [PGW^{+}20]. blame [Gre96c]. Blatant [Ste13]. Bleeding [Gre10d, RDC98]. Bleeding-Edge [Gre10d]. Block [BCC^{+}02, HFFA10, HSN^{+}23, KTK13, CG95]. Blocking [MMESG^{+}20]. Blocks [JLWL20, PSB^{+}20, ZWB19, RJHK89, VVRV95]. Blood [Alt11d]. Bloom [DKSL04, Gre06e, LK10]. Blue [CEH^{+}12, HOF^{+}12, hHH99, SWG06]. BlueField [JAS^{+}22, LKGL24]. BlueField-2 [JAS^{+}22, LKGL24]. BlueField-2/-3 [LKGL24]. Blueprint [MCV^{+}19]. Bluetooth [Eng00j]. Blurring [Das17]. Board [Alb09, Ano13g, Ano14u, Ano18u, Ano20-52, Ano21-53, Ano21-54, Ano21-55, Ano21-56, Ano21-57, Ano22-58, Ano22-59, Ano22-60, Ano22-61, Ano22-62, Ano22-63, Ano23-56, Ano23-57, Eec16e]. Bobcat [BCD^{+}11]. Body [KOKA23]. Bolt [DK18]. Bolts [Mat03d]. Boltzmann [BI17]. Bombs [Ste94a]. Bonds [MBG^{+}16]. Bonus [Gre07d]. book [Ano94c, Gre15c]. Books [Mat96a, Mat02d, Mat03e, Ano97o, Ano97q, Mat06c, Mat13c, Mat14]. Bookstore [Ano96e]. Boolean [YW94]. boom [Kan95]. Booms [Ano01a, Gre01c]. Boost [PAC^{+}14]. Booster [WHJ^{+}23]. Boosting [CJFP95, PDS^{+}13, WHJ^{+}23]. Borg [Ste98c]. Born [RJ91]. both [Gre05b]. bottleneck [DP97, Joh90b]. Bottlenecks [Lyl04]. Boundaries [NCT^{+}98]. Bounded [ATS^{+}22]. Bounding [SGK^{+}04]. Bountiful [War91a]. BoW [AFK^{+}21, FKV20]. Box [Gre13e, OFKS23]. Boxes [Eng00o]. Braille [GRS86]. Brain [KSV^{+}21, SKW^{+}23, KSLY17]. Brainwave [CFO^{+}18, FOP^{+}19]. Branch [FSR^{+}05, HCP^{+}03, MBG^{+}16, SSA16, UAN^{+}93]. Branches [KMPS06]. Breadth [Gre22d]. Breaking [Bha18, VMW^{+}19]. Breakthrough [Eng00c, Mil88a, Ano02c]. Bricks [Gre19f]. bridge [WBC^{+}95, PKB^{+}15, RNA^{+}12]. bridge/memory [WBC^{+}95]. Bridging [ACZ^{+}22, BcFP06, RSW10]. Brief [Bar21, Lee90]. briefly [Bos06b]. Briefs [Ste09a]. Bright [Bos03d]. Bringing [Ano99y, BCH^{+}23, PAM^{+}07]. Brings [Buc85]. Broad [Ano99a, Ano99t, WLF^{+}08]. Broad-Band [Ano99a]. Broad-Purpose [WLF^{+}08]. Broadband [Ano02e, Gre07d, Gre07b, Gre09b, Gre10a, Gre11b, Sam00]. Broadcast [ASK^{+}15]. Broadcast-Enabled [ASK^{+}15]. Broadcasting [SYKM11]. Broadcom [SP09]. BROOM [CCA^{+}19]. browser [Ste97d]. Browsing [ZHR15]. BTRON [SKO89, Sak87c, STK88]. BTRON/286 [SKO89]. Bubble [MTS^{+}12]. Bubble-Up [MTS^{+}12]. bubbles [Ano03e]. Buckley [Dia96a]. Buffer [KKL^{+}22, NS05, PZL06]. Buffers [ARS03, UAN^{+}93]. BugNet [NPC06]. Build [Ano13a, GSS09, MCR17, GGJ^{+}96, MIM^{+}97]. Building [Ano97f, Ano00g, BJO^{+}09, BCC^{+}02, Dia93c, Eec15a, Gre99d, Gre09b, Gre10a, JP17, LWC^{+}16, NL02, PSB^{+}20, Sak01e, SP09, SB07, ULS^{+}00, WMH^{+}10, ZWB19, CG95, Hal93, RJHK89, VVRV95]. building-block [CG95]. Built [KPP06, LHL09, MD20, NRV^{+}06]. Built-in [LHL09, NRV^{+}06]. Bulk [NRV^{+}06, OZT^{+}22]. bulk-bitwise [OZT^{+}22]. Bulldozer [BBSG11]. Bunch [AFK^{+}21, FKV20]. Bunch-of-Wires [FKV20]. Buried [STT^{+}15]. Burning [Bos04f]. Bursty [WSZS05]. Bus [AAWC94, All81, All86a, All86b, Ano84, Bor85a, Bor81, CJ85, FO89, Gil82, KKD^{+}07, Kir83b, Kir84a, Kir88b, NS81, PLK^{+}16, Pat84, Pri86, STK88, Smo88b, SB00, Ste86g, TRY^{+}09, Tau86, War91d, Bal84a, Dia95d, DM86, ES84, Fis85, OL85, SSH88, Dia96d]. Bus-Based [KKD^{+}07]. Buses [Gus84, Jam90, Kir90d, Bor85b]. Business [Gre14c, Sak87c, Ste14a, Ste14b, Gre00b, Gre01b, Sla96, Ste96c]. Business-Method [Ste14a, Ste14b]. Business-Oriented [Sak87c]. Busts [Ano01a, Gre01c]. Busy [War90b]. Butterfly [Gre23c]. Buzz [Gre15a, San97a]. Bye [Alt14b]. BYOD [DMG^{+}15]. Byte [Gus85, PCW15, Per83, Sho85]. Byte-Addressable [PCW15]. Byte-Wise [Per83, Sho85].

C [Ano92c, AH96, Mat96f, Ste91a]. C2000 [BvdGM^{+}15]. C400 [SMHB91]. C6201 [JLSM03]. cable [War91g]. Cache [AF88, BRmWH06, BJW^{+}23, BK14, CL05, cCCP00, CKD^{+}10, EWW^{+}19, ERM08, EKMW02, GHPS93, HFFA10, HNR10, HBCS04, IDI^{+}21, KK10, KMK01, KBK03, KTY24, LCWB08, LWML16, NS05, NM24, ONS^{+}23, Pre91, ROA13, RMC04, SK12, SW14, SSF^{+}14, SKJ^{+}11, SLB04a, SLB04b, TNT06, TM17, TM94b, TM94a, WGA^{+}09, XPZ^{+}19, YFDV19, ZZY97, ZZ02, HMAF90]. Cache-Conscious [ROA13]. Cache-Level [TNT06]. Cache-Miss [BRmWH06]. Cache-Only [EKMW02]. Cached [SZZ01]. Caches [CD97a, CD97b, Dog12, HKC10, JMZ^{+}11, KBK03, LH12, MBJ08, SLSO14, TSFS21, VJFG17]. Cachet [KTY24]. Caching [QJP^{+}08]. CACTI [MBJ08]. CAD [Ano92b, MM96, Sto90]. Calculating [de 84]. Calculation [ZUNN18, Sho85].

Calculations [Per83]. Calculus [PFC^{+}02b]. Calendar [Ano97b, Ano98i, Ano99d, Ano00e]. Calisto [NIJ^{+}03]. Call [Ano95a, Ano98d, Ano98e, Ano99b, Ano99c, Ano00c, Ano09c, Ano10a, Ano10b, Ano14b, Ano14c, Ano15c, Ano15d, Ano15e, Ano15f, Ano15t, Ano16a, Ano16d, Ano16e, Ano16q, Ano16s, Ano16t, Ano16r, Ano17i, Ano17k, Ano17j, Ano17l, Ano17v, Ano17w, Ano17y, Ano17x, Ano19l, Ano19m, Ano19n, Ano19k, Ano19o, Ano19p, Ano20u, Ano20s, Ano20t, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-53, Ano21t, Ano21u, Ano21s, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21r, Ano22a, Ano22b, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22-47, Ano22-48, Ano22-49, Ano22-50, Ano22-51, Ano22-52, Ano23u, Ano23v, Ano23t, Ano23-27, Ano23w, Ano23y, Ano23z, Ano23x, Ano23-38, Ano24n, AGH^{+}91, Gre96b]. Calm [Gre12b]. CAM [KYGW17, Liu02]. Camera [Ano98y, Ano99t, Fos98, SYKM11]. Camera-on-a-Chip [Ano99t]. Cameras [APS98, Kaw98]. Camp [Hal93]. CAMs [PS03]. Can [Ano96n, CB10, Gre97a, KSE^{+}22, SS82, Ste83c, Ste83d, Ste86a, Ste92b, Mat95d, MIM^{+}97, SLM^{+}97, Ste94f, FPAF02, Fre02]. Canaries [Gre04a]. cancer [Ano01c]. Cap [MAT85]. Capabilities [SIL^{+}15]. Capability [CL87, WNW^{+}16]. Capability-Based [CL87]. Capability-System [WNW^{+}16]. Capacitance [KPN^{+}20]. Capacitive [HC84]. Capacitively [KKP^{+}09]. Capacity [BJW^{+}23, HRC^{+}23, WGA^{+}09, Boa96, Hsi91]. Capping [RCC12]. CAPRA [GR92]. Caps [Sha82]. Captain [War91e]. Capturing [Kaw98]. Car [Hoe93]. Carbon [Ano98-32, Ger19, GKL^{+}22, GEH^{+}23]. Card [DVQ96, DF01, Mye89b, Sha82, Ano00m, Eng00l]. Cards [Ano96q, NM96, NFQ03, Sak01f, SIPM02, TBDL01, Tua99]. care [Alb07e]. Career [Ano13a, Ano15v, Ano16x, Ano19-30, Ano19-31, Ano20y, Ano20-58, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano24o, Ano24p, Ano17f, Ano17e, Ano17c, Ano17d]. careers [Ano97p]. Carnegie [Ano22z]. Carrizo [KBN16]. Cartridge [SCV01]. Cascade [AFK^{+}19]. Cascading [MC92]. Case [AB14, ACP95, Bea20, FAK^{+}14, HGS^{+}17, Jac03, MK10, PAC^{+}97, Sen86, SJB09, Ste86d, Ste87c, Ste89b, Ste90c, Ste90d, Ste91f, Ste91g, Ste07d, Ste08b, Ste09a, BSB^{+}92, Gre96a, Ste91h]. Case-Study [Sen86]. Catches [HSNJ21]. Catching [San97a]. Cathedral [Mat99a]. CDs [Ano96d]. ceiling [Gal97]. Celebrate [JN21a, Gre96f]. Celebrating [Ano96p, Dia96b, WG97]. Celebrities [Mat12a]. Celerity [DXT^{+}18]. Cell [ASD^{+}05, GXMZ13, MAS^{+}05, STM02, SCC^{+}05, Ste85d, Ste17c, Ste17a, Ste17b, Ste18, SV21, Ano01f, Lan87, TCD^{+}05, AP07, Ano02b, GHF^{+}06, KPP06]. Cellular [JL87]. Center [AS10, Ano15u, Ano16v, Ano16w, Ano16u, Ano17z, Ano23-39, Ano23-41, Ano23-42, Ano24o, GHLK^{+}12, OESGG^{+}21, VAFF^{+}10, Ano23-40, Ano24p]. Centers [GKL^{+}14, RC12, RTM^{+}10, SSK23, DK14, FDS^{+}17, RSW10]. Centip3De [DFG^{+}13]. Centipedes [Rob01a]. central [MIM^{+}97]. Centric [KJL^{+}10, KSK18, KCS^{+}20, RC12, Sha23a, WWR97]. Century [Ano22-72, Ano22-73, Ano23-71, Ano23-72, Ano23-73, Ano23-74, Ano23-75, IJ98, Sak99b, Sak00d, Ano14-34, Ano22-71, Ano23-70, Emm07c]. Cerebras [Lau21, Lie23]. Cerebrum [KHS^{+}23]. Certificate [Ano98p]. Certification [Ano13b]. CGPA [RAG19]. CGRAs [KOKA23]. Chain [BF02, MGP21, Gre05d]. chains [Ano02d]. Challenge [HSW^{+}89, Hur98, MC90, Sak02e, Sak01b, Ste04d]. Challenged [Gro83, Hec83a]. Challenges [AC05, AKT^{+}18, BCP04, Bor99b, Bor05, BCA99, Bos03a, Bos03b, Bos04f, BBS^{+}00, Can98, Con03, ESW97, Her93, IO16, KAC^{+}95, LCN^{+}22, MH10, Mye91a, OML^{+}07, ODH^{+}07, Pen90, SSH^{+}03, Sta01a, Sta01b, Won03, Bos04d, Bos05d, Qur19]. Challenging [Ste02a, Ste04a, Ste04b]. Champion’s [Ste06a]. Chance [Gol21]. Change [CEAY23, Gre99a, Gre18c, Gre22f, Hil87, LZY^{+}10, SWL11, SAW^{+}10, Ste93d]. Changes [Alb08, Mat99a]. Changin [Mat08b]. Changing [Cla03, Dan89]. Channel [CDBY23, DMWS13, ED18, Edd02, Gil96b, GK97, KPN^{+}20, LWML16, Sco96, WPH^{+}23, YFDV19]. Channels [CEAY23, KKP^{+}09, KPKJ08, VCD16]. Chapter [Gre10f]. Characteristics [Yi21a, Yi22a, Yi22b, Yi22c, Yi22d, Yi23a, Yi23b, Yi24b, Yi24a]. Characterization [AJC^{+}20, HE07, JLSM03, KC09, PRE11, PCLGO09, Bos06e]. Characterizing [AP07, CEAY23, JC08b, SSC^{+}22, WLY^{+}21a]. Characters [TM81]. Charge [LDL17]. Charles [Ano99q, Ano21v, BKP12]. ChARM [PGL97]. Chasing [GKL^{+}22]. Chatterji [Gre23d]. Cheap [Gre07e]. Cheaper [Eng00p]. Check [Ano01a, Del93b, PV01, SC24]. CheckMate [TLM19]. Checkpoint [ARS03]. Checkpointing [TNT06]. cheerful [Ste93d]. chemists [Ano02b]. CHERI [GBW^{+}23, WNW^{+}16]. CHERI-Based [GBW^{+}23]. Chess [hHH99]. Chicken [Gre08a]. Chief [PC93, Alb07d, Alb07e, Alb07b, Alb07a, Alb07c, Alb08, Alb09, Alb10a, Alt11a, Alt11b, Alt13c, Ano10a, Ano23-76, Ano24y, Bos03b, Bos04b, Bos06e, Bos06d, Bos06c, Bos06a, Bos06b, Dia95c, Dia98, Sak99b, Sak99a, Sak99e, Sak99d, Sak99c, Sak00c, Sak00b, Sak00a]. Children [Dia99]. Chili [YT01]. China [Ano96b, Kah93f]. Chip [AB14, ABG^{+}16, AMK17, AP07, ASX19, Ano89, Ano98-44, Ano99s, Ano99t, Ano01h, Ano03e, AOYS95, BF02, Bel96, Bos03a, Bos04c, Bos06d, BCF^{+}14, BWBJ11, Can98, Cla03, CMAS11, DGMM00, DMMD11, DRM^{+}23, Dav98, DSL^{+}18, DtEt22, DPBW19, EMYN00, EGL^{+}90a, Edw83, Eng00c, Eng00b, Eng00j, Eng00o, FBGB96, FAWR^{+}11, Fly97, Fos98, FH00, Gol96, GSVP03, GKS^{+}07, HOF^{+}12, IKN^{+}99, Joh20c, JJK^{+}11, Kah92c, Kah93i, KST04, KML04, KRD^{+}20, KBK03, KKP^{+}14, KKD^{+}07, KPV^{+}99, KTC18, KHS^{+}23, KCKP14, KPKJ08, KP07, LBD^{+}99, LHC^{+}02, Lin98, MY95, Mye83c, Mye92b, NIJ^{+}03, NCT^{+}98, OMMB13, OKN^{+}11, ODH^{+}07, PKP15, PC93, RTHA05, RGR95, SC91, SO14, SGG^{+}12, SPKJ06, SKM^{+}16, Ste85b, Ste07d, SsSMB24, TUI^{+}01, TSW^{+}01, Trö98, UBH^{+}94, WZL20, WGO^{+}14, WA11, WWZ^{+}08, WGH^{+}07, WDK^{+}20, Ano99v, Ano01c, Ano02c, Ano02d]. chip [Ano03c, Ano03d, DVQ96, FN86, HMAF90, KWM89, KSI^{+}96, LKM92, Mon97, Ste91h, TO96, IHCE07, Kul20, Kur20a, Lav02, Ste07e]. Chip- [Bos03a]. chip-layout [Ste91h]. Chip-level [Bos04c]. Chip-Package [Can98, Lin98, Trö98]. CHIPKIT [WDK^{+}20]. Chiplet [AFK^{+}21, Sha23c, WAA^{+}20, XCZ^{+}21, ZSB21]. Chiplet-Based [XCZ^{+}21]. Chips [AS95, Alt11a, Alt11d, Alt13a, Alt13d, Alt14b, Alt14c, AM08, AR16a, AR16b, Ano87a, Ano92b, Ano00i, Ano14e, Ano15j, Ano17o, AW10, BS98, BB12, CM17, DTB01, DD05, DXT^{+}18, DA23, DM88b, DM88a, Eec15c, Eec16a, Eec17a, Eec17b, Eec18d, EW23, EW24, Eng00p, FD04, For02, HW91, hHH99, HRSS11, IA11, IA13, IA22, Joh19c, Joh19a, Joh19b, Joh20c, Joh22c, Joh22e, Joh23d, Joh90b, KS11, KND02, KKS^{+}98, KZ13, KB20, KW02, KS07, KR19a, Kur21a, LNK94, Lee24b, LHL09, Mar21, Mas93, Mat97b, May12, MKAC18, MD88, NN14, NS15, Nak99, Nak00, NPY^{+}21, OYS^{+}11, PVS^{+}11, RY21, RE11, RC13, SS22, SS06, SKA^{+}14a, Ste86b, Ste86c, Ste86d, Ste90g, Ste90h, VBB14, WD03, WG97, Alb07e, Ano01h, JA96, Pri94b, Alt11c, Hoo90b, IA09, Joh19c, Jou92]. Chips [KvdW09]. Chips-III [Jou92]. Chipset [GDES08, RCC07]. Chipyard [ABG^{+}20]. Choice [Ste85f, ZV85, ZVH85]. Choices [Mye89a, SL97]. Choose [Ano16x]. Chooses [Ano96b]. Choosing [SL97]. CHOP [JMZ^{+}11]. Christmas [Mat92a]. Christos [Ste16]. Chuck [BKP12]. Chunnel [Kir91b]. CIFAR [KHS^{+}23]. CIFAR-10 [KHS^{+}23]. Cintia [CR95b]. Ciphertext [TSMS23]. Circuit [Con03, EDL^{+}04, HC84, Kid14, KP90, YBNS15, Seg97, Ste84a, Ste15a]. circuit- [Seg97]. Circuit-Level [EDL^{+}04]. Circuitry [SC24, SO02]. Circuits [AMR^{+}06, ACA^{+}20, CB10, Gre21c, Gre21e, Lin98, MFM02, Mur06, NBM^{+}06, TKM^{+}02, UTB^{+}06, VN96, Ano02c, IWM89]. CISC [Mil88b, Pit96b, Sch96]. Cisco [Ano03e]. CiSE [Ano21-68]. Citations [Yi23c]. Civil [Kah92b]. Claims [Emm06c, Ste17b, Yi22b, Yi22c, Yi23a, Yi23b, Ano95d, Ano02c, Emm05a]. Clarifying [CDBY23]. Class [BJW^{+}23, PLK^{+}16]. Classic [NM24]. Classification [Goo84, Kir84b, LK10, XLX^{+}23, YKL05]. classifier [VTVM94]. classifiers [BSB^{+}92]. Classifying [GM00]. cleanup [Mat00d, Mat05e]. click [Ste01a, SPRK04]. clicks [Gre06f]. Client [DBDF97]. Client-Server [DBDF97]. climbing [Gre05d]. Clipper [Hun87, Pri94a, SMHB91]. Clock [Del94b, MSA^{+}03, PVS^{+}11, PDT98, Cra90]. Clock-Network [PVS^{+}11]. Clockless [BY07, Cum04, Ano01e]. Closer [Ano96l, BCH^{+}23]. Closing [Gre98a]. Cloud [ABA^{+}21, Ano14n, Ano14-32, Ano14-33, Ano15g, Ano15t, Ano16q, Ano17v, BEL^{+}23, DK18, FOP^{+}19, GZC^{+}20, GLD^{+}22, Gur09, KMK^{+}19, OGLG^{+}22, PSB^{+}20, SABS20, TES^{+}18, XCZ^{+}21, ZL16]. Cloud-Optimized [SABS20]. Cloud-Scale [FOP^{+}19]. Cloud-to-Edge [PSB^{+}20]. Clouds [CCP^{+}17, KGMT17, MFN^{+}17, MMB12]. CLS [Ste14a, Ste14b]. Cluster [BDH03, KPMHB11, LCY^{+}04, RPL^{+}17, WOM01, Ano02b, GK97]. Cluster-Based [WOM01]. cluster-supercomputing [Ano02b]. Clustering [PcFH^{+}02]. Clusters [AJC^{+}20, RBKL11]. CMOS [Ano02d, BJO^{+}09, BKM^{+}82, BY17, Bos05d, CCA^{+}19, Gun06, HBd^{+}99, LBD^{+}99, MKNK83, RDJ^{+}13, STT^{+}15, STS^{+}92, WHA89, WN92, YKH^{+}19]. CMOS-Based [YKH^{+}19]. CMOS/SOS [BKM^{+}82]. CMP [HHS^{+}00, JMZ^{+}11, ZIM^{+}07]. CMPs [GSLK11, MMB^{+}08]. CMT [CCE^{+}09]. CNN [BFZ^{+}22, MKM15, WFW^{+}21]. CNN-Based [MKM15]. CNNs [SJFM19]. Co [BTK^{+}23, Lie23, PMR^{+}22, SKTO22, SKW^{+}23, YZW^{+}23, ZRB^{+}22]. Co-Design [BTK^{+}23, Lie23, PMR^{+}22, SKTO22, YZW^{+}23, ZRB^{+}22]. Co-Designed [SKW^{+}23]. CO2 [GEH^{+}23]. Coarse [AKAK^{+}18, BDV^{+}08, CSL^{+}06, LPC12, RAG19]. Coarse-Grain [CSL^{+}06]. Coarse-Grained [AKAK^{+}18, BDV^{+}08, LPC12, RAG19]. Coast [Ste07e]. Cobol [CS81]. COCOM [Kir90a]. Code [ATS^{+}22, Aug12, BCC^{+}00, DKyL^{+}17, GJLT12, HKY^{+}95, McG82, MBG^{+}16, Pal82, PO04, RNA^{+}12, SBE01, Ste85e, Ste94b, Ste06a, TATC09, ZKP^{+}23]. Code-Named [DKyL^{+}17, RNA^{+}12]. Codec [BK14, KIM^{+}09]. Codes [GXMZ13, MHP^{+}23, MT03]. Codesign [BSY^{+}10, Can98, CMR97, CGJ^{+}94, De 94, Dem94, GHY^{+}17, HDMT94, Lin98, LLLL09, TM21, Trö98, vBK98]. Coding [PP92, Kli81b, Pet92]. cofounder [Ano03d]. CogniServe [IST^{+}11]. Cognitive [AAG^{+}10, BB17, OYK^{+}17, ZRA^{+}17]. Coherence [Ber09, CSL^{+}06, GMC18, HCW^{+}04, KK10, MHW03, ONS^{+}23, SSF^{+}14, SLB04a, SLB04b, TM94b, TM94a, ZBES15]. Coherency [FRS^{+}09, Sha22]. Coherent [Gus92, War90e]. Cohesion [KJT^{+}11]. Coincidences [Mas21]. CoinTerra [BH15]. coinvention [Gre97f]. Collaboration [Ano98f, ADC00]. Collaborative [Emm07e, Emm08a, Hur97]. Collecting [Ste04c]. Collection [CHAF22, GD01, KTK13, MAK19]. Collective [ABK^{+}17, DWF^{+}21]. Collectives [TSA^{+}22]. Color [APS98, SMR18]. Column [THP^{+}19]. Column-Oriented [THP^{+}19]. Columnists [Alt12e]. Combat [LCWB08]. Combined [PKP15]. Combining [CH94, SK97, TM17, TCF96, TO96]. Come [Ano97c, MCR17, Noh19, Ste88e, Qur19]. Coming [Ste07b, Mat96b]. Comment [Ano88e, Ste89b]. Comments [Buc85, Col89, Hoo89c, Kar88a, Luu90a, ZVHL85]. Commerce [SK01]. Commercial [Gre99a, Gre15d, Gre15e, Lee21, MEB^{+}20, PS20]. Commercializing [Gre98b]. Commission [Ste95b]. Commitment [Ste08c, Ano02c]. Committee [Kir85a, Rob99e, Ano18t, Rob99c]. commodities [Gre04e]. Commodity [HcF04, ZACM14]. Common [Man09, MBG^{+}16]. commonplace [Sak00e]. Communication [AFK^{+}21, AJC^{+}20, Bos04a, Bos06d, But07, DQCL24, DWF^{+}21, DGM^{+}11, DBC^{+}98, EVM^{+}98, FKV20, GSLK11, KPK^{+}10, KZ01, KPP06, KPKJ08, KSK18, Mat11a, OKN^{+}11, RAA^{+}21, SMR07, TSA^{+}22, XYCS02, BT84, Bos05e, GK97, HP85, JKP89, JKN96, RT86, SK97, VBB95, Zha91b]. Communication-Centric [KSK18]. Communication-Enabled [RAA^{+}21]. Communications [ACDG99, CAV^{+}14, FME18, Gre05a, IHCE07, KTC18, Lea85, LS98a, NIJ^{+}03, Han96, KY91, PW96, SLM^{+}97, ZG96]. Compact [BDSC21, WKK^{+}14, IKK96]. Compaction [Liu02, SO02]. Companies [Ste85h, Yi21a, Yi21c, Yi22a, Yi22b, Yi22c, Yi22d, Yi22f, Yi22e, Yi23a, Yi23b, Yi24b, Yi24a, Ano97p]. Comparative [SMAS16]. Comparing [KAK96, NM96, PJB^{+}14]. Comparison [And82b, CBLR86, GmDT83, LCY^{+}04, PW89, Tea82, And82a, Bor85b, De 83, Eng00j, Luu90a, NN81b, mDTG81]. Comparisons [Mac84, Rys84, Smo88b]. Compatibility [Han84, Kir83b, Ste87c, Mat96f, Ste93g]. Compatible [DDG^{+}19, Eng00j]. Compatibles [Han87]. Compcon [Mye82a, Mye91b]. Competing [Cle03]. Competition [Ano99r, Gre02a, Ste89f, Ste95e]. Competitive [Gre02c, Ste02d]. Compilation [Joh22b, LBR^{+}22, WLKN22, CFM^{+}97, Ste89f]. Compiler [AS22, BCC^{+}00, KPHP04, Pen90, VCS^{+}19, WMC^{+}06, AH96]. Compiler-Based [KPHP04]. Compilers [NM22]. Compiling [AW22, AH96, CFM^{+}97, NDR^{+}22, dCMA22]. Complementary [KKH^{+}24]. Complements [Gre12c, Gre06f]. Complete [Ano97a, Ano98a, CDS07, Ano96a, EKM^{+}95]. Complete-System [CDS07]. completely [Kah93d]. Completes [Ste84e, Ano02b]. Complex [FHP00, PVB^{+}20, AO97, CG95, ESW97, MM87]. Complexity [ACG03, BAM03, HCP^{+}03, Moo03, Moo04a, Mat04a, Rit97]. Complexity-Aware [ACG03]. Compliance [Ano97-34, Ste99d]. Compliant [SLZ23]. Component [EEKS07, FSH^{+}01, STR^{+}01, Han81]. Component-Based [FSH^{+}01]. Components [ANJ^{+}04, Bor05, Mur03, Bos06a]. Composable [HLS^{+}21, Sha23b, ZWB19]. Compound [LH12]. Compound-Access [LH12]. compounds [Pri94b]. Comprehensive [MBS08, NMZ13, YBNS15, YYK^{+}20]. Compressed [MBG^{+}16, OSS^{+}24, SW14, TSFS21]. Compressing [Tho92, Ano93]. Compression [JGM^{+}20, LKGL24, OG24, BCF^{+}92]. Comprex [OG24]. COMPSAC [Ano17m, Ano20v, Ano20w]. Computation [AT93, Bos04a, Das17, DGM^{+}11, FSK^{+}22, KSV^{+}21, KGMT17, Kra96, MSS15, Smi17, SJB09, SVC01, TT12]. Computational [ANJ^{+}04, JP17, RLC^{+}13, RES^{+}13, TKM^{+}02]. computationally [FBGB96]. Computations [LSL^{+}15, RG88]. Compute [ADJK20, BBSG11, HOF^{+}12, Sha23b, SsSMB24, TSV^{+}20, VBC^{+}21, ZSS^{+}19, ZRB^{+}22, OZT^{+}22, Sha23a, SC24]. Compute-in-Memory [SsSMB24, ZSS^{+}19, ZRB^{+}22]. Computer [ABZ08, AC05, Alb10a, Ano88a, Ano88h, Ano96h, Ano14o, Ano15u, Ano16f, Ano16x, Ano16s, Ano16v, Ano16t, Ano16w, Ano16u, Ano16r, Ano17z, Ano17w, Ano17-27, Ano17y, Ano17x, Ano17-28, Ano17-29, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18m, Ano18n, Ano19g, Ano19z, Ano19u, Ano19v, Ano19w, Ano19y, Ano19x, Ano20-34, Ano20-32, Ano20-33, Ano20g, Ano20e, Ano20f, Ano20-48, Ano20-38, Ano20-39, Ano20-49, Ano20-44, Ano20-40, Ano20-50, Ano20-45, Ano20-41, Ano20-51, Ano20-46, Ano20-52, Ano20-42, Ano20-47, Ano20-43, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21k, Ano21-41, Ano21-53, Ano21-58, Ano21-36, Ano21-46, Ano21-52, Ano21-47, Ano21-42, Ano21-54, Ano21-37, Ano21-59, Ano21-48, Ano21-43, Ano21-38, Ano21-55, Ano21-49, Ano21-44, Ano21-60, Ano21-56, Ano21-39]. Computer [Ano21-50, Ano21-45, Ano21-51, Ano21-40, Ano21-57, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22l, Ano22m, Ano22-47, Ano22-41, Ano22-58, Ano22-48, Ano22-42, Ano22-59, Ano22-49, Ano22-43, Ano22-60, Ano22-55, Ano22-50, Ano22-44, Ano22-61, Ano22-51, Ano22-56, Ano22-45, Ano22-53, Ano22-62, Ano22-52, Ano22-57, Ano22-54, Ano22-46, Ano22-63, Ano23-27, Ano23y, Ano23z, Ano23-34, Ano23-47, Ano23-53, Ano23-56, Ano23-48, Ano23-54, Ano23-57, Ano23-49, Ano23-43, Ano23-35, Ano23-39, Ano23-50, Ano23-44, Ano23-36, Ano23-40, Ano23-45, Ano23-38, Ano23-51, Ano23-37, Ano23-41, Ano23-46, Ano23-52, Ano23-58, Ano23-55, Ano23-42, Ano23p, Ano24q, Ano24s, Ano24t, Ano24o, Ano24r, Ano24n, Ano24u, Ano24v, Ano24p, AF84, BBB^{+}21, BL23, Bel96, BGS89, Bos04d, Bre10, Bro17, CS15, CLM08, Cle00a, Cle00b]. Computer [De 94, DPY18, DMG00, Dwa19, Ebe03, ENSD03, Eec15e, Eec15f, Eec16b, Eec23, EI87, EIB90, ET09, Emm08b, Eng00e, FL13, FV12, FMV85, Gad07, Gon18, Gro02, GEH^{+}23, GT22, Gus84, HAC^{+}13, Hyd00, JQ17, Jim21, Joh22f, Joh23f, Kah91b, KNN^{+}90, KDH^{+}16, KSV^{+}21, KT14, Kim20, Kir89d, Kir91c, KB91, LE18, MS16, Mar14, Mas21, Mat83, Mat21a, May21, Mud10, Sak89, Sak90b, Sha96, SY06, SRU^{+}23, SSH^{+}03, Sla90c, SKW^{+}23, Ste83a, Ste91b, Ste92a, Ste08d, Ste08e, Tab84, TSV^{+}20, TSW^{+}23, TRY^{+}09, TM14, Tor12, ULS^{+}00, VW03, WWF^{+}06, Wen18, Yao85, Yi21a, Yi21c, Yi22a, Yi22b, Yi22c, Yi22d, Yi22f, Yi22e, Yi23a, Yi23b, Yi23c, Yi24b, Yi24a, Ano94c, Ano01f, Ano01h, Ano02c, Eng00j, Gil96a, Gre95c, HS85, Hsi91, Kah90c, MM87, NA84, Sak00a, Ste93d, Wil95b]. computer [vW83, Ano96c, Ano01d, Mon97, Mye85a, Ano20-71, Ano22b]. Computer-6300 [Mye85a]. Computer-Aided [De 94, Yao85]. Computer-Based [EI87]. Computer-Software-Related [Ste08d, Ste08e]. Computer-System [AF84]. computerized [Ste96c]. Computers [Ano87b, Ano88f, Ano98-32, GBD^{+}20, HLZ^{+}16, HML^{+}21, MTS^{+}12, Mat91a, MLM^{+}20, Mye82d, NAA^{+}20, Pri93b, RAA^{+}21, Sak93, Sak02g, Tab84, TSP02, AHO^{+}90, Ano97n, GP90, Gre95a, Laz89, LLC90, Pen99, Sho85, Ano19p, Ano20u, Ano20o, Ano21u, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano23w, Ano23x]. Computing [AHK^{+}14, AKAK^{+}18, AKK15, Alt12d, Alt14e, And14, ACA^{+}20, Ano94d, Ano13c, Ano14n, Ano14s, Ano15g, Ano21w, Ano23j, ACG^{+}95, BC21, BDSC21, BCN^{+}22, BR10, Bar21, BPT^{+}11, BSA21, BJ14, Bro11, CFK^{+}10, CCYT05, CSM^{+}21, Che19, CMAS11, DBDF97, Das22, Eec18a, Eec18b, Fer98a, For02, GAT^{+}22, GLN^{+}08, GS21, GHN^{+}12, Gre98e, GSS^{+}07, GGB^{+}15, GKL^{+}22, Gur09, HRK^{+}24, HGS^{+}17, HKC10, HML^{+}21, IG15, IBN^{+}21, IT15, JL11, JS18a, JGC^{+}11, JC08b, JS18b, Joh23b, Kah91c, Kah92f, Kah93f, KMN^{+}04, Kar21, KDK^{+}11, KCXmWH17, KKS^{+}23, Kir89a, Kir89c, Kra96, Kur20b, Kur21b, Kur21c, LBS^{+}11, LC18, LKJ^{+}22, Lee24b, LP21, LRC^{+}09, LNOM08, LCP^{+}11, LAT^{+}01, MBSP02, MYK^{+}10, Mat90b, Mat02a, MBP^{+}85, MSY^{+}22, MKRC97, MK10, NJZL^{+}17, NI14, NMU^{+}15, ND10, OVT90, OSS^{+}24, PLK^{+}16, Pen99, PDL08, PCDL10]. Computing [PJB^{+}14, QT21, RG85, RPL^{+}17, SVA^{+}22, Sak02a, SLC^{+}14, SJO01, SIL^{+}15, SCS^{+}09, SRL91, Sha23a, Sta01a, Sta01b, SSK23, SMT^{+}14, TMBT94, TMJ13, TAI^{+}21, TC15, TVV^{+}21, VC11, WRA^{+}14, WLD15, War91b, WB12, WGM02, WWR97, WLY^{+}21b, Wu23, WHP^{+}13, YKH^{+}19, YHT^{+}15, ZSB21, ZL16, ZUNN18, ZRA^{+}17, Ano94b, Ano99p, Ano01e, Ano02d, Ano03b, CMR97, Dia95d, Fer98b, Gon97, Gre96a, Gre96c, Ipe19, Lou91, Sak01d, Ano19-28, Ano15t, Ano16q, Ano17v, Ano20-53, Ano19a, Ano19b, Ano19c, Ano20a, Ano20b, Ano20d, Ano21t, Ano21a, Ano21b, Ano21c, Ano21j, Ano21n, Ano21o, Ano21p, Ano22c, Ano22d, Ano22e, Ano22j, Ano22k, Ano23u, Ano23c, Ano23d, Ano23j, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23m, Ano23n, Ano23o, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f]. Computing [Ano24a]. Computing-Based [ACA^{+}20]. ComputingEdge [Ano22-27, Ano20c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano22f, Ano22g, Ano22h, Ano22i, Ano23k, Ano23l]. Concept [MB15, THP^{+}19]. Concerning [Ste08a]. Concerns [CHA^{+}85a, Kar85, Ste89a, Ano01c, Mat95d, Ste99c]. Concurrency [Dea04, Yea96]. Concurrent [LHM99, Mye84c]. Conditioner [Ano97h]. Conditions [KSE^{+}22, MSS15]. Conference [Ano23a, Ano23b, KB13]. Conferences [ABZ08, Alb04, Ano14d, Ano15h, Ano15i, Ano17n, BL23, Dwa19, GT22, Joh22f, Joh23f, MS16, MRLB03, Mud10, TM14, Wen18, Ano94c, ET09, FL13, FV12, JQ17, Jim21, RG07, Tor06]. Confidentiality [ZG96]. Configurable [CCP^{+}17, FSH^{+}01, Gon00, Gon06, KPHP04, SLSO14, RH91, KSB21]. Configuration [OWK87]. Configurations [Ste86a, Gil96a, PGL97]. Configure [ACKM05]. Conflicts [Gre13f]. Conformance [AQT^{+}92]. Confrontational [Gre21a]. Confronting [Mat01a]. Congesting [OESGG^{+}21]. Congestion [CGLES^{+}23, CNC^{+}16, GQF^{+}06, Gre16a, KKP^{+}14, KM05, SLZ23]. Congestion-Aware [KKP^{+}14]. congratulations [Ano01d]. Congress [Ano19-29, Cha85b, Ste84b, Ste09b]. Connect [Ano17-47, Ano17-48]. Connected [Joh20b, LW94, Ano15-38, Ano19-39]. Connecting [FH00, Sak00e]. Connectivity [Gad07, Joh20b, Joh21a, Joh20b]. connectors [Bel93]. Conscious [ROA13, TCD^{+}05]. Consider [War90f, Ano94b]. Considerations [CGO00, Joh87, Cat88, FN86]. Considered [AW06, NMHS15]. Consistency [HCW^{+}04, LPM15, RLS11]. Consistent [MBSP02, Gil96a]. Consoles [ML21]. Consolidation [SGC94, Gre05a]. consortia [Rob01a, Upd93]. Consortium [Ano01h, Eng00f]. Constant [LHN95]. Constant-Time [LHN95]. Constrained [MLL^{+}18, SNM^{+}22, WK13]. Constraints [CDY^{+}18, HRSS11]. Construction [SO02]. Constructs [NJZL^{+}17]. Consumer [Wv92, Gol96]. consumers [Gre96e]. Consuming [Ano97g]. Consumption [HCP^{+}03, JLSM03, LS98b, Seg97, ZZ05, PGL97]. Containers [HSN^{+}23, SDG^{+}21]. Contemporary [JM98, SSLV15, De 83, mDTG81]. Content [GGB^{+}15, MC92, SML04, Ste97c, ZLBI06, Ano99w]. Content-Addressable [MC92]. Content-Aware [ZLBI06]. Content-Processing [SML04]. Contents [Ano13j, Ano14-35, Ano14-36, Ano16-42, Ano16-43, Ano17-54, Ano17-49, Ano17-50, Ano17-51, Ano17-52, Ano17-53, Ano18-39, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano19-40, Ano19-41, Ano19-43, Ano19-44, Ano20-65, Ano20-66, Ano20-67, Ano20-68, Ano20-69, Ano20-70, Ano21-69, Ano21-71, Ano21-72, Ano21-73, Ano21-74, Ano21-75, Ano22-74, Ano22-75, Ano22-76, Ano22-77, Ano22-78, Ano23-77, Ano23-78, Ano23-79, Ano23-80, Ano23-81, Ano23-82, Ano24z, Ano24-27, Ano14-37, Ano19-42, Ano19-45, Ano21-70]. Context [DMG^{+}15, HGS^{+}17, TVT19, Mat01a]. Context-Sensitive [TVT19]. Contexts [CS14]. Continual [SRA^{+}04]. Continue [Eng00m, Jam90]. Continued [Far86]. Continues [Bri94, Dai94, Dun82, Mat21c]. Continuing [Ste03b]. Continuous [MS84, RTM^{+}10, VRMC20]. Contracts [OFKS23]. Contrast [SGL93]. Contributions [LE18]. Contributors [Far91]. Control [AKK15, BdS98, EPZ02, EEJ95, HSN^{+}23, JBM95, Kir87, Kir90e, KTC18, KM05, MS84, Mye81, Pal93, PPA^{+}14, PC01, SLZ23, WM85, WHCK18, WJM^{+}05, WMC^{+}06, ZLTW13, CR95b, CDGO97, MKNK83, OTM82, PVYU94, Rob98c, Rya88, SCG95, Shl93, SM85, Tau84, Tau87, Wil84]. control-flow [PVYU94]. Control-Systems [Kir90e]. Controlled [KKL^{+}09, QJP^{+}08, SL84b]. Controller [AO97, CR95b, RGF96, THP^{+}19, TTF96, TSW^{+}01, YW88, BCF^{+}92, Cat88, DM86, GP95, LGJ95, Man86b, Man86c, NF81, RGF95, WBC^{+}95, WJR88]. Controllers [BI13, GTF97, MM09, ZMVH^{+}83c, ZVHL85, MST^{+}85, MM96, TZMVLN81, VVRV95, ZMVH^{+}83a, ZMVH^{+}83b]. Controversy [Ste84c, Ste00b]. convenient [Dia95d]. Convention [Ste88e]. conventional [TCF96, TONH96]. converge [Gre99f]. Converged [PKB^{+}15]. Convergence [Gre97b, Moo04a]. Conversation [Gre21a]. Conversion [EIB90, Jae82a, Jae82b]. converters [DFR90]. Convey [Bre10]. Convolutional [BCKY17]. Cool [Alt11d, Alt13a, Ano17o, Eec17a, EW23, IA11, IA13, IA22, Joh22e, Mas93, Nak99, OYS^{+}11, Ano14e, EW24, KKH^{+}24, Lee24b, Ano15j, IA09, Nak00]. COOL-NPU [KKH^{+}24]. Cooled [MMC^{+}22, Ano03e]. Cooling [CMAS11, KCHA21a, KCHA21b]. CoolThreads [FRS^{+}09]. Cooperation [GS99, Kah93a, McL87]. Coordinating [SKJ^{+}11]. Coordination [CWL^{+}14, Gre10e]. Coordination-Aware [CWL^{+}14]. Coordinator [Gre23d]. COPA [KSB21]. Copilot [Hoe93]. Coping [CSV02, ESW97, Gon97, KKT13]. Copper [Ano99s]. Coppermine [Ano99l, Ano99p]. Coprocessor [AT93, DKB^{+}90, HC83b, JL87, RJR88, TBD19, CPZ89, DVQ96, Kai88]. Coprocessors [BSC^{+}90, CSM^{+}21, WRA^{+}14]. Copy [Ste84b]. Copy-Protection-Defeating [Ste84b]. Copying [Ste86a, Ste91h]. Copyright [Hau88c, Kar88b, Ste84c, Ste86e, Ste87d, Ste89e, Ste04a, Ste06a, Ano91b, Ste90e, Ste93d, Ste93e, Ste96f, Ste00d, Ste02a, Ste04b, Ste91d, Ste06a]. Copyrightable [McG82]. Copyrighting [Gro83, Hec83a, Ste89f]. Copyrights [Ste91c, Ste92c]. Copywriting [Ste88a]. CORDIC [CAH86, Vac87]. cords [Eng00j]. Core [Ano16-48, Ano16-47, Ano16-46, BYM^{+}07, BJO^{+}09, BY07, CLM08, CGG^{+}21, CWS^{+}12, DXT^{+}18, DKyL^{+}17, DFG^{+}13, Edw99, EBC22, FZW^{+}12, FJL^{+}13, HMB^{+}14, HKC10, IHCE07, JJK^{+}11, KST04, LAT^{+}01, MAT^{+}18, MIM^{+}97, MB05, RHH^{+}03, SCS^{+}09, SLL^{+}18, SMS13, TKI^{+}14, WK13, XCZ^{+}21, YMA^{+}13, ZSB21, Ano16-45]. Cores [AFGM10, Bos03c, CSM^{+}21, KST12, LLT^{+}08, MBS08, MGG^{+}19, SB23, WS13, Ano00g, Ano03e, Jag97]. Cornell [Ano02b]. Corporate [Ano13d, Dia93d]. Correct [LPM15]. Correction [EDL^{+}04, Man86a, Mar84, Nel84, RGF96, SC24, Zha91a]. Corrections [Ano01a, Mac84, RTJ21, Rys84]. Correlated [WFA^{+}10]. correlators [WCH94]. Corrigendum [Sav99a]. Cortex [TKI^{+}14]. Cortex-M0 [TKI^{+}14]. Cost [BMG^{+}21, BCC^{+}02, Car93, CFRM04, Das21, Dea04, Far85, FBHN04, GALB07, Gre07e, GH88, HSP^{+}01, KDSA09, KCS^{+}20, Lea88, MBS08, MS87, Mye84c, SG01a, Sto90, UBH^{+}94, Wal97, AO97, Ano02c, DVQ96, Dia95d, DS95, GK97, Gol96, Jag97, KSI^{+}96, PGL97]. Cost-Effective [BMG^{+}21, BCC^{+}02, Das21, Far85, GH88, Lea88, Mye84c, DS95, KSI^{+}96]. Cost-Efficient [KDSA09]. Cost-Sensitive [CFRM04, Gol96]. Costs [Ano87g, CDGO97, Han96]. Cosynthesis [OHLR94]. COTS [CML^{+}23]. could [Ano02c]. Countering [RTJ20, RTJ21]. Counters [EEKS07, SIPM02]. Counting [RYK18]. Counts [FBHN04]. Couple [Alt12c]. Coupled [Kir85b, Pre91]. Course [Hyd00, Mat90c, Ano94c, Gre96e, Hal91]. Court [Ste92d, Ste06a, Ste13, Ste06b, Ano98v, Ste07c, Ste07d, Ste07e, Ste08b]. courts [Ste89e]. Cover [Ano13f, Ano14i, Ano14k, Ano14l, Ano14m, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano16o, Ano16k, Ano16l, Ano16m, Ano16n, Ano17r, Ano17s, Ano17t, Ano18b, Ano18j, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano19q, Ano19r, Ano19s, Ano19t, Ano20-31, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano24k, Ano24l, Ano14j, Ano19-39, Gil96a]. coverage [Ste04d]. Covered [Ano19y, Ano20-48, Ano20-49, Ano20-50, Ano20-51, Ano21-46, Ano21-47, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano22-56, Ano22-57, Ano23-47, Ano23-48, Ano23-49, Ano23-50, Ano23-51, Ano23-52, Ano24s]. Covert [VCD16]. Cows [Pri93b]. CPI [EEKS07]. CPU [ANJ^{+}04, Ano98g, ANM^{+}12, CGO00, Cra90, JWS^{+}19, KCHA21a, KCHA21b, Kum97, LSL^{+}15, RHH^{+}03, RYR^{+}22, Sak87b, VPV12, ZHR15]. CPUs [Alt11b, Bro11, ESG^{+}05, Has85, JAS^{+}22, Kur21a, OFKS23, Sak99d, Seg97]. cracks [Gre00c, Ste05b]. Crash [Gre02c, WN94]. Cray [Ano17-45, DVWW05]. craze [Rob98b]. CRC [AS90, Bro86, Per83, RG88, Sho85]. CRC-16 [Sho85]. Creates [Hec83b]. Creating [Ano99h, Gre23d, HO99a, Mat99b, YHHF20]. Creation [Gre10b]. Creative [Emm07a, Emm07d, Gre04b, Emm05c]. Creativity [Mat91a]. Credible [Raj94]. Critic [FSR^{+}05]. Critical [FPAF02, Fre02, Koo02, MAT^{+}18, RKA^{+}20, SKA^{+}14a, SMQP10, vBK98, Mat96b]. Cross [Ano17p, ESW97, KAK^{+}22, KGDW^{+}13, Gon97, HP81]. cross-assembler [HP81]. Cross-Development [ESW97]. Cross-Domain [KAK^{+}22]. Cross-Layer [KGDW^{+}13]. cross-platform [Gon97]. Cross-pollinate [Ano17p]. Crossbar [CD09, Cum04, GM99, MAM^{+}06, NBS^{+}18, PKP15, SFP^{+}23]. Crossbars [Ano18d, TMA18]. Crossing [WNW^{+}16]. Crowds [Gre23b]. Cruise [Pal93]. Cryogenic [BML^{+}21]. Cryptocoprocessor [HV04]. Cryptocurrency [BH15]. Cryptographic [NM96, TLYL04, TBDL01]. Cryptography [Ano97m, LSY01, SFP^{+}23, DVQ96, NM96]. cryptography-dedicated [NM96]. Cryptoprocessor [MS83]. Cryptosystem [NM24]. Cryptosystems [ESG^{+}05]. Crystal [Ano88b, DTH^{+}95]. Crystalline [TKI^{+}14]. Crystals [Ano02d]. CS [Ano16c, Ano17g, Ano96j, Ano96k, Ano14-38, Ano14-39, Ano15-41, Ano17-55, Dia96a]. CSIDC [Cle03]. Ctron [OWK87]. Cube [PFC^{+}02b]. Cubes [YW94]. CUDA [GLN^{+}08]. Cuenet [RG85]. Cup [Chr90, Joh90a]. curmudgeons [Gre96f]. Current [NRV^{+}06, Sak90b, Ste92a, SMAS16, Ano01d, BCF^{+}92, Pri94b]. Curriculum [Cle00b, Hal91]. Custom [ABG^{+}20, KP90, TZMVLN81]. Customizable [TGC^{+}20]. Customization [RT23, TVT19, BC86]. Customized [Ano00g]. Customizing [HCP^{+}03]. CUTIE [SDF^{+}23]. Cutting [Eec17b, LB00]. Cutting-Edge [Eec17b, LB00]. Cuttings [GM00]. CVPR2020 [Ano20x]. CXL [Sha23a, BEL^{+}23, BWR23, GKB^{+}23, HRC^{+}23, KKS^{+}23, KJC^{+}23]. CXL-Based [BEL^{+}23]. CXL-Enabled [BWR23]. Cyber [Ano15-33, Ano15-29]. Cybersecurity [Ano14-31, Ano15-31, Ano15-30]. Cycle [EE10, KMK^{+}19, KCKP14, Cra90, Han96]. Cycle-Exact [KMK^{+}19]. Cycles [Dia95a, Gre18c, Gre21f]. Cyclic [SC24].

Cyrix [Ano98g, Hur98].

D [Ano14o, Ano17-27, ASX19, Alt14e, Ano96o, AOYS95, BWMS19, CMAS11, DDG^{+}19, DTH^{+}95, DFG^{+}13, HRK^{+}24, IPL^{+}23, Joh19c, LXB07, LX10, MKT^{+}13, MAS^{+}07, NST97a, NST97b, PMM15, PZB^{+}19, SYW^{+}14, SCSR93, VPV12, WLF^{+}08, ZSS^{+}19]. D&I [Ano22-53, Ano22-54, Ano23-44, Ano23-46, Ano24q, Ano24r, Ano23-43, Ano23-45]. D-Integrated [WLF^{+}08]. D30V [TWN^{+}99]. D30V/MPEG [TWN^{+}99]. Daisywheel [Han85]. damages [Ste04c]. dance [Ste99e]. Dangerous [Alt13b]. Dark [Alt13b, EBS^{+}12, GHSV^{+}11, Gre11a, HFFA11, LDL17, RES^{+}13, SKS^{+}13, TS13, Tay13]. DARPA [Mat97a]. Darwin [TBD19]. Data [ADJK20, AFH16, AKK15, Alt14a, AS10, Ano14-30, Ano16-36, Ano16-37, Ano16-38, Ano19-38, BCM^{+}14, BG16, BCN^{+}22, Ber09, BK14, CGS10, CSC^{+}22, CWLS15, CHAF22, CS81, DQCL24, DK14, EV97, FG14, FSS^{+}16, FDS^{+}17, FSK^{+}22, FJB^{+}22, GKL^{+}14, Gre14a, Gre15a, GHLK^{+}12, GSLK11, GGB^{+}15, HLS^{+}21, HLIT20, Jos86, KMK01, Kir83b, Kir84a, KPR^{+}22, KSM^{+}89, KCS^{+}20, Lea85, LS22, LPC12, LKGL24, LCWB08, MA83, Mat13a, NS05, OESGG^{+}21, Pat84, PVB^{+}20, PJB^{+}14, RC12, RSW10, RTM^{+}10, SG00, SLC^{+}14, Sha23a, SRU^{+}23, SAC^{+}21, SSK23, SCH^{+}23, SMJ^{+}11, SKM23, Tho92, TT12, TKS^{+}22, VAFF^{+}10, WMH^{+}10, Wil95a, WBKR14, XYCS02, XWZ09, YHHF20, YYK^{+}20, Ano01h, Ano02e, CDGO97, DFR90, Jae82a, Jae82b, Jae82c, Jae83, KHW85, KAK96, Lou91, PVYU94, Ste89f, Ano16y, Ste84a, Ste08c, Ano19i, Ano21m, Ano22q, Ano22r]. Data [Ano22s]. Data-Center [GHLK^{+}12, OESGG^{+}21]. Data-Centric [KCS^{+}20, RC12, Sha23a]. Data-Compressing [Tho92]. Data-Driven [CSC^{+}22, KSM^{+}89, SRU^{+}23]. Data-Flow [LPC12]. Data-Intensive [CGS10, GGB^{+}15, SLC^{+}14, SAC^{+}21]. Data-Level [EV97]. Data-Oblivious [YHHF20]. Data-Parallel [WMH^{+}10, Lou91]. Data-Processing [CS81]. Data-Security [Wil95a]. Data-Triggered [TT12]. Database [AS91a, AS91b, BGRKR88, FBGB96, FTKS92, LHMH91, Mye84a, Ste91c, SMT^{+}14, WLP^{+}15, HLHR90, Hsi91, ISH^{+}91, Mat05b]. Databases [Ano92c, FM91, Kah92e, MG89, MCV^{+}14, Ano97r]. Datacenter [Alt14d, BR10, BvdGM^{+}15, CFO^{+}18, KNB14, LM16, MK10, PSP14, PCC^{+}15, RSW10, VPRS14, WAA^{+}21]. Datacenter-Scale [BR10, WAA^{+}21]. Datacenters [HSN^{+}23, MMC^{+}22]. Dataflow [CB04, CES17, FGC^{+}14, GFL^{+}17, HKS16, NGS16, ZRA^{+}20]. DataPlay [Dav02]. DataPort [Ano23-59]. Datawave [SC91]. date [KS00]. Days [Gre07b, Ano97o]. dBASE [Ste88d]. DBMS [Pap89]. DC [GA86]. DC/AC [GA86]. DDC [Kid14]. DDOS [Gre19a]. DDRx [BI13]. de-Facto [Hec83b]. Deadlock [RGK19]. Deaf [Mye83a]. Deal [Gre22c, Ste93a, Ste00b, Ste00c, Ste00a]. Dealing [Mat05a]. Death [Lah84]. Debate [Alb07c, Dun82]. Debates [Eec16b]. Debugger [CHSL17]. Debugging [CP86, GLD^{+}22, LPL86, MKOK88, NPC06, ZQL^{+}04, vW85, ESW97, EKM^{+}95, Rit97]. Debut [Ano97-27, Sca98]. DEC [Ano97i]. Decade [AC05, Del91b, Far91]. Decades [HSNJ21]. Decentralized [ZCW^{+}14, BNOv87]. Decides [Ste08b]. Decision [Ste84a, ZMVH^{+}83c, ZVHL85, MST^{+}85, ZMVH^{+}83a, ZMVH^{+}83b]. Decisions [Gre22a]. Declarative [HLHR90]. Decoder [YKG18, DKM^{+}92]. Decoding [TVT19]. decomposing [CG95]. Deconstruction [Gre04b]. Decoupled [AKK15, SW14]. decrease [JKN96].

Dedicated [Hun95, Nic91, DVQ96, KWGG95, NM96]. Deduplication [ZHZ^{+}19]. Deep [Ano97o, AJC^{+}20, CES17, DKSL04, EWW^{+}19, EPM^{+}20, FHR99, GTLY22, HRK^{+}24, IPL^{+}23, KSLY17, KSE^{+}22, KGT22, KR19b, Lie23, MSP^{+}19, MM23, MCV^{+}19, MAJ^{+}18, PEZ^{+}19, PMR^{+}22, RT23, SLL^{+}18, TAI^{+}21, TSA^{+}22, TSMS23, YKH^{+}19, YCD^{+}19, ZFW^{+}23, ZRA^{+}17, hHH99]. Deep-Learning [AJC^{+}20, MAJ^{+}18, SLL^{+}18]. Deep-Learning-Based [IPL^{+}23]. Deep-Submicrometer [YCD^{+}19]. Deep-Submicron [FHR99]. Deeply [HC02, ESW97]. DeepTools [VCS^{+}19]. Defeating [Ste84b]. Defect [TMA18, Ano18d]. Defect-Tolerant [TMA18, Ano18d]. Defense [YFDV19]. Define [Gri21]. Defined [BDV^{+}08, CN13, KKS^{+}23, LLW^{+}07, MMB12, SYY^{+}11]. Defines [Isa83, Kir83a]. Defining [BAH^{+}05, EKM^{+}95]. Definite [KW83]. Definition [Sak02a, Pet92, Sib84]. Definitions [Mat92b]. Defuzzification [RGF96, RGF95]. Defy [Goo84, Kir84b]. Degradable [GU98]. Degradation [AVU^{+}08, Bor05]. degree [Mat96f]. Déjà [Gre18d]. Delay [BF02, KBK03, PD01]. Delay-Insensitive [BF02]. delays [Ano99l, Ano99p]. Delivering [DBDF97]. Delivery [Ano98-36]. Delta [Pow94]. Delta-4 [Pow94]. Demand [ABIV06, Gre10a, KCHA21a, KCHA21b, TVT19]. demands [Ano02c, Sak00b]. Demise [Ste92f]. Democratic [GPSS83]. Democratization [Alt14a]. Democratizing [CSC^{+}22]. demos [Eng00j]. Dendrites [KHS^{+}23]. Denial [Pit96a]. DeNovoND [SKA14b]. Dense [NY22, Ipe19]. Denser [Joh19c]. Density [HKY^{+}95, Mye92b, OMMB13, Bel93, DP97]. Denver [BBTV15]. Department [Ste15b]. Dependable [Ano01a, ABC99, BFLS01, PV01, SUF^{+}12]. Dependencies [Gre21c, PVYU94]. Dependency [ED18, LS22, Ano94b]. Deployed [SKM23]. Deployment [Ano99a, LBR^{+}22]. Depth [Gre22d, IPL^{+}23]. Derek [Mor84]. describing [NM96]. description [vdDD90]. Deserve [Ano16p]. Design [ASX19, ABG^{+}20, Ano98-30, Ano98-29, Ano98-31, AS99, ASD^{+}05, BCJ^{+}20, BBB^{+}21, BAH^{+}05, BTK^{+}23, BEL^{+}23, BJW^{+}23, BGH^{+}90, BGS89, BFLS01, Bon21, Bor99a, Bor99b, Bos03a, BAM03, Bos06c, BTR02, BBS^{+}00, BGK97, CSV02, cCCP00, CWS^{+}12, Cla03, Cle03, DXT^{+}18, DGR^{+}10, DM88a, EGL^{+}90b, EGL^{+}90a, Eec15d, EPZ02, Emm08b, FRS^{+}09, FHR99, FH05, GCE^{+}21, GRD22, GH88, HHNK09, HSW^{+}89, HRSS11, Hyd00, Joh20a, Joh20c, Joh87, KNN^{+}90, KRD^{+}20, KIR19, Kli81a, KL05, Koe86, Kul20, Kur20a, Lee94, LS96, Lie23, Lin04, LYP^{+}18, LG24, LXB07, MRJ^{+}15, MT05, Mat13c, MG89, Mel89, MKRC97, Moo04a, MK10, MEB^{+}20, Mye89a, NC86, NPY^{+}21, PMM15, PKB^{+}15, PSL^{+}23, PZB^{+}19, PLBC09, PMR^{+}22, Pre91, QT21, RCR04, Red13, RSS^{+}08, SMHB91, SV03, SNC^{+}07, SKTO22, Sen86, SAW^{+}10, SRWB15, SCA^{+}12, Sim00, SBG^{+}07]. Design [SAC^{+}99, Smi96b, SGC94, STR^{+}01, SCC^{+}05, TCD^{+}05, Tay13, TCF96, UB05, WKK^{+}14, WWZ^{+}08, Won03, YZW^{+}23, ZUNN18, ZWB19, ZRB^{+}22, ZZ02, ZRA^{+}17, AKK^{+}93, Ano99v, Ano02b, Ano02d, AJR86, Bos05f, Bos06e, CH94, CM86, FHMS96, Fly97, GA86, Hea87, Jae83, Joh90b, KKT^{+}91, LDA87, Mat98b, Mat00c, Mat05c, Pap96, Seg97, Sib84, SSL82, SL97, Ste89d, Ste94f, TTF96, VVRV95, Wil95b]. Designed [SKW^{+}23, AH96]. Designer [ENSD03, Ste85f, ZV85, ZVH85, Lan87]. Designers [Ano98-38, Koe86, Ano96n, Eng00j, Gre96e]. Designing [AAWC94, ACG^{+}95, BNV^{+}15, Bor05, Bos06a, CSM^{+}21, ED18, GKL^{+}14, GM99, GEH^{+}23, Har12, HDM^{+}98, HL99, Hsu94, JBF94, KP90, KSK18, Lan96, Mat10b, MAM^{+}06, OS99, Pee87, RLC^{+}13, Sak99d, SKLY97, WBC^{+}95, ZBES15, Bos05a, RC12, Tab84]. Designs [ACG03, Alt11d, Fly97, Joh22a, KKD^{+}07, KSE^{+}22, LB00, LRC^{+}09, TC15, TK21, YBS17]. Desires [MCF^{+}85]. Desk [Dia93a, Mye92a]. Desk-Top [Dia93a, Mye92a]. Deskpro [Ano88c]. desktop [Dia95d]. Desolla [Mor84]. Destabilizing [Ano97p]. Destruct [Ano96u]. Destruction [Gre04b]. Details [Ano98c]. Detect [CML^{+}23, NRV^{+}06, CJFP95, KWGG95]. Detected [Sha82]. Detecting [LTQZ07, LDCS09, VCD16]. Detection [CYH^{+}18, FKL01, GV06, ML05, MBS08, PMS23, SGK^{+}04, SS16, TS06, WZL20]. detects [Ano01c]. Determining [Ste15a, Ste17c]. Deterministic [DLCO10, NPC06, XBH07]. Detour [Sav99a, SAA^{+}99]. Develop [Ano98q]. developed [KWGG95]. Developing [ANS96, BSC^{+}90, Chr96, HBd^{+}99, IKK96, MA94, Pri90, Sak00a, SCSR93, SBG97, TMBT94, XYT^{+}23, Rob97b]. Development [Ano99-27, ABC99, CCS21, DTS20, ESW97, Emm07e, Emm08a, Eng00k, Kah92d, LPL86, Mat01d, Mat08a, MBS92, NL02, NH81, PKR92, SPRK04, WPSR20, WDK^{+}20, Chr96, Hal93, Shl93, Vic93, Wal97, Wil84]. Developments [Ste85b, Ste86e, Ste87d, Ste92a]. Develops [Ano87d]. Device [Eng00o, MRSV11, PMS23, RT23, SB23, SKM23, ZCW^{+}14, ZRA^{+}20, Ano02d, GRS86]. Devices [Alt13a, AAC^{+}16, Ano87a, Ano88g, FHL^{+}03, Hac01, Ham00, KHL^{+}16, Pen01, RYK18, STR^{+}13, Ste86a, SKS^{+}13, WK13, WLD15, XPZ^{+}19, CJFP95, Pri94b]. DGEMM [RBKL11]. Diagnosing [Ebe03]. Diagnosis [CS08, CJFP95]. Diamond [Ano89, Gre04d, Ano01d]. diamond-wafer [Gre04d]. Diamonds [Gre95a]. Did [Ano88d, Ano98t, Gre03e, Gre07a, Joh20c]. Die-Stacked [SLSO14]. Die-Stacking [LXB07]. Dies [Dia96a, Ano01g, Ano03f, Pap96]. diet [Ano03e]. Difference [Ste85e, Gre95b]. Different [Pal82, Hal91, Rob99b, Ste90e]. Differential [MKG^{+}20]. Differentiated [Gre13b]. difficult [TCF96]. Diffractive [TMBT94]. Digital [APS98, Alt13a, Ano98y, Ano13e, CN13, DM88b, DM88a, Eic86, Eng00d, FME18, Fos98, Fra00, FGG^{+}88, GG99, Gas21, Gre10a, Gre11a, Gre13c, Gre23e, HC84, HSP^{+}01, HA96, Hun95, Jae82a, Kaw98, KW81, Klo86, KPHP04, LCS92, Mor86a, MD88, Mor88, MBK^{+}92, NN81a, NM99, NN81b, OHLR94, OW01, PS88, Pet92, Sav99b, SDF^{+}23, SP92, SAW^{+}10, ST21, SK88, Sos94, SWM^{+}20, SsSMB24, TP10, THT^{+}04, VM88, WT98, YYH98, Ano95a, Ano99w, BG81, FLRB86, Gre15c, IWM89, Jae82b, KAK96, KKT^{+}91, Mat95c, Pee87, RS90, SK97, TTF96, Ste08a]. Digital-Readout [HC84]. Digital-RF [FME18]. Digital-Signature [Eng00d]. Digital-Signatures [HA96]. Digital-to-analog [Jae82a]. Digitally [Mur06]. Digitization [Gre10b]. Dilemma [Hua89]. Dilemmas [Gre21b]. Dim [PDS^{+}13, SKS^{+}13, WS13]. Dimensional [YRC^{+}22, DGW^{+}94, Lou91, NA84]. Dimensions [Ano97j]. DINT [KHW85]. Diode [Ano97f]. Direct [Cri97, KMK01]. Directed [CHH^{+}98, CK11, LLZ^{+}04]. Direction [Gre11b]. Directions [Alb10a, Eec16b, Kni85, SVL03, VWC03, NM96]. Directory [KK10]. Dirty [Ste88d]. Disaggregated [KR19b]. Disaggregation [KJC^{+}23]. Disambiguation [SDB^{+}04]. disappearing [Gre95d]. Disassembling [Ste94b]. Disband [Ano87c]. disc [Ano02b]. Discipline [Car98]. Disciplined [SKA14b]. Disclaimers [Ste87b]. Discovered [And82b, Tea82]. Discovering [QLLG15, SPH^{+}03]. Discovery [Ano00b, Mat10b, Eng00l]. Discrete [CF90]. Discrete-Event [CF90]. Discretion [Gre22b]. Discriminating [Ste85f, ZV85, ZVH85]. Discussion [GAT^{+}22]. Disintegrate [KJL16]. Disk [AO97, HY98, MA94, MA83, Ano01f]. Dismal [Gre20c]. Dismisses [Ste06a]. Display [Ano96o, Fer98a, Ste89a, SL84b, GRS86]. Displays [Alt98, Ste88a, Ste89a, Ste89c, Ste89d, Ste89e, Ste90e]. Disqualified [Ste92b]. Distance [AKJF22]. Distillation [LP21, ZFW^{+}23]. Distinguished [Ano14a, Ano15b, Ano16b, Ano17b]. Distributed [CP86, DWF^{+}21, Dra00, DVWW05, FBC87, GM21, Gre22b, JAS^{+}22, Jos86, KHL^{+}16, KPR^{+}22, KDK^{+}89, MS87, Mye81, Pow94, RG85, SK01, SUF^{+}12, TAI^{+}21, WWR97, AGH^{+}91, Gal97, KKC93, LDA87, Mat98b]. Distributed-System [SK01]. Distribution [Dav02, Dia94a]. Dive [Lie23]. Diverge [KJMP07]. Diverge-Merge [KJMP07]. Divergence [LS22]. Diverse [Eec15c]. Diversity [Gre14b]. Division [SL97, Ste07a, ZL15]. divvying [Ste96b]. DLP [SNL^{+}03]. DLX [Ibb00]. DMA [NS81]. DMP [DLCO10]. DNA [BLC^{+}17, GAT^{+}22, KYGW17, KYG19]. DNA-Based [BLC^{+}17]. DNN [DSG^{+}22, HIP^{+}22, JAS^{+}22, KSK18, KCS^{+}20, LKJ^{+}22, MSY^{+}22]. DNNs [AHKY19, CFO^{+}18]. DNPU [SLL^{+}18]. Do [Alb07e, AAP^{+}10, Gre16e, Gre00a, Mat95d, Rob97c, Rob01b]. Doctrine [Ste92f]. Document [Dia93a]. documents [Mat99b]. Does [Gre09c, Gre16e, GSS^{+}07, Yi23c, Gre01d, Mar96, Rob97c, WHKM93a, WHKM93b]. doesn’t [Wil95b]. Dog [Gre07b]. Doing [Mat00a, Ste96c]. DOJO [TSW^{+}23]. Dollar [Gre07d]. Domain [AK24, AKT^{+}18, CYH^{+}18, KAK^{+}22, LBS^{+}11, MSA^{+}03, NKI^{+}09, NGSW17, VAD^{+}21, WNW^{+}16]. Domain-Specific [AK24, CYH^{+}18, LBS^{+}11, VAD^{+}21]. Dominated [KBK03]. Don’t [FBHN04, Gre96b, Rob99d]. DOOM [BNOv87]. Doomed [Ste14a, Ste14b]. door [Gre98a, Ste93c]. Double [RAA^{+}21, XWZ09, Ano03e]. Double-Data-Rate [XWZ09]. doubles [Reg92]. Doubling [Eng00e]. Down [Eec17c, EEKS07, Eng00j, Ste05b, Ste07b]. downturns [Gre01f]. DPUs [JAS^{+}22]. DRAF [GDN^{+}17]. Draft [Ste84e]. Drag [GGJ^{+}96]. drag-reducing [GGJ^{+}96]. Dragonfly [KDSA09, MMESGQ22]. DRAM [AMFFM^{+}16, BJO^{+}09, CBJ10, Dia96c, ERM08, GDN^{+}17, Ger19, HMAF90, Jac03, JMZ^{+}11, KOI95, LH12, NST97a, NLM^{+}19, NST97b, OGLG^{+}22, Sak97, SSY97, SZZ01, SLSO14, SKJ^{+}11, SSC^{+}22]. DRAM-Based [GDN^{+}17]. DRAMs [LCWB08]. Draw [SMR18]. Drawings [Sto90]. Dream [Kah93i]. Dreamcast [HO99a]. Dreaming [JP17]. Drive [Mye91c, Yi23c, Ano01c]. Driven [ACRV96, CSC^{+}22, DSK^{+}92, FK83, GLD^{+}22, Kha00, KSM^{+}89, KKP^{+}09, LYP^{+}18, SRU^{+}23, WMC^{+}06]. driver [FL84, IKK96]. Drives [Mye93c]. Driving [LCS92, TSV^{+}20, WHCK18, Bal84a, Wv92]. Droops [RGH^{+}10]. Drop [Gre07a]. Dropping [OESGG^{+}21, PPP01]. Drops [Ano97-32]. DSC [DKM^{+}92]. DSC-HDTV [DKM^{+}92]. DSP [BYM^{+}07, BGH^{+}90, CAV^{+}14, DMP91, DS95, Dur96, DM88a, DH90, FG00, Gol96, Joh89, KE89, KP90, LLW^{+}07, ME95, Roe86, Smi92]. DSP-Based [DMP91]. DSP32C [FGG^{+}88]. DSP56000 [Klo86]. DSPs [Ano98-30, DFR90, Lee90]. DTP [Ano94b]. Dual [KST04, MB05, JKP89]. Dual-Core [KST04, MB05]. dual-ported [JKP89]. Dual-Thread [MB05]. due [AVU^{+}08].

Dueling [QJP^{+}08]. Duplex [KG05]. Duration [IBM05]. During [All86b, GKS21]. duty [Mat96c]. DVDs [Ano96d]. DVFS [GKS21, IBM05]. DVL [OESGG^{+}21]. DVL-Lossy [OESGG^{+}21]. Dyn [Gre19a]. Dynamic [CL05, DMG^{+}15, GKS21, HRC^{+}23, HLIT20, KJMP07, MSA^{+}03, Mİ09, Tab84, WMC^{+}06, YAK18, HS92]. Dynamic-Compiler-Driven [WMC^{+}06]. Dynamically [BSC08, CO03, Dan89, SGG^{+}12]. Dynamics [GFL^{+}17]. DySER [GHN^{+}12]. Dystopian [Gre08b].

E-Booster [WHJ^{+}23]. E-business [Gre01b]. E-Commerce [SK01]. e-mail [Gre01a, Ste97a]. e6500 [BGH^{+}12]. earful [Gre01a]. Early [Gas21, Smi96a, Yi21b, Gre05b, Mar96]. Earning [Gre19c]. earnings [Ano03e]. Easier [WG92, Mat96d]. easily [KWGG95]. East [Sak91, Hoo90a, Kir90c, Sak89]. East-West [Kir90c]. Easy [MBA^{+}09, Dia95d, Pir97]. easy-to-use [Dia95d]. EasyRide [GD01]. ECC [YE11]. ECIM [LKJ^{+}22]. ECIX [Ano98u]. Eckert [Ano16c, Ano17g, Goo14, Hil19, Mud15, Wei17, LE18]. Eckert-Mauchly [Goo14]. ECL [BAC^{+}90, JBF94]. Economic [Gre08e, Gre09b, Gre16b, Gre17c, Gre20e, Gre21c, Ste09b]. Economics [Ano01a, Gre93, Gre95c, Gre95b, Gre95d, Gre96a, Gre96b, Gre96c, Gre96d, Gre96e, Gre97a, Gre97b, Gre97f, Gre97c, Gre97d, Gre97e, Gre98a, Gre98b, Gre98e, Gre98c, Gre98f, Gre99c, Gre99d, Gre99b, Gre99a, Gre99e, Gre99f, Gre00b, Gre00f, Gre00c, Gre00d, Gre00e, Gre00a, Gre01b, Gre01a, Gre01c, Gre01d, Gre01e, Gre01f, Gre02a, Gre02c, Gre02b, Gre02d, Gre02e, Gre02f, Gre03a, Gre03b, Gre03c, Gre03d, Gre03e, Gre04b, Gre04a, Gre04d, Gre04c, Gre04e, Gre04f, Gre05e, Gre05a, Gre05b, Gre05c, Gre05d, Gre05f, Gre06a, Gre06b, Gre06c, Gre06d, Gre06e, Gre06f, Gre07d, Gre07a, Gre07b, Gre07e, Gre07c, Gre07f, Gre08a, Gre08c, Gre08d, Gre08b, Gre08e, Gre09b, Gre09c, Gre09a, Gre09f]. Economics [Gre09e, Gre09d, Gre10c, Gre10d, Gre10f, Gre10e, Gre11c, Gre11d, Gre11f, Gre12a, Gre12b, Gre12c, Gre12d, Gre12e, Gre13b, Gre13c, Gre13d, Gre13e, Gre13f, Gre14a, Gre14c, Gre14d, Gre14e, Gre21a, Mat07b, WD03, WN92, Gre95a, Gre96f]. Economy [Gre20c, Gre01d]. Ecosystem [VMW^{+}19]. ECU [WHCK18]. EDA [Ano98h, STL92]. Edge [BCN^{+}22, BWBJ11, CSM^{+}21, DPBW19, Eec17b, FC22, GZC^{+}20, Gre10d, Joh22a, KNV^{+}20, LB00, MK22, NHY^{+}22, NHMM23, OGLG^{+}22, OLT^{+}23, PSL^{+}23, PSB^{+}20, PMR^{+}22, RDC98, SVA^{+}22, WL92, ZRB^{+}22, GP90, OZT^{+}22, Soo93, War91g, Ano19a, Ano19b, Ano19c, Ano20a, Ano20b, Ano23c, Ano23d, Ano24b, Ano24c]. Edge-Computing [CSM^{+}21]. Edges [Gre07c]. EDIF [Mar85]. Editor [Kir01, Sak99f, Sak01f, Tor06, Urq97, IA13, Red13, Alb04, Alb07d, Alb07e, Alb07b, Alb07a, Alb07c, Alb08, Alb09, Alb10a, Alt11a, Alt11b, Alt13c, Ano10a, Bos03b, Bos04b, Bos06e, Bos06d, Bos06c, Bos06a, Bos06b, Cas95, Cle00a, Cra00, Dia93f, Dia95c, Dia98, DH90, Emm08b, Gro92b, Gro94b, Gro02, Hoe93, Jag97, Kan95, Koo02, Lav02, Loc03, Lyl04, Mis93, Mud10, Nak99, Pen01, Rob98d, Sak90b, Sak91, Sak95, Sak99b, Sak99a, Sak99e, Sak99d, Sak99c, Sak00c, Sak00b, Sak00a, Sak00f, Sak02g, Trö98, Vei04]. Editor-in-Chief [Alb07d, Alb07b, Alb07a, Alb07c, Alt11a, Alt11b, Bos03b, Bos06e, Bos06d, Bos06c, Bos06a, Bos06b, Dia95c, Dia98, Sak99b, Sak99a, Sak99e, Sak99d, Sak99c, Sak00c, Sak00b, Sak00a]. Editorial [Alt14a, Ano97b, Ano98i, Ano99d, Ano00e, Eec16e]. Editors [AS91b, AKP96, AS05, ABZ08, AS95, AM08, ANS96, Ano23-76, Ano24y, AW10, AGJL98, ALGJ01, AJ83, BG16, BR10, BS98, BCP04, BBP09, BS84, BCA99, BAM03, CLM08, DTB01, DG89, Fag96, FL13, FG14, FD04, GS99, GR95a, HW91, Hoe92, IA09, IT15, JA96, JW99, KW02, KS07, KP07, LB00, LS96, LTL97, LK02, Mas93, MB99, MRLB03, OVT90, PNDG04, PLB06, PSP14, RDC98, RG07, Sak97, SVL03, SP92, SS06, SY06, SS05, TS13, UB05, VL00, VBB14, VN96, WD03, WG97, WT98, YT01]. Education [Ano18c, Cle00a, McK83, Nic91, Ano17k, Ano17j]. Educational [PJ91]. Edutainment [Sak99f]. EEMBC [PCLGO09]. Effect [TSMS23]. Effective [BMG^{+}21, BCC^{+}02, Das21, Far85, GH88, Laz89, Lea88, Mat11a, MSWP03, Mye84c, NRS^{+}08, SJL23, SMCT87, Yi22a, DS95, KSI^{+}96]. Effectiveness [Mat02c]. Effects [Gre23c, Ano02c, Zha91b]. Efficiency [CES17, ENSD03, HRC^{+}23]. Efficient [AMK17, AAG^{+}10, ARS03, BR21, BPT^{+}11, BNV^{+}15, BCH^{+}23, BvdGM^{+}15, CLL^{+}20, DSK^{+}92, Dea04, DWF^{+}21, FZW^{+}12, FHL^{+}03, GHS17, GQF^{+}06, GHN^{+}12, GHY^{+}17, GSS09, KSA^{+}19, KJMP07, KDSA09, KBN16, LSY01, LKJ^{+}22, LHC^{+}20, MLS^{+}16, MKG^{+}20, MH10, MBJ08, MKP06, NHMM23, PPA^{+}14, RTHA05, RSC^{+}06, RAG19, RBKL11, RPL^{+}17, SK12, SGP02, SLL^{+}18, SNM^{+}22, SO02, SRA^{+}04, STR^{+}13, SMR20, SKA14b, THP^{+}19, TNT06, UB05, VCE06, WSZS05, YKH^{+}19, YKL05, ZHR15, BG81, FL84, JKP89, Lee96, LHN95, Seg97, WN94]. Efficiently [AHKY19, CGLES^{+}23, Kra96, Yea96]. Efforts [VM88, Ano00g]. Eggers [LE18]. EIB [AP07]. EIC [Bos03c, Bos03d, Bos04c, Bos04d, Bos04e, Bos05a, Bos05b, Bos05c, Bos05e, Bos05d, Bos05f, Bos06f, Hoo91, Sak00d, Sak00e, Sak01c, Sak01a, Sak01b, Sak01d, Sak01e, Sak02c, Sak02b, Sak02d, Sak02e, Sak02a, Sak02f]. EICs [Ano01d]. Eight [FJL^{+}13, Ano03e]. Elastic [SABS20]. elect [Ano01d]. Election [Ano22-55].

Electric [Ano03b]. Electrical [Can98, HYS98, Lin92, Gre05f]. Electrical-Engineer [Lin92]. electroluminescent [Ano02b]. Electromigration [AVU^{+}08]. Electron [Ano97f, Ano98j, Ano02b]. Electronic [Alt98, Ano96e, Ano97e, Ano99i, HP85, Hoe93, KTC18, Lav02, Lea85, Mur03, SV03, SBE01, Sto94, WHCK18, Ano94b, Ste05a]. Electronic-System [SV03]. Electronics [Has94, Kir90c, Mac93, Nar19, SWM^{+}20, Ste92a, ZP93]. Electronics-Industry [Ste92a]. electropolitics [Has85]. Elegance [Moo03, Moo04a]. Element [ASD^{+}05, KNN^{+}90, NBM^{+}06, PPA^{+}14, TCD^{+}05]. elephants [Ste99e]. Eligible [Ste08d, Ste08e]. eliminate [Joh90b]. Eliminating [TT12]. Elusive [GKL^{+}22]. Embedded [AB14, Ano01a, ASD^{+}05, AGJL98, ALGJ01, BCP04, Ber09, BFLS01, BGH^{+}12, Cas95, CRV^{+}04, CR95b, CGJ^{+}94, Cum04, Dra00, EVM^{+}98, Fre02, FSH^{+}01, GALB07, GH88, GAAR88, HC02, KMN^{+}04, KG05, Koo02, KP03, LC09, LBR^{+}22, MM23, MGG^{+}19, Mon97, NKI^{+}09, PO04, PV98, PV01, PGL97, RCR04, Rea86, RSE01, SHTE08, STT^{+}15, SK02, SSY97, SCG95, SM00, SANK98, TKI^{+}14, WHP^{+}13, Ano01g, Bos04b, Cat88, DS95, ESW97, Fly97, ME95, PK88, Rob91, Rya88, TS95, Eng00f]. Embedded-Systems [SK02]. Embedding [AO97]. embodied [Ste99a, Ste99b]. Emergent [RNN^{+}16]. Emerging [Ano14s, BC21, CPS^{+}18, Che19, JLG19, JC08b, Joh19a, Joh23a, KK23, SMAS16, XZ19]. Emissions [SSK23]. emitting [Ano02c]. EMMA2 [ACLR89]. Emotion [KIS^{+}00, OS99]. Emphasizing [Yea96]. Empirical [SB00]. Employing [WHP^{+}13]. Empowering [DPY18, ZFW^{+}23]. EMT [Noh19]. EMU10K1 [Sav99b]. Emulating [MM87]. Emulation [HWG^{+}09, JLG19, LP21, Has85]. Emulators [Ste88b]. Enable [HDG^{+}22, Mye84a, MKRC97]. Enabled [ASK^{+}15, AKJF22, DJUH16, RAA^{+}21, Sak01a, BWR23]. Enabler [ACDG99]. Enabling [AK24, BDH^{+}16, CWLS15, Fly97, JUP^{+}22, Joh23a, MM09, Sha23a, SYG^{+}20, YKG18, KMPS06]. Enacts [Cha85b]. Encoder [HSR18, IKN^{+}99, KSI^{+}96]. Encoding [KGT22]. Encrypted [FSK^{+}22]. encrypting [KAK96]. Encryption [AAC^{+}16, Ano97d, Kal93, LJM^{+}23, WHJ^{+}23, ZHZ^{+}19]. encyclopedia [Ano92f]. End [DDG^{+}19, DM88b, EBS^{+}12, HcF04, Kir91b, MD88, NAA^{+}20, OW01, PNDG04, SHTE08, SJFM19, Sla91a, Ste09c, VC11, WH09, YMC^{+}12, Mat05e, WHKM93a, WHKM93b]. End-to-End [HcF04, SJFM19, YMC^{+}12]. end-user [WHKM93a, WHKM93b]. Endian [Gus85, Jam90]. endings [Sak01c]. endpoint [Gal97]. Ends [Kah93c, Ste12]. Energize [JN21a]. Energy [AAG^{+}10, Alt12d, BCH^{+}23, CES17, CHSL17, FAWR^{+}11, FHL^{+}03, GKL^{+}14, GHN^{+}12, GSS09, HCP^{+}03, HKC10, IO16, JGC^{+}11, KST12, KSA^{+}19, KJMP07, KBN16, LDL17, LKJ^{+}22, LDF^{+}13, LHC^{+}20, LZX^{+}18, LLZ^{+}04, LS98b, MLL^{+}15, MLL^{+}18, MT05, MMB^{+}08, MH10, NHMM23, PDL08, RES^{+}13, RSC^{+}06, RAG19, RBKL11, RPL^{+}17, SW14, SCA^{+}12, SP01, SLL^{+}18, STR^{+}13, TSS18, TVT19, UB05, WLD15, WB12, WA13, WMC^{+}06, ZHR15, BM19]. Energy-Aware [Alt12d, CHSL17, JGC^{+}11, WB12]. Energy-Constrained [MLL^{+}18]. Energy-Efficient [AAG^{+}10, FHL^{+}03, GHN^{+}12, GSS09, KSA^{+}19, KJMP07, KBN16, LKJ^{+}22, LHC^{+}20, MH10, NHMM23, RSC^{+}06, RAG19, RBKL11, RPL^{+}17, SLL^{+}18, STR^{+}13, UB05, ZHR15]. Energy-Error [LZX^{+}18]. Energy-Harvesting [MLL^{+}15]. Energy-Neutral [IO16]. Energy-Secure [BM19]. Enforced [NMZ13]. Enforcement [LPM15]. Engine [ANC05, EPZ02, Har12, KSLY17, RMM^{+}04, SK02, OS99, Sel18, dCMA22]. Engineer [Lin92, WG92]. Engineering [Ano14-34, Ano18t, Ano21w, Ano23j, BFK^{+}85, Buc84, Hig85, HKM^{+}85, KHHR85, Kni85, KKSV10, Lan85b, MBP^{+}85, MCF^{+}85, Ste86c, Ste86d, Ste92e, Ano92e, Ste93g, Wil95b, Ano21a, Ano21b, Ano21c, Ano22c, Ano22d, Ano22e, Ano23j, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24d, Ano24a]. Engineers [Ano98q, Mat90c, Ste92b, Ano94b]. Engines [FTKS92, FGC^{+}14, Joh89, Ste04a, SCH^{+}23, Ste02a, Ste04b]. English [Pfa94]. Enhanced [BWR23, NHY^{+}22, Lee95]. Enhancing [FCY^{+}20, TONH96]. Enjoy [Joh20d]. Enough [Gre14b]. Ensuring [ZG96, Wal97]. Enterprise [BJW^{+}23, Mat02a, dCMA22]. Enterprise-Class [BJW^{+}23]. Entertainment [HO99a, Sak99f]. Entrepreneur [Ano16f, Gre11f]. Entrepreneurial [Emm07c]. Entrepreneurs [Ano98k]. Entrepreneurship [Gre06d]. entries [Dv87]. entry [Abr83]. Envelope [Cha98]. Environment [BGH^{+}90, DMG00, FKL01, LBR^{+}22, Mat89b, MWM99, SKM23, Yao85, vW85, AGH^{+}91]. Environmental [GKL^{+}22]. Environmentally [Joh23b, Wu23]. Environments [KG05, LRC^{+}09, MSB^{+}17, NG87, SKA^{+}14a, Ven23]. EOLE [PS15]. EOS [CR95a]. EPIC [Ano03f]. Episode [Ste97d]. Epsilon [Ano17-58, Ano17-59]. ePub [Ano14f]. Equalization [DP97]. equations [KE89]. Equipment [HOHCV99]. equity [Ste94d]. Era [Ano17h, ANM^{+}12, BVZ^{+}08, DM88b, DM88a, Gre03a, Gur09, HAB^{+}09, Joh19b, KCXmWH17, MD88, ND10, VDC17, Bos05d, Gre00f, Gre05a]. Eras [Hen21a]. Ericsson [Ano98f]. ERIDANUS [AHKY19]. Erratum [Ano09d, Ano18d].

Error [Gre03a, LZX^{+}18, MBS08, RTHA05, SC24, SGK^{+}04, SS16, SMS13, TK21, WEMR04, ZLTW13, Mat96f]. error-prone [Mat96f]. Error-Rate [TK21]. Errors [Ano01a, EDL^{+}04, Gre01c, KGDW^{+}13, NRV^{+}06, SWK^{+}05, SNC^{+}07, Sha82, SSC^{+}22]. ES/9000 [SGC94]. ESDI [Ano88e]. especially [Ano94c]. Esperanto [DtEt22]. ESPRIT [Ang90, RD90]. ESSA [BM19]. Essential [Ste09b, Ste97a]. Establishing [War89a]. Estimation [IPL^{+}23, TK21]. ET-SoC-1 [DtEt22]. ETA [RMM^{+}04]. etching [Ano01c]. Ethernet [BcFP06, Gad07, HcF04, MD20, RSW10]. Ethernot [BcFP06]. Ethics [Mat13a, Ste90b, Has85]. eTRON [SK01]. EU [Ano03b]. Eudora [Ano94a]. Euler [KE89]. Europe [Ano99o, Hoe92, Kir88a, Kir88b, Kir89a, Kir89b, Kir91a, OVT90, VN10]. European [Ano10c, DG87, DG88, DGT89, DG89, GS99, HLHR90, Kir87, Kir92, LCS92, McL87, MC90, MBS92, STL92]. Evaluate [FHP00]. Evaluating [FJB^{+}22, Gil96a, LMVP05, SMS13, VPV12, MC87, War91g]. Evaluation [AS22, CJ85, DVWW05, Eec15d, GBW^{+}23, JYPP18, KW83, LYBZ04, TMA18, VW03, Ano18d]. Event [CF90, FBHN04, RGH^{+}10, SVA^{+}22]. Event-Based [SVA^{+}22]. Events [Kir85b, Ste03b]. Ever [Ano88d, Joh19c, Joh20b, Joh21a, Joh23a]. everyone [Gre95d]. Everything [Ano98n, Ste97b]. Everywhere [SV21]. Evolution [Alt12b, DKK21, DF01, DOH94, Mat03a, ZRA^{+}17, NM96, Sak01a, Eec15f]. Evolutionary [JC08a, AKK^{+}93]. evolve [Ano94b]. Evolving [Ano20y, Sla91b, Gon97]. Exa [TSW^{+}23]. Exa-Scale [TSW^{+}23]. Exact [KMK^{+}19, Mey04]. Examiners [Emm06c]. Example [Ste86d]. Examples [Kir87, RT23]. Exascale [KNB14, SIL^{+}15]. Excellence [Ano17-58, Ano17-59, Ste85h]. exceptions [Iac88]. Excitement [Smo88a]. Exciting [Pri93a]. exclusion [OL85]. exclusion/synchronization [OL85]. Execute [HKS16]. executes [FBGB96]. executing [Cra90]. Execution [BCP01, CK11, KMPS06, LBR^{+}22, MSWP03, MKP06, NPC06, RG03, RNLY23, SMQP10, TK21, UCS^{+}10, VCS^{+}19, Ven23, ERPR95]. Executive [Cro85, FK83, Hea84]. Exhaustion [Ste92f, Ste07d, Ste08b]. ExHero [TK21]. Existential [Emm08b]. Existentialist [Gre15b]. existing [NM96]. Exotic [Raj94]. Expander [KKS^{+}23]. expanders [Gre05b]. Expanding [Emm07a, GR95a, NCT^{+}98]. expands [Ano00g, Ano02c]. Expansion [Ano84, Ano02b]. expensive [Ano02d, Ste99d]. Experience [RMM^{+}04, CCD^{+}82]. Experiences [GLN^{+}08]. Experiment [Lin06]. Experimental [DMWS13, SWK^{+}05]. Experimentation [FTKS92]. Experimenting [Ano87g]. expert [KKT^{+}91]. Expertise [Gre20a, Mat83]. Experts [Ano15-34, Ano16-48, Ano16-47, Ano16-46, Ano92d, Ano16-45]. explain [Gre97a]. Explained [Mat99a]. Explaining [Ano01a, Gre01c]. Explicit [KPK^{+}10, NGS16]. Explicitly [AAP^{+}10]. Exploit [TLM19]. Exploitation [HIP^{+}22]. Exploiting [Alt13d, AML05, CEAY23, DMMD11, DJUH16, EV97, KJL16, LDL17, Rob98b, SWG06, SNL^{+}03, SW14, SPH^{+}03, FMT91]. Exploration [DGR^{+}10, MLL^{+}15, MWM99, PLBC09, RCR04, TSMS23, IKK96]. explore [Ano02b]. Explorers [Gre05b]. explores [Eng00j]. Exploring [FZW^{+}12, PSL^{+}23, SL97, ZIM^{+}07]. Expo [Mat88, Mat99c]. Exponent [LKJ^{+}22]. Exponential [Ano96f]. exponentiation [KAK96]. Exposed [TATC09]. Exposing [MFM02, TT12]. Express [Das21, KKP^{+}09, KPKJ08, OKN^{+}11, Sha22, Sha23a, Sha23b, Sha23c, SC24, LMVP05, ZCW^{+}14]. extend [Mat96f]. Extended [EKMW02, HDG^{+}22, PSL^{+}23]. Extending [Cha98, GBD^{+}20, Han96, Ano81]. extends [Ano02c]. Extensible [Gon00, Pap89]. Extension [DDHS00, GSC97, PW96, SBB^{+}17]. Extensions [NDR^{+}22, RPK00, VCS^{+}19, YHHF20, Lee96]. Extraction [CJH^{+}12, LPC12]. Extraordinary [GR95b]. Extreme [Ano96l, Ano97-30, Lin06, PMS23, SGL93, Ano01f, Mat99a]. Extreme-Ultraviolet [Ano96l, Ano97-30]. Extremely [MLL^{+}18, MH10]. Eye [YZW^{+}23]. EyeCoD [YZW^{+}23]. eyes [Wea97b].

Fab [Eng00h]. Fabric [CEH^{+}12, DXT^{+}18, GDN^{+}17, PCC^{+}15, TKM^{+}02, WGM02]. Fabrics [CNC^{+}16]. FabScalar [CWS^{+}12]. Face [BCKY17, WD03]. Face-Recognition [BCKY17]. Faces [Hur98, Mye91a]. Facilities [JGC^{+}11]. Facility [BO86, RG85]. Facing [KML04]. Facto [Hec83b, Pri94a]. Factor [ZES13, Mat96c]. Factors [Min84, MWE^{+}03]. factory [DM86]. Facts [Emm07a]. Failings [Sla90b]. Failure [KJC^{+}23, YBNS15]. Fair [Dia93b, MM09, PPBS03, PPP01, ZL15]. Fall [Gre02e, Kir90a]. Fallacy [GMM^{+}07]. Falling [Gre00c]. FaM1ly [Mat22]. Families [Yi24b, Yi24a]. Family [Als90, Bos21, BvdGM^{+}15, Mel89, OS08, Yeh07, OA81, PK88]. Famous [Gre04f]. Far [Hoo90a, Sak89, Sak91]. Far-East [Hoo90a, Sak89]. Fare [GD01]. Farewell [Sak02b]. Fast [CS14, CLMY96, DXT^{+}18, GG99, GKA^{+}16, Gre14d, GM99, LSY01, Mae87, OW01, RPE10, SG01b, WNW^{+}16, ZZY97, Abr83, DVQ96, Gre95d, Rob97d, AAG^{+}10, AH96, LNV89]. fast-track [Rob97d]. Faster [Ano01h, Eng00p, Mye93a, Sla90f]. fastest [Ano00g]. Fat [GCL^{+}20, MMESG^{+}20, VJFG17]. father [Dan96]. Fault [AF84, AGJL98, ALGJ01, BDSC21, CK98, Dra00, EVM^{+}98, EM84, FKL01, GCL^{+}20, GSVP03, GV06, Gre14d, Gro94a, Gro94b, Hum84, IEB^{+}14, JKN96, Joh84, KLD^{+}94, Kir87, Kir89a, KDK^{+}89, MS84, Pow94, PC01, Rag84, RSS^{+}08, RSE01, SB84, SKA^{+}14a, Sos94, SGC94, Str98, YW94, YNS^{+}14, YW88, AGH^{+}91, DGW^{+}94, OFG88, WJR88]. Fault-Handling [KLD^{+}94]. Fault-Tolerance [Pow94]. Fault-Tolerant [AF84, AGJL98, ALGJ01, BDSC21, CK98, EVM^{+}98, IEB^{+}14, Joh84, Kir89a, KDK^{+}89, RSS^{+}08, RSE01, SB84, SKA^{+}14a, SGC94, Str98, YW94, YNS^{+}14, YW88, JKN96, PC01, AGH^{+}91, DGW^{+}94, WJR88]. Faults [HANR13]. Faulty [AFGM10]. Favorite [Bon21, You21]. FCRAM [Ano01h]. FDDI [Jos86]. FDIV [Pri95]. FDX [SDF^{+}23]. Feasibility [AAC^{+}16]. Feast [Eec16a]. Feat [Joh23e]. Feature [RGR95, SRL91, Bor85b]. Featured [Ano20-71]. Features [Ano97-29, AAD^{+}93, FAWR^{+}11, FMN^{+}13, Spr02b, Mat96f]. Featuring [KHS^{+}23]. Federal [Ste07e, Ste06b, Ano98v, Ste07c, Ste15a]. Federated [KGT22]. Feel [Ste86f, Ste93c]. Feet [Sla90d]. Fermi [WKP11]. Fermtor [RLV85]. Fernbach [Ano17-45]. ferroelectric [DTH^{+}95]. FerroElectronics [KNV^{+}20]. Fertility [GAGV22]. Few [RT23]. FFT [Bus86, Mor86b, RFGM86, SZH82, VS87]. Fi [Gre11d]. Fiber [EKB^{+}96, Jos86, Eng00j]. Fiber-Optic [EKB^{+}96]. Field [AB14, ABG^{+}16, Alt14e, Ano87e, Eec15e, Ham00, KHS^{+}23, LJM^{+}23, Ste86a, Sti11, WHJ^{+}23]. Field-Programmable [AB14, ABG^{+}16, Ham00, KHS^{+}23, Ste86a, Sti11, WHJ^{+}23]. Field-Tests [Ano87e]. Fifty [SRU^{+}23]. Fighting [Edw83]. Figure [LKM92]. Figure-ground [LKM92]. file [Emm05b, JRHM86, Mel87, Swa19]. Filed [Ste09a]. Filers [KSR^{+}99]. Filing [Emm06f]. Filled [Sak93]. Film [KCHA21a, KCHA21b, Gre98c]. Filter [CPH90, NN81a]. filtering [NN81b]. Filters [DKSL04, LK10]. final [Pap96]. FinalFilter [SHKS19]. Finally [Noh19]. Finding [DWLN20, Ste07e]. Fine [AS91a, BYM^{+}07, BSP^{+}17, CBJ10, Dea04, HIP^{+}22, MM23, SK12]. Fine-Grain [AS91a]. Fine-Grained [BYM^{+}07, BSP^{+}17, CBJ10, Dea04, HIP^{+}22, SK12]. Fine-Tuning [MM23]. Finesses [Ste93a]. Fingerprinting [SGK^{+}04]. Fingertips [Gre20a]. Finite [LJM^{+}23]. FINN [JUP^{+}22]. Fireplane [Cha02]. Fireside [Mat95b]. FireSim [BMK^{+}21a, BMK^{+}21b, KMK^{+}19]. Firewalls [KTC18]. Firmware [War92a, TZMVLN81]. First [BH15, BY07, BBTV15, Dia99, JYPP18, Lie23, Mas21, McD21, MW19, Rag21, SNM^{+}13, XCZ^{+}21, Fly97, Ste91h, KB13, Ste90g, Ste90h]. First-Generation [BH15]. first-time-right [Fly97]. Fish [CDBY23]. Fisher [Bel13]. Fit [Ano16x, Ano16-33, Ano16-34, Ano17-43, Ano17-40, Ano17-44, Ano17-41, Ano17-42, Hen21b, Ano03e, Ano17f, Ano17e, Ano17c, Ano17d]. Five [Emm06b, Hen21a, SVC01, KAK96]. Five-Qubit [SVC01]. Fixed [Ano02e, Joh89]. Fixed-Point [Joh89]. Fixing [Ste15b]. Flash [AS10]. Flat [ZBES15]. FlatCam [YZW^{+}23]. FlatCam-Based [YZW^{+}23]. Flaw [Pri95]. flaws [Ano17u]. Fletcher [Dia96a]. Flexibility [FPAF02]. Flexible [BMG^{+}21, CKG^{+}09, CS14, DSG^{+}22, EEJ95, KSV^{+}21, KSK18, MCV^{+}19, YNS^{+}14, YE11, BCF^{+}92]. Floating [BSC^{+}90, CCG^{+}84, DKB^{+}90, DM88a, FGG^{+}88, GE86, HC83b, LKJ^{+}22, MD88, PS88, RJR88, SKL^{+}92, SK88, Ste84e, ZSB21, Iac88, KWM89, OZT^{+}22, SL97, DM88b]. Floating-Point [BSC^{+}90, CCG^{+}84, DKB^{+}90, DM88a, FGG^{+}88, GE86, HC83b, LKJ^{+}22, MD88, PS88, RJR88, SKL^{+}92, SK88, Ste84e, ZSB21, Iac88, KWM89, OZT^{+}22, SL97, DM88b]. Floppy [MA83]. Flow [LPC12, MEB^{+}20, NHY^{+}22, SL03, SRA^{+}04, TLW^{+}10, WPSR20, IWM89, PVYU94]. Flowers [Gre06e]. Flows [OESGG^{+}21]. Fluctuations [KJP^{+}13]. Fluorophore [LDL17]. fly [Sho85]. Flying [Chr96, GZC^{+}17]. FM9801 [HS99]. Focus [Ano14g, Ano14h, Ano15k, Ano15l, Ano16i, Ano16j, Ano16h, Ano16g, Ano17q, EHP^{+}07, Jos86]. Focusing [PBT06]. Footprint [GKL^{+}22, GEH^{+}23]. Force [LCS92, Wv92]. Forces [SKL^{+}92, Ano97p, Sak99e]. Forecasting [Gre99a]. Forecasts [Eng00n]. foreclosure [Gre98a]. Foreign [Kar88b]. Foreshadow [VMW^{+}19]. foresight [Gre05e]. forever [Gre95a]. Forewords [Mat12a]. Forget [Ber81]. Form [ZES13, Ano01f, Ano03c]. Formal [Rob00a, WJM^{+}05]. Format [Kir83b, Kir84a, Pat84, Ano83, Dia94a, Gre06b]. formats [KS00]. Forming [Upd93]. forms [Ste90e]. Fortran [Cro85]. Forum [Lan85a, Ste96f]. Forward [Alb10b, Ano15v, Bos03d, Eec16c, Eec17d, Joh21c, Mat98a, Mor86a, SC24, Bos06b]. Forwarding [ANC05]. Foul [Dia93b]. Foundation [IJ98]. Foundations [YHHF20]. founder [Sla96]. Foundry [Ste93a]. fountains [Ano92a]. Four [AML^{+}03, Ano17-58, Ano17-59, Gre06c, HSNJ21, TO96]. four-issue [TO96]. Four-Terabit [AML^{+}03]. Fourier [AAG^{+}10]. Fourth [HMB^{+}14]. Fourth-Generation [HMB^{+}14]. Fox [Gre20b]. FPGA [ANJ^{+}04, BFZ^{+}22, BMK^{+}21a, BMK^{+}21b, CS08, GS21, JLG19, JUP^{+}22, KMK^{+}19, KSE^{+}22, Kur21b, LNLG20, LP21, Man09, MSB^{+}17, PMM15, SAC^{+}21, TAI^{+}21, TES^{+}18, WFW^{+}21, XLX^{+}23].

FPGA-Accelerated [KMK^{+}19, LP21]. FPGA-Based [BFZ^{+}22, CS08, KSE^{+}22, LNLG20, SAC^{+}21, XLX^{+}23]. FPGAs [ABA^{+}21, ATS^{+}22, CFZ^{+}99, FDS^{+}17, GALB07, MEB^{+}20, Mye93a, OML^{+}07, SJFM19, TGC^{+}20]. fps [KII09]. FR500 [SM00]. Frame [KHS^{+}23]. Framemaker [Mat93a, Mat97c]. Framemaker-5.5 [Mat97c]. Framework [ABG^{+}20, BCN^{+}22, HLS^{+}21, JPOB20, KGT22, LYBZ04, MHW03, PBFC21, SJFM19, TGC^{+}20, WDK^{+}20]. Frameworks [Ano17l]. Framing [Ste89b]. FRAND [Ste13]. Fraud [Ste91d]. Free [CML^{+}23, Gre17f, Gre18b, Mey04, SO02, Ano01h, YMC^{+}12]. Free-p [YMC^{+}12]. Freedom [RGK19]. Freescale [BGH^{+}12]. French [Kir90b]. frequencies [SLM^{+}97]. Frequency [GKS21, Lin98, MSA^{+}03, RMC04, Sak01f, SBE01, SBJ13, RLG94]. Friend [Ano89]. Friendly [Yao85]. Friends [Mye84d]. FRM [KKY88]. Front [Ano13f, Ano14i, Ano14k, Ano14l, Ano14m, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano16o, Ano16k, Ano16l, Ano16m, Ano16n, Ano17r, Ano17s, Ano17t, Ano18j, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano19q, Ano19r, Ano19s, Ano19t, Ano20-31, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano24k, Ano24l, NAA^{+}20, OW01, Ano14j]. Front-End [NAA^{+}20, OW01]. Frontier [GBD^{+}20, Gre19f, Lav02, LZX^{+}18, MGP21]. FSK [SZP81]. FT [CWL^{+}14]. FT-Matrix [CWL^{+}14]. FT64 [WWZ^{+}08]. FTC [Ano99j, Ste02b, Ste04d, Ste05b, Ste08a, Ste08c, Ste17c, Ste17a, Ste17b, Ste18]. FTT [FPAF02]. FTT-CAN [FPAF02]. Fugaku [SKTO22]. Fuji [Mat04e]. Fujitsu [Ano03c, MAT^{+}18, YMA^{+}13, YHT^{+}15]. fulfilled [Mar96]. Full [BFZ^{+}22, KIM^{+}09, MK10, PRE11, RAA^{+}21, RPE10, TSV^{+}20, TES^{+}18, TML^{+}18, YKG18]. Full-Stack [RAA^{+}21, TML^{+}18]. Full-System [PRE11, RPE10]. Full-Throttle [MK10]. Fully [SsSMB24]. Fun [Ful91, Gre97e]. Function [BR21, Lan96, Ste84d, THC18, Vac87, Boa96, Dia93d, KKY88, LC91]. Functional [BCU^{+}99, NMU^{+}15, YNS^{+}14, AH96, WHKM93a]. Functionality [GHN^{+}12, Ste91f, Bos05a]. Functions [BWR23, KSWM90, KHS^{+}23]. Fund [Ano23-43, Ano23-45, Ano22-53, Ano22-54, Ano23-44, Ano23-46, Ano24q, Ano24r]. Funding [Gre14e, Upd93]. Furnace [HOHCV99]. Further [Ste85b, Ste87c]. Fused [NY22]. Fusing [SDG^{+}21]. Fusion [BFS12]. Futile [Mat17]. Future [Alb10a, Alt11b, Alt11e, Bor99a, Che19, Cla03, Eec15a, Fra96, Gon97, GHSV^{+}11, Gri21, GSLK11, HLZ^{+}16, HSW98, HBE^{+}10, Hoo89b, HRSS11, JUP^{+}22, JN21a, KDK^{+}11, KKD^{+}07, Kir85a, Kni85, KKS^{+}98, LZY^{+}10, Lee24a, Mat15a, MCM^{+}16, MB15, NM99, Nar19, NFQ03, PNDG04, Sak87a, Sak00f, Sak01e, Sak02c, Smo87a, Urq97, War92c, WS90, Yu96, Ano94b, Ano03e, BCF^{+}92, Dia96c, Kah91d, Mar96, Mat04b, Mat06d, Ste93g, TW00, Wea97b]. Future-Directions [Kni85]. Futurebus [SRL91, Ano91c, Bal84a, Bea90, BT84, PH91, Tau84, Tau87]. Fuzziness [Ste95a]. Fuzzy [ACG^{+}95, ACRV96, CMR97, CR95b, CS08, CDGO97, EKM^{+}95, FBGB96, GG99, GTF97, GR95a, Hun95, JBM95, Kah91e, KAC^{+}95, KKL^{+}09, MY95, NSN^{+}93, Pea95, RGF96, San97b, SU95, TTF96, TCF96, VM95, Ano95d, GP95, Kan95, LGJ95, MM96, PHC95, RGF95, VVRV95, dG95]. Fuzzy-Logic [Pea95]. Fuzzy-Logic-Based [TCF96]. Fuzzy-rule-based [SU95]. Fuzzy/Neural [San97b]. FX [CHH^{+}98].

G [DRM^{+}23]. G5 [SAC^{+}99]. GaAs [NG87, VM88]. Gabriel [BGH^{+}90, Dwa18]. gains [Hsi91]. Galapagos [TES^{+}18]. GALS [BMG^{+}21]. Game [ML21, Ste92c, LNV82]. Games [Ful91, Ste89b]. Gaming [Gre13a, Ano03d]. Gap [BcFP06, SKM23]. Garbage [MAK19]. Gas [Ano02c, Ano02b]. Gate [AB14, ABG^{+}16, KHS^{+}23, Sti11, TLW^{+}10, WHJ^{+}23]. Gate-Level [TLW^{+}10]. Gatekeeping [Gre10c]. gates [ACRV96, Gre08c, Gre08d, Mat96b, Ste94e]. gathering [Boa96]. Gating [CK11]. Gatoring [Ste02c]. Gaudi [MD20]. Gauges [PC93]. Gbps [DP97, GDES08, PDT98, ZACM14]. GDP [Gre17f]. gears [Ano03c]. Geek [Mat10a]. GeForce [MM05]. GEMM [WLY^{+}21b]. Gen [PSB^{+}20]. Gene [CEH^{+}12, HOF^{+}12, SWG06]. Gene/Q [CEH^{+}12, HOF^{+}12]. General [Bos04e, DWLN20, ESG^{+}05, ED18, EKM^{+}95, ESCB13, Gil82, LLT^{+}08, PC01, SSMI87, STS^{+}92, TKM^{+}02, ZQL^{+}04, Han96, SU95, Ste84a]. General-Purpose [DWLN20, ESG^{+}05, EKM^{+}95, ESCB13, Gil82, LLT^{+}08, STS^{+}92, TKM^{+}02, Bos04e, Han96]. Generalized [KJMP07, ZRA^{+}20]. Generally [NGSW17]. generates [Ano02d]. Generating [JLWL20, PV98]. Generation [AJK^{+}15, AS90, Ano87a, AFK^{+}19, BH15, BBS^{+}00, BML^{+}21, DKyL^{+}17, ESG^{+}05, EEL^{+}97, EBC22, FGG^{+}88, HMB^{+}14, Hol98, HL99, JPOB20, Kah91a, KSSF10, KJP^{+}13, Maj87, MYK^{+}10, PBFC21, SBJ13, SGC^{+}16, SSR21, TIT^{+}13, VE10, Web08, YMA^{+}13, YHT^{+}15, Ano01e, Ano02b, Dia96d, KHF86, Mye92c, Smo87c]. Generator [BCC^{+}00, KW81]. Generic [Tua99, WN94]. Genie [Ste92c]. Genome [ABC^{+}20]. Genomic [HLS^{+}21]. Genomics [TBD19]. Geoscience [LCP^{+}11]. Get [Ano96q, Ano98t, Ano15s, Ano16p, Ano20-34, Ano20-32, Ano20-33, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano23-34, Ano24m, Mye83a, Mye93a, Ano95c]. gets [Ste99d]. Getting [Joh19c, Moo04b]. GF100 [WKP11]. ghost [FS05, Ste05d]. GHz [Ano87d, Ano98s, Ano01c, Ano03b, Ano03c, HVS^{+}07]. Gigabit [BCF^{+}95, Gad07, HcF04]. Gigabit-per-Second [BCF^{+}95]. Gigahertz [HDM^{+}98]. GigaRing [Sco96]. Gigascale [Mei03]. give [Rob98c]. Given [Dwa18, KT14, Mar17, Ste16, Sco14]. Gives [Ste07a]. Giving [PAC^{+}14, Ste89b]. Glen [MC90]. glimpse [Kah91d]. Glitches [Ste93e]. Global [KKP^{+}09, NS05, Dia95e]. globalization [Mat05a, Pir97]. Gmicro [IKNS88, KS90, UAN^{+}93, YSMH91]. Gmicro/100 [YSMH91]. Gmicro/200 [IKNS88]. Gmicro/300 [KS90]. Gmicro/500 [UAN^{+}93]. Gnat [Ste98c]. Go [CB10, Gre03e, Ano14-38, Ano14-39, Ano15-41, Ano17-55]. goals [Ano17f, Ano17e, Ano17c, Ano17d, Pap96]. Godson [FZW^{+}12, HWG^{+}09]. Godson-3 [HWG^{+}09]. Godson-T [FZW^{+}12]. Going [Alt13a, Mat05b, Nar19, Ste91g, Ano94b, Mat03f]. Gold [Gre23a, Gre24a, Kir89c]. Golden [CML^{+}23, DPY18, Mas21]. Golden-Free [CML^{+}23]. Goldstrike [BH15]. Good [Alt14b, Hau88a, Mor86b, RFGM86, SRJ^{+}91, Joh90b, Rob00e]. Good-Bye [Alt14b]. Goodbye [Eec18c]. Goods [Gre13c]. Google [BDH03, Gre09c, Gre10f, Gre22c, NPY^{+}21, RTM^{+}10]. Google-Wide [RTM^{+}10]. Gordon [CGS10, Gre15f, Mye84d]. Got [Smo87d, Ano17u]. GPS [Eng00j]. GPU [ANM^{+}12, Bur20, CWLS15, CGG^{+}21, Cho23, DKK21, FD17, FSBA12, LSL^{+}15, ND10, PEZ^{+}19, RTJ20, RTJ21, RBKL11, SSF^{+}14, SCYY11, SSC^{+}22, SKC^{+}23, TAI^{+}21, VPV12, VC11, WAA^{+}21, WKP11]. GPU-Aware [SKC^{+}23]. GPU-FPGA [TAI^{+}21]. GPUs [AMK17, AKT^{+}18, Alt11b, Bro11, FDS^{+}17, KDK^{+}11, Kur21a, LJM^{+}23, ZZNT^{+}23]. Grain [AS91a, CSL^{+}06, CKG^{+}09]. Grained [AKAK^{+}18, BYM^{+}07, BSP^{+}17, BDV^{+}08, CBJ10, Dea04, HIP^{+}22, LPC12, RAG19, SK12, SKM^{+}16]. Grandmaster [hHH99]. Graph [AAC^{+}23, AMK17, Ano17l, MCV^{+}14, OYK^{+}17, RBB21, CK95]. Graph-Based [RBB21]. Graphics [Ano20g, Ano20e, Ano20f, Ano21k, Ano22l, Ano22m, AOYS95, DKK21, Han87, Joh89, KBN16, LNOM08, MMG^{+}99, UBH^{+}94, Pri90]. Graphs [ED18, ZRA^{+}20]. Grateful [Alt14b, Eec18c]. Gravitating [NBS^{+}18]. Gray [BUMV95]. Gray-Scale [BUMV95]. Great [All86b]. Greater [Ste91a]. Greater-Than-Software [Ste91a]. Green [Mat09d]. GreenDroid [GHSV^{+}11]. grew [Rob99e]. Griffin [OS08]. Ground [Alb07d, LKM92]. Group [Ste84e, JKN96, Rob00b, WWR97]. Groups [Smo88c, Rob01d]. Grow [Ano00o, Eng00l]. Growth [Ano88b, Eng00n, Gre16b, Gre22d, IJ98, Kah91b, Ano02c, Hsi91]. GT [Das21, Sha22]. GT/s [Das21, Sha22]. Guardband [LDF^{+}13]. guardedly [Ste93d]. Guest [IA13, Red13, Sak99f, Sak01f, Urq97, AS91b, AKP96, AS05, ABZ08, Alb04, AS95, AM08, ANS96, AW10, AGJL98, ALGJ01, AJ83, BG16, BR10, BS98, BCP04, BBP09, BS84, BCA99, BAM03, Cas95, CLM08, Cle00a, Cra00, DTB01, DG89, Dia93f, DH90, Emm08b, Fag96, FL13, FG14, FD04, GS99, GR95a, Gro92b, Gro94b, Gro02, HW91, Hoe93, Hoe92, IA09, IT15, Jag97, JA96, JW99, Kan95, Koo02, KW02, KS07, KP07, LB00, Lav02, LS96, LTL97, LK02, Loc03, Lyl04, Mas93, MB99, Mis93, MRLB03, Mud10, Nak99, OVT90, PNDG04, Pen01, PLB06, PSP14, RDC98, Rob98d, RG07, Sak90b, Sak91, Sak95, Sak97, Sak00f, Sak02g, SVL03, SP92, SS06, SY06, SS05, TS13, Tor06, Trö98, UB05, VL00, VBB14, Vei04]. Guest [VN96, WD03, WG97, WT98, YT01].

Guidance [Mat21a, NNS^{+}93]. Guide [Ano98b, Ano98c, Eng00i, Fra94, Mat13b, SJO01]. Guided [SNM^{+}22]. Guidelines [ASX19]. GX [Pri90].

h [WHCK18]. H.264 [HSR18]. H.264/SVC [HSR18]. H.324 [Gol96]. H100 [Cho23]. Habana [MD20]. HALO [SKW^{+}23]. HAMLeT [AFH16]. Hand [Ano15h, Ano15i, Ano17n]. Handbook [Mat99c]. Handheld [SYW^{+}14, VW03, ZES13, Seg97]. Handicapped [AJ83, Mye82d, GRP83, HP85]. Handling [KLD^{+}94]. Hands [Hen21b, Sch91b]. Happening [ECY^{+}12, HSX18, Smo88c, Ano94c]. happens [Gre04c]. Happy [Mat99c, Mat00b]. Hard [Ano00d, Eng00l, UCS^{+}10]. Hardening [Ano87b]. Hardware [AF82, ADJK20, ABIV06, Alt12c, ACKM05, BCJ^{+}20, BC20, BTK^{+}23, BSY^{+}10, BMV^{+}08, BMM15, BSB^{+}92, BLW02, CSC^{+}22, CKG^{+}09, CGJ^{+}94, DTS20, De 94, Dem94, DPT^{+}21, DF01, Eec18c, ECK^{+}22, FN86, FSBA12, GZC^{+}20, Gro94a, GHY^{+}17, HLS^{+}21, HCW^{+}04, Hun95, HSR18, INKM05, JPOB20, Joh20a, Joh23c, Kal97, KAC^{+}95, KS18, KTC18, KCS^{+}20, Lie23, LLLL09, LP89, LSBM17, MAK19, MSS15, ML05, MRJ^{+}15, MNU^{+}15, MM23, MCC^{+}07, MCV^{+}19, NMZ13, NRS^{+}08, OT97, OHLR94, PSL^{+}23, PBFC21, PVB^{+}20, PFC^{+}02a, PFC^{+}02b, PMR^{+}22, PP92, RPE10, SG01a, SWM87, SNC^{+}07, SL03, SML04, Sch91b, SDB^{+}04, Spr02a, SKW^{+}23, SMR20, Ste83d, Ste84a, Ste85c, Ste86a, Ste87e, Sti19, SKA14b, SV21, TM94b, TM94a, TBDL01, TLM19, TATC09, VRMC20, VCD16, WPSR20, WBKR14, XBH07, YBS17, YFDV19, YZW^{+}23, Ano92b, CMR97]. hardware [CDGO97, DBDF97, FBGB96, ISH^{+}91, KKC93, KKT^{+}91, Ste83c, Ste89f, TZMVLN81, dG95]. Hardware-Accelerated [ML05]. Hardware-Agnostic [ECK^{+}22]. Hardware-Assisted [KTC18]. Hardware-Aware [MM23, TLM19]. Hardware-Based [SML04]. Hardware-Enforced [NMZ13]. Hardware-Level [INKM05]. Hardware-Software [BSY^{+}10, CGJ^{+}94, De 94, Dem94, Kal97, LLLL09, MCC^{+}07, CMR97]. Hardware/Software [Lie23, PMR^{+}22, SG01a, Ano92b, KKT^{+}91]. Harlan [Ano14o, Ano17-27, Ano18k]. Harmful [AW06, NMHS15]. Harry [Ste88d]. Harsh [Alt14b, KSE^{+}22, SKA^{+}14a, VBB14]. HArtes [BSY^{+}10]. Hartley [LNV89]. Harvesting [MLL^{+}15, MLL^{+}18]. HASE [Ibb00]. Hash [ZHZ^{+}19]. Haswell [HMB^{+}14]. Hauling [Ste95b]. HBM2 [KKL^{+}22]. HBM2-PIM [KKL^{+}22]. HC [Bre10]. HC-1 [Bre10]. HD [GDES08, KIM^{+}09]. HDL [Ano96r]. HDTV [DKM^{+}92, Kah93e, Mye91a, RT92]. Head [MMESG^{+}20, Yu96]. Head-of-Line [MMESG^{+}20]. Health [ZL16]. Healthcare [Rob99a]. Healthy [Alb07c, Gre09e]. Heap [SSMI87]. Hear [Ste07d]. heard [Eng00g]. Hearing [WMSH09]. Heart [CJFP95]. Heat [Joh20c, Ano02d]. Heavy [KLD^{+}94, Mat96c]. Heavy-duty [Mat96c]. Heavy-Ion [KLD^{+}94]. Heidelberg [MSB87]. Height [HK82]. Heightened [Ano01c]. Heights [Ano16-48, Ano16-47, Ano16-46, Ano16-45]. Helix [CJH^{+}12]. Help [Eng00j, Mat91b, Mat98d]. Helper [WCW^{+}04]. Helpful [Gre23d]. Helps [DF01]. here [Ano94c, Mat06d, Rob01c]. Hermes [Kir92]. Hertzbleed [WPH^{+}23]. HeteroGen [ONS^{+}23]. Heterogeneity [Eec15b]. Heterogeneous [Alt11d, AMFFM^{+}16, BSY^{+}10, BNV^{+}15, BSC08, DK14, ECK^{+}22, EK16, IST^{+}11, IT15, KHL^{+}16, KCXmWH17, KKS^{+}23, LSL^{+}15, LBS^{+}11, LKJ^{+}22, MRSV11, MKT^{+}13, NMU^{+}15, NGS16, ONS^{+}23, SAR10, SSLV15, SIL^{+}15, Sha23a, SLL^{+}18, SLB04a, SLB04b, SV21, TAI^{+}21, WHCK18, XYCS02, AGH^{+}91, SPT^{+}92, WWR97]. Heuristic [Den83]. Hewlett [Ano01g, Ste93a]. Hexagon [CAV^{+}14]. Hidden [ML21]. Hiding [War91f, Yea96]. Hierarchical [ACLR89, CF90, GM00, HY98, Kli81b, LHC^{+}02, PVS17, OFG88]. Hierarchies [MH08]. Hierarchy [CKD^{+}10, Sha23a, CG95]. High [Alt14d, Ano98k, ACLR89, AT93, AJC^{+}20, BFZ^{+}22, BAH^{+}05, BJW^{+}23, BDH^{+}16, Bos03c, Bos05b, BTR02, BJ14, BGH^{+}12, Car93, CRV^{+}04, Cha85b, CCYT05, CCE^{+}09, CDS^{+}15, CGMV99, CS08, CD09, CS14, CMAS11, Cum04, DQCL24, Das21, Dav98, Dia96d, Dia96c, For02, Gal97, GV97, GCL^{+}20, GRD22, Gre07e, GBW^{+}23, Gun06, HSP^{+}01, HKY^{+}95, HV04, HYS98, Hua89, HML^{+}21, JGF98, JBM95, JL87, Jos86, Kah93c, KMG^{+}03, KCXmWH17, KL05, Lin98, LLW^{+}07, LLLL09, LCP^{+}11, LCY^{+}04, MKAC18, MSB^{+}17, MM09, NG87, NFQ03, OMMB13, PKL13, PMM15, PNDG04, PKP15, PP82, PcFH^{+}02, PLB06, PSP14, Qua00, QJP^{+}08, RG03, RSW10, RC13, RBKL11, RKA^{+}20, SSLV15, SHTE08, Sak02a, Sch84, SDB^{+}04, SBJ13, SLM^{+}97, SHS85, Ste85h, SYY^{+}11, TP10, TRY^{+}09, TMJ13, TIT^{+}13, VC11, WAA^{+}20, WH09, WHCK18, WFW^{+}21]. High [WEMR04, XYT^{+}23, Yeh07, YHT^{+}15, ZZY97, Ano81, Ano96n, Ano03b, Bel93, DP97, Fis85, GP95, Iac88, Ipe19, Jag97, Kli81b, Man86b, Man86c, Pet92, TO96, Wv92, vdDD90, MHW94]. High-Associativity [ZZY97]. High-Availability [Qua00]. High-Bandwidth [Das21, TIT^{+}13, WAA^{+}20].

High-Bandwidth-Density [OMMB13]. high-definition [Pet92]. high-density [Bel93]. High-End [PNDG04, SHTE08, VC11, WH09]. High-Frequency [Lin98, SBJ13]. High-ILP [SDB^{+}04]. High-Integrity [MKAC18]. High-Level [CS14, GRD22, KCXmWH17, SSLV15, SHS85, Ano81, Kli81b, Man86c, Wv92, vdDD90]. High-Level-Language [Sch84, Man86b, Man86c]. High-Performance [ACLR89, AT93, AJC^{+}20, BFZ^{+}22, BAH^{+}05, BDH^{+}16, Bos03c, BGH^{+}12, Car93, CRV^{+}04, CCYT05, CCE^{+}09, CGMV99, CS08, CMAS11, Cum04, Dav98, For02, GV97, GBW^{+}23, Hua89, HML^{+}21, JGF98, Jos86, LLW^{+}07, LCP^{+}11, MM09, NFQ03, PKL13, PcFH^{+}02, PLB06, QJP^{+}08, RG03, RSW10, Sak02a, TMJ13, WHCK18, WFW^{+}21, WEMR04, XYT^{+}23, Yeh07, YHT^{+}15, Ano03b, Fis85, Ipe19, Jag97, TO96]. High-Quality [GCL^{+}20]. High-Radix [PKP15]. High-Reliability [Das21]. High-Speed [Alt14d, BJ14, DQCL24, Gal97, Gun06, HSP^{+}01, HYS98, JBM95, JL87, KL05, LLLL09, LCY^{+}04, PMM15, PSP14, SLM^{+}97, TP10, TRY^{+}09, Dia96c, DP97, GP95, MHW94]. High-Tech [Ano98k, Cha85b, Kah93c]. High-Temperature [MSB^{+}17]. High-Throughput [CDS^{+}15, CD09, HV04, NG87, RKA^{+}20, SYY^{+}11]. high-visibility [Ano96n]. Higher [RMC04]. highest [AAW^{+}96]. Highlights [AR16b]. Highly [DWF^{+}21, Gro94a, KSR^{+}99, RBKL11, SBG97, GDLT86]. highway [Gre96b, Mat96b]. hijacking [Ste05b]. Him [Gre15f]. Historical [Bos21, Yi21a, Yi22a, Yi22b, Yi22c, Yi22d, Yi23a, Yi23b, Yi24b, Yi24a]. History [Alt11f, Ano88a, Ano20d, Ano21j, Ano22j, Ano22k, Ano23m, Ano23n, Ano23o, Ano24e, Ano24f, Bar21, CCS21, FHMS96, Fer98a, Gas21, Hen21a, HL06, JW20, NS05, NH81, Rag21, TK21, Van21, de 84, Dan96, Gre15c, Mat05c]. History-Aware [TK21]. History-Based [HL06]. Hitachi [Ano03b]. Hits [Wil95a]. HLL [Laz89]. HLP [Ste91a]. HM [LDA87]. HM-Nucleus [LDA87]. Hold [Emm07e]. Holds [Ano99j, Jae82c, Ste06b]. Holiday [Mat01b]. Hollywood [Gre98c]. Holographic [Ano01h]. Holography [Kah92c]. Home [FH00, Joh20d, Wil95a, Ste07b]. Homebrewers [Ano87c]. homogeneous [WWR97, LDA87]. Homomorphic [LJM^{+}23, WHJ^{+}23]. Honest [Gre11c]. Honesty [Gre13e]. Hopfield [VJ89]. Hopper [Cho23]. Horizon [Sak02d, ZRA^{+}17]. Horizontally [PMM15]. Horus [KO05]. HOST [Ano20-35, Ano20-36, Ano20-37]. Hot [Alb07b, Alt12a, Alt13d, Alt14c, AR16a, AR16b, Ano00i, Ano17o, BS98, BBP09, BCN95, CM17, Eec15c, Eec16a, Eec16b, Eec17a, Eec17b, GGS22, GG16, HW91, Joh19a, Joh22e, Joh90b, JA96, KB20, Lyl04, Mas93, Ste90g, Ste90h, YT01, Alb07e, AS95, Alt11c, Alt12a, AM08, AW10, BB12, DTB01, DD05, DA23, Eec18d, FD04, GH20, GV21, HGPT12, Hoo90b, Joh22c, Joh23d, Jou92, KvdW09, KZ13, KW02, KS07, KR19a, Kur21a, Lev23, LS24, LK02, Loc03, Mat97b, NN14, NS15, RY21, RE11, SS22, SS06, SS05, WD03]. Hotmetal [Ano96g]. Hotmetal-Pro-3.0 [Ano96g]. Hottest [LTL97]. House [Ano13a, Ano13d, Ano13b, Ano13c, Ano13e, Ano13g, Ano13i, Ano14e, Ano14h, Ano14p, Ano14-29, Ano14-33, Ano14-34, Ano15j, Ano15c, Ano15f, Ano15l, Ano15s, Ano15t, Ano15u, Ano15v, Ano15b, Ano15-31, Ano15-29, Ano15-30, Ano15-32, Ano15-40, Ano15-34, Ano15-39, Ano15-35, Ano15-36, Ano16f, Ano16d, Ano16e, Ano16i, Ano16j, Ano16p, Ano16q, Ano16s, Ano16v, Ano16t, Ano16w, Ano16y, Ano16-37, Ano16-38, Ano16-41, Ano16-48, Ano17k, Ano17l, Ano17n, Ano17q, Ano17v, Ano17-27, Ano17y, Ano17-28, Ano17-39, Ano17-45, Ano17-48, Sak90a, Ano14-39, Ano17f, Ano17u, Ano17-44, Ano17-55]. HP [Han84, Kum97]. HPC [Ano18l, KL08, LNLG20, MAM^{+}06, SABS20]. HPC2002 [Ano03b]. HPS [MBG^{+}16]. HPVM [ECK^{+}22]. Hub [FRS^{+}09, MIM^{+}97]. Huge [BJW^{+}23]. Hughes [Ano87d]. Human [KHS^{+}23, WMSH09]. Humphrey [Ano23-83, Ano23-84, Ano22-40]. hundreds [SLM^{+}97]. hundredth [Pri94b]. Hung [Gre00d]. Hurdle [Kah93f]. Hush [Gre17a]. Hush-Hush [Gre17a]. HW [ZRB^{+}22]. Hybrid [ANJ^{+}04, BPT^{+}11, Bro11, HRK^{+}24, KJT^{+}11, PPO^{+}04, STR^{+}13, SWM^{+}20, XPZ^{+}19]. Hybrids [FSR^{+}05]. Hydra [HHS^{+}00]. hype [Gre97b]. Hypercard [MG88]. Hypercube [CF90, FTKS92, HMS^{+}86, LW94]. hyperlinks [Ste01f]. Hyperscale [SD21]. Hyperthreading [KM03]. HyperTransport [Ano01h]. Hyundai [Ano99k].

I-DVFS [GKS21]. I/O [Ano84, BJG^{+}19, BMS16, Ber09, DP97, HSP^{+}01, HSW98, OMMB13, WAA^{+}20]. I/Os [KMD^{+}13]. i486 [Cra90]. I860 [Atk91, KM89]. IA [Ano97w, BCC^{+}00, HMR^{+}00, KKL^{+}00, RDJ^{+}13, SCV01]. IA-32 [RDJ^{+}13]. IA-64 [Ano97w, BCC^{+}00, HMR^{+}00, KKL^{+}00, SCV01]. IAA [MLL^{+}18]. IBM [Ano96h, Ano98l, Ano01e, Ano01g, Ano02b, Ano02c, Ano03b, Ano03c, BWBJ11, Bus86, CEH^{+}12, Eng00j, HOF^{+}12, Har21, hHH99, JW20, KST04, KSSF10, OB91, RSS^{+}08, RMFG85, STKS17, SBJ13, SAC^{+}99, SGC94, STSM21, TSW^{+}01, Wea97a, Web08, dCMA22]. IBM-PC-based [RMFG85]. IC [Ano87d, Ano99w, Cla03, Koe86, STS^{+}92]. iCFP [HNR10]. iCore [RHH^{+}03]. ICs [DKM^{+}92, Mye93b, Soo93]. IDCT [RT92]. Idea [Hau88a, SRJ^{+}91, Ste88e]. Ideal [KPKJ08]. ideas [Ano17p, Bos06f]. Identification [Sak01f, SBE01]. ideology [Gre15c]. Idiomatic [WLKN22]. IEC [KZ01]. IEC/IEEE [KZ01]. IEE [Ano19d, Ano19e]. IEEE [Ano16c, Ano17g, Ano23a, Ano23b, Bel13, All86b, Ano96r, Ano98c, Ano99e, Ano99w, Ano00e, Ano01b, Ano02a, Ano02e, Ano03a, Ano05, Ano06, Ano07, Ano14n, Ano14o, Ano14p, Ano14r, Ano14q, Ano14s, Ano15t, Ano15u, Ano16q, Ano16x, Ano16s, Ano16v, Ano16t, Ano16w, Ano16u, Ano16r, Ano16y, Ano17v, Ano17z, Ano17w, Ano17-27, Ano17y, Ano17x, Ano17-28, Ano17-29, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18m, Ano18n, Ano18t, Ano19z, Ano19u, Ano19v, Ano19w, Ano19y, Ano19x, Ano19-27, Ano19-28, Ano19-29, Ano20-48, Ano20-38, Ano20-39, Ano20-49, Ano20-44, Ano20-40, Ano20-50, Ano20-45, Ano20-41, Ano20-51, Ano20-46, Ano20-52, Ano20-42, Ano20-47, Ano20-43, Ano20-53, Ano20-57, Ano20-54, Ano20-55, Ano20-56, Ano21-41, Ano21-53, Ano21-58, Ano21-36, Ano21-46, Ano21-52]. IEEE [Ano21-47, Ano21-42, Ano21-54, Ano21-37, Ano21-59, Ano21-48, Ano21-43, Ano21-38, Ano21-55, Ano21-49, Ano21-44, Ano21-60, Ano21-56, Ano21-39, Ano21-50, Ano21-45, Ano21-51, Ano21-40, Ano21-57, Ano22-47, Ano22-41, Ano22-58, Ano22-48, Ano22-42, Ano22-59, Ano22-49, Ano22-43, Ano22-60, Ano22-55, Ano22-50, Ano22-44, Ano22-61, Ano22-51, Ano22-56, Ano22-45, Ano22-53, Ano22-62, Ano22-52, Ano22-57, Ano22-54, Ano22-46, Ano22-63, Ano22-64, Ano23-27, Ano23y, Ano23z, Ano23-47, Ano23-53, Ano23-56, Ano23-48, Ano23-54, Ano23-57, Ano23-49, Ano23-43, Ano23-35, Ano23-39, Ano23-50, Ano23-44, Ano23-36, Ano23-40, Ano23-45, Ano23-38, Ano23-51, Ano23-37, Ano23-41, Ano23-46, Ano23-52, Ano23-58, Ano23-55, Ano23-42, Ano23-59, Ano23-60, Ano23-61, Ano23-62, Ano23-63, Ano24q, Ano24s, Ano24t, Ano24o, Ano24r, Ano24n, Ano24u, Ano24v, Ano24p]. IEEE [Bal84a, Bel12, BT84, Buc84, Dia94b, Dia95d, Dia96d, ES84, Eng00j, Fis85, Gro83, Hec83a, JC84, Kir01, KZ01, NS81, OL85, Pit91, RSW10, Rob97c, Rob99c, SRL91, Smo87c, Smo88b, SK88, SB00, Ste91e, Ste01e, Ste07a, Ste08c, Ste09b, Ste15b, Tau84, Tau87, War91c, Alt13a, Ano19o, Ano19p, Ano19f, Ano19g, Ano19h, Ano19i, Ano20u, Ano20-34, Ano20-32, Ano20-33, Ano20d, Ano20g, Ano20e, Ano20f, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21t, Ano21u, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n]. IEEE [Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano23u, Ano23w, Ano23x, Ano23-34, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano24m, Ano24e, Ano24f, Ano24g, Ano24h]. IEEE-1394 [SB00]. IEEE-488 [NS81]. IEEE-USA [Ste09b]. IETF [Eng00j]. If [Ano94c, MCR17, Ste08d, Ste08e]. iFlow [OG01]. II [Ang90, AQT^{+}92, Ano98-33, HW91, Jae82b, Kir85a, Man86c, Ste83d, Ste89d, Ste08e, Yi22a, Yi22e, ZMVH^{+}83a]. III [Ano99w, Ano99-28, HL99, Jae82c, Jou92, Nak00, RPK00, Ste89e, Yi22b, ZMVH^{+}83b, NCT^{+}98]. Illegal [Ste84a, Ste02c]. Illinois [CFK^{+}10]. ILLIXR [HDG^{+}22]. illustrates [Gre96a]. ILP [SNL^{+}03, SDB^{+}04, SZZ01]. iMac [Ano98m]. Image [Ano97h, CG95, Dur96, KIR19, KII09, MSP^{+}19, BCF^{+}92]. Image-Sensor-Based [MSP^{+}19]. Images [Kaw98, CG95]. Imagine [KDK^{+}01]. Imaging [Alt98, NHY^{+}22, OW01, SCYY11, WT98]. Imec [Ano98f]. Imitation [Gre04c]. Immersion [MMC^{+}22]. Immersion-Cooled [MMC^{+}22]. Impact [Bos06c, BSC08, Eec15d, KGDW^{+}13, MMESGQ22, Mar96, MCM^{+}16, UTB^{+}06, Won03, Bos06f, BTHS92, Sak99a]. Impacts [Mas21]. Impaired [LMC^{+}83]. impairment [HC83a]. impatience [Gre00f]. Imperative [LPC12]. implantable [CJFP95]. ImplantBench [JC08b]. Implement [LDL17]. Implementable [GSP02]. Implementation [ABG^{+}20, AT93, CPZ89, EGL^{+}90b, EAA85, GE86, KKY88, LNV89, LHN95, PS15, SC24, SL97, AB83, BCF^{+}92, BG81, BSB^{+}92, CM86, DKM^{+}92, FL84, KE89, NN81a, RMFG85, SMHB91, SMCT87, VS87, VJ89]. Implementations [IKK96, MC95, OFW99, PJB^{+}14, Jag97, SL97]. Implemented [SZH82, SZP81]. Implementing [ACRV96, BAC^{+}90, DMP91, GU98, GM99, KSM99, KPV^{+}99, LBS^{+}11, LM16, MMG^{+}99, RPK00, WE93, ZUNN18]. Implications [Alt13e, CEP^{+}17, DK18, GZC^{+}20, HLZ^{+}16, HKC10, MRSV11, PCDL10, Ste87c, WS13]. Important [MB99]. Imports [Noy85]. Impressive [Mat90a]. Improve [KBH^{+}08, AO97, Ano01c, CFM^{+}97, GK97, TTF96]. Improved [CGS10, LLSS05, Mac93, Tau87, Han81]. Improvement [Kah90b]. improves [Ano01h]. Improving [Ano91a, HRC^{+}23, PW96, Tab91, WK13, ZP93]. IMS [HMSS87]. In-Cache [EWW^{+}19]. In-Hardware [ADJK20]. In-Kernel [TM17]. In-Memory [ACA^{+}20, Das22, DSG^{+}22, FHL^{+}17, GAT^{+}22, GTLY22, HABHW^{+}18, KIR19, KKL^{+}22, MSY^{+}22, OSS^{+}24, PJB^{+}14, YKH^{+}19]. In-Network [LPKP22, OG24]. In-NIC [TM17]. In-Order [HNR10]. In-Package [WAA^{+}20]. in-situ [PHC95]. Inappropriate [Ste89a, Ste89c, Ste89d, Ste89e, Ste90e]. Inaugural [Bel12]. Incentives [Gre23d, ZL15]. Incidental [MLL^{+}18]. Incoherent [HBCS04]. Incomplete [Alt13d]. Inconspicuous [Cho21]. incorporate [IKK96]. increase [JKN96]. Increased [Eng00h, Ano01f]. increases [Ano01h]. Increasing [DSG^{+}22, ERM08, MTS^{+}12, Mye93b]. Increasingly [Eec15c, MB99, ESW97]. Independent [Dun81, HE07, Ste84e, Chr96, CCG^{+}84]. Index [Ano96a, Ano97a, Ano98a, Ano99e, Ano00a, Ano01b, Ano02a, Ano03a, Ano04a, Ano05, Ano06, Ano07, Ano08, Ano09b]. Index-Complete [Ano97a]. India [Kah93f]. Indirect [NAJE22]. Individual [Har12, TUI^{+}01]. Individual-Based [Har12]. indoor [SLM^{+}97]. Inductive [MKT^{+}13]. Industrial [Gre98e, Kir88b, KWGG95, Ste93f, Yi23c, Wil84]. Industrial-Property [Ste93f]. Industries [Gre02a, Gre02f, Kir90c]. Industry [Ano98h, Ano98t, ADC00, Bel96, Eec17b, Eng00m, JN21b, LCS92, MRC^{+}20, SV03, Ste92a, Ano99w, Gre98c, Kah93a, McL87, Mon87, Sla96]. industry-oriented [Mon87]. industry-standard [Ano99w]. Inference [AHKY19, DPT^{+}21, DPBW19, EKM^{+}95, MD20, MSY^{+}22, MY95, NSN^{+}93, OGLG^{+}22, RCK^{+}21, RAG19, SB23, WFW^{+}21, ACRV96, dG95, SDF^{+}23]. InfiniBand [Ano00i, CGLES^{+}23, Edd02, LMVP05, MMESG^{+}20, WPM03]. InfiniBridge [Edd02]. Inflation [Gre22e]. Influential [Bro17, Gon18, KT14, Mar14, Tor12]. informal [Rob01d]. Informatics [Kir89c]. Information [Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano19z, Ano21-52, Ano23-53, Ano23-54, Ano23-55, Ano24t, Ano24u, Dav02, FO89, Hac01, IWM89, Mil87, Pal93, Pen01, STM02, Ste94c, TLW^{+}10, AHO^{+}90, Ano16-35, Boa96, Gre93, Mat96b, Mat05d, McL87, Gre99b]. Information-Flow [TLW^{+}10]. information-gathering [Boa96]. Information-Processing [Mil87]. Informed [Sav99a, SAA^{+}99]. Infrastructure [Gre01b, PSB^{+}20, RTM^{+}10, Gre93, Gre19e]. Infringement [Ste85e, Ano91b, Ste96f, Ste00d, Ste04c, Ste04e, Ste05a]. infringing [Ste96f]. Infusion [BdS98]. initial [Han96, Pap96]. injuries [Gre96d]. Ink [TM81]. Innovation [CGG^{+}21, Dia93e, Emm07b, Gre07c, WD03, Yi23c]. Innovations [Bre10, Emm05c, Emm05d, Emm05a, Emm06e, Emm06b, Emm06a, Emm06f, Emm06c, Emm06d, Emm07a, Emm07b, Emm07c, Emm07d, Emm07e, Emm08a, Ing99, Sha23c]. Innovative [Gre02a, Gre96a]. Innovativeness [Gre09e]. Input [GSP02, HSN^{+}23, PKP15, SGP02, NA84]. Input-Output [PKP15]. Input-Queued [GSP02, SGP02]. Insensitive [BF02]. Insertion [QJP^{+}08]. Insider [Gre17b]. Insiders [Gre15b]. Insights [BCM^{+}14, GAT^{+}22, KKSV10, Wei17]. Inspection [DKSL04, KWGG95, VCK^{+}13]. Inspection-Resistant [VCK^{+}13]. inspiration [GGJ^{+}96]. Inspiring [Cho21]. Instant [Mat92b]. Instantaneous [GKS21]. Instruction [Bre10, CKG^{+}09, Cre82, CSC^{+}05, DS94, EV97, Fai82a, Fai82b, HCP^{+}16, MSWP03, NMU^{+}15, NT89, RCA07, Sch84, Sim97, Smi82, Ste87c, WRA^{+}14, ERPR95, FMT91, Lee96, MC87, MM87, TONH96, WHKM93b]. Instruction-Grain [CKG^{+}09]. Instruction-Level [EV97, RCA07]. Instruction-Set [NMU^{+}15]. instructional [RH91]. Instructions [LSY01, PPA^{+}14, Cra90, TO96]. instrument [SSL82, Gas21]. Instrumentation [DTS20, Jae82c]. Instruments [FLRB86, Chr96]. Integer [BFZ^{+}22, Mae87]. Integrals [KW83]. Integrated [AAC^{+}23, ABG^{+}20, BCU^{+}99, Bos05c, DMG00, Edd02, FMN^{+}13, Gre21c, Gre21e, MBH95, PCDL10, WLF^{+}08, YKH^{+}19, GRP83, KKT^{+}91]. Integrating [Ano97h, CDS07, JMZ^{+}11, Mur03, NST97a, NST97b, SLB04a, SLB04b]. Integration [AO97, ASX19, Alt14e, ANM^{+}12, BWMS19, Bos03a, CGO00, DDG^{+}19, GCE^{+}21, JWS^{+}19, Lau21, Mei03, MAS^{+}07, PLK^{+}16, SB07, TES^{+}18, Trö98, VRMC20, ZSS^{+}19, KHW85]. Integrity [KTY24, MKAC18]. Intel [AAC^{+}23, Ano97i, Ano97-32, Ano98-33, Ano99l, Ano99m, Ano99p, Ano99w, Ano99-28, Ano01c, Ano02c, Ano03b, Ano03c, Ano03d, Ano03e, AFK^{+}19, BCC^{+}00, BDH^{+}16, BCC^{+}02, BvdGM^{+}15, Col21, DKyL^{+}17, EAA85, Eng00i, HMB^{+}14, HSNJ21, HF81, KM89, LPKP22, NH81, PW96, PC93, PK88, RCC07, RMM^{+}04, RNA^{+}12, RYR^{+}22, RMBK81, Rya88, Sla90a, SGC^{+}16, Ste87c, Ste93a, Ste00b, Yu96, ZES13]. Intel-Intergraph [Ste00b]. Intelect [Ano14t]. intellectual [Ano98z, Dav93, Rob00d, Ste94f]. Intelligence [AK24, Cai89, FC22, FHL^{+}17, HRK^{+}24, Joh22a, KNV^{+}20, MK22, NHY^{+}22, SMM^{+}22, Ano23a, Ano23b]. Intelligent [BG02, Eec18a, GM00, Joh19b, KMD^{+}13, Pal93, PAC^{+}97, Sak90a, CR95b, GRS86]. Intelligent-Memory [BG02]. Intensive [CGS10, GGB^{+}15, SLC^{+}14, SAC^{+}21, FBGB96]. Inter [AFK^{+}21]. Inter-Chiplet [AFK^{+}21]. interact [Ste90e]. Interaction [Bel93, CLM08, FBHN04, Mat00c, War90g]. Interactions [Bha20, Kal97, Mas21]. Interactive [CP86, MW19, vW85, MM96]. Interchiplet [FKV20]. intercommunication [Mar85]. Interconnect [ANS96, BF02, BMG^{+}21, BPUH06, Cha02, Das21, FD17, Gal97, HVS^{+}07, JGF98, KND02, KL05, LNLG20, Lin04, MB99, Mei03, Sha23c, TIT^{+}13, TSA^{+}22, XLW^{+}12, XWZ09, AIH^{+}12]. Interconnect-Aware [TSA^{+}22]. Interconnected [KL08, CK95]. Interconnection [CMB22, CEH^{+}12, ED18, GQF^{+}06, GKS^{+}07, Her93, Mac93, Mis93, ODH^{+}07, RGK19, SB07, VL00, VPRS14, WGH^{+}07]. Interconnections [Mye84a, TRY^{+}09, War91b]. Interconnects [Alt13e, Alt14d, Ano00i, Ano17o, AJC^{+}20, BBP09, BCN95, Eec16b, Eec17a, GAT^{+}22, GGS22, GG16, GH20, GV21, Gun06, HAC^{+}13, HGPT12, Joh22e, Joh23a, KB13, KK23, KSR^{+}99, KNB14, KM05, KP07, Lee24a, Lev23, LS24, LTL97, LCY^{+}04, Loc03, Lyl04, MBJ08, PLB06, PSP14, SS05, TMJ13, Alt12a, LK02]. Interest [Ano85, Ano86b]. Interesting [Yi21b]. interests [Ano97t, Wil97]. Interface [Ano96m, Ano96s, Ano02e, CN13, CGO00, DRM^{+}98, Eck82, FKV20, Gil82, HKS16, Jos86, LSBM17, MCC^{+}07, MBH95, MKT^{+}13, PH91, War90e, War92b, Dan89, Dia94b, Iac88, JC84, Mat98b, Gus92]. Interfaces [BDF^{+}95, CLMY96, DJUH16, KSV^{+}21, KOI95, Mar21, SF18, SKW^{+}23, Ste89a, WBHv98, Lan96, Ste89c, Ste89d, Ste89e, Ste90e]. Interfacing [Ful91]. Interference [MMESGQ22]. Intergraph [Ano98v, Ste00b]. Interjob [MMESGQ22]. Interleaving [LTQZ07]. Intermediate [MLM^{+}20, NM22]. Intermediate-Scale [MLM^{+}20]. Intermittently [CHSL17, XPZ^{+}19]. International [Bro17, Gon18, KT14, Mar14, Rob98e, Rob01b, Ste93b, Ste95b, Tor12, Wal97, SRU^{+}23]. International-Trade [Ste93b]. Internationalization [Pir97]. Internet [GAGV22, MK22, Sav99a, AKJF22, AAC^{+}16, Ano95c, Ano99j, Ano99n, Ano99p, cCCP00, EK16, FJB^{+}22, Fra94, Gre98b, Gre00e, Gre01e, Gre02f, Gre03d, Gre03e, Gre07a, Gre08b, Gre11e, Gre15d, Gre15e, KHL^{+}16, Loc03, Mat95d, Mon97, Pfa94, RK16, RNN^{+}16, SAA^{+}99]. Interpolation [LWB09]. Interposer [KJL16]. Interprocess [US23]. interprocessor [JKP89, RT86, Zha91b]. Interrupt [SG01a]. interruptions [WE93]. Interrupts [Kir85b, MV96]. Intertwined [Mye91a]. Interview [Gre23d]. Intra [HSR18]. Intradisk [GSS09]. Intravenous [BdS98]. introduces [Ano01g]. Introducing [AH96, Cra00, Dia95c, FAWR^{+}11, Hac01, HMR^{+}00, KM89, MB15, Nak99, SSH88, SM00]. Introduction [AS91b, AKP96, AS05, ABZ08, Alb04, AS95, AM08, ANS96, AW10, AGJL98, ALGJ01, AJ83, BR10, BS98, BCP04, Ber86, BBP09, BS84, BCN95, BCA99, BAM03, Cas95, CLM08, Cle00a, Cra00, DTB01, DG89, Dem94, Dia93f, DH90, Emm08b, Fag96, FL13, FD04, GS99, GR95a, Gro92b, Gro94b, Gro02, HW91, Hoe93, Hoe92, HL86, HF84, Hun87, IA09, Jag97, Jou92, JW99, Kni85, Koo02, KW02, KS07, KP07, LB00, Lav02, LS96, LTL97, LK02, Loc03, Lyl04, Mas93, MB99, Mis93, Mon87, MRLB03, Mud10, Nak99, Nic84, OVT90, PNDG04, Pen01, PFC^{+}02a, PLB06, PP92, RDC98, Rob98d, RG07, Sak89, Sak90b, Sak91, Sak95, Sak97, Sak99f, Sak00f, Sak01f, Sak02g, SVL03, SP92, SS06, SY06, SS05, Tor06, Trö98, UB05]. Introduction [Urq97, VL00, Vei04, VN96, WD03, WG97, WT98, YT01, BG16, FG14, IA13, IT15, JA96, Kan95, PSP14, Red13, TS13, VBB14]. Introspection [MAS^{+}07]. Intrusion [TS06]. Invariants [LTQZ07]. invented [Ste01f]. Inventing [Emm07c]. Inventions [Emm05c]. Inventors [Gre04f]. inverted [CK95]. inverted-graph [CK95]. inverter [GA86]. Investigate [Ste08a]. Investigated [Ano98j]. Investigators [Mat07a]. investments [Ste94d]. Invisible [Sak02g, YYH98, Mat96e]. Invited [Emm07e]. Inviting [Ste98e]. IOCost [HSN^{+}23]. Ion [KLD^{+}94]. IoT [CEP^{+}17, GZC^{+}17, IO16, MLL^{+}18, Mar21, OG24, XPZ^{+}19, YBS17]. IOV [ZCW^{+}14]. IP [ANC05, Ano99w, Ano00g, CM04, Emm07e, Emm08a, GSC97, MFM02, SL03, SML04, Ste99a, Ste99b, Ste00b, Ste00c, Ste00a]. IP-Development [Emm07e, Emm08a]. IP-related [Ste00b, Ste00c, Ste00a]. IPC [AW06]. IrDA [Eng00j]. Irony [Gre14e]. Irregularity [DWLN20]. irresponsible [Wil95b]. ISA [AMFFM^{+}16, Kah92a, MMB^{+}08, YHHF20]. ISCA [HCPS03]. ISDN [Ano87e, Kah92b]. Isn’t [Hau88c, Ste15b, Ste97b]. ISOBlue [BCN^{+}22]. Isolating [OESGG^{+}21]. Isolation [US23]. ISSCC99 [Ano99w]. Issue [ACG03, Ano15-35, Ano15-36, Ano22b, AW22, BL23, Bor85a, Cas15, Das22, DA23, EW23, EW24, FC22, GGS22, GT22, Hoo90a, IA22, Joh22e, Kar21, KB13, KK23, Lee21, Lev23, LS24, MK22, RM23, Sak89, SS22, Sim97, Ven23, Wu23, Ano95a, TO96, Sak91]. Issued [Yi22f, Yi22e]. Issues [Alt13f, Bos03c, Bos04f, CD97b, Eec16c, FHR99, FH05, Jac03, Mat89a, Ste93b, Ste08a, Wes89, CT95, Gon97, Mat96b, Sla96, Ste89d]. Italy [Mat21c]. Itanium [Ano99m, AK00, Cra00, Eng00i, MS03, MB05, Qua00, RMC04, SCV01, SA00]. ITC [Ste95b]. Iterative [MMCH18]. ITProfessional [Ano20t, Ano21s]. ITRON [Mon87, TS95, TS91]. Itron-MP [TS91]. Itself [Ano98t]. ITT [MAT85]. IV [Jae83, Ste89a, Yi22c]. Ivy [PKB^{+}15]. Iwarp [PSW91]. iWatcher [ZQL^{+}04]. IX [Mat97b, Yi24a].

J [KHS^{+}23, SDF^{+}23]. J/Frame [KHS^{+}23]. J/Inference [SDF^{+}23]. Jackendoff [Mat13b]. Jaded [Gre98c]. Japan [Ano97-27, Kah90a, Kah92d, Kah93a, Kah93e, Kah93g, Kah93h, Sak89, Sak95]. Japanese [Mat90b, Sak90b, TM81]. Japanese-Language [Mat90b]. Java [Ano97p, Ano97q, Ano00m, CO03, CFM^{+}97, Eng00l, Fla99, Gon97, Hac01, Mat96f, Mon97, OT97, Pir97, Rit97, Rob98a, Sak01a, Urq97, WWR97]. Java-Centric [WWR97]. Java-enabled [Sak01a]. JavaBeans [Wea97a]. Javaone [San97a]. Javaone-97 [San97a]. Javastation [Ano96i]. JAZiO [HSP^{+}01]. Jeffries [Jef84]. Jersey [Ste06b]. Jet [TM81]. Jini [Edw99, Mat99d]. JN [Mon97]. Job [Alt13d, Ano14g, Ano14h, Ano15k, Ano15l, Ano16i, Ano16j, Ano16h, Ano16g, Ano17q, Ano17-30, Ano19-32]. Jobs [Ano13g, Ano14u, Ano18u, Ano20-52, Ano21-53, Ano21-54, Ano21-55, Ano21-56, Ano21-57, Ano22-58, Ano22-59, Ano22-60, Ano22-61, Ano22-62, Ano22-63, Ano23-56, Ano23-57, Gre11f]. John [Ano99q]. Join [Rob00b, SKL^{+}92]. Joining [Hau88c]. Joint [Ano98p, CHAF22, Ano03b, SM85]. Joseph [Bel13]. Josephson [HYM^{+}90]. Josephson-Technology [HYM^{+}90]. Josh [Bel13]. Journal [Ano97e, Ano98-37, Ano20-34, Ano20-32, Ano20-33, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano23-34, Ano23p]. Journey [ABC^{+}20, Gre11d, May21]. Joy [Ano03d]. Jrpm [CO03]. JTRON [Hac01]. Juki [Han85]. Jumping [Gre03b]. junk [Ste97a]. Just [CFM^{+}97, FBHN04]. Justice [Ste15b].

K5 [Ano96b, Chr96]. Kabini [BCF^{+}14]. Kanji [TM81]. Kao [Ano99q]. Kaya [Eec23]. Kbit [HM93]. Kbps [Ano97c]. Kbyte [ASD^{+}05]. Keep [Ano15v, Ano19-30, Ano19-31, Ano20-58]. Keeping [War90c]. Ken [Ano17-45]. Kennedy [Ano17-45]. Kernel [MNU^{+}15, OWK87, TS91, TM17, LDA87]. Kernel-to-User-Mode [MNU^{+}15]. Kerr [SSB95]. Kerr-type [SSB95]. Key [AKP96, ACDG99, Ano97m, ESG^{+}05, Fan96, Joh22b, LLL^{+}16, LP21, Ano97t, Ano03e, Bos06e, DVQ96, Wil97]. Key-Value [LLL^{+}16]. keyboard [NKPC83]. Keyword [RT23]. killer [CJFP95]. Kilo [CSC^{+}05, FSBA12]. Kilo-Instruction [CSC^{+}05]. KiloCore [BSP^{+}17]. Kinds [Ste08d, Ste08e]. Kinect [SO14]. Kirk [War91e]. Klessydra [CSM^{+}21]. Klessydra-T [CSM^{+}21]. KMDS [KKT^{+}91]. Knights [SGC^{+}16]. KnightShift [WA13]. knockoff [Ste96a]. know [Gre00a]. Knowing [Moo03]. knowledge [Ano17-46]. Kozyrakis [Ste16]. Kremlin [GJLT12]. Kunpeng [XCZ^{+}21].

L1 [LCWB08]. L3 [RMC04]. Lab [Sch91b]. Laboratory [LMC^{+}83, HS85, SSL82]. Labs [MD20]. Lake [RYR^{+}22, AFK^{+}19]. LAN [Ano01h, DM86, STK88, SLM^{+}97]. Lances [Buc87]. Landing [SGC^{+}16]. Landscape [Eec15c, Tay13]. Lane [Joh21b]. Lanes [Gre14d]. Language [Bal84b, Bal84c, CS81, DQCL24, KGS^{+}19, Mat90b, Mye83b, PP82, Sch84, SNM^{+}22, SHS85, Ano99w, AH96, Man86b, Man86c, SMCT87]. Language-Guided [SNM^{+}22]. Languages [LBS^{+}11, Mat99c, Ano81, HLHR90]. LANs [Ano96v]. Laplace [TKS^{+}22]. laptops [Ano99p]. Large [ADJK20, Alt11f, BJG^{+}19, CMB22, DQCL24, Dav98, Far85, FM91, HAC^{+}13, IST^{+}11, JL11, JGC^{+}11, KDSA09, KO05, KKSV10, LHMH91, LH12, Mae87, MBJ08, MSWP03, PVS^{+}11, PCC^{+}15, RNN^{+}16, Sak02d, XLX^{+}23, ZIM^{+}07, ZRA^{+}20, AKK^{+}93, Mat96f, Yea96]. Large-Scale [Alt11f, BJG^{+}19, Far85, HAC^{+}13, IST^{+}11, JL11, JGC^{+}11, KDSA09, KO05, KKSV10, PCC^{+}15, XLX^{+}23, ZIM^{+}07, AKK^{+}93]. Larger [RMC04, MIM^{+}97]. Larrabee [SCS^{+}09]. Laser [Ano02d, CAH86, Ano92a]. Lasers [Ano87a]. Last [Gol21, Gre16a, IDI^{+}21, Ste09d, Ste85g, SKJ^{+}11]. Last-Level-Cache [SKJ^{+}11]. LastLayer [VRMC20]. late [Bos05d, Gre05b]. Latency [BRmWH06, BTK^{+}23, BJW^{+}23, CSV02, DMMD11, Das21, DGM^{+}11, GAR^{+}06, LWB09, LM16, MKP06, RKA^{+}20, SB07, SZZ01, Sha22, SGK^{+}04, SRA^{+}04, BD94, VBB95, Yea96, Zha91b]. Latency-Critical [RKA^{+}20]. latency-hiding [Yea96]. Latency-Tolerant [GAR^{+}06]. Lateral [NNS^{+}93]. Latest [MAT^{+}18]. Lattice [SFP^{+}23]. Lattice-Based [SFP^{+}23]. Lattices [Ano97m]. launches [Ano03b, Ano03d]. Launching [Del91b]. laurels [Ano96k]. Law [FS05, Gre15f, Mat83, Ste83b, Ste83c, Ste83d, Ste83a, Ste84a, Ste84b, Ste84c, Ste84d, Ste85b, Ste85c, Ste85d, Ste85e, Ste86a, Ste86f, Ste86b, Ste86c, Ste86d, Ste86e, Ste87a, Ste87c, Ste87b, Ste87d, Ste87e, Ste88e, Ste88a, Ste88b, Ste88c, Ste88d, Ste89c, Ste89d, Ste89e, Ste89a, Ste89b, Ste89f, Ste90e, Ste90a, Ste90b, Ste90c, Ste90d, Ste90f, Ste91b, Ste91a, Ste91c, Ste91h, Ste91d, Ste91e, Ste91f, Ste91g, Ste92a, Ste92b, Ste92c, Ste92d, Ste92e, Ste92f, Ste93c, Ste93d, Ste93e, Ste93a, Ste93f, Ste93b, Ste93g, Ste94b, Ste94d, Ste94c, Ste94a, Ste94e, Ste94f, Ste95a, Ste95b, Ste95c, Ste95d, Ste95e, Ste96a, Ste96b, Ste96d, Ste96e, Ste96c, Ste96f, Ste97a, Ste97b, Ste97c, Ste97d, Ste97f, Ste97e, Ste98c, Ste98e, Ste98a, Ste98f, Ste98b, Ste98d, Ste99a, Ste99b, Ste99e, Ste99c, Ste99d, Ste00b, Ste00c, Ste00a, Ste00d, Ste01a]. Law [Ste01b, Ste01d, Ste01c, Ste01e, Ste01f, Ste02a, Ste02b, Ste02c, Ste02d, Ste03a, Ste03b, Ste04a, Ste04b, Ste04c, Ste04d, Ste04e, Ste05d, Ste05b, Ste05c, Ste05a, Ste06a, Ste06b, Ste07a, Ste07b, Ste07c, Ste07d, Ste07e, Ste08a, Ste08c, Ste08b, Ste08d, Ste08e, Ste09a, Ste09c, Ste09b, Ste09d, Ste12, Yi22f, Eec17c, Eec17e, Gre12f, Gre17c]. Laws [Ano99j, Ano99n, Ano99p, Dav93]. Layer [Gre14a, KGDW^{+}13, WLF^{+}08]. Layered [BLW02]. Layerwise [DSG^{+}22]. Layout [ZKP^{+}23, Ste91h]. lazy [Ano97o]. Lead [Ano01h, Pri94b]. Lead-free [Ano01h]. Leaders [Alt14e, JN21b]. Leadership [Ano17-29, Mat03b, Zsc84]. Leading [Ano16-48, Ano16-47, Ano16-46, Yi21c, Ano16-45]. Leakage [AMR^{+}06]. Leaking [TSFS21]. Leap [Mil89]. learn [Ano94c]. Learned [Pri95]. Learning [ATS^{+}22, ACA^{+}20, AJC^{+}20, BJG^{+}19, BTK^{+}23, BCH^{+}23, CDBY23, DSL^{+}18, DPY18, EPM^{+}20, EP19, GTLY22, Gre19c, IPL^{+}23, IO16, Joh19d, KSA^{+}19, KRD^{+}20, KKH^{+}24, KSE^{+}22, KGT22, Kur20b, KR19b, Lie23, LH20, LYP^{+}18, MSP^{+}19, Mİ09, Mat02b, MRC^{+}20, MCV^{+}19, MAJ^{+}18, NM99, NHMM23, NMF^{+}23, PEZ^{+}19, PMS23, PFC^{+}02a, PFC^{+}02b, PMR^{+}22, RT23, SVA^{+}22, SLL^{+}18, SNM^{+}22, SKM23, TAI^{+}21, TSA^{+}22, WZL20, WAA^{+}21, YKH^{+}19, ZFW^{+}23, ZRA^{+}17, Ano03e, CT95, Ipe19, OZT^{+}22, PHC95]. Learning-Based [NMF^{+}23, WZL20]. Left [AS22, Ste93e]. Legacies [Gre23c]. Legacy [ATS^{+}22, Mar21, Web21]. Legal [HA96, Mac98, Ste87a, Ste89a, Ste91a, Ste03a, Ste89c, Ste89d, Ste89e, Ste90e]. Legend [Ano96b]. Legislating [Gre06d]. Legislation [Eng00d]. Legislative [Ste86c]. Lego [Dia99]. Length [PPP01, Yi22a, CCG^{+}84]. Less [Ano97g, Ano15s, KST12, Ano02d]. Lesson [Gre07e]. Lessons [Bos04b, HAWC^{+}11, Pow94, Pri95, Mat02b]. Let [Gre97c]. Letter [Far87, Kir01]. Letters [Ano19g, Del92, Hoo90d, KS00, Mar98, Par00, Wha97, Wil03]. Level [BMR^{+}06, Bos03a, CJH^{+}12, CDS07, CS14, Dun81, EDL^{+}04, EV97, EE08, FZW^{+}12, GRD22, HNR10, INKM05, Jac03, KCXmWH17, MT05, MBG^{+}16, NPC06, PP82, PLBC09, RCA07, RSC^{+}06, SSLV15, Sch84, SHS85, SKJ^{+}11, SsSMB24, TNT06, TLW^{+}10, WBHv98, Ano81, Bos04c, Kli81b, KSI^{+}96, Kra96, Man86b, Man86c, Rit97, Seg97, Wv92, vdDD90].

Level-Independent [Dun81]. levels [FMT91, OFG88]. Leveraging [BMR^{+}06, MMESG^{+}20, YFDV19]. levy [Ste07b]. LG [Ano99k]. Liability [Ste87b, Ste98d, Ste04e]. liable [Ste96f]. Libraries [Ste85d, SKC^{+}23, Ano03e]. Library [Ano96e, DWF^{+}21, MBS92, Ano14-38, Ano14-39, Ano15-41, Ano17-55]. Licensable [BY07]. License [Ste93a, Ste97f]. licenses [Jag97]. Licensing [Ste99a, Ste99b, Rob00d, Ste94e]. lies [Ste05b]. Life [Dia95a, Dia98, Gre12e, Mat09a, WG92, Ano94b, Han96, Mat96d]. Life-Cycles [Dia95a]. Lifeguards [CKG^{+}09]. Lifetime [SABR05]. Light [Ano02c, Kir91b, YYH98, Ano02b, DTH^{+}95]. Light-emitting [Ano02c]. Lightweight [KHL^{+}16]. Like [Ano88d, Gre24b, Gre98c]. Limitations [HYS98]. Limited [BGK97, DVQ96]. Limiting [CDGO97]. Limits [Mye92b, NBS^{+}18, PDS^{+}13, Gre00b]. Line [Ano98g, CJFP95, DDG^{+}19, FH00, MMESG^{+}20, SIPM02, DO84]. Linear [HGS^{+}17, Ipe19]. Lines [Das17, Gre14d, GT83]. Lineup [Ano98l]. Link [ANJ^{+}04, AFK^{+}21, PPBS03, Sha23a, Sha23b, SC24, SLM^{+}97]. Links [EKB^{+}96, KKP^{+}09, OMMB13]. Linpack [RBKL11]. Linu [Gre12f]. Linux [Eng00f]. liquid [Ano03e, DTH^{+}95]. liquid-cooled [Ano03e]. Lisa [Ste89b]. list [Ano97s, LLC90, Rob97a]. listing [Ano96a]. Listings [Ano97a, Ano98a, Ano01b, Ano00a]. Lite [NY22]. Lithography [Ano88g, Ano96l, Ano01e, Ano01f, Ano01h]. Little [Gus85, Mat03e, Mye83b, Rob99e, Gre08a]. Little-Endian [Gus85]. Live [WPSR20]. LiveHD [WPSR20]. lives [Mat95b]. LIW [PSW91]. Llano [BFS12]. lnspired [VN96]. Load [AKJF22, ACKM05, GAR^{+}06, KCAR18, RTJ20, RTJ21, SMR07]. Load-Balanced [ACKM05]. Load-to-Use [RTJ20, RTJ21]. Local [BCF^{+}95, Mye82b, Mye82c, RMBK81]. Local-Area-Network [BCF^{+}95]. Local-Network [Mye82b, Mye82c]. Locality [SG00, SW14]. Localized [KM05]. log [WN94]. Logarithm [Mae87]. Logic [Ano18d, ABK^{+}17, CMR97, CDGO97, GT83, Ham00, IGH^{+}99, JL87, LDL17, LM16, MSS15, Pea95, PZB^{+}19, PFC^{+}02a, PFC^{+}02b, PDL08, Ste86a, TTF96, TCF96, TMA18, WS13, YBNS15, Ano95d, GP95, Lan87, LGJ95]. Logic-on-Memory [PZB^{+}19]. Logical [BDSC21, MG89, Ste85f, ZV85, ZVH85, Dan89]. Loh [Dwa18]. Loihi [DSL^{+}18]. Long [AML^{+}03, Gre08c, Gre08d, IBM05, KYG19, Ste85g, Gre07f]. Long-Term [IBM05]. Longtime [Ano96j]. Look [CCS21, Lie23, NHMM23, Ste86f, Ste94a, ZZ05, Gre98c, Rob99b, Ste93c]. Look-Ahead [ZZ05]. Looking [Ano17-30, Ano19-32, Bos03d, Bos06b, Eec16c, Eec17d, Gre97d, Joh21c, Mat98a, Mat07c, Sak87a]. looks [Yu96]. Lookup [CM04, YKL05]. Loop [CK11]. Loop-Directed [CK11]. loses [Ste01a]. Losses [Kar88b]. Lossless [LKGL24]. Lossy [LKGL24, OESGG^{+}21, OESGG^{+}21]. Love [Kir90b]. Low [Ano17-57, ASD^{+}05, BFZ^{+}22, BJW^{+}23, BCKY17, BS17, BCD^{+}11, BGH^{+}12, Car93, CCA^{+}19, CL05, CDY^{+}18, CR95b, CEP^{+}17, CJFP95, Das21, Dea04, DRB^{+}12, DPT^{+}21, Eec17e, EDL^{+}04, GDN^{+}17, GZC^{+}17, GALB07, HRK^{+}24, HSP^{+}01, HKY^{+}95, KSLY17, KOKA23, KTY24, LM16, LAT^{+}01, MBS08, MS87, NKDN95, NIJ^{+}03, OKH^{+}12, OMMB13, PO04, RC13, SCA^{+}12, Sha22, SBG^{+}07, SCC^{+}05, Sto90, SYY^{+}11, UBH^{+}94, VBB95, WAA^{+}20, WFW^{+}21, WGA^{+}09, YBS17, Yeh07, ZZ02, Ano02b, DVQ96, Dia95d, Eng00j, Fly97, FN94, GK97, Jag97, Kra96, Lan96, Sak99d]. low- [Eng00j]. Low-Bit [BFZ^{+}22]. Low-Cost [Car93, Dea04, GALB07, HSP^{+}01, MBS08, MS87, Sto90, UBH^{+}94, DVQ96, Dia95d, GK97, Jag97]. Low-Energy [SCA^{+}12]. Low-Latency [Das21, Sha22, VBB95]. low-level [Kra96]. Low-Overhead [KTY24]. Low-Power [ASD^{+}05, BCKY17, BCD^{+}11, BGH^{+}12, CL05, CR95b, CJFP95, DRB^{+}12, EDL^{+}04, GDN^{+}17, GZC^{+}17, HRK^{+}24, HKY^{+}95, KSLY17, KOKA23, LAT^{+}01, NKDN95, NIJ^{+}03, OKH^{+}12, OMMB13, PO04, Sha22, SBG^{+}07, SCC^{+}05, SYY^{+}11, WAA^{+}20, Yeh07, ZZ02, Fly97, FN94, Jag97, Lan96, Sak99d]. Low-Precision [DPT^{+}21]. Low-Voltage [CCA^{+}19, WGA^{+}09, Ano02b, FN94]. low-voltage/low-power [FN94]. lower [Ano02c]. LPDDR5 [KKL^{+}22]. LPDDR5-PIM [KKL^{+}22]. LSFQ [BFZ^{+}22]. LSI [Tab84, AR83, Ano02c, KKS^{+}98, Pee87, SSY97, Tab84]. Lunch [Gre18b].

M0 [TKI^{+}14]. M3 [RBGZ19]. M32R [NST97a, NST97b]. M32R/D [NST97a, NST97b]. M5 [BDH^{+}06]. M7 [AJK^{+}15]. Mac [Ano98r, Ano98-38]. MacChesney [Ano99q]. Machine [AF82, ACA^{+}20, BJG^{+}19, BTK^{+}23, BCH^{+}23, CDBY23, DPY18, EP19, Gre19c, Joh19d, KSA^{+}19, KRD^{+}20, Kur20b, LL03, LH20, LYP^{+}18, Mİ09, MRC^{+}20, NMF^{+}23, PMS23, SWL90, WAA^{+}21, ZL16, Ano03e, Boa96, FS05, HS92, Ipe19, OZT^{+}22, Ste05d, BNOv87, Mon97, OT97]. Machine-Learning [DPY18]. Machine-Learning-Assisted [CDBY23]. machine-vision [Boa96]. Machines [AS91b, BMS16, BI17, de 84, WWR97]. MacInTax [Mat95c]. Macintosh [LS98b, Mat89a, Mat89b, Mat93b, Mat97c, Wes89]. MacWorld [Mat99c, Mat88]. Made [CCS21, MBA^{+}09, Ano95d]. Madhavani [Gre12e]. MAESTRO [KCS^{+}20]. Magazine [Ano24h, Ano24g, RJ91]. Magazines [Ano13e]. magic [Hin88].

Magnetic [YW88]. Magnification [Vac87]. Magnitude [AB83]. mail [Gre01a, Ste97a]. Main [Cri97, DRB^{+}12, JWS^{+}19, LZY^{+}10, YE11, ZHZ^{+}19, KSI^{+}96, Swa19]. Mainframe [JW20, SBJ13, Web08, Web21]. Mainframes [Hen21c, Gre95d]. Mainstream [CB10, CJH^{+}12, Sti11, Dia00]. Maintain [LDF^{+}13, Zsc84, Mat96f]. Maintaining [Ber09, SIPM02]. MAJC [TCC^{+}00]. Major [Ano16s, Ano16t, Ano16r, Ano17w, Ano17y, Ano17x, Ano23t, SL97]. Make [WG92]. makes [Ano02b, Ano02d, Gre96a, Mat96d]. Making [BCH^{+}23, CJH^{+}12, Mat01c, Pir97, Rob00c, Sak02g, WFA^{+}10]. Malaysia [Kah93b]. Malicious [SWL11]. Malthus [Gre03c]. man [Fer98b]. Manage [Ano19-38, Mye84a]. Management [Ano18t, BBE^{+}11, CK98, Dia93a, FAWR^{+}11, FMN^{+}13, GQF^{+}06, KC09, LDF^{+}13, LLZ^{+}04, LLSS05, Mİ09, MMB12, Mil90, MW19, NMC^{+}08, RNA^{+}12, SBG^{+}07, TSS18, TVT19, WBHv98, WJM^{+}05, YFDV19, ZHPR17, CM86, Kai88]. Managers [KHHR85]. Managing [Ano99f, GKL^{+}14, Gre12c, Mat01d, Mat03c, Moo03, Moo04a]. Manipulating [BK14]. Manipulators [EEJ95]. Manticore [ZSB21]. mantras [Mat95c]. Manufacturers [Ste87b, Ste95b]. Manufacturing [HOHCV99, KWGG95]. Many [BYM^{+}07, BJO^{+}09, CLM08, FZW^{+}12, Hen21b, HKC10, LLT^{+}08, Mat03e, SCS^{+}09, WK13, Mat06c, Rob99f]. Many-Core [BYM^{+}07, BJO^{+}09, CLM08, FZW^{+}12, HKC10, SCS^{+}09, WK13]. Manycore [DSL^{+}18, MFN^{+}17]. ManySim [ZIM^{+}07]. Map [Ano87f]. Map1000A [BLO00]. Mapped [BDF^{+}95]. Mapping [KYG19, KMG^{+}03, MM96, SHS85, Dv87]. Mappings [KCS^{+}20]. MapReduce [PJB^{+}14]. Maps [RGR95]. march [Gre05c]. Margin [ZHPR17]. marker [Ano01c]. Market [Ano00g, Cas95, Gon99, Gre10d, Gre16c, Mye93a, Mye93c, Rob98d, Sak02d, Ano02c, Ano03d, Gre95c, Gre97f, Hal93, MKRC97, Sak99e]. Marketing [Smi96a, Ste89b]. Markets [Gre93, Gre02b, Gre22e, Gol96, Gre05f]. Marriage [Gre14c]. Mars [KDK^{+}89]. Marshaling [SMJ^{+}11]. Marvell [SSR21]. Mass [Gre10d]. MASSC [Tua99]. Massive [ASK^{+}15, Ano88h]. Massively [But07, DGM^{+}11, ROA13, Lou91]. Masthead [Ano09e, Ano09f, Ano10d, Ano10e, Ano11, Ano13h, Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano16-31, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano18v, Ano18w, Ano18x, Ano18y, Ano19-33, Ano19-34, Ano19-35, Ano19-36, Ano19-37, Ano20-59, Ano20-60, Ano20-61, Ano20-62, Ano20-63, Ano20-64, Ano21-61, Ano21-62, Ano21-63, Ano21-64, Ano21-65, Ano21-66, Ano21-67, Ano22-65, Ano22-66, Ano22-67, Ano22-68, Ano22-69, Ano22-70, Ano23-64, Ano23-65, Ano23-66, Ano23-67, Ano23-68, Ano23-69, Ano24w, Ano24x]. Material [Ano87b, Ano01h, Pri94b, Ste96f]. materials [Hal91, SSB95]. Mathematica [Mat91b]. Mathematical [And82a, ACG^{+}88, KW83, KHW85, KHF86]. matrices [RJHK89]. Matrix [CWL^{+}14, NY22, dCMA22]. Matter [Gre11a]. Matters [Ano14-27, Ano14-28, Ano14-29]. Maturing [DH90]. maturity [Gre05c]. Mauchly [Ano16c, Ano17g, Goo14, Hil19, LE18, Mud15, Wei17]. Maurer [Ano99q]. Maurice [Dwa18, KT14, Mar17, Sco14, Ste16]. MAX [Lee96]. MAX-2 [Lee96]. may [Ano01c, Pri94b, Joh20b]. Maze [JP17]. Mbits [SLM^{+}97]. MBus [PLK^{+}16]. MC68010 [MM83]. MC68020 [MMM84, MR85, Mac84, Rys84]. MC68020-Based [MR85]. MC68060 [CEM^{+}95]. MC6809 [NS81, SL84a]. MC68332 [JGB^{+}89, McD21]. MC68824 [DM86]. MC68851 [CM86]. MC68881 [HC83b]. MC68HC11 [GA86, Sib84]. McEliece [NM24]. MCM [Ano97z, Dav98]. MCU [Dan96]. Meaning [Mat13b, Mat13b]. Means [Joh23e, VC11]. Measure [Gil96a]. measurement [VS87]. Measurements [War90a, KKC93]. Measuring [Ano97j, DMWS13, MWE^{+}03, Bos06f]. Mechanics [Emm06f]. Mechanism [CHAF22, Mor84, YMC^{+}12]. Mechanisms [DSK^{+}92, KLD^{+}94, OL85]. MEDEA [Bor99a, GS99]. Medfield [ZES13]. Media [DDHS00, KDK^{+}01, LS96, TONH96, Ano95a, Ano98z, Han96, Lee96]. media-processing [TONH96]. mediaDSP [SP09]. Mediaprocessor [BLO00, THT^{+}04, Han96]. Mediaprocessors [KMG^{+}03, KMK01, Mou96]. Medical [CS08, FO89, Gre19c, SCYY11]. Medium [Pap89]. Meet [Ano92d, Bos03b, DPT^{+}21, Mat22, PGL97]. Meeting [Kir85a]. Meetings [Far88a]. Meets [Gre03c, JUP^{+}22, KCXmWH17]. Mega [OYS^{+}11]. Mega-Arrays [OYS^{+}11]. Megacells [Sto86]. Melco [Kah92c]. mellifluous [Gre05c]. Mellon [Ano22z]. Meltdown [HMR^{+}19]. Members [Eec16e]. Membership [Ano13i, Ano14-27, Ano14-28, Ano14-29, Ano16-33, Ano16-34, Ano17-43, Ano17-40, Ano17-44, Ano17-41, Ano17-42, Ano18z, Ano18-27, Ano18-28, Ano18-33, Ano17-46]. Memcapacitive [OSS^{+}24]. Memoization [THC18]. Memories [AF88, Bon21, CL05, CEAY23, Gro92a, Gro92b, Har21, JLG19, Joh22d, Kat97, MHP^{+}23, MC92, SCSR93, Ano98-29, Qur19]. Memory [AAC^{+}23, ADF^{+}10, AFH16, Aki18, Alt13a, AAK^{+}06, ACA^{+}20, Ara00, AMFFM^{+}16, BEL^{+}23, Bha17, Bha18, BDF^{+}95, BMV^{+}08, BWR23, BNV^{+}15, BG02, CL04, Che19, cCCP00, CKD^{+}10, Cri97, CSC^{+}05, DD05, Das17, Das22, DSG^{+}22, DRB^{+}12, DLCO10, DVWW05, EGL^{+}90a, Eng00e, EKMW02, FSS^{+}16, FHL^{+}17, FHL^{+}03, FSBA12, GKA^{+}16, GHS17, GAT^{+}22, GTLY22, Ger19, Gil96b, GV97, GKB^{+}23, GGB^{+}15, HRC^{+}23, HABHW^{+}18, HCU^{+}07, HKS16, HL06, JM98, JWS^{+}19, Joh21b, KZS^{+}22, KJL^{+}10, KJT^{+}11, KMK01, KPMHB11, KLM^{+}15, KHL^{+}16, KCXmWH17, KIR19, KKL^{+}22, KKS^{+}23, KGDW^{+}13, KL05, KGS^{+}19, KTY24, KR19b, KJC^{+}23, LZY^{+}10, LKJ^{+}22, LHL09, LG24, LPM15, LSBM17, MM83, MHW03, MCC^{+}07, MSY^{+}22, Mil90, MBH95, MKP06, MM09, NMZ13, NAJE22, OSS^{+}24, PSL^{+}23, PCW15, PZB^{+}19, Pre91, PJB^{+}14, PVS17, RRP^{+}08, Rob92, RLS11, SWL11, SDB^{+}04, SZZ01]. Memory [Sha22, Sha23a, SNM^{+}13, SAC^{+}21, SB23, SsSMB24, TS91, THP^{+}19, TM94b, TM94a, TSW^{+}01, TML^{+}18, US23, VCK^{+}13, VMW^{+}19, WH09, WLY^{+}21a, WBHv98, WWZ^{+}08, WHKM93b, XZ19, XBH07, YKH^{+}19, YE11, YMC^{+}12, ZSS^{+}19, ZRB^{+}22, ZHZ^{+}19, Ano95b, Ano01h, Ano02d, BD94, CM86, HMAF90, HM93, Kai88, OZT^{+}22, Sol19, Swa19, WBC^{+}95, GK97]. Memory-Integrated [MBH95]. Memory-Oriented [PSL^{+}23]. Memristive [BI17, HABHW^{+}18, Ipe19, YKG18]. Memristor [Ano18d, Chu18, JS18b, SFP^{+}23, TMA18]. Memristor-Based [JS18b]. Memristors [Aki18, Eec18e]. MEMS [Ano01c, Ano02e, TP10]. MemScale [DRB^{+}12]. MEMTI [DPBW19]. mensch [Gre99c]. mentally [HP85]. Merasa [UCS^{+}10]. Merced [Ano98w]. Merge [KJMP07]. Merges [Ano99k]. Merging [DFR90, DVQ96]. Merwin [Ano14a, Ano15b, Ano16b, Ano17-29, Ano17b, Ano18-29]. Mesh [HVS^{+}07, LHL09]. Mesh-Based [LHL09].

Meshes [LSL^{+}15]. Mesoscale [GFL^{+}17]. Message [Alb07d, Alb07e, Alb07b, Alb07a, Alb07c, Bos03b, Bos03c, Bos03d, Bos04b, Bos04c, Bos04d, Bos04e, Bos05a, Bos05b, Bos05c, Bos05e, Bos05d, Bos05f, Bos06e, Bos06d, Bos06c, Bos06a, Bos06b, Bos06f, DSK^{+}92, Dia98, Sak99b, Sak99a, Sak99e, Sak99d, Sak99c, Sak00c, Sak00a, Sak00d, Sak00e, Sak01c, Sak01a, Sak01b, Sak01d, Sak01e, Sak02c, Sak02b, Sak02d, Sak02e, Sak02a, Sak02f, SL84b, Tal93, XLW^{+}12, Sak00b]. Message-Driven [DSK^{+}92]. Message-Passing [XLW^{+}12]. Message-Routing [Tal93]. messages [VBB95]. Messaging [Gre09d]. Meta [NHMM23, Sko83]. Meta-assemblers [Sko83]. Meta-Reasoning [NHMM23]. Metaclasses [Ano98z]. Metadata [KTY24]. MetaE2RL [NHMM23]. Metaflow [PSS^{+}91]. metal [IWM89]. metal-oxide [IWM89]. metaphors [Gre19d]. Metaphysics [Emm08b]. MetaTM [RRP^{+}08]. MetaTM/TxLinux [RRP^{+}08]. Method [KOKA23, PBT06, SHTE08, Ste14a, Ste14b, KAK96]. Methodologies [DXT^{+}18]. Methodology [ED18, KL08, LHC^{+}02, SCC^{+}05, XYT^{+}23, RS90]. methods [Ste96c]. Metric [Kir91a]. Metrics [EE08]. Mflops [Gil96a]. MHADBOR [AKJF22]. MHz [Ano96k, Ano97-31, JBF94, NG87, RHH^{+}03, WHKM93a, WHKM93b]. Mica [HC02]. mice [Ste99e]. Micon [BGS89]. Micro [Ano91b, Ano94d, Ano95b, Ano95c, Ano95d, Ano96l, Ano96k, Ano96m, Ano96n, Ano97l, Ano97m, Ano97k, Ano97n, Ano97o, Ano97p, Ano97r, Ano97q, Ano97s, Ano97t, Ano98t, Ano98u, Ano98s, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99n, Ano99l, Ano99m, Ano99o, Ano99p, Ano99r, Ano99q, Ano99s, Ano99t, Ano99w, Ano99u, Ano99v, Ano99x, Ano99y, Ano00f, Ano00g, Ano01a, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02b, Ano02c, Ano02d, Ano02e, Ano03b, Ano03c, Ano03e, Ano04b, Ano04c, Ano04d, Ano04e, Col21, Dia93c, Dia93d, Dia95d, Dia95e, Dia96a, Dia96d, Dia96c, Dia99, Dia00, Emm05c, Emm05d, Emm05a, Emm06e, Emm06b, Emm06a, Emm06f, Emm06c, Emm06d, Emm07a, Emm07b, Emm07c, Emm07d, Emm07e, Emm08a]. Micro [Eng00a, Eng00c, Eng00b, Eng00e, Eng00d, Eng00f, Eng00h, Eng00i, Eng00j, Eng00k, Eng00l, Eng00m, Eng00o, Eng00n, Eng00p, Eng00g, Fer98a, Fer98b, Fla99, FS05, Gon99, Gre93, Gre95a, Gre95c, Gre95b, Gre95d, Gre96a, Gre96b, Gre96c, Gre96d, Gre96e, Gre96f, Gre97a, Gre97b, Gre97f, Gre97c, Gre97d, Gre97e, Gre98a, Gre98b, Gre98e, Gre98c, Gre98f, Gre99c, Gre99d, Gre99b, Gre99a, Gre99e, Gre99f, Gre00b, Gre00f, Gre00c, Gre00d, Gre00e, Gre00a, Gre01b, Gre01a, Gre01c, Gre01d, Gre01e, Gre01f, Gre02a, Gre02c, Gre02b, Gre02d, Gre02e, Gre02f, Gre03a, Gre03b, Gre03c, Gre03d, Gre03e, Gre04b, Gre04a, Gre04d, Gre04c, Gre04e, Gre04f, Gre05e, Gre05a, Gre05b, Gre05c, Gre05d, Gre05f, Gre06a, Gre06b]. Micro [Gre06c, Gre06d, Gre06e, Gre06f, Gre07d, Gre07a, Gre07b, Gre07e, Gre07c, Gre07f, Gre08a, Gre08c, Gre08d, Gre08b, Gre08e, Gre09b, Gre09c, Gre09a, Gre09f, Gre09e, Gre09d, Gre10d, Gre10f, Gre10e, Gre11c, Gre11d, Gre12a, Gre12b, Gre12c, Gre12d, Gre12e, Gre13b, Gre13c, Gre13d, Gre13e, Gre13f, Gre14a, Gre14c, Gre14d, Gre14e, Hur97, IJ98, LCN^{+}22, Mat95b, Mat95c, Mat95d, Mat96a, Mat96c, Mat96e, Mat96b, Mat96d, Mat96f, Mat97a, Mat97b, Mat97c, Mat97d, Mat98b, Mat98c, Mat98d, Mat99b, Mat99a, Mat99c, Mat99d, Mat99e, Mat99f, Mat00a, Mat00b, Mat00c, Mat00d, Mat00e, Mat01a, Mat01b, Mat01c, Mat01d, Mat01e, Mat01f, Mat02a, Mat02b, Mat02d, Mat02c, Mat03a, Mat03b, Mat03c, Mat03e, Mat03d, Mat03f]. Micro [Mat04a, Mat04b, Mat04c, Mat04e, Mat04d, Mat05a, Mat05b, Mat05d, Mat05c, Mat05e, Mat06d, Mat06a, Mat06c, Mat06b, Mat07a, Mat07b, Mat07c, Mat07d, Mat08b, Mat08a, Mat09a, Mat09b, Mat09d, Mat09c, Mat09e, Mat10b, Mat10c, Mat10d, Mat11a, Mat12a, Mat12b, Mat13a, Mat13b, Pit95, Pri94a, Rob97a, Rob97b, Rob97c, Rob97e, Rob97d, Rob98b, Rob98e, Rob98c, Rob99b, Rob99a, Rob99c, Rob99e, Rob99d, Rob99f, Rob00a, Rob00e, Rob00b, Rob00c, Rob00d, Rob01a, Rob01b, Rob01d, Rob01c, Smo87a, Ste83b, Ste83c, Ste83d, Ste83a, Ste84a, Ste84b, Ste84c, Ste84d, Ste85b, Ste85c, Ste85d, Ste85e, Ste86a, Ste86f, Ste86b, Ste86c, Ste86d, Ste86e, Ste87a, Ste87c, Ste87b, Ste87d, Ste87e, Ste88e, Ste88a, Ste88b, Ste88c, Ste88d, Ste89c, Ste89d, Ste89e, Ste89a, Ste89b, Ste89f, Ste90e]. Micro [Ste90a, Ste90b, Ste90c, Ste90d, Ste90f, Ste91b, Ste91a, Ste91c, Ste91h, Ste91d, Ste91e, Ste91f, Ste91g, Ste92a, Ste92b, Ste92c, Ste92d, Ste92e, Ste92f, Ste93c, Ste93d, Ste93e, Ste93a, Ste93f, Ste93b, Ste93g, Ste94b, Ste94d, Ste94c, Ste94a, Ste94e, Ste94f, Ste95a, Ste95b, Ste95c, Ste95d, Ste95e, Ste96a, Ste96b, Ste96d, Ste96e, Ste96c, Ste96f, Ste97a, Ste97b, Ste97c, Ste97d, Ste97f, Ste97e, Ste98c, Ste98e, Ste98a, Ste98f, Ste98b, Ste98d, Ste99a, Ste99b, Ste99e, Ste99c, Ste99d, Ste00b, Ste00c, Ste00a, Ste00d, Ste01a, Ste01b, Ste01d, Ste01c, Ste01e, Ste01f, Ste02a, Ste02b, Ste02c, Ste02d, Ste03a, Ste03b, Ste04a, Ste04b, Ste04c, Ste04d, Ste04e, Ste05d, Ste05b, Ste05c, Ste05a, Ste06a, Ste06b, Ste07a, Ste07b, Ste07c, Ste07d, Ste07e, Ste08a, Ste08c, Ste08b, Ste08d, Ste08e, Ste08f, Ste09a, Ste09c, Ste09b, Ste09d, Ste12, TW00]. Micro [Wea97a, Wea97b, Wil95b, Wil97, Yi22f, BCM^{+}14, MB15, MBTS16, Alb04, Alt12b, Ano98c, Ano99e, Ano00e, Ano01b, Ano02a, Ano03a, Ano05, Ano06, Ano07, CS81, Dia95c, Hoo90c, MRLB03, RG07, Tor06, Alt13a, Ano19o, Ano19h]. MICRO-46 [BCM^{+}14]. Micro-Cobol [CS81]. Micro-Ops [Col21]. Micro-RISC [Gon99]. Micro-UAVs [LCN^{+}22]. microactuator [Lan96]. Microarchitectural [Kha00, LPM15, PZL06, RGH^{+}10, SYG^{+}20, Bos06f]. Microarchitecture [AS05, Alb04, Alt12f, BMR^{+}06, BBS^{+}00, DKyL^{+}17, FRB^{+}18, HE07, HS99, KM03, MT05, MS03, MRLB03, MWM99, Red13, RNA^{+}12, SA00, TSW^{+}23, TKS^{+}22, UTB^{+}06, Yi21b, Bos04c, Bos06e, Pap96, RG07, Tor06]. Microarchitecture-Independent [HE07]. Microarchitecture-Level [MT05]. Microarchitectures [PD01, YHHF20, Bos05c]. Microbenchmark [LCY^{+}04]. Microchain [Ano02e]. Microcode [Eng00k, Ste85c, Ste87c, TVT19, VWC03, Abr83]. Microcoding [Mar86]. Microcompilation [Man86c]. Microcomputer [All84, Dun81, Ful91, Lea88, LMC^{+}83, Nic84, SZH82, WM85, Dan89, ES84, FLRB86, GA86, Han81, Hea84, MKNK83, NF81, SM85, SZP81, UBL^{+}82]. Microcomputer-Based [LMC^{+}83, WM85, NF81, SM85]. Microcomputer-Implemented [SZH82, SZP81]. Microcomputers [Kli81a, McK83, Far84, Kli81b, NN81a, NN81b]. Microcomputing [AJ83]. Microcontroller [Cas95, CDGO97, Fan96, JGB^{+}89, MKRC97, STT^{+}15, CH94, ME95]. Microcontroller-Based [Cas95, ME95]. Microcontrollers [AT09, Dea04, Her00, Rag21]. MicroCourses [Ano86a]. MicroDesign [Sla96]. Microdisplay [Dia00]. Microelectronics [ACDG99, Ano99o, GS99, Hoe92, Sak95, SVL03, Sak99c]. Microfluidics [SWM^{+}20]. Microfluidics-Based [SWM^{+}20]. Microlasers [Ano98-27].

Microlithography [Won03].

Micromachines [Ano88g, Kah93c]. Micromouse [Lan85a]. Micromyths [Ste87a]. Micron [HBd^{+}99, Ano02d]. Micronet [vW83]. Micropascal [Man86c]. Microprocessor [AF88, AA93, And82a, ANUN98, AAD^{+}93, Atk91, Bal84b, BAM^{+}93, Bon21, Bor81, CCS21, CS81, CL87, CES^{+}11, Dia96b, Eec15c, Fag96, Fag21, Fai82b, FHR99, FH05, Gol21, GAAR88, HK82, HMS^{+}86, Hen21a, Hsu94, Isa83, IDI^{+}21, JN21b, Joh21c, JN21a, Joh84, Kes99, Kir83a, KS90, KM89, KSM^{+}89, Lan85a, Lee94, LX10, MSA^{+}03, Man92, Mar96, Mil90, MKOK88, MS83, Mye83a, NST97a, Nic91, NH81, NST97b, OS08, PSW91, Phi85, PJ91, Pre21, Rea86, RSS^{+}08, Roe86, SBJ13, Sla96, SAC^{+}99, Smi96b, SDC94, Ste21, SL84b, SM00, TKM^{+}02, Van21, War91c, WEMR04, Web08, Web21, WMC^{+}06, YSMH91, You21, AB83, Ano83, Ano96p, Ano01g, BKM^{+}82, Dan96, DA92, DS95, ERPR95, GDLT86, Gre96f, Hsi91, JC84, JBF94, JA96, KKT^{+}91, Mat96e, MC87, OA81, RH91, Sib84, SL97]. microprocessor [UAN^{+}93, Yea96, Yu96]. Microprocessor-Based [HK82, HMS^{+}86, Joh84, MS83, Mye83a, Hsi91, KKT^{+}91]. Microprocessor-Controlled [SL84b]. Microprocessors [Ano98s, BBS^{+}00, BDJS07, CGMV99, CBLR86, Eec17b, Goo84, GmDT83, Hen96, Hen21c, Her00, Hua89, JM98, Kir84b, LWC^{+}16, LCP^{+}11, LSZ82, Maj87, Mor86a, Mye81, Mye83c, Mye84b, Mye84a, Mye84c, SWK^{+}05, Smi96a, SV21, VM88, Yu96, Ano81, Bos05a, De 83, Far84, Lee95, NM96, Sak00d, mDTG81]. microprogram [OTM82]. microprogrammable [LLC90]. Microprogrammed [BCP01]. microprogramming [Man86b]. Microring [OMMB13]. Micros [Hum84]. Microscale [PLK^{+}16]. microscope [Ano02b, SYG^{+}20]. microsensor [Lan96]. Microservices [GZC^{+}20]. Microsoft [Ano97r, Gre00c, Mar98, Mat93b, Ste94e, Ste95c, Ste98a, Ste13]. Microstandards [Hil87, RT86, Smo86b, Smo87b, Ste86h, Bor85b, Smo87c, Buc85]. Microsystems [Bel96, Mur03, Ano03d]. Microtransducer [HC84]. MicroUnity [Han96]. Microvias [Hol98]. Mid [Smi96b]. Mid-1990s [Smi96b]. Middle [Ste21]. Middle-Aged [Ste21]. might [Ano94b]. Migration [FGC^{+}14]. Milan [Mat21c]. Mile [Gre16a]. Milestones [Gri21, Ano00i]. Military [Kah92b]. Millennium [Ano96d, Her00, Sak00c]. millimeter [SLM^{+}97]. millimeter-wave [SLM^{+}97]. Millionth [Ano99s]. Millip3De [SYW^{+}14]. Mills [Ano14o, Ano17-27, Ano18k]. Mimicking [Boa96]. Mimicry [SJL23]. MIN [SJL23]. MindStorms [Dia99]. minimal [Lee96]. Mining [BH15, FBGB96, LLT^{+}08]. Minitel [Kir90b]. MIPS [HYM^{+}90, MWV92, MBG^{+}16, RJR88, Yea96]. Mirage [Ste89b]. Mire [Mac98]. Misapplied [Gre19d]. Miscellany [Mat92c, Mat10c, Mat12b]. Misconceptions [Ste91f]. misconduct [Ste00c]. Misdeed [Ste94b]. Miss [BRmWH06]. Missed [Mat19]. Misses [HNR10]. Missing [ANJ^{+}04]. Mission [MAT^{+}18, Fer98b]. Mission-Critical [MAT^{+}18]. MissMap [LH12]. MITI [Kah92f]. Mitigating [Bha17, NAA^{+}20, SSC^{+}22]. Mitigation [GXMZ13, WZL20]. Mitsubishi [Ano03b]. Mix [Gre19f, MC87, Rob99d]. Mixed [GXMZ13, KSA^{+}19, LCS92, SKM^{+}16, SJB09, WFW^{+}21, DFR90, RS90]. Mixed-Grained [SKM^{+}16]. Mixed-Low-Precision [WFW^{+}21]. Mixed-Signal [KSA^{+}19, LCS92, SJB09, DFR90]. Mixing [Alb07a]. MKS [Mat93c, Mat97d]. MKS-Toolkit-6.1 [Mat97d]. ML [DtEt22, GLD^{+}22, Joh20c, LBR^{+}22, Maa20, ZRB^{+}22]. ML-Driven [GLD^{+}22]. ML-HW [ZRB^{+}22]. MLPerf [MRC^{+}20, RCK^{+}21]. mm [Ano02c, HOHCV99]. MMX [PW96]. Mobile [ACDG99, Alt13a, Ano99p, Ano00b, Ano14-32, Ano14-33, BBC^{+}15, Cas15, CAV^{+}14, Dav02, GSC97, GHSV^{+}11, Hac01, HIP^{+}22, IPL^{+}23, KCHA21a, KCHA21b, KIM^{+}09, OKH^{+}12, ZHR15, Eng00l, FMN^{+}13]. MOD [NKPC83]. Mode [MNU^{+}15, NS81, ZZ02, ZHZ^{+}19]. Model [ADJK20, BVZ^{+}08, BK14, DQCL24, DPT^{+}21, FCY^{+}20, Gre23e, Han85, Ibb00, KJL^{+}10, KJT^{+}11, NMU^{+}15, NL02, PD01, PC01, SSLV15, SGC94, TML^{+}18, WM85, WPM03, War90d, Han81, SSL82, vdDD90, Ano88c]. Model-80 [Ano88c]. Model-Based [NL02, PC01]. Modeling [Ano15-35, BDH^{+}06, BCA99, Bos06c, Bos21, BBS^{+}00, BDJS07, Can98, GEH^{+}23, IN87, JPOB20, JLSM03, SRWB15, SY06, WPM03, WLY^{+}21a, Bos05d]. Models [ANJ^{+}04, LWK94, LPM15, LSBM17, RT23, SAR10, SNM^{+}13, Ste87e, WMH^{+}10, ZRB^{+}22]. Modems [Ano97c, Tho92, Wal97]. Moderate [LS22]. Modern [Gre23e, HGS^{+}17, HL06, MHP^{+}23, MTS^{+}12, SAC^{+}21, Tab84, DP97, Gre04d]. Modes [DRB^{+}12]. Modiac [NC86]. modified [NKPC83]. Modular [LAT^{+}01, PLK^{+}16, Tab84, WLKN22, YW94, KAK96, SSL82]. Modulation [WM85, TTF96]. Modulators [ZLTW13, DTH^{+}95]. module [Bel93, SPT^{+}92]. Modules [AMFFM^{+}16, BS93, Ano83, HDMT94]. Modulo [MMCH18]. molasses [Ano92a]. Molecular [SWM^{+}20]. molecules [Ano02b]. Mom [Gre09d]. momentum [Gre06a]. Mona [Ste89b]. Money [Gre09c]. Monitor [SL03]. Monitoring [CGLES^{+}23, Ebe03, GAGV22, LP89, MKAC18, Spr02a, Spr02b, ZL16, ZLTW13]. Monocular [IPL^{+}23]. Monolithic [ASX19, BJO^{+}09, BWMS19, CS13, DDG^{+}19, JWS^{+}19, PZB^{+}19, ZSS^{+}19].

Monolithically [YKH^{+}19]. Monopoly [Ano97k, Sla91a, Gre97c]. Monotonic [Gaf91]. Monsoon [ADC00]. Montecito [MB05]. Montgomery [KAK96]. Moore [BKP12, Ano17h, BBB^{+}21, Eec17c, Eec17e, Gre03c, Gre06a, Gre15f, Gre17c, KCXmWH17, VDC17]. MOPED [GSLK11]. Mops [PSW91]. Morello [GBW^{+}23]. MOS [Mea96]. MOSFET [FN94]. MOSIS [Wea97b]. MosquitoNet [CB96]. most [KAK96, Mat96f]. Mothballing [CK11]. motion [KE89]. Motivating [TSS18]. Motivation [JYPP18]. motor [HC83a]. Motorola [Als90, Ano97u, Ano00g, DA92, Fan96, Far84, Gol21, Klo86, MMM84, MF85, McD21, Sib84, Ste12]. Mount [Mat04e]. Mountain [FD04]. Mounted [SP01]. Mouse [Mat91c, Gre99e]. Mouse-Trak [Mat91c]. mousetrap [Par00]. mousetrapping [Ste01c]. Moustache [Ste89b]. Movement [SCH^{+}23]. Movidius [IG15]. Moving [ADJK20, Alb10b, Ano15v, TSP02]. MP [TS91]. MP3 [AML05]. MPA [MBA^{+}09]. Mpact [Kal97]. MPC105 [WBC^{+}95]. MPEG [Ano97-28, IKN^{+}99, KSM99, KSI^{+}96, TWN^{+}99]. MPEG-2 [Ano97-28, IKN^{+}99, KSI^{+}96]. MPEG-4 [KSM99]. MPI [Ano03b, SKC^{+}23, TSA^{+}22]. Mpixel [RT92]. Mpixel/s [RT92]. MRAM [Ano01g]. MRF [NBM^{+}06]. MS [Mat93d]. MS-DOS [Mat93d]. MSI [Pee87]. MTA [Mat97a]. mu [CCD^{+}82, KE89, Eic86, STK88]. Mu-Btron [STK88]. Mu-Pd77230 [Eic86]. Much [Gre03d, Gre09c, Gre13d, Mat05d]. Multi [Ano16-48, Ano16-47, Ano16-46, CML^{+}23, DK18, KAK^{+}22, PEZ^{+}19, RBKL11, SLL^{+}18, Ano16-45, GDLT86]. Multi-Acceleration [KAK^{+}22]. Multi-Core [Ano16-48, Ano16-47, Ano16-46, SLL^{+}18, Ano16-45]. Multi-GPU [PEZ^{+}19, RBKL11]. multi-microprocessor [GDLT86]. Multi-PCB [CML^{+}23].

Multi-Tenancy [DK18]. Multiapplication [Tua99]. Multibit [SMS13]. Multibug [CP86]. Multibus [AQT^{+}92, Kir85a]. Multibus-II [Kir85a]. Multichip [BS93, Bel93]. Multichiplet [ZZNT^{+}23]. Multiclock [BMK^{+}21a, BMK^{+}21b]. Multicluster [CFRM04]. Multicomputer [DSK^{+}92]. Multicomputers [PSW91, Tal93, CK95, Zha91b]. Multicore [ASK^{+}15, Ano10c, Ber09, BSY^{+}10, BBE^{+}11, BSC08, BVZ^{+}08, EBS^{+}12, Goo21, GHF^{+}06, Har12, HAB^{+}09, HWG^{+}09, KJL16, KKD^{+}07, KBH^{+}08, KC09, LC09, LHL09, Mİ09, MBA^{+}09, MKT^{+}13, NMC^{+}08, NKI^{+}09, OKN^{+}11, PGW^{+}20, RAA^{+}21, SAR10, SP09, SMQP10, SMJ^{+}11, UCS^{+}10, VN10]. Multicores [AMK17, ADJK20, AAP^{+}10, KP07]. Multidevice [FCY^{+}20]. Multidimensional [SSA16]. Multidrop [TRY^{+}09]. Multigoal [NHMM23]. Multihop [KCKP14]. multilayer [CT95]. Multilevel [BJW^{+}23, KMN^{+}04, LHM99, NM22, TM17, Ano01f, dG95]. Multimatch [YKL05]. Multimedia [ANUN98, CAV^{+}14, HC99, KMN^{+}04, KSM99, KBN16, NKDN95, Ran97, SSY97, SANK98, TWN^{+}99, UBH^{+}94, Ano99-27, Gol96, Lee95, PW96, TO96]. Multimicrocomputer [FMV85, FK83]. Multimicrocomputer-Based [FMV85]. Multimicroprocessor [AF84, CCD^{+}82]. Multimicroprocessor-Based [AF84]. Multimodal [SMR20]. Multiobjective [MM23]. Multipass [BRmWH06]. Multiple [AH96, GXMZ13, MSA^{+}03, PFC^{+}02a, PFC^{+}02b, WPO^{+}07]. Multiple-Cell [GXMZ13]. Multiple-Clock-Domain [MSA^{+}03]. Multiple-Stack [AH96]. Multiple-Valued [PFC^{+}02a, PFC^{+}02b]. Multiples [Gre03c]. Multiplexed [BUMV95, Jam90, SK97]. multiplexers [Jae82c]. Multiplication [KAK96]. Multiplier [LYP^{+}18]. Multipliers [LZX^{+}18]. Multiprocessing [ABG^{+}16, CJ85, DLCO10, Joh86, KO05]. Multiprocessor [AW06, ACLR89, CD97b, Eck82, EMYN00, Har12, KMAC03, KPP06, LP89, NC86, NIJ^{+}03, Pre91, RLV85, SC91, SLB04a, SLB04b, TS91, YW88, HS85, Hea87, OL85, SSL82, SMCT87, TGF88, WJR88, LDA87]. Multiprocessors [AAW^{+}96, BO86, GSVP03, Kir83b, Kir85b, Kir89b, KL05, MHW03, RTHA05, SKM^{+}16, TM94b, TM94a, WA11, ZL15, AKK^{+}93]. Multiprogram [EE08]. Multirate [CPH90]. Multiresolution [HLIT20]. Multiservice [Yun01]. Multisocket [FRS^{+}09]. Multistandard [KIM^{+}09]. Multitask [DPBW19]. Multitasking [SHTE08, Sch91b]. Multitenant [MFN^{+}17]. Multithreaded [Ano98-28, BGH^{+}12, BBSG11, CSM^{+}21, EHP^{+}07, KST04, KML04, KAO05, RCC12, ROA13, SUF^{+}12]. Multithreading [EEL^{+}97, RG03, WCW^{+}04]. mundane [Mat95c]. Museum [Ing99, SJO01]. Music [STK88, BG81]. Must [SAW^{+}10, Sak99a]. MUTABOR [Kai88]. mutual [OL85]. mW [Kra96, SDF^{+}23]. MXT [TSW^{+}01]. My [Bon21, Mat92a, May21, You21]. MyCS [Ano18-30, Ano16-32, Ano17-39, Ano17-37, Ano17-38]. Myoelectric [KB91]. Myriad [IG15]. Myrinet [BCF^{+}95, CMC98, DBC^{+}98]. Mystery [MGP21, Gre04d]. Mythology [Ste87a]. myths [Rob97e].

N [Bel12, Ste08c]. N-Data [Ste08c]. N1 [PSB^{+}20]. NAE [Ano99q]. Name [HABHW^{+}18, Mil88b]. Named [DKyL^{+}17, Gre15f, RNA^{+}12]. Naming [Ano97k]. NanoBridge [MSB^{+}17]. NanoBridge-Based [MSB^{+}17]. Nanometer [BDJS07]. Nanometer-Scale [BDJS07]. Nanoscale [AMR^{+}06, NBM^{+}06, PDL08, PCDL10, WLD15].

Nanoscale-Integrated [PCDL10]. Nanotube [Ger19]. nanotubes [Ano02c]. nanowires [Eng00g]. Napster [Ste00d]. National [Ano98x, Zsc84]. native [Ano95a]. Navigate [Ano00d, Eng00l]. Navigating [BBB^{+}21]. Navigation [BTK^{+}23, IKK96]. Near [AKK15, AMFFM^{+}16, BCM^{+}14, BG16, CB10, DFG^{+}13, Fai82a, FSS^{+}16, Gon97, HFFA10, KKT13, KZS^{+}22, KCXmWH17, PJB^{+}14, RPL^{+}17, SVA^{+}22, SAC^{+}21, Smi82, Ano94b]. Near-Data [BCM^{+}14, BG16, PJB^{+}14]. Near-DRAM [AMFFM^{+}16]. Near-Memory [FSS^{+}16, KZS^{+}22, KCXmWH17, SAC^{+}21]. Near-Optimal [Fai82a, HFFA10, Smi82]. Near-Sensor [SVA^{+}22]. Near-Threshold [AKK15, CB10, DFG^{+}13, KKT13, RPL^{+}17]. NEC [Cho21, Eic86, KE89, Ste87c]. Need [Ano20y, AAP^{+}10]. Needed [Joh21a, Mat83, Noy85]. Needs [CPS^{+}18, Rob99a, Sla90f, AH96, Shl93]. Negative [KPN^{+}20]. Neocortex [Smi17]. Neon [MMG^{+}99]. Neoverse [PSB^{+}20]. NePSim [LYBZ04]. Net [Ano96u, DMP91, MBK^{+}92, Mye93c, Ste96b, Gre06c, Ste96f, Ste96e]. Netburst [KM03]. NetDistiller [ZFW^{+}23]. NetFPGA [ZACM14]. Nets [SKLY97]. Network [AKJF22, AP07, ASX19, Ano87f, Ano96h, BAH^{+}05, BDF^{+}95, BCF^{+}95, BCKY17, BLW02, BUMV95, CB04, CDS^{+}15, CES17, CB96, CMC98, CJFP95, CG95, CGO00, CLMY96, DMMD11, DJUH16, Ebe03, EPZ02, FH00, FHL^{+}03, Gal97, GSC97, Gil96b, Gre09a, GHY^{+}17, HRK^{+}24, HGPT12, Hor95, IHCE07, JGM^{+}20, KML04, KKP^{+}14, KZ01, KPP06, KTC18, KCKP14, KSB21, LPKP22, LYBZ04, Lyl04, MKG^{+}20, MBH95, Mon97, MBL^{+}02, Mye82b, Mye82c, OGLG^{+}22, OG24, OSS^{+}24, PVS^{+}11, PNDG04, PcFH^{+}02, PC01, Rag84, RCBL00, RKA^{+}20, RMBK81, San97b, SLC^{+}14, SPRK04, SF18, SsSMB24, SKC^{+}23, TLYL04, TSMS23, WHA89, WZL20, WBHv98, YRC^{+}22, ZCW^{+}14, ZZNT^{+}23, ZLBI06, Ano95b, BSB^{+}92, GK97, JRHM86, KWGG95, LC91, Mel87, PHC95, SSB95, Ste94f, UBL^{+}82, VJ89, VTVM94, ZG96, vW83, BWBJ11, GK97]. Network-Assisted [SKC^{+}23]. Network-Attached [RCBL00]. Network-Facing [KML04]. Network-on-Chip [ASX19, DMMD11, KKP^{+}14, WZL20]. Network-on-Wafer [ZZNT^{+}23]. Networked [BDH^{+}06]. Networking [Ano18-32, FMV85, Gre15c, KND02, Mil86, VAFF^{+}10]. Networks [AB14, AK24, BJO^{+}09, BG02, CLL^{+}20, CMB22, CGLES^{+}23, DQCL24, DGT89, Dur96, ED18, EWW^{+}19, EPM^{+}20, For02, Fre02, GQF^{+}06, GHRS89, GR95b, GKS^{+}07, HC02, Hoo89a, IBN^{+}21, Jos86, Koo02, LHL09, MMESGQ22, MM23, Mur89, MCH^{+}94, OESGG^{+}21, ODH^{+}07, RGK19, Rüc02, SB07, SPKJ06, TPV89, WGO^{+}14, YTR^{+}98, YCD^{+}19, BTHS92, Gre15c, RJHK89, VBB95, Wil95b, vdDD90, ACP95]. Neumann [Dor86, Mar86, NGS16, Wil86]. Neumann/Explicit [NGS16]. Neural [BCKY17, BG02, BUMV95, CLL^{+}20, CDS^{+}15, CES17, CG95, DLR02, DGT89, Dur96, EWW^{+}19, EPM^{+}20, ESCB13, FCY^{+}20, GHRS89, GR95a, GR95b, GHY^{+}17, HRK^{+}24, Hoo89a, IBN^{+}21, JGM^{+}20, Kah92c, KKH^{+}24, Kir89c, LNK94, MKG^{+}20, MHW94, MM23, MCC^{+}94, MBK^{+}92, Mur89, MCH^{+}94, Mye93c, OGLG^{+}22, OSS^{+}24, Rüc02, San97b, SJB09, SsSMB24, TSMS23, TPV89, WHA89, YRC^{+}22, YCD^{+}19, BSB^{+}92, BTHS92, KWGG95, PHC95, RJHK89, SSB95, Ste94f, VJ89, VTVM94]. Neural-Net [Mye93c]. Neuro [CR95b, KKL^{+}09, VVRV95]. Neuro-Fuzzy [CR95b, KKL^{+}09, VVRV95]. Neurocomputing [Ang90, Mil87]. Neurocontrol [NNS^{+}93]. Neuromorphic [DSL^{+}18, Eec18b, KHS^{+}23, SVA^{+}22].

Neuroprocessor [SK97]. Neutral

[Dia94a, IO16]. neutrality [Gre06c]. Never [Ste12]. New-Generation [Ano87a, MYK^{+}10, YMA^{+}13]. Newcache [LWML16]. newer [Bos04d, LHN95]. News [Ano91b, Ano95b, Ano96l, Ano96k, Ano96m, Ano96p, Ano97l, Ano97m, Ano97k, Ano98t, Ano98u, Ano98s, Ano98v, Ano98w, Ano98x, Ano98y, Ano98-32, Ano98-33, Ano98-35, Ano98-34, Ano98-36, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99n, Ano99l, Ano99m, Ano99o, Ano99p, Ano99r, Ano99q, Ano99s, Ano99t, Ano99w, Ano99u, Ano99v, Ano00g, Ano00i, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02b, Ano02c, Ano02d, Ano02e, Ano03b, Ano03c, Ano03d, Ano03e, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Dia96a, Eng00a, Eng00c, Eng00b, Eng00e, Eng00d, Eng00f, Eng00h, Eng00i, Eng00j, Eng00k, Eng00l, Eng00m, Eng00o, Eng00n, Eng00p, Eng00g, Mat97a, Mat97b, Mye91b, Ste08f]. newsgroup [Ste96f]. Newton [KE89, NBS^{+}18]. Newton-Euler [KE89]. Next [AC05, AJK^{+}15, Ano01e, Ano02b, AFK^{+}19, BBS^{+}00, BML^{+}21, Cri97, ESG^{+}05, Eec17c, EEL^{+}97, EBC22, Gre10f, Hol98, KSSF10, Kir90a, Lav02, MGP21, Mye89a, PSB^{+}20, Sak02e, SSR21, TIT^{+}13, Web08, YHT^{+}15]. Next-Gen [PSB^{+}20]. Next-Generation [AJK^{+}15, BBS^{+}00, BML^{+}21, ESG^{+}05, EEL^{+}97, EBC22, KSSF10, SSR21, TIT^{+}13, Web08, YHT^{+}15, Ano01e, Ano02b]. Nexus [BCJ^{+}20]. Niagara [KAO05]. NIC [TM17, ZCW^{+}14]. NIC-Switching [ZCW^{+}14]. Nightmail [Aud95]. nightmares [Gre06c]. NISQ [BSA21]. NIST [Ano99r, Ano02b]. nitrogen [Ano01f]. nm [ABG^{+}16, Ano01h, Ano03c, CCA^{+}19, FME18, KBN16, Man09, PAM^{+}07, RDJ^{+}13, SDF^{+}23, TKI^{+}14, XCZ^{+}21]. NN [MKG^{+}20]. No [Ano92e, Gre16c, Joh23e, Mat09b, Mye90, Ste85e, Ste92d, Gre05a, MIM^{+}97, Ste06b].

NoC [OML^{+}07, WZL20]. NoCs [PLBC09, PAM^{+}07, XWZ09]. Node [DSK^{+}92, WN94]. node-crash [WN94]. Nodes [EK16]. Noise [RKK^{+}11, ZRB^{+}22]. Noise-Robust [ZRB^{+}22]. Noisy [MLM^{+}20]. Nominations [Ano15f, Ano16a, Ano16s, Ano16t, Ano16r, Ano17k, Ano17j, Ano17w, Ano17y, Ano17x, Ano17-45, Ano19l, Ano19m, Ano22a, Ano22-40, Ano23t]. Nominees [Ano15c, Ano16d, Ano16e, Ano17i, Ano17k, Ano17j]. Non [KCAR18, Lah84]. Non-Death [Lah84]. Non-Speculative [KCAR18]. Nonblocking [NLM^{+}19]. Noncontact [Sak01f]. Noncontiguous [SKC^{+}23]. Nondeterminism [SKA14b]. Nonelectronic [Mur03]. Nonlinear [KHS^{+}23, Lan96, YCD^{+}19, SSB95]. Nonliteral [Ste90d]. Nonuniform [HFFA10, KBK03, MRSV11]. Nonvolatile [Che19, DPBW19, KLM^{+}15, KTY24, MLL^{+}15, MLS^{+}16, MLL^{+}18, PCW15, WLY^{+}21a, YMC^{+}12, ZHZ^{+}19, Swa19]. Norm [Gre17a]. normal [KHF86]. Northbridge [CH07, OS08, RCC07]. NoSQ [SMR07]. NoSQL [TM17]. Notation [Ber81, Dun81, Dun82]. Note [Kah93i, Joh90b, Ste93d]. Notebooks [Ano98-35]. nothing [Ste95a]. Notification [CNC^{+}16]. Notoriety [Emm07b]. Novel [GXMZ13, Mey04, MEB^{+}20, Sha23b, XPZ^{+}19]. NP [SPRK04]. NP-Click [SPRK04]. NPU [KKH^{+}24, PPO^{+}04]. NS16000 [HF84]. NS32081 [GE86]. NT [Mat97d]. Nucleus [LDA87]. NUMA [BMS16, KSR^{+}99]. NUMA-Aware [BMS16]. Numeric [SG00]. Numerical [AT93, KWM89]. Nuts [Mat03d]. NVIDIA [RTJ21, Bur20, CGG^{+}21, Cho23, LNOM08, RTJ20, BBTV15]. NVLink [FD17].

O [Ano84, BJG^{+}19, BMS16, Ber09, DP97, HSP^{+}01, HSW98, OMMB13, WAA^{+}20]. OASIS [UBL^{+}82]. Obituary [Ano03f, Mor84]. Object [Ano92f, BNOv87, CYH^{+}18, IPL^{+}23, KKL^{+}09, OKH^{+}12, Ste94b, Ano83, Ano97r, Kai88]. Object-Oriented [BNOv87, Kai88]. Object-Recognition [OKH^{+}12]. object-relational [Ano97r]. objects [Mat98b]. Oblivious [YHHF20]. Observations [KBH^{+}08]. Obstacles [Kah93f]. obviousness [Emm06b]. OCOLOS [ZKP^{+}23]. Octocore [MYK^{+}10]. Odd [Alt12c]. Odds [Kah93c]. Odometers [WKK^{+}14]. OEM [MKRC97]. Off [Ano97-32, Ano99j, PH91, WGA^{+}09]. Off-the-Shelf [PH91]. offer [Mar96]. office [Ste89e, Ste91d]. Official [Ano98o]. Offload [DJUH16]. Offload-Enabled [DJUH16]. Offloading [ABK^{+}17]. Offs [AF88, FHP00, Pap96, SMHB91]. Often [SRJ^{+}91, Gre97e]. Okay [Ste07a]. OKs [Ano03b]. Old [Bos03b, Mat06a, Mat06b, LHN95, Mar96, Mat04c]. OLTP [KAV99]. OMIs [Hur97]. Omni [BDH^{+}16]. Omni-Path [BDH^{+}16]. On-Chip [AP07, Bos06d, DSL^{+}18, DPBW19, Fly97, GKS^{+}07, KBK03, KKD^{+}07, KPKJ08, KP07, ODH^{+}07, PKP15, SPKJ06, WWZ^{+}08, WGH^{+}07, HMAF90, TO96]. On-Demand [KCHA21a, KCHA21b, TVT19]. On-Device [PMS23, RT23, SKM23]. On-Line [CJFP95, DO84]. Onboard [SSK23]. Once [NHMM23]. One [Ano99s, Ano17-46, CFZ^{+}99, Chr90, Fer98b, Gre11f, Joh90a, KTC18, LLL^{+}16, LSZ82, McD21, Sel18, Ste09d, Ano94c, Cra90, Pri94b, Ste01a, SO14]. One-Bit [LSZ82]. one-click [Ste01a]. one-hundredth [Pri94b]. One-Millionth [Ano99s]. One-Time [CFZ^{+}99]. One-Time-Programmable [KTC18]. Ongoing [ABC^{+}20]. Online [Ano98-37, Ano01a, Ano15-35, Gre13e, KKH^{+}24, KKSV10, PV01, ZKP^{+}23, Ano98-31]. Only [Ano97q, EKMW02, NHMM23, RCA07, SLZ23]. Ons [Ste92c]. Onto [OSS^{+}24, Ano03e, MBA^{+}09, MM96, Ste02b]. Open [Ano88e, Ano99w, Ano14p, Ano19-30, Ano19-31, Ano20-34, Ano20-32, Ano20-33, Ano20-58, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano23-34, Ano23p, AFK^{+}21, BCJ^{+}20, BC20, CCA^{+}19, CN13, DXT^{+}18, Far87, GV97, GCE^{+}21, Gre15c, Gre16d, HCP^{+}16, HDG^{+}22, JPOB20, KTI^{+}15, LG24, MEB^{+}20, PVB^{+}20, PGW^{+}20, SK02, Sch91a, TGC^{+}20, Urq97, Uss91, War91c, War91d, WDK^{+}20, ZWB19, Gre11e, VAD^{+}21]. Open-Letter [Far87]. Open-Source [BC20, CCA^{+}19, DXT^{+}18, GCE^{+}21, JPOB20, MEB^{+}20, PGW^{+}20, TGC^{+}20, WDK^{+}20]. Open-Standard [GV97]. OpenCL [CS14]. OpenFPGA [TGC^{+}20]. OpenMP [Ano03b]. OpenPiton [BCJ^{+}20]. operas [Gre95b]. Operating [AHK^{+}14, And14, AT09, CR95a, CLM08, FSH^{+}01, Gre95b, Gre23e, HL86, MMB^{+}08, RRP^{+}08, Rea86, RDJ^{+}13, Sak87c, Ste84d, TGE95, vW83, JC84, Mon87, Upd93, WJR88]. Operating-Systems [HL86]. Operation [CCA^{+}19, EDL^{+}04, WGA^{+}09]. Operations [AS91a, ABK^{+}17, JL87, THP^{+}19, Kra96]. Opportunistic [AKJF22, GV06]. Opportunities [AS91b, AC05, AKT^{+}18, Ano20y, BCP04, HAWC^{+}11, IO16, Joh23a, Joh23c, LCN^{+}22, Mei03, MH10, SSH^{+}03, SD21]. Ops [Col21]. Opteron [CH07, CKD^{+}10, KMAC03, KO05]. Optic [EKB^{+}96]. Optical [Alt13e, Ano01f, GAT^{+}22, Kah91c, KB13, KKD^{+}07, KL05, LNK94, LHN95, MA94, PDL08, SLC^{+}14, SSB95, STR^{+}13, TMBT94, TRY^{+}09, TMJ13, TIT^{+}13, WAA^{+}20, WCH94, YTR^{+}98, Ano92a, Lou91, RLG94]. Optical-Disk [MA94]. Optically [CK95, KL08]. Optics [Ano02e, TMBT94, Eng00j, LHN95]. Optimal [Fai82a, HFFA10, Smi82].

Optimists [Gre16d]. Optimization

[AML05, Kid14, KOKA23, KAV99, PMM15, PVS^{+}11, PSL^{+}23, SWG06, SW14, TLYL04, TATC09, WWZ^{+}08, WAA^{+}21]. Optimizations [CWLS15, YKH^{+}19, ZKP^{+}23]. Optimize [CES17, OESGG^{+}21, Boa96]. Optimized [BMK^{+}21a, BMK^{+}21b, CAV^{+}14, HSR18, RGF96, SLC^{+}14, SABS20, RGF95, Rya88]. Optimizer [KKL^{+}00]. Optimizing [Ano97w, DPBW19, Dra00, GTF97, GHLK^{+}12, JAS^{+}22, NM22, PEZ^{+}19, CDGO97]. Options [Ano98-38, Ano16-33, Ano16-34, Ano17-43, Ano17-40, Ano17-44, Ano17-41, Ano17-42, Ano19-30, Ano19-31, Ano20-58, LBD^{+}99]. Optisim [KL08]. optoelectronic [BUMV95]. Opts [Han84]. Oracle [AJK^{+}15, FJL^{+}13, GJLT12]. Oral [Ste09a]. Order [Ano98v, CCA^{+}19, Gre11e, HNR10]. Ordered [JSY^{+}16]. Ordering [CL04, Gus85, KCAR18, LSBM17]. Organic [Ano88f, Ano02d, Pri94b]. Organization [DA92, Ano94c]. Organized [Gre18c]. Organizing [Dia93d, RGR95]. Oriented [BNOv87, PSL^{+}23, PHB15, Sak87c, THP^{+}19, Kai88, Mon87]. Origin [Col21]. OS-X [Ano98r]. Oscillators [TP10]. Other [Alt14c, War92b]. Our [Ano20x, Eec16d, Gre09d, Gre20a, Mye84d, Alt14e, Ano97n, Gre97f, Gre97e, Mat95b]. Out-of-Order [CCA^{+}19]. Outlier [SS16]. outlines [Mat96b, Sla96]. Outperforms [Ano88c]. Output [HSN^{+}23, PKP15, HP85]. Outsider [Ano18-31, Wil96]. Outsiders [Gre15b]. Outsourcing [Gre05d]. Outstanding [LE18]. Overclocking [MMC^{+}22]. overcome [DP97]. Overcoming [CSC^{+}05, DGM^{+}11, Emm06b]. Overflow [PZL06]. Overhead [KTY24, JKN96]. Overheads [SMS13]. overlapped [Dv87]. overlapping [Fur88]. Overtake [Ano96d]. Overturns [Ste84a]. Overview [HCU^{+}07, HYS98, Kir87, Koe86, Lee90, NJZL^{+}17, SKO89, VM88, YBS17, OA81]. Owns [Alt11b, Ste84c]. Oxide [STT^{+}15, TKI^{+}14, IWM89]. Oxymoron [Gre06d].

p [RGF96, Sav99a, YMC^{+}12]. P1014 [Fis85]. P1073 [FO89]. P1296 [RT86]. P1394 [Dia95d]. P1754 [War91c]. P694 [Bal84b]. P854 [Ste84e]. P896 [All81]. P959 [Ano84]. PA [Kum97]. PA-8000 [Kum97]. PA7100 [AAD^{+}93]. Package [Can98, Lin98, Sha23c, Trö98, WAA^{+}20, Ano01h]. Packages [Han87, Hol98, Jef84]. Packaging [Ano98-34, Ano98n, Far85, Has94, Her93, Mis93, JBF94]. Packard [Ste93a]. Packet [AML^{+}03, BLW02, BJ14, DMMD11, DKSL04, LL03, LK10, MIM^{+}97, OESGG^{+}21, RMM^{+}04, XLX^{+}23, YTR^{+}98, YKL05, ZBH^{+}00]. Packet-Switched [YTR^{+}98]. Packets [GM00, PPP01]. PACMAN [RNLY23]. Page [Mat21b]. paged [CM86]. Paging [GHS17]. painless [Mat95d]. pair [War91g]. PAL20RA10 [BC86]. Palm [Ano15h, Ano15i, Ano17n]. Palms [Hen21b]. Palmtop [Mye92a]. PAM [Das21, Sha22]. PAM-4 [Das21, Sha22]. Pandemics [Gre20c]. Panel [GAT^{+}22, HCPS03]. Paper [Ano99i, Bro17, Gon18, KT14, Lah84, Mar14, Tor12]. Paperback [Ste90c, Ste90d, Ste91f, Ste91g]. Papers [Ano09c, Ano10b, Ano14c, Ano15d, Ano15t, Ano16q, Ano17l, Ano17v, Ano19o, Ano19p, Ano20u, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano21u, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22-47, Ano22-48, Ano22-49, Ano22-50, Ano22-51, Ano22-52, Ano23-27, Ano23w, Ano23y, Ano23z, Ano23x, Ano23-38, Ano24n, Mat87, RY21, YT01, Yi23c, Ano00c, Ano14b, Ano15e]. Paradigm [Mil87, WMH^{+}10]. Paradigms [Bos03b, Mat08b, Ste97e]. Paradox [Gre18d, Gre96c, Gre04d, Gre04e]. Paragon [DK14]. Parallel [AFH16, AS90, AHO^{+}90, Ano17l, AAP^{+}10, ACG^{+}95, BSP^{+}17, But07, CFK^{+}10, DLR02, DKSL04, DGM^{+}11, ECK^{+}22, EKB^{+}96, FBGB96, GLN^{+}08, GSP02, Gro94a, HCW^{+}04, JBM95, KTK13, KNN^{+}90, KDK^{+}11, KII09, Lea88, LBS^{+}11, LHN95, MA94, MT03, Mye84a, OVT90, PZK^{+}18, RPL^{+}17, SAR10, SHTE08, SWL90, SKL^{+}92, WMH^{+}10, XLX^{+}23, Dia94b, FMT91, Hsi91, Kah90c, Lou91, OTM82, SMCT87]. Parallel-Readout [MA94]. Parallelism [CJH^{+}12, DD05, EV97, FZW^{+}12, FCY^{+}20, GHN^{+}12, GSS09, JSY^{+}16, Lee96, MM09, PDS^{+}13, SWG06, TCC^{+}00, TT12, FMT91]. Parallelism-Aware [MM09]. Parallelization [GJLT12, LHC^{+}12, PEZ^{+}19]. Parallelizing [Aug12, CO03, MBA^{+}09, AAW^{+}96]. Parameter [UTB^{+}06]. Parametric [KKT13]. paranoid [Ano97q]. Pareto [LZX^{+}18]. Paris [Kir85a]. park [NF81]. Part [CD97a, CD97b, EGL^{+}90a, Gre08d, Gre15d, Gre15e, Sta01a, Sta01b, Ste97d, Ste04a, Ste04b, Ste17c, Ste17a, Ste17b, Ste18, Ste90g, Ste90h, SLB04a, SLB04b, TM94b, TM94a, WHKM93a, WHKM93b, Yi24b, EGL^{+}90b, PFC^{+}02a, PFC^{+}02b, Ste83c, Ste83d, Ste99b, Ste00b, Ste00c, Ste00a, Ste02a, Ste08d, Ste08e, Ste14a, Ste14b, Yi22a, Yi22b, Yi22c, Yi22d, Yi22e, Yi23a, Yi23b, Yi24a, ZMVH^{+}83c]. Partha [Sco14]. Parthasarathy [Sco14]. partially [Joh90b]. Participant [Dan96]. participants [Ste98e]. participation [Dia95e]. Parting [Moo03]. Partitionable [SC24]. Partitioned [PMM15]. Partitioning [CMR97, CFRM04, NKI^{+}09, SK12, VM95, WBKR14]. Partitions [MMESGQ22]. Partners [Ano02d]. Partnerships [Eng00m]. Parts [PH91]. Party [Emm07e, Gre24b]. Pascal [FD17]. Passing [XLW^{+}12]. Past [Alt11e, Chu18, Gri21, Hoo89b, Mat95e, Mor86a, WS90, Ano01d]. Patching [SNC^{+}07]. Patent [Ano99t, Emm06f, Emm06c, Sla90b, Ste93a, Ste07d, Ste09b, Yi21a, Yi22a, Yi22b, Yi22c, Yi22d, Yi23a, Yi23b, Yi24b, Yi24a, Emm05a, Emm06a, Emm06d, Ste01a, Ste04c, Ste04d, Ste05a]. patentable [Emm05d]. patented [Ste98b]. Patentees [Ste07a]. Patenting [Ste96d, Ste96c, Yi21a, Yi22a, Yi22b, Yi22c, Yi22d, Yi23a, Yi23b, Yi24b, Yi24a]. Patents [Alt14d, Emm05b, Ste90a, Ste90f, Ste93b, Ste96e, Ste03a, Ste08d, Ste08e, Ste14a, Ste14b, Yi21b, Yi21c, Yi22f, Yi22e, Yi23c, Emm06e, Ste95d]. Path [BDH^{+}16, Lau21, Abr83]. Path2SL [MMESG^{+}20]. Pathologies [BMV^{+}08]. Pathways [Ano18-31]. Patients [CJFP95]. Patt [Bel12]. Pattern [Ano15-36, Rob92, WHA89, BSB^{+}92, RLG94]. Pattern-Addressable [Rob92]. Patterns [Mat08a, PZK^{+}18, WSZS05]. Patterson [Pri93a]. Pax [Kah90c]. Payment [DVQ96]. Payoff [Gre12a]. pays [Gre96d]. PC [RMFG85, Ano98l, Ano98t, Bus86, Dia94b, Gol96, Gre98c, Han87, Hig85, JBM95, Jef84, Mat92c, Mon97, Mor88, Ran97, Ste05b]. PC-Based [Mor88]. PCB [CML^{+}23]. PCI [ZCW^{+}14, Das21, Gil96b, GK97, LMVP05, OKN^{+}11, Sha22, SC24, WBC^{+}95]. PCI-based [GK97]. PCMCIA [War92b]. PCs [Ano99p, Gre00e, Mat21b]. PCs/laptops [Ano99p]. PD77230 [KE89, Eic86]. PDAs [Eng00j]. Peach [OKN^{+}11]. PEFL [KGT22]. Penalties [Ste92e]. Penalty [Bur96, Pit95]. Pensando [GM21]. Pentium [Ano03d, AA93, Ano98-33, Ano99w, Ano99-28, Ano03b, BM95, Pap96, Pri95, RPK00, Spr02b]. Pentium-II [Ano98-33]. Pentium-III [Ano99-28]. people [HC83a]. PEPPHER [BPT^{+}11]. Per-Thread [EE10]. Perceived [SMR18]. Perceived-Color [SMR18]. Percent [Gre20f]. perceptrons [CT95]. perfect [Sak01d]. Perform [MSS15]. Performance [AS22, AF88, Ano18d, ACLR89, AAD^{+}93, Atk91, AT93, AJC^{+}20, BcFP06, BFZ^{+}22, BCU^{+}99, BAH^{+}05, BDH^{+}16, BMV^{+}08, Bos03c, BPUH06, BGH^{+}12, BBSG11, Car93, CRV^{+}04, CCYT05, CCE^{+}09, CDS07, CGMV99, CGF18, CGG^{+}21, Cho23, CS08, CMAS11, Cum04, DD05, Dav98, Dia96d, DVWW05, Eec15d, ECY^{+}12, EEKS07, EE08, FD17, For02, FGC^{+}14, GLD^{+}22, GHPS93, GV97, GBW^{+}23, HO99b, HL99, Hua89, HKC10, HML^{+}21, HcF04, IN87, JRHM86, JLG19, JGF98, Jos86, KMG^{+}03, KK10, KBH^{+}08, KCS^{+}20, LNV89, LLZ^{+}04, LLW^{+}07, LCP^{+}11, LCY^{+}04, LMVP05, MR85, MT03, MRC^{+}20, Mel87, MRSV11, MKAC18, MKOK88, MCV^{+}14, Mor86b, MBK^{+}92, MM09, NFQ03, PKL13, PcFH^{+}02, PLB06, QJP^{+}08, RG03, RCK^{+}21, RSW10, RFGM86, RC13, RBKL11, Sak00c, Sak02a, SWG06, Spr02a, Spr02b, SZH82, TMJ13, TMA18, WHCK18, WFW^{+}21]. Performance [WEMR04, WJM^{+}05, WMC^{+}06, XYT^{+}23, Yeh07, YHT^{+}15, AO97, Ano03b, BM95, Bos05a, Bos05b, CJFP95, CFM^{+}97, DBDF97, De 83, Fis85, Gil96a, GK97, Hsi91, Iac88, Ipe19, Jae83, Jag97, KKC93, MC87, NN81b, OL85, OB91, Pap96, PW96, PGL97, SZP81, TO96, WHKM93a, WHKM93b]. Performance-Directed [LLZ^{+}04]. Performance-Monitoring [Spr02a, Spr02b]. Peripheral [Sch91b, LC91, NA84]. Peripherals [All84, Nic84]. Perish [Smo86a]. Permutation [LSY01]. Persistency [KPR^{+}22, KGS^{+}19, PCW15]. Persistent [KJC^{+}23, Sol19, US23]. Person [Chr90, Joh90a]. Personal [EI87, EIB90, Kir89d, Kir91c, Mat02c, MAT85, Mye82d, Mye85a, Ond96, Sha96, LLC90]. Personal-Computer [Kir89d, Sha96]. Personalized [KZS^{+}22]. Perspective [AAW^{+}96, Bos21, Dan96, GMC18, Mat94, Gre97d]. Perspectives [LKGL24]. Pervasive [Ano16-40, Ano16-39, Ano20-53, Ano21t, Ano23u, Joh22a, Mat01e]. Petaflop [SB23]. Petascale [HGPT12, MYK^{+}10]. Petri [DMP91, SKLY97]. Petri-Net [DMP91]. PFS [Mye85b]. Phase [CEAY23, LZY^{+}10, SWL11, Ano02b]. Phase-Change [CEAY23, LZY^{+}10, SWL11]. Phases [IBM05, SPH^{+}03]. Phi [SGC^{+}16]. Phillippe [Ste95c]. Philosophy [Kli81a]. Phone [FH00, Ste17c, Ste17a, Ste17b, Ste18]. Phones [Ano97-27, STM02]. Photobit [Ano99t]. Photonic [HAC^{+}13, KNB14, OMMB13, PLBC09, SB07, ZZNT^{+}23]. Photonics [BJO^{+}09, Gun06]. Photoshop [Ano98z]. PHY [Sha22]. Phylogenetics [ABA^{+}21]. Physical [NBS^{+}18, PVS^{+}11]. physically [HP85]. Pi [Ano17-58, Ano17-59]. PIA [Han81]. Picks [ABZ08, Alb04, Alt12e, Alt13c, Alt14f, Ano19o, BL23, CS15, Dwa19, Eec15e, Eec16e, Eec17f, Eec18f, ET09, FL13, FV12, GT22, HGPT12, JQ17, Jim21, Joh19f, Joh20d, Joh22f, Joh23f, Kim20, Kur21d, MS16, MRLB03, Mud10, PM11, RG07, TM14, Tor06, Wen18]. Picojava [OT97, HO99b]. Picojava-I [OT97]. pie [Ste96b]. Piepho [Luu90a]. Piezoelectrics [SP01]. Piles [Ste02b]. PILOT [Ano91c]. PIM [KKL^{+}22, LS22, THP^{+}19]. PIM-Based [LS22]. Pinnacle [TSW^{+}01]. Pioneer [Alt11c, Ano03f]. Pipelined [SC24, SsSMB24, TK21, XWZ09, Gal97, Iac88, WE93]. Pipelines [BRmWH06, SRA^{+}04, WHKM93a]. Pipelining [KKL^{+}09]. Piracy [Hau88b, Kar88a, Ste88c]. Pirates [Edw83, Kar88b, Ste97c]. PISQ [BSA21]. Piton [MFN^{+}17]. PIUMA [LPKP22]. PivotPoint [Cum04]. Pixel [KII09]. Pixel-Parallel [KII09]. PLA [Ano91b].

Placement

[BR21, CWLS15, HFFA10, ZRA^{+}20]. Plagiarism [Alt13f]. Plain [Pfa94]. Plaintext [TSMS23]. Plan [SRJ^{+}91, War92c]. Planet [BDH03, KGMT17]. planning [Ano94c]. Plasticine [PZK^{+}18]. plastics [Ano02b]. Platform [ABG^{+}16, Ano00m, BYM^{+}07, DMG^{+}15, EEL^{+}97, Gre13f, HC02, MAS^{+}05, MBSP02, Man09, MBA^{+}09, NIJ^{+}03, PSB^{+}20, SK02, SP09, SMR20, SWM^{+}20, Eng00l, Gon97, GBW^{+}23]. Platforms [BEL^{+}23, BSY^{+}10, Gre98e, Gre09a, Gre13b, JMZ^{+}11, MM23]. Play [NM99, Gre97c]. playing [Gre96e]. PlayStation [Ano03d]. PLDs [CH94]. Plod [ACG^{+}88]. Pluton [Mat21b]. Pod [OZT^{+}22, WLF^{+}08]. Pod-racing [OZT^{+}22]. Podcast [Ano22-71, Ano22-72, Ano22-73, Ano23-70, Ano23-71, Ano23-72, Ano23-73, Ano23-74, Ano23-75]. Poetry [Gre09d]. Point [BSC^{+}90, CCG^{+}84, DKB^{+}90, Del93b, DM88a, FGG^{+}88, GE86, HC83b, Joh89, LKJ^{+}22, MD88, PS88, RJR88, SKL^{+}92, SK88, Ste84e, ZSB21, Iac88, KWM89, OZT^{+}22, SL97, DM88b]. Pointer [RNLY23]. Pointers [Mey04]. Polarized [CMB22]. Policies [SKJ^{+}11]. Policy [Gre02a, Gre11c, Gre17d, Gre21b, SJL23, Ste89a, Ste15b, Wet86, Zsc84, Gre19d, Kir01, Ste89d, Ste90e, Ste01e]. pollinate [Ano17p]. Polymorphous [SNL^{+}03, WGM02]. Polyp [MSB87]. Pooled [HRC^{+}23]. Pooling [GKB^{+}23]. Pools [BEL^{+}23]. Pop [Ste04a, Ste04b]. Pop-Ups [Ste04a, Ste04b]. popular [KAK96]. Porcupines [Ste88b]. Portability [SSLV15]. Portable [CWLS15, Has94, LS98a, MKRC97, Sto94, Str98, THT^{+}04, Dia95d, Seg97]. portable-computing [Dia95d]. portables [Ano98-30, Ano98-29, Ano98-31]. Portal [KFF00]. ported [JKP89]. Position [LS22, Ste99b]. Positioning [VWC03]. POSIX [IJ98]. POSIX/UNIX [IJ98].

Possibilities [Sak02c]. Possibility [Ano88f]. Possible [Ano98-32, NM96]. Post [Ano17h, BBB^{+}21, KCXmWH17, VDC17]. Post-Moore [Ano17h, BBB^{+}21, KCXmWH17, VDC17]. Postquantum [MGG^{+}19, NM24]. posts [Ste96f]. postscript [Ste00a]. Pot [Mat99d, Mat99e]. Potential [HSW98, IG15, ML21, Ste07c]. Pourri [Mat99d, Mat99e]. Power [ACG03, AMR^{+}06, Alt12d, Ano97g, Ano98-36, Ano17-57, ASD^{+}05, BCKY17, BAM03, BBS^{+}00, BDJS07, BS17, BWBJ11, BCD^{+}11, BGH^{+}12, BvdGM^{+}15, CL05, CDS07, CDY^{+}18, CR95b, CHAF22, CEP^{+}17, CJFP95, CBJ10, CK11, DD05, DRB^{+}12, Eec15b, Eec17e, ERM08, EDL^{+}04, ECY^{+}12, Fla99, FMN^{+}13, GDN^{+}17, GZC^{+}17, HRK^{+}24, HKY^{+}95, JLSM03, KK10, Kid14, KSLY17, KPN^{+}20, KOKA23, LAT^{+}01, LYBZ04, MLS^{+}16, MKG^{+}20, MKP06, Mye89a, NKDN95, NIJ^{+}03, OKH^{+}12, OMMB13, OYS^{+}11, PO04, PRE11, RTHA05, RCC12, RC13, RNA^{+}12, SWG06, Seg97, Sha22, SBG^{+}07, SCC^{+}05, SYY^{+}11, TCD^{+}05, VW03, WAA^{+}20, WPM03, WS13, WK13, WPH^{+}23, WJM^{+}05, WSZS05, YBS17, Yeh07, ZZ02, ZZ05, ZHPR17, Ano02c, Bos04b, Bos05b, Bos05e, Fly97, FN94, Jag97, Kra96, Lan96, PGL97, Sak99d, Bos21]. Power- [ACG03, Alt12d]. Power-Aware [BBS^{+}00]. Power-Conscious [TCD^{+}05]. Power-Constrained [WK13]. Power-Efficient [BvdGM^{+}15, MKG^{+}20, MKP06, RTHA05, WSZS05]. power-Lessons [Bos04b]. Power-Management [FMN^{+}13, RNA^{+}12]. Power-Performance [WJM^{+}05]. POWER10 [STSM21, SCH^{+}23]. Power4 [BTR02]. Power5 [KST04]. Power6 [RSS^{+}08]. Power7 [FAWR^{+}11, KSSF10, LDF^{+}13]. Power9 [STKS17]. Powered [CHSL17, GKL^{+}14, XPZ^{+}19]. PowerPC [AAWC94, Ano96f, Ano03c, BAM^{+}93, DOH94, DS94, DDHS00, Mat94, MWM99, PVYU94, SDC94, SF95]. PowerPC-604 [SDC94]. powertrain [HDMT94]. Practical [CD09, GLD^{+}22, PS15, SSA16, WFA^{+}10, YFDV19, YMC^{+}12]. Practicality [PBT06]. Practices [SMM^{+}22]. Praised [Smi86a, Smi86b]. Pre [Bos06c, LDL17]. Pre-Charge [LDL17]. Pre-Silicon [Bos06c]. Prealignment [KYG19]. Precise [MV96, Iac88, WE93]. Precisely [Chr91]. Precision [CT95, CHAF22, DPT^{+}21, HIP^{+}22, LCN^{+}22, WFW^{+}21, YAK18]. precluding [BD94]. Predicated [KMPS06]. Predication [KJMP07]. predict [Ano02c]. Predicting [BD94, HRSS11, RGH^{+}10, TW00]. Prediction [FSR^{+}05, Gre98f, MBG^{+}16, PS15, SSA16, YBNS15]. Predictions [Alt13b, Gre08a, IBM05, ZZ02]. Predictive [Ano16-40, Ano16-39]. Predictor [HCP^{+}03, SJB09]. Predicts [Pri93a]. preempted [Ste97f]. Prefetching [KST12, NS05, WFA^{+}10]. Prefix [ANC05, CM04]. Prepare [Ano17-47, Ano17-48]. Preparing [Dia95e, HC99]. preposterous [Ano95d]. Present [Bor99a, Gon97, Gri21, Hoo89b, Kni85, WS90]. Presenting [Sak91]. presents [Mat96b]. Preserving [Bha17, Ven23]. President [Ano01d, Eng00j]. President-elect [Ano01d]. Presilicon [Bos05d, Bos21]. Pretext [Ste00c]. Prevailing [Gre12d]. Preventing [AVU^{+}08, Ste01e, Kir01]. Preview [DRM^{+}23]. Price [Eng00j, Gre02d, Gre07a, Ste15b, Mor84]. Pricing [Gre01e]. Primer [ABC^{+}20]. Printer [Han85]. Printing [TM81]. Prints [Ste89b]. priorities [Bos04d]. Priority [Kah93i]. Privacy [Ano99j, Ano99n, Ano19-27, Ano22-71, Ano22-72, Ano22-73, Ano23-70, Ano23-71, Ano23-72, Ano23-73, Ano23-74, Ano23-75, DCMS20, Joh23c, KGT22, Lea85, LG24, Ven23, Ano99p, Mat95d, Ano19j, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21l, Ano22n, Ano22o, Ano22p, Ano23q, Ano23r, Ano23s, Ano24m, Ano24g]. Privacy-Encoding-Based [KGT22]. Privacy-Preserving [Ven23]. private [Gar93, ZG96]. Privileges [Gre17b]. prize [Ano99q]. Pro [Ano96g, Pap96]. PRO3 [PPO^{+}04]. Probabilistic [NBM^{+}06, WLD15]. probes [Ano01c]. Problem [BM85, Gre20b, Hoo89a, Moo03, VPV12, Bal84a]. Problem-Solving [BM85, Hoo89a]. Problems [CD97a, LG24, Maa20, Mat90b, Mye84c, VL00, BD94, Dur96, LHN95, SCG95, WCH94]. procedure [AGH^{+}91]. Process [Ano87e, Ano97v, Buc84, HBd^{+}99, Kid14, Kir87, LCWB08, MS84, MB15, NPY^{+}21, Rob98a, Emm05c]. Process-Control [Kir87, MS84]. processes [Ano01c, LC91]. Processing [APS98, ARS03, AKK15, Ano10c, Ano17l, AF84, AMFFM^{+}16, BCM^{+}14, BG16, BBC^{+}15, BB17, BDV^{+}08, BCF^{+}14, BLW02, BJ14, BvdGM^{+}15, CWL^{+}14, CS81, CEP^{+}17, DSK^{+}92, DKK21, DDHS00, Dur96, DM88b, DM88a, Fet95, GAR^{+}06, GU98, Goo21, GHF^{+}06, HABHW^{+}18, HOHCV99, JYPP18, KNN^{+}90, KYGW17, KZS^{+}22, KDK^{+}01, KIR19, KKL^{+}22, KKH^{+}24, KBN16, LCS92, LL03, LS96, LHC^{+}20, LKGL24, Mil87, MCC^{+}94, Mor86a, MD88, NG87, NM24, NY22, OLT^{+}23, PPA^{+}14, PKR92, PP92, RMM^{+}04, SG01a, SP92, SML04, SKL^{+}92, TONH96, VWC03, WSM^{+}10, WLP^{+}15, AHO^{+}90, Ano92b, Ano95a, BTHS92, DO84, EKM^{+}95, FMT91, Gol96, Han96, Lee96, RMFG85, SPT^{+}92, Wv92, OGLG^{+}22]. Processing-in-DRAM [OGLG^{+}22]. Processing-in-Storage [KYGW17]. Processor [AAC^{+}23, AO97, AJK^{+}15, AML05, Ano97-31, Ano98-33, Ano99m, AFK^{+}19, ASD^{+}05, ACRV96, AOYS95, BH15, BJO^{+}09, BY07, BBTV15, BSP^{+}17, BCKY17, BCA99, Bos03c, Bos21, BWBJ11, BGK97, BCD^{+}11, BGH^{+}12, BvdGM^{+}15, But07, BML^{+}21, Cat88, CCA^{+}19, CCE^{+}09, CS08, CKD^{+}10, CAH86, DSK^{+}92, DLR02, DSL^{+}18, DMWS13, EGL^{+}90b, EGL^{+}90a, Eic86, EKM^{+}95, FZW^{+}12, FJL^{+}13, FOP^{+}19, Fra00, FRB^{+}18, FGG^{+}88, FMN^{+}13, GG99, Gas21, Gon00, Gon06, GR92, HMB^{+}14, HIP^{+}22, HRK^{+}24, HO99b, HYM^{+}90, HMR^{+}19, HSW98, HHNK09, HVS^{+}07, HWG^{+}09, IPL^{+}23, KST04, KSSF10, KML04, KMAC03, KJMP07, KJP^{+}13, KKP^{+}14, Klo86, KII09, KAO05, KHS^{+}23, KPHP04, LKJ^{+}22, Lin06, LXB07, LSZ82, LYBZ04, MLL^{+}15, MLS^{+}16, MAS^{+}05, MYK^{+}10, MAT^{+}18, MHW94, MFN^{+}17, McL93, MS03, MB05, MD20, Mey04, Mil88c, MC95, MWV92, Mor86b, MBG^{+}16, NSN^{+}93, NGS16]. Processor [OG01, OW01, PS88, PZB^{+}19, PVS17, Qua00, RPK00, RMM^{+}04, RFGM86, RKA^{+}20, RDJ^{+}13, RBGZ19, RMC04, STKS17, SCV01, SWM87, SSMI87, SNC^{+}07, Sav99b, SKLY97, SZZ01, SA00, SLL^{+}18, SK88, SKW^{+}23, STR^{+}01, STSM21, SCC^{+}05, SVC01, STS^{+}92, SHKS19, SCH^{+}23, SSB20, SSR21, SUF^{+}12, SANK98, SMS13, TCD^{+}05, TWN^{+}99, YNS^{+}14, Yeh07, YMA^{+}13, YHT^{+}15, ZWB19, ZLBI06, ZZ05, AKK^{+}93, Ano96k, Ano01c, Ano03c, AH96, BCF^{+}92, BM95, Chr96, DVQ96, Dur96, FL84, HS92, ISH^{+}91, Jag97, KY91, KBW95, Laz89, LLC90, OTM82, PK88, Rob91, RT92, Sak99a, TO96, VTVM94, WHKM93a, WHKM93b, GHSV^{+}11, WGH^{+}07]. Processor-Based [ZLBI06]. Processor-System [PZB^{+}19]. Processor-to-DRAM [BJO^{+}09]. processor/controller [BCF^{+}92]. Processors [Ano01a, Ano17-57, AS99, BCP01, BSC08, BS17, CB04, CRV^{+}04, CDY^{+}18, cCCP00, CFRM04, Cra00, CSC^{+}05, DtEt22, Eec17e, EEL^{+}97, FAK^{+}14, GAR^{+}06, GH88, Gro92a, Gro92b, GHLK^{+}12, HNR10, HL06, JW20, KJL16, KP03, LC09, MLL^{+}18, MH10, MBK^{+}92, NKI^{+}09, OKH^{+}12, PKL13, PNDG04, PO04, PV98, PV01, RCR04, RKK^{+}11, ROA13, SP09, SDB^{+}04, SPRK04, SKL^{+}92, ST21, Sla90f, SYY^{+}11, TLYL04, Vei04, WK13, WMSH09, WPO^{+}07, XPZ^{+}19, XYT^{+}23, ZHPR17, Bos04e, DFR90, SU95, WE93]. Procreation [Ste88b]. Produce [Gre23d]. Producing [Mat87]. Product [Ano91a, Ano97x, Ano97y, Ano98-39, Ano98-40, Ano98-41, Ano98-42, Ano98-43, Ano99-29, Ano99-30, Ano99-31, Ano99-32, Ano00j, Ano00k, Ano00l, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano02f, Ano02g, SBE01, SGC^{+}16, Tab91, BC86, Dia99, Pap96, Ste98d, Wal97]. Production [Eng00b, Min84, RKK^{+}11, Ano01c, Ano02c, Ano03d]. Productive [Alb07c, BPT^{+}11, SPRK04, WPSR20]. productivity [Gre96c]. Products [Ano98-30, Ano98-29, Ano98-31, Ano99z, Ano99-27, JW99, Kir87, Lee21, PS20, AQT^{+}92, Seg97, Ste04c, Yu96]. Professional [Ste90b, Ano20s, Ano20r, Ano20p, Ano20q, Ano21q, Ano21r, Ano23v, Ano24i, Ano24j]. professionals [Ano94b]. Profile [CHH^{+}98, KSI^{+}96]. Profile-Directed [CHH^{+}98]. Profiles [Bea90]. Profiling [AJC^{+}20, KDH^{+}16, RTM^{+}10]. profound [Mat95c]. Program [Ano98p, Ano13d, DCMS20, Kah92f, RGH^{+}10, SPH^{+}03, CFM^{+}97, MF85, Ste93d]. Programmability [CGF18, MT03]. Programmable [AAC^{+}23, AB14, ABG^{+}16, Ano98y, ABK^{+}17, BCF^{+}92, BI13, BS93, CFZ^{+}99, FME18, Ham00, HV04, JGM^{+}20, KSA^{+}19, KTC18, KHS^{+}23, LL03, LPL86, Lee90, LM16, MKM15, SNC^{+}07, SP09, Ste86a, Sti11, WHJ^{+}23, ZBH^{+}00, ZUNN18, ZMVH^{+}83c, ZVHL85, GDLT86, MST^{+}85, Man86b, Man86c, ZMVH^{+}83a, ZMVH^{+}83b].

Programmed [Ste86a]. Programmers [AAP^{+}10, Sha82]. Programming [ANJ^{+}04, Ano93, AAP^{+}10, BVZ^{+}08, ECK^{+}22, KMK01, KAK^{+}22, LNV82, Mat93e, Mat99a, Mat99c, Mat02d, OS99, Rit97, SAR10, SSLV15, Tab84, WMH^{+}10, Yao85, KWM89]. Programs [AAP^{+}10, CO03, Dun81, ESCB13, LPC12, SMR18, Ste84b, TKM^{+}02, AAW^{+}96, Hea84]. Progress [Kah92b, MLS^{+}16, RGK19]. Project [Ang90, Ano98p, Ano99r, Bon21, Kah91a, Mat01f, CCD^{+}82, CFO^{+}18, DBC^{+}98, FOP^{+}19, RD90, Sak87d, Ste99b]. projected [Ano01c]. Projecting [JC08b]. Projects [Ano10c, Mat03c, Sak89, Smo87a, Ano97s, Ano99u, Gus92, Rob97a]. Prolegomena [Dog12, LX10, VC11, Gur09]. Prolog [CPZ89]. Prominence [Ano18-31]. promise [Mar96]. Promises [Ano88h, Ste86h]. Promising [OML^{+}07]. prone [Mat96f]. Propagate [Koo88]. Proper [Hec83b]. Properties [BMR^{+}06, CM04, SHKS19, WGO^{+}14]. Property [Ste93f, Ano98z, Dav93, Rob00d, Ste94f]. Prophet [FSR^{+}05]. Prophet/Critic [FSR^{+}05]. Proponents [Pit96a]. Proportionality [WA13]. Proposal [Ano03e, Ste83a]. Proposals [Ano22b]. Proposed [Ano84, Ano98x, CCG^{+}84, Ano81, Ano83, Ano00g, Bal84a, BT84, ES84, JC84, Reg92, Tau84]. Proprietary [HCP^{+}16, Ste85d]. Prosecuting [Emm06d]. Prosecution [Yi22a, Yi23a, Yi23b]. Prospects [TS14, WCH94]. protect [Ste94f]. Protected [Ste86a]. Protecting [SWL11, Sla90c, Ste86b, Ste89f, Ste93f, Ste93c]. Protection [Hau88b, Kar88a, Mat83, PZL06, Ste83a, Ste84b, Ste85e, Ste88c, Ste89a, Ste07e, SMS13, VMW^{+}19, WNW^{+}16, YMC^{+}12, YYK^{+}20, OFG88, Ste89c, Ste89d, Ste89e, Ste90e, Ste96a]. Protection-Domain [WNW^{+}16]. Protects [Ano87b, Ano99w]. Protein [Ano02d].

Protocol [Bea90, BLW02, FPAF02, KSB21, SABS20, SB00, BT84]. Protocols [CMC98, HBCS04, ONS^{+}23, Sha22, SLB04a, SLB04b, ZBES15]. Prototype [SCH^{+}23, TNT06, LGJ95]. Prototypes [Ano97z]. Prototyping [Ham00, OML^{+}07, TGC^{+}20, ME95]. Provable [WGO^{+}14]. proven [Mat02b]. provider [Ste96f]. provides [Ano96a]. Providing [WWR97, Wil95b]. Provisioning [NMF^{+}23]. Provocative [Ano99w]. Pruning [LK10, RAG19]. PS [Ano88c]. PS/2 [Ano88c]. PTO [Ste95d]. Public [AKP96, BEL^{+}23, ESG^{+}05, Fan96, Gre13c, Gre14e, KMK^{+}19, DVQ96, Gar93]. Public-Key [AKP96, ESG^{+}05, Fan96, DVQ96]. Publication [GPSS83, Hec83b, Ano16-35]. Publications [Ano23-76, Ano24y]. Publish [Smo86a]. Published [Ano20-34, Ano20-32, Ano20-33, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano23-34, Ano24m]. Publisher [Ano96j, Ano96k, Ano99-33]. Publisher-2000 [Ano99-33]. Publishing [Ano14p, Ano94b, Mat95d, Mat96c]. PULPv2 [RPL^{+}17]. Pulse [HK82, Mur89, MCH^{+}94, WM85, SK97, TTF96]. Pulse-Height [HK82]. Pulse-Width [WM85]. Punch [KFF00]. PurpleDrop [SWM^{+}20]. Purpose [DWLN20, ESG^{+}05, EKM^{+}95, ESCB13, Gil82, LLT^{+}08, MD20, STS^{+}92, TKM^{+}02, WLF^{+}08, Bos04e, Han96, SU95]. Purpose-Built [MD20]. Push [LNV82]. Push-over [LNV82]. Putting [AFGM10, Dia99]. PVCoherence [ZBES15]. PWM [TTF96]. PWRficient [Yeh07]. Pygmalion [Ang90]. PyMTL3 [JPOB20]. PyRTL [DTS20]. Python [ACZ^{+}22, JPOB20].

Q [CEH^{+}12, HOF^{+}12]. Q100 [WLP^{+}15].

QoS [CRV^{+}04]. QsNet [BAH^{+}05]. Quadrics [PcFH^{+}02]. Qualcomm [Ste06b, Ste17c, Ste17a, Ste17b, Ste18]. Quality [Ano23-83, Ano23-84, DK14, Dia92, FJB^{+}22, GCL^{+}20, Kah90b]. Quality-of-Service-Aware [DK14]. Quanta [Ste08b]. Quantitative [DMWS13]. Quantization [BFZ^{+}22, EPM^{+}20]. Quantized [CNC^{+}16]. Quantum [Ano20-57, Ano20-54, Ano20-55, Ano20-56, Ano22-64, Ano23-60, Ano23-61, Ano23-62, Ano23-63, BC21, BDSC21, BSA21, FRB^{+}18, GAT^{+}22, GBD^{+}20, HML^{+}21, Kar21, Kur21c, LP21, Mil89, MLM^{+}20, NM22, QT21, RBB21, RAA^{+}21, SVC01, TM21, Ano02d, Eng00j]. quarter [Ano03e]. Qubit [SVC01]. Qubits [BDSC21]. queries [FBGB96]. Questions [Gre16d, Ste85a, Ste03a]. Queue [ACG03, Kah93i, SMR07, SKJ^{+}11]. Queued [GSP02, PKP15, SGP02]. Queues [MC95]. Queuing [MW19]. Queuing-First [MW19]. Quick [Ano97z]. Quick-Turn [Ano97z]. quickly [MKRC97]. Quill [MF85]. Qutrits [GBD^{+}20]. QVGA [KII09]. QVGA-Size [KII09].

R [BKP12, Eec15f, Mat13b, FMN^{+}13]. R&D [Ano99u, VM88]. R-Mobile [FMN^{+}13]. R. [Luu90a]. R10000 [Yea96]. R3010 [RJR88]. R3TOS [IEB^{+}14]. R3TOS-Based [IEB^{+}14]. R4000 [MWV92]. Race [MSS15, XBH07]. racetrack [OZT^{+}22]. racing [OZT^{+}22]. Radiation [KLD^{+}94]. Radio [AAG^{+}10, BDV^{+}08, Ebe03, LLW^{+}07, MMB12, Sak01f, SBE01, SYY^{+}11]. RadioML [JUP^{+}22]. Radios [CN13]. Radix [CCG^{+}84, PKP15, Ste84e]. Radix- [CCG^{+}84]. Radix-Independent [Ste84e]. Rainbow [Ano22-71, Ano22-72, Ano22-73, Ano23-70, Ano23-71, Ano23-72, Ano23-73, Ano23-74, Ano23-75]. Raising [Gal97, Ste89d]. Rajwar [KT14]. RAM [KMD^{+}13, PAC^{+}97, XPZ^{+}19].

Ramakrishna [Bel12, Bel13]. Ramakrishnan [Ano16a]. Rambus [Cri97, Ste02b, Ste03b, Ste07b, Ste09c, Ste09d]. RAMP [WPO^{+}07]. RAMs [GXMZ13, JKP89, Nic88]. RAND [Ste07a, Ste08a, Ste15a, Ste15b]. random [KHF86]. Randomized [SGP02]. Ranganathan [Sco14]. Range [GKA^{+}16, Gre12f, RDJ^{+}13]. RAP [Dia95c]. Rapid [Ham00, WDK^{+}20, VCS^{+}19]. Rapidly [Mye93b, Gon97]. RAS [SLSO14]. RASSA [KYG19]. Rate [Gaf91, TK21, WEMR04, XWZ09, ZLTW13, Reg92]. ratios [AAW^{+}96]. Rau [Ano16a, Bel12, Bel13, Ano03f]. Ravi [KT14]. Raw [TKM^{+}02]. Ray [Ano88g, Ano97-33]. Razor [EDL^{+}04]. RDMA [SLZ23]. Re [RC12]. reach [Dia00, MKRC97]. React [CGLES^{+}23]. Reactive [CWB94, HFFA10]. Read [KYG19, Ano94c]. Reader [Ano85, Ano86b, Eec16d, Mat93f, Ste98a]. Readers [Ste85a]. Reading [Mat01b, Ano99w, Mat95b]. Readout [HC84, MA94]. Ready [Sti11, Ano03d, Dia96d]. Real [AT09, BTK^{+}23, Bos06c, CR95a, CR95b, CWB94, CFO^{+}18, Cle03, Cro85, DLR02, Dea04, EPZ02, FBC87, FOP^{+}19, HABHW^{+}18, Hum84, JW99, KE89, Kah92f, KKL^{+}09, KDK^{+}89, LPL86, ML05, MAS^{+}05, Mat97e, MBP^{+}85, MLM^{+}20, OKH^{+}12, PP92, RCR04, Rea86, RSE01, SK02, SRL91, SUF^{+}12, TS91, TGE95, ULS^{+}00, UCS^{+}10, Dur96, EKM^{+}95, Hea84, Hea87, RLG94, RH91, Yea96]. Real-System [MLM^{+}20]. Real-Time [AT09, BTK^{+}23, CR95a, CR95b, CWB94, Cro85, DLR02, Dea04, EPZ02, FBC87, FOP^{+}19, KKL^{+}09, KDK^{+}89, LPL86, ML05, MAS^{+}05, MBP^{+}85, OKH^{+}12, PP92, RCR04, Rea86, RSE01, SK02, SRL91, SUF^{+}12, TS91, TGE95, UCS^{+}10, KE89, Hea84, Hea87, RLG94, RH91]. Real-World [Cle03, Dur96, RH91, Yea96]. Reality [GMM^{+}07, HDG^{+}22, Kah93h, KKP^{+}14, LHC^{+}20, PSL^{+}23, SsSMB24]. Realization [IKNS88]. Realizing [KSWM90, War90d]. Really [Pal82, Ste91g, Ste96e]. rear [Ano99y]. Reason [Mil88c]. Reasonable [Ste17c]. Reasoning [NHMM23]. Rebuttal [Smo87d]. Receiver [PDT98, SZP81]. Receives [Bel12, Bel13, LE18, Ano01d]. Receiving [Hil19]. recessions [Gre01f]. Recipient [Goo14, Wei17]. Recognition [Ano15-36, Ano16p, BCKY17, HA96, HHNK09, IPL^{+}23, IST^{+}11, KKL^{+}09, KIR19, OKH^{+}12, TUI^{+}01, DO84, RLG94]. Recognizing [Alt14e]. Recommendation [DPT^{+}21, DtEt22, KZS^{+}22]. Reconfigurability [SKM^{+}16]. Reconfigurable [AHK^{+}14, AKAK^{+}18, Alt14e, And14, BLW02, BJ14, FGC^{+}14, GFL^{+}17, GDN^{+}17, GALB07, NI14, OYS^{+}11, PZK^{+}18, PCC^{+}15, SL03, SK97, SMT^{+}14, SYY^{+}11, TS14, WS13, WA11, GP95, OTM82, PHC95]. Reconfigurable-Computing [SMT^{+}14]. Reconfiguration [CS14, PC01]. Reconfiguring [CFZ^{+}99, DGW^{+}94]. ReconOS [AHK^{+}14]. record [Wha97]. recorded [AAW^{+}96]. Recorder [XBH07]. Recording [NPC06]. Recovery [ARS03, Ano01a, GSVP03, PV01, PDT98, RCA07, Ste09b, WN94]. Recurring [RGH^{+}10]. recycles [Dia98]. Red [YT01]. Redefining [ANM^{+}12]. Redesigning [Swa19]. Reduce [HCP^{+}03, MMESG^{+}20, ZZ05, AO97, Ano02c]. Reduced [Sch84, WRA^{+}14, MM87]. Reducing [ERM08, MMESGQ22, Rit97, RC13, Seg97, Wal97, WEMR04, GGJ^{+}96, Han96]. Reduction [AMR^{+}06, CBJ10, GGJ^{+}96, Kid14, LPKP22, SZZ01, VE10]. Redundancy [NBM^{+}06, SC24]. Redundant [TT12]. Reengineering [Dia93f]. Referee [CHA^{+}85a, Kar85]. Reference [Fla99]. References [Yi22d].

Refining [Pap96]. Reflections [Goo14, Hil19, Ste88d]. Reform [Ste09b]. Refresh [ERM08, NLM^{+}19, SWL11]. Refueling [AVU^{+}08]. refusals [Ste00b, Ste00c, Ste00a]. Reg [MM23]. Reg-TuneV2 [MM23]. Regime [Tay13]. Region [CSL^{+}06]. RegionScout [CSL^{+}06]. Register [RS93, Sim00, Fur88]. Registration [Lin92, Rob99c]. Regression [LB07, MM23, WL92, ZUNN18]. Regression-Based [MM23]. Regular [Rag84, Kra96]. Reimagining [NMU^{+}15]. Reinforcement [EPM^{+}20, NHMM23, SNM^{+}22]. Reinforcer [NBM^{+}06]. Reintegrate [KJL16]. Reinventing [Emm07c, Par00]. relate [WHKM93a, WHKM93b]. Related [AKT^{+}18, Ste08d, Ste08e, Gus92, Ste00b, Ste00c, Ste00a]. relates [Dan96]. Relational [AS91a, MG89, Mye84a, Ano97r, ISH^{+}91]. Relationship [Yi23a, Yi23b]. release [Ano94b, Ano03e]. Releases [Eng00i]. ReLeQ [EPM^{+}20]. Reliability [Alt13f, BJW^{+}23, BTR02, BDJS07, CPS^{+}18, Con03, Das21, GMM^{+}07, INKM05, LDF^{+}13, LLSS05, Qua00, Red13, SABR05, YE11, ZRA^{+}17, JKN96, Wil84, ZP93, AS05]. Reliability-Aware [Red13, AS05]. Reliability/The [ZRA^{+}17]. Reliable [BR21, Bor05, GKS^{+}05, Hor95, MLS^{+}16, MKAC18, NRS^{+}08, PV98, RG03, SBG97, WRA^{+}14, Bos06a, KWM89]. Relying [Sak99e]. Relyzer [HANR13]. ReMAP [WA11]. Remembering [Alt11c]. Remembrance [Chu18]. Remote [Gre21d, WPH^{+}23, AGH^{+}91]. Renaissance [KB20]. Renaming [Sim00]. render [Ano02b]. Rendering [HRK^{+}24]. Renesas [Van21]. Renewable [GKL^{+}14]. Rental [Pit91, Ste91e]. Rentals [Ste91b]. Reorder [ARS03]. Reordering [KCAR18]. Reorganization [AFH16]. Repairing [BCP01]. Repetitive [Gre96d]. Replace [Ger19]. Replacing [LCWB08]. Replay [NPC06, SYG^{+}20, XBH07]. Replays [Bha18]. Replica [CK98]. Reply [And82b, Ano91a, Ano00n, Dai94, Fai82a, Joh90a, Kar85, Kir83a, Kir84a, Kir84b, Mac84, Mat89a, Pit96b, RFGM86, Smi85, Smi86b, Ste88c, Ste91e, Uss91, ZVHL85, ZVH85]. Report [All81, Bal84b, BM19, Jef84, Kah90c, Kah91e, Kah91d, Kah92f, Kah93f, Kah93h, Kir85a, Kah93d, Kah93g, Far88b]. Reported [Mye84b]. Representation [NM22]. Representative [JC08b]. Reprinted [Jef84]. Requests [LLL^{+}16]. Requirement [Ste08e]. Requirements [BFK^{+}85, BSB^{+}92, PGL97]. ReRAM [JWS^{+}19]. ReRAM-Based [JWS^{+}19]. Research [Alb10a, And14, Ano88g, Ano99o, Ano19-38, ADC00, Eec16b, Hil19, HDG^{+}22, Kah92c, Kah92d, Kah93e, KB13, Kir89b, KZ13, ODH^{+}07, Shl93, Smi17, WPO^{+}07, Yi23c, ZACM14, Ano01e, Bos04d]. Research-and-Development [Kah92d]. Researchers [Ano02d]. Reshaping [Che19]. Residue [MHP^{+}23]. resigns [Ano03d]. Resilience [KGDW^{+}13, SS16]. Resiliency [GCL^{+}20, HANR13]. Resilient [CCA^{+}19, MSY^{+}22, PKL13, SKM^{+}16]. resist [Ano96n]. Resistance [Mat17, Soo93]. Resistant [VCK^{+}13]. Resistive [GGB^{+}15, KYGW17, KYG19]. Resolution [Ano97-33, PVYU94]. resolve [Ano03e]. Resonance [LDL17, PDL08, WLD15]. Resonant [ZLTW13]. Resource [BMK^{+}21a, BMK^{+}21b, BBE^{+}11, Mİ09, MMB12, NMF^{+}23, NMC^{+}08, SNM^{+}22, Sla96, SRA^{+}04]. Resource-Constrained [SNM^{+}22]. Resource-Efficient [SRA^{+}04]. Resource-Optimized [BMK^{+}21a, BMK^{+}21b]. Resources [Ano16x, MMESG^{+}20]. Responds [Ste98a]. Response [Eec15b, Ste86g]. Responsive [SUF^{+}12]. Responsiveness [RLC^{+}13]. Restraints [Ste98f, Gre05a]. Restricted [Ste91b]. restrictions [Ste97f]. Results [And82b, OML^{+}07, Tea82]. Resurfaces [Ano99i]. retarded [HP85]. Retargetable [NM22]. Rethinking [AS10, ERM08, JKN96, SAW^{+}10]. retina [Boa96]. Retinomorphic [Boa96]. Retires [Ano96j, Ano96k]. Retrospective [mHP18, SRU^{+}23, Vic93]. Return [Gre98c, War91a]. Reunifying [Kir90c]. Reusable [Fly97, JLWL20, WDK^{+}20]. Reuse [KCS^{+}20, MRJ^{+}15]. Rev [Ano88a]. Rev-1 [Ano88a]. Reverse [FGC^{+}14, Ste86c, Ste86d, Ste92e, Ste93g, Ano92e]. Review [Ano95c, Ano97n, Ano97o, Ano97p, Ano97r, Ano97q, Ano98z, Ano99x, Ano99y, CHA^{+}85a, Fla99, Hec83b, Kar85, Mat95b, Mat95c, Mat95d, Mat96a, Mat96c, Mat96e, Mat96b, Mat96d, Mat96f, Mat97c, Mat97d, Mat98b, Mat98c, Mat98d, Mat99b, Mat99a, Mat99c, Mat99d, Mat99e, Mat99f, Mat00a, Mat00b, Mat00c, Mat00d, Mat00e, Mat01a, Mat01b, Mat01c, Mat01d, Mat01e, Mat01f, Mat02a, Mat02b, Mat02d, Mat02c, Mat03a, Mat03b, Mat03c, Mat03e, Mat03d, Mat03f, Mat04a, Mat04b, Mat04c, Mat04e, Mat04d, Mat05a, Mat05b, Mat05d, Mat05c, Mat05e, Mat06d, Mat06a, Mat06c, Mat06b, Mat07a, Mat07b, Mat07c, Mat07d, Mat08b, Mat08a, Mat09a, Mat09b, Mat09d, Mat09c, Mat09e, Mat10b, Mat10c, Mat10d, Mat11a, Mat12a, Mat12b, Mat13a, Mat13b]. Review [Tab84, Yi22f, Yi22e, Gre15c]. reviewed [Mat13c, Mat14]. Reviewers [Ano12, Ano15a, Ano17a]. revisionism [Gre15c]. Revisited [Bor85a, Bro86, Emm06c, Ste87c, Ste11, Ste96d]. Revisiting [BVZ^{+}08, MHP^{+}23]. Revival [LWB09]. Revizor [OFKS23]. Revolution [DPY18, Gre00a, Gre09f, Joh22b, Sam00]. Rewriting [AS99]. RF [ASK^{+}15, Ano98-36, Ano01c, CN13, FME18, JUP^{+}22]. RF-Digital [CN13]. RFSoC [FME18]. RIBs [PKL13]. Richard [Ano14a, Ano15b, Ano16b, Ano17-29, Ano17b]. Riches [Eec16a]. Ride [Gre02e, NF81]. Riding [Dia95b]. Rigel [JJK^{+}11]. Right [Gre02d, SL97, Ano17f, Ano17e, Ano17c, Ano17d, Fly97, Mat00a, Moo04b, Ste97b]. Rights [Ste85b, Ste85d, Ste93f]. Rinda [ISH^{+}91]. Ring [LW94, JKN96]. Ring-Connected [LW94]. rips [Mat96f]. RISC [Kum97, Luu90a, OB91, AO97, AS22, ANUN98, AH96, Bur96, CGMV99, Col89, DXT^{+}18, DA92, DtEt22, DS95, Fur88, Gon99, Hen96, Hoo89c, HWG^{+}09, Hua89, Joh90b, Laz89, LWC^{+}16, Mel89, Mil88b, Mil88d, MBG^{+}16, NKDN95, NG87, PGW^{+}20, PW89, Pit95, Pit96a, Pit96b, Rob91, Sch96, Sla90e, SB23, SDC94, SANK98, TONH96, WE93, XYT^{+}23, ZSB21]. RISC-V [AS22, DXT^{+}18, DtEt22, LWC^{+}16, PGW^{+}20, SB23, XYT^{+}23, ZSB21]. RISC-V/Tensor [DtEt22]. RISC/DSP [DS95]. RISCs [Mye92c, PP82, Pen90]. RISCy [Smi92]. Risk [Ano16-41, CS18]. Risk-Based [Ano16-41]. Risky [Gre22a]. RMP [JKN96]. RNN [RAG19]. Road [FH05, Mat96b, Mye89b]. Roadside [Pal93]. Rob [Ano03f]. Robert [Ano99q]. Robohelp [Mat98e]. Robohelp-5.5 [Mat98e]. Robot [Ano88d, EEJ95, WM85, SM85]. Robotic [Mye81]. Robots [Ano00b, KSLY17, Eng00l]. Robust [Bos06d, EPZ02, WKK^{+}14, ZRB^{+}22, Bos05a]. Rock [Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano15-31, Ano15-29, Ano15-30, Ano15-32, Ano15-37, Ano16-36, Ano16-37, Ano16-38, Ano16-40, Ano16-39, Ano16-41, CCE^{+}09]. Rockstars [Ano15-33]. Role [ANM^{+}12, Hoo91, ZHR15, Dan96, Rob97b]. roles [Gar93]. Rollback [TNT06]. rollout [Ano03c]. ROM [STT^{+}15]. Rome [Mat21c]. ROMed [McG82, Pal82]. Ronnie [Gre23d]. Room [Gre06e, Ano99w]. Roomware [TSP02]. Roomware-Moving [TSP02]. Root [And82b, SL97, Tea82]. Rosetta [Gre16c]. rotating [Dv87]. Round [AML^{+}03]. Round-Trip [AML^{+}03]. Route [Trö98]. Router [PD01, SIPM02, WOM01]. Routers [WPM03, WH09]. Routing [CMB22, Den83, Liu02, Sav99a, SAA^{+}99, Tal93, Gal97]. Row [SsSMB24]. Row-Pipelined [SsSMB24]. Royalties [Ste15a, Ste07b]. Royalty [Ste17c]. RRAM [MSY^{+}22, THP^{+}19, YKH^{+}19]. RRAM- [YKH^{+}19]. RRAM-Based [MSY^{+}22]. RST [Pre91]. RTL [Bea20]. RTX [Bur20]. Rule [ACRV96, XLX^{+}23, SU95]. Rule-Driven [ACRV96]. Rules [Ste84a]. run [Yea96]. Runahead [MSWP03, MKP06, NAJE22]. Running [AHKY19, KFF00]. Runs [Gre23b, Mye83b, Ano03c]. Runtime [CK11, SHKS19, VCS^{+}19, CFM^{+}97]. Rush [Gre24a, Gre23a, Kir89c]. Russel [Gre15c]. Russell [Gre15c]. RV [KHF86].

S [Ano23-83, Ano23-84, Luu90a, Das21, RT92, SDF^{+}23, Kir84a, Pat84, SAC^{+}99, Sha22]. S-100 [Kir84a, Pat84]. S/390 [SAC^{+}99]. Sacrifices [Mye90]. Safe [BdS98]. Safety [AKT^{+}18, FPAF02, NMZ13, SNM^{+}13, SKA^{+}14a, ZRA^{+}17, vBK98, ZP93]. Safety-Critical [FPAF02, SKA^{+}14a, vBK98]. Safety-First [SNM^{+}13]. Safety-Related [AKT^{+}18]. Saga [Ste03b, Ste09c, Ste07b]. Sage [GLD^{+}22]. Said [Mye90]. sails [Gre04a]. Sakamura [Ano01d]. Sam [War90c]. same [Gre96e]. Sample [Jae82c]. Sample-and-holds [Jae82c]. Sampling [LB07, PBT06, VCE06, WWF^{+}06]. Samsung [Ano02c, RBGZ19]. Sandy [RNA^{+}12]. SANs [Ano99f]. SARC [KPK^{+}10, KK10, RCJ^{+}10]. Save [LDF^{+}13, MMB^{+}08, RES^{+}13]. Saving [Bos04b]. say [Ano02d]. Says [Mye84d]. SBCs [Ano98-29]. Sbus [War91d]. SC-49 [Fan96]. Scalability [TCC^{+}00]. Scalable [ARS03, AFK^{+}19, BDH^{+}16, BCC^{+}02, BPUH06, CNC^{+}16, DWF^{+}21, For02, GLD^{+}22, GAR^{+}06, GQF^{+}06, GKS^{+}05, HWG^{+}09, KJL^{+}10, KL05, KOKA23, KP03, LSL^{+}15, MKM15, MRSV11, MKT^{+}13, SK12, SDB^{+}04, SABS20, SBB^{+}17, War90e, ZBES15, ACRV96, Gal97, Hsi91, Gus92, IHCE07]. Scale [Alt11f, BJG^{+}19, BR10, Bar21, BDJS07, BUMV95, CFO^{+}18, DPT^{+}21, Far85, FAK^{+}14, FOP^{+}19, Gre17e, GHLK^{+}12, HLZ^{+}16, HAC^{+}13, IST^{+}11, JL11, JGC^{+}11, KDH^{+}16, KMK^{+}19, KDSA09, KO05, KKSV10, Lau21, MTS^{+}12, MLM^{+}20, NAA^{+}20, OG24, PCC^{+}15, RSC^{+}22, RNN^{+}16, TSW^{+}23, VAFF^{+}10, VJFG17, WAA^{+}21, XLX^{+}23, ZIM^{+}07, AKK^{+}93, TS95]. Scale-Out [FAK^{+}14, GHLK^{+}12, KMK^{+}19, VAFF^{+}10, VJFG17]. Scaleout [Sha23b]. Scales [FJL^{+}13]. Scaling [BY17, Bor99b, Cho23, EBS^{+}12, FD04, FGC^{+}14, GKS21, HRSS11, KK10, MSA^{+}03, Mea96, MD20, MCV^{+}14, WA13, YAK18]. SCALPS [DVQ96]. scanner [Ano95b]. scanners [HP85]. Scanning [LLLL09, TS06]. Scavenging [SP01]. Scenarios [MLL^{+}18]. Scene [Kir88b, Sak90b]. Scenes [SGL93]. Scheduler [GSP02, GM99, KKP^{+}14, MAM^{+}06, ZBH^{+}00]. Schedulers [HL06]. Scheduling [AMK17, BSC08, CWB94, CD09, DK14, Gaf91, KPMHB11, LH12, MNU^{+}15, MM09, MMCH18, RSE01, RBB21, ROA13, SGP02, MIM^{+}97]. Scheme [AKJF22, ANC05, CL05, JKP89, Tau87]. Schemes [ZZY97]. Scholarship [Ano15-40, Ano17-29]. SCI [Ano91c, EKB^{+}96]. Science [Ano21a, Ano21b, Ano21c, Ano21w, Ano22c, Ano22d, Ano22e, Ano23j, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24d, Ano24a, Gre23d, Ano92c, Hin88]. Scientific [DGM^{+}11, IG15, Mye84c, WWZ^{+}08, Ipe19]. scientists [Ano94b]. Scorpio [Sel18]. Screen [Ste88a, Ste89a, Ste89c, Ste89d, Ste89e, Ste90e]. script [DO84]. Sculpture [Ano99h]. SDAARC [EKMW02]. SDOs [Rob00a, Rob01a]. Se [Ste84a]. Search [Ano14g, Ano14h, Ano15k, Ano15l, Ano16i, Ano16j, Ano16h, Ano16g, Ano17q, BDH03, FCY^{+}20, HIP^{+}22, KSLY17, Ste04a, Ste85h, HM93, Sak01d, Ste02a, Ste04b]. Searching [Gil96a, PS03, ISH^{+}91]. SeaStar [BPUH06]. Second [BCF^{+}95, FGG^{+}88, Has85, LLL^{+}16, Mye92c, SGC^{+}16, Dia96d, SLM^{+}97]. Second-Generation [FGG^{+}88, SGC^{+}16, Mye92c, Dia96d]. Second-sourcing [Has85]. Secret [Gre12e]. Secrets [TSFS21]. Section [SMQP10, Ano96a]. sector [Gar93]. Secure [Joh19e, KTC18, KTY24, LWML16, ST19, TLW^{+}10, WZL20, WHJ^{+}23, YHHF20, ZHZ^{+}19, DVQ96, BM19]. Secures [Ano99t]. Security [AKP96, Ano15-33, Ano15-29, Ano16-41, Ano19-27, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21l, Ano22n, Ano22o, Ano22p, Ano22-71, Ano22-72, Ano22-73, Ano23q, Ano23r, Ano23s, Ano23-70, Ano23-71, Ano23-72, Ano23-73, Ano23-74, Ano23-75, Ano24g, DK18, DMWS13, Eec16d, Gon97, GBW^{+}23, GSS^{+}07, HMR^{+}19, Joh23c, KTC18, MGP21, MGG^{+}19, Ond96, SWL11, Sti19, SHKS19, SMAS16, TUI^{+}01, TSS18, TVT19, TA16, TLM19, US23, Ven23, WGO^{+}14, Wil95a, WHP^{+}13, YBS17, ZL16, Ano99-27, Ano01c, Wil95b, Ano19j]. Security&Privacy [Ano24h]. Security-Aware [TSS18]. see [Rob00b]. Seek [Ano23-76, Ano24y, Mat04d]. seeks [Mat96f]. Seemingly [Cas95]. Sees [Ste07c]. Sega [HO99a]. Segregation [ANC05, LKM92]. Seismic [BBB^{+}21]. Selected [KB13, KZ13]. Selecting [PGL97, Sak99a]. selection [HC83a]. Selections [Eec17f]. Self [Ano96u, BCP01, GALB07, IO16, LHL09, RGR95, TSV^{+}20, YNS^{+}14]. Self-Destruct [Ano96u]. Self-Driving [TSV^{+}20]. Self-Learning [IO16]. Self-Organizing [RGR95]. Self-Reconfigurable [GALB07]. Self-Repairing [BCP01]. Self-Tuning [YNS^{+}14]. Selfish [Ano97t, Wil97]. selling [Ste96e]. Semantic [MCV^{+}14]. Semantics [PCW15]. Semaphore [Lun85]. Semicon [Ano99k]. Semiconductor [Ano99w, Kat97, Ste07d, TKI^{+}14, Ano00i, Ano01c, Ano03b, IWM89]. Semiconductors [Gre23d]. Semicustom [Ste86b, AJR86]. sending [Ste97a]. Sensing [PCDL10, SVA^{+}22, SMR20]. Sensitive [CFRM04, TVT19, Gol96]. Sensitivity [CL05]. Sensor [Ano97h, BCH^{+}23, EK16, FJB^{+}22, MSP^{+}19, SVA^{+}22, SO14, Ano02b]. Sensornet [HHNK09]. Sensors [IKK96, NRV^{+}06, SCA^{+}12, WKK^{+}14, WHP^{+}13, Ano02c]. Sensory [SJO01]. Sensory-Augmented [SJO01]. SEP [Ste17c]. Sequence [KYGW17, TZMVLN81]. sequences [Hal91]. Sequential [Aug12, BVZ^{+}08, CO03, GJLT12]. Serial [Dia96d, EWW^{+}19, KMD^{+}13, SB00, Dia95d]. SerialExpress [JGF98]. Series [GA21, VBB14]. Server [AK00, CNC^{+}16, DGMM00, DBDF97, GKS^{+}05, IST^{+}11, JMZ^{+}11, KSSF10, KKSV10, LLL^{+}16, LRC^{+}09, PKB^{+}15, SGG^{+}12, SSR21, TIT^{+}13, JRHM86]. Server-on-a-Chip [SGG^{+}12]. Server/Workstation [DGMM00]. Servers [BCC^{+}02, FRS^{+}09, Gad07, HFFA11, KMAC03, MAT^{+}18, RCC12, VJFG17, YMA^{+}13, GK97]. Service [Ano14a, Ano15b, Ano16b, Ano21-58, Ano21-59, Ano21-60, Ano24v, DK14, HRC^{+}23, Ano99w, WN94, Ano17b, Ano23-58]. Services [Eng00k, FSS^{+}16, GM21, KKSV10, LM16, MW19, PCC^{+}15, STM02, XCZ^{+}21, XLW^{+}12, Ano98-29, Ano19-29]. Serving [CFO^{+}18]. Session [Emm07e, Emm08a]. Set [Ano00m, AOYS95, Bre10, DGMM00, DS94, Eng00o, Fai82a, Fai82b, FBGB96, FH00, NMU^{+}15, NT89, PKR92, QJP^{+}08, Sch84, Smi82, Ste09a, UBH^{+}94, WRA^{+}14, Ano03c, Eng00l, FN86, Lee96, MM87, WHKM93b]. Set-Dueling-Controlled [QJP^{+}08]. Set-Top [Eng00o]. Sets [Cre82, HCP^{+}16, Ste87c, XLX^{+}23, TONH96]. Setters [Ste07a]. Setting [Ste94c, Ste03a, Ste13, Wha97, FS05, Gar93, Ste98e, Ste05d, Upd93]. severe [HC83a]. sexy [Ano96n]. Seymour [Ano17-45]. SGML [Ano97p]. SGX [VMW^{+}19]. SH [BHM^{+}00]. SH-5 [BHM^{+}00]. SH3 [HKY^{+}95]. SH4 [ANUN98]. shapes [CG95, Gre97f]. shaping [Mat95b]. Share [Ano19-38]. Shared [DLCO10, DVWW05, KHL^{+}16, KL05, KCKP14, MHW03, MM09, TS91, TM94b, TM94a]. Shared-Memory [DLCO10, DVWW05, KL05, MHW03, TS91, TM94b, TM94a]. Sharing [Ano87g, ZL15]. Shedding [YYH98]. Shelf [PH91]. Shepherd [Gre20b]. Sherwood [Mar17]. shielded [War91g]. Shift [BBB^{+}21]. Shifting [Bos04d, RS93]. Shipped [Ano99s]. Ships [Ano97u]. Shoe [SP01]. Shoe-Mounted [SP01]. Shooting [Gre96e]. Short [Kah93i, Ste94a]. Shortages [Gre21e]. shortening [Rit97]. Shortfalls [Gre01f]. Shot [ZRA^{+}20]. Should [EHP^{+}07, Ste84b, Ste96f, Ste98b, Ano94c, Gre96f, Mat95d]. Show [Mat04d]. Shrimp [DBDF97, DBC^{+}98]. Shrink [Ste97f]. Shrink-wrap [Ste97f]. Shuttle [Kir92]. SIA [Eng00n]. Side [CEAY23, CDBY23, DMWS13, KPN^{+}20, LWML16, WPH^{+}23]. Side-Channel [CDBY23, DMWS13, KPN^{+}20, LWML16, WPH^{+}23]. Sides [Gre17e]. Sidney [Ano17-45]. Siemens [Ano98-34]. Signal [Ano97h, AF84, CWL^{+}14, DM88b, DM88a, Eic86, Fra00, FGG^{+}88, Gas21, HSP^{+}01, KSA^{+}19, KW81, Klo86, KB91, KPHP04, LCS92, Mor86a, MD88, MBK^{+}92, NG87, PS88, PKR92, SP92, ST21, SK88, SJB09, WSM^{+}10, Ano92b, Ano95a, BTHS92, DFR90, FLRB86, RMFG85, Wv92]. Signal-Processing [AF84, DM88a, Mor86a, MD88, NG87, Wv92]. Signal-Switching [HSP^{+}01]. Signaling [DP97, Das21, HYS98, PDT98, Sha22]. signals [Ste98b]. Signature [Eng00d, LLLL09]. Signature-based [LLLL09]. Signatures [HA96, TATC09]. Signed [Gre22c]. significantly [TONH96]. signing [KAK96]. Silent [SFG^{+}22]. Silicon [ACZ^{+}22, Alt13b, Ano02b, BJO^{+}09, Bos06c, Cai89, CS13, EBS^{+}12, FD04, GHSV^{+}11, HFFA11, HAC^{+}13, KKS^{+}98, LWK94, OMMB13, PDS^{+}13, RES^{+}13, STT^{+}15, STR^{+}13, SKS^{+}13, TP10, TS13, Tay13, WKK^{+}14, Ano01h, Ano02c, Ano03b, DTH^{+}95, Pri94b, MC90]. Silicon-on-Thin-Buried-Oxide [STT^{+}15]. silicon/ferroelectric [DTH^{+}95]. Silk [Eng00a]. Silver [Ano02b, MF85]. SIMD [RPK00]. SimFlex [WWF^{+}06]. similar [Gre05f]. Simple [FHP00, MBS08, ZQL^{+}04, CG95, KSI^{+}96, Rob00c]. Simplify [Mar21]. Simplifying [HCW^{+}04, Wal97]. SimPoint [VCE06]. SiMul [LYP^{+}18]. Simulating [BO86, GGC^{+}11, LC91]. Simulation [ABG^{+}20, Bea20, BMK^{+}21a, BMK^{+}21b, Can98, CF90, DMP91, ENSD03, GKS06, Har12, HBE^{+}10, Ibb00, JPOB20, KMK^{+}19, KL08, LHM99, MBK^{+}92, OHLR94, RPE10, SY06, WWF^{+}06, ESW97, RS90, UBL^{+}82, vdDD90]. Simulations [AW03, Kha00, Pap96]. Simulator [BCU^{+}99, BDH^{+}06, LYBZ04]. Simulators [CDS07, NMHS15]. Simultaneous [EEL^{+}97, IGH^{+}99]. SimWattch [CDS07]. Singapore [Kah93b]. Single [AMK17, Ano97f, Ano98-36, Ano99-33, AMFFM^{+}16, CMAS11, EMYN00, EHP^{+}07, Eng00o, Gol96, JJK^{+}11, KPV^{+}99, KHS^{+}23, KCKP14, LBD^{+}99, LLL^{+}16, Mat04e, MMB^{+}08, Mye83c, NIJ^{+}03, SC91, ZRA^{+}20, Mon97]. Single-Chip [AMK17, CMAS11, EMYN00, Eng00o, Gol96, JJK^{+}11, KPV^{+}99, KHS^{+}23, LBD^{+}99, Mye83c, NIJ^{+}03, SC91, Mon97]. Single-Cycle [KCKP14]. Single-Electron [Ano97f]. Single-ISA [AMFFM^{+}16, MMB^{+}08]. Single-Shot [ZRA^{+}20]. Single-Sourcing [Ano99-33]. Single-Threaded [EHP^{+}07]. Single-Unit [Ano98-36]. Sips [Mat97e]. Sirius [HLZ^{+}16]. Situ [WKK^{+}14, ZFW^{+}23, PHC95]. Situational [AMK17]. Six [Gre19e]. Size [KII09, MCV^{+}14, Fur88, Pri94b]. Skiing [Rob99d]. Skills [Ano20y, Emm07d]. Skullduggery [Ste01b, Ste02b, Ste07c, Ste09c, Ste09d, Ste11, Ste12, Ste17c, Ste17a, Ste17b, Ste18, Ste01d, Ste05c, Ste07b]. Skunk [Gre16e]. Sky [GZC^{+}17]. Skylake [DKyL^{+}17]. Slack [DMMD11]. SLDRAM [GV97]. Slice [Gre22b]. Slicing [Ano87g]. Slickedit [Ano96t]. Slope [SKS^{+}13]. Slot [Hur98]. Slot-1 [Hur98]. Slotcars [McK83]. Slouching [Gre08b]. Slowing [Eec17c]. Slump [Sak01e]. Smaky [Kir89d]. Small [AT09, LLT^{+}08, Pap89, TUI^{+}01, TS95]. small-scale [TS95]. Smaller [Eng00p]. smallest [Ano02c]. Smart [Ano96q, Ano97-27, DF01, EMYN00, FJB^{+}22, HC84, Joh22d, KGT22, MK22, NM96, NFQ03, Sak01f, SMM^{+}22, SCA^{+}12, SF18, TBDL01, Tua99, DVQ96, KCKP14]. SmartNIC [DRM^{+}23]. Smartphone [ZES13]. Smell [Ste86f]. SMP [Cha98]. SMT [CRV^{+}04, KKS^{+}23]. soap [Gre95b]. SoC [DtEt22, Ano00g, CSV02, Sak02c, GMC18, GCE^{+}21, Lin04, PSB^{+}20, XCZ^{+}21]. Soccer [Gre09d]. Social [Ano18-32]. Society [Ano14o, Ano17-27, Ano17y, Ano17-29, Ano18t, Ano20-52, Ano21-53, Ano21-58, Ano21-54, Ano21-59, Ano21-55, Ano21-56, Ano21-57, Ano22-58, Ano22-59, Ano22-60, Ano22-61, Ano22-53, Ano22-62, Ano22-54, Ano22-63, Ano23-56, Ano23-57, Ano23-43, Ano23-44, Ano23-40, Ano23-46, Ano23-58, Ano24q, Ano24r, Ano24v, Ano24p, Mar96, Ano96c, Ano01d, Ano15u, Ano16x, Ano16s, Ano16v, Ano16t, Ano16w, Ano16u, Ano16r, Ano17z, Ano17w, Ano17x, Ano17-28, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18m, Ano18n, Ano19z, Ano19u, Ano19v, Ano19w, Ano19y, Ano19x, Ano20-48, Ano20-38, Ano20-39, Ano20-49, Ano20-44, Ano20-40, Ano20-50, Ano20-45, Ano20-41, Ano20-51, Ano20-46, Ano20-42, Ano20-47, Ano20-43, Ano21-41, Ano21-36, Ano21-46, Ano21-52, Ano21-47, Ano21-42, Ano21-37, Ano21-48, Ano21-43, Ano21-38]. Society [Ano21-49, Ano21-44, Ano21-60, Ano21-39, Ano21-50, Ano21-45, Ano21-51, Ano21-40, Ano22-47, Ano22-41, Ano22-48, Ano22-42, Ano22-49, Ano22-43, Ano22-55, Ano22-50, Ano22-44, Ano22-51, Ano22-56, Ano22-45, Ano22-52, Ano22-57, Ano22-46, Ano23-27, Ano23y, Ano23z, Ano23-47, Ano23-53, Ano23-48, Ano23-54, Ano23-49, Ano23-35, Ano23-39, Ano23-50, Ano23-36, Ano23-45, Ano23-38, Ano23-51, Ano23-37, Ano23-41, Ano23-52, Ano23-55, Ano23-42, Ano24s, Ano24t, Ano24o, Ano24n, Ano24u, Ano20-34, Ano20-32, Ano20-33, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano23-34, Ano23p]. Socket [Ano96m, Ano96s]. Sockets [FJL^{+}13, ZG96]. SoCs [ABG^{+}20, McD21, PGW^{+}20]. soda [MIM^{+}97, ACZ^{+}22, LLW^{+}07]. Soft [NRV^{+}06, SWK^{+}05, SGK^{+}04, SSC^{+}22, SMS13, WEMR04, CMR97]. Soft-Error [SMS13, WEMR04]. Soft-Error-Detection [SGK^{+}04]. SoftSig [TATC09]. Software [ABIV06, Alt12c, AAW^{+}96, And82a, Ano14-34, Ano15-34, Ano23-83, Ano23-84, Bea20, BTK^{+}23, BSY^{+}10, BMM15, BDV^{+}08, Bus86, BM85, CGJ^{+}94, CN13, De 94, Dem94, DF01, ECY^{+}12, GZC^{+}20, Gon06, Gre18b, GHY^{+}17, HCW^{+}04, Hea87, HKM^{+}85, HAB^{+}09, Joh90b, KW83, Kah90c, Kah91e, Kah91d, Kah93d, Kal97, KST12, KKS^{+}23, LPL86, LSY01, Lie23, LLW^{+}07, LLLL09, MAS^{+}05, Mat90a, Mat96d, Mat03c, Mat08a, Mat09c, Mat83, MCC^{+}07, MMB12, MCV^{+}19, Mor86b, NRS^{+}08, OHLR94, OFG88, PMR^{+}22, RCA07, RFGM86, RPE10, SG01a, SPRK04, SKW^{+}23, Ste83d, Ste83a, Ste84a, Ste84c, Ste85c, Ste86a, Ste86f, Ste86e, Ste87d, Ste87e, Ste89b, Ste90a, Ste90f, Ste91b, Ste91a, Ste08d, Ste08e, Ste14a, Ste14b, Str98, SBG97, SYY^{+}11, TKM^{+}02, TATC09, VRMC20, Wal97, ZQL^{+}04, Ano92b, Ano92e, Ano98-29, ACG^{+}88, CMR97, FL84]. software [Gre97d, HF81, KHW85, KHF86, Kah93a, KKT^{+}91, Pir97, SS82, Ste83c, Ste93e, Ste95d, Ste96d, Ano14q, Buc84, Pit91, Ste91e, VBC^{+}21]. Software-Configurable [Gon06]. Software-Defined [BDV^{+}08, CN13, KKS^{+}23, LLW^{+}07, MMB12, SYY^{+}11]. Software-Exposed [TATC09]. Software-Hardware [GHY^{+}17, OHLR94, Ste84a]. Software-Only [RCA07]. Soggy [Joh90b]. SOI [NFQ03]. Soil [GAGV22]. Solicited [Ano17-45]. Solid [Alb07d, Alt11e]. Solution [Del91c, DMG^{+}15, For02, HSR18, SLSO14, SABR05, TSV^{+}20, Bal84a]. Solutions [CD97a, CPS^{+}18, JP17, Won03, Ano99-27, LHN95]. Solvent [Ano98q]. Solving [BM85, GFL^{+}17, Hoo89a, Lyl04, VL00]. Some [Alt11f, Kir85a, Lei98]. Sometimes [SRJ^{+}91]. Sonic [SYW^{+}14]. soon [Mat06d, Pri94a]. Sorry [War91e]. Sorting [LHN95, PS03, ISH^{+}91]. SOS [BKM^{+}82]. Source [BC20, CCA^{+}19, DXT^{+}18, GCE^{+}21, JPOB20, MEB^{+}20, Pal82, PGW^{+}20, Ste85e, Ste06a, TGC^{+}20, WDK^{+}20, Ano02b, Pri94b, SL84a]. Sourcing [Ano99-33, Mat04e, Has85]. Sovereignty [BCN^{+}22]. SP [MKM15]. SP-CNN [MKM15]. Space [AF84, DGR^{+}10, Kir92, NBM^{+}06, RCR04, Sim00, Ano01f, IKK96, RLG94, WCH94].

Space-Based [AF84]. space-frequency [RLG94]. Space-Shuttle [Kir92]. Spacetime [Smi17]. Span [RD90]. Sparc [FJL^{+}13, CCE^{+}09, AJK^{+}15, BSC^{+}90, BAC^{+}90, DKB^{+}90, KAO05, SGG^{+}12]. SPARC64 [MAT^{+}18, MYK^{+}10, YMA^{+}13, YHT^{+}15]. Sparcle [AKK^{+}93]. Spare [PKL13]. Sparse [NY22, Ipe19]. Sparsity [HIP^{+}22]. Spatial [LB07, PPA^{+}14, SW14, STM02, WLKN22, DTH^{+}95]. Spatial-Information [STM02]. Speak [JN21b]. Speaking [Chr91]. Speaks [Ste15a, Mar98]. Spearmints [KKC93]. SPEC [Ano03b, HCPS03, Ano97-28]. Spec92 [GHPS93]. SPECfp [AAW^{+}96]. SpecHLS [GRD22]. Special [Ano97-29, Ano15-35, Ano15-36, Ano22b, AW22, BL23, Bon21, Cas15, Das22, Del92, DA23, EW23, EW24, FC22, GGS22, GT22, IA22, Joh22e, Kah92f, Kah93f, Kah93g, Kah93h, Kar21, KB13, KK23, Lee21, Lev23, LS24, MK22, RM23, Sak89, Sak91, SS22, SRL91, VBB14, Ven23, Wu23, Ano95a, Bor85b]. Specialization [GHN^{+}12, MCV^{+}19, NDR^{+}22, NGSW17, SV21]. Specialized [CLL^{+}20, FAK^{+}14, KSV^{+}21, ISH^{+}91]. Specializing [BJG^{+}19, KGMT17]. specially [KWGG95]. Specific [AK24, CYH^{+}18, JL87, Koe86, LBS^{+}11, Vei04, VAD^{+}21, Ano02c, Bos04e]. Specification [Das21, PVB^{+}20, SKO89, TS91]. Spectre [HMR^{+}19]. Spectrograph [Mor88]. spectroscopy [Ano01c]. Spectrum [Gre09f]. Speculation [NRS^{+}08, OFKS23, RSC^{+}06]. Speculative [GRD22, HBCS04, KCAR18, MT03, RNLY23, YYK^{+}20]. Speculatively [YYK^{+}20]. Speech [EI87, EIB90, Mor88, HP85]. Speed [Alt14d, BJ14, BWBJ11, DQCL24, Gal97, Gun06, HSP^{+}01, HYS98, JBM95, JL87, KPP06, KL05, LLLL09, LCY^{+}04, PMM15, PSP14, SLM^{+}97, TP10, TRY^{+}09, Ano01h, Ano02e, Ano03b, DP97, Dia96c, GP95, KAK96, MHW94, Mat93f]. speedAI240 [SB23]. Speeding [Ste89b]. SpeedLog [WN94]. Speeds [Ano88h, Ano96f, TONH96, FBGB96, SLM^{+}97]. Spent [Mat92a]. Sphere [Sti19]. Spider [Gal97]. Spiking [OSS^{+}24]. Spillovers [Gre11b]. SPIN [RGK19]. Spiritual [Ano94d]. Split [TS06]. Splitter [SL03]. Sponsors [Ano20x]. Spotting [RT23]. Spring [Mye82a]. springs [Joh90b]. Sprinting [RLC^{+}13, RES^{+}13]. spurious [Rob97e]. spyware [Ste05b]. sqrt [And82a]. Square [And82b, SL97, Tea82]. Square-Root-X [And82b, Tea82]. Squeaks [Lan85a]. Squeezed [NHMM23]. Squeezed-Edge [NHMM23]. SR [ZCW^{+}14]. SR-IOV [ZCW^{+}14]. SRAM [ASD^{+}05, SCA^{+}12, TKI^{+}14, YBNS15, ZSS^{+}19]. SRAMs [LCWB08]. SSBLT [Reg92]. SSDStreamer [BJG^{+}19]. SSI [Pee87]. Stack [ADF^{+}10, AH96, BJG^{+}19, Bea90, Mat91b, PZL06, RAA^{+}21, STR^{+}01, TES^{+}18, TML^{+}18, Gre01d]. Stack-Based [PZL06]. Stacked [DFG^{+}13, LX10, SLSO14, YRC^{+}22, Ano95b]. Stacking [HSX18, LXB07]. Stalls [NAA^{+}20, RTJ20, RTJ21]. Stand [GSS^{+}07]. Standard [Ano84, Ano88e, Ano96r, Ano02e, AMFFM^{+}16, Bal84c, CS13, CCG^{+}84, Cri97, Gar93, GV97, Jos86, KSM99, MRC^{+}20, MD20, Mye82b, Mye82c, Rob98e, Smo86a, Ste03a, Ste13, Tho92, War90f, War91c, War91f, Ano81, Ano83, Ano99w, Ano00i, BC86, Dia94b, Dia95d, ES84, Fis85, FS05, GK97, JC84, Mar85, Pri94a, RT86, Reg92, Ste98e, Ste99d, Ste05d, Ste05c, TZMVLN81, Upd93, Ano97d, Ste07a, Ste08a]. Standard-Setting [Ste13, FS05, Ste05d, Upd93]. Standardization [Ano96v, Car98, Gre10e, STL92, Ste01b, Ste02b, Ste02d, Ste05c, Ste07c, Ste09d, Ste11, Ste12, Ste17c, Ste17a, Ste17b, Ste18, Dav93, Dia96d, Ste01d, Ste07b]. Standards [All86b, Ano97s, Ano98x, Ano15f, Bor85a, Bor81, BS84, Buc84, Buc87, Dia92, Dia93f, Dia93c, Dia93d, Dia93e, Dia95d, Dia95e, Dia96d, Dia96c, Gre10d, Gro83, Hec83a, Hec83b, HAB^{+}09, IJ98, Kal93, Lei98, Mye84d, RSW10, Rob97a, Rob97b, Rob97c, Rob97e, Rob97d, Rob98d, Rob98b, Rob98e, Rob98c, Rob99b, Rob99a, Rob99c, Rob99e, Rob99d, Rob99f, Rob00a, Rob00e, Rob00b, Rob00c, Rob00d, Rob01a, Rob01b, Rob01d, Rob01c, Smo87a, Smo88a, Ste94c, Ste08c, Ste15a, War89b, War92a, BCF^{+}92, Eng00j, Gre93, Gre15c, Gus92, Hal93, Kir01, Smo87c, Ste99a, Ste99b, Ste00c, Ste01e, Vic93]. Standing [Alb07d]. Stanford [CFK^{+}10, HHS^{+}00]. Starfire [Cha98]. Stars [Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano15-31, Ano15-29, Ano15-30, Ano15-32, Ano15-37, Ano16-36, Ano16-37, Ano16-38, Ano16-40, Ano16-39, Ano16-41]. Start [KLM^{+}15, ADC00]. Start-up [KLM^{+}15]. StarT-Voyager [ADC00]. Starting [Rob98e, TM82]. starts [Mat96f]. Startup [Ano15-37, VCE06]. Stat [Jef84]. State [Eec15e, LL03]. States [CHA^{+}85a, Kar85, LDL17, ZHPR17, Gar93, Ste91b, Ste92a, Zsc84]. Static [GXMZ13]. Statistical [ENSD03, WWF^{+}06]. Statistics [SIPM02]. Status [All81, All84, Bal84b, Kni85]. Status-Report [All81, Bal84b]. Stay [Ano15-38, Ano19-39, Rob01c]. STC [Ano14r, Ano15-39]. Std [Dia94b, Dia95d, Dia96d]. STEAM [GKS06]. Steep [SKS^{+}13]. Steep-Slope [SKS^{+}13]. Stepping [Sak00f]. Steps [Ano96l]. Steve [Ano01d, Gre11f]. Stick [Ara00]. Sticking [Ste95c]. Still [Kaw98, Kir91c, Pre21, Alb07e, Rob00a]. Stimulus [Gre09b]. Stochastic [NJZL^{+}17]. Stone [Gre16c]. stop [SS82]. Storage [BLC^{+}17, Dav02, DPBW19, GKS06, Gur09, GSS09, HLIT20, KYGW17, LLZ^{+}04, RCBL00, SFG^{+}22, Sto94, SF95, Ano01h, Ano02b]. Store [GAR^{+}06, KCAR18, SMR07]. Store-Load [SMR07]. Storing [BK14]. Story [Gol21, Kir89d, Lau21, Van21, BC86, Eng00g, FHMS96]. straight [Wha97]. Strained [Ano01h]. Strategies [Ano16-48, Ano16-47, Ano16-46, KMG^{+}03, LB07, PEZ^{+}19, SG01a, Ano16-45, CR95b, Emm06b, LNV82]. Strategy [Ano98x, Gre98e, Lun85, MK10, Gre99c]. Stream [MCH^{+}94, NDR^{+}22, RCR04, WWZ^{+}08, ZG96, SK97]. Stream-Based [NDR^{+}22]. Streaming [JUP^{+}22, RPK00]. Streams [KDK^{+}01, PVB^{+}20]. stress [Gre96d]. Stressmark [KJP^{+}13]. stretch [Ste07b]. Stretches [Mor86b, RFGM86]. String [ATS^{+}22, TS06]. Stripes [Gre19c]. Strong [SLSO14]. StrongARMing [LS98a]. Stronger [CCS21]. Structure [DWLN20, Eec15f, FMV85, Gre13a, Nic88, SHS85, Boa96, HF81, MKNK83]. Structured [AJR86, Man86b]. Structures [Bor81, PVB^{+}20, CDGO97]. STT [YYK^{+}20, XPZ^{+}19]. STT-RAM-Based [XPZ^{+}19]. Student [Ano15-40, Ano17-29, Ano17-58, Ano17-59, Ano18-33]. Students [Cle03, LMC^{+}83]. studies [CPZ89, RH91]. Study [HGS^{+}17, KGDW^{+}13, MLM^{+}20, SWK^{+}05, Sen86, Smi86a, Smi86b, Smo88c, SJB09, SZH82, BSB^{+}92, Gre96a, OL85, SZP81]. Study-Groups [Smo88c]. Studying [Ano97-30, Jac03]. Stuff [Alb07b, BS98]. Subject [Ano97a, Ano98a, Ano00a, Ano01b, Ano94c, Ano96a]. Submicrometer [YCD^{+}19]. Submicron [Ano97j, FHR99]. submissions [Ano98c]. Subscribe [Ano21-68]. Subsetting [JC08a]. Substitution [LHC^{+}12]. Substrate [Car93, VBC^{+}21]. Substrates [Hol98, Bel93]. Subsystem [CKD^{+}10, Pri86, WHKM93b]. Subsystems [WH09]. Subthreshold [CB10]. subtractive [BG81]. Subword [Lee96]. Success [Joh22b, LCS92, Ste85g, Joh90b]. Successful [GS99, Lau21]. Sue [Ste17c, Ste17a, Ste17b, Ste18]. Sues [Ste08c]. Suit [Ste85e, Ste06a]. Suite [GHPS93, MRC^{+}20, Ano03b, PCLGO09]. SUME [ZACM14]. Summary [Ano97x, Ano97y, Ano98-39, Ano98-40, Ano98-41, Ano98-42, Ano98-43, Ano99-29, Ano99-30, Ano99-31, Ano99-32, Ano00j, Ano00k, Ano00l, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano02f, Ano02g]. Summer [Mat00d, Ano97o]. Summit [Ano15-34]. Sun [Ano03d, Cha02, FRS^{+}09]. Supercharging [Emm07d]. Supercomputer [DM88b, GGC^{+}11, HMS^{+}86, Kir89b, MD88, MAT85, MBK^{+}92, SKTO22, Ano00g]. Supercomputers [AK24]. Supercomputing [EVM^{+}98, Kah93g, Ano02b]. Superconducting [FRB^{+}18]. Superconductor [IBN^{+}21]. Superconductors [TVV^{+}21]. SuperEnc [IKN^{+}99]. SuperH [BHM^{+}00]. Superhighway [Ste94c]. Superscalar [CWS^{+}12, CEM^{+}95, ERPR95, Sim97, Sla89, SANK98, DA92, UAN^{+}93, Yea96]. Supersmart [Mye89b]. Supplemental [TBDL01]. suppliers [Ano02c]. Supply [ABIV06, MGP21]. Support [Ano97-31, Ano99n, BFK^{+}85, BB17, HKM^{+}85, INKM05, KSWM90, KHHR85, Kni85, KGS^{+}19, MBP^{+}85, MCF^{+}85, MKOK88, PP82, Pir97, Ste98a, TA16, ZUNN18, ZSS^{+}19, ZQL^{+}04, Ano99p, KKC93]. Support/Privacy [Ano99n, Ano99p]. supporter [Mar98]. Supporting [AML^{+}03, BMS16, CR95b, Fly97, LS22, LH12, Mon97, UCS^{+}10, Kai88, Lee96, TO96, WN94]. supports [Dia95d]. Supreme [Ste08b, Ste07d]. Sure [KY91]. surface [GGJ^{+}96]. Surprises [Lei98]. surprising [Pri94b]. Survey [Ano85, Ano86b, Eec16d, Fet95, FSH^{+}01, GAGV22, Gro92a, Kal93, YBS17]. survive [Ano97q]. Survives [Hoo90c]. Surviving [LDCS09, Sak01e]. Susan [LE18]. Sustainable [Ano21n, Ano21o, Ano21p, CHAF22, Eec23, GEH^{+}23, Joh23b, OLT^{+}23, Wu23, Ano19-28]. SVC [HSR18]. SWAP [LHC^{+}12]. Swarm [JSY^{+}16]. Sweat [Ste94d]. SWICH [TNT06]. Swiss [Kir89d]. Switch [AML^{+}03, ACKM05, Cum04, Edd02, GKS21, KPV^{+}99, MIM^{+}97, STT^{+}15, SGP02, Yun01, ZBH^{+}00, ZLBI06]. Switched [YTR^{+}98]. Switches [GSP02, PKP15]. Switching [DMG^{+}15, HSP^{+}01, KSI^{+}96, KM05, MFM02, ZCW^{+}14]. SymbiFlow [MEB^{+}20]. Symbiosis [DF01]. symbols [Lan87]. Symmetric [KO05]. Symposium [Bro17, Gon18, HW91, KT14, Mar14, SRU^{+}23, Tor12, Ste90g, Ste90h]. Symptoms [Gre09e]. Synapses [YCD^{+}19]. synaptic [RJHK89]. Synchronization [ADJK20, But07, KPK^{+}10, MT03, OL85]. Synchronized [RGK19]. Synchronous [CB04, Lin04]. SyncLink [Dia96c]. Synergistic [ASD^{+}05, GHF^{+}06, TCD^{+}05]. Syntactic [SWM87]. Syntax [SHS85]. Synthesis [Ano18d, CFRM04, CS14, EI87, GRD22, KCXmWH17, KIS^{+}00, Lan96, ONS^{+}23, PVS^{+}11, TCC^{+}00, TLM19, TMA18, BG81, Wv92]. Synthesizable [RHH^{+}03]. Synthesizing [ATS^{+}22]. synthetic [MC87]. System [AHK^{+}14, ABG^{+}16, AB06, Ano98-28, Ano99v, Ano01h, AF84, BdS98, Bel96, BFK^{+}85, BGS89, Boa96, BCKY17, BLC^{+}17, Bos03a, BM19, BTR02, BCF^{+}14, BWBJ11, CR95a, CO03, CDS07, CFRM04, Cla03, CL87, CES^{+}11, DRM^{+}23, Dav98, DFG^{+}13, EI87, EE08, FBC87, FKL01, Fos98, FCY^{+}20, GBW^{+}23, GA21, GR92, GGJ^{+}96, GD01, HKM^{+}85, Hor95, IN87, IKK96, Jac03, Jae83, Joh23a, Joh87, KY91, KSWM90, KMK^{+}19, KK23, Kir91a, KGDW^{+}13, Kni85, KL08, Koe86, KKS^{+}98, KAV99, KR19b, LHM99, LP89, ML05, MA94, MBP^{+}85, MCF^{+}85, MLM^{+}20, NCT^{+}98, NL02, OHLR94, OKN^{+}11, PLK^{+}16, PZB^{+}19, PLBC09, PRE11, Pre91, RRP^{+}08, Rea86, RNN^{+}16, RPE10, Sak87c, SK01, SV03, SKTO22, SML04, SO14, Sha23c, Sla90b, Ste83a, Ste84d, Ste91c, SL84b, STS^{+}92, SsSMB24, Trö98, TGE95, VM95]. System [WM85, Wal97, WKK^{+}14, WNW^{+}16, WMSH09, WWZ^{+}08, WWF^{+}06, Yao85, YZW^{+}23, Zha91b, CCD^{+}82, CH94, CDGO97, DKM^{+}92, ES84, Han96, HP85, HS85, Hea84, Joh90b, KKT^{+}91, Mon87, PGL97, Rit97, RH91, SSH88, Seg97, SM85, Ste93e, TGF88, WJR88, Ber86, HLHR90, MSB87, Mat09d, OB91, PJ91, SB84]. System-in-package [Ano01h]. System-Level [Bos03a, EE08, PLBC09, Seg97]. System-on-a-Chip [Bel96, Ano99v]. System-on-Chip [ABG^{+}16]. System-on-Chip-Level [SsSMB24]. System-on-Silicon [KKS^{+}98]. System/6000 [OB91]. Systematically [TGE95]. Systems [AKP96, AAG^{+}10, Alt11f, And14, AT09, Ano87a, Ano98-44, Ano02e, ABC99, AS99, AGJL98, ALGJ01, BBB^{+}21, BCP04, BPT^{+}11, Ber09, BMG^{+}21, Bha20, BMK^{+}21a, BMK^{+}21b, BBE^{+}11, BDH^{+}06, BDH^{+}16, BFLS01, Bor05, Cas15, Cas95, CRV^{+}04, Che19, CK98, CR95b, CGJ^{+}94, CLM08, CWB94, CML^{+}23, CS81, Cle03, CHSL17, CP86, CMAS11, Cum04, DKB^{+}90, DWF^{+}21, Dra00, DM88a, Ebe03, Eec23, ECK^{+}22, FK83, FPAF02, Fet95, FSH^{+}01, GZC^{+}20, GALB07, GR95a, Gro94a, GGB^{+}15, GEH^{+}23, GKS06, GSS09, Her00, HSW98, HAC^{+}13, HL86, HcF04, HDG^{+}22, IEB^{+}14, Jag97, JL11, Joh19a, Joh19b, Joh84, KND02, KG05, KDSA09, KLM^{+}15, KKS^{+}23, Kir90e, KBH^{+}08, KHHR85, KL08, KDK^{+}89, KO05, KP03, KTY24, Kur20b, LWK94, LHMH91, LC09, LHC^{+}02, LLZ^{+}04, Lin98, LH20, LBR^{+}22, Maa20, MR85, Mat97c, MD20, MS87]. Systems [MMB^{+}08, Mye81, NMF^{+}23, OKH^{+}12, OW01, OYK^{+}17, PVS^{+}11, PNDG04, Pap89, PGL97, RSE01, RBB21, Rit97, SK02, San97b, SNM^{+}22, SSH^{+}03, Sos94, Sto94, Str98, SLB04a, SLB04b, SUF^{+}12, SMJ^{+}11, Tab84, TP10, TS91, Tal93, TMJ13, TIT^{+}13, TA16, TS14, ULS^{+}00, VPV12, VM88, VC11, WLY^{+}21a, WHP^{+}13, XYCS02, YBS17, Ano02c, Ano03e, AGH^{+}91, BM95, Boa96, Bos04b, Bos06a, DS95, ESW97, EKM^{+}95, Fly97, Gre95b, Han96, Hea87, JC84, KKC93, KKT^{+}91, ME95, Mel87, Pee87, Shl93, SLM^{+}97, Ste05a, SU95, Swa19, TS95, VS87, Wil84, ZG96, vW83, DVQ96, Lav02, Tab84]. Systems-Design [DM88a]. Systolic [AHKY19, JLWL20, MCC^{+}94, MM96, dG95].

T [BMM15, CSM^{+}21, FZW^{+}12, SK02]. T-Engine [SK02]. T4 [SGG^{+}12]. T414 [NT89]. T5 [FJL^{+}13]. T800 [HMSS87]. Table [AS22, Ano13j, Ano14-37, Ano14-35, Ano14-36, Ano16-42, Ano16-43, Ano17-54, Ano17-49, Ano17-50, Ano17-51, Ano17-52, Ano17-53, Ano18-39, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano19-40, Ano19-41, Ano19-42, Ano19-43, Ano19-44, Ano19-45, Ano20-65, Ano20-66, Ano20-67, Ano20-68, Ano20-69, Ano20-70, Ano21-69, Ano21-70, Ano21-71, Ano21-72, Ano21-73, Ano21-74, Ano21-75, Ano22-74, Ano22-75, Ano22-76, Ano22-77, Ano22-78, Ano23-77, Ano23-78, Ano23-79, Ano23-80, Ano23-81, Ano23-82, Ano24z, Ano24-27, Liu02]. Tackling [Dur96]. tactics [Gre00b]. Tag [Mey04]. Tag-Free [Mey04]. Tail [MW19, Gre07f]. Taint [YYK^{+}20]. Taiwan [Kah91b, Kah92a]. Take [Ano14-38, Ano14-39, Ano15-41, Ano16-48, Ano16-47, Ano16-46, Ano17-55, Joh90b, Mat21b, Ano16-45]. Takeda [Ano01d]. Taking [MGP21]. Talent [Emm07a].

Talisman [Ran97]. Taller [Joh19c]. Target [EK16, LS96]. Targeting [Eng00j]. Task [BSP^{+}17, FK83, KJL^{+}10, KKL^{+}09, FBGB96, FMT91]. Task-Centric [KJL^{+}10]. Task-Driven [FK83]. Task-Parallel [BSP^{+}17]. tasks [TCF96]. Taste [Ste86f]. Taxonomy [Maa20]. TCAM [ANC05, CM04, YKL05]. TCAM-Based [ANC05]. TCAMS [WSZS05, SG01b]. TCN [SDF^{+}23]. TCN-CUTIE [SDF^{+}23]. TCO [GHLK^{+}12]. TCP [MFM02, SL03, SML04]. TCP/IP [SL03, SML04]. Tea [Chr90, Joh90a]. teach [Ano94c]. Teaching [DMG00, Hyd00]. team [FHMS96]. Tech [Ano98k, Ano17-30, Ano19-32, Cha85b, Kah93c]. TechIgnite [Ano16-44, Ano17-56]. Technical [Ano98-37, Gre16b, Mat87, Mat10d, Mat13c, Mat83, Ste89d, Gre96f, Sak99a, Ste94f]. Techniques [AR83, Ano01a, MA83, PV01, Sim00, VE10, WJM^{+}05, CMR97, Pet92, Yea96]. Techno [Gre16d]. Techno-Optimists [Gre16d]. Technological [Gre18d, Zsc84]. Technologies [BC21, Che19, GHRS89, Has94, KJL16, Koe86, LCS92, LNLG20, LWK94, LXB07, MCM^{+}16, PCW15, SYKM11, SMAS16, TIT^{+}13, TC15, XZ19, Gre99f, Mat01e]. Technology [ANS96, Ano88g, Ano96o, Ano01h, Ano18t, Ano20-71, BMG^{+}21, Bor99b, Car93, Cri97, DRM^{+}23, Dav02, Dia95b, Eng00a, Eng00c, FRS^{+}09, Gre02b, Gre17d, Gre18e, Gre20c, Gre21b, Gre22e, Gre22f, HSP^{+}01, HYM^{+}90, Hen21a, Ing99, JW99, Kah92b, KKD^{+}07, KGDW^{+}13, KM03, LZY^{+}10, Mat07d, Mat11b, Mea96, Mis93, Mye93c, NFQ03, NKI^{+}09, OFW99, PW96, PZB^{+}19, PcFH^{+}02, Sak97, SDF^{+}23, Ste98f, Ste85h, WAA^{+}20, WN92, Ano92f, Ano01c, Ano01f, Ano02d, DP97, Far84, FN94, Gre95c, Gre97a, Gre97e, GGJ^{+}96, Jae82a, Jae82b, Jae82c, Jae83, Mat95b, Mat01a, McL87, Sak99c, SK97, Sla96, Vic93].

Technology-Based [KGDW^{+}13]. Teeth [Smo87d, Ste01a]. Tel [Gre18e]. Telecommunication [MS87]. Telecommunications [Fra96]. Telematics [Kir90b]. Telephony [Gre02c]. Tells [Ste09b, Ste13, FHMS96]. Temperature [HAWC^{+}11, KC09, MSY^{+}22, MSB^{+}17, SPKJ06, SSH^{+}03, SBG^{+}07]. Temperature-Aware [HAWC^{+}11, SPKJ06, SSH^{+}03]. Temperature-Resilient [MSY^{+}22]. Temporal [PVS17, THC18, TVV^{+}21]. Temporally [BUMV95]. TEMS [Ano18t]. Ten [Alt13c, Gre16d, mHP18]. Ten-Year [mHP18]. Tenancy [DK18]. Tensor [CGG^{+}21, JYPP18, DtEt22]. TensorFlow [NY22]. Tera [Mat97a, MIM^{+}97]. Terabit [AML^{+}03, Yun01]. terabits [MIM^{+}97]. Teraflops [HVS^{+}07]. Teraflops/W [SB23]. TeraPHY [WAA^{+}20]. Term [AS99, IBM05, Yi22a]. Terminals [EMYN00, HC99]. Ternary [GGB^{+}15, Liu02, PS03, SDF^{+}23]. Tesla [LNOM08, TSV^{+}20, TSW^{+}23]. Test [LHC^{+}02, LHL09, MB15, MBTS16, Sak02f, WDK^{+}20]. testability [AJR86, WL92]. Testbed [HDG^{+}22]. Tested [Ano87f]. Testing [AR83, KJP^{+}13, OFKS23, TGE95, AQT^{+}92, JBF94]. Tests [Ano87e, Ano03e]. Tetrahedral [LSL^{+}15]. Texas [FLRB86, Gas21]. Text [EIB90, PAC^{+}14, HC83a]. Text-to-Speech [EIB90]. Texture [Dog12]. TFP [Hsu94]. Thank [Ano20x]. theft [SS82]. Their [Alt13e, Ste86a, Won03, NM96]. Them [Alt13d, Smo87d, CG95, Rob01b]. Theme [Alt13f, Eec16c]. Themes [Alt14c, Del92, Eec17d, Mat95e, Mat04c]. Theory [PMS23, RGK19, Kah91e]. There [Cai89, CDBY23, Gre15f, LX10, War91f, Ano95c, Gre00b]. Thermal [BDJS07, CPS^{+}18, GKS06, KC09, LLSS05, Soo93]. Thermoelectric [KCHA21a, KCHA21b]. These [Joh20d]. Theta [HM93]. They’re [Rob00a]. Thin [KCHA21a, KCHA21b, STT^{+}15]. Thin-Film [KCHA21a, KCHA21b]. Things [Chu18, GAGV22, MK22, AKJF22, AAC^{+}16, EK16, FJB^{+}22, KHL^{+}16, RK16, RNN^{+}16]. Think [Ano88d]. Thinking [Loc03, Mat05c, Mat07d, Mat09b]. Third [HL99, SBJ13]. Third-Generation [HL99, SBJ13]. Thought [Luu90b, Mat13b, Pat90, Gre95d]. Thoughts [Eec17f, FH05, Kir85a, Lei98, Moo03, Mud15, Pea95]. Thousand [Gre06e]. Thread [ADJK20, BSC08, CJH^{+}12, EE10, FZW^{+}12, KG05, KPMHB11, KBH^{+}04, MB05, RSC^{+}06, ROA13]. Thread-Based [KG05]. Thread-Level [CJH^{+}12, FZW^{+}12, RSC^{+}06]. Threaded [EHP^{+}07, SGG^{+}12]. Threads [LPC12, TT12, WCW^{+}04]. Threat [WZL20]. Threats [Joh23c, US23]. Three [Gre19b, Lou91, YRC^{+}22, De 83, Jag97, NA84, SM85]. Three-Dimensional [YRC^{+}22, Lou91, NA84]. three-joint [SM85]. Threshold [AKK15, CB10, DFG^{+}13, KKT13, RPL^{+}17]. Thrives [Ano98h, Gre19f]. Throttle [MK10]. Throughput [CCYT05, CDS^{+}15, CD09, DSG^{+}22, HV04, HKC10, NG87, PHB15, RKA^{+}20, SLC^{+}14, SYY^{+}11, WK13]. Throughput-Optimized [SLC^{+}14]. Throughput-Oriented [PHB15]. ThruChip [MKT^{+}13]. thud [Mat96c]. Thumb [SCG95]. ThunderX3 [SSR21]. Thwarting [LWML16]. Ti [ZHPR17, Ano97-32, Bon21, JLSM03]. Ti-States [ZHPR17]. Tianhe [XLW^{+}12]. Tianhe-1A [XLW^{+}12]. Ticks [Del94b]. Tie [Ste84a, Ste92a]. Tie-in [Ste84a, Ste92a]. Tied [Ste83d, Ste83c]. Tiered [DXT^{+}18]. Tiering [KKS^{+}23]. TigerSHARC [FG00]. Tightly [Kir85b, Pre91]. Tile [WGH^{+}07]. Tilting [Ste94e]. Time [AT09, BTK^{+}23, BR21, CR95a, CFZ^{+}99, CR95b, CWB94, CFO^{+}18, CFM^{+}97, Cro85, DLR02, Dea04, EPZ02, FBC87, FOP^{+}19, FGC^{+}14, Gre22f, JN21a, KKL^{+}09, KDK^{+}89, KTC18, LPL86, LHN95, ML05, MAS^{+}05, MBSP02, Mat03e, MBP^{+}85, MB15, MBTS16, NJZL^{+}17, Noh19, OKH^{+}12, PP92, RCR04, Rea86, RSE01, Sak02f, SK02, SRL91, SK97, Ste88e, Ste94a, SUF^{+}12, TS91, TGE95, UCS^{+}10, Yi21b, Yi22a, Yi23a, Yi23b, EKM^{+}95, Fly97, Hea84, Hea87, KE89, RLG94, Rit97, RH91]. Time-Based [NJZL^{+}17]. Time-Efficient [BR21]. Time-Multiplexed [SK97]. Time-Triggered [MBSP02]. Timed [Kah93i]. Timely [GPSS83]. Times [AML^{+}03, Ano97-33, Pri93a, PW96]. Timeshared [CJ85]. Timing [EDL^{+}04, MKAC18, SKA^{+}14a, VCD16, WPH^{+}23, XYCS02, YFDV19, ZHPR17]. Timothy [Mar17]. Tiny [Ano88h, Ano02d, Ano03e, BTK^{+}23, Joh23e, PMS23, RT23, ZFW^{+}23, MIM^{+}97]. TinyIREE [LBR^{+}22]. TinyML [Joh23e, RM23, ZRB^{+}22]. TLP [SNL^{+}03]. TM [FSBA12]. TMS320C25 [CPH90, FLRB86]. TMS320C25-Based [CPH90]. TMS320C30 [PS88]. TMS320C82 [Gol96]. TMS34010 [GAAR88]. TMS390C602A [DKB^{+}90]. Tnet [Hor95]. Tobus [SSH88]. Today [Ano21-68, Cla03, Sla96]. Tofu [AIH^{+}12]. Token [MHW03, DM86, JKN96]. Tokens [Ond96]. Tolerance [Dra00, EM84, Gro94a, Gro94b, Hum84, Kir87, MS84, MKP06, Pow94, Rag84, RTHA05, Sos94, SRA^{+}04, OFG88]. Tolerant [AF84, AGJL98, ALGJ01, BDSC21, CK98, EVM^{+}98, GAR^{+}06, IEB^{+}14, Joh84, Kir89a, KDK^{+}89, KJC^{+}23, LWB09, RSS^{+}08, RSE01, SB84, SKA^{+}14a, SGC94, Str98, TMA18, YW94, YNS^{+}14, YW88, AGH^{+}91, Bos05f, DGW^{+}94, JKN96, PC01, WJR88, Ano18d]. Tolerating [BRmWH06, HNR10]. Too [Gre03d, Gre09c, Mat05d, Ste92b, Rob99f, Ste99d]. Tool [Ano00m, BM85, DMP91, Eng00l, GTF97, GH88, GEH^{+}23, MG89, MG88, PGL97, San97b, Ano01c, MM96]. Toolbox [ENSD03]. Toolchain [ACZ^{+}22]. Toolkit [Mat93c, Mat97d]. Tools [FHP00, Hoo89a, KFF00, Mat15b, Nic91, TCF96, WG92, Ano92b, Ano94b, Ano95c, Ano98-30, Ano98-31, Ano99-27, Mat01f]. Top [ABZ08, Alb04, Alt12e, Alt13c, Alt14f, Ano19o, Ano20-71, BL23, CS15, Dia93a, Dwa19, Eec15e, Eec16e, Eec17f, Eec18f, ET09, Eng00o, EEKS07, FL13, FV12, GT22, HGPT12, JQ17, Jim21, Joh19f, Joh20d, Joh22f, Joh23f, Kim20, Kur21d, MS16, MRLB03, Mud10, Mye92a, PM11, RG07, TM14, Tor06, Wen18]. Top-Down [EEKS07]. TOp/s/W [SDF^{+}23]. Topics [Alt12a, Ano14s, Mat06a]. Topologies [MRSV11, PC01, CK95]. Topology [KDSA09, VPRS14]. Total [KCAR18]. Tour [Fra94, Mat21c]. Tower [War92d]. Tower-of-Babel [War92d]. Toy [MG88]. TPUv2 [NPY^{+}21]. TPUv3 [NPY^{+}21]. Trace [DCMS20, Kha00]. Trace-Driven [Kha00]. Tracing [MAK19]. Track [Mye82b, Rob97d]. Tracking [CSL^{+}06, PDT98, TLW^{+}10, TKS^{+}22, YZW^{+}23, YYK^{+}20]. Trade [AF88, FHP00, Ste93b, Ste95b, Wet86, Pap96, SMHB91]. Trade-Offs [AF88, FHP00, Pap96, SMHB91]. Trade-Policy [Wet86]. Tradeoffs [BEL^{+}23]. Trading [WGA^{+}09]. Traditional [LCP^{+}11]. Train [Kir90d, Kir90e, KZ01]. trainable [KWGG95]. Training [DQCL24, DWF^{+}21, HIP^{+}22, JAS^{+}22, KJC^{+}23, LKJ^{+}22, MD20, NPY^{+}21, PEZ^{+}19, SKM23]. Trak [Mat91c]. Transactional [ADF^{+}10, AAK^{+}06, BMV^{+}08, FSBA12, HCW^{+}04, HCU^{+}07, MCC^{+}07, RG03, RRP^{+}08]. Transactions [Ano14s, Ano16y, Ano19p, Ano19i, Ano20u, Ano20o, Ano21u, Ano21m, Ano21n, Ano21o, Ano21p, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22q, Ano22r, Ano22s, Ano23w, Ano23x, Ano24m, GP90, Ano19-28]. Transceiver [GDES08, IGH^{+}99]. Transfer [CHAF22, LDL17, MA83, PDL08, WLD15, Ano02e, Reg92]. Transfer-Based [WLD15]. Transfers [SKC^{+}23]. Transform [LNV89]. Transformations [WLKN22]. Transformer [WMH^{+}10]. Transforming [PO04, SP92]. Transforms [SMR18, AAG^{+}10]. Transient [GSVP03, GV06, HANR13, Sos94]. Transient-Fault [GSVP03, GV06]. Transistency [LSBM17]. Transistor [Bor05, KPN^{+}20, RC13, Ano01h, Ano03b]. Transistors [DDG^{+}19, Kid14, Ano03b]. Transition [MNU^{+}15, Moo03, Moo04a]. Transition-Aware [MNU^{+}15]. Translation [Bha17, Bha18, PBFC21, PHB15, RLS11, SL84a]. Translations [GKA^{+}16, SDG^{+}21]. Translator [CHH^{+}98, Mye83b]. Transmission [GT83, War90d]. Transmission-Lines [GT83]. Transmissions [Gre20d]. Transmitter [DP97]. Transmitters [STR^{+}13]. Transnational [Ste05a]. Transparent [ZG96]. Transponders [GD01]. Transport [CMC98, Sav99a, SAA^{+}99, SABS20]. Transputer [NT89, Tal93, HMSS87]. Transputer-Based [Tal93]. Transputer-T414 [NT89]. Transputers [Kah92e, WS90]. traps [Gre05e]. Traversals [KCKP14]. Tree [PMM15, WHJ^{+}23]. Tree-Based [PMM15]. Trees [GCL^{+}20, MMESG^{+}20]. Trends [AS91b, All84, Ano20-71, BY17, Bos03c, Car93, Con03, Fra00, Kat97, Lee94, MBS92, PC93, Sak88, SVL03, WN92, Won03, Bos04c, Gre19e]. Trial [Smo86a]. Trial-Use [Smo86a]. TriCheck [TML^{+}18]. Triggered [MBSP02, PPA^{+}14, TT12]. Triggers [Gre20d]. Trimming [CAH86]. Trip [AML^{+}03]. TRIPS [GKS^{+}07, SNL^{+}03]. Tristate [FKL01]. Trojans [CML^{+}23]. Trolls [Emm06c]. TRON [KWM89, SSH88, Sak87b, Sak87a, Sak87d, Sak90a]. troublesome [Mat96f]. Trucking [Gre18a]. True [McD21, Ano95d, Ste05b]. Truly [Alb07e]. Trump [Gre17d]. TRUSS [GKS^{+}05]. Trust [KSE^{+}22]. Trusted [GSS^{+}07]. Truth [Rob97e]. TSA [WZL20]. TSA-NoC [WZL20]. TSMC [Ano03b]. Tunable [RLV85]. TuneV2 [MM23]. Tuning [MM23, Pap96, PGL97, YNS^{+}14]. Tuple [LK10]. Turing [Bur20, RTJ20, RTJ21]. Turn [Ano97z, Ste03b]. Turning [Hig85, WPH^{+}23]. Turns [Ano96c, KvdW09, Ste04d]. Tutorial [Col89, Gus84, Hoo89c, Jae82a, Jae82b, Jae82c, Jae83, Pri89, RG88]. TV [Ste08a, Pet92]. tweezers [Ano92a]. Twenty [Gre15d, Gre15e]. Twin [VPRS14]. twisted [War91g]. twisted-pair [War91g]. Twitter [Mat09e]. Two [Gre17e, Gre20f, KSI^{+}96, Mat13c, MBG^{+}16, RYK18, ZZY97, DGW^{+}94, Fur88, MKRC97, Rob99f]. Two-chip [KSI^{+}96]. two-dimensional [DGW^{+}94]. Two-Level [MBG^{+}16]. two-size [Fur88]. TX [GDLT86]. TX1 [MKOK88]. TxLinux [RRP^{+}08]. Tydi [PVB^{+}20]. type [SSB95]. tyranny [Ste97e].

U.S. [Gre22f]. U2 [FMN^{+}13]. UAV

[CHAF22]. UAVs [LCN^{+}22]. Ubiquitous [CFK^{+}10, FHL^{+}03, Gre06f, SCA^{+}12, STM02, TSP02]. UCIe [Sha23c]. UCX [TSA^{+}22]. uGEMM [WLY^{+}21b]. ugly [Rob00e]. ULSI [Rüc02]. Ultimate [Del91c, QT21, RNN^{+}16]. Ultra [Ano17-57, BS17, CDY^{+}18, CEP^{+}17, Eec17e, FD17, LM16, RNN^{+}16, SCA^{+}12, TUI^{+}01, WRA^{+}14, YBS17]. Ultra-Large-Scale [RNN^{+}16]. Ultra-Low [CDY^{+}18]. Ultra-Low-Latency [LM16]. Ultra-Low-Power [Ano17-57, BS17, CEP^{+}17, Eec17e, YBS17]. Ultra-Performance [FD17]. Ultra-Reduced [WRA^{+}14]. Ultraefficient [ZSB21]. Ultrafast [Ano88g]. Ultralow [OYS^{+}11, SB07]. Ultralow-Latency [SB07]. Ultralow-Power [OYS^{+}11]. Ultrasound [NHY^{+}22, SYW^{+}14, SCYY11]. UltraSPARC [HL99, NCT^{+}98, TO96]. UltraSPARC-III [HL99, NCT^{+}98]. Ultraviolet [Ano96l, Ano97-30, Ano01f]. UMTS [Ste05c]. Unary [WLY^{+}21b]. Unbiased [CNC^{+}16]. Unbounded [AAK^{+}06]. Uncertain [BMM15, WD03, BMM15]. Uncertainty [Gre98f, MT05, TKS^{+}22]. Uncle [War90c]. Uncomfortable [Gre20e]. Uncompressed [GDES08]. Unconscious [Mat13b]. Uncovering [DK18]. Underclocked [KST12]. Undergraduate [Cle00b]. Underground [GAGV22]. Understand [KCS^{+}20]. Understanding [CEAY23, NAA^{+}20, RCK^{+}21, SSLV15, SD21, War92d, Ano98z]. Undervolting [KSE^{+}22]. undiminished [DBDF97]. Unexpected [Gre98f]. Unfair [Ste02c, Ste01c]. unFRANDly [Ste06b]. Unification [SWL90]. Unified [AAC^{+}23, HO99a, LNOM08, San97b]. uniform [KHF86]. Unifying [GHN^{+}12, WLKN22]. Unigraphics [Ano00h, Eng00l]. Uniprocessors [CD97a]. Unique [Fai82b]. Unit [Ano98-36, BBC^{+}15, BCP01, BCF^{+}14, DKK21, GE86, HABHW^{+}18, JYPP18, KKH^{+}24, KBN16, KIS^{+}00, WHCK18, WLP^{+}15, YNS^{+}14, ZUNN18, CM86]. United [Gar93, Ste91b, Ste92a, Zsc84]. Units [CK11, KTC18, LKGL24, MKAC18, Mil90, CH94, WHKM93a]. Universal [RBB21, Sha23c, Ano83, HP81]. University [ADC00, Ano22z]. University-Industry [ADC00]. UNIX [IJ98, Mat97c, Hin88, Man92, Mel87, YMA^{+}13, ZG96]. Unix-based [Mel87]. Unlike [Mat96f]. Unlimited [Cas95, Ano17-46]. Unlocking [JSY^{+}16]. Unnecessary [NGSW17]. Unobserved [Ste92f]. Unorthodox [Gro02]. unpatented [Ste04c]. Unreliable [Bor05, WK13, Bos06a]. Unresolved [Ste03a]. Unstable [MLS^{+}16]. Unstructured [LSL^{+}15]. Untitled [Ano00n, Del94a]. Unveiling [GZC^{+}20]. Unveils [Ano96h, Ano99m, Ano03b]. Upcoming [Eec17d]. Updatable [XLX^{+}23]. Update [Ano98w, DBC^{+}98, FS05, Ste01b, Ste08b, War89b, Ste05c]. Updates [Ste09d]. Updating [SG01b, Ste00b]. upgradability [Wal97]. Upgrading [Mat95d]. Upholds [Ste07e]. Upon [Hil19]. Ups [Ste04a, Ano03b, Ste04b]. Upsets [GXMZ13]. Upsilon [Ano17-58, Ano17-59]. urban [Rob97e]. USA [Ste09b]. Usage [BPT^{+}11]. Use [Gre02f, Hac01, HCPS03, RTJ20, RTJ21, Smo86a, Ano00g, Dia95d, HS85, RH91]. Useful [MSS15, Ste08e]. User [BFK^{+}85, CDS07, DS94, MNU^{+}15, Mat13b, MCF^{+}85, Ste89a, WBHv98, Abr83, DBDF97, Ste89c, Ste89d, Ste89e, Ste90e, Ste96f, WHKM93a, WHKM93b]. User-Level [CDS07, WBHv98]. Users [Mye90]. Uses [KTK13]. Using [ABA^{+}21, ATS^{+}22, ACKM05, AS99, AHKY19, CM04, CMR97, CSC^{+}22, CES17, CK11, CFM^{+}97, FHP00, GFL^{+}17, GSC97, GK97, Gol96, GRD22, GGJ^{+}96, GKS06, GSS09, Ham00, HYM^{+}90, JAS^{+}22, KLD^{+}94, Kid14, KIR19, KSR^{+}99, KBH^{+}08, KWGG95, KP90, KPKJ08, LHM99, LLT^{+}08, LWB09, LK10, LHN95, MMESGQ22, MKG^{+}20, MTS^{+}12, MIM^{+}97, MD20, MC95, MKRC97, MMB^{+}08, NS05, NRV^{+}06, OML^{+}07, PFC^{+}02a, PFC^{+}02b, PDL08, RGH^{+}10, RLG94, SKLY97, SB07, Sha23b, STM02, SLM^{+}97, SYY^{+}11, TKI^{+}14, TSA^{+}22, TTF96, VVRV95, WHJ^{+}23, XYT^{+}23, YBNS15, ZCW^{+}14, ZSS^{+}19, ZIM^{+}07, AML05, BJ14, CK95, CS14, DKSL04, JKP89, LKM92, MLL^{+}18, PS03, Rit97, SK97, SSB95, VBB95]. Utilization [MTS^{+}12]. Utilizing [KTC18, RES^{+}13]. UWB [Eng00j].

v [Sla90a, Ste12, Ste14a, Ste14b, AS22, Ber86, DXT^{+}18, GDES08, LWC^{+}16, PGW^{+}20, SB23, Ste90e, XYT^{+}23, Yi22d, ZSB21]. V&V [MKAC18]. V-System [Ber86]. V.42bis [Tho92]. V/Tensor [DtEt22]. V15 [Ano00h, Eng00l]. V2.0 [Ano96t]. V20 [Cho21]. V60 [KKY88]. V60/V70 [KKY88]. V70 [KKY88]. V80 [KSWM90]. V830R [SANK98]. V830R/AV [SANK98]. Vacation [Mat92a]. Validate [KLD^{+}94]. Validated [RPE10]. Validating [GBW^{+}23, Kha00]. Validation [ABC99, BFLS01, BCA99, HO99b]. Validation-Based [ABC99]. Value [CL04, Gre10b, Gre17f, LLL^{+}16, MAJ^{+}18, PS15, Gre05d]. Value-Based [CL04, MAJ^{+}18]. Valued [PFC^{+}02a, PFC^{+}02b]. Values [PMS23]. Vantage [SK12]. vaporware [Ste95c]. Variability [AW03, Bor05, LCWB08, RC13]. Variable [LWB09, PPP01]. Variable-Length [PPP01]. variables [KHF86]. Variation [Bos05f, GR95b, KKT13, KC09, LWB09]. Variation-Tolerant [LWB09, Bos05f]. Variations [UTB^{+}06]. Variety [Gre14b, CR95b, Gil96a]. Varifocal [HLIT20]. Various [KHS^{+}23]. Vast [Mye84a]. VAX [Abr83]. VAX-11 [Abr83]. VAX-11/780 [Abr83]. VAXes [Hen21b]. Vector [AT93, CSM^{+}21, KP03, KBH^{+}04, KIS^{+}00, LSZ82, NAJE22, NDR^{+}22, SBB^{+}17, ZUNN18, Dur96]. Vector-Thread [KBH^{+}04]. Vectors [KTK13]. Vehicle [Mye93b, NNS^{+}93, Shl93]. Vehicles [KTI^{+}15, SSK23]. Velocity [IKK96]. Velox [ADF^{+}10]. venture [Ano03b, Ano03c]. Verification [EGL^{+}90a, JPOB20, KTY24, LHM99, SKA^{+}14a, STR^{+}01, TML^{+}18, TLM19, ZBES15]. Verify [AS99]. Verifying [HS99, LPM15]. Verilog [Ano96r].

Versatile [HHNK09, LW94]. Version [Mat93f, Mye85b]. Version-B [Mye85b]. Versus [KSV^{+}21, Mar86, Pit96b, Sch96, Ste98c, Bos04a, Bos04e, FDS^{+}17, HCP^{+}16, Ste95a]. vertebrate [Boa96]. Vertical [Fet95, HOHCV99, Nar19]. Very [Alt11f, JL11, LH12]. VHDL [KP90]. VI [AS95, Yi23a]. via [LTQZ07, NM22, PPA^{+}14, Ste96e, TLM19, WCW^{+}04, YZW^{+}23, ZFW^{+}23, ZKP^{+}23]. viable [Ano03b]. Vibrant [Eec15a]. Vicarious [Ste04e]. Video [HSR18, IKN^{+}99, KIM^{+}09, LHC^{+}20, LLT^{+}08, ML21, Nic88, PP92, RSC^{+}22, SC91, SP09, Ste89b, DKM^{+}92, KSI^{+}96, Pet92]. Video-Mining [LLT^{+}08]. Videoconferencing [Gol96]. View [All86b, Ano94d, Ano96n, Ano97t, Dia99, Dia00, Fer98a, Fer98b, Gre12d, Hur97, IJ98, Pit95, Sla90d, TW00, VN10, Wea97a, Wea97b, Wil95b, Wil96, Wil97, Ano95d, Kah93g, Pri94a, WWR97]. Viewpoint [Bos21]. VII [Yi23b]. VIII [Yi24b]. VIIIfx [MYK^{+}10]. villages [Ano94b]. Violating [Ste08c]. Violation [Ste07c, Ste07e, Ste13, Ste06b]. Violations [LTQZ07, LDCS09]. Virtual [Ano96m, Ano96s, Ano99h, BDSC21, BMS16, Bha17, BDF^{+}95, CD97a, CD97b, CMC98, DRM^{+}98, GKA^{+}16, Gre99d, JM98, Kah93h, KG05, KKP^{+}09, KPKJ08, LHC^{+}20, MM83, ME95, MMESGQ22, MH08, OT97, STR^{+}01, SKJ^{+}11, SsSMB24, VMW^{+}19, WCW^{+}04, YBNS15, ZL16, Ano99w, RH91, Mon97]. Virtual-Address [CD97a, CD97b]. Virtual-Memory-Mapped [BDF^{+}95]. Virtualization [DMG^{+}15, GHS17]. Virtualized [BR21, YE11]. Virtuous [Gre21f]. Virulent [Gre99e]. Virus [LLLL09]. VIS [TONH96]. visibility [Ano96n]. Vision [BBC^{+}15, Boa96, GZC^{+}17, KII09, OKH^{+}12, PMR^{+}22, RCK^{+}21, SGL93, BCF^{+}92, HS92, Mat96b, Mat98b, SPT^{+}92]. Visions [LG24]. Visiting [Mat97b]. Visual [CEP^{+}17, DPBW19, IO16, KWGG95, SCS^{+}09, LC91, Ano96t]. Visual-Slickedit-V2.0 [Ano96t]. Visualization [VPV12]. Visually [LMC^{+}83, GRP83]. Vital [Alt11e, Gre20f]. Vitality [Gre16c]. VLIW [Ano00g, Ano03f, BLO00, Sla89]. VLIW/EPIC [Ano03f]. VLIWs [ST21]. VLSI [Sak87b, ACRV96, AJR86, BTHS92, CT95, CPZ89, Con03, DP97, DGT89, DM86, EM84, GHRS89, GGJ^{+}96, HF81, IN87, IKK96, KWM89, KWGG95, Laz89, LHMH91, LC91, LKM92, MKNK83, MM96, Mur89, MCH^{+}94, Pee87, RJHK89, Sib84, TPV89, VJ89, vdDD90]. VME [Fis85, Pri86]. VME64 [Reg92]. VMEbus [AQT^{+}92, Hea87]. Voice [WMSH09]. Vol [Ano03a, Ano05, Jef84, RGF96, Sav99a]. Volta [CGF18]. Voltage [AKK15, CCA^{+}19, GKS21, KJP^{+}13, LWB09, MSA^{+}03, RGH^{+}10, RKK^{+}11, RDJ^{+}13, WGA^{+}09, Ano02b]. voltage/low [FN94]. Voltages [KKT13]. Volume [Ano96a, Ano00a, Ano01b, Ano02a, Ano06, Ano07, Mye93c, Tab84]. Volunteer [Ano21-58, Ano21-59, Ano21-60, Ano23-58, Ano24v, Dia96a]. Voting [Gre08e]. Voyager [ADC00]. VP [AT93]. VPR [MEB^{+}20]. VR [YRC^{+}22]. VRTX [Rea86]. vs [Ano97i, Dav98, EHP^{+}07, GSS^{+}07, Gus85, Kah92b, Pee87, Ste87c]. VSI [Ano97t, Wil97]. Vu [Gre18d]. VulHunter [QLLG15]. Vulnerabilities [GSS^{+}07, HMR^{+}19, QLLG15]. Vulnerability [MWE^{+}03].

W [JBF94, SDF^{+}23, SB23]. W. [Luu90a]. Wafer [Ano87g, HOHCV99, Lau21, ZZNT^{+}23, Ano02c, Gre04d]. Wafer-Scale [Lau21]. Wagging [Gre07f]. Walking [LZX^{+}18, Ste00d]. Wall [Bha17, Bha18, CSC^{+}05, Eec15b, Kir90a, WS13, WA13].

Wally [Gre12e]. Wan [Fra96]. want [Ano94c, Rob97d]. Wants [Smo86a]. War [Bri94, Dai94, Dav93]. Warehouse [Bar21, HLZ^{+}16, KDH^{+}16, LRC^{+}09, MTS^{+}12, NAA^{+}20, RSC^{+}22]. Warehouse-Computing [LRC^{+}09]. Warehouse-Scale [Bar21, HLZ^{+}16, KDH^{+}16, MTS^{+}12, NAA^{+}20, RSC^{+}22]. Warp [SMR20]. Warpage [Ano97v]. Wars [All86a, All86b, Jam90, Ste86g, Tau86, Gre06b, Ste97d]. Was [Kir91c]. Watch [Ano16-48, Ano16-47, Ano16-45, Ano16-46, Ano22a, Ste99e]. watchword [Kah93a]. Watermarks [YYH98]. Waters [Gre20e]. Watts [Ano23-83, Ano23-84]. Wave [Ano87a, Mye89a, XWZ09, SLM^{+}97]. Wave-Pipelined [XWZ09]. Waveguides [CS13]. Wavelength [ZLTW13]. Waves [Dia95b]. Way [Alt12f, Ano97r, AK00, Cai89, Kir91a, KAO05]. WE32100 [FN86]. WE32200 [HSW^{+}89]. Wealth [Gre08c, Gre08d]. Wear [SWL11]. Wearable [Fer98b, Pen99, Pen01, Sta01a, Sta01b]. Wearables [Ano15-32]. WearARM [LAT^{+}01]. Wearing [SJO01]. Web [Ano00d, BDH03, Dia95c, Eng00l, KFF00, Mat98c, Ste99b, Ste99c, ZHR15]. Webworks [Ano99-33]. Week [Ano20-57, Ano23-63, Ano20-54, Ano20-55, Ano20-56, Ano22-64, Ano23-60, Ano23-61, Ano23-62]. Weights [BUMV95, MKG^{+}20]. Weird [Ste03b]. Welcome [Alb09]. Welcomed [Mat89a, Wes89]. Welcoming [Eec16e, Sak99b]. Well [Mat15c]. Were [ST21, Mat03f]. West [Kir90c, Ste07e]. Wheels [SSK23]. Where [Ano16x, EHP^{+}07, Gre03e, Gre19f, GSS^{+}07, Mat03f]. wherever [Ano14-38, Ano14-39, Ano15-41, Ano17-55]. Which [Alt12f, Gre02f, Mat96f, SLM^{+}97]. While [Ano87g, Joh20d, Han96]. whips [Gre04a]. White [Del94b]. Who [Alt11b, Gre96d, Gre15f, Sla90f, Smo86a, Ste84c, Ste01f, Wil95b]. Whole [GGC^{+}11].

Whose [Ste88e]. Wi [Gre11d]. Wi-Fi [Gre11d]. Wide [RTM^{+}10, RDJ^{+}13, SK01]. Wide-Area [SK01]. Wide-Voltage-Operating [RDJ^{+}13]. Width [WM85, TTF96]. Wikipedia [Gre07f]. Wilkes [Dwa18, KT14, Mar17, Sco14, Ste16]. Will [Ano96u, Ger19, Joh22b, MCR17, Ano97n, Mat06d, Sak00e]. William [Ano01g]. windmill [Ste94e]. Windmills [Smo87d]. Windows [Mat93b, MSWP03, RS93, Fur88, Ano96g, Ano96t, Ano99-33, Fra94, Mat93e, Mat93f, Mat95d, Mat97c, Mat97d, Mat98d, Mat00e, Sca98, ZG96]. Windows-95 [Mat97d]. Windows-98 [Sca98]. Windows-NT [Mat97d]. Winners [MB15, MBTS16]. Winning [Mud15]. Wins [Ano98v, HSNJ21, Ste98a]. Winsocking [Ste95e]. Winwriters [Mat99e]. Wire [AVU^{+}08, BMR^{+}06, BWBJ11, GT83, KBK03, NL02, War90g, Ano02d]. Wire-Delay [KBK03]. Wire-OR [GT83]. Wire-Speed [BWBJ11]. Wire-to-Wire [War90g]. Wireless [ASK^{+}15, Ano96v, Ano00o, Ano01h, Ano02e, CB96, CHAF22, EK16, Eng00l, GSC97, GDES08, Gon99, HC02, SLM^{+}97, Ano00g, Ano01c, Gre05f]. Wires [AFK^{+}21, Lee24a, FKV20]. WISC [Mil88b]. WISCs [Koo88]. Wisdom [Gre23b, Mat99f]. Wise [Ano96q, Hau88c, Per83, Sho85]. Wish [KMPS06]. Wishful [Mat09b]. Within [ML21, RD90, Rob91]. Without [Gre23b, Hec83b, Ste13, Ano99p, Chr96, Gre18b, SMR07]. Witness [SFG^{+}22]. woes [Gre96c]. Wonk [Gre11c]. won’t [Mat95d]. Word [CCG^{+}84, DO84, Mat93b, Gre99e, Mat93b]. Word-length-independent [CCG^{+}84]. Words [Bri94, Dai94, Emm07a, Mat99f, Dav93]. Work [AFGM10, Gre21d, Joh20d, Mat09a, Mat15a, Ano02d, Gre96a, Mat01c]. Working [Mat98c, Rob01d, Ste84e, Ano02c].

Workload [AW03, Bos06e, HE07, IBM05, KKL^{+}09, SWG06, VE10]. Workload-Aware [KKL^{+}09]. Workloads [AMK17, AW06, AJC^{+}20, EE08, FAK^{+}14, KML04, KAV99, PJB^{+}14, RCC12, WAA^{+}21, ZRA^{+}17, dCMA22]. Works [Gre16e, Ano02d]. Workshop [BM19, BCM^{+}14]. Workstation [DGMM00, Hig85, JGF98, Kni85, Lan85b, UBH^{+}94, GRP83, Mar85, RMFG85]. Workstations [ACP95]. World [Ano16-48, Ano19-29, Cle03, GR95a, Gre99d, HO99a, Hum84, Kah92f, Mar21, Sak93, SP92, Ano00g, Ano16-45, Dur96, Rob00b, RH91, Yea96, Ano16-47, Ano16-46]. Worm [ML05]. Would [Ste13, Gre98c]. wrap [Ste97f]. Wrappers [BLW02]. Wringing [DCMS20]. Write [AAP^{+}10, Mye85b, SKJ^{+}11, ZHZ^{+}19, Emm06a, HP81]. Writing [Emm05a, Mat90c, Mat10d, Mat15c, Ano92c, HC83a]. WTL3170 [BSC^{+}90]. WTL3170/3171 [BSC^{+}90]. Wu [Luu90a]. WWW [Ano95c].

X [Sel18, And82b, Ano88g, Ano97-33, Ano98r, GA21, NL02, Tea82, YMA^{+}13]. X-by-Wire [NL02]. X-Ray [Ano97-33]. X-Ray-Lithography [Ano88g]. X1 [DVWW05]. x86 [BCD^{+}11, HWG^{+}09, RPE10, SCS^{+}09, WPH^{+}23, Chr96]. Xbox [AB06, GA21, Mat21b, SO14, Sel18]. XCRYPT [SFP^{+}23]. xDSPcore [KPHP04]. Xeon [SGC^{+}16, Ano01c, AFK^{+}19, RMM^{+}04]. XIfx [YHT^{+}15]. XII [MAT^{+}18]. XIX [Ano15j]. XL [KKL^{+}22]. XMOS [May12]. XS1 [May12]. Xtensa [Gon00]. XVIII [Ano14e].

Y2K [Ste98d]. Yale [Bel12]. Yang [KAK^{+}22]. Year [Ano97-34, Dia96a, Hen21a, mHP18, Kur21d, Mat99c, Mat05e, Mil86, Mye91c, War90b, Mat98d, Mat00b]. Year-end [Mat05e]. Years [Alt13c, Eec15a, Gre15d, Gre15e, IDI^{+}21, Rag21, SRU^{+}23, Ste85g, Mar96, Yu96]. yield [AAW^{+}96]. Yin [KAK^{+}22]. Yin-Yang [KAK^{+}22]. You’d [Ano88d]. You’re [Emm07e, Ano94c].

Z [JW20]. z10 [Web08]. Z80 [Lun85, Pre21, SL84a]. Z80000 [Phi85]. zEC12 [SBJ13]. Zen [SSB20, EBC22]. zEnterprise [CES^{+}11]. Zero [CL05]. Zero-Sensitivity [CL05]. Zip [Gre19f]. ZNET [UBL^{+}82]. Zvi [Gre01a].

References

Alpert:1993:APM

[AA93] D. Alpert and D. Avnon. Architecture of the Pentium microprocessor. IEEE Micro, 13(3):11–21, May/June 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ambrosin:2016:FAB

[AAC^{+}16] Moreno Ambrosin, Arman Anzanpour, Mauro Conti, Tooska Dargahi, Sanaz Rahimi Moosavi, Amir M. Rahmani, and Pasi Liljeberg. On the feasibility of attributebased encryption on Internet of Things devices. IEEE Micro, 36(6):25–35, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060025-
abs.html.

Aananthakrishnan:2023:IPI

[AAC^{+}23] Sriram Aananthakrishnan, Shamsul Abedin, Vincent Cavé, Fabio Checconi, Kristof Du Bois, Stijn Eyerman, Joshua B. Fryman, Wim Heirman, Jason Howard, Ibrahim Hur, Samkit Jain, Marek M. Landowski, Kevin Ma, Jarrod A. Nelson, Robert Pawlowski, Fabrizio Petrini, Sebastian Szkoda, Sanjaya Tayal, Jesmin Jahan Tithi, and Yves Vandriessche. The Intel programmable and integrated unified memory architecture graph analytics processor. IEEE Micro, 43(5):78– 87, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Asprey:1993:PFP

[AAD^{+}93] T. Asprey, G. S. Averill, E. DeLano, R. Mason, B. Weiner, and J. Yetter. Performance features of the PA7100 microprocessor. IEEE Micro, 13(3):22–35, May/ June 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Airoldi:2010:EEF

[AAG^{+}10] Roberto Airoldi, Omer Anjum, Fabio Garzia, Alexander Wyglinski, and Jari Nurmi. Energy-efficient Fast Fourier Transforms for cognitive radio systems. IEEE Micro, 30(6): 66–76, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ananian:2006:UTM

[AAK^{+}06] C. Scott Ananian, Krste Asanovic, Bradley C. Kuszmaul, Charles E. Leiserson, and Sean Lie. Unbounded transactional memory. IEEE Micro, 26(1):59–69, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arvind:2010:PMD

[AAP^{+}10] Arvind, David August, Keshav Pingali, Derek Chiou, Resit Sendag, and Joshua J. Yi. Programming multicores: Do applications programmers need to write explicitly parallel programs? IEEE Micro, 30(3):19–33, May/June 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Amarasinghe:1996:MSP

[AAW^{+}96] Saman P. Amarasinghe, Jennifer M. Anderson, Christopher S. Wilson, Shih-Wei Liao, Brian R. Murphy, Robert S. French, Monica S. Lam, and Mary W. Hall. Multiprocessors from a software perspective — automatically parallelizing benchmark programs to yield the highest SPECfp ratios recorded. IEEE Micro, 16(3):52–61, May/ June 1996. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Allen:1994:DPB

[AAWC94] M. S. Allen, M. Alexander, C. Wright, and J. Chang. Designing the PowerPC 60X bus. IEEE Micro, 14(5):42– 51, September/October 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Adams:1983:MAM

[AB83] W. Thomas Adams and John Brady. Magnitude approximations for microprocessor implementation. IEEE Micro, 3(5):27–31, September/October 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Andrews:2006:XSA

[AB06] Jeff Andrews and Nick Baker. Xbox 360 system architecture. IEEE Micro, 26(2):25– 37, March/April 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Abdelfattah:2014:CEN

[AB14] Mohamed S. Abdelfattah and Vaughn Betz. The case for embedded networks on chip on field-programmable gate arrays. IEEE Micro, 34(1):80– 89, January/February 2014.

CODEN IEMIDZ. ISSN 0272-1732.

Alachiotis:2021:APU

[ABA^{+}21] Nikolaos Alachiotis, Andreas Brokalakis, Vasilis Amourgianos, Sotiris Ioannidis, Pavlos Malakonakis, and Tasos Bokalidis. Accelerating phylogenetics using FPGAs in the cloud. IEEE Micro, 41(4):24– 30, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arlat:1999:VBD

[ABC99] Jean Arlat, Jérome Boué, and Yves Crouzet. Validationbased development of dependable systems. IEEE Micro, 19(4):66–79, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4066.
pdf.

Alser:2020:AGA

[ABC^{+}20] M. Alser, Z. Bingöl, D. S. Cali, J. Kim, S. Ghose, C. Alkan, and O. Mutlu. Accelerating genome analysis: a primer on an ongoing journey. IEEE Micro, 40(5):65– 75, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ahmad:2016:NMS

[ABG^{+}16] Sagheer Ahmad, Vamsi Boppana, Ilya Ganusov, Vinod Kathail, Vidya Rajagopalan, and Ralph Wittig. A 16nm multiprocessing systemon-chip field-programmable gate array platform. IEEE Micro, 36(2):48–62, March/ April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2016/02/mmi2016020048-
abs.html.

Amid:2020:CID

[ABG^{+}20] A. Amid, D. Biancolin, A. Gonzalez, D. Grubb, S. Karandikar, H. Liew, A. Magyar, H. Mao, A. Ou, N. Pemberton, P. Rigge, C. Schmidt, J. Wright, J. Zhao, Y. S. Shao, K. Asanovič, and B. Nikolič. Chipyard: Integrated design, simulation, and implementation framework for custom SoCs. IEEE Micro, 40(4):10– 21, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Alastruey:2006:SDH

[ABIV06] Jesus Alastruey, Jose Luis Briz, Pablo Ibáñez, and Victor Viñals. Software demand, hardware supply. IEEE Micro, 26(4):72–82, July/August 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arap:2017:OCO

[ABK^{+}17] Omer Arap, Lucas R. B.

Brasilino, Ezra Kissel, Alexander Shroyer, and Martin Swany. Offloading collective operations to programmable logic. IEEE Micro, 37(5):52– 60, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050052-
abs.html.

Abrahamson:1983:FEP

[Abr83] David M. Abrahamson. A fast entry path into user microcode on the VAX-11/780. IEEE Micro, 3(6):40–43, November/December 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Adve:2008:GEI

[ABZ08] Sarita Adve, David Brooks, and Craig Zilles. Guest Editors’ introduction: Top picks from the computer architecture conferences of 2007. IEEE Micro, 28(1):8–11, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2008/01/mmi2008010008.
pdf.

Agerwala:2005:CAC

[AC05] Tilak Agerwala and Siddhartha Chatterjee. Computer architecture: Challenges and opportunities for the next decade. IEEE Micro, 25(3): 58–69, May/June 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ankit:2020:CAM

[ACA^{+}20] A. Ankit, I. Chakraborty, A. Agrawal, M. Ali, and K. Roy. Circuits and architectures for in-memory computing-based machine learning accelerators. IEEE Micro, 40(6):8–22, November/December 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Addra:1999:MMC

[ACDG99] Modeste Addra, Dominique Castel, Jacques Dulongpont, and Pierre Genest. Microelectronics in mobile communications: a key enabler. IEEE Micro, 19(5):62– 70, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5062.
pdf;   http://www.computer.
org/micro/mi1999/m5062abs.
htm.

Argon:1988:MSP

[ACG^{+}88] Elvira Argon, I.-Lok Chang, Gamini Gunaratna, David K. Kahaner, and Martin A. Reed. Mathematical software:

Plod. IEEE Micro, 8(4):56– 61, July/August 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ascia:1995:DPF

[ACG^{+}95] G. Ascia, V. Catania, B. Giacalone, M. Russo, and L. Vita. Designing for parallel fuzzy computing. IEEE Micro, 15(6):62, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Abella:2003:PCA

[ACG03] Jaume Abella, Ramon Canal, and Antonio González. Powerand complexity-aware issue queue designs. IEEE Micro, 23(5):50–58, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/05/m5050abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/05/m5050.
pdf.

Arekapudi:2005:UHC

[ACKM05] Srikanth Arekapudi, Shang-Tse Chuang, Isaac Keslassy, and Nick McKeown. Using hardware to configure a load-balanced switch. IEEE Micro, 25(1):70–78, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2005/01/m1070.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2005/01/m1070.pdf.

Appiani:1989:EHP

[ACLR89] Enrico Appiani, Bruno Conterno, Vildo Luperini, and Leonardo Roncarolo. EMMA2, a high-performance hierarchical multiprocessor. IEEE Micro, 9(1):42–56, January/ February 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anderson:1995:CNN

[ACP95] T. E. Anderson, D. E. Culler, and D. A. Patterson. A case for NOW (Networks of Workstations). IEEE Micro, 15 (1):54–64, January/February 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ascia:1996:RDV

[ACRV96] Giuseppe Ascia, Vincenzo Catania, Marco Russo, and Lorenzo Vita. Rule-driven VLSI fuzzy processor — implementing a scalable inference processor with analog fuzzy gates. IEEE Micro, 16(3):62–74, May/June 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Agostini:2022:BPS

[ACZ^{+}22] Nicolas Bohm Agostini, Serena Curzel, Jeff Jun Zhang, Ankur Limaye, Cheng Tan, Vinay Amatya, Marco Minutoli, Vito Giovanni Castellana, Joseph Manzano, David Brooks, Gu-Yeon Wei, and Antonino Tumeo. Bridging Python to silicon: The SODA toolchain. IEEE Micro, 42(5): 78–88, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arvind:2000:MSV

[ADC00] Arvind, Anton T. Dahbura, and Alejandro Caro. From Monsoon to StarT-Voyager: University-industry collaboration in research. IEEE Micro, 20(3):75–84, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3075.
pdf;   http://www.computer.
org/micro/mi2000/m3075abs.
htm.

Afek:2010:VTM

[ADF^{+}10] Yehuda Afek, Ulrich Drepper, Pascal Felber, Christof Fetzer, Vincent Gramoli, Michael Hohmuth, Etienne Riviere, Per Stenstrom, Osman Unsal, Walther Maldonado Moreira, Derin Harmanci, Patrick Marlier, Stephan Diestelhorst, Martin Pohlack, Adrian Cristal, Ibrahim Hur, Aleksandar Dragojevic, Rachid Guerraoui, Michal Kapalka, Sasa Tomic, Guy Korland, Nir Shavit, Martin Nowack, and Torvald Riegel. The Velox transactional memory stack. IEEE Micro, 30(5):76– 87, September/October 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ahmad:2020:HMC

[ADJK20] M. Ahmad, H. Dogan, J. A. Joao, and O. Khan. Inhardware moving compute to data model to accelerate thread synchronization on large multicores. IEEE Micro, 40(1):83–92, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

AboElNaga:1982:HAM

[AF82] N. M. Abo El Naga and J. A. Field. A hardware Algol machine. IEEE Micro, 2(4):37– 47, October/December 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Armstrong:1984:FTM

[AF84] Cedric V. W. Armstrong and Eli T. Fathi. A faulttolerant multimicroprocessorbased computer-system for space-based signal-processing. IEEE Micro, 4(6):54–65, November/December 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Alpert:1988:PTO

[AF88] Donald B. Alpert and Michael J.

Flynn. Performance tradeoffs for microprocessor cache memories. IEEE Micro, 8 (4):44–54, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ansari:2010:PFC

[AFGM10] Amin Ansari, Shuguang Feng, Shantanu Gupta, and Scott Mahlke. Putting faulty cores to work. IEEE Micro, 30(6): 36–45, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Akin:2016:HAP

[AFH16] Berkin Akin, Franz Franchetti, and James C. Hoe. HAM-LeT architecture for parallel data reorganization in memory. IEEE Micro, 36(1):14–23, January/February 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/01/mmi2016010014-
abs.html.

Arafa:2019:CLN

[AFK^{+}19] M. Arafa, B. Fahim, S. Kottapalli, A. Kumar, L. P. Looi, S. Mandava, A. Rudoff, I. M. Steiner, B. Valentine, G. Vedaraman, and S. Vora.

Cascade Lake: Next generation Intel Xeon scalable processor. IEEE Micro, 39(2):29– 36, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ardalan:2021:OIC

[AFK^{+}21] S. Ardalan, R. Farjadrad, M. Kuemerle, K. Poulton, S. Subramaniam, and B. Vinnakota. An open inter-chiplet communication link: Bunch of Wires (BoW). IEEE Micro, 41 (1):54–60, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Aschmann:1991:ARP

[AGH^{+}91] Hans-Ruedi R. Aschmann, Niklaus Giger, Elisabeth Hoepli, Peter Janak, and Hubert Kirrmann. Alphorn: a remote procedure call environment for fault-tolerant, heterogeneous, distributed systems. IEEE Micro, 11 (5):16–19, 60–67, September/October 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Avresky:1998:GEI

[AGJL98] Dimiter R. Avresky, Karl E. Grosspietsch, Barry W. Johnson, and Fabrizio Lombardi. Guest Editors’ introduction: Embedded fault-tolerant systems. IEEE Micro, 18(5):8– 11, September/October 1998.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5008.
pdf.

Assmann:1996:CCM

[AH96] Claus Assmann and Andreas Huth. Compiling C on a multiple-stack architecture — introducing Fast, a RISC processor designed for functional language needs, and its C compiler. IEEE Micro, 16(5): 60–67, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Agne:2014:ROS

[AHK^{+}14] Andreas Agne, Markus Happe, Ariane Keller, Enno Lubbers, Bernhard Plattner, Marco Platzner, and Christian Plessl. ReconOS: An operating system approach for reconfigurable computing. IEEE Micro, 34(1):60–71, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Asgari:2019:EER

[AHKY19] B. Asgari, R. Hadidi, H. Kim, and S. Yalamanchili. ERI-DANUS: Efficiently running inference of DNNs using systolic arrays. IEEE Micro, 39(5):46–54, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

America:1990:PCA

[AHO^{+}90] Pierre H. M. America, Ben J. A. Hulshof, Eddy A. M. Odijk, Frans Sijstermans, Rob A. H. van Twist, and Rogier H. H. Wester. Parallel computers for advanced information processing. IEEE Micro, 10(6):12–15, 61–75, November/December 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ajima:2012:TI

[AIH^{+}12] Yuichiro Ajima, Tomohiro Inoue, Shinya Hiramoto, Toshiyuki Shimizu, and Yuzo Takagi. The Tofu Interconnect. IEEE Micro, 32(1):21– 31, January/February 2012. CODEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Aylor:1983:GEI

[AJ83] J. H. Aylor and E. L. Johnson. Guest Editors’ introduction — microcomputing to aid the handicapped. 1. IEEE Micro, 3(3):6–7, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Awan:2020:CPC

[AJC^{+}20] A. A. Awan, A. Jain, C. Chu, H. Subramoni, and D. K. Panda. Communication profiling and characterization of deep-learning workloads on clusters with highperformance interconnects. IEEE Micro, 40(1):35–43, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Aingaran:2015:MON

[AJK^{+}15] Kathirgamar Aingaran, Sumti Jairath, Georgios Konstadinidis, Serena Leung, Paul Loewenstein, Curtis McAllister, Stephen Phillips, Zoran Radovic, Ram Sivaramakrishnan, David Smentek, and Thomas Wicki. M7: Oracle’s next-generation Sparc processor. IEEE Micro, 35(2):36– 45, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020036-
abs.html.

Aylor:1986:SDT

[AJR86] James H. Aylor, Barry W.

Johnson, and Bruce J. Rector. Structured design for testability in semicustom VLSI. IEEE Micro, 6(1):51–58, January/February 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Aono:2000:AWI

[AK00] Fumio Aono and Masayuki Kimura. The AzusA 16way Itanium server. IEEE Micro, 20(5):54–60, September/October 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2000/
pdf/m5054.pdf;   http://
www.computer.org/micro/
mi2000/m5054abs.htm.

Abts:2024:EAI

[AK24] Dennis Abts and John Kim. Enabling artificial intelligence supercomputers with domainspecific networks. IEEE Micro, 44(2):41–49, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Akbari:2018:TAC

[AKAK^{+}18] Omid Akbari, Mehdi Kamal, Ali Afzali-Kusha, Massoud Pedram, and Muhammad Shafique. Toward approximate computing for coarse-grained reconfigurable architectures. IEEE Micro, 38(6):63–72, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08487001-abs.
html.

Akinwande:2018:MMA

[Aki18]  Deji Akinwande.  Mem-

ory, memristors, and atomristors. IEEE Micro, 38(5):50– 52, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

mi/2018/05/mmi2018050050.
html.

Adil:2022:MAE

[AKJF22] Muhammad Adil, Muhammad Khurram Khan, Mona Jamjoom, and Ahmed Farouk. MHADBOR: AI-enabled administrativedistance-based opportunistic load balancing scheme for an agriculture Internet of Things network. IEEE Micro, 42 (1):41–50, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Agarwal:1993:SEP

[AKK^{+}93] A. Agarwal, J. Kubiatowicz, D. Kranz, B. H. Lim, D. Yeung, G. D’Souza, and M. Parkin. Sparcle: an evolutionary processor design for large-scale multiprocessors. IEEE Micro, 13(3):48– 61, May/June 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Akturk:2015:DCD

[AKK15] Ismail Akturk, Nam Sung Kim, and Ulya R. Karpuzcu. Decoupled control and data processing for approximate near-threshold voltage computing. IEEE Micro, 35 (4):70–78, July/August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040070-
abs.html.

Abdelguerfi:1996:GEI

[AKP96] Mahdi Abdelguerfi, Burton S. Kaliski, Jr., and Wayne Patterson. Guest Editors’ introduction: Public-key security systems. IEEE Micro, 16(3): 10–13, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Alcaide:2018:SRC

[AKT^{+}18] Sergi Alcaide, Leonidas Kosmidis, Hamid Tabani, Carles Hernandez, Jaume Abella, and Francisco J. Cazorla. Safety-related challenges and opportunities for GPUs in the automotive domain. IEEE Micro, 38(6):46–55, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/06/08486955-abs.
                                        html.

Albonesi:2004:GEI

[Alb04] David H. Albonesi. Guest Editor’s introduction: Micro’s top picks from microarchitecture conferences. IEEE Micro, 24(6):8–9, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/06/m6008.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2004/06/m6008.
                                        htm.

Albonesi:2007:ECMd

[Alb07a] David H. Albonesi. Editorin-Chief’s message: Mixing it up. IEEE Micro, 27(4):3– 4, July/August 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/04/mmi2007040003.
pdf.

Albonesi:2007:ECMc

[Alb07b] David H. Albonesi. Editorin-Chief’s message: More hot stuff. IEEE Micro, 27(3): 4–5, May/June 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/03/mmi2007030004.
pdf.

Albonesi:2007:ECMe

[Alb07c] David H. Albonesi. Editorin-Chief’s message: Productive and healthy debate. IEEE Micro, 27(6):6, November/December 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/06/mmi2007060006.
pdf.

Albonesi:2007:ECMa

[Alb07d] David H. Albonesi. Editorin-Chief’s message: Standing on solid ground. IEEE Micro, 27(1):5–6, January/ February 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2007/01/m1005.pdf.

Albonesi:2007:ECMb

[Alb07e] David H. Albonesi. Editor in Chief’s message: Truly “hot” chips—do we still care? IEEE Micro, 27(2): 4–5, March/April 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2007/02/m2004.pdf.

Albonesi:2008:ECC

[Alb08] David H. Albonesi. From the Editor in Chief: Changes ahead. IEEE Micro, 28(5): 4, September/October 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Albonesi:2009:ECW

[Alb09] David H. Albonesi. From the Editor in Chief: Welcome aboard. IEEE Micro, 29(5):2– 5, September/October 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Albonesi:2010:ECF

[Alb10a] David H. Albonesi. From the Editor in Chief: Future directions in computer architecture research. IEEE Micro, 30 (3):5, May/June 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Albonesi:2010:MF

[Alb10b] David H. Albonesi. Moving forward. IEEE Micro, 30 (6):4–5, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Avresky:2001:GEI

[ALGJ01] Dimiter R. Avresky, Fabrizio Lombardi, Karl E. Grosspietsch, and Barry W. Johnson. Guest Editors’ introduction: Fault-tolerant embedded systems. IEEE Micro, 21(5):12– 15, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5012abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m5012.
pdf.

Allison:1981:SRP

[All81] Andrew A. Allison. Statusreport on the P896 backplane bus. IEEE Micro, 1(1):67–69, 71–82, January/March 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Allison:1984:MPS

[All84] Andrew Allison. Microcomputer peripherals — status and trends. IEEE Micro, 4 (5):9–17, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Allison:1986:BW

[All86a]  A. Allison.  Bus wars.

IEEE Micro, 6(6):5, November/December 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Allison:1986:ISD

[All86b] A. Allison. IEEE standards during the great bus wars

— another view. IEEE Micro, 6(6):82–83, November/December 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Alsup:1990:MFA

[Als90]  Mitch Alsup.  Motorola’s

88000 family architecture. IEEE Micro, 10(3):48–66, May/June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Alt:1998:DEI

[Alt98] Paul M. Alt. Displays for electronic imaging. IEEE Micro, 18(6):42–53, November/December 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1998/
                                        pdf/m6042.pdf;   http://

www.computer.org/micro/
mi1998/m6042abs.htm.

Altman:2011:ECB

[Alt11a] Erik R. Altman. From the Editor-in-Chief: Big chips and beyond. IEEE Micro, 31(4): 2, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2011:ECC

[Alt11b] Erik R. Altman. From the Editor-in-Chief: CPUs and GPUs: Who owns the future? IEEE Micro, 31(5):2– 3, September/October 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2011:HCR

[Alt11c] Erik R. Altman. Hot Chips and remembering a pioneer. IEEE Micro, 31(2):3, March/ April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2011:NBC

[Alt11d] Erik R. Altman. New blood, cool chips, and heterogeneous designs. IEEE Micro, 31(6):2– 3, November/December 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2011:SPV

[Alt11e] Erik R. Altman. A solid past, a vital future. IEEE Micro, 31(1):4–5, January/February 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2011:VLS

[Alt11f] Erik R. Altman. Very largescale systems and some history. IEEE Micro, 31(3): 2–3, May/June 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2012:HIH

[Alt12a] Erik R. Altman. Hot Interconnects and hot topics. IEEE Micro, 32(1):2–3, January/February 2012. CO-DEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2012:ME

[Alt12b] Erik R. Altman. Micro evolution. IEEE Micro, 32(2): 2, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2012:OCH

[Alt12c] Erik R. Altman. The odd couple: Hardware and software. IEEE Micro, 32(4): 2, July/August 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2012:PEA

[Alt12d] Erik R. Altman. Powerand energy-aware computing. IEEE Micro, 32(5):

2, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2012:TPC

[Alt12e] Erik R. Altman. Top picks, columnists, and artists. IEEE Micro, 32(3):2, May/June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2012:WWM

[Alt12f] Erik R. Altman. Which way microarchitecture? IEEE Micro, 32(6):2, November/ December 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2013:CCM

[Alt13a] Erik R. Altman. Cool chips, mobile devices, memory, and IEEE Micro going digital. IEEE Micro, 33(6):2, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Altman:2013:DSD

[Alt13b] Erik R. Altman. Dark silicon and dangerous predictions. IEEE Micro, 33(5):4– 5, September/October 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2013:ECT

[Alt13c] Erik R. Altman. From the Editor in Chief: Ten years of top picks. IEEE Micro, 33 (3):2, May/June 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2013:HCI

[Alt13d] Erik R. Altman. Hot chips and the incomplete job of exploiting them. IEEE Micro, 33(2):4–5, March/April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2013:OIT

[Alt13e] Erik R. Altman. Optical interconnects and their implications. IEEE Micro, 33(1):2, January/February 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2013:RTI

[Alt13f] Erik R. Altman. Reliability, theme issues, and plagiarism. IEEE Micro, 33(4): 2, July/August 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2014:BDD

[Alt14a] Erik R. Altman. Big data and democratization [Editorial]. IEEE Micro, 34(4):2– 3, July/August 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040002-
abs.html.

Altman:2014:HCG

[Alt14b]  Erik R. Altman.  Harsh

chips, but a grateful good-bye. IEEE Micro, 34(6):4, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060004-
abs.html.

Altman:2014:HCO

[Alt14c] Erik R. Altman. Hot chips and other themes. IEEE Micro, 34(2):2–3, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Altman:2014:PHS

[Alt14d] Erik R. Altman. Patents and high-speed datacenter interconnects. IEEE Micro, 34(5):4–5, September/ October 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050004-
abs.html.

Altman:2014:RCI

[Alt14e] Erik R. Altman. Reconfigurable computing, 3D integration, and recognizing leaders in our field. IEEE Micro, 34(1):2–3, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Altman:2014:TP

[Alt14f] Erik R. Altman. Top picks from 2013. IEEE Micro, 34(3):

2–3, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Amirtharajah:2008:GEI

[AM08] Raj Amirtharajah and John Mashey. Guest Editors’ introduction: Hot Chips 19. IEEE Micro, 28(2):7– 9, March/April 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2008/02/mmi2008020007.
pdf.

Asghari-Moghaddam:2016:NDA

[AMFFM^{+}16] Hadi Asghari-Moghaddam, Amin Farmahini-Farahani, Katherine Morrow, Jung Ho Ahn, and Nam Sung Kim. Near-DRAM acceleration with single-ISA heterogeneous processing in standard memory modules. IEEE Micro, 36(1):24–34, January/ February 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/01/mmi2016010024-
abs.html.

Ahmad:2017:ESS

[AMK17] Masab Ahmad, Christopher J. Michael, and Omer Khan. Efficient situational scheduling of graph workloads on singlechip multicores and GPUs. IEEE Micro, 37(1):30–40, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010030-
abs.html.

Abel:2003:FTP

[AML^{+}03] François Abel, Cyriel Minkenberg, Ronald P. Luijten, Mitchell Gusat, and Ilias Iliadis. A four-terabit packet switch supporting long roundtrip times. IEEE Micro, 23 (1):10–24, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1010.
pdf;   http://www.computer.
org/micro/mi2003/m1010abs.
htm.

Anguita:2005:MOE

[AML05] Mancia Anguita and J. Manuel Martinez-Lechado. MP3 optimization exploiting processor architecture and using better algorithms. IEEE Micro, 25(3):81–92, May/June 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Agarwal:2006:LPA

[AMR^{+}06] Amit Agarwal, Saibal Mukhopadhyay, Arijit Raychowdhury, Kaushik Roy, and Chris H. Kim. Leakage power analysis and reduction for nanoscale circuits. IEEE Micro, 26 (2):68–80, March/April 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Akhbarizadeh:2005:PSS

[ANC05] Mohammad J. Akhbarizadeh, Mehrdad Nourani, and Cyrus D. Cantrell. Prefix segregation scheme for a TCAM-based IP forwarding engine. IEEE Micro, 25(4):48–63, July/August 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Andrews:1982:MMS

[And82a] M. Andrews. Mathematical microprocessor software: a sqrt(x) comparison. IEEE Micro, 2(3):63–79, July/ September 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Andrews:1982:SRX

[And82b] M. Andrews. Square-root-X comparison — new results discovered — reply. IEEE Micro, 2(4):5–6, October/December 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Andrews:2014:OSR

[And14] David Andrews. Operating systems research for reconfigurable computing. IEEE Micro, 34(1):54–58, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Angeniol:1990:PEI

[Ang90] Bernard Angeniol. Pygmalion: ESPRIT II Project 2059, neurocomputing. IEEE Micro, 10(6):28–31, 99–102, November/December 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Andrews:2004:PMH

[ANJ^{+}04] David Andrews, Douglas Niehaus, Razali Jidin, Michael Finley, Wesley Peck, Michael Frisbie, Jorge Ortiz, Ed Komp, and Peter Ashenden. Programming models for hybrid FPGA–CPU computational components: a missing link. IEEE Micro, 24(4):42– 53, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4042abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/04/m4042.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/04/
m4042.pdf.

Arora:2012:RRC

[ANM^{+}12] Manish Arora, Siddhartha Nath, Subhra Mazumdar, Scott B. Baden, and Dean M. Tullsen. Redefining the role of the CPU in the era of CPU–GPU integration. IEEE Micro, 32(6):4–16, November/December 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1981:PSE

[Ano81]  Anonymous.  A proposed

standard for extending highlevel languages for microprocessors. IEEE Micro, 1(2): 70–75, April/June 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1983:MUF

[Ano83] Anonymous. The microprocessor universal format for object modules — proposed standard. IEEE Micro, 3 (4):48–66, July/August 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1984:PEB

[Ano84] Anonymous. P959 I/O expansion bus — proposed standard. IEEE Micro, 4(3):33– 54, May/June 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1985:RIS

[Ano85] Anonymous. Reader interest survey. IEEE Micro, 5(5): 85, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1986:M

[Ano86a] Anonymous. MicroCourses. IEEE Micro, 6(3):78, May/ June 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1986:RIS

[Ano86b] Anonymous. Reader interest survey. IEEE Micro, 6(6):80, November/December 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1987:CLN

[Ano87a] Anonymous. Chips — lasers = new-generation wave devices systems. IEEE Micro, 7(1):81, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1987:HMP

[Ano87b] Anonymous. Hardening material protects computers. IEEE Micro, 7(1):82, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1987:HD

[Ano87c] Anonymous. Homebrewers disband. IEEE Micro, 7(1):82, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1987:HDG

[Ano87d] Anonymous. Hughes develops 18-GHz IC. IEEE Micro, 7(1):80–83, January/February 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1987:IFT

[Ano87e] Anonymous. ISDN field-tests in process. IEEE Micro, 7(1): 81, January/February 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1987:MNT

[Ano87f] Anonymous. Map network tested. IEEE Micro, 7(1):82, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1987:WSS

[Ano87g] Anonymous. Wafer sharing — slicing costs while experimenting. IEEE Micro, 7(1):82, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:CHR

[Ano88a] Anonymous. Computer history, rev-1. IEEE Micro, 8(4): 84, July/August 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:CG

[Ano88b] Anonymous. Crystal growth. IEEE Micro, 8(4):83, July/ August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:DOP

[Ano88c] Anonymous. Deskpro 386-20 outperforms PS/2 Model-80. IEEE Micro, 8(4):82, July/August 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:DYE

[Ano88d] Anonymous. Did you ever think you’d like a robot? IEEE Micro, 8(4):83–84, July/August 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:ESO

[Ano88e] Anonymous. ESDI standard open for comment. IEEE Micro, 8(4):84, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:OCP

[Ano88f] Anonymous. Organic computers a possibility. IEEE Micro, 8(4):83, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:TRU

[Ano88g] Anonymous. Technology research — ultrafast devices, X-ray-lithography, and micromachines. IEEE Micro, 8(3): 94–95, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1988:TCP

[Ano88h] Anonymous. Tiny computer promises massive speeds. IEEE Micro, 8(4):82, July/ August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1989:DCB

[Ano89] Anonymous. Diamond — a chip’s best friend. IEEE Micro, 9(2):96–??, March/April 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1991:IPR

[Ano91a] Anonymous. Improving the product — reply. IEEE Micro, 11(2):2, March/April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1991:MNP

[Ano91b] Anonymous. Micro news: PLA copyright infringement. IEEE Micro, 11(6):4–??, November/December 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1991:PSF

[Ano91c] Anonymous. PILOT, SCI, Futurebus+. IEEE Micro, 11(4): 40–??, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1992:AFL

[Ano92a] Anonymous. Atomic fountains; laser tweezers; optical molasses. IEEE Micro, 12(4): 88–??, July/August 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1992:CCT

[Ano92b] Anonymous. Chips; CAD tools; signal processing hardware/software. IEEE Micro, 12(4):90–??, July/August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1992:DCS

[Ano92c]  Anonymous.  Databases;

C++; science writing. IEEE Micro, 12(4):6–??, July/ August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1992:ME

[Ano92d] Anonymous. Meet the experts. IEEE Micro, 12(3): 72–??, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1992:NMS

[Ano92e]  Anonymous.  No more

software reverse engineering? IEEE Micro, 12(3): 3–??, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1992:OET

[Ano92f] Anonymous. Object encyclopedia technology. IEEE Micro, 12(3):74–??, May/June 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1993:PC

[Ano93] Anonymous. Programming and compressing. IEEE Micro, 13(4):61–??, July/August 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1994:E

[Ano94a] Anonymous. Eudora. IEEE Micro, 14(6):5–6, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1994:HYC

[Ano94b] Anonymous. How’s your computing life going to evolve in the near future? you might consider electronic villages and DTP tools that release engineers and scientists from dependency on publishing professionals. IEEE Micro, 14(2):5–??, March/April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1994:IYW

[Ano94c] Anonymous. If you want to learn about computer organization, here’s one book you should read, especially if you’re planning to teach a course on the subject. also, what’s happening to conferences? IEEE Micro, 14(3): 2–??, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1994:MVS

[Ano94d]  Anonymous.  Micro view:

Spiritual computing. IEEE Micro, 14(6):80–??, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1995:CAA

[Ano95a] Anonymous. Call for articles — August 1996 special issue on media processing: native and digital signal processing approaches. IEEE Micro, 15(6):60–??, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1995:MNS

[Ano95b] Anonymous. Micro news: 3D stacked memory, network attack scanner. IEEE Micro, 15(3):75–??, May/June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1995:MRI

[Ano95c] Anonymous. Micro review:

The Internet, WWW, and tools to get you there. IEEE Micro, 15(3):72–??, May/June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1995:MVC

[Ano95d]  Anonymous.  Micro view:

Are claims made for fuzzy logic preposterous or true? IEEE Micro, 15(4):80–??, July/August 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:AIV

[Ano96a] Anonymous. 1996 annual index — volume 16 — this section provides a complete author and subject listing. IEEE Micro, 16(6):68–73, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:CLC

[Ano96b] Anonymous. China legend chooses K5. IEEE Micro, 16 (6):80, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:CST

[Ano96c] Anonymous. Computer Society turns 50. IEEE Micro, 16 (1):4, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:DOC

[Ano96d]  Anonymous.  DVDs to

overtake CDs at millennium. IEEE Micro, 16(6):80, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:ELB

[Ano96e] Anonymous. Electronic library, bookstore. IEEE Micro, 16(5):76, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:ESP

[Ano96f]  Anonymous.  Exponential

speeds for PowerPC. IEEE Micro, 16(6):3, November/ December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:HPW

[Ano96g] Anonymous. Hotmetal-Pro-3.0 for Windows. IEEE Micro, 16(5):9, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:IUN

[Ano96h] Anonymous. IBM unveils network computer. IEEE Micro, 16(5):76, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:JA

[Ano96i] Anonymous. Javastation arrives. IEEE Micro, 16(6):80, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:LCP

[Ano96j] Anonymous. Longtime CS publisher retires. IEEE Micro, 16(6):80, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:MNM

[Ano96k] Anonymous. Micro news: 533-MHz processor, laurels for the 8051, CS publisher retires. IEEE Micro, 16(6): 3, 80, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:MNE

[Ano96l] Anonymous. Micro news: Extreme-ultraviolet lithography 2 steps closer. IEEE Micro, 16(5):2–??, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:MNV

[Ano96m] Anonymous. Micro news: Virtual socket interface alliance. IEEE Micro, 16(5):2– ??, September/October 1996.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:MVC

[Ano96n]  Anonymous.  Micro view:

Can designers resist the allure of sexy, high-visibility apps? IEEE Micro, 16(5): 80, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:NDT

[Ano96o] Anonymous. New 3D display technology. IEEE Micro, 16 (5):4–??, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:NCA

[Ano96p] Anonymous. News — celebrating the 25th anniversary of the microprocessor. IEEE Micro, 16(1):3, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:SCG

[Ano96q] Anonymous. Smart cards get wise. IEEE Micro, 16(1):4, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:VHB

[Ano96r] Anonymous. Verilog HDL becomes IEEE standard.

IEEE Micro, 16(1):4, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:VSI

[Ano96s] Anonymous. Virtual socket interface alliance. IEEE Micro, 16(5):4, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:VSV

[Ano96t] Anonymous. Visual-Slickedit-V2.0 for Windows 95. IEEE Micro, 16(5):9–??, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:WNS

[Ano96u] Anonymous. Will the net selfdestruct. IEEE Micro, 16(6): 3–??, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:WLS

[Ano96v] Anonymous. Wireless LANs standardization. IEEE Micro, 16(5):76, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:AIC

[Ano97a] Anonymous. 1997 annual index-complete author and subject listings. IEEE Micro, 17(6):75–??, November/December 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1997/
pdf/m6075.pdf.

Anonymous:1997:EC

[Ano97b] Anonymous. 1998 editorial calendar. IEEE Micro, 17(5): 11–??, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5011.
pdf.

Anonymous:1997:KMC

[Ano97c] Anonymous. 56-Kbps modems to come. IEEE Micro, 17(1):6, January/February 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:AES

[Ano97d] Anonymous. Advanced Encryption Standard. IEEE Micro, 17(1):6, January/ February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:AEJ

[Ano97e] Anonymous. All-electronic journal. IEEE Micro, 17(4): 79, July/August 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:BSE

[Ano97f]  Anonymous.  Building a

single-electron diode. IEEE Micro, 17(5):6, September/October 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:CLP

[Ano97g] Anonymous. Consuming less power. IEEE Micro, 17(5): 72, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:IIS

[Ano97h] Anonymous. Integrating image sensor, signal conditioner. IEEE Micro, 17(5): 72, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:IVD

[Ano97i] Anonymous. Intel vs DEC. IEEE Micro, 17(3):6, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://pascal.computer.
                                        org/mi/books/mi1997/pdf/
                                        m3005.pdf.

Anonymous:1997:MSD

[Ano97j] Anonymous. Measuring submicron dimensions. IEEE Micro, 17(4):79, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:MNM

[Ano97k] Anonymous. Micro news: a monopoly on naming. IEEE Micro, 17(4):2, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4002.
pdf.

Anonymous:1997:MNA

[Ano97l] Anonymous. Micro news:

ACM’s 50th. IEEE Micro, 17 (2):3, March/April 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:MNL

[Ano97m] Anonymous. Micro news:

Lattices — a new key for cryptography. IEEE Micro, 17(4):2, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4002.
pdf.

Anonymous:1997:MRB

[Ano97n] Anonymous. Micro review:

Bending computers to our will. IEEE Micro, 17(5):74– 76, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5074.
pdf.

Anonymous:1997:MRD

[Ano97o] Anonymous. Micro review: Deep and ambitious books for lazy summer days. IEEE Micro, 17(4):7–8, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4007.
pdf.

Anonymous:1997:MRJ

[Ano97p] Anonymous. Micro review: Java, SGML: Destabilizing forces for companies and careers? IEEE Micro, 17(3):3–5, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3003.
pdf.

Anonymous:1997:MRO

[Ano97q] Anonymous. Micro review: Only the paranoid survive; recent Java books. IEEE Micro, 17(2):4–77, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:MRM

[Ano97r]  Anonymous.  Micro re-

view: The Microsoft way; object-relational databases.

IEEE Micro, 17(1):4–6, January/February 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:MSA

[Ano97s] Anonymous. Micro standards:

Active projects list. IEEE Micro, 17(2):70–75, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:MVV

[Ano97t] Anonymous. Micro view: VSI alliance: Selfish interests the key? IEEE Micro, 17(1):80– ??, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:MSB

[Ano97u] Anonymous. Motorola ships 2 billionth 68HC05. IEEE Micro, 17(3):6, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:NPB

[Ano97v] Anonymous. New process battles warpage. IEEE Micro, 17(1):78, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:OI

[Ano97w] Anonymous. Optimizing the IA-64. IEEE Micro, 17(5): 6, September/October 1997.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:PSa

[Ano97x] Anonymous. Product summary. IEEE Micro, 17 (4):79–??, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4079.
pdf.

Anonymous:1997:PSb

[Ano97y] Anonymous. Product summary. IEEE Micro, 17(5):77– ??, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5077.
pdf.

Anonymous:1997:QTM

[Ano97z]  Anonymous.  Quick-turn

MCM prototypes. IEEE Micro, 17(3):6, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:SPD

[Ano97-27] Anonymous. Smart phones debut in Japan. IEEE Micro, 17(2):2–3, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:SBM

[Ano97-28] Anonymous. Spec benchmark for MPEG-2. IEEE Micro, 17(4):79, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:SF

[Ano97-29] Anonymous. Special features.

IEEE Micro, 17(6):??, November/December 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:SEU

[Ano97-30] Anonymous. Studying extremeultraviolet. IEEE Micro, 17(5):6, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:SMP

[Ano97-31] Anonymous. Support for 533-MHz processor. IEEE Micro, 17(1):78, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:TDI

[Ano97-32] Anonymous. TI drops off Intel bandwagon. IEEE Micro, 17 (1):6, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:XRR

[Ano97-33] Anonymous. X-ray resolution times 10. IEEE Micro, 17(1): 6–??, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:YC

[Ano97-34] Anonymous. Year 2000 compliance. IEEE Micro, 17(5):6– ??, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:AIC

[Ano98a]  Anonymous.  1998 an-

nual index — complete author and subject listings. IEEE Micro, 18(6):77–??, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6077.
pdf.

Anonymous:1998:AG

[Ano98b] Anonymous. Author guide. IEEE Micro, 18(2):85–??, March/April 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:AGD

[Ano98c] Anonymous. Author guide: Details for submissions to IEEE Micro. IEEE Micro, 18(6):85–??, November/December 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m6085.pdf.

Anonymous:1998:CAa

[Ano98d] Anonymous. Call for articles. IEEE Micro, 18 (2):13–??, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2013.
pdf.

Anonymous:1998:CAb

[Ano98e] Anonymous. Call for articles. IEEE Micro, 18(6): 2–??, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6002.
pdf;   http://dlib.computer.
org/mi/books/mi1998/pdf/
m6011.pdf;   http://dlib.
computer.org/mi/books/mi1998/
pdf/m6031.pdf.

Anonymous:1998:CIE

[Ano98f] Anonymous. Collaboration for Imec and Ericsson. IEEE Micro, 18(4):83, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:CAC

[Ano98g] Anonymous. Cyrix adds to CPU line. IEEE Micro, 18 (3):76, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:EIT

[Ano98h] Anonymous. EDA industry thrives. IEEE Micro, 18(3): 77, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:EC

[Ano98i] Anonymous. Editorial calendar. IEEE Micro, 18 (3):84–??, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m3084.
                                        pdf.

Anonymous:1998:EBI

[Ano98j] Anonymous. Electron behavior investigated. IEEE Micro, 18(3):76, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:HTE

[Ano98k] Anonymous. High-tech entrepreneurs. IEEE Micro, 18 (3):77, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:INP

[Ano98l] Anonymous. IBM’s new PC lineup. IEEE Micro, 18(3): 76–77, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:I

[Ano98m] Anonymous. The iMac. IEEE Micro, 18(3):76, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:PE

[Ano98n] Anonymous. Is packaging everything? IEEE Micro, 18 (3):76, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:O

[Ano98o]  Anonymous.  It’s official.

IEEE Micro, 18(3):76, May/ June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:JPC

[Ano98p] Anonymous. Joint project certificate program. IEEE Micro, 18(4):83, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:AED

[Ano98q] Anonymous. Los Alamos engineers develop solvent. IEEE Micro, 18(4):83, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:MX

[Ano98r]  Anonymous.  Mac OS-X.

IEEE Micro, 18(3):76, May/ June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:MNG

[Ano98s] Anonymous. Micro news: 1-GHz microprocessors. IEEE Micro, 18(2):2, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2002.
pdf.

Anonymous:1998:MND

[Ano98t] Anonymous. Micro news: Did the PC industry get ahead of itself? IEEE Micro, 18(3):75, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3075.
pdf.

Anonymous:1998:MNE

[Ano98u] Anonymous. Micro news: ECIX adoption. IEEE Micro, 18(2):2–3, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi1998/pdf/m2002.
pdf.

Anonymous:1998:MNI

[Ano98v] Anonymous. Micro news: Intergraph wins Federal Court order. IEEE Micro, 18(3): 75, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m3075.pdf.

Anonymous:1998:MNM

[Ano98w] Anonymous. Micro news:

Merced update. IEEE Micro, 18(2):2, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2002.
pdf.

Anonymous:1998:MNN

[Ano98x] Anonymous. Micro news:

National standards strategy proposed. IEEE Micro, 18(3):75, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3075.
pdf.

Anonymous:1998:MNP

[Ano98y] Anonymous. Micro news:

Programmable digital camera. IEEE Micro, 18(3):75, May/ June 1998. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3075.
pdf.

Anonymous:1998:MRM

[Ano98z] Anonymous. Micro review: Metaclasses; intellectual property; understanding the new media; Adobe Photoshop 5. IEEE Micro, 18(6):9–??, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6009.
pdf.

Anonymous:1998:M

[Ano98-27]  Anonymous.  Microlasers.

IEEE Micro, 18(4):83, July/ August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:MS

[Ano98-28] Anonymous. Multithreaded system. IEEE Micro, 18(3): 76, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:NPDb

[Ano98-29]  Anonymous.  New prod-

ucts: Design services, portables, SBCs, software, memories. IEEE Micro, 18(4): 84–??, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4084.
pdf.

Anonymous:1998:NPDa

[Ano98-30]  Anonymous.  New prod-

ucts: Design tools, portables, and DSPs. IEEE Micro, 18(3):81–??, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3081.
pdf.

Anonymous:1998:NPDc

[Ano98-31] Anonymous. New products:

Design tools; portables, online tools. IEEE Micro, 18(6): 4–??, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6004.
pdf.

Anonymous:1998:NCC

[Ano98-32] Anonymous. News: Carbon computers possible? IEEE Micro, 18(4):82, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4082.
pdf.

Anonymous:1998:NIN

[Ano98-33]  Anonymous.  News: In-

tel’s new Pentium-II processor. IEEE Micro, 18 (4):82, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m4082.
                                        pdf.

Anonymous:1998:NNP

[Ano98-34] Anonymous. News: New packaging from Siemens. IEEE Micro, 18(4):82, July/ August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1998/
                                        pdf/m4082.pdf.

Anonymous:1998:NN

[Ano98-35] Anonymous. News on notebooks. IEEE Micro, 18(2): 3, March/April 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:NSU

[Ano98-36] Anonymous. News: Singleunit RF power delivery. IEEE Micro, 18(4):82, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m4082.
                                        pdf.

Anonymous:1998:OTJ

[Ano98-37] Anonymous. Online technical journal. IEEE Micro, 18 (3):77, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:OMD

[Ano98-38] Anonymous. Options for Mac designers. IEEE Micro, 18(4): 83, July/August 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:PSa

[Ano98-39] Anonymous. Product summary. IEEE Micro, 18 (2):87–??, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2087.
pdf.

Anonymous:1998:PSb

[Ano98-40] Anonymous. Product summary. IEEE Micro, 18 (3):85–??, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3085.
pdf.

Anonymous:1998:PSc

[Ano98-41] Anonymous. Product summary. IEEE Micro, 18 (4):86–??, July/August 1998.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4086.
pdf.

Anonymous:1998:PSd

[Ano98-42] Anonymous. Product summary. IEEE Micro, 18(5):86– ??, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5086.
pdf.

Anonymous:1998:PSe

[Ano98-43] Anonymous. Product summary. IEEE Micro, 18 (6):88, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6088.
pdf.

Anonymous:1998:SC

[Ano98-44] Anonymous. Systems on a chip. IEEE Micro, 18(4): 83, July/August 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1999:BBD

[Ano99a] Anonymous. Broad-band deployment act. IEEE Micro, 19(4):81, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1999:CAa

[Ano99b] Anonymous. Call for articles. IEEE Micro, 19(1): 29, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1029.
pdf;   http://dlib.computer.
org/mi/books/mi1999/pdf/
m1049.pdf.

Anonymous:1999:CAb

[Ano99c] Anonymous. Call for articles.

IEEE Micro, 19(3):72, May/ June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3072.
pdf.

Anonymous:1999:EC

[Ano99d] Anonymous. Editorial calendar. IEEE Micro, 19(1): 88, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1088.
pdf.

Anonymous:1999:IMA

[Ano99e] Anonymous. IEEE Micro 1999 annual index. IEEE Micro, 19(6):75–83, November/December 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1999/
                                        pdf/m6075.pdf.

Anonymous:1999:MS

[Ano99f]  Anonymous.  Managing

SANs. IEEE Micro, 19(4): 81, July/August 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1999:MNA

[Ano99g] Anonymous. Micro news: AMD announces new Athlon. IEEE Micro, 19(6):2, November/December 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1999/
                                        pdf/m6002.pdf.

Anonymous:1999:MNC

[Ano99h] Anonymous. Micro news: Creating virtual sculpture. IEEE Micro, 19(1):10–11, January/February 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1999/
                                        pdf/m1010.pdf.

Anonymous:1999:MNE

[Ano99i] Anonymous. Micro news: Electronic paper resurfaces. IEEE Micro, 19(6):3, November/December 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.

computer.org/mi/books/mi1999/
pdf/m6002.pdf.

Anonymous:1999:MNF

[Ano99j] Anonymous. Micro news:

FTC holds off on Internet privacy laws. IEEE Micro, 19(4):80–81, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4080.
pdf.

Anonymous:1999:MNH

[Ano99k] Anonymous. Micro news:

Hyundai merges with LG Semicon. IEEE Micro, 19 (6):3, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6002.
pdf.

Anonymous:1999:MNIb

[Ano99l] Anonymous. Micro news: Intel delays Coppermine. IEEE Micro, 19(4):81, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4080.
pdf.

Anonymous:1999:MNIc

[Ano99m] Anonymous. Micro news: Intel unveils Itanium processor. IEEE Micro, 19(6): 2, November/December 1999.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6002.
pdf.

Anonymous:1999:MNIa

[Ano99n] Anonymous. Micro news: Internet backbone support/ privacy laws. IEEE Micro, 19(4):80, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4080.
pdf.

Anonymous:1999:MNMa

[Ano99o] Anonymous. Micro news: Microelectronics research in Europe. IEEE Micro, 19(1): 10, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1010.
pdf.

Anonymous:1999:MNMb

[Ano99p] Anonymous. Micro news: Mobile computing without PCs/ laptops; Intel delays Coppermine; Internet backbone support/privacy laws. IEEE Micro, 19(4):80, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi1999/pdf/m4080.
pdf.

Anonymous:1999:MNNb

[Ano99q] Anonymous. Micro news:

NAE prize awarded [to John MacChesney, Charles Kao, and Robert Maurer]. IEEE Micro, 19(6):2– 3, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6002.
pdf.

Anonymous:1999:MNNa

[Ano99r] Anonymous. Micro news:

NIST project competition. IEEE Micro, 19(1):10, January/February 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1999/
pdf/m1010.pdf.

Anonymous:1999:MNO

[Ano99s] Anonymous. Micro news:

One-millionth copper chip shipped. IEEE Micro, 19 (6):2, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6002.
pdf.

Anonymous:1999:MNPa

[Ano99t] Anonymous. Micro news:

Photobit secures broad cameraon-a-chip patent. IEEE Micro, 19(1):10, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1010.
pdf.

Anonymous:1999:MNR

[Ano99u] Anonymous. Micro news: R&D projects awarded. IEEE Micro, 19(6):3, November/ December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1999/
pdf/m6002.pdf.

Anonymous:1999:MNS

[Ano99v] Anonymous. Micro news: System-on-a-chip design. IEEE Micro, 19(1):10–11, January/February 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1999/
pdf/m1010.pdf.

Anonymous:1999:MNPb

[Ano99w] Anonymous. Micro news: The provocative ISSCC99; Pentium III and “Intel Inside”; semiconductor IP service; open industry-standard language; a virtual reading room; IEEE 1394 IC protects digital content. IEEE Micro, 19(2):9, 82, March/ April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2009.
pdf.

Anonymous:1999:MRA

[Ano99x] Anonymous. Micro review:

Adobe Acrobat 4. IEEE Micro, 19(5):11, September/October 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1999:MRB

[Ano99y] Anonymous. Micro review:

Bringing up the rear. IEEE Micro, 19(3):3–4, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3003.
pdf.

Anonymous:1999:NP

[Ano99z]  Anonymous.  New prod-

ucts. IEEE Micro, 19(5):84– 86, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5084.
pdf.

Anonymous:1999:NPD

[Ano99-27]  Anonymous.  New prod-

ucts: Development tools; security solutions; multimedia. IEEE Micro, 19(1):84– 85, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m1084.
                                        pdf.

Anonymous:1999:PII

[Ano99-28] Anonymous. Pentium-III and Intel Inside. IEEE Micro, 19 (2):9, 82, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1999:PSa

[Ano99-29] Anonymous. Product summary. IEEE Micro, 19(1):86– 87, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m1086.
                                        pdf.

Anonymous:1999:PSb

[Ano99-30] Anonymous. Product summary. IEEE Micro, 19 (2):87, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m2087.
                                        pdf.

Anonymous:1999:PSc

[Ano99-31] Anonymous. Product summary. IEEE Micro, 19(4): 87–88, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4087.
pdf.

Anonymous:1999:PSd

[Ano99-32] Anonymous. Product summary. IEEE Micro, 19(5): 87, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5087.
pdf.

Anonymous:1999:SSW

[Ano99-33] Anonymous. Single-sourcing Webworks Publisher-2000 for Windows. IEEE Micro, 19(4): 85, July/August 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2000:AIV

[Ano00a] Anonymous. Annual index — volume 20 author and subject listings. IEEE Micro, 20(6): 85–95, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6085.
pdf.

Anonymous:2000:BDA

[Ano00b] Anonymous. Biomechanical discovery affects mobile applications, robots. IEEE Micro, 20(3):87, May/June 2000.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2000:CP

[Ano00c] Anonymous. Call for papers. IEEE Micro, 20(2):47, March/ April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m2047.
                                        pdf.

Anonymous:2000:HNW

[Ano00d] Anonymous. Hard to navigate Web. IEEE Micro, 20 (3):87, May/June 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2000:IME

[Ano00e] Anonymous. IEEE Micro editorial calendar. IEEE Micro, 20(2):70, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m2070.
                                        pdf.

Anonymous:2000:MB

[Ano00f] Anonymous. Micro bits. IEEE Micro, 20(4):3, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2000:MNM

[Ano00g] Anonymous. Micro news: Motorola expands IP and SOC efforts; market benefits again; customized VLIW cores proposed; building the world’s fastest supercomputer; advancing wireless use. IEEE Micro, 20(1):4– 5, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1004.
pdf.

Anonymous:2000:NBU

[Ano00h] Anonymous. New benchmark for Unigraphics V15. IEEE Micro, 20(3):86, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2000:NHI

[Ano00i] Anonymous. News: Hot interconnects, hot chips, InfiniBand standard, semiconductor milestones. IEEE Micro, 20(6):2–3, November/December 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2000/
pdf/m6002.pdf.

Anonymous:2000:PSa

[Ano00j] Anonymous. Product summary. IEEE Micro, 20(3): 88, May/June 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.

                                        computer.org/mi/books/mi2000/
                                        pdf/m3088.pdf.

Anonymous:2000:PSb

[Ano00k] Anonymous. Product summary. IEEE Micro, 20(5): 88, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m5088.
                                        pdf.

Anonymous:2000:PSc

[Ano00l] Anonymous. Product summary. IEEE Micro, 20 (6):96, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m6096.
                                        pdf.

Anonymous:2000:TSJ

[Ano00m] Anonymous. Tool set for the Java Card platform. IEEE Micro, 20(3):87, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2000:UR

[Ano00n] Anonymous. Untitled — reply. IEEE Micro, 20(1):3, January/February 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2000:WAG

[Ano00o] Anonymous. Wireless applications grow. IEEE Micro, 20 (3):86, May/June 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2001:COC

[Ano01a]  Anonymous.  Corrections:

“Online Check and Recovery Techniques for Dependable Embedded Processors” and “Micro Economics: Explaining Booms, Busts, and Errors”. IEEE Micro, 21 (6):2, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6002abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m6002.
pdf;   http://dlib.computer.
org/mi/books/mi2001/pdf/
m6002corr.pdf. See [PV01,

Gre01c].

Anonymous:2001:IMA

[Ano01b] Anonymous. IEEE Micro 2001 annual index, volume 21: Author and subject listings. IEEE Micro, 21(6): 80–88, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6080abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m6080.
pdf.

Anonymous:2001:MNH

[Ano01c] Anonymous. Micro news: Heightened security concerns may drive chip technology; chip detects cancer marker; new spectroscopy tool probes semiconductor etching processes; RF MEMS projected to improve wireless; 2.0-GHz [Intel Xeon] processor in production. IEEE Micro, 21(5): 11, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/m5011abs.htm;
                                         http://dlib.computer.org/
                                        mi/books/mi2001/pdf/m5011.
                                        pdf.

Anonymous:2001:MNM

[Ano01d] Anonymous. Micro news: Micro congratulations to current and past EICs: Sakamura receives Takeda Award; Steve Diamond is Computer Society President-elect. IEEE Micro, 21(6):3, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/m6003abs.htm;
                                         http://dlib.computer.org/
                                        mi/books/mi2001/pdf/m6003.
                                        pdf.

Anonymous:2001:MNN

[Ano01e] Anonymous. Micro news: Next-generation lithography; new research at IBM; clockless computing; micro bits. IEEE Micro, 21(2):11, 196, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2011.
pdf.

Anonymous:2001:MNO

[Ano01f] Anonymous. Micro news:

Optical computer; multilevel cell technology; increased disk space; new form of nitrogen; extreme ultraviolet lithography. IEEE Micro, 21(3):7, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3007.
pdf.

Anonymous:2001:MNWa

[Ano01g] Anonymous. Micro news:

William Hewlett dies; IBM introduces MRAM; embedded microprocessor benchmark. IEEE Micro, 21(1): 6–7, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1006.
pdf.

Anonymous:2001:MNWb

[Ano01h] Anonymous. Micro news:

Wireless LAN chip; holographic data storage; systemin-package FCRAM; strained silicon increases chip speed; 157-nm lithography; material improves computer memory; HyperTransport Technology Consortium; faster transistor; lead-free chips; micro bits. IEEE Micro, 21 (4):11, 68, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m4011.
                                        pdf;   m4011abs.htm.

Anonymous:2001:PSa

[Ano01i] Anonymous. Product summary. IEEE Micro, 21(1):94– 95, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m1094.
                                        pdf.

Anonymous:2001:PSb

[Ano01j] Anonymous. Product summary. IEEE Micro, 21 (2):93, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m2093.
                                        pdf.

Anonymous:2001:PSc

[Ano01k] Anonymous. Product summary. IEEE Micro, 21(3): 80, May/June 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.

computer.org/mi/books/mi2001/
pdf/m3080.pdf.

Anonymous:2001:PSd

[Ano01l] Anonymous. Product summary. IEEE Micro, 21 (4):72, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4072.
pdf;   m4072abs.htm.

Anonymous:2001:PSe

[Ano01m] Anonymous. Product summary. IEEE Micro, 21(5): 88, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5088abs.htm;
http://dlib.computer.org/
mi/books/mi2001/pdf/m5088.
pdf.

Anonymous:2002:IMA

[Ano02a] Anonymous. IEEE Micro, 2002 annual index, volume

22. IEEE Micro, 22(6): 78–80, November/December 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/www.computer.org/micro/
mi2002/m6078.pdf.

Anonymous:2002:MNIa

[Ano02b] Anonymous. Micro news:

IBM’s Cell completes design phase; silver molecules render electroluminescent light source; next-generation disc storage; IBM electron microscope; Cornell clustersupercomputing expansion; silicon makes low-voltage gas sensor; NIST chemists explore plastics; micro bits. IEEE Micro, 22(5):9–11, September/October 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2002/
pdf/m5009.pdf.

Anonymous:2002:MNIb

[Ano02c] Anonymous. Micro news: Intel expands 300-mm wafer production; IBM claims smallest working computer circuits; 802.11b chip suppliers predict growth, market effects; Samsung extends systems LSI commitment; light-emitting silicon breakthrough; gas sensors could reduce power demands, lower cost; application-specific nanotubes; micro bits. IEEE Micro, 22(6):6–6, 74, November/December 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2002/
pdf/m6006.pdf;   http://
www.computer.org/micro/
mi2002/m6006abs.htm.

Anonymous:2002:MNL

[Ano02d] Anonymous. Micro news: Laser design makes less expensive chip; protein chains into a wire; crystals and computing; researchers say quantum design works; partners work for 0.9-micron CMOS; tiny technology generates heat; organic memory device; micro bits. IEEE Micro, 22 (4):12–13, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4012.
pdf.

Anonymous:2002:MNO

[Ano02e] Anonymous. Micro news: Optics speed data transfer; microchain MEMS; micro bits [IEEE Standard 802.16 (air interface for fixed broadband wireless access systems)]. IEEE Micro, 22(1): 90, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1090abs.htm;
 http://dlib.computer.org/
mi/books/mi2002/pdf/m1090.
pdf.

Anonymous:2002:PSa

[Ano02f] Anonymous. Product summary. IEEE Micro, 22(1): 96, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1096abs.htm;
 http://dlib.computer.org/

                                        mi/books/mi2002/pdf/m1096.
                                        pdf.

Anonymous:2002:PSb

[Ano02g] Anonymous. Product summary. IEEE Micro, 22(4): 84–85, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m4084.
                                        pdf.

Anonymous:2003:IMA

[Ano03a] Anonymous. 2003 IEEE Micro annual index, vol. 23. IEEE Micro, 23(6):136–139, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/06/m6136.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2003/06/m6136.
                                        htm.

Anonymous:2003:MNIa

[Ano03b] Anonymous. Micro news: IBM ups the ante in silicon transistor speed; new benchmark suite based on highperformance computing applications, MPI and OpenMP [SPEC HPC2002]; EU OKs Hitachi, Mitsubishi Electric semiconductor joint venture; Intel launches Pentium 4 at 3.06 GHz; TSMC unveils viable 25nm transistors. IEEE Micro, 23(1): 6–6, 87, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1006.
pdf.

Anonymous:2003:MNIb

[Ano03c] Anonymous. Micro news: Intel gears up 90-nm processor, chip set rollout; IBM PowerPC runs up to 2.5 GHz; AMD, Fujitsu form chip venture. IEEE Micro, 23(2):6, March/April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m2006.
pdf.

Anonymous:2003:NAL

[Ano03d] Anonymous. News: AMD launches Athlon 64; Intel Pentium 4 aimed at gaming market; Sun Microsystems cofounder [Bill Joy] resigns; PlayStation 3 chip ready for production. IEEE Micro, 23(5):85–??, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2003/05/m5085.pdf.

Anonymous:2003:NIE

[Ano03e] Anonymous. News: Intel’s earnings double in quarter; tiny bubbles key to future liquid-cooled systems; chip diet tests Cisco’s resolve; Intel to release machine learning libraries; proposal in to fit eight Alpha cores onto a chip; micro bits. IEEE Micro, 23(4):6–7, July/August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2003/04/m4006.pdf.

Anonymous:2003:ORR

[Ano03f] Anonymous. Obituary: Rob Rau: 1951–2002: Pioneer of VLIW/EPIC architecture dies. IEEE Micro, 23(1): 7, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m1006.
                                        pdf.

Anonymous:2004:AI

[Ano04a]  Anonymous.  2004 an-

nual index. IEEE Micro, 24(6):135–144, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/06/m6135.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2004/06/m6135.
                                        htm.

Anonymous:2004:MNa

[Ano04b] Anonymous. Micro news. IEEE Micro, 24(2):76–77, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/02/m2076.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/02/m2076.pdf.

Anonymous:2004:MNb

[Ano04c] Anonymous. Micro news.

IEEE Micro, 24(3):72, May/ June 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/dl/
mags/mi/2004/03/m3072.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/03/m3072.
pdf.

Anonymous:2004:MNc

[Ano04d] Anonymous. Micro news.

IEEE Micro, 24(4):82, July/ August 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4082.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/04/m4082.
htm.

Anonymous:2004:MNd

[Ano04e] Anonymous. Micro news.

IEEE Micro, 24(6):129, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/06/m6129.pdf;
http://csdl.computer.org/

                                        dl/mags/mi/2004/06/m6129.
                                        htm.

Anonymous:2004:N

[Ano04f] Anonymous. News. IEEE Micro, 24(1):6, January/ February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/01/m1006.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2004/01/m1006.
                                        htm.

Anonymous:2005:IMA

[Ano05] Anonymous. IEEE Micro 2005 annual index, vol. 25. IEEE Micro, 25(6):92–103, November/December 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2005/06/m6092.pdf.

Anonymous:2006:IMA

[Ano06] Anonymous. IEEE Micro 2006 annual index, volume 26. IEEE Micro, 26(6):56–67, November/December 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2006/06/m6056.pdf.

Anonymous:2007:IMA

[Ano07] Anonymous. IEEE Micro 2007 annual index, volume 27. IEEE Micro, 27(6):i1–i12, November/December 2007.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/06/mmi20070600i1.
pdf.

Anonymous:2008:AI

[Ano08] Anonymous. Annual index.

IEEE Micro, 28(6):0, November/December 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2009:A

[Ano09a] Anonymous. Advert. IEEE Micro, 29(3):1, May/June 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2009:AI

[Ano09b] Anonymous. Annual index.

IEEE Micro, 29(6):1, November/December 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2009:CP

[Ano09c] Anonymous. Call for papers. IEEE Micro, 29(4): 4, July/August 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2009:E

[Ano09d] Anonymous. Erratum. IEEE Micro, 29(6):4, November/ December 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2009:Ma

[Ano09e] Anonymous. Masthead. IEEE Micro, 29(4):1, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2009:Mb

[Ano09f]  Anonymous.  Masthead.

IEEE Micro, 29(5):1, September/October 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2010:CAE

[Ano10a] Anonymous. Call for applications for Editor in Chief. IEEE Micro, 30(1):3, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2010:CP

[Ano10b] Anonymous. Call for papers. IEEE Micro, 30(6): 1, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2010:EMP

[Ano10c] Anonymous. European multicore processing projects. IEEE Micro, 30(5):98–101, September/October 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2010:Ma

[Ano10d] Anonymous. Masthead. IEEE Micro, 30(5):3–4, September/October 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2010:Mb

[Ano10e]  Anonymous.  Masthead.

IEEE Micro, 30(6):3, November/December 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2011:M

[Ano11]  Anonymous.  Masthead.

IEEE Micro, 31(1):3, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2012:R

[Ano12] Anonymous. 2012 reviewers. IEEE Micro, 32(6):62, November/December 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2013:BYC

[Ano13a] Anonymous. Build your career house advertisement. IEEE Micro, 33(6):25, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2013:CHA

[Ano13b]  Anonymous.  Certification

house advertisement. IEEE Micro, 33(1):41, January/ February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2013:CNH

[Ano13c] Anonymous. Computing now house advertisement. IEEE Micro, 33(6):75, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2013:CAP

[Ano13d] Anonymous. Corporate affiliate program house advertisement. IEEE Micro, 33(1):79, January/February 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2013:DMH

[Ano13e] Anonymous. Digital magazines house advertisement. IEEE Micro, 33(6):37, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2013:FC

[Ano13f] Anonymous. Front cover. IEEE Micro, 33(6):c1, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2013:JBH

[Ano13g]  Anonymous.  Jobs board

house advertisement. IEEE Micro, 33(6):c3, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2013:M

[Ano13h] Anonymous. Masthead. IEEE Micro, 33(6):1, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2013:MHA

[Ano13i]  Anonymous.  Membership

house advertisement. IEEE Micro, 33(6):c4, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2013:TC

[Ano13j] Anonymous. Table of contents. IEEE Micro, 33 (6):c2, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:RMD

[Ano14a] Anonymous. 2015 Richard E.

Merwin Distinguished Service Award. IEEE Micro, 34(5): c4, September/October 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi20140500c4.
pdf.

Anonymous:2014:CPa

[Ano14b] Anonymous. Call for papers.

IEEE Micro, 34(1):93, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:CPb

[Ano14c] Anonymous. Call for papers. IEEE Micro, 34(5): c2, September/October 2014.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi20140500c2.
pdf.

Anonymous:2014:C

[Ano14d]  Anonymous.  Conferences.

IEEE Micro, 34(1):79, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:CCX

[Ano14e] Anonymous. COOL Chips XVIII house advertisement. IEEE Micro, 34(6):c4, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi20140600c4.
pdf.

Anonymous:2014:EA

[Ano14f] Anonymous. ePub [advertisement]. IEEE Micro, 34(1):41, January/February 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:FYJa

[Ano14g]  Anonymous.  Focus on

your job search [advertisement]. IEEE Micro, 34(4): 15, July/August 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040015.
pdf.

Anonymous:2014:FYJb

[Ano14h] Anonymous. Focus on your job search house advertisement. IEEE Micro, 34(6): 5, November/December 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060005.
pdf.

Anonymous:2014:FCa

[Ano14i] Anonymous. Front cover.

IEEE Micro, 34(2):c1, March/ April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:FCb

[Ano14j] Anonymous. [Front cover].

IEEE Micro, 34(3):c1, May/ June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:FCc

[Ano14k] Anonymous. Front cover.

IEEE Micro, 34(4):c1, July/ August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi20140400c1.
pdf.

Anonymous:2014:FCd

[Ano14l] Anonymous. Front cover.

IEEE Micro, 34(5):c1, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/

mi/2014/05/mmi20140500c1.
pdf.

Anonymous:2014:FCe

[Ano14m] Anonymous. Front cover. IEEE Micro, 34(6):c1, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi20140600c1.
pdf.

Anonymous:2014:ICC

[Ano14n]  Anonymous.  IEEE cloud

computing [advertisement]. IEEE Micro, 34(1):c3, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:ICS

[Ano14o]  Anonymous.  IEEE Com-

puter Society Harlan D. Mills Award. IEEE Micro, 34(5): c3, September/October 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi20140500c3.
pdf.

Anonymous:2014:IOA

[Ano14p] Anonymous. IEEE open access publishing house advertisement. IEEE Micro, 34(5): 51, September/October 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050051.
pdf.

Anonymous:2014:ISA

[Ano14q]  Anonymous.  IEEE Soft-

ware [advertisement]. IEEE Micro, 34(5):72, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050072.
pdf.

Anonymous:2014:IS

[Ano14r] Anonymous. IEEE STC 2014.

IEEE Micro, 34(1):7, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:ITE

[Ano14s] Anonymous. IEEE transactions on emerging topics in computing. IEEE Micro, 34(4):27, July/August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040027.
pdf.

Anonymous:2014:IA

[Ano14t]  Anonymous.  Intelect ad-

vertisement. IEEE Micro, 34(6):c2, November/ December 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi20140600c2.
pdf.

Anonymous:2014:JBA

[Ano14u] Anonymous. Jobs board [advertisement]. IEEE Micro, 34(1):96, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:Ma

[Ano14v] Anonymous. Masthead. IEEE Micro, 34(1):1, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:Mb

[Ano14w] Anonymous. Masthead. IEEE Micro, 34(2):1, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:Mc

[Ano14x]  Anonymous.  Masthead.

IEEE Micro, 34(4):1, July/ August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/04/mmi2014040001.
                                        pdf.

Anonymous:2014:Md

[Ano14y]  Anonymous.  [Masthead].

IEEE Micro, 34(5):3, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/05/mmi2014050003.
                                        pdf.

Anonymous:2014:Me

[Ano14z]  Anonymous.  Masthead.

IEEE Micro, 34(6):3, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060003.
pdf.

Anonymous:2014:MMAa

[Ano14-27]  Anonymous.  Membership

matters [advertisement]. IEEE Micro, 34(1):59, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:MMAb

[Ano14-28]  Anonymous.  Membership

matters [advertisement]. IEEE Micro, 34(4):43, July/August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2014/04/mmi2014040043.
pdf.

Anonymous:2014:MMH

[Ano14-29]  Anonymous.  Membership

matters [house advertisement]. IEEE Micro, 34(2):c4, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:RSB

[Ano14-30] Anonymous. Rock stars of big data analytics [advertisement]. IEEE Micro, 34(5): 1, September/October 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/05/mmi2014050001.
                                        pdf.

Anonymous:2014:RSC

[Ano14-31] Anonymous. Rock stars of cybersecurity. IEEE Micro, 34(4):c4, July/August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/04/mmi20140400c4.
                                        pdf.

Anonymous:2014:RSMa

[Ano14-32] Anonymous. Rock stars of mobile cloud [advertisement]. IEEE Micro, 34(1):c4, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:RSMb

[Ano14-33] Anonymous. Rock stars of mobile cloud [house advertisement]. IEEE Micro, 34(2):c3, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:SEC

[Ano14-34] Anonymous. Software engineering for the 21st Century [house advertisement]. IEEE Micro, 34(2):21, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:TCa

[Ano14-35] Anonymous. Table of contents. IEEE Micro, 34(1):c2, January/February 2014. CO-DEN IEMIDZ. ISSN 0272-

Anonymous:2014:TCb

[Ano14-36] Anonymous. Table of contents. IEEE Micro, 34(2):c2, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Anonymous:2014:TC

[Ano14-37] Anonymous. Table of contents. IEEE Micro, 34(5): 2, September/October 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050002.
pdf.

Anonymous:2014:TCLa

[Ano14-38]  Anonymous.  Take the

CS Library wherever you go! [advertisement]. IEEE Micro, 34(5):31, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050031.
pdf.

Anonymous:2014:TCLb

[Ano14-39]  Anonymous.  Take the

CS Library wherever you go! House advertisement. IEEE Micro, 34(6):1, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060001.
pdf.

Anonymous:2015:R

[Ano15a] Anonymous. 2014 reviewers. IEEE Micro, 35(1):62–63, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010062.
html.

Anonymous:2015:RMD

[Ano15b] Anonymous. 2016 Richard E. Merwin Distinguished Service Award house advertisement. IEEE Micro, 35(5): 31, September/October 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050031.
pdf.

Anonymous:2015:CNH

[Ano15c] Anonymous. Call for nominees house advertisement. IEEE Micro, 35(5):41, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050041.
pdf.

Anonymous:2015:CPa

[Ano15d] Anonymous. Call for papers. IEEE Micro, 35(3): 71, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.

computer.org/csdl/mags/
mi/2015/03/mmi2015030071.
pdf.

Anonymous:2015:CPb

[Ano15e] Anonymous. Call for papers. IEEE Micro, 35(4): 25, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040025.
pdf.

Anonymous:2015:CSA

[Ano15f] Anonymous. Call for standards award nominations house advertisement. IEEE Micro, 35(5):15, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050015.
pdf.

Anonymous:2015:CC

[Ano15g] Anonymous. Cloud computing. IEEE Micro, 35(3): 33, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030033.
pdf.

Anonymous:2015:CPYa

[Ano15h] Anonymous. Conferences in the palm of your hand. IEEE Micro, 35(3):47, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/03/mmi2015030047.
pdf.

Anonymous:2015:CPYb

[Ano15i]  Anonymous.  Conferences

in the palm of your hand. IEEE Micro, 35(4):59, July/ August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040059.
pdf.

Anonymous:2015:CCX

[Ano15j]  Anonymous.  Cool Chips

XIX house advertisement. IEEE Micro, 35(6):c4, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi20150600c4.
pdf.

Anonymous:2015:FYJa

[Ano15k] Anonymous. Focus on your job search. IEEE Micro, 35 (3):23, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030023.
pdf.

Anonymous:2015:FYJb

[Ano15l] Anonymous. Focus on your job search house advertisement. IEEE Micro, 35(4): c4, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi20150400c4.
pdf.

Anonymous:2015:FCa

[Ano15m] Anonymous. Front cover.

IEEE Micro, 35(1):c1, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi20150100c1.
pdf.

Anonymous:2015:FCb

[Ano15n] Anonymous. Front cover.

IEEE Micro, 35(2):c1, March/ April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/02/mmi20150200c1.
pdf.

Anonymous:2015:FCc

[Ano15o] Anonymous. Front cover.

IEEE Micro, 35(3):c1, May/ June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/03/mmi20150300c1.
pdf.

Anonymous:2015:FCd

[Ano15p] Anonymous. Front cover. IEEE Micro, 35(4):c1, July/ August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/04/mmi20150400c1.
                                        pdf.

Anonymous:2015:FCe

[Ano15q] Anonymous. Front cover. IEEE Micro, 35(5):c1, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        mi/2015/05/mmi20150500c1.
                                        pdf.

Anonymous:2015:FCf

[Ano15r] Anonymous. Front cover. IEEE Micro, 35(6):c1, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        mi/2015/06/mmi20150600c1.
                                        pdf.

Anonymous:2015:GML

[Ano15s]  Anonymous.  Get more,

for less house advertisement. IEEE Micro, 35(2): c4, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/

mi/2015/02/mmi20150200c4.
pdf.

Anonymous:2015:ICC

[Ano15t] Anonymous. IEEE Cloud Computing call for papers house advertisement. IEEE Micro, 35(2):35, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/02/mmi2015020035.
pdf.

Anonymous:2015:ICS

[Ano15u]  Anonymous.  IEEE Com-

puter Society: Be at the center of it all house advertisement. IEEE Micro, 35(6):37, November/December 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060037.
pdf.

Anonymous:2015:KYC

[Ano15v] Anonymous. Keep your career moving forward house advertisement. IEEE Micro, 35(5): 85, September/October 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050085.
pdf.

Anonymous:2015:Ma

[Ano15w]  Anonymous.  Masthead.

IEEE Micro, 35(1):1, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/01/mmi2015010001.
                                        pdf.

Anonymous:2015:Mb

[Ano15x] Anonymous. Masthead. IEEE Micro, 35(2):1, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //www.computer.org/csdl/
                                        mags/mi/2015/02/mmi2015020001.
                                        pdf.

Anonymous:2015:Mc

[Ano15y]  Anonymous.  [Masthead].

IEEE Micro, 35(3):1, May/ June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //www.computer.org/csdl/
                                        mags/mi/2015/03/mmi2015030001.
                                        pdf.

Anonymous:2015:Md

[Ano15z]  Anonymous.  Masthead.

IEEE Micro, 35(4):1, July/ August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/04/mmi2015040001.
                                        pdf.

Anonymous:2015:Me

[Ano15-27]  Anonymous.  Masthead.

IEEE Micro, 35(5):1, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050001.
pdf.

Anonymous:2015:Mf

[Ano15-28]  Anonymous.  Masthead.

IEEE Micro, 35(6):1, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060001.
pdf.

Anonymous:2015:RSCa

[Ano15-29] Anonymous. Rock stars of cyber security house advertisement. IEEE Micro, 35(2): 15, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020015.
pdf.

Anonymous:2015:RSCb

[Ano15-30] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Micro, 35(4): 69, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040069.
pdf.

Anonymous:2015:RSC

[Ano15-31] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Micro, 35(5): 51, September/October 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        mi/2015/05/mmi2015050051.
                                        pdf.

Anonymous:2015:RSW

[Ano15-32] Anonymous. Rock stars of wearables house advertisement. IEEE Micro, 35(4): 79, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/04/mmi2015040079.
                                        pdf.

Anonymous:2015:RCS

[Ano15-33] Anonymous. Rockstars of cyber security. IEEE Micro, 35 (3):c4, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/03/mmi20150300c4.
                                        pdf.

Anonymous:2015:SES

[Ano15-34] Anonymous. Software experts summit house advertisement. IEEE Micro, 35(5): c4, September/October 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi20150500c4.
pdf.

Anonymous:2015:SIO

[Ano15-35] Anonymous. Special issue on online behavioral analysis and modeling house advertisement. IEEE Micro, 35(2): 77, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020077.
pdf.

Anonymous:2015:SIP

[Ano15-36] Anonymous. Special issue on pattern recognition house advertisement. IEEE Micro, 35(2):67, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020067.
pdf.

Anonymous:2015:SRS

[Ano15-37] Anonymous. Startup rock stars. IEEE Micro, 35(1):c4, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi20150100c4.
pdf.

Anonymous:2015:SC

[Ano15-38]  Anonymous.  Stay con-

nected. IEEE Micro, 35(3): 83, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/03/mmi2015030083.
                                        pdf.

Anonymous:2015:SHA

[Ano15-39] Anonymous. STC house advertisement. IEEE Micro, 35 (3):c3, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/03/mmi20150300c3.
                                        pdf.

Anonymous:2015:SAS

[Ano15-40] Anonymous. Student award and scholarship house advertisement. IEEE Micro, 35(5): c3, September/October 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        mi/2015/05/mmi20150500c3.
                                        pdf.

Anonymous:2015:TCL

[Ano15-41]  Anonymous.  Take the

CS Library wherever you go! IEEE Micro, 35(3): 101, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.

computer.org/csdl/mags/
mi/2015/03/mmi2015030101.
pdf.

Anonymous:2016:BRR

[Ano16a] Anonymous. 2017 B. Ramakrishnan Rau Award call for nominations. IEEE Micro, 36(5):64, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050064.
pdf.

Anonymous:2016:RMA

[Ano16b] Anonymous. 2017 Richard E. Merwin Award for distinguished service. IEEE Micro, 36(4):27, July/August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040027.
pdf.

Anonymous:2016:AIC

[Ano16c]  Anonymous.  ACM–IEEE

CS Eckert–Mauchly Award. IEEE Micro, 36(6):c3, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi20160600c3.
pdf.

Anonymous:2016:CNHa

[Ano16d] Anonymous. Call for nominees: House advertisement. IEEE Micro, 36(2): 21, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2016/02/mmi2016020021.
                                        pdf.

Anonymous:2016:CNHb

[Ano16e] Anonymous. Call for nominees house advertisement. IEEE Micro, 36(3):c2, May/June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2016/03/mmi20160300c2.
                                        pdf.

Anonymous:2016:CEA

[Ano16f] Anonymous. Computer entrepreneur award: House advertisement. IEEE Micro, 36(2):c2, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2016/02/mmi20160200c2.
                                        pdf.

Anonymous:2016:FYJd

[Ano16g]  Anonymous.  Focus on

your job search. IEEE Micro, 36(6):59, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060059.
pdf.

Anonymous:2016:FYJc

[Ano16h]  Anonymous.  Focus on

your job search [advertisement]. IEEE Micro, 36(4): 69, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040069.
pdf.

Anonymous:2016:FYJa

[Ano16i] Anonymous. Focus on your job search: House advertisement. IEEE Micro, 36(2): 63, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020063.
pdf.

Anonymous:2016:FYJb

[Ano16j] Anonymous. Focus on your job search: House advertisement. IEEE Micro, 36(2): 72, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020072.
pdf.

Anonymous:2016:FCa

[Ano16k] Anonymous. Front cover. IEEE Micro, 36(1):c1, January/February 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/01/mmi20160100c1.
                                        pdf.

Anonymous:2016:FCb

[Ano16l] Anonymous. Front cover. IEEE Micro, 36(2):c1, March/ April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //www.computer.org/csdl/
                                        mags/mi/2016/02/mmi20160200c1.
                                        pdf.

Anonymous:2016:FCc

[Ano16m] Anonymous. Front cover. IEEE Micro, 36(4):c1, July/ August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/04/mmi20160400c1.
                                        pdf.

Anonymous:2016:FCd

[Ano16n] Anonymous. Front cover. IEEE Micro, 36(5):c1, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/05/mmi20160500c1.
                                        pdf.

Anonymous:2016:FC

[Ano16o] Anonymous. Front cover.

IEEE Micro, 36(6):c1, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi20160600c1.
pdf.

Anonymous:2016:GRY

[Ano16p] Anonymous. Get the recognition you deserve: House advertisement. IEEE Micro, 36 (2):5, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020005.
pdf.

Anonymous:2016:ICC

[Ano16q] Anonymous. IEEE Cloud Computing call for papers house advertisement. IEEE Micro, 36(1):c4, January/ February 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/01/mmi20160100c4.
pdf.

Anonymous:2016:ICSf

[Ano16r]  Anonymous.  IEEE Com-

puter Society 2016 call for major award nominations. IEEE Micro, 36(4):c4, July/ August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi20160400c4.
pdf.

Anonymous:2016:ICSa

[Ano16s] Anonymous. IEEE Computer Society 2016 call for major award nominations: House advertisement. IEEE Micro, 36 (2):1, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020001.
pdf.

Anonymous:2016:ICSc

[Ano16t] Anonymous. IEEE Computer Society 2016 call for major award nominations house advertisement. IEEE Micro, 36 (3):71, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030071.
pdf.

Anonymous:2016:ICSe

[Ano16u]  Anonymous.  IEEE Com-

puter Society: Be at the center of it all [advertisement]. IEEE Micro, 36(4): 37, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

mi/2016/04/mmi2016040037.
pdf.

Anonymous:2016:ICSb

[Ano16v]  Anonymous.  IEEE Com-

puter Society be at the center of it all: House advertisement. IEEE Micro, 36(2): 47, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020047.
pdf.

Anonymous:2016:ICSd

[Ano16w]  Anonymous.  IEEE Com-

puter Society: Be at the center of it all house advertisement. IEEE Micro, 36(3): 127, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030127.
pdf.

Anonymous:2016:ICS

[Ano16x] Anonymous. IEEE Computer Society is where you choose the resources that fit your career. IEEE Micro, 36(6): 1, November/December 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060001.
pdf.

Anonymous:2016:ITB

[Ano16y] Anonymous. IEEE Transactions on Big Data house advertisement. IEEE Micro, 36 (3):c4, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/03/mmi20160300c4.
                                        pdf.

Anonymous:2016:Ma

[Ano16z]  Anonymous.  Masthead.

IEEE Micro, 36(1):1, January/February 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/01/mmi2016010001.
                                        pdf.

Anonymous:2016:Mb

[Ano16-27] Anonymous. Masthead. IEEE Micro, 36(2):3, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //www.computer.org/csdl/
                                        mags/mi/2016/02/mmi2016020003.
                                        pdf.

Anonymous:2016:Mc

[Ano16-28] Anonymous. Masthead. IEEE Micro, 36(3):5, May/June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2016/03/mmi2016030005.
                                        pdf.

Anonymous:2016:Md

[Ano16-29]  Anonymous.  [Masthead].

IEEE Micro, 36(4):3, July/ August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040003.
pdf.

Anonymous:2016:Me

[Ano16-30]  Anonymous.  Masthead.

IEEE Micro, 36(5):3, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050003.
pdf.

Anonymous:2016:M

[Ano16-31]  Anonymous.  Masthead.

IEEE Micro, 36(6):3, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060003.
pdf.

Anonymous:2016:Mf

[Ano16-32] Anonymous. myCS. IEEE Micro, 36(5):c2, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi20160500c2.
pdf.

Anonymous:2016:NMOa

[Ano16-33]  Anonymous.  New mem-

bership options for a better fit. IEEE Micro, 36(5): c3, September/October 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/05/mmi20160500c3.
                                        pdf.

Anonymous:2016:NMOb

[Ano16-34] Anonymous. New membership options for a better fit. IEEE Micro, 36(6):c2, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/06/mmi20160600c2.
                                        pdf.

Anonymous:2016:PI

[Ano16-35] Anonymous. [Publication information]. IEEE Micro, 36 (3):c1, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/03/mmi20160300c1.
                                        pdf.

Anonymous:2016:RSB

[Ano16-36]  Anonymous.  Rock stars

of big data [advertisement]. IEEE Micro, 36(4):c3, July/ August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi20160400c3.
pdf.

Anonymous:2016:RSBa

[Ano16-37] Anonymous. Rock stars of big data: House advertisement. IEEE Micro, 36(2): c4, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi20160200c4.
pdf.

Anonymous:2016:RSBb

[Ano16-38] Anonymous. Rock stars of big data house advertisement. IEEE Micro, 36(3): c3, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi20160300c3.
pdf.

Anonymous:2016:RSPb

[Ano16-39] Anonymous. Rock stars of pervasive, predictive analytics. IEEE Micro, 36(5): 1, September/October 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050001.
pdf.

Anonymous:2016:RSPa

[Ano16-40] Anonymous. Rock stars of pervasive, predictive analytics [advertisement]. IEEE Micro, 36(4):57, July/August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/04/mmi2016040057.
                                        pdf.

Anonymous:2016:RSR

[Ano16-41] Anonymous. Rock stars of risk-based security: House advertisement. IEEE Micro, 36(2):c3, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2016/02/mmi20160200c3.
                                        pdf.

Anonymous:2016:TCa

[Ano16-42] Anonymous. Table of contents. IEEE Micro, 36(5): 2, September/October 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/05/mmi2016050002.
                                        pdf.

Anonymous:2016:TCb

[Ano16-43] Anonymous. Table of contents. IEEE Micro, 36(6): 2, November/December 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.

computer.org/csdl/mags/
mi/2016/06/mmi2016060002.
pdf.

Anonymous:2016:T

[Ano16-44]  Anonymous.  TechIgnite.

IEEE Micro, 36(6):c4, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi20160600c4.
pdf.

Anonymous:2016:WWLc

[Ano16-45]  Anonymous.  Watch the

world’s leading experts take multi-core strategies to new heights. IEEE Micro, 36(5): c4, September/October 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi20160500c4.
pdf.

Anonymous:2016:WWLd

[Ano16-46]  Anonymous.  Watch the

World’s leading experts take multi-core strategies to new heights. IEEE Micro, 36 (6):64, November/December 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2016/06/mmi2016060064.
pdf.

Anonymous:2016:WWLb

[Ano16-47]  Anonymous.  Watch the

World’s leading experts take multi-core strategies to new heights [advertisement]. IEEE Micro, 36(4):88, July/August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2016/04/mmi2016040088.
pdf.

Anonymous:2016:WWLa

[Ano16-48]  Anonymous.  Watch the

world’s leading experts take multi-core strategies to new heights house advertisement. IEEE Micro, 36(3):132, May/ June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

https://www.computer.org/
csdl/mags/mi/2016/03/mmi2016030132.
pdf.

Anonymous:2017:R

[Ano17a] Anonymous. 2016 reviewers. IEEE Micro, 37(1):77, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010077.
html.

Anonymous:2017:RMA

[Ano17b] Anonymous. 2018 Richard E. Merwin Award for Distinguished Service. IEEE Micro, 37(5):c4, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi20170500c4.
pdf.

Anonymous:2017:AYCc

[Ano17c] Anonymous. Achieve your career goals with the fit that’s right for you. IEEE Micro, 37(5):1, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050001.
pdf.

Anonymous:2017:AYCd

[Ano17d] Anonymous. Achieve your career goals with the fit that’s right for you. IEEE Micro, 37(6):1, November/ December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060001.
pdf.

Anonymous:2017:AYCb

[Ano17e] Anonymous. Achieve your career goals with the fit that’s right for you [advertisement]. IEEE Micro, 37(4): 1, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        mi/2017/04/mmi2017040001.
                                        pdf.

Anonymous:2017:AYCa

[Ano17f] Anonymous. Achieve your career goals with the fit that’s right for you. House advertisement. IEEE Micro, 37 (3):1, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030001.
                                        pdf.

Anonymous:2017:AIC

[Ano17g]  Anonymous.  ACM–IEEE

CS Eckert–Mauchly Award. IEEE Micro, 37(1):73, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/01/mmi2017010073.
                                        pdf.

Anonymous:2017:APM

[Ano17h] Anonymous. Architectures for the post-Moore era. IEEE Micro, 37(4):c1, July/ August 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi20170400c1.
                                        pdf.

Anonymous:2017:CN

[Ano17i] Anonymous. Call for nominees. IEEE Micro, 37(1):41, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010041.
pdf.

Anonymous:2017:CNEb

[Ano17j] Anonymous. Call for nominees Education Awards nominations. IEEE Micro, 37(4): c3, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi20170400c3.
pdf.

Anonymous:2017:CNEa

[Ano17k] Anonymous. Call for nominees: Education Awards nominations house advertisement. IEEE Micro, 37(3): c3, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi20170300c3.
pdf.

Anonymous:2017:CPA

[Ano17l] Anonymous. Call for papers:

Advances in parallel graph processing: Algorithms, architectures, and application frameworks house advertisement. IEEE Micro, 37(1):c4, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/01/mmi20170100c4.
                                        pdf.

Anonymous:2017:C

[Ano17m]  Anonymous.  COMPSAC

2018. IEEE Micro, 37(6):c4, November/December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/06/mmi20170600c4.
                                        pdf.

Anonymous:2017:CPYa

[Ano17n] Anonymous. Conferences in the palm of your hand house advertisement. IEEE Micro, 37(3):129, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030129.
                                        pdf.

Anonymous:2017:CCH

[Ano17o] Anonymous. Cool chips and hot interconnects. IEEE Micro, 37(5):c1, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/05/mmi20170500c1.
                                        pdf.

Anonymous:2017:CPYb

[Ano17p] Anonymous. Cross-pollinate your ideas. IEEE Micro, 37(6):c3, November/ December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi20170600c3.
pdf.

Anonymous:2017:FYJ

[Ano17q] Anonymous. Focus on your job search house advertisement. IEEE Micro, 37(3): 79, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030079.
pdf.

Anonymous:2017:FCa

[Ano17r] Anonymous. Front cover.

IEEE Micro, 37(1):c1, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi20170100c1.
pdf.

Anonymous:2017:FCb

[Ano17s] Anonymous. Front cover.

IEEE Micro, 37(2):c1, March/ April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

https://www.computer.org/
csdl/mags/mi/2017/02/mmi20170200c1.
pdf.

Anonymous:2017:FCc

[Ano17t] Anonymous. Front cover. IEEE Micro, 37(3):c1, May/ June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/mi/2017/03/mmi20170300c1.
                                        pdf.

Anonymous:2017:GFH

[Ano17u] Anonymous. got flaws? House advertisement. IEEE Micro, 37(3):115, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030115.
                                        pdf.

Anonymous:2017:ICC

[Ano17v] Anonymous. IEEE Cloud Computing call for papers house advertisement. IEEE Micro, 37(3):105, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2017/03/mmi2017030105.
                                        pdf.

Anonymous:2017:ICSa

[Ano17w]  Anonymous.  IEEE Com-

puter Society 2017 call for major award nominations. IEEE Micro, 37(1):80, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.

computer.org/csdl/mags/
mi/2017/01/mmi2017010080.
pdf.

Anonymous:2017:ICSd

[Ano17x]  Anonymous.  IEEE Com-

puter Society 2017 call for major award nominations. IEEE Micro, 37(4):9, July/ August 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040009.
pdf.

Anonymous:2017:ICSc

[Ano17y] Anonymous. IEEE Computer Society 2017 call for major award nominations house advertisement. IEEE Micro, 37 (3):c4, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi20170300c4.
pdf.

Anonymous:2017:ICS

[Ano17z] Anonymous. IEEE Computer Society: Be at the center of it all. IEEE Micro, 37(6):96, November/December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060096.
pdf.

Anonymous:2017:ICSb

[Ano17-27]  Anonymous.  IEEE Com-

puter Society Harlan D. Mills Award house advertisement. IEEE Micro, 37(3): 87, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030087.
                                        pdf.

Anonymous:2017:ICSe

[Ano17-28]  Anonymous.  IEEE Com-

puter Society house advertisement. IEEE Micro, 37(4): 29, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi2017040029.
                                        pdf.

Anonymous:2017:ICSf

[Ano17-29] Anonymous. IEEE Computer Society Richard E. Merwin Student Leadership Scholarship. IEEE Micro, 37(4): 39, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi2017040039.
                                        pdf.

Anonymous:2017:LBT

[Ano17-30]  Anonymous.  Looking for

the BEST tech job for you? IEEE Micro, 37(4):

19, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040019.
pdf.

Anonymous:2017:Ma

[Ano17-31]  Anonymous.  Masthead.

IEEE Micro, 37(1):3, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010003.
pdf.

Anonymous:2017:Mb

[Ano17-32] Anonymous. Masthead. IEEE Micro, 37(2):3, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2017/02/mmi2017020003.
pdf.

Anonymous:2017:Mc

[Ano17-33] Anonymous. Masthead. IEEE Micro, 37(3):3, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2017/03/mmi2017030003.
pdf.

Anonymous:2017:Md

[Ano17-34]  Anonymous.  Masthead.

IEEE Micro, 37(4):2, July/ August 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi2017040002.
                                        pdf.

Anonymous:2017:Mf

[Ano17-35]  Anonymous.  Masthead.

IEEE Micro, 37(5):2, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/05/mmi2017050002.
                                        pdf.

Anonymous:2017:Mg

[Ano17-36]  Anonymous.  Masthead.

IEEE Micro, 37(6):2, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/06/mmi2017060002.
                                        pdf.

Anonymous:2017:Me

[Ano17-37] Anonymous. myCS. IEEE Micro, 37(4):63, July/August 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2017/04/mmi2017040063.
                                        pdf.

Anonymous:2017:Mh

[Ano17-38] Anonymous. myCS. IEEE Micro, 37(6):39, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060039.
pdf.

Anonymous:2017:MHA

[Ano17-39] Anonymous. myCS house advertisement. IEEE Micro, 37 (3):51, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030051.
pdf.

Anonymous:2017:NMOa

[Ano17-40] Anonymous. New membership options for a better fit. IEEE Micro, 37(1):c2, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi20170100c2.
pdf.

Anonymous:2017:NMOc

[Ano17-41]  Anonymous.  New mem-

bership options for a better fit. IEEE Micro, 37(4): c2, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi20170400c2.
pdf.

Anonymous:2017:NMOd

[Ano17-42]  Anonymous.  New mem-

bership options for a better fit. IEEE Micro, 37(5): c2, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/05/mmi20170500c2.
                                        pdf.

Anonymous:2017:NMO

[Ano17-43] Anonymous. New membership options for a better fit. IEEE Micro, 37(6):c2, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/06/mmi20170600c2.
                                        pdf.

Anonymous:2017:NMOb

[Ano17-44] Anonymous. New membership options for a better fit. House advertisement. IEEE Micro, 37(3):c2, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi20170300c2.
                                        pdf.

Anonymous:2017:NSS

[Ano17-45]  Anonymous.  Nominations

are solicited for the Seymour Cray, Sidney Fernbach & Ken Kennedy awards house advertisement. IEEE Micro, 37(1): c3, January/February 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi20170100c3.
pdf.

Anonymous:2017:OMU

[Ano17-46]  Anonymous.  One mem-

bership. Unlimited knowledge. IEEE Micro, 37(5): c3, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi20170500c3.
pdf.

Anonymous:2017:PC

[Ano17-47] Anonymous. Prepare to connect. IEEE Micro, 37(4): 51, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040051.
pdf.

Anonymous:2017:PCH

[Ano17-48]  Anonymous.  Prepare to

connect house advertisement. IEEE Micro, 37(3): 39, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030039.
pdf.

Anonymous:2017:TCa

[Ano17-49] Anonymous. Table of contents. IEEE Micro, 37(1):2, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/01/mmi2017010002.
                                        pdf.

Anonymous:2017:TCb

[Ano17-50] Anonymous. Table of contents. IEEE Micro, 37(2): 2, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/02/mmi2017020002.
                                        pdf.

Anonymous:2017:TCc

[Ano17-51] Anonymous. Table of contents. IEEE Micro, 37(3): 2, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030002.
                                        pdf.

Anonymous:2017:TCd

[Ano17-52] Anonymous. Table of contents. IEEE Micro, 37(4): 3, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi2017040003.
                                        pdf.

Anonymous:2017:TCe

[Ano17-53] Anonymous. Table of contents. IEEE Micro, 37(5): 3, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050003.
pdf.

Anonymous:2017:TC

[Ano17-54] Anonymous. Table of contents. IEEE Micro, 37(6): 3, November/December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060003.
pdf.

Anonymous:2017:TCL

[Ano17-55] Anonymous. Take the CS Library wherever you go! House advertisement. IEEE Micro, 37(3):132, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030132.
pdf.

Anonymous:2017:T

[Ano17-56]  Anonymous.  TechIgnite.

IEEE Micro, 37(1):1, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        mi/2017/01/mmi2017010001.
                                        pdf.

Anonymous:2017:ULP

[Ano17-57]  Anonymous.  Ultra-low-

power processors. IEEE Micro, 37(6):c1, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/06/mmi20170600c1.
                                        pdf.

Anonymous:2017:UPEa

[Ano17-58]  Anonymous.  Upsilon Pi

Epsilon Student Excellence Award: Up to four $1,000 awards. IEEE Micro, 37(4): c4, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi20170400c4.
                                        pdf.

Anonymous:2017:UPEb

[Ano17-59]  Anonymous.  Upsilon Pi

Epsilon Student Excellence Award: Up to four $1,000 awards. IEEE Micro, 37(5): 43, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/05/mmi2017050043.
                                        pdf.

Anonymous:2018:BAA

[Ano18a] Anonymous. Babbage Award ad. IEEE Micro, 38(4): c2, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi20180400c2.
pdf.

Anonymous:2018:C

[Ano18b] Anonymous. Cover. IEEE Micro, 38(6):C1, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08584173.pdf.

Anonymous:2018:EAA

[Ano18c] Anonymous. Education award ad. IEEE Micro, 38(4): 1, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040001.
pdf.

Anonymous:2018:EDT

[Ano18d]  Anonymous.  Erratum to

Defect-Tolerant Logic Synthesis for Memristor Crossbars with Performance Evaluation. IEEE Micro, 38(6):85, November/December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.

                                        computer.org/csdl/mags/
                                        mi/2018/06/08585392-abs.

html. See [TMA18].

Anonymous:2018:FCa

[Ano18e] Anonymous. Front cover. IEEE Micro, 38(1):c1, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/01/mmi20180100c1.
                                        pdf.

Anonymous:2018:FCb

[Ano18f] Anonymous. Front cover. IEEE Micro, 38(2):c1, March/ April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/mi/2018/02/mmi20180200c1.
                                        pdf.

Anonymous:2018:FCc

[Ano18g] Anonymous. Front cover. IEEE Micro, 38(3):c1, May/ June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/mi/2018/03/mmi20180300c1.
                                        pdf.

Anonymous:2018:FCd

[Ano18h] Anonymous. Front cover. IEEE Micro, 38(4):c1, July/ August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

mi/2018/04/mmi20180400c1.
pdf.

Anonymous:2018:FCe

[Ano18i] Anonymous. Front cover.

IEEE Micro, 38(5):c1, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi20180500c1.
pdf.

Anonymous:2018:FC

[Ano18j] Anonymous. Front cover.

IEEE Micro, 38(6):C1, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08584167.pdf.

Anonymous:2018:HMA

[Ano18k] Anonymous. Harlan Mills Award ad. IEEE Micro, 38(4): c4, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi20180400c4.
pdf.

Anonymous:2018:HAA

[Ano18l]  Anonymous.  HPC award

ad. IEEE Micro, 38(2): c2, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        mi/2018/02/mmi20180200c2.
                                        pdf.

Anonymous:2018:ICSf

[Ano18m]  Anonymous.  IEEE Com-

puter Society. IEEE Micro, 38(6):C3, November/ December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/06/08584180.pdf.

Anonymous:2018:ICSg

[Ano18n]  Anonymous.  IEEE Com-

puter Society. IEEE Micro, 38(6):C4, November/ December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/06/08584146.pdf.

Anonymous:2018:ICSa

[Ano18o]  Anonymous.  IEEE Com-

puter Society information. IEEE Micro, 38(1):c3, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/01/mmi20180100c3.
                                        pdf.

Anonymous:2018:ICSb

[Ano18p] Anonymous. IEEE Computer Society information. IEEE Micro, 38(2):5, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2018/02/mmi2018020005.
pdf.

Anonymous:2018:ICSc

[Ano18q] Anonymous. IEEE Computer Society information. IEEE Micro, 38(3):c3, May/June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2018/03/mmi20180300c3.
pdf.

Anonymous:2018:ICSd

[Ano18r]  Anonymous.  IEEE Com-

puter Society information. IEEE Micro, 38(4):5, July/ August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040005.
pdf.

Anonymous:2018:ICSe

[Ano18s] Anonymous. IEEE Computer Society information. IEEE Micro, 38(5):c3, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi20180500c3.
pdf.

Anonymous:2018:ITE

[Ano18t] Anonymous. IEEE Technology and Engineering Management Society TEMS Administrative Committee. IEEE Micro, 38(6):1, November/ December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/06/08584168.pdf.

Anonymous:2018:JBA

[Ano18u]  Anonymous.  Jobs board

ad. IEEE Micro, 38(2): 1, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/02/mmi2018020001.
                                        pdf.

Anonymous:2018:Mb

[Ano18v] Anonymous. Masthead. IEEE Micro, 38(2):4, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2018/02/mmi2018020004.
                                        pdf.

Anonymous:2018:Mc

[Ano18w] Anonymous. Masthead. IEEE Micro, 38(3):1, May/June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2018/03/mmi2018030001.
                                        pdf.

Anonymous:2018:Md

[Ano18x]  Anonymous.  Masthead.

IEEE Micro, 38(4):4, July/ August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040004.
pdf.

Anonymous:2018:Me

[Ano18y]  Anonymous.  Masthead.

IEEE Micro, 38(5):1, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050001.
pdf.

Anonymous:2018:MAa

[Ano18z]  Anonymous.  Membership

ad. IEEE Micro, 38(2): c4, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/02/mmi20180200c4.
pdf.

Anonymous:2018:MAb

[Ano18-27]  Anonymous.  Membership

ad. IEEE Micro, 38(3): c2, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi20180300c2.
pdf.

Anonymous:2018:MAd

[Ano18-28]  Anonymous.  Membership

ad. IEEE Micro, 38(5): c4, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/05/mmi20180500c4.
                                        pdf.

Anonymous:2018:MAAa

[Ano18-29] Anonymous. Merwin Award ad. IEEE Micro, 38(4): c3, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/04/mmi20180400c3.
                                        pdf.

Anonymous:2018:MAc

[Ano18-30] Anonymous. MyCS ad. IEEE Micro, 38(3):c4, May/June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2018/03/mmi20180300c4.
                                        pdf.

Anonymous:2018:OPP

[Ano18-31] Anonymous. Outsider pathways to prominence. IEEE Micro, 38(2):81–84, March/ April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/mi/2018/02/mmi2018020081.
                                        html.

Anonymous:2018:SNA

[Ano18-32] Anonymous. Social networking ad. IEEE Micro, 38(5): c2, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi20180500c2.
pdf.

Anonymous:2018:SMA

[Ano18-33] Anonymous. Student membership ad. IEEE Micro, 38(2):c3, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/02/mmi20180200c3.
pdf.

Anonymous:2018:TCa

[Ano18-34] Anonymous. Table of contents. IEEE Micro, 38(1):2–3, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010002.
pdf.

Anonymous:2018:TCb

[Ano18-35] Anonymous. Table of contents. IEEE Micro, 38(2):2– 3, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/02/mmi2018020002.
pdf.

Anonymous:2018:TCc

[Ano18-36] Anonymous. Table of contents. IEEE Micro, 38(3): 2–3, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/03/mmi2018030002.
                                        pdf.

Anonymous:2018:TCd

[Ano18-37] Anonymous. Table of contents. IEEE Micro, 38(4):2– 3, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/04/mmi2018040002.
                                        pdf.

Anonymous:2018:TCe

[Ano18-38] Anonymous. Table of contents. IEEE Micro, 38(5):2– 3, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/05/mmi2018050002.
                                        pdf.

Anonymous:2018:TC

[Ano18-39] Anonymous. Table of contents. IEEE Micro, 38(6):2– 3, November/December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/06/08584179.pdf.

Anonymous:2019:CEa

[Ano19a]  Anonymous.  Computing

Edge. IEEE Micro, 39(4): C4, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CEb

[Ano19b]  Anonymous.  Computing

Edge. IEEE Micro, 39(5): C4, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CEc

[Ano19c]  Anonymous.  Computing

Edge. IEEE Micro, 39(6):C4, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:IAb

[Ano19d]  Anonymous.  IEE An-

nals. IEEE Micro, 39 (5):117, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:IAc

[Ano19e]  Anonymous.  IEE An-

nals. IEEE Micro, 39 (5):118, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:IAa

[Ano19f]  Anonymous.  IEEE An-

nals. IEEE Micro, 39(4): 26, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ICA

[Ano19g] Anonymous. IEEE Computer Architecture [Letters]. IEEE Micro, 39(4):65, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:IM

[Ano19h]  Anonymous.  IEEE Mi-

cro. IEEE Micro, 39 (5):113, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ITB

[Ano19i] Anonymous. IEEE Transactions on Big Data. IEEE Micro, 39(4):54, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:SP

[Ano19j] Anonymous. Security & Privacy. IEEE Micro, 39(6): 5, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CAb

[Ano19k] Anonymous. Call for articles. IEEE Micro, 39(4): 5, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CANa

[Ano19l] Anonymous. Call for award nominations. IEEE Micro, 39(1):C2, January/February 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CANb

[Ano19m] Anonymous. Call for award nominations. IEEE Micro, 39(2):C2, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CAa

[Ano19n] Anonymous. Call for awards.

IEEE Micro, 39(4):C2, July/ August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CPIa

[Ano19o] Anonymous. Call for papers: IEEE Micro top picks 2019. IEEE Micro, 39 (5):112, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:CPIb

[Ano19p] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 39 (6):81, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:FCa

[Ano19q] Anonymous. Front cover. IEEE Micro, 39(3):C1, May/ June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:FCb

[Ano19r] Anonymous. Front cover. IEEE Micro, 39(4):C1, July/ August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:FCc

[Ano19s] Anonymous. Front cover. IEEE Micro, 39(5):C1, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:FCd

[Ano19t] Anonymous. Front cover. IEEE Micro, 39(6):C1, November/December 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ICSb

[Ano19u] Anonymous. IEEE Computer Society. IEEE Micro, 39(3): C3, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ICSc

[Ano19v] Anonymous. IEEE Computer Society. IEEE Micro, 39(4): C3, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ICSd

[Ano19w] Anonymous. IEEE Computer Society. IEEE Micro, 39(5): C3, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ICSf

[Ano19x] Anonymous. IEEE Computer Society. IEEE Micro, 39 (6):C3, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ICSe

[Ano19y]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 39 (6):37, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ICSa

[Ano19z] Anonymous. IEEE Computer Society information. IEEE Micro, 39(2):C3, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ISP

[Ano19-27] Anonymous. IEEE security privacy. IEEE Micro, 39(4): 43, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:ITS

[Ano19-28] Anonymous. IEEE TRANS-ACTIONS ON SUSTAIN-ABLE COMPUTING. IEEE Micro, 39(4):7, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:IWC

[Ano19-29] Anonymous. IEEE World Congress on Services 2019. IEEE Micro, 39(3):C2, May/ June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:KYCa

[Ano19-30] Anonymous. Keep your career options open. IEEE Micro, 39(5):C2, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:KYCb

[Ano19-31] Anonymous. Keep your career options open. IEEE Micro, 39 (6):C2, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:LBT

[Ano19-32] Anonymous. Looking for the BEST tech job for you? IEEE Micro, 39(3):C4, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:Ma

[Ano19-33]  Anonymous.  [Masthead].

IEEE Micro, 39(2):1, March/ April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:Mb

[Ano19-34] Anonymous. Masthead. IEEE Micro, 39(3):1, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:Mc

[Ano19-35] Anonymous. Masthead. IEEE Micro, 39(4):1, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:Md

[Ano19-36]  Anonymous.  Masthead.

IEEE Micro, 39(5):1, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:Me

[Ano19-37]  Anonymous.  Masthead.

IEEE Micro, 39(6):1, November/December 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:SMY

[Ano19-38] Anonymous. Share and manage your research data. IEEE Micro, 39(5):125, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:SCB

[Ano19-39] Anonymous. Stay connected [back cover]. IEEE Micro, 39(2):C4, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:TCa

[Ano19-40] Anonymous. Table of contents. IEEE Micro, 39(1):2–3, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:TCb

[Ano19-41] Anonymous. Table of contents. IEEE Micro, 39(2):2– 3, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:TCc

[Ano19-42] Anonymous. Table of contents. IEEE Micro, 39(3): 2–3, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:TCd

[Ano19-43] Anonymous. Table of contents. IEEE Micro, 39(4):2– 3, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:TCe

[Ano19-44] Anonymous. Table of contents. IEEE Micro, 39(5):2– 3, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2019:TCf

[Ano19-45] Anonymous. Table of contents. IEEE Micro, 39(6):2– 3, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:CEa

[Ano20a] Anonymous. Computing Edge.

IEEE Micro, 40(1):C4, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:CEb

[Ano20b]  Anonymous.  Computing

Edge. IEEE Micro, 40(2): C4, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:C

[Ano20c] Anonymous. ComputingEdge. IEEE Micro, 40(6):C4, November/December 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IAH

[Ano20d] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 40(6):58, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICGa

[Ano20e]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Micro, 40(1):66, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICGb

[Ano20f]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Micro, 40(2): 5, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICG

[Ano20g]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Micro, 40(3): 90, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ISPa

[Ano20h] Anonymous. IEEE Security & Privacy. IEEE Micro, 40(1): 43, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ISPb

[Ano20i] Anonymous. IEEE Security & Privacy. IEEE Micro, 40(2): 52, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ISPc

[Ano20j] Anonymous. IEEE Security & Privacy. IEEE Micro, 40 (3):115, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ISPd

[Ano20k] Anonymous. IEEE Security & Privacy. IEEE Micro, 40(4): 92, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ISPe

[Ano20l] Anonymous. IEEE Security & Privacy. IEEE Micro, 40(4): 111, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ISPf

[Ano20m] Anonymous. IEEE Security & Privacy. IEEE Micro, 40(5): 55, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ISPg

[Ano20n] Anonymous. IEEE Security & Privacy. IEEE Micro, 40 (6):84, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ITC

[Ano20o] Anonymous. IEEE Transactions on Computers. IEEE Micro, 40(1):34, January/ February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IPa

[Ano20p] Anonymous. IT Professional. IEEE Micro, 40(1):56, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IPb

[Ano20q] Anonymous. IT Professional.

IEEE Micro, 40(2):71, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IP

[Ano20r]  Anonymous.  IT Profes-

sional. IEEE Micro, 40(4): 9, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICAa

[Ano20s] Anonymous. IT Professional:

Call for articles. IEEE Micro, 40(5):45, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICAb

[Ano20t] Anonymous. ITProfessional:

Call for articles. IEEE Micro, 40(6):66, November/December 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:CPI

[Ano20u] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Micro, 40(5): 86, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Ca

[Ano20v]  Anonymous.  COMPSAC

2020. IEEE Micro, 40(1):C2, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Cb

[Ano20w]  Anonymous.  COMPSAC

2020. IEEE Micro, 40(2): 73, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:CTY

[Ano20x]  Anonymous.  CVPR2020:

Thank you to our sponsors! IEEE Micro, 40 (6):C2, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ECO

[Ano20y] Anonymous. Evolving career opportunities need your skills. IEEE Micro, 40(4): 131, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:FCa

[Ano20z] Anonymous. Front cover. IEEE Micro, 40(1):C1, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:FCb

[Ano20-27] Anonymous. Front cover.

IEEE Micro, 40(2):C1, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:FCc

[Ano20-28] Anonymous. Front cover.

IEEE Micro, 40(3):C1, May/ June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:FCd

[Ano20-29] Anonymous. Front cover.

IEEE Micro, 40(4):C1, July/ August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:FCe

[Ano20-30] Anonymous. Front cover.

IEEE Micro, 40(5):C1, September/October 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:FC

[Ano20-31] Anonymous. Front cover.

IEEE Micro, 40(6):C1, November/December 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:GPNa

[Ano20-32] Anonymous. Get published in the new IEEE Open Journal of the Computer Society.

IEEE Micro, 40(3):117, May/ June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:GPNb

[Ano20-33] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 40(4): 133, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:GPN

[Ano20-34] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 40(6):C3, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Ha

[Ano20-35] Anonymous. HOST 2020. IEEE Micro, 40(2):C2, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Hb

[Ano20-36] Anonymous. HOST 2020. IEEE Micro, 40(3):116, May/ June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Hc

[Ano20-37] Anonymous. HOST 2020. IEEE Micro, 40(4):132, July/ August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSb

[Ano20-38] Anonymous. IEEE Computer Society. IEEE Micro, 40(1): C3, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSc

[Ano20-39] Anonymous. IEEE Computer Society. IEEE Micro, 40(2): C3, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSf

[Ano20-40] Anonymous. IEEE Computer Society. IEEE Micro, 40(3): C3, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSi

[Ano20-41] Anonymous. IEEE Computer Society. IEEE Micro, 40(4): C3, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSm

[Ano20-42] Anonymous. IEEE Computer Society. IEEE Micro, 40(5): C3, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSo

[Ano20-43] Anonymous. IEEE Computer Society. IEEE Micro, 40 (6):85, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSe

[Ano20-44] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 40(3):55, May/June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSh

[Ano20-45] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 40(4):21, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSk

[Ano20-46]  Anonymous.  IEEE Com-

puter Society: Call for papers. IEEE Micro, 40(5): 36, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSn

[Ano20-47]  Anonymous.  IEEE Com-

puter Society: Call for papers. IEEE Micro, 40(6):22, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSa

[Ano20-48]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 40(1):24, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSd

[Ano20-49]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 40(3):C2, May/ June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSg

[Ano20-50]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 40(4): C2, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSj

[Ano20-51]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 40(5): C2, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:ICSl

[Ano20-52] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 40(5):87, September/October 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IPC

[Ano20-53] Anonymous. IEEE Pervasive Computing: Call for articles. IEEE Micro, 40 (6):48, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IQWb

[Ano20-54] Anonymous. IEEE Quantum Week. IEEE Micro, 40(3): C4, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IQWc

[Ano20-55] Anonymous. IEEE Quantum Week. IEEE Micro, 40(4): C4, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IQWd

[Ano20-56] Anonymous. IEEE Quantum Week. IEEE Micro, 40(5): C4, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:IQWa

[Ano20-57] Anonymous. IEEE Quantum Week 2020. IEEE Micro, 40(2):72, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:KYC

[Ano20-58] Anonymous. Keep your career options open. IEEE Micro, 40(1):93, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Ma

[Ano20-59]  Anonymous.  Masthead.

IEEE Micro, 40(1):1, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Mb

[Ano20-60] Anonymous. Masthead. IEEE Micro, 40(2):1, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Mc

[Ano20-61] Anonymous. Masthead. IEEE Micro, 40(3):1, May/June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Md

[Ano20-62] Anonymous. Masthead. IEEE Micro, 40(4):1, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Me

[Ano20-63]  Anonymous.  Masthead.

IEEE Micro, 40(5):1, September/October 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:Mf

[Ano20-64]  Anonymous.  Masthead.

IEEE Micro, 40(6):1, November/December 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:TCa

[Ano20-65] Anonymous. Table of contents. IEEE Micro, 40(1):2–3, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:TCb

[Ano20-66] Anonymous. Table of contents. IEEE Micro, 40(2):2– 3, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:TCc

[Ano20-67] Anonymous. Table of contents. IEEE Micro, 40(3): 2–3, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:TCd

[Ano20-68] Anonymous. Table of contents. IEEE Micro, 40(4):2– 3, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:TCe

[Ano20-69] Anonymous. Table of contents. IEEE Micro, 40(5):2– 3, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:TCf

[Ano20-70] Anonymous. Table of contents. IEEE Micro, 40(6):2– 3, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2020:TTT

[Ano20-71] Anonymous. Top technology trends for 2020 featured in Computer. IEEE Micro, 40(1):14, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:CSEa

[Ano21a] Anonymous. Computing in Science & Engineering. IEEE Micro, 41(3):79, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:CSEb

[Ano21b] Anonymous. Computing in Science & Engineering. IEEE Micro, 41(4):23, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:CSEc

[Ano21c] Anonymous. Computing in Science & Engineering. IEEE Micro, 41(5):32, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Ca

[Ano21d] Anonymous. ComputingEdge. IEEE Micro, 41(1):C2, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Cb

[Ano21e] Anonymous. ComputingEdge. IEEE Micro, 41(2):C2, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Cc

[Ano21f] Anonymous. ComputingEdge. IEEE Micro, 41(3):C2, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Cd

[Ano21g] Anonymous. ComputingEdge. IEEE Micro, 41(4):C2, July/ August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Ce

[Ano21h] Anonymous. ComputingEdge. IEEE Micro, 41(5):C2, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Cf

[Ano21i] Anonymous. ComputingEdge.

IEEE Micro, 41(6):C2, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:IAH

[Ano21j] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 41(6):167, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICG

[Ano21k]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Micro, 41 (6):147, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ISP

[Ano21l] Anonymous. IEEE Security & Privacy. IEEE Micro, 41 (6):106, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ITB

[Ano21m] Anonymous. IEEE Transactions on Big Data. IEEE Micro, 41(6):96, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ITSa

[Ano21n] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Micro, 41(2): 21, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ITSb

[Ano21o] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Micro, 41 (5):132, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ITSc

[Ano21p] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Micro, 41 (6):70, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:IP

[Ano21q] Anonymous. IT Professional. IEEE Micro, 41(2):83, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:IPC

[Ano21r]  Anonymous.  IT Profes-

sional: Call for articles. IEEE Micro, 41(6):15, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICA

[Ano21s] Anonymous. ITProfessional:

Call for articles. IEEE Micro, 41(5):123, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:CAI

[Ano21t] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Micro, 41 (6):108, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:CPI

[Ano21u] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 41 (6):67, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:CBA

[Ano21v] Anonymous. Charles Babbage Award. IEEE Micro, 41(5): 14, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:CSEd

[Ano21w] Anonymous. Computing in science & engineering. IEEE Micro, 41(6):51, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:FCa

[Ano21x] Anonymous. Front cover. IEEE Micro, 41(1):C1, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:FCb

[Ano21y] Anonymous. Front cover. IEEE Micro, 41(2):C1, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:FCc

[Ano21z] Anonymous. Front cover. IEEE Micro, 41(3):C1, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:FCd

[Ano21-27] Anonymous. Front cover. IEEE Micro, 41(4):C1, July/ August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:FCe

[Ano21-28] Anonymous. Front cover. IEEE Micro, 41(5):C1, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:FCf

[Ano21-29] Anonymous. Front cover.

IEEE Micro, 41(6):C1, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:GPNa

[Ano21-30] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 41(1):C3, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:GPNb

[Ano21-31] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 41(2): C3, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:GPNc

[Ano21-32] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 41(3):C3, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:GPNd

[Ano21-33] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 41(4): C3, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:GPNe

[Ano21-34] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 41(5): C3, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:GPNf

[Ano21-35] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 41(6):C3, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSd

[Ano21-36] Anonymous. IEEE Computer Society. IEEE Micro, 41(1): 85, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSj

[Ano21-37] Anonymous. IEEE Computer Society. IEEE Micro, 41(3):

108, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSn

[Ano21-38] Anonymous. IEEE Computer Society. IEEE Micro, 41(4): 84, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSt

[Ano21-39] Anonymous. IEEE Computer Society. IEEE Micro, 41 (5):129, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSx

[Ano21-40] Anonymous. IEEE Computer Society. IEEE Micro, 41 (6):183, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSa

[Ano21-41]  Anonymous.  IEEE Com-

puter Society: Call for papers. IEEE Micro, 41(1):68, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSh

[Ano21-42] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 41(3):70, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSm

[Ano21-43] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 41(4):83, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSq

[Ano21-44] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 41(5):112, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSv

[Ano21-45]  Anonymous.  IEEE Com-

puter Society: Call for papers. IEEE Micro, 41(6):12, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSe

[Ano21-46]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 41(1):C4, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSg

[Ano21-47]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 41(2):

C4, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSl

[Ano21-48]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 41(3):C4, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSp

[Ano21-49]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 41(4): C4, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSu

[Ano21-50]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 41(5): C4, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSw

[Ano21-51]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 41 (6):178, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSf

[Ano21-52] Anonymous. IEEE Computer Society information. IEEE Micro, 41(2):84, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSb

[Ano21-53]  Anonymous.  IEEE Com-

puter Society Jobs Board. IEEE Micro, 41(1):82, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSi

[Ano21-54] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 41(3):102, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSo

[Ano21-55] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 41(4):85, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSs

[Ano21-56] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 41(5):124, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSy

[Ano21-57] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 41(6):C4, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSc

[Ano21-58]  Anonymous.  IEEE Com-

puter Society Volunteer Service Awards. IEEE Micro, 41(1):84, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSk

[Ano21-59]  Anonymous.  IEEE Com-

puter Society Volunteer Service Awards. IEEE Micro, 41 (3):109, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:ICSr

[Ano21-60]  Anonymous.  IEEE Com-

puter Society: Volunteer service awards. IEEE Micro, 41 (5):113, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Ma

[Ano21-61]  Anonymous.  Masthead.

IEEE Micro, 41(1):1, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Mb

[Ano21-62] Anonymous. Masthead. IEEE Micro, 41(2):1, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Mc

[Ano21-63] Anonymous. Masthead. IEEE Micro, 41(3):1, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Md

[Ano21-64] Anonymous. Masthead. IEEE Micro, 41(4):1, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Me

[Ano21-65] Anonymous. Masthead. IEEE Micro, 41(4):1, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Mf

[Ano21-66]  Anonymous.  Masthead.

IEEE Micro, 41(5):1, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:Mg

[Ano21-67]  Anonymous.  [Masthead].

IEEE Micro, 41(6):1, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:SCT

[Ano21-68] Anonymous. Subscribe to CiSE today! IEEE Micro, 41(2):85, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:TCa

[Ano21-69] Anonymous. Table of contents. IEEE Micro, 41(1):2–3, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:TCb

[Ano21-70] Anonymous. Table of contents. IEEE Micro, 41(2):2– 3, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:TCc

[Ano21-71] Anonymous. Table of contents. IEEE Micro, 41(3): 2–3, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:TCd

[Ano21-72] Anonymous. Table of contents. IEEE Micro, 41(4):2– 3, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:TCe

[Ano21-73] Anonymous. Table of contents. IEEE Micro, 41(4):2– 3, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:TCf

[Ano21-74] Anonymous. Table of contents. IEEE Micro, 41(5):2– 3, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:TCg

[Ano21-75] Anonymous. Table of contents. IEEE Micro, 41(6):2– 4, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:AWA

[Ano22a] Anonymous. AI’s 10 to Watch Award: Call for nominations. IEEE Micro, 42(3):66, May/ June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CCS

[Ano22b] Anonymous. Computer: Call for special issue proposals. IEEE Micro, 42(3):39, May/ June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CSEa

[Ano22c]  Anonymous.  Computing

in Science & Engineering. IEEE Micro, 42(1):24, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CSEb

[Ano22d] Anonymous. Computing in Science & Engineering. IEEE Micro, 42(2):25, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CSEc

[Ano22e] Anonymous. Computing in Science & Engineering. IEEE Micro, 42(6):54, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Ca

[Ano22f] Anonymous. ComputingEdge.

IEEE Micro, 42(1):C2, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Cb

[Ano22g] Anonymous. ComputingEdge.

IEEE Micro, 42(2):C2, March/ April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Cc

[Ano22h] Anonymous. ComputingEdge. IEEE Micro, 42(3):C2, May/ June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Cd

[Ano22i] Anonymous. ComputingEdge. IEEE Micro, 42(4):C2, July/ August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:IAHa

[Ano22j] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 42(1):140, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:IAHb

[Ano22k] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 42(6):140, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICGa

[Ano22l]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Micro, 42(1):136, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICGb

[Ano22m]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Micro, 42 (6):133, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ISPa

[Ano22n] Anonymous. IEEE Security & Privacy. IEEE Micro, 42(1): 88, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ISPb

[Ano22o] Anonymous. IEEE Security & Privacy. IEEE Micro, 42(2): 84, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ISPc

[Ano22p] Anonymous. IEEE Security & Privacy. IEEE Micro, 42 (6):114, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ITBa

[Ano22q] Anonymous. IEEE Transactions on Big Data. IEEE Micro, 42(1):60, January/ February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ITBb

[Ano22r] Anonymous. IEEE Transactions on Big Data. IEEE Micro, 42(2):67, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ITBc

[Ano22s] Anonymous. IEEE Transactions on Big Data. IEEE Micro, 42(6):106, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CPIa

[Ano22t] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 42(1):50, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CPIb

[Ano22u] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 42(2): 34, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CPIc

[Ano22v] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 42(3): 65, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CPId

[Ano22w] Anonymous. Call for papers:

IEEE Transactions on Computers. IEEE Micro, 42(4): 132, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CPIe

[Ano22x] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 42 (5):127, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CPIf

[Ano22y] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 42 (6):95, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:CMU

[Ano22z] Anonymous. Carnegie Mellon University. IEEE Micro, 42(4):77, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Ce

[Ano22-27] Anonymous. ComputingEdge.

IEEE Micro, 42(6):C2, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:FCa

[Ano22-28] Anonymous. Front cover. IEEE Micro, 42(1):C1, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:FCb

[Ano22-29] Anonymous. Front cover. IEEE Micro, 42(2):C1, March/ April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:FCc

[Ano22-30] Anonymous. Front cover. IEEE Micro, 42(3):C1, May/ June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:FCd

[Ano22-31] Anonymous. Front cover. IEEE Micro, 42(4):C1, July/ August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:FCe

[Ano22-32] Anonymous. Front cover. IEEE Micro, 42(5):C1, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:FCf

[Ano22-33] Anonymous. Front cover. IEEE Micro, 42(6):C1, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:GPNa

[Ano22-34] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 42(1):C3, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:GPNb

[Ano22-35] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 42(2): C3, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:GPNc

[Ano22-36] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 42(3):C3, May/ June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:GPNd

[Ano22-37] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 42(4): C3, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:GPNe

[Ano22-38] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 42(5): C3, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:GPNf

[Ano22-39] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 42(6):C3, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:HAN

[Ano22-40] Anonymous. The Humphrey Award nominations. IEEE Micro, 42(3):77, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSb

[Ano22-41] Anonymous. IEEE Computer Society. IEEE Micro, 42(1): 137, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSe

[Ano22-42] Anonymous. IEEE Computer Society. IEEE Micro, 42(2): 85, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSh

[Ano22-43] Anonymous. IEEE Computer Society. IEEE Micro, 42(3): 85, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSl

[Ano22-44] Anonymous. IEEE Computer Society. IEEE Micro, 42(4): 133, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSp

[Ano22-45] Anonymous. IEEE Computer Society. IEEE Micro, 42 (5):128, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSv

[Ano22-46] Anonymous. IEEE Computer Society. IEEE Micro, 42 (6):141, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSa

[Ano22-47]  Anonymous.  IEEE Com-

puter Society: Call for papers. IEEE Micro, 42(1):7, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSd

[Ano22-48] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 42(2):5, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSg

[Ano22-49] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 42(3):38, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSk

[Ano22-50] Anonymous. IEEE Computer Society: Call for papers. IEEE Micro, 42(4):36, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSn

[Ano22-51]  Anonymous.  IEEE Com-

puter Society: Call for papers. IEEE Micro, 42(5): 8, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSs

[Ano22-52]  Anonymous.  IEEE Com-

puter Society: Call for papers. IEEE Micro, 42(6): 5, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSq

[Ano22-53] Anonymous. IEEE Computer Society D&I Fund. IEEE Micro, 42(5):129, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSu

[Ano22-54] Anonymous. IEEE Computer Society D&I Fund. IEEE Micro, 42(6):134, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSj

[Ano22-55] Anonymous. IEEE Computer Society election. IEEE Micro, 42(4):26, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSo

[Ano22-56]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 42 (5):118, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSt

[Ano22-57]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 42 (6):115, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSc

[Ano22-58]  Anonymous.  IEEE Com-

puter Society Jobs Board. IEEE Micro, 42(1):C4, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSf

[Ano22-59] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 42(2):C4, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSi

[Ano22-60] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 42(3):C4, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSm

[Ano22-61] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 42(4):C4, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSr

[Ano22-62] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 42(5):C4, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:ICSw

[Ano22-63] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 42(6):C4, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:IQW

[Ano22-64] Anonymous. IEEE Quantum Week. IEEE Micro, 42(5): C2, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Ma

[Ano22-65]  Anonymous.  Masthead.

IEEE Micro, 42(1):1, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Mb

[Ano22-66] Anonymous. Masthead. IEEE Micro, 42(2):1, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Mc

[Ano22-67] Anonymous. Masthead. IEEE Micro, 42(3):1, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Md

[Ano22-68] Anonymous. Masthead. IEEE Micro, 42(4):1, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Me

[Ano22-69]  Anonymous.  Masthead.

IEEE Micro, 42(5):1, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:Mf

[Ano22-70]  Anonymous.  Masthead.

IEEE Micro, 42(6):1, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:RCSa

[Ano22-71] Anonymous. Over the rainbow: 21st Century security & privacy podcast. IEEE Micro, 42(4):115, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:RCSb

[Ano22-72] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Micro, 42(5):33, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:RCSc

[Ano22-73] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Micro, 42(6):87, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:TCa

[Ano22-74] Anonymous. Table of contents. IEEE Micro, 42(1):2–3, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:TCb

[Ano22-75] Anonymous. Table of contents. IEEE Micro, 42(2):2– 3, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:TCc

[Ano22-76] Anonymous. Table of contents. IEEE Micro, 42(3): 2–3, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:TCd

[Ano22-77] Anonymous. Table of contents. IEEE Micro, 42(5):2– 3, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2022:TCe

[Ano22-78] Anonymous. Table of contents. IEEE Micro, 42(6):2– 3, November/December 2022.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICAa

[Ano23a] Anonymous. 2023 IEEE Conference on Artificial Intelligence. IEEE Micro, 43(2): 140, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICAb

[Ano23b] Anonymous. 2023 IEEE Conference on Artificial Intelligence. IEEE Micro, 43(3): C2, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CEa

[Ano23c]  Anonymous.  Computing

Edge. IEEE Micro, 43(5): C4, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CEb

[Ano23d]  Anonymous.  Computing

Edge. IEEE Micro, 43(6):C4, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CSEa

[Ano23e] Anonymous. Computing in Science & Engineering. IEEE Micro, 43(2):8, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CSEb

[Ano23f] Anonymous. Computing in Science & Engineering. IEEE Micro, 43(3):49, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CSEc

[Ano23g] Anonymous. Computing in Science & Engineering. IEEE Micro, 43(4):44, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CSEd

[Ano23h] Anonymous. Computing in Science & Engineering. IEEE Micro, 43(5):100, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CSEe

[Ano23i] Anonymous. Computing in Science & Engineering. IEEE Micro, 43(6):73, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CSC

[Ano23j] Anonymous. Computing in Science Computing in Science & Engineering Engineering. IEEE Micro, 43(1):66, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Ca

[Ano23k] Anonymous. ComputingEdge. IEEE Micro, 43(1):C4, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Cb

[Ano23l] Anonymous. ComputingEdge. IEEE Micro, 43(2):C4, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IAHa

[Ano23m] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 43(2):98, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IAHb

[Ano23n] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 43(3):88, May/ June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IAHc

[Ano23o] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 43(4): 52, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IOJ

[Ano23p]  Anonymous.  IEEE Open

Journal of the Computer Society. IEEE Micro, 43(3): 73, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ISPa

[Ano23q] Anonymous. IEEE Security & Privacy. IEEE Micro, 43 (1):6, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ISPb

[Ano23r] Anonymous. IEEE Security & Privacy. IEEE Micro, 43 (2):5, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ISPc

[Ano23s] Anonymous. IEEE Security & Privacy. IEEE Micro, 43(4): 79, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CMA

[Ano23t] Anonymous. Call for 2023 major awards nominations. IEEE Micro, 43(3):84, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CAIa

[Ano23u] Anonymous. Call for articles: IEEE Pervasive Computing. IEEE Micro, 43(4): 87, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CAIb

[Ano23v] Anonymous. Call for articles: IT Professional. IEEE Micro, 43(4):97, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CPIb

[Ano23w] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 43(2): 119, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CPIe

[Ano23x] Anonymous. Call for papers: IEEE Transactions on Computers. IEEE Micro, 43(4): 70, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CPIc

[Ano23y] Anonymous. Call for papers for IEEE Computer Society. IEEE Micro, 43(3):6, May/ June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CPId

[Ano23z] Anonymous. Call for papers for IEEE Computer Society. IEEE Micro, 43(4): 5, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:CPIa

[Ano23-27] Anonymous. Call for papers: IEEE Computer Society. IEEE Micro, 43(2):95, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:FCa

[Ano23-28] Anonymous. Front cover.

IEEE Micro, 43(1):C1, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:FCb

[Ano23-29] Anonymous. Front cover.

IEEE Micro, 43(2):C1, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:FCc

[Ano23-30] Anonymous. Front cover.

IEEE Micro, 43(3):C1, May/ June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:FCd

[Ano23-31] Anonymous. Front cover.

IEEE Micro, 43(4):C1, July/ August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:FCe

[Ano23-32] Anonymous. Front cover. IEEE Micro, 43(5):C1, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:FCf

[Ano23-33] Anonymous. Front cover. IEEE Micro, 43(6):C1, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:GPN

[Ano23-34] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. IEEE Micro, 43(2):96, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSi

[Ano23-35] Anonymous. IEEE Computer Society. IEEE Micro, 43(3): 85, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSm

[Ano23-36] Anonymous. IEEE Computer Society. IEEE Micro, 43(4): 124, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSr

[Ano23-37] Anonymous. IEEE Computer Society. IEEE Micro, 43(5): 97, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSp

[Ano23-38]  Anonymous.  IEEE Com-

puter Society — call for papers. IEEE Micro, 43(5): 7, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSj

[Ano23-39] Anonymous. IEEE Computer Society career center. IEEE Micro, 43(3):C3, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSn

[Ano23-40] Anonymous. IEEE Computer Society Career Center. IEEE Micro, 43(4):C3, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSs

[Ano23-41]  Anonymous.  IEEE Com-

puter Society career center. IEEE Micro, 43(5): C3, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSx

[Ano23-42] Anonymous. IEEE Computer Society career center. IEEE Micro, 43(6):C3, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSh

[Ano23-43] Anonymous. IEEE Computer Society D&I fund. IEEE Micro, 43(3):83, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSl

[Ano23-44] Anonymous. IEEE Computer Society D&I Fund. IEEE Micro, 43(4):118, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSo

[Ano23-45] Anonymous. IEEE Computer Society D&I fund. IEEE Micro, 43(5):C2, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSt

[Ano23-46] Anonymous. IEEE Computer Society D&I Fund. IEEE Micro, 43(6):C2, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSa

[Ano23-47]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 43(1):C2, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSd

[Ano23-48]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 43(2): C2, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSg

[Ano23-49]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 43(3):8, May/ June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSk

[Ano23-50]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 43(4): 18, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSq

[Ano23-51]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 43(5): 44, September/October 2023.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSu

[Ano23-52]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 43 (6):27, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSb

[Ano23-53]  Anonymous.  IEEE Com-

puter Society information. IEEE Micro, 43(1):89, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSe

[Ano23-54] Anonymous. IEEE Computer Society information. IEEE Micro, 43(2):141, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSw

[Ano23-55] Anonymous. IEEE Computer Society information. IEEE Micro, 43(6):125, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSc

[Ano23-56]  Anonymous.  IEEE Com-

puter Society Jobs Board.

IEEE Micro, 43(1):C3, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSf

[Ano23-57] Anonymous. IEEE Computer Society Jobs Board. IEEE Micro, 43(2):C3, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ICSv

[Ano23-58]  Anonymous.  IEEE Com-

puter Society Volunteer Service awards. IEEE Micro, 43 (6):28, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:ID

[Ano23-59] Anonymous. IEEE Data-Port. IEEE Micro, 43(5): 77, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IQWa

[Ano23-60] Anonymous. IEEE Quantum Week. IEEE Micro, 43(3): C4, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IQWb

[Ano23-61] Anonymous. IEEE Quantum Week. IEEE Micro, 43(4):

C2, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IQWc

[Ano23-62] Anonymous. IEEE Quantum Week. IEEE Micro, 43(4): C4, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:IQWd

[Ano23-63] Anonymous. IEEE quantum week. IEEE Micro, 43(5): 76, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Ma

[Ano23-64]  Anonymous.  Masthead.

IEEE Micro, 43(1):1, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Mb

[Ano23-65] Anonymous. Masthead. IEEE Micro, 43(2):1, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Mc

[Ano23-66] Anonymous. Masthead. IEEE Micro, 43(3):1, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Md

[Ano23-67] Anonymous. Masthead. IEEE Micro, 43(4):1, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Me

[Ano23-68]  Anonymous.  Masthead.

IEEE Micro, 43(5):1, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:Mf

[Ano23-69]  Anonymous.  Masthead.

IEEE Micro, 43(6):1, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:RCSa

[Ano23-70] Anonymous. Over the rainbow: 21st Century security & privacy podcast. IEEE Micro, 43(1):88, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:RCSb

[Ano23-71] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Micro, 43(2):38, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:RCSc

[Ano23-72] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Micro, 43(3):17, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:RCSd

[Ano23-73] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Micro, 43(4):27, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:RCSe

[Ano23-74] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Micro, 43(5):54, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:RCSf

[Ano23-75] Anonymous. Over the rainbow: 21st century security & privacy podcast. IEEE Micro, 43(6):39, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:PSE

[Ano23-76]  Anonymous.  Publications

seek 2025 Editors in Chief. IEEE Micro, 43(6):102, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:TCa

[Ano23-77] Anonymous. Table of contents. IEEE Micro, 43(1):2–3, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:TCb

[Ano23-78] Anonymous. Table of contents. IEEE Micro, 43(2):2– 3, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:TCc

[Ano23-79] Anonymous. Table of contents. IEEE Micro, 43(3): 2–3, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:TCd

[Ano23-80] Anonymous. Table of contents. IEEE Micro, 43(4):2– 3, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:TCe

[Ano23-81] Anonymous. Table of contents. IEEE Micro, 43(5):2– 3, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:TCf

[Ano23-82] Anonymous. Table of contents. IEEE Micro, 43(6):2– 3, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:WHSa

[Ano23-83]  Anonymous.  Watts S.

Humphrey Software Quality Award. IEEE Micro, 43(3): 30, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2023:WHSb

[Ano23-84]  Anonymous.  Watts S.

Humphrey Software Quality Award. IEEE Micro, 43(4): 36, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:CSEb

[Ano24a] Anonymous. Computing in Science & Engineering. IEEE Micro, 44(2):77, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:CEa

[Ano24b] Anonymous. Computing Edge. IEEE Micro, 44(1):C4, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:CEb

[Ano24c]  Anonymous.  Computing

Edge. IEEE Micro, 44(2): C4, March/April 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:CSEa

[Ano24d]  Anonymous.  Computing

in Science & Engineering. IEEE Micro, 44(1):74, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:IAHa

[Ano24e] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 44(1):58, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:IAHb

[Ano24f] Anonymous. IEEE Annals of the History of Computing. IEEE Micro, 44(2):70, March/ April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ISP

[Ano24g] Anonymous. IEEE Security & Privacy magazine. IEEE Micro, 44(2):59, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ISM

[Ano24h]  Anonymous.  IEEE Secu-

rity&Privacy magazine. IEEE Micro, 44(1):37, January/ February 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:IPa

[Ano24i] Anonymous. IT Professional. IEEE Micro, 44(1):78, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:IPb

[Ano24j] Anonymous. IT Professional. IEEE Micro, 44(2):80, March/ April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:FCa

[Ano24k] Anonymous. Front cover. IEEE Micro, 44(1):C1, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:FCb

[Ano24l] Anonymous. Front cover. IEEE Micro, 44(2):C1, March/ April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:GPN

[Ano24m] Anonymous. Get published in the new IEEE Transactions on Privacy. IEEE Micro, 44(2):19, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSf

[Ano24n] Anonymous. IEEE Computer Society — call for papers. IEEE Micro, 44(2):7, March/ April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSd

[Ano24o]  Anonymous.  IEEE Com-

puter Society career center. IEEE Micro, 44(1):C3, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSi

[Ano24p] Anonymous. IEEE Computer Society Career Center. IEEE Micro, 44(2):C3, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSa

[Ano24q]  Anonymous.  IEEE Com-

puter Society D&I Fund. IEEE Micro, 44(1):C2, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSe

[Ano24r] Anonymous. IEEE Computer Society D&I Fund. IEEE Micro, 44(2):C2, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSb

[Ano24s]  Anonymous.  IEEE Com-

puter Society has you covered! IEEE Micro, 44(1):5, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSc

[Ano24t]  Anonymous.  IEEE Com-

puter Society information. IEEE Micro, 44(1):69, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSg

[Ano24u] Anonymous. IEEE Computer Society information. IEEE Micro, 44(2):60, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:ICSh

[Ano24v]  Anonymous.  IEEE Com-

puter Society Volunteer Service Awards. IEEE Micro, 44(2):71, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:Ma

[Ano24w]  Anonymous.  Masthead.

IEEE Micro, 44(1):1, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:Mb

[Ano24x] Anonymous. Masthead. IEEE Micro, 44(2):1, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:PSE

[Ano24y]  Anonymous.  Publications

seek 2025 Editors in Chief. IEEE Micro, 44(1):75, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:TCa

[Ano24z] Anonymous. Table of contents. IEEE Micro, 44(1):2–3, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2024:TCb

[Ano24-27] Anonymous. Table of contents. IEEE Micro, 44(2):2– 3, March/April 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anderson:1996:GEI

[ANS96] Thomas E. Anderson, Kathleen M. Nichols, and Vivian Shen. Guest Editors’ introduction: Developing interconnect technology. IEEE Micro, 16(1):10–11, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arakawa:1998:SRM

[ANUN98] Fumio Arakawa, Osamu Nishii, Kunio Uchiyama, and Norio Nakagawa. SH4 RISC multimedia microprocessor. IEEE Micro, 18 (2):26–34, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2026.
pdf;   http://www.computer.
org/micro/mi1998/m2026abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24– 26, 1997.

Adams:1997:PID

[AO97] Lyle Adams and Michael Ou. Processor integration in a disk controller: Embedding a RISC processor in a complex ASIC to reduce cost and improve performance. IEEE Micro, 17(4):44–48, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4044.
pdf.

Awaga:1995:GPC

[AOYS95] Makoto Awaga, Tatsushi Ohtsuka, Hideki Yoshizawa, and Shigeru Sasaki. 3D graphics processor chip set. IEEE Micro, 15(6):37–45, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ainsworth:2007:CCE

[AP07] Thomas William Ainsworth and Timothy Mark Pinkston. Characterizing the Cell EIB on-chip network. IEEE Micro, 27(5):6–14, September/October 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Adams:1998:CPD

[APS98] Jim Adams, Ken Parulski, and Kevin Spaulding. Color processing in digital cameras. IEEE Micro, 18(6): 20–30, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6020.
pdf;   http://www.computer.
org/micro/mi1998/m6020abs.
htm.

Adams:1992:CTV

[AQT^{+}92] Marcus Adams, Yi Qian, Jacek Tomaszunas, Josef Burtscheidt, Edgar Kaiser, and Csaba Juhasz. Conformance testing of VMEbus and Multibus II products. IEEE Micro, 12(1):57– 64, January/February 1992.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Abadir:1983:LTT

[AR83] Magdy S. Abadir and Hassan K. Reghbati. LSI testing techniques. IEEE Micro, 3(1):34–51, January/February 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Amirtharajah:2016:HC

[AR16a] Rajeevan Amirtharajah and Behnam Robatmili. Hot chips 27. IEEE Micro, 36(2):6– 7, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2016/02/mmi2016020006-
                                        abs.html.

Amirtharajah:2016:HCH

[AR16b] Rajeevan Amirtharajah and Behnam Robatmili. Hot chips 27 highlights. IEEE Micro, 36 (2):64–69, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2016/02/mmi2016020064.
                                        html.

Araki:2000:MS

[Ara00] Shigeo Araki. The memory stick. IEEE Micro, 20 (4):40–46, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4040.
pdf;   http://www.computer.
org/micro/mi2000/m4040abs.
htm.

Akkary:2003:CPR

[ARS03] Haitham Akkary, Ravi Rajwar, and Srikanth T. Srinivasan. Checkpoint processing and recovery: An efficient, scalable alternative to reorder buffers. IEEE Micro, 23(6):11–19, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6011abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/06/m6011.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6011.pdf.

Albertengo:1990:PCG

[AS90] Guido Albertengo and Riccardo Sisto. Parallel CRC generation. IEEE Micro, 10(5): 63–71, September/October 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Abdelguerfi:1991:FGA

[AS91a] M. Abdelguerfi and A. K.

Sood. A fine-grain architecture for relational database aggregation operations. IEEE Micro, 11(6):35–43, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Abdelguerfi:1991:GEI

[AS91b] M. Abdelguerfi and A. K. Sood. Guest Editors’ introduction: Database machines — trends and opportunities. IEEE Micro, 11(6):6– 7, November/December 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Alpert:1995:GEI

[AS95] Donald Alpert and Alan Jay Smith. Guest Editors’ introduction: Hot Chips VI. IEEE Micro, 15(2):8–9, March/ April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arvind:1999:UTR

[AS99] Arvind and Xiaowei W. Shen. Using term rewriting systems to design and verify processors. IEEE Micro, 19 (3):36–46, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m3036.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m3036abs.
                                        htm.

Adve:2005:GEI

[AS05] Sarita V. Adve and Pia Sanda. Guest Editors’ introduction:

Reliability-aware microarchitecture. IEEE Micro, 25(6):8– 9, November/December 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/06/m6008.pdf.

Andersen:2010:RFD

[AS10] David G. Andersen and Steven Swanson. Rethinking flash in the data center. IEEE Micro, 30(4):52– 54, July/August 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Adit:2022:PLT

[AS22] Neil Adit and Adrian Sampson. Performance left on the table: An evaluation of compiler autovectorization for RISC-V. IEEE Micro, 42(5): 41–48, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Asano:2005:LPD

[ASD^{+}05] Toru Asano, Joel Silberman, Sang H. Dhong, Osamu Takahashi, Michael White, Scott Cottier, Takaaki Nakazato, Atsushi Kawasumi, and Hiroshi Yoshihara. Low-power design approach of 11FO4 256-kbyte embedded SRAM for the synergistic processor element of a cell processor. IEEE Micro, 25(5):30– 38, September/October 2005.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Abadal:2015:BEM

[ASK^{+}15] Sergi Abadal, Benny Sheinman, Oded Katz, Ofer Markish, Danny Elad, Yvan Fournier, Damian Roca, Mauricio Hanzich, Guillaume Houzeaux, Mario Nemirovsky, Eduard Alarcon, and Albert Cabellos-Aparicio. Broadcastenabled massive multicore architectures: A wireless RF approach. IEEE Micro, 35(5):52–61, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        mi/2015/05/mmi2015050052-
                                        abs.html.

Akgun:2019:NCD

[ASX19] I. Akgun, D. Stow, and Y. Xie. Network-on-chip design guidelines for monolithic 3-D integration. IEEE Micro, 39(6): 46–53, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Awaga:1993:BVC

[AT93] Makoto Awaga and Hiromasa Takahashi. The μVP 64-bit vector coprocessor: a new implementation of highperformance numerical computation. IEEE Micro, 13(5): 24–36, September/October 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anh:2009:RTO

[AT09] Tran Nguyen Bao Anh and Su-Lim Tan. Real-time operating systems for small microcontrollers. IEEE Micro, 29(5):30–45, September/October 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Atkins:1991:PIM

[Atk91] Mark Atkins. Performance and the I860 microprocessor. IEEE Micro, 11(5):24–27, 72– 78, September/October 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Angstadt:2022:SLS

[ATS^{+}22] Kevin Angstadt, Tommy Tracy, Kevin Skadron, Jean-Baptiste Jeannin, and Westley Weimer. Synthesizing legacy string code for FPGAs using bounded automata learning. IEEE Micro, 42(5):70– 77, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Auden:1995:N

[Aud95]  W. H. Auden.  Night-

mail. IEEE Micro, 15(2):4– 5, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

August:2012:PSC

[Aug12] David I. August. Parallelizing sequential code. IEEE Micro, 32(4):6–7, July/August 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Abella:2008:RPW

[AVU^{+}08] Jaume Abella, Xavier Vera, Osman S. Unsal, Oguz Ergin, Antonio González, and James W. Tschanz. Refueling: Preventing wire degradation due to electromigration. IEEE Micro, 28(6):37–46, November/December 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Alameldeen:2003:AWV

[AW03] Alaa R. Alameldeen and David A. Wood. Addressing workload variability in architectural simulations. IEEE Micro, 23(6):94–98, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6094abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/06/m6094.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6094.pdf.

Alameldeen:2006:ICH

[AW06] Alaa R. Alameldeen and David A. Wood. IPC considered harmful for multiprocessor workloads. IEEE Micro, 26(4):8–17, July/August 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Asanovic:2010:GEI

[AW10] Krste Asanović and Ralph Wittig. Guest Editors’ introduction: Hot Chips 21. IEEE Micro, 30(2):5–6, March/ April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Araujo:2022:SIC

[AW22] Guido Araujo and Lucas Wanner. Special issue on compiling for accelerators. IEEE Micro, 42(5):6–8, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brown:1990:ISE

[BAC^{+}90] Emil W. Brown, Anant Agrawal, Trevor Creary, Michael F. Klein, David Murata, and Joseph Petolino. Implementing Sparc in ECL. IEEE Micro, 10(1):10–22, January/February 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Beecroft:2005:QDH

[BAH^{+}05] Jon Beecroft, David Addison, David Hewson, Moray McLaren, Duncan Roweth, Fabrizio Petrini, and Jarek Nieplocha. QsNet^{II}: Defining high-performance network design. IEEE Micro, 25(4):34– 47, July/August 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Balakrishnan:1984:PIF

[Bal84a] R. V. Balakrishnan. The proposed IEEE 896 Futurebus — a solution to the bus driving problem. IEEE Micro, 4 (4):23–27, July/August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Baldwin:1984:SRP

[Bal84b] G. Baldwin. Status-report P694 microprocessor assembly language. IEEE Micro, 4(1): 83, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Baldwin:1984:TAL

[Bal84c]  Geoff Baldwin.  Towards

an assembly language standard. IEEE Micro, 4(4):81– 85, July/August 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Becker:1993:PM

[BAM^{+}93] Michael C. Becker, Michael S.

Allen, Charles R. Moore, John S. Muhich, and David P. Tuttle. The PowerPC 601 microprocessor. IEEE Micro, 13(5):54–68, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bose:2003:GEI

[BAM03] Pradip Bose, David H. Albonesi, and Diana Marculescu. Guest Editors’ introduction: Power and complexity aware design. IEEE Micro, 23(5):8–11, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/05/m5008.pdf.

Barroso:2021:BHW

[Bar21] L. A. Barroso. A brief history of warehouse-scale computing. IEEE Micro, 41 (2):78–83, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Baum:2012:HC

[BB12] Allen Baum and Bevan Bass.

Hot Chips 23. IEEE Micro, 32(2):6–7, March/April 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bose:2017:ASC

[BB17] Pradip Bose and Alper Buyuktosunoglu. Architectural support for cognitive processing. IEEE Micro, 37(1):6–7, January/ February 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010006-
abs.html.

Banerjee:2021:NSS

[BBB^{+}21] Anindya Banerjee, Sankar Basu, Erik Brunvand, Pinaki Mazumder, Rance Cleaveland, Gurdip Singh, Margaret Martonosi, and Fernanda Pembleton. Navigating the seismic shift of post-Moore computer systems design. IEEE Micro, 41(6):162–167, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barry:2015:AVP

[BBC^{+}15] Brendan Barry, Cormac Brick, Fergal Connor, David Donohoe, David Moloney, Richard Richmond, Martin O’Riordan, and Vasile Toma. Always-on vision processing unit for mobile applications. IEEE Micro, 35(2):56– 66, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.

computer.org/csdl/mags/
mi/2015/02/mmi2015020056-
abs.html.

Bini:2011:RMM

[BBE^{+}11] Enrico Bini, Giorgio Buttazzo, Johan Eker, Stefan Schorr, Raphael Guerra, Gerhard Fohler, Karl-Erik Arzen, Vanessa Romero Segovia, and Claudio Scordino. Resource management on multicore systems: The ACTORS approach. IEEE Micro, 31(3): 72–81, May/June 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bergman:2009:GEI

[BBP09] Keren Bergman, Ron Brightwell, and Fabrizio Petrini. Guest Editors’ introduction: Hot interconnects. IEEE Micro, 29(4):5–7, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brooks:2000:PAM

[BBS^{+}00] David M. Brooks, Pradip Bose, Stanley E. Schuster, Hans Jacobson, Prabhakar N. Kudva, Alper Buyuktosunoğlu, John-David Wellman, Victor Zyuban, Manish Gupta, and Peter W. Cook. Power-aware microarchitecture: Design and modeling challenges for nextgeneration microprocessors. IEEE Micro, 20(6):26–44, November/December 2000.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6026.
pdf;   http://www.computer.
org/micro/mi2000/m6026abs.
htm.

Butler:2011:BAM

[BBSG11] Michael Butler, Leslie Barnes, Debjit Das Sarma, and Bob Gelinas. Bulldozer: An approach to multithreaded compute performance. IEEE Micro, 31(2):6–15, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Boggs:2015:DNF

[BBTV15] Darrell Boggs, Gary Brown, Nathan Tuck, and K S Venkatraman. Denver: Nvidia’s first 64-bit ARM processor. IEEE Micro, 35(2):46– 55, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020046-
abs.html.

Baker:1986:PSC

[BC86] Marc A. Baker and Vincent J. Coli. The PAL20RA10 story — the customization of a standard product. IEEE Micro, 6(5):45–60, September/October 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bao:2020:AOS

[BC20] Y. Bao and T. E. Carlson.

Agile and open-source hardware. IEEE Micro, 40(4):6– 9, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Baker:2021:ETQ

[BC21] Jonathan M. Baker and Frederic T. Chong. Emerging technologies for quantum computing. IEEE Micro, 41(5):41– 47, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bose:1999:GEI

[BCA99] Pradip Bose, Thomas M.

Conte, and Todd M. Austin. Guest Editors’ introduction: Challenges in processor modeling and validation. IEEE Micro, 19(3):9–14, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3009.
pdf;   http://www.computer.
org/micro/mi1999/m3009abs.
htm.

Bharadwaj:2000:IIC

[BCC^{+}00] Jay Bharadwaj, William Y.

Chen, Weihaw Chuang, Gerolf Hoflehner, Kishore Menezes, Kalyan Muthukumar, and Jim Pierce. The Intel IA-64 compiler code generator. IEEE Micro, 20(5):44– 53, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5044.
pdf;   http://www.computer.
org/micro/mi2000/m5044abs.
htm.

Briggs:2002:IBB

[BCC^{+}02] Fayé Briggs, Michel Cekleov, Ken Creta, Manoj Khare, Steve Kulick, Akhilesh Kumar, Lily Pao Looi, Chitra Natarajan, Sivakumar Radhakrishnan, and Linda Rankin. Intel 870: a building block for cost-effective, scalable servers. IEEE Micro, 22(2):36–47, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2036.
pdf;   http://www.computer.
org/micro/mi2002/m2036abs.
htm.

Burgess:2011:BAL

[BCD^{+}11] Brad Burgess, Brad Cohen, Marvin Denman, Jim Dundas, David Kaplan, and Jeff Rupley. Bobcat: AMD’s lowpower x86 processor. IEEE Micro, 31(2):16–25, March/ April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bailey:1992:PVP

[BCF^{+}92] Doug Bailey, Matt Cressa, Jan Fandrianto, Doug Neubauer, Hedley Rainnie, and Chi-Shin Wang. Programmable vision processor/controller for flexible implementation of current and future image compression standards. IEEE Micro, 12(5): 33–39, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Boden:1995:MGP

[BCF^{+}95] N. J. Boden, D. Cohen, R. E.

Felderman, A. E. Kulawik, C. L. Seitz, J. N. Seizovic, and Wen-King K. Su. Myrinet — a gigabit-per-second localarea-network. IEEE Micro, 15 (1):29–36, January/February 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bouvier:2014:KAa

[BCF^{+}14] Dan Bouvier, Brad Cohen, Walter Fry, Sreekanth Godey, and Michael Mantor. Kabini: An AMD accelerated processing unit system on a chip. IEEE Micro, 34(2):22– 33, March/April 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Balaji:2006:BEE

[BcFP06] Pavan Balaji, Wu chun Feng, and Dhabaleswar K. Panda. Bridging the Ethernet–Ethernot performance gap. IEEE Micro, 26(3):24–40, May/June 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

                                        /csdl.computer.org/comp/
                                        mags/mi/2006/03/m3024.pdf.

Brehler:2023:MML

[BCH^{+}23] Marius Brehler, Lucas Camphausen, Benjamin Heidebroek, Dennis Krön, Henri Gründer, and Simon Camphausen. Making machine learning more energy efficient by bringing it closer to the sensor. IEEE Micro, 43(6): 11–18, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Balkind:2020:ONO

[BCJ^{+}20] J. Balkind, T. Chang, P. J. Jackson, G. Tziantzioulis, A. Li, F. Gao, A. Lavrov, G. Chirkov, J. Tu, M. Shahrad, and D. Wentzlaff. OpenPiton at 5: A nexus for open and agile hardware design. IEEE Micro, 40(4):22–31, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bong:2017:LPC

[BCKY17] Kyeongryeol Bong, Sungpill Choi, Changhyeon Kim, and Hoi-Jun Yoo. Lowpower convolutional neural network processor for a facerecognition system. IEEE Micro, 37(6):30–38, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

mi/2017/06/mmi2017060030-
abs.html.

Balasubramonian:2014:NDP

[BCM^{+}14] Rajeev Balasubramonian, Jichuan Chang, Troy Manning, Jaime H. Moreno, Richard Murphy, Ravi Nair, and Steven Swanson. Near-data processing: Insights from a MICRO-46 Workshop. IEEE Micro, 34 (4):36–42, July/August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040036-
abs.html.

Borrill:1995:HII

[BCN95] Paul L. Borrill, David E.

Culler, and Kathleen M. Nichols. Hot interconnects — introduction. IEEE Micro, 15(1):9–10, January/February 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Balmos:2022:IAF

[BCN^{+}22] Andrew D. Balmos, Fabio A.

Castiblanco, Aaron J. Neustedter, James V. Krogmeier, and Dennis R. Buckmaster. ISOBlue Avena: a framework for agricultural edge computing and data sovereignty. IEEE Micro, 42(1):78–86, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Benso:2001:SRE

[BCP01] Alfredo Benso, Silvia Chiusano, and Paolo Prinetto. A self-repairing execution unit for microprogrammed processors. IEEE Micro, 21(5):16– 22, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5016abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m5016.
pdf.

Bechini:2004:GEI

[BCP04] Alessio Bechini, Thomas M. Conte, and Cosimo Antonio Prete. Guest Editors’ introduction: Opportunities and challenges in embedded systems. IEEE Micro, 24(4):8– 9, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4008.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/04/m4008.
htm.

Bechem:1999:IFP

[BCU^{+}99] Candice Bechem, Jonathan Combs, Noppanunt Utamaphethai, Bryan Black, R. D. Shawn Blanton, and John Paul Shen. An integrated functional performance simulator. IEEE Micro, 19(3):26–35, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3026.
pdf;   http://www.computer.
org/micro/mi1999/m3026abs.
htm.

Boland:1994:PPP

[BD94] Keith Boland and Apostolos Dollas. Predicting and precluding problems with memory latency. IEEE Micro, 14 (4):59–67, July/August 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Blumrich:1995:VMM

[BDF^{+}95] Matthias A. Blumrich, Cezary Dubnicki, Edward W. Felten, Kai Li, and Malena R. Mesarina. Virtual-memory-mapped network interfaces. IEEE Micro, 15(1):21–28, January/February 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barroso:2003:WSP

[BDH03] Luiz André Barroso, Jeffrey Dean, and Urs Hölzle. Web search for a planet: The Google Cluster Architecture. IEEE Micro, 23 (2):22–28, March/April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m2022.
pdf;   http://www.computer.
org/micro/mi2003/m2022abs.
htm.

Binkert:2006:MSM

[BDH^{+}06] Nathan L. Binkert, Ronald G. Dreslinski, Lisa R. Hsu, Kevin T. Lim, Ali G. Saidi, and Steven K. Reinhardt. The M5 simulator: Modeling networked systems. IEEE Micro, 26(4):52–60, July/August 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Birrittella:2016:ESH

[BDH^{+}16] Mark S. Birrittella, Mark Debbage, Ram Huggahalli, James Kunz, Tom Lovett, Todd Rimmer, Keith D. Underwood, and Robert C. Zak. Enabling scalable highperformance systems with the Intel omni-path architecture. IEEE Micro, 36(4):38– 47, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/04/mmi2016040038-
                                        abs.html.

Brooks:2007:PTR

[BDJS07] David Brooks, Robert P. Dick, Russ Joseph, and Li Shang. Power, thermal, and reliability modeling in nanometer-scale microprocessors. IEEE Micro, 27(3):49–62, May/June 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barros:1998:SAI

[BdS98] Edna Barros and Marcus V. D. dos Santos. A safe, accurate intravenous infusion control system. IEEE Micro, 18(5):12–21, September/October 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m5012.pdf;   http://
www.computer.org/micro/
mi1998/m5012abs.htm.

Baker:2021:VLQ

[BDSC21] Jonathan M. Baker, Casey Duckering, David I. Schuster, and Frederic T. Chong. Virtual logical qubits: a compact architecture for fault-tolerant quantum computing. IEEE Micro, 41(3):95–101, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bougard:2008:CGA

[BDV^{+}08] Bruno Bougard, Bjorn De Sutter, Diederik Verkest, Liesbet Van der Perre, and Rudy Lauwereins. A coarse-grained array accelerator for softwaredefined radio baseband processing. IEEE Micro, 28 (4):41–50, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Beasley:1990:FPS

[Bea90] H. Beasley. The Futurebus+ protocol stack and profiles. IEEE Micro, 10(3):2– ??, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Beamer:2020:CAS

[Bea20] S. Beamer. A case for accelerating software RTL simulation. IEEE Micro, 40(4):112– 119, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Belopolsky:1993:IMM

[Bel93] Yakov Belopolsky. Interaction of multichip module substrates with high-density connectors. IEEE Micro, 13 (2):36–44, March/April 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bell:1996:SCM

[Bel96] G. Bell. The system-on-a-chip, microsystems computer industry. IEEE Micro, 16 (6):52, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Belgard:2012:YPR

[Bel12]  Rich Belgard.  Yale N.

Patt receives the inaugural IEEE B. Ramakrishna Rau Award. IEEE Micro, 32 (4):68–69, July/August 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Belgard:2013:AJJ

[Bel13] Rich Belgard. Awards: Joseph A. (Josh) Fisher receives the 2012 IEEE B. Ramakrishna Rau Award. IEEE Micro, 33(5):60–61, September/October 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Berger:2023:DTC

[BEL^{+}23] Daniel S. Berger, Daniel Ernst, Huaicheng Li, Pantea Zardoshti, Monish Shah, Samir Rajadnya, Scott Lee, Lisa Hsu, Ishwar Agarwal, Mark D. Hill, and Ricardo Bianchini. Design tradeoffs in CXL-based memory pools for public cloud platforms. IEEE Micro, 43(2):30–38, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Berman:1981:FAN

[Ber81] T. S. Berman. Forget algebraic notation. IEEE Micro, 1(4):3, October/December 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Berglund:1986:IVS

[Ber86] Eric J. Berglund. An introduction to the V-System. IEEE Micro, 6(4):35–52, July/August 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Berg:2009:MDC

[Ber09]  Thomas B. Berg.  Main-

taining I/O data coherence in embedded multicore systems. IEEE Micro, 29(3): 10–19, May/June 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bainbridge:2002:CDI

[BF02] John Bainbridge and Steve Furber. Chain: a delayinsensitive chip area interconnect. IEEE Micro, 22(5):16– 23, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5016.
pdf;   http://www.computer.
org/micro/mi2002/m5016abs.
htm.

Bezanson:1985:ESS

[BFK^{+}85] Llewellyn Bezanson, Louis G. Fields, Donald O. Knight, Michael J. Merritt, Bruce R. Millard, Donald S. Miller, and Peter R. Rony. Engineering support system user requirements. IEEE Micro, 5(5):36– 51, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bondavalli:2001:DVE

[BFLS01] Andrea Bondavalli, Alessandro Fantechi, Diego Latella, and Luca Simoncini. Design validation of embedded dependable systems. IEEE Micro, 21(5):52–62, September/October 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
m5052abs.htm;   http://
dlib.computer.org/mi/books/
mi2001/pdf/m5052.pdf.

Branover:2012:AFA

[BFS12] Alexander Branover, Denis Foley, and Maurice Steinman. AMD Fusion APU: Llano. IEEE Micro, 32(2):28– 37, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bao:2022:LLB

[BFZ^{+}22] Zhenshan Bao, Guohang Fu, Wenbo Zhang, Kang Zhan, and Junnan Guo. LSFQ: A low-bit full integer quantization for high-performance FPGA-based CNN acceleration. IEEE Micro, 42(2):8– 15, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bass:1981:EDI

[BG81] Steven C. Bass and Thomas W.

Goeddel. The efficient digital implementation of subtractive music synthesis. IEEE Micro, 1(3):24–37, July/September 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Buddefeld:2002:IMA

[BG02] Jürgen Büddefeld and Karl E. Grosspietsch. Intelligentmemory architecture for artificial neural networks. IEEE Micro, 22(3):32–40, May/ June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m3032.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m3032abs.
                                        htm.

Balasubramonian:2016:NDP

[BG16] Rajeev Balasubramonian and Boris Grot. Near-data processing [Guest Editors’ introduction]. IEEE Micro, 36(1): 4–5, January/February 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/01/mmi2016010004.
                                        html.

Bier:1990:GDE

[BGH^{+}90] Jeffrey C. Bier, Edwin E. Goei, Wai H. Ho, Philip D. Lapsley, Maureen P. O’Reilly, Gilbert C. Sih, and Edward A. Lee. Gabriel — a design environment for DSP. IEEE Micro, 10(5):28–45, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burgess:2012:EFL

[BGH^{+}12] David Burgess, Edmund Gieske, James Holt, Thomas Hoy, and Gary Whisenhunt. e6500: Freescale’s lowpower, high-performance multithreaded embedded processor. IEEE Micro, 32(5):26– 36, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burger:1997:LBA

[BGK97] Doug Burger, James R. Goodman, and Alain Kägi. Limited bandwidth to affect processor design. IEEE Micro, 17(6): 55–62, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m6055.
pdf;   http://www.computer.
org/micro/mi1997/m6055abs.
htm.

Bergsten:1988:ADA

[BGRKR88] Bjorn Bergsten, Ruben Gonzalez-Rubio, Brigitte Kerherve, and Jean Rohmer. An advanced database accelerator. IEEE Micro, 8(5):47–63, September/October 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Birmingham:1989:MSC

[BGS89] William P. Birmingham, Anurag P. Gupta, and Daniel P. Siewiorek. The Micon system for computer design. IEEE Micro, 9(5):61–67, September/October 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barkatullah:2015:GCF

[BH15] Javed Barkatullah and Timo Hanke. Goldstrike 1: Coin-Terra’s first-generation cryptocurrency mining processor for Bitcoin. IEEE Micro, 35 (2):68–76, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/02/mmi2015020068-
                                        abs.html.

Bhattacharjee:2017:PVM

[Bha17] Abhishek Bhattacharjee. Preserving virtual memory by mitigating the address translation wall. IEEE Micro, 37(5):6–10, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/05/mmi2017050006-
                                        abs.html.

Bhattacharjee:2018:BAT

[Bha18] Abhishek Bhattacharjee. Breaking the address translation wall by accelerating memory replays. IEEE Micro, 38(3): 69–78, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

mi/2018/03/mmi2018030069-
abs.html.

Bhattacharjee:2020:BSI

[Bha20]  A. Bhattacharjee.  Biol-

ogy and systems interactions. IEEE Micro, 40(5): 64, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Biswas:2000:SBS

[BHM^{+}00] Prasenjit Biswas, Atsushi Hasegawa, Srinivas Mandaville, Mark Debbage, Andy Sturges, Fumio Arakawa, Yasuhiko Saito, and Kunio Uchiyama. SH-5 — the 64-bit SuperH architecture. IEEE Micro, 20(4): 28–39, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4028.
pdf;   http://www.computer.
org/micro/mi2000/m4028abs.
htm.

Bojnordi:2013:PDC

[BI13] Mahdi Nazm Bojnordi and Engin Ipek. Programmable DDRx controllers. IEEE Micro, 33(3):106–115, May/June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bojnordi:2017:MBM

[BI17] Mahdi Nazm Bojnordi and Engin Ipek. The memristive Boltzmann machines.

IEEE Micro, 37(3):22–29, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030022-
abs.html.

Brebner:2014:HSP

[BJ14] Gordon Brebner and Weirong Jiang. High-speed packet processing using reconfigurable computing. IEEE Micro, 34(1):8–18, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Bae:2019:SSS

[BJG^{+}19] J. Bae, H. Jang, J. Gong, W. Jin, S. Kim, J. Jang, T. J. Ham, J. Jeong, and J. W. Lee. SSDStreamer: Specializing I/O stack for largescale machine learning. IEEE Micro, 39(5):73–81, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Batten:2009:BMC

[BJO^{+}09] Christopher Batten, Ajay Joshi, Jason Orcutt, Anatol Khilo, Benjamin Moss, Charles W. Holzwarth, Miloš A. Popović, Hanqing Li, Henry I. Smith, Judy L. Hoyt, Franz X. Kärtner, Rajeev J. Ram, Vladimir Stojanović, and Krste Asanović. Building many-core processor-to-DRAM networks with monolithic CMOS silicon photonics. IEEE Micro, 29(4):8– 21, July/August 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Berger:2023:ECM

[BJW^{+}23] Deanna Berger, Christian Jacobi, Craig R. Walters, Robert J. Sonnelitter, Mike Cadigan, and Matthias Klein. Enterprise-class multilevel cache design: Low latency, huge capacity, and high reliability. IEEE Micro, 43(1):58– 66, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bui:2014:CCM

[BK14] Van Bui and Martha A. Kim.

The cache and codec model for storing and manipulating data. IEEE Micro, 34(4):28– 35, July/August 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040028-
abs.html.

Best:1982:AAC

[BKM^{+}82] D. W. Best, C. E. Kress, N. M.

Mykris, J. D. Russell, and W. J. Smith. An advancedarchitecture CMOS/SOS microprocessor. IEEE Micro, 2(3):10–26, July/September 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burger:2012:CRC

[BKP12] Doug Burger, Stephen W. Keckler, and Mark Papermaster. Charles R. (Chuck) Moore (1961–2012). IEEE Micro, 32(4):3–5, July/August 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Batten:2023:SIT

[BL23] Christopher Batten and Jae W. Lee. Special issue on top picks from the 2022 computer architecture conferences. IEEE Micro, 43(4):6–10, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bornholt:2017:TDB

[BLC^{+}17] James Bornholt, Randolph Lopez, Douglas M. Carmean, Luis Ceze, Georg Seelig, and Karin Strauss. Toward a DNA-based archival storage system. IEEE Micro, 37(3): 98–104, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030098-
                                        abs.html.

Basoglu:2000:MVM

[BLO00] Chris Basoglu, Woobin Lee, and John Setel O’Donnell. The Map1000A VLIW mediaprocessor. IEEE Micro, 20 (2):48–59, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2048.
pdf;   http://www.computer.
org/micro/mi2000/m2048abs.
htm.

Braun:2002:PWL

[BLW02] Florian Braun, John Lockwood, and Marcel Waldvogel. Protocol wrappers for layered network packet processing in reconfigurable hardware. IEEE Micro, 22(1):66– 74, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1066abs.htm;
 http://dlib.computer.org/
mi/books/mi2002/pdf/m1066.
pdf.

Byles:1985:STA

[BM85] T. Byles and W. Myers.

Software tool AIDS problemsolving. IEEE Micro, 5(1):67– 71, January/February 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bekerman:1995:PAP

[BM95] Michael Bekerman and Avi Mendelson. A performance analysis of Pentium processor systems. IEEE Micro, 15(5): 72–83, September/October 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bose:2019:ESS

[BM19] P. Bose and S. Mukhopadhyay. Energy-Secure System Architectures (ESSA): A workshop report. IEEE Micro, 39(4):27–34, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bertozzi:2021:CEF

[BMG^{+}21] D. Bertozzi, G. Miorandi, A. Ghiribaldi, W. Burleson, G. Sadowski, K. Bhardwaj, W. Jiang, and S. M. Nowick. Cost-effective and flexible asynchronous interconnect technology for GALS systems. IEEE Micro, 41(1):69– 81, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Biancolin:2021:AFRa

[BMK^{+}21a] David Biancolin, Albert Magyar, Sagar Karandikar, Alon Amid, Borivoje Nikoli , Jonathan Bachrach, and Krste Asanovi . Accessible, FPGA resource-optimized simulation of multiclock systems in FireSim. IEEE Micro, 41 (4):58–66, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Biancolin:2021:AFRb

[BMK^{+}21b] David Biancolin, Albert Magyar, Sagar Karandikar, Alon Amid, Borivoje Nikoli , Jonathan Bachrach, and Krste Asanovi . Accessible, FPGA resource-optimized simulation of multiclock systems in FireSim. IEEE Micro, 41 (4):58–66, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Byun:2021:NGC

[BML^{+}21] Ilkwon Byun, Dongmoon Min, Gyuhyeon Lee, Seongmin Na, and Jangwoo Kim. A nextgeneration cryogenic processor architecture. IEEE Micro, 41(3):80–86, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bornholt:2015:UAU

[BMM15] James Bornholt, Todd Mytkowicz, and Kathryn S. McKinley. Uncertain<T>: Abstractions for uncertain hardware and software. IEEE Micro, 35 (3):132–143, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030132-
abs.html.

Balasubramonian:2006:LWP

[BMR^{+}06] Rajeev Balasubramonian, Naveen Muralimanohar, Karthik Ramani, Liqun Cheng, and John B. Carter. Leveraging wire properties at the microarchitecture level. IEEE Micro, 26(6):40–52, November/December 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Banerjee:2016:SNA

[BMS16] Amitabha Banerjee, Rishi Mehta, and Zach Shen. Supporting NUMA-aware I/O in virtual machines. IEEE Micro, 36(4):28–36, July/ August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/04/mmi2016040028-
                                        abs.html.

Bobba:2008:PPH

[BMV^{+}08] Jayaram Bobba, Kevin E. Moore, Haris Volos, Luke Yen, Mark D. Hill, Michael M. Swift, and David A. Wood. Performance pathologies in hardware transactional memory. IEEE Micro, 28(1):32–41, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bronnenberg:1987:DDO

[BNOv87] Wim J. H. J. Bronnenberg, Loek Nijman, Eddy A. M. Odijk, and Rob A. H. van Twist. DOOM — a Decentralized Object-Oriented Machine. IEEE Micro, 7(5):52– 69, September/October 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bolotin:2015:DEH

[BNV^{+}15] Evgeny Bolotin, David Nellans, Oreste Villa, Mike O’Connor, Alex Ramirez, and Stephen W. Keckler. Designing efficient heterogeneous memory architectures. IEEE Micro, 35(4):60– 68, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040060-
abs.html.

Butler:1986:FSM

[BO86] James M. Butler and A. Yavuz Oruc. A facility for simulating multiprocessors. IEEE Micro, 6(5):32–44, September/October 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Boahen:1996:RVS

[Boa96]  Kwabena A. Boahen.  A

retinomorphic vision system — mimicking the structure and function of the vertebrate retina to optimize the information-gathering capacity of machine-vision systems. IEEE Micro, 16(5):30– 39, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bondi:2021:SMM

[Bon21]  James O. Bondi.  Spe-

cial memories from my favorite TI microprocessor design project. IEEE Micro, 41 (6):155, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Borrill:1981:MBS

[Bor81] Paul L. Borrill. Microprocessor bus structures and standards. IEEE Micro, 1(1):84–95, January/March 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Reprinted from Microprocessors — Software, Firmware and Hardware, 1980.

Borill:1985:BBS

[Bor85a] P. L. Borill. The 32-bit bus standards issue revisited. IEEE Micro, 5(5):76– 84, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Borrill:1985:MSF

[Bor85b] Paul L. Borrill. MicroStandards special feature — a comparison of 32-bit buses. IEEE Micro, 5(6):71–79, November/December 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Borel:1999:DAM

[Bor99a] Joseph Borel. Design automation in MEDEA: Present and future. IEEE Micro, 19(5):71– 79, September/October 1999.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5071.
pdf;   http://www.computer.
org/micro/mi1999/m5071abs.
htm.

Borkar:1999:DCT

[Bor99b] Shekhar Borkar. Design challenges of technology scaling. IEEE Micro, 19(4): 23–29, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4023.
pdf;   http://www.computer.
org/micro/mi1999/m4023abs.
htm.

Borkar:2005:DRS

[Bor05] Shekhar Borkar. Designing reliable systems from unreliable components: The challenges of transistor variability and degradation. IEEE Micro, 25(6):10–16, November/December 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bose:2003:DIC

[Bos03a] Pradip Bose. Design and integration: Chip- and systemlevel challenges. IEEE Micro, 23(3):5, May/June 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.

                                        computer.org/dl/mags/mi/
                                        2003/03/m3005.pdf.

Bose:2003:ECM

[Bos03b]  Pradip Bose.  Editor-in-

Chief’s message: Adapting old paradigms to meet new challenges. IEEE Micro, 23(4): 5, July/August 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/04/m4005.pdf.

Bose:2003:EMI

[Bos03c] Pradip Bose. EIC’s message: Issues and trends in highperformance processor cores. IEEE Micro, 23(2):5, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m2005.
                                        pdf.

Bose:2003:EML

[Bos03d] Pradip Bose. EIC’s message: Looking forward to bright new beginnings. IEEE Micro, 23(1):5–7, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m1005.
                                        pdf;   http://www.computer.
                                        org/micro/mi2003/m1005abs.
                                        htm.

Bose:2004:CVC

[Bos04a]  Pradip Bose.  Communi-

cation versus computation.

IEEE Micro, 24(5):5, September/October 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/05/m5005.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/05/m5005.
htm.

Bose:2004:ECM

[Bos04b] Pradip Bose. Editor in Chief’s message: Saving powerlessons from embedded systems. IEEE Micro, 24(4):5– 6, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4005.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/04/m4005.
htm.

Bose:2004:EMCa

[Bos04c] Pradip Bose. EIC’s message:

Chip-level microarchitecture trends. IEEE Micro, 24(2): 5, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/02/m2005.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/02/m2005.
htm.

Bose:2004:EMCb

[Bos04d] Pradip Bose. EIC’s message: Computer architecture research: Shifting priorities and newer challenges. IEEE Micro, 24(6):5, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/06/m6005.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2004/06/m6005.
                                        htm.

Bose:2004:EMG

[Bos04e] Pradip Bose. EIC’s message: General-purpose versus application-specific processors. IEEE Micro, 24 (3):5, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/03/m3005.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2004/03/m3005.
                                        htm.

Bose:2004:NCB

[Bos04f]  Pradip Bose.  New chal-

lenges and burning issues. IEEE Micro, 24(1):5, January/February 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/01/m1005.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2004/01/m1005.
                                        htm.

Bose:2005:EMD

[Bos05a] Pradip Bose. EIC’s message: Designing microprocessors with robust functionality and performance. IEEE Micro, 25(6):5, November/December 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/06/m6005.pdf.

Bose:2005:EMH

[Bos05b] Pradip Bose. EIC’s message: High performance at affordable power. IEEE Micro, 25(5):5, September/October 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/05/m5005.pdf.

Bose:2005:EMI

[Bos05c] Pradip Bose. EIC’s message: Integrated microarchitectures. IEEE Micro, 25(3): 5–6, May/June 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/03/m3005.pdf.

Bose:2005:EMPb

[Bos05d] Pradip Bose. EIC’s message: Presilicon modeling: challenges in the late CMOS era. IEEE Micro, 25(4):5– 6, July/August 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.

computer.org/comp/mags/
mi/2005/04/m4005.pdf.

Bose:2005:EMPa

[Bos05e] Pradip Bose. EIC’s message: The “power” of communication. IEEE Micro, 25(1):5, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/01/m1005.pdf;
http://csdl.computer.org/
dl/mags/mi/2005/01/m1005.
htm.

Bose:2005:EMV

[Bos05f] Pradip Bose. EIC’s message: Variation-tolerant design. IEEE Micro, 25(2): 5, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2005.pdf;
http://csdl.computer.org/
comp/mags/mi/2005/02/m2005abs.
htm.

Bose:2006:ECMd

[Bos06a]  Pradip Bose.  Editor-in-

Chief’s message: Designing reliable systems with unreliable components. IEEE Micro, 26(5):5–6, September/October 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/05/m5005.pdf.

Bose:2006:ECMe

[Bos06b]  Pradip Bose.  Editor-in-

Chief’s message: Looking briefly back, and then forward . . .. IEEE Micro, 26(6):8– 9, November/December 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/06/m6008.pdf.

Bose:2006:ECMc

[Bos06c] Pradip Bose. Editor-in-Chief’s message: Pre-silicon modeling and analysis: Impact on real design. IEEE Micro, 26 (4):3, July/August 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/04/m4003.pdf.

Bose:2006:ECMb

[Bos06d] Pradip Bose. Editor-in-Chief’s message: Robust on-chip communication. IEEE Micro, 26 (3):5, May/June 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/03/m3005.pdf.

Bose:2006:ECMa

[Bos06e]  Pradip Bose.  Editor-in-

Chief’s message: Workload characterization: a key aspect of microarchitecture design. IEEE Micro, 26(2):5– 6, March/April 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/02/m2005.pdf.

Bose:2006:EMM

[Bos06f] Pradip Bose. EIC’s message: Measuring the impact of microarchitectural ideas. IEEE Micro, 26(1):5–6, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/01/m1005.pdf.

Bose:2021:PPF

[Bos21] Pradip Bose. The POWER processor family: A historical perspective from the viewpoint of presilicon modeling. IEEE Micro, 41(6): 71–77, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Benkner:2011:PEP

[BPT^{+}11] Siegfried Benkner, Sabri Pllana, Jesper Larsson Träff, Philippas Tsigas, Uwe Dolinsky, Cédric Augonnet, Beverly Bachmayer, Christoph Kessler, David Moloney, and Vitaly Osipov. PEPPHER: Efficient and productive usage of hybrid computing systems. IEEE Micro, 31(5):28– 41, September/October 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brightwell:2006:SIB

[BPUH06] Ron Brightwell, Kevin T.

Pedretti, Keith D. Underwood, and Trammell Hudson. SeaStar interconnect: Balanced bandwidth for scalable performance. IEEE Micro, 26(3):41–57, May/June 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barroso:2010:GEI

[BR10] Luiz Andre Barroso and Parthasarathy Ranganathan. Guest Editors’ introduction: Datacenter-scale computing. IEEE Micro, 30(4): 6–7, July/August 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

BenHaim:2021:RTE

[BR21] R. Ben Haim and O. Rottenstreich. Reliable and timeefficient virtualized function placement. IEEE Micro, 41 (1):45–53, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brewer:2010:ISI

[Bre10] Tony M. Brewer. Instruction set innovations for the Convey HC-1 computer. IEEE Micro, 30(2):70–79, March/April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Briggs:1994:WWC

[Bri94] B. Briggs. The war of the words continues. IEEE Micro, 14(2):2–??, March/April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barnes:2006:TCM

[BRmWH06] Ronald D. Barnes, Shane Ryoo, and Wen mei W. Hwu. Tolerating cache-miss latency with multipass pipelines. IEEE Micro, 26(1):40–47, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bronson:1986:CR

[Bro86] R. Bronson. CRC revisited. IEEE Micro, 6(4): 79, July/August 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brooks:2011:CGH

[Bro11] David Brooks. CPUs, GPUs, and hybrid computing. IEEE Micro, 31(5):4–6, September/October 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brooks:2017:ISC

[Bro17] David Brooks. 2017 International Symposium on Computer Architecture Influential Paper Award. IEEE Micro, 37(6):90–91, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060090.
html.

Borrill:1984:SGE

[BS84] P. Borrill and R. G. Stewart.

Standards — Guest Editors’ introduction. IEEE Micro, 4(4):3–6, July/August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burman:1993:PMM

[BS93] Surendra Burman and Naveed A.

Sherwani. Programmable multichip modules. IEEE Micro, 13(2):28–35, March/April 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Baum:1998:GEI

[BS98] Allen J. Baum and Alan Jay Smith. Guest Editors’ introduction: Hot chips—hot stuff. IEEE Micro, 18(2): 11–13, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2011.
pdf.

Brooks:2017:ULP

[BS17] David Brooks and John Sartori. Ultra-low-power processors. IEEE Micro, 37(6):16–19, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060016.
html.

Bertels:2021:QCN

[BSA21] Koen Bertels, Aritra Sarkar, and Imran Ashraf. Quantum computing from NISQ to PISQ. IEEE Micro, 41(5):24– 32, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Boser:1992:HRN

[BSB^{+}92] Bernhard E. Boser, Eduard Sackinger, Jane Bromley, Yann leCun, and Lawrence D. Jackel. Hardware requirements for neural network pattern classifiers: a case study and implementation. IEEE Micro, 12(1):32–40, January/February 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Birman:1990:DWS

[BSC^{+}90] Mark Birman, Allen Samuels, George Chu, Ting Chuk, Larry Hu, John McLeod, and John Barnes. Developing the WTL3170/3171 Sparc floating-point coprocessors. IEEE Micro, 10(1):55– 64, January/February 1990.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bower:2008:IDH

[BSC08] Fred A. Bower, Daniel J.

Sorin, and Landon P. Cox. The impact of dynamically heterogeneous multicore processors on thread scheduling. IEEE Micro, 28(3):17– 25, May/June 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bohnenstiehl:2017:KFG

[BSP^{+}17] Brent Bohnenstiehl, Aaron Stillmaker, Jon Pimentel, Timothy Andreas, Bin Liu, Anh Tran, Emmanuel Adeagbo, and Bevan Baas. KiloCore: A fine-grained 1,000-processor array for task-parallel applications. IEEE Micro, 37(2):63– 69, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020063-
abs.html.

Bertels:2010:HHS

[BSY^{+}10] Koen Bertels, Vlad-Mihai Sima, Yana Yankova, Georgi Kuzmanov, Wayne Luk, Gabriel Coutinho, Fabrizio Ferrandi, Christian Pilato, Marco Lattuada, Donatella Sciuto, and Andrea Michelotti. HArtes: Hardwaresoftware codesign for heterogeneous multicore platforms. IEEE Micro, 30(5):88– 97, September/October 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Borrill:1984:ACP

[BT84] Paul Borrill and John Theus. An advanced communication protocol for the proposed IEEE 896 Futurebus. IEEE Micro, 4(4):42– 56, July/August 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brauch:1992:AVN

[BTHS92] Jeff Brauch, Simon M. Tam, Mark A. Holler, and Arthur L. Shmurun. Analog VLSI neural networks for impact signal processing. IEEE Micro, 12(6):34–45, November/December 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Behnam:2023:HSC

[BTK^{+}23] Payman Behnam, Jianming Tong, Alind Khare, Yangyu Chen, Yue Pan, Pranav Gadikar, Abhimanyu Bambhaniya, Tushar Krishna, and Alexey Tumanov. Hardware software co-design for realtime latency accuracy navigation in tiny machine learning applications. IEEE Micro, 43(6):93–101, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bossen:2002:PSD

[BTR02] Douglas C. Bossen, Joel M.

Tendler, and Kevin Reick. Power4 system design for high reliability. IEEE Micro, 22 (2):16–24, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2016.
pdf;   http://www.computer.
org/micro/mi2002/m2016abs.
htm.

Buckley:1984:ISE

[Buc84] Fletcher J. Buckley. The IEEE Software engineering standards process. IEEE Micro, 4(4):57–62, July/August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Buckley:1985:AMB

[Buc85]  F. J. Buckley.  August

Microstandards brings comments. IEEE Micro, 5(5):4, 93, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Buckley:1987:SL

[Buc87] F. J. Buckley. Standards and lances. IEEE Micro, 7(1):2, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burns:1995:ONN

[BUMV95] D. C. Burns, I. Underwood, A. F. Murray, and D. G. Vass. An optoelectronic neural network with temporally multiplexed gray-scale weights. IEEE Micro, 15(3): 49–51, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burgess:1996:WRP

[Bur96] B. Burgess. What RISC penalty? IEEE Micro, 16 (1):79–80, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burgess:2020:RNT

[Bur20] J. Burgess. RTX on the NVIDIA Turing GPU. IEEE Micro, 40(2):36–44, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Busigin:1986:FSI

[Bus86] A. Busigin. FFT software for the IBM PC. IEEE Micro, 6 (1):6, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Butts:2007:STC

[But07] Mike Butts. Synchronization through communication in a massively parallel processor array. IEEE Micro, 27(5):32– 40, September/October 2007.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Burres:2015:IAC

[BvdGM^{+}15] Bradley Burres, Johan van de Groenendaal, Praveen Mosur, Jonathan Robinson, Ian Steiner, Yi-Feng Liu, Sin S. Tan, Erik McShane, Belliappa Kuttanna, and Sridhar Lakshmanamurthy. Intel Atom C2000 processor family: Power-efficient datacenter processing. IEEE Micro, 35 (2):26–34, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020026-
abs.html.

Bridges:2008:RSP

[BVZ^{+}08] Matthew J. Bridges, Neil Vachharajani, Yun Zhang, Thomas Jablin, and David I. August. Revisiting the sequential programming model for the multicore era. IEEE Micro, 28(1):12–20, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brown:2011:IPE

[BWBJ11] Jeffrey D. Brown, Sandra Woodward, Brian M. Bass, and Charles L. Johnson. IBM Power Edge of Network processor: a wire-speed system on a chip. IEEE Micro, 31 (2):76–85, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bishop:2019:MDI

[BWMS19] M. D. Bishop, H.-P. Wong, S. Mitra, and M. M. Shulaker. Monolithic 3-D integration. IEEE Micro, 39(6): 16–27, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Boles:2023:CEE

[BWR23] David Boles, Daniel Waddington, and David A. Roberts. CXL-Enabled enhanced memory functions. IEEE Micro, 43 (2):58–65, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bink:2007:AFL

[BY07] Arjan Bink and Richard York. ARM996HS: The first licensable, clockless 32-bit processor core. IEEE Micro, 27(2):58– 68, March/April 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bohr:2017:CST

[BY17] Mark T. Bohr and Ian A. Young. CMOS scaling trends and beyond. IEEE Micro, 37(6):20–29, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060020-
abs.html.

Baas:2007:AFG

[BYM^{+}07] Bevan Baas, Zhiyi Yu, Michael Meeuwsen, Omar Sattari, Ryan Apperson, Eric Work, Jeremy Webb, Michael Lai, Tinoosh Mohsenin, Dean Truong, and Jason Cheung. AsAP: a fine-grained manycore platform for DSP applications. IEEE Micro, 27(2):34– 45, March/April 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Curtis:1986:CPL

[CAH86] T. W. Curtis, Paul Allison, and James A. Howard. A CORDIC processor for laser trimming. IEEE Micro, 6(3): 61–71, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Caianiello:1989:TSW

[Cai89] Eduardo R. Caianiello. Is there a silicon way to intelligence? IEEE Micro, 9(6): 75–76, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cangellaris:1998:EMS

[Can98] Andreas C. Cangellaris. Electrical modeling and simulation challenges in chip-package codesign. IEEE Micro, 18 (4):50–59, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4050.
pdf;   http://www.computer.
org/micro/mi1998/m4050abs.
htm.

Carey:1993:TLC

[Car93]  David H. Carey.  Trends

in low-cost, high-performance substrate technology. IEEE Micro, 13(2):19–27, March/ April 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cargill:1998:SAD

[Car98] Carl F. Cargill. Standardization — art or discipline. IEEE Micro, 18(3):18–24, May/ June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3018.
pdf;   http://www.computer.
org/micro/mi1998/m3018abs.
htm.

Castelli:1995:GEI

[Cas95] Gianluigi Castelli. Guest Editor’s introduction: The seemingly unlimited market for microcontroller-based embedded systems. IEEE Micro, 15 (5):6–8, September/October 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cascaval:2015:SIM

[Cas15]  Calin Cascaval.  Special

issue on mobile systems. IEEE Micro, 35(1):4–5, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010004-
abs.html.

Cates:1988:PAC

[Cat88] Ron Cates. Processor architecture considerations for embedded controller applications. IEEE Micro, 8(3):28– 38, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Codrescu:2014:HDA

[CAV^{+}14] Lucian Codrescu, Willie Anderson, Suresh Venkumanhanti, Mao Zeng, Erich Plondke, Chris Koob, Ajay Ingle, Charles Tabony, and Rick Maule. Hexagon DSP: An architecture optimized for mobile multimedia and communications. IEEE Micro, 34 (2):34–43, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Cheshire:1996:WNM

[CB96] Stuart Cheshire and Mary Baker. A wireless network in MosquitoNet. IEEE Micro, 16 (1):44–52, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Carlstrom:2004:SDA

[CB04] Jakob Carlström and Thomas Bodén. Synchronous dataflow architecture for network processors. IEEE Micro, 24(5): 10–18, September/October 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/mi/2004/05/m5010.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2004/05/m5010.
                                        pdf.

Calhoun:2010:CSN

[CB10] Benton H. Calhoun and David Brooks. Can subthreshold and near-threshold circuits go mainstream? IEEE Micro, 30 (4):80–85, July/August 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cooper-Balis:2010:FGA

[CBJ10] Elliott Cooper-Balis and Bruce Jacob. Fine-grained activation for power reduction in DRAM. IEEE Micro, 30(3): 34–47, May/June 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cooper:1986:BCB

[CBLR86] Thayne C. Cooper, Wayne D. Bell, Frank C. Lin, and Norm J. Rasmussen. A benchmark comparison of 32-bit microprocessors. IEEE Micro, 6 (4):53–58, July/August 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Celio:2019:BOS

[CCA^{+}19] C. Celio, P. Chiu, K. Asanovic, B. Nikolic, and D. Patterson. BROOM: An open-source out-of-order processor with resilient low-voltage operation in 28-nm CMOS. IEEE Micro, 39(2):52–60, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chiueh:2000:CMD

[cCCP00] Tzi cker C. Chiueh and Prashant Pradhan. Cache memory design for Internet processors. IEEE Micro, 20 (1):28–33, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1028.
pdf.

Civera:1982:MPE

[CCD^{+}82] P. Civera, G. Conte, D. Del Corso, F. Gregoretti, and E. Pasero. The mu * project: an experience with a multimicroprocessor system. IEEE Micro, 2(2):38–50, April/June 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chaudhry:2009:RHP

[CCE^{+}09] Shailender Chaudhry, Robert Cypher, Magnus Ekman, Martin Karlsson, Anders Landin, Sherman Yip, Håkan Zeffer, and Marc Tremblay. Rock: a high-performance SPARC CMT processor. IEEE Micro, 29(2):6–16, March/April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cody:1984:PRW

[CCG^{+}84] William J. Cody, Jr., Jerome T. Coonen, David M. Gay, K. Hanson, David G. Hough, William Kahan, Richard Karpinski, John F. Palmer, Frederic N. Ris, and David Stevenson. A proposed radixand word-length-independent standard for floating-point arithmetic. IEEE Micro, 4 (4):86–100, July/August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Caulfield:2017:CC

[CCP^{+}17] Adrian M. Caulfield, Eric S. Chung, Andrew Putnam, Hari Angepat, Daniel Firestone, Jeremy Fowers, Michael Haselman, Stephen Heil, Matt Humphrey, Puneet Kaur, Joo-Young Kim, Daniel Lo, Todd Massengill, Kalin Ovtcharov, Michael Papamichael, Lisa Woods, Sitaram Lanka, Derek Chiou, and Doug Burger. Configurable clouds. IEEE Micro, 37(3):52–61, May/ June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

https://www.computer.org/
csdl/mags/mi/2017/03/mmi2017030052-
abs.html.

Christie:2021:WMU

[CCS21] Dave Christie, Mike Clark, and Mike Schulte. What made us stronger: An inside look back at the history of AMD microprocessor development. IEEE Micro, 41(6): 29–36, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chaudhry:2005:HPT

[CCYT05] Shailender Chaudhry, Paul Caprioli, Sherman Yip, and Marc Tremblay. Highperformance throughput computing. IEEE Micro, 25(3): 32–45, May/June 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cekleov:1997:VACa

[CD97a] Michel Cekleov and Michel Dubois. Virtual-address caches, part 1: Problems and solutions in uniprocessors. IEEE Micro, 17(5):64– 71, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5064.

                                        pdf;   http://www.computer.
                                        org/micro/mi1997/m5064abs.
                                        htm.

Cekleov:1997:VACb

[CD97b] Michel Cekleov and Michel Dubois. Virtual-address caches, part 2: Multiprocessor issues. IEEE Micro, 17(6): 69–74, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1997/pdf/m6069.
                                        pdf;   http://www.computer.
                                        org/micro/mi1997/m6069abs.
                                        htm.

Chrysos:2009:PHT

[CD09] Nikos Chrysos and Giorgos Dimitrakopoulos. Practical high-throughput crossbar scheduling. IEEE Micro, 29 (4):22–35, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cook:2023:TAB

[CDBY23] Jack Cook, Jules Drean, Jonathan Behrens, and Mengjia Yan. There’s always a bigger fish: a clarifying analysis of a machine-learning-assisted side-channel attack. IEEE Micro, 43(4):28–36, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Costa:1997:FLM

[CDGO97] Alessandra Costa, Alessandro De Gloria, Fabrizio Giudici, and Mauro Olivieri. Fuzzy logic microcontroller: Limiting system costs for fuzzy control applications by optimizing data structures and hardware. IEEE Micro, 17(1):66– 74, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chen:2007:SIC

[CDS07] Jianwei Chen, Michel Dubois, and Per Stenström. SimWattch: Integrating complete-system and user-level performance and power simulators. IEEE Micro, 27(4):34–48, July/ August 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chen:2015:HTN

[CDS^{+}15] Tianshi Chen, Zidong Du, Ninghui Sun, Jia Wang, Chengyong Wu, Yunji Chen, and Olivier Temam. A highthroughput neural network accelerator. IEEE Micro, 35(3): 24–32, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030024-
abs.html.

Cherupalli:2018:BPA

[CDY^{+}18] Hari Cherupalli, Henry Duwe, Weidong Ye, Rakesh Kumar, and John Sartori. Bespoke processors for applications with ultra-low area and power constraints. IEEE Micro, 38 (3):32–39, May/June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/03/mmi2018030032-
                                        abs.html.

Chowdhuryy:2023:UCS

[CEAY23] Md Hafizul Islam Chowdhuryy, Rickard Ewetz, Amro Awad, and Fan Yao. Understanding and characterizing side channels exploiting phase-change memories. IEEE Micro, 43(5):8–15, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chen:2012:IBG

[CEH^{+}12] Dong Chen, Noel A. Eisley, Philip Heidelberger, Robert M. Senger, Yutaka Sugawara, Sameer Kumar, Valentina Salapura, David Satterfield, Burkhard Steinmacher-Burow, and Jeffrey Parker. The IBM Blue Gene/Q interconnection fabric. IEEE Micro, 32(1):32–43, January/February 2012. CO-DEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Circello:1995:SAM

[CEM^{+}95] Joe Circello, Greg Edgington, Dan McCarthy, James Gay, David Schimke, Steven Sullivan, Richard Duerden, Chris Hinds, Danny Marquette, Lal Sood, Al Crouch, and Daniel Chow. The superscalar architecture of the MC68060. IEEE Micro, 15(2):10–21, March/ April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Chua:2017:VIU

[CEP^{+}17] Vui Seng Chua, Julio Zamora Esquivel, Anindya S. Paul, Thawee Techathamnukool, Carlos Flores Fajardo, Nilesh Jain, Omesh Tickoo, and Ravi Iyer. Visual IoT: Ultralow-power processing architectures and implications. IEEE Micro, 37(6):52–61, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060052-
abs.html.

Curran:2011:ZSM

[CES^{+}11] Brian W. Curran, Lee E.

Eisen, Eric M. Schwarz, Pak kin Mak, James Warnock, Patrick J. Meaney, and Michael Fee. The zEnterprise 196 system and microprocessor. IEEE Micro, 31(2):26– 40, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chen:2017:UDO

[CES17] Yu-Hsin Chen, Joel Emer, and Vivienne Sze. Using dataflow to optimize energy efficiency of deep neural network accelerators. IEEE Micro, 37(3): 12–21, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030012-
                                        abs.html.

Chamberlain:1990:HDE

[CF90] Roger D. Chamberlain and Mark A. Franklin. Hierarchical discrete-event simulation on hypercube architectures. IEEE Micro, 10(4):10– 20, July/August 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Catanzaro:2010:UPC

[CFK^{+}10] Bryan Catanzaro, Armando Fox, Kurt Keutzer, David Patterson, Bor-Yiing Su, Marc Snir, Kunle Olukotun, Pat Hanrahan, and Hassan Chafi. Ubiquitous parallel computing from Berkeley, Illinois, and Stanford. IEEE Micro, 30 (2):41–55, March/April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cramer:1997:CJJ

[CFM^{+}97] Timothy Cramer, Richard Friedman, Terrence Miller, David Seherger, Robert Wilson, and Mario Wolczko. Compiling Java just in time: Using runtime compilation to improve Java program performance. IEEE Micro, 17 (3):36–43, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3036.
pdf.

Chung:2018:SDR

[CFO^{+}18] Eric Chung, Jeremy Fowers, Kalin Ovtcharov, Michael Papamichael, Adrian Caulfield, Todd Massengill, Ming Liu, Daniel Lo, Shlomi Alkalay, Michael Haselman, Maleen Abeydeera, Logan Adams, Hari Angepat, Christian Boehn, Derek Chiou, Oren Firestein, Alessandro Forin, Kang Su Gatlin, Mahdi Ghandi, Stephen Heil, Kyle Holohan, Ahmad El Husseini, Tamas Juhasz, Kara Kagi, Ratna Kovvuri, Sitaram Lanka, Friedel van Megen, Dima Mukhortov, Prerak Patel, Brandon Perez, Amanda Rapsang, Steven Reinhardt, Bita Rouhani, Adam Sapek, Raja Seera, Sangeetha Shekar, Balaji Sridharan, Gabriel Weisz, Lisa Woods, Phillip Yi Xiao, Dan Zhang, Ritchie Zhao, and Doug Burger. Serving DNNs in real time at datacenter scale with Project Brainwave. IEEE Micro, 38 (2):8–20, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/02/mmi2018020008-
                                        abs.html.

Chu:2004:CSP

[CFRM04] Michael L. Chu, Kevin C. Fan, Rajiv A. Ravindran, and Scott A. Mahlke. Costsensitive partitioning in an architecture synthesis system for multicluster processors. IEEE Micro, 24(3):10– 20, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/03/m3010.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/03/m3010.pdf.

Chen:1999:ROT

[CFZ^{+}99] Xiao-Tao T. Chen, Wenyi Y. Feng, Jun Zhao, Fred J. Meyer, and Fabrizio Lombardi. Reconfiguring onetime programmable FPGAs. IEEE Micro, 19(6):53–63, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m6053.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m6053abs.
                                        htm.

Cosatto:1995:NNA

[CG95] Eric Cosatto and Hans Peter Graf. A neural network accelerator for image analysis — analyzing complex images by decomposing them into a hierarchy of simple, “buildingblock” shapes. IEEE Micro, 15(3):32–38, May/June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Choquette:2018:VPP

[CGF18] Jack Choquette, Olivier Giroux, and Denis Foley. Volta: Performance and programmability. IEEE Micro, 38(2):42– 52, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/02/mmi2018020042-
abs.html.

Choquette:2021:NAT

[CGG^{+}21] J. Choquette, W. Gandhi, O. Giroux, N. Stam, and R. Krashinsky. NVIDIA A100 Tensor Core GPU: Performance and innovation. IEEE Micro, 41(2):29–35, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chiodo:1994:HSC

[CGJ^{+}94] Massimiliano Chiodo, Paolo Giusto, Attila Jurecska, Harry C. Hsieh, Alberto Sangiovanni-Vincentelli, and Luciano Lavagno. Hardware-software codesign of embedded systems. IEEE Micro, 14(4):26– 36, July/August 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cascajo:2023:MIN

[CGLES^{+}23] Alberto Cascajo, Gabriel Gomez-Lopez, Jesus Escudero-Sahuquillo, Pedro Javier Garcia, David E. Singh, Francisco Alfaro-Cortés, Francisco J. Quiles, and Jesus Carretero. Monitoring InfiniBand networks to react efficiently to congestion. IEEE Micro, 43(2):120–130, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Choquette:1999:HPR

[CGMV99] Jack Choquette, Mayank Gupta, Dominic McCarthy, and Jack Veenstra. Highperformance RISC microprocessors. IEEE Micro, 19 (4):48–55, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m4048.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m4048abs.
                                        htm.

Cranor:2000:ACC

[CGO00] Charles D. Cranor, R. Gopalakrishnan, and Peter Z. Onufryk. Architectural considerations for CPU and network interface integration. IEEE Micro, 20 (1):18–26, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1018.
pdf;   http://www.computer.
org/micro/mi2000/m1018abs.
htm.

Caulfield:2010:GIA

[CGS10] Adrian M. Caulfield, Laura M.

Grupp, and Steven Swanson. Gordon: An improved architecture for dataintensive applications. IEEE Micro, 30(1):121–130, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Clapp:1994:CMU

[CH94] Alan E. Clapp and Thomas L.

Harman. Combining microcontroller units and PLDs for best system design. IEEE Micro, 14(2):70–78, March/April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Conway:2007:AON

[CH07] Pat Conway and Bill Hughes.

The AMD Opteron Northbridge architecture. IEEE Micro, 27(2):10–21, March/April 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cain:1985:RRS

[CHA^{+}85a] T. Cain, D. Hannum, J. Aylor, B. Stewart, and D. Jaeger.

Review referee states concerns. IEEE Micro, 5(5):5– ??, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chafee:1985:CEH

[Cha85b]  J. H. Chafee.  Congress

enacts its high-tech agenda. IEEE Micro, 5(1):3–6, January/February 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Charlesworth:1998:SES

[Cha98] A. Charlesworth. Starfire — extending the SMP envelope. IEEE Micro, 18(1):39– 49, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Charlesworth:2002:SFI

[Cha02]  Alan Charlesworth.  The

Sun Fireplane interconnect. IEEE Micro, 22(1):36–45, January/February 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2002/
                                        m1036abs.htm;   http://
                                        dlib.computer.org/mi/books/
                                        mi2002/pdf/m1036.pdf.

Chien:2022:UAJ

[CHAF22] Wei-Che Chien, Mohammad Mehedi Hassan, Ahmed Alsanad, and Giancarlo Fortino. UAV assisted joint wireless power transfer and data collection mechanism for sustainable precision agriculture in 5G. IEEE Micro, 42(1):25–32, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chen:2019:RFC

[Che19]  Y. Chen.  Reshaping fu-

ture computing systems with emerging nonvolatile memory technologies. IEEE Micro, 39 (1):54–57, January/February 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chernoff:1998:FPD

[CHH^{+}98] Anton Chernoff, Mark Herdeg, Ray Hookway, Chris Reeve, Norman Rubin, Tony Tye, S. Bharadwaj Yadavalli, and John Yates. FX!32: a profile-directed binary translator. IEEE Micro, 18(2): 56–64, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2056.
pdf;   http://www.computer.
org/micro/mi1998/m2056abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24– 26, 1997.

Chodorek:2021:NVI

[Cho21] Robert Ryszard Chodorek.

NEC V20: Inspiring, inconspicuous. IEEE Micro, 41(6):158–159, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Choquette:2023:NHH

[Cho23] Jack Choquette. NVIDIA Hopper H100 GPU: Scaling performance. IEEE Micro, 43(3):9–17, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chrzaszcz:1990:OPC

[Chr90] J. Chrzaszcz. One person’s cup of tea is . . .. IEEE Micro, 10(1):93, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chrzaszcz:1991:PS

[Chr91] J. R. Chrzaszcz. Precisely speaking. IEEE Micro, 11(2): 2, March/April 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Christie:1996:DAK

[Chr96] Dave Christie. Developing the AMD-K5 architecture: Flying without instruments: the independent development on the x86 processor. IEEE Micro, 16(2):16–26, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Colin:2017:EAD

[CHSL17] Alexei Colin, Graham Harvey, Alanson P. Sample, and Brandon Lucia. An energy-aware debugger for intermittently powered systems. IEEE Micro, 37(3):116– 125, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030116-
abs.html.

Chua:2018:MRT

[Chu18] Leon Chua. Memristor: Remembrance of things past. IEEE Micro, 38(5):7–12, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050007.
html.

Corrigan:1985:EBM

[CJ85] Brian E. Corrigan and Everett L. Johnson. An evaluation of 8085-based multiprocessing on a timeshared bus. IEEE Micro, 5(3):11– 21, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Coggins:1995:LPN

[CJFP95] Richard Coggins, Marwan Jabri, Barry Flower, and Stephen Pickard. A low-power network for on-line diagnosis of heart patients — boosting performance of implantable devices that detect killer arrhythmias. IEEE Micro, 15 (3):18–25, May/June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Campanoni:2012:HME

[CJH^{+}12] Simone Campanoni, Timothy M. Jones, Glenn Holloway, Gu-Yeon Wei, and David Brooks. Helix: Making the extraction of thread-level parallelism mainstream. IEEE Micro, 32(4):8–18, July/August 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chamberlain:1995:OIM

[CK95] Roger D. Chamberlain and Robert R. Krchnavek. Optically interconnected multicomputers using invertedgraph topologies. IEEE Micro, 15(2):59–69, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cherif:1998:RMF

[CK98] Adel Cherif and Takuya Katayama. Replica management for fault-tolerant systems. IEEE Micro, 18(5):54– 65, September/October 1998.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5054.
pdf;   http://www.computer.
org/micro/mi1998/m5054abs.
htm.

Court:2011:LDM

[CK11] Craig A. Court and Paul H. J.

Kelly. Loop-directed mothballing: Power gating execution units using runtime loop analysis. IEEE Micro, 31(6): 29–38, November/December 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Conway:2010:CHM

[CKD^{+}10] Pat Conway, Nathan Kalyanasundharam, Gregg Donley, Kevin Lepak, and Bill Hughes. Cache hierarchy and memory subsystem of the AMD Opteron processor. IEEE Micro, 30(2):16–29, March/April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chen:2009:FHA

[CKG^{+}09] Shimin Chen, Michael Kozuch, Phillip B. Gibbons, Michael Ryan, Theodoros Strigkos, Todd C. Mowry, Olatunji Ruwase, Evangelos Vlachos, Babak Falsafi, and Vijaya Ramachandran. Flexible hardware acceleration for instruction-grain lifeguards. IEEE Micro, 29(1):62–72, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Corsini:1987:ACB

[CL87] Paolo Corsini and Lanfranco Lopriore. The architecture of a capability-based microprocessor system. IEEE Micro, 7(3):35–51, May/June 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cain:2004:MOV

[CL04] Harold W. Cain and Mikko H. Lipasti. Memory ordering: a value-based approach. IEEE Micro, 24(6):110–117, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/06/m6110.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/06/m6110.pdf.

Chang:2005:DZS

[CL05] Yen-Jen Chang and Feipei Lai. Dynamic zero-sensitivity scheme for low-power cache memories. IEEE Micro, 25 (4):20–32, July/August 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Claasen:2003:SCC

[Cla03] Theo A. C. M. Claasen. System on a chip: Changing IC design today and in the future. IEEE Micro, 23(3): 20–26, May/June 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/03/m3020abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/03/m3020.
pdf.

Clements:2000:GEI

[Cle00a] Alan Clements. Guest Editor’s introduction: Computer architecture education. IEEE Micro, 20(3):10–12, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3010.
pdf.

Clements:2000:UCC

[Cle00b] Alan Clements. The undergraduate curriculum in computer architecture. IEEE Micro, 20(3):13–22, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3013.
pdf;   http://www.computer.
org/micro/mi2000/m3013abs.
htm.

Clements:2003:CCS

[Cle03]  Alan Clements.  CSIDC:

Competing students design real-world systems. IEEE Micro, 23(4):78–80, July/ August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2003/04/m4078.pdf.

Cai:2020:AAE

[CLL^{+}20] H. Cai, J. Lin, Y. Lin, Z. Liu, K. Wang, T. Wang, L. Zhu, and S. Han. AutoML for architecting efficient and specialized neural networks. IEEE Micro, 40(1):75– 82, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cho:2008:GEI

[CLM08] Sangyeun Cho, Tao Li, and Onur Mutlu. Guest Editors’ introduction: Interaction of many-core computer architecture and operating systems. IEEE Micro, 28(3): 2–5, May/June 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2008/03/mmi2008030002.
                                        pdf.

Culler:1996:AFN

[CLMY96] David E. Culler, Lok Tin Liu, Richard P. Martin, and Chad O. Yoshikawa. Assessing fast network interfaces. IEEE Micro, 16(1):35– 43, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cohen:1986:DIM

[CM86] Brad Cohen and Ralph Mc-Garity. The design and implementation of the MC68851 paged memory management unit. IEEE Micro, 6(2):13– 28, March/April 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

C:2004:TAI

[CM04] Ravikumar V. C. and Rabi N.

Mahapatra. TCAM architecture for IP lookup using prefix properties. IEEE Micro, 24 (2):60–69, March/April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/02/m2060abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/02/m2060.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/
m2060.pdf.

Chin:2017:HC

[CM17] Bryan Chin and Subhasish Mitra. Hot chips 28. IEEE Micro, 37(2):5– 6, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020005.
html.

Coskun:2011:ASC

[CMAS11] Ayse K. Coskun, Jie Meng, David Atienza, and Mohamed M. Sabry. Attaining single-chip, high-performance computing through 3D systems with active cooling. IEEE Micro, 31(4):63–75, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Camarero:2022:PRL

[CMB22] Cristóbal Camarero, Carmen Martínez, and Ramón Beivide. Polarized routing for large interconnection networks. IEEE Micro, 42(2):61– 67, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chun:1998:VNT

[CMC98] B. N. Chun, A. M. Mainwaring, and D. E. Culler. Virtual network transport protocols for Myrinet. IEEE Micro, 18 (1):53–63, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chowdhury:2023:GFA

[CML^{+}23] Animesh Basak Chowdhury, Anushree Mahapatra, Yang Liu, Prashanth Krishnamurthy, Farshad Khorrami, and Ramesh Karri. A goldenfree approach to detect Trojans in COTS multi-PCB systems. IEEE Micro, 43(5):64– 76, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Catania:1997:AFL

[CMR97] Vincenzo Catania, Michele Malgeri, and Marco Russo. Applying fuzzy logic to codesign partitioning — using soft computing techniques to automate partitioning for hardware-software codesign. IEEE Micro, 17 (3):62–70, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3062.
pdf.

Cooklev:2013:ORD

[CN13] Todor Cooklev and Akinori Nishihara. An open RFdigital interface for softwaredefined radios. IEEE Micro, 33(6):47–55, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Chrysos:2016:UQC

[CNC^{+}16] Nikolaos Chrysos, Fredy Neeser, Rolf Clauberg, Daniel Crisan, Kenneth M. Valk, Claude Basso, Cyriel Minkenberg, and Mitch Gusat. Unbiased quantized congestion notification for scalable server fabrics. IEEE Micro, 36(6):50–58, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/06/mmi2016060050-
                                        abs.html.

Chen:2003:JSD

[CO03] Michael K. Chen and Kunle Olukotun. The Jrpm system for dynamically parallelizing sequential Java programs. IEEE Micro, 23(6):26–35, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/06/m6026abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/06/m6026.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2003/06/
                                        m6026.pdf.

Colwell:1989:RTC

[Col89] R. P. Colwell. RISC tutorial comments. IEEE Micro, 9(3):4, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Colwell:2021:OIM

[Col21] Robert P. Colwell. The origin of Intel’s micro-ops. IEEE Micro, 41(6):37–41, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Constantinescu:2003:TCV

[Con03] Cristian Constantinescu. Trends and challenges in VLSI circuit reliability. IEEE Micro, 23 (4):14–19, July/August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/04/m4014abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/04/m4014.
pdf.

Corsini:1986:MID

[CP86] Paolo Corsini and Cosimo Antonio Prete. Multibug: Interactive debugging in distributed systems. IEEE Micro, 6(3):26–33, May/June 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chassaing:1990:TBM

[CPH90] Rulph Chassaing, Wayne A.

Peterson, and Darrell W. Horning. A TMS320C25based multirate filter. IEEE Micro, 10(5):54–62, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chang:2018:EAT

[CPS^{+}18] Norman Chang, Stephen Pan, Karthik Srinivasan, Zhigang Feng, Wenbo Xia, Tim Pawlak, and David Geb. Emerging ADAS thermal reliability needs and solutions.

IEEE Micro, 38(1):66–81, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010066-
abs.html.

Civera:1989:ISV

[CPZ89] Pierluigi Civera, Gianluca Piccinini, and Maurizio Zamboni. Implementation studies for a VLSI Prolog coprocessor. IEEE Micro, 9(1):10–23, January/February 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Castelli:1995:ERT

[CR95a] Gianluigi Castelli and Gianluca Ragazzini. EOS — a real-time operating system adapts to application architectures. IEEE Micro, 15(5):41– 49, September/October 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chiaberge:1995:CNF

[CR95b] Marcello Chiaberge and Leonardo M. Reyneri. Cintia: a neuro-fuzzy real-time controller for lowpower embedded systems — supporting a variety of intelligent control strategies. IEEE Micro, 15(3):40–47, May/ June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Crawford:1990:ICE

[Cra90]  John H. Crawford.  The

i486 CPU: executing instructions in one clock cycle. IEEE Micro, 10(1):27–36, January/February 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Crawford:2000:GEI

[Cra00] John H. Crawford. Guest Editor’s introduction: Introducing the Itanium processors. IEEE Micro, 20(5):9– 11, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5009.
pdf;   http://www.computer.
org/micro/mi2000/m5009abs.
htm.

Crenshaw:1982:MIS

[Cre82] J. W. Crenshaw. More on instruction sets. IEEE Micro, 2(4):6, October/December 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Crisp:1997:DRT

[Cri97]  Richard Crisp.  Direct

Rambus technology: The next main memory standard. IEEE Micro, 17(6): 18–28, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi1997/pdf/m6018.
pdf;   http://www.computer.
org/micro/mi1997/m6018abs.
htm.

Crowl:1985:RTF

[Cro85] Daniel A. Crowl. A real-time Fortran executive. IEEE Micro, 5(4):48–66, July/August 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cazorla:2004:QHP

[CRV^{+}04] Francisco J. Cazorla, Alex Ramirez, Mateo Valero, Peter M. W. Knijnenburg, Rizos Sakellariou, and Enrique Fernández. QoS for high-performance SMT processors in embedded systems. IEEE Micro, 24(4):24– 31, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4024abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/04/m4024.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/04/
m4024.pdf.

Chu:1981:MCD

[CS81] Yaohan Chu and Paul L. Schapiro. Micro-Cobol — a data-processing language for microprocessor systems. IEEE Micro, 1(4):43–55, October/December 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chowdhury:2008:HPF

[CS08] Shubhajit Roy Chowdhury and Hiranmay Saha. A highperformance FPGA-based fuzzy processor architecture for medical diagnosis. IEEE Micro, 28(5):38–52, September/October 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chang:2013:MSW

[CS13] Chia-Ming Chang and Olav Solgaard. Monolithic silicon waveguides in standard silicon. IEEE Micro, 33(1):32– 40, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Coole:2014:FFH

[CS14] James Coole and Greg Stitt.

Fast, flexible high-level synthesis from OpenCL using reconfiguration contexts. IEEE Micro, 34(1):42–53, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Ceze:2015:TPC

[CS15] Luis Ceze and Karin Strauss.

The 2014 top picks in computer architecture. IEEE Micro, 35(3):5–9, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/03/mmi2015030005.
html.

Cui:2018:AR

[CS18] Weilong Cui and Timothy Sherwood. Architectural risk. IEEE Micro, 38(3):116– 125, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030116-
abs.html.

Cristal:2005:KIP

[CSC^{+}05] Adrian Cristal, Oliverio J. Santana, Francisco Cazorla, Marco Galluzzi, Tanausu Ramirez, Miquel Pericas, and Mateo Valero. Kiloinstruction processors: Overcoming the memory wall. IEEE Micro, 25(3):48–57, May/June 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chandra:2022:DDD

[CSC^{+}22] Ranveer Chandra, Manohar Swaminathan, Tusher Chakraborty, Jian Ding, Zerina Kapetanovic, Peeyush Kumar, and Deepak Vasisht. Democratizing datadriven agriculture using affordable hardware. IEEE Micro, 42(1):69–77, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cantin:2006:CGC

[CSL^{+}06] Jason F. Cantin, James E. Smith, Mikko H. Lipasti, Andreas Moshovos, and Babak Falsafi. Coarse-grain coherence tracking: Region-Scout and region coherence arrays. IEEE Micro, 26(1):70– 79, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cheikh:2021:KDV

[CSM^{+}21] A. Cheikh, S. Sordillo, A. Mastrandrea, F. Menichelli, G. Scotti, and M. Olivieri. Klessydra-T: Designing vector coprocessors for multithreaded edge-computing cores. IEEE Micro, 41(2):64– 71, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Carloni:2002:CLS

[CSV02] Luca P. Carloni and Alberto L. Sangiovanni-Vincentelli. Coping with latency in SOC design. IEEE Micro, 22(5):24– 35, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5024.
pdf;   http://www.computer.
org/micro/mi2002/m5024abs.
htm.

Cairns:1995:PIL

[CT95] Graham Cairns and Lionel Tarassenko. Precision issues for learning with analog VLSI multilayer perceptrons. IEEE Micro, 15(3):

54–56, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cummings:2004:PCC

[Cum04] Uri Cummings. PivotPoint: Clockless crossbar switch for high-performance embedded systems. IEEE Micro, 24 (2):48–59, March/April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/02/m2048abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2004/02/m2048.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2004/02/
                                        m2048.pdf.

Chou:1994:SRR

[CWB94] Pai Chou, Elizabeth A. Walkup, and Gaetano Borriello. Scheduling for reactive real-time systems. IEEE Micro, 14(4):37–47, July/August 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chen:2014:FMC

[CWL^{+}14] Shuming Chen, Yaohua Wang, Sheng Liu, Jianghua Wan, Haiyan Chen, Hengzhu Liu, Kai Zhang, Xiangyuan Liu, and Xi Ning. FT-matrix: A coordination-aware architecture for signal processing. IEEE Micro, 34(6):64–73, November/December 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060064-
abs.html.

Chen:2015:EPO

[CWLS15] Guoyang Chen, Bo Wu, Dong Li, and Xipeng Shen. Enabling portable optimizations of data placement on GPU. IEEE Micro, 35(4):16–24, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040016-
abs.html.

Choudhary:2012:FAS

[CWS^{+}12] Niket K. Choudhary, Salil V.

Wadhavkar, Tanmay A. Shah, Hiran Mayukh, Jayneel Gandhi, Brandon H. Dwiel, Sandeep Navada, Hashem H. Najafabadi, and Eric Rotenberg. FabScalar: Automating superscalar core design. IEEE Micro, 32(3):48–59, May/ June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chin:2018:DSA

[CYH^{+}18] Ting-Wu Chin, Chia-Lin Yu, Matthew Halpern, Hasan Genc, Shiao-Li Tsao, and Vijay Janapa Reddi. Domainspecific approximation for object detection. IEEE Micro, 38(1):31–40, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/01/mmi2018010031-
                                        abs.html.

Diefendorff:1992:OMS

[DA92] Keith Diefendorff and Michael Allen. Organization of the Motorola 88110 superscalar RISC microprocessor. IEEE Micro, 12(2):40–63, March/ April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamant:2023:SIH

[DA23] Ron Diamant and Krste Asanovic. Special issue on Hot Chips 34. IEEE Micro, 43 (3):7–8, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Daivs:1994:WWC

[Dai94] G. G. Daivs. The war of the words continues — reply. IEEE Micro, 14(2):79, March/ April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dancea:1989:DCL

[Dan89] Ioan Dancea. Dynamically changing the logical behavior of a microcomputer interface. IEEE Micro, 9(2):39– 51, March/April 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Daniels:1996:PPF

[Dan96] R. Gary Daniels. A participant’s perspective — the “father” of the 6805 MCU relates his role in the history of the microprocessor. IEEE Micro, 16(6):21–31, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Das:2017:BLB

[Das17] Reetuparna Das. Blurring the lines between memory and computation. IEEE Micro, 37(6):13–15, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060013-
abs.html.

DasSharma:2021:PES

[Das21] D. Das Sharma. PCI Express 6.0 specification: a low-latency, high-bandwidth, high-reliability, and costeffective interconnect with 64.0 GT/s PAM-4 signaling. IEEE Micro, 41(1):23–29, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Das:2022:SIM

[Das22] Reetuparna Das. Special issue on in-memory computing. IEEE Micro, 42(1):87–88, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Davis:1993:WWI

[Dav93] G. Gervaise Davis III. War of the words: intellectual property laws and standardization. IEEE Micro, 13(6): 19–27, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Davidson:1998:LCV

[Dav98] Evan E. Davidson. Large chip vs. MCM for a highperformance system. IEEE Micro, 18(4):33–41, July/ August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m4033.pdf;   http://
www.computer.org/micro/
mi1998/m4033abs.htm.

Davies:2002:DMI

[Dav02]  David H. Davies.  Data-

Play’s mobile information distribution and storage technology. IEEE Micro, 22 (2):8–15, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2008.
pdf;   http://www.computer.
org/micro/mi2002/m2008abs.
htm.

Dubnicki:1998:SPU

[DBC^{+}98] C. Dubnicki, A. Bilas, Y. Q.

Chen, S. N. Damianakis, and K. Li. Shrimp Project update — Myrinet communication. IEEE Micro, 18(1):50– 52, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Damianakis:1997:CSC

[DBDF97] Stefanos N. Damianakis, Angelos Bilas, Cezary Dubnicki, and Edward W. Felten. Client-server computing on Shrimp — delivering almost undiminished hardware performance to user applications. IEEE Micro, 17(1):8– 18, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

deCarvalho:2022:CIM

[dCMA22] João P. L. de Carvalho, José E. Moreira, and José Nelson Amaral. Compiling for the IBM Matrix Engine for enterprise workloads. IEEE Micro, 42(5):34–40, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dangwal:2020:TWP

[DCMS20] D. Dangwal, W. Cui, J. McMahan, and T. Sherwood. Trace wringing for program trace privacy. IEEE Micro, 40(3):

108–115, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dally:2005:HCP

[DD05] Bill Dally and Keith Diefendorff. Hot Chips 16: Power, parallelism, and memory performance. IEEE Micro, 25(2): 8–9, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2008.pdf;
http://csdl.computer.org/
comp/mags/mi/2005/02/m2008abs.
htm.

Datta:2019:BEL

[DDG^{+}19] S. Datta, S. Dutta, B. Grisafe, J. Smith, S. Srinivasa, and H. Ye. Back-end-of-line compatible transistors for monolithic 3-D integration. IEEE Micro, 39(6):8–15, November/December 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diefendorff:2000:AEP

[DDHS00] Keith Diefendorff, Pradeep K. Dubey, Ron Hochsprung, and Hunter Scales. AltiVec extension to PowerPC accelerates media processing. IEEE Micro, 20(2):85–95, March/ April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi2000/pdf/m2085.
pdf.

DePrycker:1983:PCT

[De 83] Martin De Prycker. A performance comparison of three contemporary 16-bit microprocessors. IEEE Micro, 3 (2):26–37, March/April 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

deSollaPrice:1984:HCM

[de 84] Derek de Solla Price. A history of calculating machines. IEEE Micro, 4(1):22–52, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

DeMicheli:1994:CAH

[De 94] Giovanni De Micheli. Computeraided hardware-software codesign. IEEE Micro, 14(4):10– 16, July/August 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dean:2004:ERT

[Dea04] Alexander G. Dean. Efficient real-time fine-grained concurrency on low-cost microcontrollers. IEEE Micro, 24 (4):10–22, July/August 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4010abs.htm;

                                         http://csdl.computer.org/
                                        dl/mags/mi/2004/04/m4010.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2004/04/
                                        m4010.pdf.

Delcorso:1991:BA

[Del91a] D. Delcorso. Bioengineering applications. IEEE Micro, 11 (5):2–3, September/October 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1991:LD

[Del91b] D. Delcorso. Launching the 2nd decade. IEEE Micro, 11(1):2–4, January/February 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1991:US

[Del91c] D. Delcorso. The ultimate solution. IEEE Micro, 11(2): 6–7, March/April 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1992:LST

[Del92] D. Delcorso. Letters and special themes. IEEE Micro, 12(6):2, November/December 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1993:BAA

[Del93a] D. Delcorso. Best article award for 1992. IEEE Micro, 13(4):2, July/August 1993.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1993:CP

[Del93b] D. Delcorso. Check point 2. IEEE Micro, 13(1):2–3, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1994:U

[Del94a]  D. Delcorso.  Untitled.

IEEE Micro, 14(4):2–3, July/ August 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1994:WBC

[Del94b] D. Delcorso. White beards and clock ticks. IEEE Micro, 14(6):2–3, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Demicheli:1994:HSC

[Dem94]  G. Demicheli.  Hardware-

software codesign — introduction. IEEE Micro, 14(4):8– 9, July/August 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dennis:1983:HRA

[Den83] Philip C. Dennis. A heuristic routing algorithm. IEEE Micro, 3(2):48–54, March/April 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dhem:2001:HSS

[DF01] Jean-François Dhem and Nathalie Feyt. Hardware and software symbiosis helps smart card evolution. IEEE Micro, 21(6):14–25, November/December 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
m6014abs.htm;   http://
dlib.computer.org/mi/books/
mi2001/pdf/m6014.pdf.

Dreslinski:2013:CCS

[DFG^{+}13] Ronald G. Dreslinski, David Fick, Bharan Giridhar, Gyouho Kim, Sangwon Seo, Matthew Fojtik, Sudhir Satpathy, Yoonmyung Lee, Daeyeon Kim, Nurrachman Liu, Michael Wieckowski, Gregory Chen, Dennis Sylvester, David Blaauw, and Trevor Mudge. Centip3De: A 64-core, 3D stacked near-threshold system. IEEE Micro, 33(2):8–16, March/ April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Davis:1990:MDC

[DFR90] Henry Davis, Robert Fine, and Denis Regimbal. Merging data converters and DSPs for mixed-signal processors. IEEE Micro, 10(5):17–27, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1987:EAA

[DG87] D. Delcorso and K. E.

Grosspietsch. European approaches for advanced architectures. IEEE Micro, 7(5):4– 5, September/October 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1988:EAA

[DG88] D. Delcorso and K. E.

Grosspietsch. European approaches for advanced architectures. IEEE Micro, 8(5):8– 9, September/October 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1989:GEI

[DG89] D. Delcorso and K. E.

Grosspietsch. Guest Editors’ introduction — European approaches for advanced architectures. 2. IEEE Micro, 9 (1):9, January/February 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

deSalvador:1995:MSA

[dG95] Luis de Salvador and Julio Gutierrez. A multilevel systolic approach for fuzzy inference hardware. IEEE Micro, 15(5):61–71, September/October 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dror:2011:OCL

[DGM^{+}11] Ron O. Dror, J. P. Grossman, Kenneth M. Mackenzie, Brian Towles, Edmond Chow, John K. Salmon, Cliff Young, Joseph A. Bank, Brannon Batson, Martin M. Deneroff, Jeffrey S. Kuskin, Richard H. Larson, Mark A. Moraes, and David E. Shaw. Overcoming communication latency barriers in massively parallel scientific computation. IEEE Micro, 31(3):8–19, May/June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dahlen:2000:SWC

[DGMM00] Eric Dahlen, Jennifer Gustin, Susan Meredith, and Doug Moran. The 82460GX server/workstation chip set. IEEE Micro, 20(6):69–75, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6069.
pdf;   http://www.computer.
org/micro/mi2000/m6069abs.
htm.

Desmet:2010:AAD

[DGR^{+}10] Veerle Desmet, Sylvain Girbal, Alex Ramirez, Augusto Vega, and Olivier Temam. ArchExplorer for automatic design space exploration. IEEE Micro, 30(5):5– 15, September/October 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Delcorso:1989:EAV

[DGT89] D. Delcorso, K. E. Grosspietsch, and P. Treleaven. European approaches to VLSI neural networks. IEEE Micro, 9 (6):5–7, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Davis:1994:RFT

[DGW^{+}94] Nathaniel J. Davis IV, F. Gail Gray, Joseph A. Wegner, Shannon E. Lawson, Vinay Murthy, and Tennis S. White. Reconfiguring fault-tolerant two-dimensional array architectures. IEEE Micro, 14 (2):60–69, March/April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dyer:1990:GEI

[DH90] Stephen A. Dyer and Richard J.

Higgins. Guest Editor’s introduction: The maturing of DSP. IEEE Micro, 10(5):11– 13, September/October 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1992:QS

[Dia92] S. L. Diamond. Quality standards. IEEE Micro, 12(6):3– 5, November/December 1992.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1993:DTD

[Dia93a] S. L. Diamond. Desk-top document management. IEEE Micro, 13(3):4–6, May/June 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1993:FFF

[Dia93b]  S. L. Diamond.  Fair is

foul, and foul is fair. IEEE Micro, 13(1):71–72, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1993:MSB

[Dia93c] S. L. Diamond. Micro standards: Building standards. IEEE Micro, 13(4):59–60, July/August 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1993:MSO

[Dia93d] S. L. Diamond. Micro standards: Organizing the corporate standards function. IEEE Micro, 13(3):81–83, May/ June 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1993:SI

[Dia93e] S. L. Diamond. Standards and innovation. IEEE Micro, 13 (2):77–78, March/April 1993.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1993:GEI

[Dia93f] Stephen L. Diamond. Guest Editor’s introduction: Reengineering standards. IEEE Micro, 13(6):8–9, November/December 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1994:AND

[Dia94a] S. L. Diamond. Architecture Neutral Distribution Format (ANDF). IEEE Micro, 14(6): 73–76, November/December 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1994:NPP

[Dia94b] S. L. Diamond. A new PC parallel interface standard (IEEE Std 1284-1994). IEEE Micro, 14(4):3, 78, July/August 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1995:LC

[Dia95a] S. L. Diamond. Life-cycles.

IEEE Micro, 15(1):4–5, January/February 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1995:RWT

[Dia95b] S. L. Diamond. Riding the waves of technology. IEEE Micro, 15(3):5–6, May/June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1995:ECI

[Dia95c] Stephen L. Diamond. From the Editor-in-Chief — introducing Micro Web and μRAP. IEEE Micro, 15(6):2, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1995:MSI

[Dia95d] Stephen L. Diamond. Micro standards: IEEE Std P1394: This convenient, easyto-use serial bus standard supports low-cost desktop and portable-computing applications. IEEE Micro, 15(2):81– 83, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1995:MSP

[Dia95e] Stephen L. Diamond. Micro standards: Preparing for global participation. IEEE Micro, 15(1):76–77, January/February 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1996:MNF

[Dia96a] S. Diamond. Micro news: Fletcher Buckley, 30-year CS volunteer, dies. IEEE Micro, 16(5):2–??, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1996:CM

[Dia96b] S. L. Diamond. Celebrating the microprocessor. IEEE Micro, 16(6):2–3, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1996:MSS

[Dia96c] S. L. Diamond. Micro standards: SyncLink: High-speed DRAM for the future. IEEE Micro, 16(6):74–75, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1996:MSI

[Dia96d] Stephen L. Diamond. Micro standards: IEEE Std 1394-1955 High Performance Serial Bus is ready for second-generation standardization. IEEE Micro, 16(3): 75–78, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diamond:1998:ECM

[Dia98] S. L. Diamond. Editor-in-Chief’s message: Life recycles. IEEE Micro, 18(6): 3, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6003.
pdf.

Diamond:1999:MVP

[Dia99] Stephen L. Diamond. Micro view: Putting children first: the Lego MindStorms product. IEEE Micro, 19(6): 88, 87, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6088.
pdf.

Diamond:2000:MVM

[Dia00] Stephen L. Diamond. Micro view: Microdisplay applications reach the mainstream. IEEE Micro, 20 (1):80, 79, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1080.

pdf. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

DiGirolamo:2016:EOE

[DJUH16] Salvatore Di Girolamo, Pierre Jolivet, Keith D. Underwood, and Torsten Hoefler. Exploiting offload-enabled network interfaces. IEEE Micro, 36(4): 6–17, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040006-
abs.html.

Delimitrou:2014:QSA

[DK14] Christina Delimitrou and Christos Kozyrakis. Qualityof-service-aware scheduling in heterogeneous data centers with paragon. IEEE Micro, 34(3):17–30, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Delimitrou:2018:USI

[DK18] Christina Delimitrou and Christos Kozyrakis. Uncovering the security implications of cloud multi-tenancy with Bolt. IEEE Micro, 38(3): 86–97, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030086-
abs.html.

Darley:1990:TFP

[DKB^{+}90] Merrick Darley, Bill Kronlage, David Bural, Bob Churchill, David Pulling, Paul Wang, Rick Iwamoto, and Larry Yang. The TMS390C602A floating-point coprocessor for Sparc systems. IEEE Micro, 10(3):36–47, May/June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dally:2021:EGP

[DKK21] William J. Dally, Stephen W. Keckler, and David B. Kirk. Evolution of the graphics processing unit (GPU). IEEE Micro, 41(6):42–51, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Duardo:1992:AII

[DKM^{+}92] Obed Duardo, Scott C. Knauer, John N. Mailhot, Kalyan Mondal, and Tommy C. Poon. Architecture and implementation of ICs for a DSC-HDTV video decoder system. IEEE Micro, 12(5): 22–27, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dharmapurikar:2004:DPI

[DKSL04] Sarang Dharmapurikar, Praveen Krishnamurthy, Todd S. Sproull, and John W. Lockwood. Deep packet inspection using parallel bloom filters. IEEE Micro, 24(1):52– 61, January/February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1052abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/01/m1052.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/01/
m1052.pdf.

Doweck:2017:IGI

[DKyL^{+}17] Jack Doweck, Wen-Fu Kao, Allen Kuan yu Lu, Julius Mandelblat, Anirudha Rahatekar, Lihu Rappoport, Efraim Rotem, Ahmad Yasin, and Adi Yoaz. Inside 6thgeneration Intel core: New microarchitecture code-named Skylake. IEEE Micro, 37 (2):52–62, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020052-
abs.html.

Devietti:2010:DDS

[DLCO10] Joseph Devietti, Brandon Lucia, Luis Ceze, and Mark Oskin. DMP: Deterministic shared-memory multiprocessing. IEEE Micro, 30(1):40–49, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Danese:2002:PNP

[DLR02] Giovanni Danese, Francesco Leporati, and Stefano Ramat. A parallel neural processor for real-time applications. IEEE Micro, 22(3):20–31, May/ June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m3020.
pdf;   http://www.computer.
org/micro/mi2002/m3020abs.
htm.

Dirvin:1986:MTB

[DM86] Rhonda Alexis Dirvin and Arthur R. Miller. The MC68824 token bus controller: VLSI for the factory LAN. IEEE Micro, 6(3):15– 25, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dyer:1988:FPD

[DM88a] S. A. Dyer and L. R. Morris. Floating-point digital signalprocessing chips — a new era for DSP systems-design. IEEE Micro, 8(6):10–12, November/December 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dyer:1988:AFP

[DM88b] Stephen A. Dyer and L. Robert Morris. Afterword: Floating-Point digital signal processing chips, the end of the supercomputer era? IEEE Micro, 8(6):86–??, November/December 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Djordjevic:2000:IET

[DMG00] Jovan Djordjevic, Aleksandar Milenkovic, and Nenad Grbanovic. An integrated environment for teaching computer architecture. IEEE Micro, 20(3):66–74, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3066.
pdf;   http://www.computer.
org/micro/mi2000/m3066abs.
htm.

Dong:2015:VSB

[DMG^{+}15] YaoZu Dong, JunJie Mao, HaiBing Guan, Jian Li, and Yu Chen. A virtualization solution for BYOD with dynamic platform context switching. IEEE Micro, 35(1):34–43, January/ February 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010034-
abs.html.

Das:2011:ANC

[DMMD11] Reetuparna Das, Onur Mutlu, Thomas Moscibroda, and Chita R. Das. Aérgia: a network-on-chip exploiting packet latency slack. IEEE Micro, 31(1):29–41, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

DiStefano:1991:IDB

[DMP91] Antonella Di Stefano, Orazio Mirabella, and Fabio Presente. Implementing a DSPbased Petri-net simulation tool. IEEE Micro, 11(2):20– 23, 56–64, March/April 1991.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Demme:2013:QEA

[DMWS13] John Demme, Robert Martin, Adam Waksman, and Simha Sethumadhavan. A quantitative, experimental approach to measuring processor sidechannel security. IEEE Micro, 33(3):68–77, May/June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Doster:1984:WPL

[DO84] Wolfgang Doster and Richard Oed. Word processing with on-line script recognition. IEEE Micro, 4(5):36– 43, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Doggett:2012:PTC

[Dog12] Michael Doggett. Prolegomena: Texture caches. IEEE Micro, 32(3):136–141, May/ June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diefendorff:1994:EPA

[DOH94] Keith Diefendorff, Rich Oehler, and Ron Hochsprung. Evolution of the PowerPC architecture. IEEE Micro, 14(2):34– 49, March/April 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dorcey:1986:WNN

[Dor86] C. T. Dorcey. When is a von Neumann not a von Neumann? IEEE Micro, 6(1):6, January/February 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dally:1997:TEG

[DP97] William J. Dally and John Poulton. Transmitter equalization for 4-Gbps signaling — applying the density and speed of modern VLSI technology to overcome the I/O bottleneck with high-speed signaling. IEEE Micro, 17 (1):48–56, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Donato:2019:MOC

[DPBW19] M. Donato, L. Pentecost, D. Brooks, and G. Wei. MEMTI: Optimizing on-chip nonvolatile storage for visual multitask inference at the edge. IEEE Micro, 39(6): 73–81, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Deng:2021:LPH

[DPT^{+}21] Zhaoxia Deng, Jongsoo Park, Ping Tak Peter Tang, Haixin Liu, Jie Yang, Hector Yuen, Jianyu Huang, Daya Khudia, Xiaohan Wei, Ellie Wen, Dhruv Choudhary, Raghuraman Krishnamoorthi, Carole-Jean Wu, Satish Nadathur, Changkyu Kim, Maxim Naumov, Sam Naghshineh, and Mikhail Smelyanskiy. Lowprecision hardware architectures meet recommendation model inference at scale. IEEE Micro, 41(5):93–100, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dean:2018:NGA

[DPY18] Jeff Dean, David Patterson, and Cliff Young. A new golden age in computer architecture: Empowering the machine-learning revolution. IEEE Micro, 38(2):21– 29, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/02/mmi2018020021-
abs.html.

Dai:2024:HSD

[DQCL24] Liuyao Dai, Hao Qi, Weicong Chen, and Xiaoyi Lu. High-speed data communication with advanced networks in large language model training. IEEE Micro, 44(2):31– 40, March/April 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Draber:2000:OFT

[Dra00] Silke Draber. Optimizing fault tolerance in embedded distributed systems. IEEE Micro, 20(4):76–87, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4076.
pdf;   http://www.computer.
org/micro/mi2000/m4076abs.
htm.

Deng:2012:ALP

[DRB^{+}12] Qingyuan Deng, Luiz Ramos, Ricardo Bianchini, David Meisner, and Thomas F. Wenisch. Active low-power modes for main memory with MemScale. IEEE Micro, 32 (3):60–69, May/June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dunning:1998:VIA

[DRM^{+}98] Dave Dunning, Greg Regnier, Gary McAlpine, Don Cameron, Bill Shubert, Frank Berry, Anne Marie Merritt, Ed Gronke, and Chris Dodd. The virtual interface architecture. IEEE Micro, 18 (2):66–76, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2066.
pdf;   http://www.computer.
org/micro/mi1998/m2066abs.
htm.

Dastidar:2023:AGA

[DRM^{+}23] Jaideep Dastidar, David Riddoch, Jason Moore, Steven Pope, and Jim Wesselkamper. The AMD 400-G adaptive SmartNIC system on chip: a technology preview. IEEE Micro, 43(3):40–49, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Diefendorff:1994:PUI

[DS94] K. Diefendorff and E. Silha. The PowerPC user instruction set architecture. IEEE Micro, 14(5):30–41, September/October 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dolle:1995:CER

[DS95] Michael Dolle and Manfred Schlett. A cost-effective RISC/DSP microprocessor for embedded systems. IEEE Micro, 15(5):32–40, September/October 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

DelaParra:2022:ITM

[DSG^{+}22] Cecilia De la Parra, Taha Soliman, Andre Guntoro, Akash Kumar, and Norbert Wehn. Increasing throughput of in-memory DNN accelerators by flexible layerwise DNN approximation. IEEE Micro, 42(6):17–24, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dally:1992:MDP

[DSK^{+}92] William J. Dally, J. A. Stuart Fiske, John S. Keen, Richard A. Lethin, Michael D. Noakes, Peter R. Nuth, Roy E. Davison, and Gregory A. Fyler. The message-driven processor — a multicomputer processing node with efficient mechanisms. IEEE Micro, 12 (2):23–39, March/April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips III, Stanford University, 1992.

Davies:2018:LNM

[DSL^{+}18] Mike Davies, Narayan Srinivasa, Tsung-Han Lin, Gautham Chinya, Yongqiang Cao, Sri Harsha Choday, Georgios Dimou, Prasad Joshi, Nabil Imam, Shweta Jain, Yuyun Liao, Chit-Kwan Lin, Andrew Lines, Ruokun Liu, Deepak Mathaikutty, Steven McCoy, Arnab Paul, Jonathan Tse, Guruguhanathan Venkataramanan, Yi-Hsin Weng, Andreas Wild, Yoonseok Yang, and Hong Wang. Loihi: A neuromorphic manycore processor with on-chip learning. IEEE Micro, 38(1):82–99, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010082-
abs.html.

Dally:2001:GEI

[DTB01] William J. Dally, Marc Tremblay, and Allen J. Baum. Guest Editors’ introduction: Hot Chips 12. IEEE Micro, 21(2):13–15, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m2013.
                                        pdf.

Ditzel:2022:AMR

[DtEt22] David R. Ditzel and the Esperanto team. Accelerating ML recommendation with over 1,000 RISC-V/Tensor processors on Esperanto’s ET-SoC-1 chip. IEEE Micro, 42(3):31–38, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Drabik:1995:SFL

[DTH^{+}95] Timothy J. Drabik, Albert H. Titus, Mark A. Handschy, David Banas, Stephen D. Gaalema, and David J. Ward. 2D silicon/ferroelectric liquid crystal spatial light modulators. IEEE Micro, 15(4):67– 76, July/August 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dangwal:2020:AHD

[DTS20] D. Dangwal, G. Tzimpragos, and T. Sherwood. Agile hardware development and instrumentation with PyRTL. IEEE Micro, 40(4):76–84, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Duncan:1981:LIN

[Dun81] Fraser George Duncan. Levelindependent notation for microcomputer programs. IEEE Micro, 1(2):47–52, April/June 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Duncan:1982:DCA

[Dun82] F. G. Duncan. Debate continues on algebraic notation. IEEE Micro, 2(1):5, January/March 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Duranton:1996:IPN

[Dur96] Marc Duranton. Image processing by neural networks: Tackling real-world image analysis problems with a 12-DSP vector processor. IEEE Micro, 16(5):12–19, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dekker:1987:AAM

[Dv87] G. J. Dekker and A. J. van de Goor. AMORE — address mapping with overlapped rotating entries. IEEE Micro, 7(3):22–34, May/June 1987.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dhem:1996:SSC

[DVQ96] Jean-François Dhem, Daniel Veithen, and Jean-Jacques Quisquater. SCALPS: Smart Card for Limited Payment Systems: Merging a processor and a coprocessor on a fast, secure, low-cost chip dedicated to public-key cryptography. IEEE Micro, 16(3): 42–51, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dunigan:2005:PEC

[DVWW05] Thomas H. Dunigan, Jr., Jeffrey S. Vetter, James B. White III, and Patrick H. Worley. Performance evaluation of the Cray X1 distributed shared-memory architecture. IEEE Micro, 25(1):30–40, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1030.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1030.pdf.

Dwarkadas:2018:MWA

[Dwa18] Sandhya Dwarkadas. 2018 Maurice Wilkes Award given to Gabriel Loh. IEEE Micro, 38(6):83–84, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08585397-abs.
html.

Dwarkadas:2019:TPC

[Dwa19] S. Dwarkadas. Top picks in computer architecture from conferences in 2018. IEEE Micro, 39(3):6–10, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dong:2021:AAH

[DWF^{+}21] Jianbo Dong, Shaochuang Wang, Fei Feng, Zheng Cao, Heng Pan, Lingbo Tang, Pengcheng Li, Hao Li, Qianyuan Ran, Yiqun Guo, Shanyuan Gao, Xin Long, Jie Zhang, Yong Li, Zhisheng Xia, Liuyihan Song, Yingya Zhang, Pan Pan, Guohui Wang, and Xiaowei Jiang. ACCL: Architecting highly scalable distributed training systems with highly efficient collective communication library. IEEE Micro, 41(5):85– 92, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dadu:2020:TGP

[DWLN20] V. Dadu, J. Weng, S. Liu, and T. Nowatzki. Towards general-purpose acceleration: Finding structure in irregularity. IEEE Micro, 40(3): 37–46, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Davidson:2018:COS

[DXT^{+}18] Scott Davidson, Shaolin Xie, Christopher Torng, Khalid Al-Hawai, Austin Rovinski, Tutu Ajayi, Luis Vega, Chun Zhao, Ritchie Zhao, Steve Dai, Aporva Amarnath, Bandhav Veluri, Paul Gao, Anuj Rao, Gai Liu, Rajesh K. Gupta, Zhiru Zhang, Ronald Dreslinski, Christopher Batten, and Michael Bedford Taylor. The Celerity open-source 511-core RISC-V tiered accelerator fabric: Fast architectures and design methodologies for fast chips. IEEE Micro, 38(2):30– 41, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/02/mmi2018020030-
                                        abs.html.

El-Ayat:1985:IAI

[EAA85] Khaled A. El-Ayat and Rakesh K. Agarwal. The Intel 80386 — architecture and implementation. IEEE Micro, 5(6):4–22, November/December 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Evers:2022:ANG

[EBC22] Mark Evers, Leslie Barnes, and Mike Clark. The AMD next-generation “Zen 3” core. IEEE Micro, 42(3): 7–12, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eberle:2003:RNM

[Ebe03] Hans Eberle. A radio network for monitoring and diagnosing computer systems. IEEE Micro, 23(1):60– 65, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1060.
pdf;   http://www.computer.
org/micro/mi2003/m1060abs.
htm.

Esmaeilzadeh:2012:DSE

[EBS^{+}12] Hadi Esmaeilzadeh, Emily Blem, Renee St. Amant, Karthikeyan Sankaralingam, and Doug Burger. Dark silicon and the end of multicore scaling. IEEE Micro, 32(3): 122–134, May/June 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eckert:1982:MI

[Eck82] K. Eckert. A multiprocessor interface. IEEE Micro, 2 (4):67–70, October/December 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ejjeh:2022:HHA

[ECK^{+}22] Adel Ejjeh, Aaron Councilman, Akash Kothari, Maria Kotsifakou, Leon Medvinsky, Abdul Rafae Noor, Hashim Sharif, Yifan Zhao, Sarita Adve, Sasa Misailovic, and Vikram Adve. HPVM: Hardware-agnostic programming for heterogeneous parallel systems. IEEE Micro, 42(5):108–117, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Esmaeilzadeh:2012:WHP

[ECY^{+}12] Hadi Esmaeilzadeh, Ting Cao, Xi Yang, Stephen M. Blackburn, and Kathryn S. McKinley. What is happening to power, performance, and software? IEEE Micro, 32(3): 110–121, May/June 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ebrahimi:2018:GMD

[ED18] Masoumeh Ebrahimi and Masoud Daneshtalab. A general methodology on designing acyclic channel dependency graphs in interconnection networks. IEEE Micro, 38(3): 79–85, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030079-
abs.html.

Eddington:2002:IIC

[Edd02] Chris Eddington. InfiniBridge: An InfiniBand channel adapter with integrated switch. IEEE Micro, 22 (2):48–56, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2048.
pdf;   http://www.computer.
org/micro/mi2002/m2048abs.
htm.

Ernst:2004:RCL

[EDL^{+}04] Dan Ernst, Shidhartha Das, Seokwoo Lee, David Blaauw, Todd Austin, Trevor Mudge, Nam Sung Kim, and Krisztián Flautner. Razor: Circuit-level correction of timing errors for low-power operation. IEEE Micro, 24(6):10–20, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6010.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6010.pdf.

Edwards:1983:FCP

[Edw83] D. Edwards. Fighting the chip pirates. IEEE Micro, 3(6):5– 6, November/December 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Edwards:1999:CJ

[Edw99]  W. K. Edwards.  Core

Jini. IEEE Micro, 19(5):10– 11, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eyerman:2008:SLP

[EE08] Stijn Eyerman and Lieven Eeckhout. System-level performance metrics for multiprogram workloads. IEEE Micro, 28(3):42–53, May/June 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eyerman:2010:PTC

[EE10] Stijn Eyerman and Lieven Eeckhout. Per-thread cycle accounting. IEEE Micro, 30 (1):71–80, January/February 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eeckhout:2015:BYT

[Eec15a] Lieven Eeckhout. Building on 35 years toward a vibrant future. IEEE Micro, 35(1):2–3, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/01/mmi2015010002-
                                        abs.html.

Eeckhout:2015:HRP

[Eec15b] Lieven Eeckhout. Heterogeneity in response to the power wall. IEEE Micro, 35(4):2– 3, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040002-
abs.html.

Eeckhout:2015:HCI

[Eec15c]  Lieven Eeckhout.  Hot

chips in an increasingly diverse microprocessor landscape. IEEE Micro, 35(2): 2–3, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020002-
abs.html.

Eeckhout:2015:PEI

[Eec15d]  Lieven Eeckhout.  Perfor-

mance evaluation and its impact on design. IEEE Micro, 35(6):2–3, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060002-
abs.html.

Eeckhout:2015:SCAa

[Eec15e]  Lieven Eeckhout.  The

state of the computer architecture field and its top picks. IEEE Micro, 35(3): 2–4, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030002-
abs.html.

Eeckhout:2015:SCAb

[Eec15f] Lieven Eeckhout. The structure of computer architecture (R)evolution. IEEE Micro, 35(5):2–3, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        mi/2015/05/mmi2015050002-
                                        abs.html.

Eeckhout:2016:HCA

[Eec16a]  Lieven Eeckhout.  Hot

chips: The annual feast of riches. IEEE Micro, 36(2): 4, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2016/02/mmi2016020004-
                                        abs.html.

Eeckhout:2016:HID

[Eec16b] Lieven Eeckhout. Hot interconnects and debates on computer architecture research directions. IEEE Micro, 36(4): 2, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/04/mmi2016040002.
                                        html.

Eeckhout:2016:LFT

[Eec16c] Lieven Eeckhout. Looking forward to the 2016 theme issues. IEEE Micro, 36(1):2–3, January/February 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/01/mmi2016010002-
abs.html.

Eeckhout:2016:SOR

[Eec16d] Lieven Eeckhout. Security and our reader survey. IEEE Micro, 36(5):4–5, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050004-
abs.html.

Eeckhout:2016:TPW

[Eec16e] Lieven Eeckhout. Top picks and welcoming new Editorial Board members. IEEE Micro, 36(3):2–4, May/June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030002-
abs.html.

Eeckhout:2017:CCH

[Eec17a]  Lieven Eeckhout.  From

cool chips to hot interconnects. IEEE Micro, 37(5):4– 5, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050004.
html.

Eeckhout:2017:HCI

[Eec17b]  Lieven Eeckhout.  Hot

chips: Industry and academia cutting-edge microprocessors. IEEE Micro, 37(2):4, March/ April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/mi/2017/02/mmi2017020004.
                                        html.

Eeckhout:2017:MLS

[Eec17c] Lieven Eeckhout. Is Moore’s Law slowing down? What’s next? IEEE Micro, 37(4): 4–5, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi2017040004.
                                        html.

Eeckhout:2017:LFU

[Eec17d] Lieven Eeckhout. Looking forward to upcoming themes. IEEE Micro, 37(1):4–5, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/01/mmi2017010004-
                                        abs.html.

Eeckhout:2017:MLU

[Eec17e] Lieven Eeckhout. Moore’s Law and ultra-low-power processors. IEEE Micro, 37(6):4– 5, November/December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060004.
html.

Eeckhout:2017:TTP

[Eec17f] Lieven Eeckhout. Thoughts on the top picks selections. IEEE Micro, 37(3): 4–5, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030004.
html.

Eeckhout:2018:ACI

[Eec18a] Lieven Eeckhout. Approximate computing, intelligent computing. IEEE Micro, 38 (4):6–7, July/August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040006.
html.

Eeckhout:2018:ACN

[Eec18b] Lieven Eeckhout. Automotive computing, neuromorphic computing, and beyond. IEEE Micro, 38(1):4–5, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010004.
html.

Eeckhout:2018:HAG

[Eec18c]  Lieven Eeckhout.  Hard-

ware acceleration and a grateful goodbye. IEEE Micro, 38(6):4–5, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/06/08585393-abs.
                                        html.

Eeckhout:2018:HC

[Eec18d] Lieven Eeckhout. Hot Chips 29. IEEE Micro, 38(2):6– 7, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/02/mmi2018020006.
                                        html.

Eeckhout:2018:MM

[Eec18e] Lieven Eeckhout. Memristors and more. IEEE Micro, 38(5): 4, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/05/mmi2018050004.
                                        html.

Eeckhout:2018:TP

[Eec18f]  Lieven Eeckhout.  Top

picks. IEEE Micro, 38(3): 4, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

mi/2018/03/mmi2018030004.
html.

Eeckhout:2023:KCA

[Eec23] Lieven Eeckhout. Kaya for computer architects: Toward sustainable computer systems. IEEE Micro, 43(1):9–18, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Enriquez:1995:FCR

[EEJ95] Angel L. Enriquez, Adel H. Eltimsahy, and Mohsin M. Jamali. Flexible control for robot manipulators. IEEE Micro, 15(6):55–60, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eyerman:2007:TAA

[EEKS07] Stijn Eyerman, Lieven Eeckhout, Tejas Karkhanis, and James E. Smith. A top-down approach to architecting CPI component performance counters. IEEE Micro, 27(1):84– 93, January/February 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eggers:1997:SMP

[EEL^{+}97] Susan J. Eggers, Joel S.

Emer, Henry M. Levy, Jack L. Lo, Rebecca L. Stamm, and Dean M. Tullsen. Simultaneous multithreading: a platform for next-generation processors. IEEE Micro, 17(5): 12–19, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5012.
pdf;   http://www.computer.
org/micro/mi1997/m5012abs.
htm.

Edenfield:1990:PPM

[EGL^{+}90a] Robin W. Edenfield, Michael G. Gallup, William B. Ledbetter, Jr., Ralph C. McGarity, Eric E. Quintana, and Russel A. Reininger. The 68040 processor: Part 2, memory design and chip verification. IEEE Micro, 10(3): 22–35, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Edenfield:1990:PPD

[EGL^{+}90b] Robin W. Edenfield, Michael G. Gallup, William B. Ledbetter, Jr., Ralph C. McGarity, Eric E. Quintana, and Russell A. Reininger. The 68040 processor: Part I, design and implementation. IEEE Micro, 10(1):66–78, January/February 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emer:2007:STV

[EHP^{+}07] Joel Emer, Mark D. Hill, Yale N. Patt, Joshua J. Yi, Derek Chiou, and Resit Sendag. Single-threaded vs. multithreaded: Where should we focus? IEEE Micro, 27(6): 14–24, November/December 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

El-Imam:1987:PCB

[EI87] Yousif A. El-Imam. A personal computer-based speech analysis and synthesis system. IEEE Micro, 7(3):4–21, May/ June 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

El-Imam:1990:TSC

[EIB90] Yousif A. El-Imam and Karima Banat. Text-to-speech conversion on a personal computer. IEEE Micro, 10(4):62– 74, July/August 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eichen:1986:NMP

[Eic86]  Bill Eichen.  NEC’s Mu-

Pd77230 digital signal processor. IEEE Micro, 6(6): 60–69, November/December 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Engel:2016:HWS

[EK16] Andreas Engel and Andreas Koch. Heterogeneous wireless sensor nodes that target the Internet of Things. IEEE Micro, 36(6):8–15, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060008-
abs.html.

Engebretsen:1996:PFO

[EKB^{+}96] David R. Engebretsen, Daniel M. Kuchta, Richard C. Booth, John D. Crow, and Wayne G. Nation. Parallel fiber-optic SCI links. IEEE Micro, 16 (1):20–26, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eichfeld:1995:GPF

[EKM^{+}95] Herbert Eichfeld, Martin Klimke, Manfred Menke, Jürgen Nolles, and Thomas Künemund. A generalpurpose fuzzy inference processor — defining, debugging, and processing complete fuzzy systems in real time. IEEE Micro, 15(3): 12–17, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eschmann:2002:SEC

[EKMW02] Frank Eschmann, Bernd Klauer, Ronald Moore, and Klaus Waldschmidt. SDAARC: An extended cache-only memory architecture. IEEE Micro, 22(3):62–70, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m3062.
pdf;   http://www.computer.
org/micro/mi2002/m3062abs.
htm.

Emmerson:1984:FTA

[EM84] Richard Emmerson and Michael J.

McGowan. Fault tolerance achieved in VLSI. IEEE Micro, 4(6):34–43, November/December 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emma:2005:MIWb

[Emm05a] Phil Emma. Micro innovations: Writing the claims for a patent. IEEE Micro, 25(6): 79–81, November/December 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emma:2005:PFF

[Emm05b] Phil Emma. Patents: To file or not to file? IEEE Micro, 25(5):79–81, September/October 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emma:2005:MII

[Emm05c] Philip G. Emma. Micro innovations: Inventions and the creative process. IEEE Micro, 25(3):96, 93–95, May/ June 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emma:2005:MIWa

[Emm05d]  Phillip G. Emma.  Mi-

cro innovations: What is patentable? IEEE Micro, 25(4):7–9, July/August 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emma:2006:MIH

[Emm06a] Phil Emma. Micro innovations: How to write a patent. IEEE Micro, 26(1):144, 143, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emma:2006:MIF

[Emm06b]  Philip G. Emma.  Micro

innovations: Five strategies for overcoming obviousness. IEEE Micro, 26(6):72, 70–71, November/December 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Emma:2006:MIPa

[Emm06c] Philip G. Emma. Micro innovations: Patent claims revisited: Examiners and trolls. IEEE Micro, 26(3):96, 94– 95, May/June 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Emma:2006:MIPb

[Emm06d] Philip G. Emma. Micro innovations: Prosecuting your patent. IEEE Micro, 26(5):88, 87, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Emma:2006:MIB

[Emm06e] Philip G. Emma. Micro innovations: The best patents of all. IEEE Micro, 26(2):84, 83, March/April 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Emma:2006:MIM

[Emm06f] Philip G. Emma. Micro innovations: The mechanics of filing a patent. IEEE Micro, 26 (4):88, 87, July/August 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Emma:2007:MIA

[Emm07a] Philip Emma. Micro innovations: Arcane facts and new words: Expanding your creative talent. IEEE Micro, 27(3):112, 110– 111, May/June 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Emma:2007:MII

[Emm07b] Philip Emma. Micro innovations: Innovation or notoriety? IEEE Micro, 27(4):

64, 62–63, July/August 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Emma:2007:MIR

[Emm07c] Philip Emma. Micro innovations: Reinventing entrepreneurial inventing for the 21st Century. IEEE Micro, 27(1):136, 134–135, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Emma:2007:MIS

[Emm07d] Philip Emma. Micro innovations: Supercharging your creative skills. IEEE Micro, 27(2):88, 86–87, March/April 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://

bell.computer.org/dlcomments/

.

Emma:2007:MIY

[Emm07e] Philip Emma. Micro innovations: You’re invited to a party! (How to hold a collaborative IP-development session). IEEE Micro, 27(6):64, 62–63, November/December 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Emma:2008:MIC

[Emm08a] Philip Emma. Micro innovations: a collaborative IPdevelopment session. IEEE Micro, 28(1):112, 110–111, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Emma:2008:GEI

[Emm08b]  Philip G. Emma.  Guest

Editor’s introduction: Existential architectures: The metaphysics of computer design. IEEE Micro, 28(6):4– 6, November/December 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Edahiro:2000:SCM

[EMYN00] Masato Edahiro, Satoshi Matsushita, Masakazu Yamashina, and Naoki Nishi. A single-chip multiprocessor for smart terminals. IEEE Micro, 20(4):12–20, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4012.
pdf;   http://www.computer.
org/micro/mi2000/m4012abs.
htm.

English:2000:MNA

[Eng00a]  Marie English.  Micro

news: Advancing Silk technology. IEEE Micro, 20(4): 2, July/August 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

English:2000:MNCb

[Eng00b] Marie English. Micro news: Chip production. IEEE Micro, 20(4):2, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

English:2000:MNCa

[Eng00c]  Marie English.  Micro

news: Chip technology breakthrough. IEEE Micro, 20(3): 3–3, 86–87, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m3003.
                                        pdf.

English:2000:MNDb

[Eng00d]  Marie English.  Micro

news: Digital-signature legislation. IEEE Micro, 20(4): 4, July/August 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

English:2000:MNDa

[Eng00e] Marie English. Micro news: Doubling computer memory. IEEE Micro, 20(4): 3, July/August 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

English:2000:MNE

[Eng00f] Marie English. Micro news: Embedded Linux Consortium begun. IEEE Micro, 20(3):3, May/June 2000.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3003.
pdf.

English:2000:MNY

[Eng00g] Marie English. Micro news:

Have you heard the story about bent nanowires? IEEE Micro, 20(4):4, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4002.
pdf.

English:2000:MNIa

[Eng00h] Marie English. Micro news:

Increased fab activity in 2000. IEEE Micro, 20(3):3, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3003.
pdf.

English:2000:MNIb

[Eng00i] Marie English. Micro news:

Intel releases Itanium guide. IEEE Micro, 20(3):3, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3003.
pdf.

English:2000:MNIc

[Eng00j] Marie English. Micro news: IrDA and Bluetooth: down with cords!; targeting lowk; help for designers; IBM demos quantum computer; GPS compatible with UWB?; chip benefits fiber optics; IEEE President’s awards; price comparison for PDAs; IETF explores standards. IEEE Micro, 20(5):3–4, 85– 86, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5003.
pdf.

English:2000:MNM

[Eng00k] Marie English. Micro news: Microcode development services. IEEE Micro, 20(4): 2, July/August 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

English:2000:MNN

[Eng00l] Marie English. Micro news: New benchmark for Unigraphics V15; wireless applications grow; tool set for the Java Card platform; biomechanical discovery affects mobile applications, robots; hard to navigate Web. IEEE Micro, 20(3):3–3, 86–87, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi2000/pdf/m3003.
pdf.

English:2000:MNP

[Eng00m] Marie English. Micro news:

Partnerships continue to aid industry. IEEE Micro, 20(4): 2, July/August 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

English:2000:MNSb

[Eng00n] Marie English. Micro news:

SIA forecasts growth. IEEE Micro, 20(4):2, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

English:2000:MNSa

[Eng00o] Marie English. Micro news:

Single-chip device for settop boxes. IEEE Micro, 20(3):3, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3003.
pdf.

English:2000:MNSc

[Eng00p]  Marie English.  Micro

news: Smaller, faster, cheaper chips. IEEE Micro, 20(4):3– 4, July/August 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eeckhout:2003:SSA

[ENSD03] Lieven Eeckhout, Sebastien Nussbaum, James E. Smith, and Koen De Bosschere. Statistical simulation: Adding efficiency to the computer designer’s toolbox. IEEE Micro, 23(5):26–38, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/05/m5026abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/05/m5026.
pdf.

Esmaeilzadeh:2019:MLA

[EP19] H. Esmaeilzadeh and J. Park. Machine learning acceleration. IEEE Micro, 39(5):6– 7, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Elthakeb:2020:RRL

[EPM^{+}20] A. T. Elthakeb, P. Pilligundla, F. Mireshghallah, A. Yazdanbakhsh, and H. Esmaeilzadeh. ReLeQ : A reinforcement learning approach for automatic deep quantization of neural networks. IEEE Micro, 40(5):37–45, September/October 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ellims:2002:DAR

[EPZ02] Michael Ellims, Stephen Parker, and James Zurlo. Design and analysis of a robust real-time engine control network. IEEE Micro, 22(4):20–27, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4020.
pdf;   http://www.computer.
org/micro/mi2002/m4020abs.
htm.

Emma:2008:RRI

[ERM08] Philip G. Emma, William R.

Reohr, and Mesut Meterelliyoz. Rethinking refresh: Increasing availability and reducing power in DRAM for cache applications. IEEE Micro, 28(6):47–56, November/December 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Edmondson:1995:SIE

[ERPR95] John H. Edmondson, Paul Rubinfeld, Ronald Preston, and Vidya Rajagopalan. Superscalar instruction execution in the 21164 Alpha microprocessor. IEEE Micro, 15(2):33– 43, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Elsmore:1984:PIM

[ES84] Tim Elsmore and William R. Shields. The proposed IEEE 1000 microcomputer system bus standard. IEEE Micro, 4 (4):72–80, July/August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Esmaeilzadeh:2013:NAG

[ESCB13] Hadi Esmaeilzadeh, Adrian Sampson, Luis Ceze, and Doug Burger. Neural acceleration for general-purpose approximate programs. IEEE Micro, 33(3):16–27, May/ June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eberle:2005:ANG

[ESG^{+}05] Hans Eberle, Sheueling Shantz, Vipul Gupta, Nils Gura, Leonard Rarick, and Lawrence Spracklen. Accelerating nextgeneration public-key cryptosystems on general-purpose CPUs. IEEE Micro, 25(2):52– 59, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2005/02/m2052abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2005/02/m2052.
                                        pdf.

Earnshaw:1997:CCD

[ESW97] Richard W. Earnshaw, Lee D. Smith, and Kevin Welton. Challenges in crossdevelopment: Coping with debugging and simulation for increasingly complex deeply embedded systems. IEEE Micro, 17(4):28–36, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4028.
pdf.

Emer:2009:TPC

[ET09] Joel Emer and Dean Tullsen.

Top picks from the 2008 Computer Architecture Conferences. IEEE Micro, 29(1): 6–9, January/February 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Espasa:1997:EIL

[EV97] Roger Espasa and Mateo Valero. Exploiting instructionlevel and data-level parallelism. IEEE Micro, 17(5):20– 27, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5020.
pdf;   http://www.computer.
org/micro/mi1997/m5020abs.
htm.

Efthivoulidis:1998:FTC

[EVM^{+}98] Giorgos Efthivoulidis, Evangelos A. Verentziotis, Apostolos N. Meliones, Theodora A. Varvarigou, Antonios Kontizas, Geert Deconinck, and Vincenzo De Florio. Faulttolerant communication in embedded supercomputing. IEEE Micro, 18(5):42–52, September/October 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1998/
                                        pdf/m5042.pdf;   http://
                                        www.computer.org/micro/
                                        mi1998/m5042abs.htm.

Egawa:2023:SIC

[EW23] Ryusuke Egawa and Yasutaka Wada. Special issue on cool chips. IEEE Micro, 43(1):40– 41, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Egawa:2024:SIC

[EW24] Ryusuke Egawa and Yasutaka Wada. Special issue on COOL chips. IEEE Micro, 44(1):6–7, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eckert:2019:NCB

[EWW^{+}19] C. Eckert, X. Wang, J. Wang, A. Subramaniyan, D. Sylvester, D. Blaauw, R. Das, and R. Iyer. Neural cache: Bitserial in-cache acceleration of deep neural networks. IEEE Micro, 39(3):11–19, May/ June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Faggin:1996:GEI

[Fag96] Federico Faggin. Guest Editors’ introduction — the microprocessor. IEEE Micro, 16 (6):7–9, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Faggin:2021:BM

[Fag21] Federico Faggin. The birth of the microprocessor. IEEE Micro, 41(6):16–19, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fairclough:1982:NOI

[Fai82a] D. A. Fairclough. The nearoptimal instruction set — reply. IEEE Micro, 2(3):6–7, July/September 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fairclough:1982:UMI

[Fai82b] D. A. Fairclough. A unique microprocessor instruction set. IEEE Micro, 2(2):8– 18, April/June 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ferdman:2014:CSP

[FAK^{+}14] Michael Ferdman, Almutaz Adileh, Onur Kocberber, Stavros Volos, Mohammad Alisafaee, Djordje Jevdjic, Cansu Kaynak, Adrian Daniel Popescu, Anastasia Ailamaki, and Babak Falsafi. A case for specialized processors for scale-out workloads. IEEE Micro, 34(3):31–42, May/ June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Fancher:1996:MSP

[Fan96] C. H. Fancher. Motorola’s SC-49 — a public-key microcontroller. IEEE Micro, 16 (3):18, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farrell:1984:ATM

[Far84] James J. Farrell III. The advancing technology of Motorola’s microprocessors and microcomputers. IEEE Micro, 4(5):55–63, September/October 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farrell:1985:LSC

[Far85] James J. Farrell III. Largescale cost-effective packaging. IEEE Micro, 5(3):5–10, May/ June 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farrell:1986:BC

[Far86]  J. Farrell.  Benchmarks

(continued…). IEEE Micro, 6(5):3, September/October 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farrell:1987:OLA

[Far87] J. Farrell. An open-letter to Austin. IEEE Micro, 7(2):3– 91, March/April 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farrell:1988:MA

[Far88a] J. Farrell. Meetings in Anaheim. IEEE Micro, 8(4): 2, July/August 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farrell:1988:AR

[Far88b] J. J. Farrell. The Adams Report. IEEE Micro, 8(3): 3–4, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farrell:1991:DC

[Far91] J. J. Farrell. A decade of contributors. IEEE Micro, 11(1): 8–9, January/February 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Floyd:2011:IAE

[FAWR^{+}11] Michael Floyd, Malcolm Allen-Ware, Karthick Rajamani, Bishop Brock, Charles Lefurgy, Alan J. Drake, Lorena Pesantez, Tilman Gloekler, Jose A. Tierno, Pradip Bose, and Alper Buyuktosunoglu. Introducing the adaptive energy management features of the Power7 chip. IEEE Micro, 31(2):60– 75, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fathi:1987:DSR

[FBC87] Eli T. Fathi, Eloi Bosse, and Jean Caseault. A distributed system for real-time applications. IEEE Micro, 7(6): 21–28, November/December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ferrari:1996:ACS

[FBGB96] Alberto Ferrari, Antonella Bellettini, Roberto Guerrieri, and Giorgio Baccarani. An ASIC chip set for parallel fuzzy database mining — parallel hardware that executes fuzzy queries speeds up a computationally intensive task. IEEE Micro, 16(6): 60–67, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fields:2004:ICW

[FBHN04] Brian A. Fields, Rastislav Bodík, Mark D. Hill, and Chris J. Newburn. Interaction cost: For when event counts just don’t add up. IEEE Micro, 24(6):57–61, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6057.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6057.pdf.

Falcao:2022:SIA

[FC22] Gabriel Falcao and Joseph R.

Cavallaro. Special issue on artificial intelligence at the edge. IEEE Micro, 42(6):6– 8, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fu:2020:EMP

[FCY^{+}20] C. Fu, H. Chen, Z. Yang, F. Koushanfar, Y. Tian, and J. Zhao. Enhancing model parallelism in neural architecture search for multidevice system. IEEE Micro, 40(5): 46–55, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Flynn:2004:GEI

[FD04] Michael Flynn and Pradeep Dubey. Guest Editors’ introduction: Hot Chips 15— scaling the silicon mountain. IEEE Micro, 24(2):7– 9, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/02/m2007.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/02/m2007.
htm.

Foley:2017:UPP

[FD17] Denis Foley and John Danskin. Ultra-performance Pascal GPU and NVLink interconnect. IEEE Micro, 37(2): 7–17, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/02/mmi2017020007-
                                        abs.html.

Falsafi:2017:FVG

[FDS^{+}17] Babak Falsafi, Bill Dally, Desh Singh, Derek Chiou, Joshua J. Yi, and Resit Sendag. FP-GAs versus GPUs in data centers. IEEE Micro, 37(1):60– 72, January/February 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/01/mmi2017010060-
                                        abs.html.

Ferrero:1998:MVH

[Fer98a] J. L. Ferrero. Micro view: History of computing on display. IEEE Micro, 18 (3):88–??, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m3088.
                                        pdf.

Ferrero:1998:MVW

[Fer98b] J. L. Ferrero. Micro view: Wearable computing: One man’s mission. IEEE Micro, 18(5):88, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5088.
pdf.

Fet:1995:VPS

[Fet95] Yakov I. Fet. Vertical processing systems — a survey. IEEE Micro, 15(1):65–75, January/February 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fridman:2000:TDA

[FG00] Jose Fridman and Zvi Greenfield. The TigerSHARC DSP architecture. IEEE Micro, 20 (1):66–76, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1066.
pdf;   http://www.computer.
org/micro/mi2000/m1066abs.
htm.

Falsafi:2014:BDG

[FG14] Babak Falsafi and Boris Grot.

Big data [Guest Editors’ introduction]. IEEE Micro, 34(4): 4–5, July/August 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040004-
abs.html.

Fu:2014:SRT

[FGC^{+}14] Haohuan Fu, Lin Gan, Robert G. Clapp, Huabin Ruan, Oliver Pell, Oskar Mencer, Michael Flynn, Xiaomeng Huang, and Guangwen Yang. Scaling reverse time migration performance through reconfigurable dataflow engines. IEEE Micro, 34(1):30–40, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Fuccio:1988:DAS

[FGG^{+}88] Michael L. Fuccio, Renato N. Gadenz, Craig J. Garen, Joan M. Huser, Benjamin Ng, Steven P. Pekarich, and Kreg D. Ulery. The DSP32C: AT&T’s second-generation floating-point digital signal processor. IEEE Micro, 8(6): 30–48, November/December 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Frank:2000:CHP

[FH00] Edward H. Frank and Jack Holloway. Connecting the home with a phone line network chip set. IEEE Micro, 20(2):27–37, 39, March/ April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2027.
pdf;   http://www.computer.
org/micro/mi2000/m2027abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Flynn:2005:MDI

[FH05] Michael J. Flynn and Patrick Hung. Microprocessor design issues: Thoughts on the road ahead. IEEE Micro, 25(3): 16–31, May/June 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fryman:2003:EEN

[FHL^{+}03] Joshua B. Fryman, Chad M.

Huneycutt, Hsien-Hsin (Sean) Lee, Kenneth M. Mackenzie, and David E. Schimmel. Energy-efficient network memory for ubiquitous devices. IEEE Micro, 23(5):60– 70, September/October 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/05/m5060abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/05/m5060.
pdf.

Finkbeiner:2017:MI

[FHL^{+}17] Tim Finkbeiner, Glen Hush, Troy Larsen, Perry Lea, John Leidel, and Troy Manning. In-memory intelligence. IEEE Micro, 37(4):30– 38, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

mi/2017/04/mmi2017040030-
abs.html.

Faggin:1996:HDT

[FHMS96] Federico Faggin, Marcian E. Hoff, Jr., Stanley Mazor, and Masatoshi Shima. The history of the 4004 — the 4004 design team tells its story. IEEE Micro, 16(6):10–20, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Flynn:2000:UST

[FHP00] Michael J. Flynn, Patrick Hung, and Armita Peymandoust. Using simple tools to evaluate complex architectural trade-offs. IEEE Micro, 20(4):67–75, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4067.
pdf;   http://www.computer.
org/micro/mi2000/m4067abs.
htm.

Flynn:1999:DSM

[FHR99] Michael J. Flynn, Patrick Hung, and Kevin W. Rudd. Deep-submicron microprocessor design issues. IEEE Micro, 19(4):11–22, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4011.
pdf;   http://www.computer.

org/micro/mi1999/m4011abs.
htm.

Fischer:1985:IPS

[Fis85]  Wayne Fischer.  IEEE

P1014 — a standard for the high-performance VME bus. IEEE Micro, 5(1):31–41, January/February 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fizza:2022:ESD

[FJB^{+}22] Kaneez Fizza, Prem Prakash Jayaraman, Abhik Banerjee, Dimitrios Georgakopoulos, and Rajiv Ranjan. Evaluating sensor data quality in Internet of Things smart agriculture applications. IEEE Micro, 42 (1):51–60, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Feehrer:2013:OST

[FJL^{+}13] John Feehrer, Sumti Jairath, Paul Loewenstein, Ram Sivaramakrishnan, David Smentek, Sebastian Turullols, and Ali Vahidsafa. The Oracle Sparc T5 16-core processor scales to eight sockets. IEEE Micro, 33 (2):48–57, March/April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fathi:1983:ETD

[FK83] Eli T. Fathi and Moshe Krieger. An executive for taskdriven multimicrocomputer systems. IEEE Micro, 3(5): 32–41, September/October 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Feng:2001:FDT

[FKL01] Wenyi Feng, Farzin Karimi, and Fabrizio Lombardi. Fault detection in a tristate system environment. IEEE Micro, 21(5):77–85, September/October 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
m5077abs.htm;   http://
dlib.computer.org/mi/books/
mi2001/pdf/m5077.pdf.

Farjadrad:2020:BWB

[FKV20] R. Farjadrad, M. Kuemerle, and B. Vinnakota. A Bunchof-Wires (BoW) interface for interchiplet communication. IEEE Micro, 40(1):15–24, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Furht:1984:ESD

[FL84] Borivoje Furht and Peter Lee. An efficient software driver for Am9511 arithmetic processor implementation. IEEE Micro, 4(3):7–19, May/June 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Falsafi:2013:GEI

[FL13] Babak Falsafi and Gabriel H.

Loh. Guest Editors’ introduction: Top picks from the 2012 Computer Architecture Conferences. IEEE Micro, 33(3): 4–7, May/June 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Flanagan:1999:MRJ

[Fla99] D. Flanagan. Micro review:

Java power reference. IEEE Micro, 19(5):10–11, September/October 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Frantz:1986:TIT

[FLRB86] Gene A. Frantz, Kun-Shan Lin, Jay B. Reimer, and Jon Bradley. The Texas Instruments TMS320C25 digital signal microcomputer. IEEE Micro, 6(6):10–28, November/December 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Flynn:1997:AER

[Fly97] David Flynn. AMBA: Enabling reusable on-chip designs: Supporting the firsttime-right design of low-power embedded systems. IEEE Micro, 17(4):20–27, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4020.
pdf.

Faudemay:1991:AAL

[FM91] Pascal Faudemay and Mongia Mhiri. An associative accelerator for large databases. IEEE Micro, 11(6):22–34, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Farley:2018:APN

[FME18] Brendan Farley, John Mc-Grath, and Christophe Erdmann. An all-programmable 16-nm RFSoC for digital-RF communications. IEEE Micro, 38(2):61–71, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2018/02/mmi2018020061-
abs.html.

Fukuoka:2013:PMF

[FMN^{+}13] Kazuki Fukuoka, Noriaki Maeda, Koji Nii, Masaki Fujigaya, Noriaki Sakamoto, Takao Koike, Takahiro Irita, Kohei Wakahara, Tsugio Matsuyama, Keiji Hasegawa, Toshiharu Saito, Akira Fukuda, Kaname Teranishi, Takeshi Kataoka, and Toshihiro Hattori. Power-management features of R-Mobile U2, an integrated application processor and baseband processor. IEEE Micro, 33(6):

26–36, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Fukuda:1991:TAP

[FMT91] Akira Fukuda, Kazuaki Murakami, and Shinji Tomita. Toward advanced parallel processing: exploiting parallelism at task and instruction levels. IEEE Micro, 11(4):16–19, 50– 61, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Faro:1985:MBS

[FMV85] Alberto Faro, Orazio Mirabella, and Lorenzo Vita. A multimicrocomputer-based structure for computer networking. IEEE Micro, 5(2):53– 66, March/April 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fuccio:1986:HAC

[FN86] Michael L. Fuccio and Benjamin Ng. Hardware architecture considerations in the WE32100 chip set. IEEE Micro, 6(2):29–46, March/April 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Foty:1994:MTL

[FN94] Daniel P. Foty and Edward J.

Nowak. MOSFET technology for low-voltage/low-power applications. IEEE Micro, 14(3):

68–77, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Franklin:1989:PMI

[FO89] David F. Franklin and David V. Ostler. The P1073 medical information bus. IEEE Micro, 9(5):52–60, September/October 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fowers:2019:IPB

[FOP^{+}19] J. Fowers, K. Ovtcharov, M. K. Papamichael, T. Massengill, M. Liu, D. Lo, S. Alkalay, M. Haselman, L. Adams, M. Ghandi, S. Heil, P. Patel, A. Sapek, G. Weisz, L. Woods, S. Lanka, S. K. Reinhardt, A. M. Caulfield, E. S. Chung, and D. Burger. Inside Project Brainwave’s cloud-scale, realtime AI processor. IEEE Micro, 39(3):20–28, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Forsell:2002:SHP

[For02]  Martti Forsell.  A scal-

able high-performance computing solution for networks on chips. IEEE Micro, 22(5): 46–55, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5046.

pdf;   http://www.computer.
org/micro/mi2002/m5046abs.
htm.

Fossum:1998:DCS

[Fos98] Eric R. Fossum. Digital camera system on a chip. IEEE Micro, 18(3):8–15, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3008.
pdf;   http://www.computer.
org/micro/mi1998/m3008abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24– 26, 1997.

Ferreira:2002:FCP

[FPAF02] Joaquim Ferreira, Paulo Pedreiras, Luís Almeida, and José Alberto Fonseca. The FTT-CAN protocol for flexibility in safety-critical systems. IEEE Micro, 22(4): 46–55, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4046.
pdf;   http://www.computer.
org/micro/mi2002/m4046abs.
htm.

Fraase:1994:WIT

[Fra94]  M. Fraase.  The Win-

dows Internet tour guide. IEEE Micro, 14(6):5, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fraser:1996:FWT

[Fra96] Alexander G. Fraser. Future Wan telecommunications. IEEE Micro, 16(1):53–57, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Frantz:2000:DSP

[Fra00] Gene Frantz. Digital signal processor trends. IEEE Micro, 20(6):52–59, November/December 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2000/
                                        pdf/m6052.pdf;   http://
                                        www.computer.org/micro/
                                        mi2000/m6052abs.htm.

Fu:2018:MSQ

[FRB^{+}18] X. Fu, M. A. Rol, C. C. Bultink, J. van Someren, N. Khammassi, I. Ashraf, R. F. L. Vermeulen, J. C. de Sterke, W. J. Vlothuizen, R. N. Schouten, C. G. Almudever, L. DiCarlo, and K. Bertels. A microarchitecture for a superconducting quantum processor. IEEE Micro, 38(3):40– 47, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

mi/2018/03/mmi2018030040-
abs.html.

Fredriksson:2002:CCE

[Fre02] Lars-Berno Fredriksson. CAN for critical embedded automotive networks. IEEE Micro, 22(4):28–35, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4028.
pdf;   http://www.computer.
org/micro/mi2002/m4028abs.
htm.

Feehrer:2009:CHD

[FRS^{+}09] John Feehrer, Paul Rotker, Milton Shih, Paul Gingras, Peter Yakutis, Stephen Phillips, and John Heath. Coherency hub design for multisocket Sun servers with CoolThreads technology. IEEE Micro, 29(4):36–47, July/ August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fromm:2005:MLU

[FS05] Jeffery B. Fromm and Robert A.

Skitol. Micro law: Update on the antitrust ghost in the standard-setting machine. IEEE Micro, 25(5):77– 79, September/October 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fung:2012:KTH

[FSBA12] Wilson W. L. Fung, Inderpreet Singh, Andrew Brownsword, and Tor M. Aamodt. Kilo TM: Hardware transactional memory for GPU architectures. IEEE Micro, 32(3): 7–16, May/June 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Friedrich:2001:SCC

[FSH^{+}01] L. Fernando Friedrich, John Stankovic, Marty Humphrey, Michael Marley, and John Haskins, Jr. A survey of configurable, component-based operating systems for embedded applications. IEEE Micro, 21(3):54–68, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3054.
pdf;   http://www.computer.
org/micro/mi2001/m3054abs.
htm.

Feldmann:2022:AAC

[FSK^{+}22] Axel Feldmann, Nikola Samardzic, Aleksandar Krastev, Srinivas Devadas, Ron Dreslinski, Chris Peikert, and Daniel Sanchez. An architecture to accelerate computation on encrypted data. IEEE Micro, 42 (4):59–68, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Falcon:2005:BBP

[FSR^{+}05] Ayose Falcón, Jared Stark, Alex Ramirez, Konrad Lai, and Mateo Valero. Better branch prediction through prophet/critic hybrids. IEEE Micro, 25(1):80–89, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1080.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1080.pdf.

Falsafi:2016:NMD

[FSS^{+}16] Babak Falsafi, Mircea Stan, Kevin Skadron, Nuwan Jayasena, Yunji Chen, Jinhua Tao, Ravi Nair, Jaime Moreno, Naveen Muralimanohar, Karthikeyan Sankaralingam, and Cristian Estan. Near-memory data services. IEEE Micro, 36(1):6– 13, January/February 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/01/mmi2016010006.
html.

Frieder:1992:EHD

[FTKS92] Ophir Frieder, Vijaykumar A.

Topkar, Ramesh K. Karne, and Arun K. Sood. Experimentation with hypercube database engines. IEEE Micro, 12(1):42–56, January/February 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fulcher:1991:FGM

[Ful91]  John A. Fulcher.  Fun

and games and microcomputer interfacing. IEEE Micro, 11(1):18–21, 75–78, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Furht:1988:RAT

[Fur88] Borivoje Furht. A RISC architecture with two-size, overlapping register windows. IEEE Micro, 8(2):67–80, March/ April 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Faraboschi:2012:TPC

[FV12] Paolo Faraboschi and T. N. Vijaykumar. Top picks from the 2011 Computer Architecture Conferences. IEEE Micro, 32(3):3–6, May/June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fan:2012:GEM

[FZW^{+}12] Dongrui Fan, Hao Zhang, Da Wang, Xiaochun Ye, Fenglong Song, Guojie Li, and Ninghui Sun. Godson-T: An efficient many-core processor exploring thread-level parallelism. IEEE Micro, 32(2):38– 47, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gabbay:1986:DDA

[GA86] David Gabbay and Joseph Appelbaum. The design of a DC/AC inverter with the MC68HC11 microcomputer. IEEE Micro, 6(1):16–23, January/February 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grossman:2021:XSX

[GA21] M. Grossman and J. Andrews.

The Xbox Series X system architecture. IEEE Micro, 41 (2):22–28, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Guttag:1988:TEM

[GAAR88] Karl M. Guttag, Thomas M.

Albers, Michael D. Asal, and Kevin G. Rose. The TMS34010 — an embedded microprocessor. IEEE Micro, 8(3):39–52, May/June 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

GadelRab:2007:GEC

[Gad07] Serag GadelRab. 10-gigabit Ethernet connectivity for computer servers. IEEE Micro, 27(3):94–105, May/June 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gafford:1991:RMS

[Gaf91] J. D. Gafford. Rate monotonic scheduling. IEEE Micro, 11(3):34–??, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ghosh:2022:SFM

[GAGV22] Debjani Ghosh, Akash Anand, Satya Sankalp Gautam, and Ankit Vidyarthi. Soil fertility monitoring with internet of underground things: a survey. IEEE Micro, 42(1):8–16, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Galles:1997:SHS

[Gal97] Mike Galles. Spider: a highspeed network interconnect — raising the bandwidth ceiling with a scalable, pipelined interconnect for distributed endpoint routing. IEEE Micro, 17 (1):34–39, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gonzalez:2007:SRE

[GALB07] Ivan Gonzalez, Estanislao Aguayo, and Sergio Lopez-Buedo. Self-reconfigurable embedded systems on lowcost FPGAs. IEEE Micro, 27 (4):49–57, July/August 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Garcia:1993:SSU

[Gar93]  D. Linda Garcia.  Stan-

dard setting in the United States: public and private sector roles. IEEE Micro, 13(6): 28–35, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gandhi:2006:SLS

[GAR^{+}06] Amit Gandhi, Haitham Akkary, Ravi Rajwar, Srikanth T. Srinivasan, and Konrad Lai. Scalable load and store processing in latency-tolerant processors. IEEE Micro, 26 (1):30–39, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gass:2021:EHT

[Gas21] Wanda Gass. Early history of Texas Instrument’s digital signal processor. IEEE Micro, 41(6):129–130, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ganguly:2022:IDQ

[GAT^{+}22] Amlan Ganguly, Sergi Abadal, Ishan Thakkar, Natalie Enright Jerger, Marc Riedel, Masoud Babaie, Rajeev Balasubramonian, Abu Sebastian, Sudeep Pasricha, and Baris Taskin. Interconnects for DNA, quantum, in-memory, and optical computing: Insights from a panel discussion. IEEE Micro, 42(3):40– 49, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gokhale:2020:EFQ

[GBD^{+}20] P. Gokhale, J. M. Baker, C. Duckering, F. T. Chong, N. C. Brown, and K. R. Brown. Extending the frontier of quantum computers with qutrits. IEEE Micro, 40(3): 64–72, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grisenthwaite:2023:AME

[GBW^{+}23] Richard Grisenthwaite, Graeme Barnes, Robert N. M. Watson, Simon W. Moore, Peter Sewell, and Jonathan Woodruff. The Arm Morello Evaluation Platform validating CHERI-based security in a high-performance system. IEEE Micro, 43(3):50– 57, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Giri:2021:AIO

[GCE^{+}21] Davide Giri, Kuan-Lin Chiu, Guy Eichler, Paolo Mantovani, and Luca P. Carloni. Accelerator integration for opensource SoC design. IEEE Micro, 41(4):8–14, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gliksberg:2020:HQF

[GCL^{+}20] J. Gliksberg, A. Capra, A. Louvet, P. J. García, and D. Sohier. High-quality fault resiliency in fat trees. IEEE Micro, 40(1):44–49, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gyger:2001:EAT

[GD01] Thomas Gyger and Olivier Desjeux. EasyRide: Active transponders for a fare collection system. IEEE Micro, 21(6):36–42, November/December 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
m6036abs.htm;   http://
dlib.computer.org/mi/books/
mi2001/pdf/m6036.pdf.

Gilbert:2008:GUW

[GDES08] Jeffrey M. Gilbert, Chinh H.

Doan, Sohrab Emami, and C. Bernard Shung. A 4-Gbps uncompressed wireless HD A/V transceiver chipset. IEEE Micro, 28(2):56–64, March/April 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gaudiot:1986:THP

[GDLT86] Jean-Luc Gaudiot, Michel Dubois, Liang-Teh Lee, and Nadim G. Tohme. The TX 16: a highly programmable multi-microprocessor architecture. IEEE Micro, 6(5):18– 31, September/October 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gao:2017:DLP

[GDN^{+}17] Mingyu Gao, Christina Delimitrou, Dimin Niu, Krishna T. Malladi, Hongzhong Zheng, Bob Brennan, and Christos Kozyrakis. DRAF: A low-power DRAM-based reconfigurable acceleration fabric. IEEE Micro, 37(3):70– 78, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030070-
abs.html.

Gavrielov:1986:NFP

[GE86] Moshe Gavrielov and Lev Epstein. The NS32081 floatingpoint unit — architecture and implementation. IEEE Micro, 6(2):6–12, March/April 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gupta:2023:ACF

[GEH^{+}23] Udit Gupta, Mariam Elgamal, Gage Hills, Gu-Yeon Wei, Hsien-Hsin S. Lee, David Brooks, and Carole-Jean Wu. Architectural CO2 footprint tool: Designing sustainable computer systems with an architectural carbon modeling tool. IEEE Micro, 43(4):107– 117, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gervasi:2019:WCN

[Ger19]  B. Gervasi.  Will carbon

nanotube memory replace DRAM? IEEE Micro, 39 (2):45–51, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gan:2017:SMA

[GFL^{+}17] Lin Gan, Haohuan Fu, Wayne Luk, Chao Yang, Wei Xue, and Guangwen Yang. Solving mesoscale atmospheric dynamics using a reconfigurable dataflow architecture. IEEE Micro, 37(4):40–50, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040040-
abs.html.

Gabrielli:1999:FDF

[GG99] Alessandro Gabrielli and Enzo Gandolfi. A fast digital fuzzy processor. IEEE Micro, 19 (1):68–79, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1068.
pdf;   http://www.computer.
org/micro/mi1999/m1068abs.
htm.

Grant:2016:HI

[GG16] Ryan E. Grant and Ada Gavrilovska. Hot interconnects 23. IEEE Micro, 36(4): 4–5, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040004.
html.

Guo:2015:RTC

[GGB^{+}15] Qing Guo, Xiaochen Guo, Yuxin Bai, Ravi Patel, Engin Ipek, and Eby G. Friedman. Resistive ternary content addressable memory systems for data-intensive computing. IEEE Micro, 35(5):62– 71, September/October 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050062-
abs.html.

Gonzalez:2011:SWS

[GGC^{+}11] Juan Gonzalez, Judit Gimenez, Marc Casas, Miquel Moreto, Alex Ramirez, Jesus Labarta, and Mateo Valero. Simulating whole supercomputer applications. IEEE Micro, 31(3): 32–45, May/June 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gupta:1996:AVS

[GGJ^{+}96] Bhusan Gupta, Rodney Goodman, Fukang Jiang, Yu-Chong Tai, Steev Tung, and Chih-Ming Ho. Analog VLSI system for active drag reduction — using a biological inspiration to build a drag-reducing surface with VLSI technology. IEEE Micro, 16(5):53– 59, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ghosh:2022:SIH

[GGS22] Sayan Ghosh, Ryan E. Grant, and Min Si. Special issue on hot interconnects. IEEE Micro, 42(2):35–36, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grimes:1988:EPA

[GH88] J. Grimes and J. Hootman.

Embedded processors — another cost-effective design tool. IEEE Micro, 8(3):8– 9, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grant:2020:HI

[GH20] R. E. Grant and K. Hamidouche. Hot Interconnects 26. IEEE Micro, 40(1):6–7, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gschwind:2006:SPC

[GHF^{+}06] Michael Gschwind, H. Peter Hofstee, Brian Flachs, Martin Hopkins, Yukio Watanabe, and Takeshi Yamazaki. Synergistic processing in Cell’s multicore architecture. IEEE Micro, 26(2):10–24, March/April 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grot:2012:ODC

[GHLK^{+}12] Boris Grot, Damien Hardy, Pejman Lotfi-Kamran, Babak Falsafi, Chrysostomos Nicopoulos, and Yiannakis Sazeides. Optimizing data-center TCO with scale-out processors. IEEE Micro, 32(5):52– 63, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Govindaraju:2012:DUF

[GHN^{+}12] Venkatraman Govindaraju, Chen-Han Ho, Tony Nowatzki, Jatin Chhugani, Nadathur Satish, Karthikeyan Sankaralingam, and Changkyu Kim. DySER: Unifying functionality and parallelism specialization for energy-efficient computing. IEEE Micro, 32(5): 38–51, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gee:1993:CPS

[GHPS93] Jeffrey D. Gee, Mark D. Hill, Dionisios N. Pnevmatikatos, and Alan Jay Smith. Cache performance of the Spec92 benchmark suite. IEEE Micro, 13(4):17–27, July/August 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Goser:1989:VTA

[GHRS89] Karl Goser, Ulrich Hilleringmann, Ulrich Rückert, and Klaus Schumacher. VLSI technologies for artificial neural networks. IEEE Micro, 9(6):28–44, November/December 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gandhi:2017:APE

[GHS17] Jayneel Gandhi, Mark D. Hill, and Michael M. Swift. Agile paging for efficient memory virtualization. IEEE Micro, 37(3):80–86, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030080-
abs.html.

Goulding-Hotta:2011:GMA

[GHSV^{+}11] Nathan Goulding-Hotta, Jack Sampson, Ganesh Venkatesh, Saturnino Garcia, Joe Auricchio, Po-Chao Huang, Manish Arora, Siddhartha Nath, Vikram Bhatt, Jonathan Babb, Steven Swanson, and Michael Bedford Taylor. The GreenDroid Mobile Application Processor: An architecture for silicon’s dark future. IEEE Micro, 31(2):86– 95, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Guo:2017:SHC

[GHY^{+}17] Kaiyuan Guo, Song Han, Song Yao, Yu Wang, Yuan Xie, and Huazhong Yang. Softwarehardware codesign for efficient neural network acceleration. IEEE Micro, 37(2):18– 25, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020018-
abs.html.

Gilbert:1982:GPI

[Gil82]  Richard Gilbert.  The

general-purpose interface bus. IEEE Micro, 2(1):41–51, January/March 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Giladi:1996:EMM

[Gil96a] Ran Giladi. Evaluating the Mflops measure: Searching for a consistent performance measure to cover a variety of computer configurations. IEEE Micro, 16(4):69– 75, July/August 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gillett:1996:MCN

[Gil96b] Richard B. Gillett. Memory channel network for PCI. IEEE Micro, 16(1):12–18, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Garcia:2012:KOS

[GJLT12] Saturnino Garcia, Donghwan Jeon, Christopher Louie, and Michael Bedford Taylor. The Kremlin Oracle for sequential code parallelization. IEEE Micro, 32(4):42– 53, July/August 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gillett:1997:UMC

[GK97] Richard Gillett and Richard Kaufmann. Using the Memory Channel Network: Using a cluster of standard PCI-based servers with a low-cost network to improve communication performance. IEEE Micro, 17(1):19–25, January/February 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gandhi:2016:RTF

[GKA^{+}16] Jayneel Gandhi, Vasileios Karakostas, Furkan Ayar, Adrian Cristal, Mark D. Hill, Kathryn S. McKinley, Mario Nemirovsky, Michael M. Swift, and Osman S. Unsal. Range translations for fast virtual memory. IEEE Micro, 36 (3):118–126, May/June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030118-
abs.html.

Gouk:2023:MPC

[GKB^{+}23] Donghyun Gouk, Miryeong Kwon, Hanyeoreum Bae, Sangwon Lee, and Myoungsoo Jung. Memory pooling with CXL. IEEE Micro, 43 (2):48–57, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Goiri:2014:DMD

[GKL^{+}14] Inigo Goiri, William Katsak, Kien Le, Thu D. Nguyen, and Ricardo Bianchini. Designing and managing data centers powered by renewable energy. IEEE Micro, 34(3):8– 16, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Gupta:2022:CCE

[GKL^{+}22] Udit Gupta, Young Geun Kim, Sylvia Lee, Jordan Tse, Hsien-Hsin S. Lee, Gu-Yeon Wei, David Brooks, and Carole-Jean Wu. Chasing carbon: The elusive environmental footprint of computing. IEEE Micro, 42(4):37– 47, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gold:2005:TRS

[GKS^{+}05] Brian T. Gold, Jangwoo Kim, Jared C. Smolens, Eric S. Chung, Vasileios Liaskovitis, Eriko Nurvitadhi, Babak Falsafi, James C. Hoe, and Andreas G. Nowatzyk. TRUSS: a reliable, scalable server architecture. IEEE Micro, 25(6): 51–59, November/December 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gurumurthi:2006:UST

[GKS06] Sudhanva Gurumurthi, Youngjae Kim, and Anand Sivasubramaniam. Using STEAM for thermal simulation of storage systems. IEEE Micro, 26 (4):43–51, July/August 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gratz:2007:CIN

[GKS^{+}07] Paul Gratz, Changkyu Kim, Karthikeyan Sankaralingam, Heather Hanson, Premkishore Shivakumar, Stephen W. Keckler, and Doug Burger. On-chip interconnection networks of the TRIPS chip. IEEE Micro, 27(5):41–50, September/October 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gendler:2021:DIF

[GKS21] Alex Gendler, Ernest Knoll, and Yiannakis Sazeides. I-DVFS: Instantaneous frequency switch during dynamic voltage and frequency scaling. IEEE Micro, 41(5):76– 84, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gan:2022:PSM

[GLD^{+}22] Yu Gan, Mingyu Liang, Sundar Dev, David Lo, and Christina Delimitrou. Practical and scalable ML-driven cloud performance debugging with Sage. IEEE Micro, 42 (4):27–36, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Garland:2008:PCE

[GLN^{+}08] Michael Garland, Scott Le Grand, John Nickolls, Joshua Anderson, Jim Hardwick, Scott Morton, Everett Phillips, Yao Zhang, and Vasily Volkov. Parallel computing experiences with CUDA. IEEE Micro, 28(4):13–27, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gupta:1999:DIF

[GM99] Pankaj Gupta and Nick McKeown. Designing and implementing a fast crossbar scheduler. IEEE Micro, 19 (1):20–28, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1020.
pdf;   http://www.computer.
org/micro/mi1999/m1020abs.
htm.

Gupta:2000:CPH

[GM00] Pankaj Gupta and Nick McKeown. Classifying packets with hierarchical intelligent cuttings. IEEE Micro, 20 (1):34–41, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1034.
pdf;   http://www.computer.
org/micro/mi2000/m1034abs.
htm.

Galles:2021:PDS

[GM21] M. Galles and F. Matus. Pensando distributed services architecture. IEEE Micro, 41 (2):43–49, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Giri:2018:ACS

[GMC18] Davide Giri, Paolo Mantovani, and Luca P. Carloni. Accelerators and coherence: An SoC perspective. IEEE Micro, 38(6):36–45, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.

                                        computer.org/csdl/mags/
                                        mi/2018/06/08525419.pdf.

Gupta:1983:ACB

[GmDT83] Amar Gupta and Hoo min D. Toong. An architectural comparison of 32-bit microprocessors. IEEE Micro, 3(1): 9–22, January/February 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gonzalez:2007:RFR

[GMM^{+}07] Antonio González, Scott Mahlke, Shubu Mukherjee, Resit Sendag, Derek Chiou, and Joshua J. Yi. Reliability: Fallacy or reality? IEEE Micro, 27(6):36–45, November/December 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Golston:1996:SCH

[Gol96] Jeremiah Golston. Singlechip H.324 videoconferencing — using the TMS320C82 DSP for multimedia processing in cost-sensitive consumer and PC markets. IEEE Micro, 16 (4):21–33, July/August 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Goldman:2021:LCM

[Gol21]  Murray Goldman.  Last

chance: The Motorola microprocessor story. IEEE Micro, 41(6):148–149, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gong:1997:JSP

[Gon97]  Li Gong.  Java security:

Present and near future: Coping with the rapidly evolving security issues of crossplatform computing. IEEE Micro, 17(3):14–19, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://pascal.computer.
org/mi/books/mi1997/pdf/
m3014.pdf.

Gonzales:1999:MRA

[Gon99] David Ruimy Gonzales. Micro-RISC architecture for the wireless market. IEEE Micro, 19(4):30–37, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4030.
pdf;   http://www.computer.
org/micro/mi1999/m4030abs.
htm.

Gonzalez:2000:XCE

[Gon00] Ricardo E. Gonzalez. Xtensa — a configurable and extensible processor. IEEE Micro, 20(2):60–70, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2060.
pdf;   http://www.computer.

                                        org/micro/mi2000/m2060abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Gonzalez:2006:SCP

[Gon06]  Ricardo E. Gonzalez.  A

software-configurable processor architecture. IEEE Micro, 26(5):42–51, September/October 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gonzalez:2018:ISC

[Gon18] Antonio Gonzalez. 2018 International Symposium on Computer Architecture Influential Paper Award. IEEE Micro, 38 (4):76–77, July/August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/04/mmi2018040076.
                                        html.

Goodman:1984:MDC

[Goo84] J. R. Goodman. Microprocessors defy classification. IEEE Micro, 4(3):3–6, May/June 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Goodman:2014:REM

[Goo14] James Goodman. Reflections from the 2013 Eckert-Mauchly Award recipient. IEEE Micro, 34(3):149–151, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Goodacre:2021:BAM

[Goo21] John Goodacre. The birth of Arm multicore processing. IEEE Micro, 41(6):150–152, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Govers:1990:EAT

[GP90] Francis X. Govers III and Michael J. Pierson. On the edge — analysis of transactions and computers. IEEE Micro, 10(5):73–75, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Guo:1995:HSR

[GP95] S. W. Guo and L. Peters. A high-speed, reconfigurable fuzzy logic controller. IEEE Micro, 15 (6):65, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gustavson:1983:PTD

[GPSS83] D. B. Gustavson, T. Pittman, M. Smolin, and R. G. Stewart. Publication is timely and democratic. IEEE Micro, 3(3):58–60, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Garcia:2006:ESC

[GQF^{+}06] Pedro J. García, Francisco J.

Quiles, José Flich, José Duato, Ian Johnson, and Finbar Naven. Efficient, scalable congestion management for interconnection networks. IEEE Micro, 26(5):52–66, September/October 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grosspietsch:1992:APS

[GR92] Karl E. Grosspietsch and Ralf Reetz. The associative processor system CAPRA: Architecture and applications. IEEE Micro, 12(6):58–67, November/December 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Graf:1995:GEI

[GR95a] Hans Peter Graf and Leonardo M. Reyneri. Guest Editors’ introduction: The expanding world of neural and fuzzy systems. IEEE Micro, 15(3): 10–11, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Graf:1995:NNE

[GR95b] Hans Peter Graf and Leonardo M. Reyneri. Neural networks — extraordinary variation. IEEE Micro, 15(3):48–59, May/ June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gorius:2022:SSA

[GRD22] Jean-Michel Gorius, Simon Rokicki, and Steven Derrien. SpecHLS: Speculative accelerator design using high-level synthesis. IEEE Micro, 42(5): 99–107, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1993:MEM

[Gre93] Shane Greenstein. Micro economics: Markets, standards, and the information infrastructure. IEEE Micro, 13(6): 36–51, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1995:MED

[Gre95a] Shane Greenstein. Micro economics: Diamonds are forever; computers are not. IEEE Micro, 15(4):4–5, July/August 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1995:MEO

[Gre95b] Shane Greenstein. Micro economics: Operating systems, soap operas: What’s the difference? IEEE Micro, 15 (6):3–4, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1995:MEM

[Gre95c] Shane Greenstein. Micro economics: The market and computer technology. IEEE Micro, 15(3):8–9, 78, May/June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1995:MEW

[Gre95d] Shane Greenstein. Micro economics: Why aren’t the mainframes disappearing as fast as everyone thought? IEEE Micro, 15(5):4–5, September/October 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1996:MEC

[Gre96a]  Shane Greenstein.  Micro

economics: a case study illustrates innovative computing and what makes it work. IEEE Micro, 16(1):68–69, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1996:MED

[Gre96b] Shane Greenstein. Micro economics: Don’t call it a highway! IEEE Micro, 16(6): 78–79, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1996:MEP

[Gre96c] Shane Greenstein. Micro economics: Is the productivity paradox to blame for your computing woes? IEEE Micro, 16(4):2–3, July/August 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1996:MER

[Gre96d] Shane Greenstein. Micro economics: Repetitive stress injuries: Who pays? IEEE Micro, 16(5):72–73, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1996:MES

[Gre96e] Shane Greenstein. Micro economics: Shooting for par: Are designers and consumers playing on the same course? IEEE Micro, 16(3):2, 79, May/June 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1996:MEW

[Gre96f] Shane Greenstein. Micro economics: Why should technical curmudgeons celebrate the microprocessor’s anniversary? IEEE Micro, 16(2): 3–4, March/April 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1997:MECa

[Gre97a] S. Greenstein. Micro economics: Can biology explain the economics of technology? IEEE Micro, 17 (4):3–5, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

                                        /pascal.computer.org/mi/
                                        books/mi1997/pdf/m4003.
                                        pdf.

Greenstein:1997:MECb

[Gre97b] S. Greenstein. Micro economics: Convergence hype. IEEE Micro, 17(5):7–8, September/October 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1997/
                                        pdf/m5007.pdf.

Greenstein:1997:MELa

[Gre97c] S. Greenstein. Micro economics: Let’s play Monopoly. IEEE Micro, 17(3):73–74, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://

                                        pascal.computer.org/mi/
                                        books/mi1997/pdf/m3073.
                                        pdf.

Greenstein:1997:MELb

[Gre97d] S. Greenstein. Micro economics: Looking at software from an academic perspective. IEEE Micro, 17 (6):6–7, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1997/pdf/m6006.
                                        pdf.

Greenstein:1997:MEW

[Gre97e] S. Greenstein. Micro economics: Why fun is so often associated with our technology. IEEE Micro, 17(2): 5–6, March/April 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1997:MEH

[Gre97f]  Shane Greenstein.  Mi-

cro economics: How coinvention shapes our market. IEEE Micro, 17(1):2–3, January/February 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1998:MECa

[Gre98a] S. Greenstein. Micro economics: Closing the door on foreclosure. IEEE Micro, 18(2):4–5, 83–84, March/ April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2004.
pdf.

Greenstein:1998:MECb

[Gre98b] S. Greenstein. Micro economics: Commercializing the Internet. IEEE Micro, 18 (6):6–7, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6006.
pdf.

Greenstein:1998:MER

[Gre98c] S. Greenstein. Micro economics: Return of the jaded: What would a Hollywood film about the PC industry look like? IEEE Micro, 18 (5):4–5, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m5004.
                                        pdf.

Greenstein:1998:HHN

[Gre98d] S. Greenstein. To have and to have not. IEEE Micro, 18 (1):76–84, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:1998:MEI

[Gre98e]  Shane Greenstein.  Micro

economics: Industrial economics and strategy: Computing platforms. IEEE Micro, 18(3):43–53, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m3043.
                                        pdf;   http://www.computer.
                                        org/micro/mi1998/m3043abs.
                                        htm.

Greenstein:1998:MEU

[Gre98f]  Shane Greenstein.  Mi-

cro economics: Uncertainty, prediction, and the unexpected. IEEE Micro, 18 (4):76–77, July/August 1998.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4076.
pdf.

Greenstein:1999:MEF

[Gre99a] S. Greenstein. Micro economics: Forecasting commercial change. IEEE Micro, 19(3):6–7, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3006.
pdf.

Greenstein:1999:MEBc

[Gre99b] Shane Greenstein. Micro economics: Banking on the Information Age. IEEE Micro, 19 (5):5–6, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5005.
pdf.

Greenstein:1999:MEBa

[Gre99c] Shane Greenstein. Micro economics: Bill, adopt a mensch strategy! IEEE Micro, 19(2):8, 83–84, March/ April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2008.
pdf.

Greenstein:1999:MEBb

[Gre99d] Shane Greenstein. Micro economics: Building the virtual world. IEEE Micro, 19(4):5–6, 86, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m4005.
                                        pdf.

Greenstein:1999:MEV

[Gre99e] Shane Greenstein. Micro economics: Virulent word of mouse. IEEE Micro, 19 (6):6–8, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m6006.
                                        pdf.

Greenstein:1999:MEW

[Gre99f]  Shane Greenstein.  Micro

economics: When technologies converge. IEEE Micro, 19(1):8–9, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m1008.
                                        pdf.

Greenstein:2000:MER

[Gre00a] Shane Greenstein. Micro economics: a revolution? how do you know? IEEE Micro, 20(2):5–7, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2005.
pdf.

Greenstein:2000:MEA

[Gre00b]  Shane Greenstein.  Micro

economics: Aggressive business tactics: are there limits? IEEE Micro, 20(1):13– 14, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1013.
pdf.

Greenstein:2000:MEF

[Gre00c]  Shane Greenstein.  Micro

economics: Falling through the cracks at Microsoft. IEEE Micro, 20(5):7–7, 87– 87, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5007.
pdf;   http://www.computer.
org/micro/mi2000/m5007abs.
htm.

Greenstein:2000:MEH

[Gre00d]  Shane Greenstein.  Micro

economics: Hung up on AT&T? IEEE Micro, 20 (4):5–6, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

                                        books/mi2000/pdf/m4005.
                                        pdf.

Greenstein:2000:MEP

[Gre00e] Shane Greenstein. Micro economics: PCs, the Internet, and you. IEEE Micro, 20 (6):6–7, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m6006.
                                        pdf.

Greenstein:2000:MEE

[Gre00f]  Shane Greenstein.  Mi-

cro economics: The era of impatience. IEEE Micro, 20(3):8–9, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m3008.
                                        pdf.

Greenstein:2001:MEEa

[Gre01a]  Shane Greenstein.  Micro

economics: An earful about Zvi’s e-mail. IEEE Micro, 21 (4):8–10, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m4008.
                                        pdf;   m4008abs.htm.

Greenstein:2001:MEB

[Gre01b] Shane Greenstein. Micro economics: E-business infrastructure. IEEE Micro, 21(6):

70–71, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6070abs.htm;
http://dlib.computer.org/
mi/books/mi2001/pdf/m6070.
pdf.

Greenstein:2001:MEEb

[Gre01c]  Shane Greenstein.  Mi-

cro economics: Explaining booms, busts, and errors. IEEE Micro, 21(5): 6–7, 10, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5006abs.htm;
http://dlib.computer.org/
mi/books/mi2001/pdf/m5006.

pdf. See corrections [Ano01a].

Greenstein:2001:MEH

[Gre01d]  Shane Greenstein.  Mi-

cro economics: How does the new economy stack up? IEEE Micro, 21(1):11– 12, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1011.
pdf.

Greenstein:2001:MEP

[Gre01e]  Shane Greenstein.  Micro

economics: Pricing Internet access. IEEE Micro, 21 (2):5–6, March/April 2001.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2005.
pdf.

Greenstein:2001:MES

[Gre01f]  Shane Greenstein.  Micro

economics: Shortfalls, downturns, and recessions. IEEE Micro, 21(3):6, 77, May/ June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3006.
pdf.

Greenstein:2002:MECa

[Gre02a]  Shane Greenstein.  Micro

economics: Competition policy for innovative industries. IEEE Micro, 22(2): 4–5, 72, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2004.
pdf;   http://www.computer.
org/micro/mi2002/m2004abs.
htm.

Greenstein:2002:MEM

[Gre02b]  Shane Greenstein.  Mi-

cro economics: Markets for technology. IEEE Micro, 22(3):4–5, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi2002/pdf/m3004.
pdf;   http://www.computer.
org/micro/mi2002/m3004abs.
htm.

Greenstein:2002:MECb

[Gre02c] Shane Greenstein. Micro economics: The crash in competitive telephony. IEEE Micro, 22(4):8–9, 88, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4008.
pdf;   http://www.computer.
org/micro/mi2002/m4008abs.
htm.

Greenstein:2002:MEP

[Gre02d] Shane Greenstein. Micro economics: The price is not right. IEEE Micro, 22(5):12– 13, 96, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5012.
pdf.

Greenstein:2002:MER

[Gre02e] Shane Greenstein. Micro economics: The ride before the fall. IEEE Micro, 22(1): 4–5, 91, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1004abs.htm;
 http://dlib.computer.org/
mi/books/mi2002/pdf/m1004.
pdf.

Greenstein:2002:MEW

[Gre02f]  Shane Greenstein.  Mi-

cro economics: Which industries use the Internet? IEEE Micro, 22(6):70–72, November/December 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m6070.
                                        pdf.

Greenstein:2003:MEE

[Gre03a]  Shane Greenstein.  Micro

economics: An era of error. IEEE Micro, 23(3):68– 69, May/June 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2003/03/m3068.pdf.

Greenstein:2003:MEJ

[Gre03b]  Shane Greenstein.  Mi-

cro economics: Jumping on bandwagons. IEEE Micro, 23(5):75–77, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2003/05/m5075.pdf.

Greenstein:2003:MEM

[Gre03c] Shane Greenstein. Micro economics: Moore meets Malthus in multiples. IEEE Micro, 23 (4):8–10, July/August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2003/04/m4008.pdf.

Greenstein:2003:MET

[Gre03d]  Shane Greenstein.  Micro

economics: Too much Internet backbone? IEEE Micro, 23(2):78–80, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m2078.
pdf.

Greenstein:2003:MEW

[Gre03e] Shane Greenstein. Micro economics: Where did the Internet go? IEEE Micro, 23 (1):84–86, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1084.
pdf;   http://www.computer.
org/micro/mi2003/m1084abs.
htm.

Greenstein:2004:MECb

[Gre04a]  Shane Greenstein.  Mi-

cro economics: Canaries, whips, and sails. IEEE Micro, 24(6):6–7, 129, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6006.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6006.pdf.

Greenstein:2004:MECa

[Gre04b] Shane Greenstein. Micro economics: Creative destruction and deconstruction. IEEE Micro, 24(5):83–85, September/October 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/05/m5083.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/05/m5083.pdf.

Greenstein:2004:MEI

[Gre04c]  Shane Greenstein.  Micro

economics: Imitation happens. IEEE Micro, 24(3): 67–69, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/03/m3067.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/03/m3067.pdf.

Greenstein:2004:MED

[Gre04d] Shane Greenstein. Micro economics: The diamond-wafer paradox: a modern mystery. IEEE Micro, 24(4):79– 81, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/04/m4079.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/04/m4079.pdf.

Greenstein:2004:MEP

[Gre04e] Shane Greenstein. Micro economics: The paradox of commodities. IEEE Micro, 24 (2):73–75, March/April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/02/m2073.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/02/m2073.pdf.

Greenstein:2004:MEW

[Gre04f]  Shane Greenstein.  Micro

economics: Why inventors are not famous. IEEE Micro, 24(1):76–78, January/February 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/01/m1076.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/01/m1076.pdf.

Greenstein:2005:MEC

[Gre05a] Shane Greenstein. Micro economics: Communications consolidation after an era of no restraints. IEEE Micro, 25(2): 72, 70–71, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2072abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2005/02/m2072.
pdf.

Greenstein:2005:MEE

[Gre05b]  Shane Greenstein.  Micro

economics: Explorers and expanders, both early and late. IEEE Micro, 25(4):77– 79, July/August 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2005:MEM

[Gre05c] Shane Greenstein. Micro economics: Not a mellifluous march to maturity. IEEE Micro, 25(1):104, 102–103, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2005/01/m1104.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2005/01/m1104.pdf.

Greenstein:2005:MEO

[Gre05d]  Shane Greenstein.  Micro

economics: Outsourcing and climbing a value chain. IEEE Micro, 25(5):84, 83, September/October 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2005:MEA

[Gre05e] Shane Greenstein. Micro economics: The anatomy of foresight traps. IEEE Micro, 25(3):10–12, May/June 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2005:MEW

[Gre05f] Shane Greenstein. Micro economics: Wireless access and electrical markets: Becoming similar? IEEE Micro, 25 (6):6–7, November/December 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2006:MEA

[Gre06a]  Shane Greenstein.  Micro

economics: Andy’s acceleration and Moore’s momentum. IEEE Micro, 26(2): 7, 81–82, March/April 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2006:MEFa

[Gre06b] Shane Greenstein. Micro economics: Format wars all over again. IEEE Micro, 26(1):7, 140, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2006:MEFb

[Gre06c]  Shane Greenstein.  Micro

economics: Four nightmares for net neutrality. IEEE Micro, 26(6):12–13, November/December 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2006:MEL

[Gre06d]  Shane Greenstein.  Micro

economics: Legislating entrepreneurship: An oxymoron? IEEE Micro, 26(4): 4, 86, July/August 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2006:MER

[Gre06e] Shane Greenstein. Micro economics: Room for a thousand flowers to bloom. IEEE Micro, 26(3):6, 93, May/June 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2006:MEU

[Gre06f] Shane Greenstein. Micro economics: Ubiquitous clicks and complements. IEEE Micro, 26 (5):7–8, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2007:MEDa

[Gre07a] Shane Greenstein. Micro economics: Did the price of the Internet drop? IEEE Micro, 27(3):6–7, May/June 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2007:MEDb

[Gre07b]  Shane Greenstein.  Mi-

cro economics: Dog days for broadband. IEEE Micro, 27(5):112, 111, September/October 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Greenstein:2007:MEI

[Gre07c] Shane Greenstein. Micro economics: Innovation at the edges. IEEE Micro, 27(6): 8–10, November/December 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2007:MEB

[Gre07d] Shane Greenstein. Micro economics: The 15-billion-dollar broadband bonus. IEEE Micro, 27(4):5, 58, July/August 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2007:MEH

[Gre07e] Shane Greenstein. Micro economics: The high cost of a cheap lesson. IEEE Micro, 27(1):7, 132–133, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2007:MEW

[Gre07f] Shane Greenstein. Micro economics: Wagging Wikipedia’s long tail. IEEE Micro, 27(2): 6, 79, March/April 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2008:MEC

[Gre08a] Shane Greenstein. Micro economics: Chicken Little predictions. IEEE Micro, 28(6):2– 3, November/December 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2008:MES

[Gre08b] Shane Greenstein. Micro economics: Slouching toward a dystopian Internet. IEEE Micro, 28(5):6–7, September/October 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2008:MELa

[Gre08c]  Shane Greenstein.  Micro

economics: The long arc behind Bill Gates’ wealth. IEEE Micro, 28(1):4–7, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2008:MELb

[Gre08d] Shane Greenstein. Micro economics: The long arc behind Bill Gates’ wealth, part 2. IEEE Micro, 28(2):2– 5, March/April 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2008:MEV

[Gre08e] Shane Greenstein. Micro economics: Voting and economic asymmetry. IEEE Micro, 28(4):2–3, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2009:MEN

[Gre09a] Shane Greenstein. Micro economics: a network of platforms. IEEE Micro, 29(6):2– 3, November/December 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2009:MEB

[Gre09b] Shane Greenstein. Micro economics: Building broadband as economic stimulus. IEEE Micro, 29(2):2–3, March/ April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2009:MED

[Gre09c] Shane Greenstein. Micro economics: Does Google have too much money? IEEE Micro, 29 (5):6–7, September/October 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2009:MESb

[Gre09d] Shane Greenstein. Micro economics: Soccer mom messaging is the poetry of our age. IEEE Micro, 29(4):2– 3, July/August 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2009:MESa

[Gre09e] Shane Greenstein. Micro economics: Symptoms of healthy innovativeness. IEEE Micro, 29(1):3–5, January/February 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2009:MER

[Gre09f] Shane Greenstein. Micro economics: The revolution in spectrum allocation. IEEE Micro, 29(3):4–6, May/June 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2010:BBA

[Gre10a] Shane Greenstein. Building broadband ahead of digital demand. IEEE Micro, 30(6):6– 8, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2010:DVC

[Gre10b]  Shane Greenstein.  Dig-

itization and value creation. IEEE Micro, 30(4):4– 5, July/August 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2010:GE

[Gre10c] Shane Greenstein. Gatekeeping economics. IEEE Micro, 30(5):102–104, September/October 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2010:MEB

[Gre10d] Shane Greenstein. Micro economics: Bleeding-edge mass market standards. IEEE Micro, 30(2):2–4, March/April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2010:MES

[Gre10e] Shane Greenstein. Micro economics: Standardization and coordination. IEEE Micro, 30 (3):6–7, May/June 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2010:MEN

[Gre10f] Shane Greenstein. Micro economics: The next chapter at Google. IEEE Micro, 30(1): 4–7, January/February 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2011:DDM

[Gre11a] Shane Greenstein. Digital dark matter. IEEE Micro, 31(1):128, January/February 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2011:DBS

[Gre11b] Shane Greenstein. The direction of broadband spillovers. IEEE Micro, 31(2):104, 103, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2011:MEH

[Gre11c] Shane Greenstein. Micro economics: An honest policy wonk. IEEE Micro, 31(4):80, 79, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2011:MEW

[Gre11d]  Shane Greenstein.  Micro

economics: The wi-fi journey. IEEE Micro, 31(5):80, 79, September/October 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2011:OIO

[Gre11e] Shane Greenstein. The Open Internet order. IEEE Micro, 31(3):88, 87, May/June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2011:SJE

[Gre11f]  Shane Greenstein.  Steve

Jobs and the economics of one entrepreneur. IEEE Micro, 31(6):64–65, November/December 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2012:MEB

[Gre12a] Shane Greenstein. Micro economics: a big payoff. IEEE Micro, 32(2):64, C3, March/ April 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2012:MEC

[Gre12b]  Shane Greenstein.  Mi-

cro economics: Calm economics. IEEE Micro, 32(4): 72, July/August 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2012:MEM

[Gre12c] Shane Greenstein. Micro economics: Managing complements. IEEE Micro, 32(6): 64, 63, November/December 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2012:MEP

[Gre12d]  Shane Greenstein.  Micro

economics: The prevailing view. IEEE Micro, 32(5): 80, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2012:MES

[Gre12e] Shane Greenstein. Micro economics: The secret life of Wally Madhavani. IEEE Micro, 32(3):142–143, May/June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2012:RLL

[Gre12f] Shane Greenstein. The range of Linus’ Law. IEEE Micro, 32 (1):72, 71, January/February 2012. CODEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2013:GS

[Gre13a] Shane Greenstein. Gaming structure. IEEE Micro, 33(1): 88, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2013:MEDa

[Gre13b] Shane Greenstein. Micro economics: Differentiated platforms. IEEE Micro, 33(3): 120, May/June 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2013:MEDb

[Gre13c]  Shane Greenstein.  Micro

economics: Digital public goods. IEEE Micro, 33(5):62– 63, September/October 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2013:MEH

[Gre13d]  Shane Greenstein.  Mi-

cro economics: How much Apache? IEEE Micro, 33 (6):80, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Greenstein:2013:MEO

[Gre13e] Shane Greenstein. Micro economics: The online honesty box. IEEE Micro, 33(2):62– 63, March/April 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2013:PCM

[Gre13f]  Shane Greenstein.  Plat-

form conflicts [micro economics]. IEEE Micro, 33 (4):78–79, July/August 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2014:BDL

[Gre14a] Shane Greenstein. Baking the data layer [micro economics]. IEEE Micro, 34(4): 56–c3, July/August 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040056-
abs.html.

Greenstein:2014:EVD

[Gre14b] Shane Greenstein. Enough variety and diversity? IEEE Micro, 34(5):70–71, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050070-
abs.html.

Greenstein:2014:MEA

[Gre14c] Shane Greenstein. Micro economics: The academic and business marriage. IEEE Micro, 34(3):152–c3, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Greenstein:2014:MEF

[Gre14d] Shane Greenstein. Micro economics: The fault lines along fast lanes. IEEE Micro, 34(2): 64, March/April 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Greenstein:2014:MEI

[Gre14e] Shane Greenstein. Micro economics: The irony of public funding. IEEE Micro, 34 (1):94–95, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Greenstein:2015:BBB

[Gre15a] Shane Greenstein. Behind the buzz of behavioral data. IEEE Micro, 35(2):88–c3, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020088-
abs.html.

Greenstein:2015:IOE

[Gre15b] Shane Greenstein. Insiders, outsiders, and an existentialist. IEEE Micro, 35(6):72–c3, November/December 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060072-
abs.html.

Greenstein:2015:NSR

[Gre15c]  Shane Greenstein.  Net-

working standards and Russell’s revisionism (review of, “Open standards and the digital age: history, ideology, and networks, Russel, A.; 2014) [book review]. IEEE Micro, 35(1):64– c3, January/February 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010064.
html.

Greenstein:2015:TYCa

[Gre15d] Shane Greenstein. Twenty years of the commercial Internet, part 1. IEEE Micro, 35 (3):150–152, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030150-
abs.html.

Greenstein:2015:TYCb

[Gre15e] Shane Greenstein. Twenty years of the commercial Internet, part 2. IEEE Micro, 35(5):86–88, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050086-
abs.html.

Greenstein:2015:WGM

[Gre15f]  Shane Greenstein.  Who

is Gordon Moore, and why is there a law named for him? IEEE Micro, 35(4):80– c3, July/August 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/04/mmi2015040080-
                                        abs.html.

Greenstein:2016:CLM

[Gre16a] Shane Greenstein. Congestion on the last mile. IEEE Micro, 36(6):62–63, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/06/mmi2016060062-
                                        abs.html.

Greenstein:2016:EGT

[Gre16b] Shane Greenstein. Economic growth from technical advance. IEEE Micro, 36(3): 130–131, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/03/mmi2016030130-
                                        abs.html.

Greenstein:2016:NRS

[Gre16c]  Shane Greenstein.  No

Rosetta Stone for market vitality. IEEE Micro, 36(5):62– 63, September/October 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/05/mmi2016050062-
                                        abs.html.

Greenstein:2016:TOQ

[Gre16d]  Shane Greenstein.  Ten

open questions for technooptimists. IEEE Micro, 36 (4):86–87, July/August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040086-
abs.html.

Greenstein:2016:WDS

[Gre16e]  Shane Greenstein.  What

does a skunk works do? IEEE Micro, 36(2):70–71, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020070-
abs.html.

Greenstein:2017:HHN

[Gre17a] Shane Greenstein. The hushhush norm. IEEE Micro, 37(6):92–95, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060092-
abs.html.

Greenstein:2017:IP

[Gre17b]  Shane Greenstein.  In-

sider privileges. IEEE Micro, 37(5):70–72, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/05/mmi2017050070-
                                        abs.html.

Greenstein:2017:MLE

[Gre17c] Shane Greenstein. Moore’s Law and economic architectures. IEEE Micro, 37(4):82– 84, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi2017040082-
                                        abs.html.

Greenstein:2017:TPT

[Gre17d] Shane Greenstein. Technology policy and the Trump Administration. IEEE Micro, 37(1):78–79, January/ February 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/01/mmi2017010078-
                                        abs.html.

Greenstein:2017:TSS

[Gre17e] Shane Greenstein. Two sides to scale. IEEE Micro, 37(3): 130–131, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030130-
                                        abs.html.

Greenstein:2017:VFG

[Gre17f] Shane Greenstein. The value of free in GDP. IEEE Micro, 37(2):106–107, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2017/02/mmi2017020106-
abs.html.

Greenstein:2018:AAT

[Gre18a] Shane Greenstein. Adjusting to autonomous trucking. IEEE Micro, 38(3):126– 128, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030126.
html.

Greenstein:2018:FSF

[Gre18b] Shane Greenstein. Free software without a free lunch or free beer. IEEE Micro, 38(5):94–96, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050094.
html.

Greenstein:2018:OCC

[Gre18c] Shane Greenstein. Organized for cycles of change. IEEE Micro, 38(6):86–88, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08585395-abs.
html.

Greenstein:2018:PTD

[Gre18d] Shane Greenstein. The paradox of technological déjà vu. IEEE Micro, 38(1):118–120, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/01/mmi2018010118.
                                        html.

Greenstein:2018:TT

[Gre18e] Shane Greenstein. The technology tel. IEEE Micro, 38 (4):78–80, July/August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/04/mmi2018040078.
                                        html.

Greenstein:2019:ADD

[Gre19a] S. Greenstein. The aftermath of the Dyn DDOS attack. IEEE Micro, 39(4):66– 68, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2019:ATA

[Gre19b] S. Greenstein. Antitrust in three acts. IEEE Micro, 39(6): 82–84, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2019:ESM

[Gre19c] S. Greenstein. Earning stripes in medical machine learning. IEEE Micro, 39(5): 126–128, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2019:MMA

[Gre19d] S. Greenstein. Misapplied metaphors in AI policy. IEEE Micro, 39(3):118–120, May/ June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2019:SIT

[Gre19e] S. Greenstein. Six infrastructure trends. IEEE Micro, 39 (1):70–72, January/February 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2019:WFT

[Gre19f] S. Greenstein. Where the frontier thrives: Bricks, mix, and zip. IEEE Micro, 39(2):62– 64, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2020:EOF

[Gre20a] S. Greenstein. Expertise at our fingertips. IEEE Micro, 40 (2):74–76, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2020:FSP

[Gre20b] S. Greenstein. The fox and shepherd problem. IEEE Micro, 40(6):86–88, November/December 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2020:PDT

[Gre20c] S. Greenstein. Pandemics and the dismal technology economy. IEEE Micro, 40(3):118– 120, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2020:TTA

[Gre20d]  S. Greenstein.  Triggers,

transmissions, and adjustments. IEEE Micro, 40(5):88– 90, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2020:UEW

[Gre20e] S. Greenstein. Uncomfortable economic waters. IEEE Micro, 40(4):134–136, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2020:VTP

[Gre20f] S. Greenstein. The vital two percent. IEEE Micro, 40 (1):94–96, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2021:ECC

[Gre21a] S. Greenstein. The economics of confrontational conversation. IEEE Micro, 41(2):86– 88, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2021:TPD

[Gre21b] S. Greenstein. Technology policy dilemmas in the new administration. IEEE Micro, 41 (1):86–88, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2021:EDI

[Gre21c] Shane Greenstein. Economic dependencies in integrated circuits. IEEE Micro, 41(5): 130–132, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2021:RW

[Gre21d] Shane Greenstein. Remote work. IEEE Micro, 41(3):110– 112, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2021:SIC

[Gre21e] Shane Greenstein. Shortages of integrated circuits. IEEE Micro, 41(4):86–88, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2021:VC

[Gre21f]  Shane Greenstein.  Vir-

tuous cycles. IEEE Micro, 41(6):184–186, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2022:ARD

[Gre22a] Shane Greenstein. Archetypes of risky decisions. IEEE Micro, 42(5):130–132, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2022:DDS

[Gre22b] Shane Greenstein. Distributed discretion by the slice. IEEE Micro, 42(6):142–144, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2022:GAS

[Gre22c] Shane Greenstein. Google and Apple signed a deal. IEEE Micro, 42(1):138–140, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2022:GBD

[Gre22d] Shane Greenstein. Growth from breadth and depth. IEEE Micro, 42(3):86–88, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2022:ITM

[Gre22e]  Shane Greenstein.  Infla-

tion and technology markets. IEEE Micro, 42(4):134– 136, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2022:TCU

[Gre22f] Shane Greenstein. Time for a change in U.S. antitrust for technology? IEEE Micro, 42 (2):86–88, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2023:AGR

[Gre23a] Shane Greenstein. The AI Gold Rush. IEEE Micro, 43(6):126–128, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2023:BRW

[Gre23b] Shane Greenstein. Bank runs without the wisdom of the crowds. IEEE Micro, 43(3): 86–88, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2023:BEL

[Gre23c] Shane Greenstein. Butterfly effects and legacies. IEEE Micro, 43(2):142–144, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2023:IRC

[Gre23d]  Shane Greenstein.  Inter-

view with Ronnie Chatterji, coordinator for the creating helpful incentives to produce semiconductors and science act. IEEE Micro, 43(5): 98–100, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2023:MDO

[Gre23e] Shane Greenstein. The modern digital operating model. IEEE Micro, 43(1):90–92, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2024:AGR

[Gre24a] Shane Greenstein. After the gold rush. IEEE Micro, 44 (1):76–78, January/February 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Greenstein:2024:PL

[Gre24b] Shane Greenstein. Party like it’s 1999? IEEE Micro, 44 (2):78–80, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grisenthwaite:2021:MDA

[Gri21] Richard Grisenthwaite. The milestones that define Arm’s past, present, and future. IEEE Micro, 41(6):58–67, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grogono:1983:CIS

[Gro83] P. Grogono. Copyrighting of IEEE standards challenged. IEEE Micro, 3(6):4, November/December 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grosspietsch:1992:APM

[Gro92a] K. E. Grosspietsch. Associativeprocessors and memories — a survey. IEEE Micro, 12(3): 12–19, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grosspietsch:1992:GEI

[Gro92b] K. E. Grosspietsch. Guest Editor’s introduction: Associative processors and memories. IEEE Micro, 12(3): 10–11, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grosspietsch:1994:FTH

[Gro94a] K. E. Grosspietsch. Fault tolerance in highly parallel hardware systems. IEEE Micro, 14 (1):60–68, January/February 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grosspietsch:1994:GEI

[Gro94b] Karl E. Grosspietsch. Guest Editor’s introduction: Fault tolerance. IEEE Micro, 14(1): 6–7, January/February 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grosspietsch:2002:GEI

[Gro02] Karl E. Grosspietsch. Guest Editor’s introduction: Unorthodox computer architectures. IEEE Micro, 22 (3):8–9, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m3008.
pdf;   http://www.computer.
org/micro/mi2002/m3008abs.
htm.

Grossner:1983:IWV

[GRP83] Clifford P. Grossner, Thiruvengadam Radhakrishnan, and Andy Pospiech. An integrated workstation for the visually handicapped. IEEE Micro, 3(3):8–16, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grossner:1986:IBD

[GRS86] Clifford P. Grossner, Thiruvengadam Radhakrishnan, and Alex Schena. An intelligent Braille display device. IEEE Micro, 6(3):43– 51, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gotz:1999:GEI

[GS99] Dieter Gotz and Anton Sauer. Guest Editors’ introduction: MEDEA: a successful European cooperation in microelectronics. IEEE Micro, 19(5):12– 15, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5012.
pdf.

Gokhale:2021:FC

[GS21] Maya Gokhale and Lesley Shannon. FPGA computing. IEEE Micro, 41(4):6– 7, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Geiger:1997:WNE

[GSC97] Robert L. Geiger, James D.

Solomon, and Kenneth J. Crisler. Wireless network extension using mobile IP. IEEE Micro, 17(6):63–68, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

                                        books/mi1997/pdf/m6063.
                                        pdf;   http://www.computer.
                                        org/micro/mi1997/m6063abs.
                                        htm.

Gu:2011:MAD

[GSLK11] Junli Gu, Yihe Sun, Steven S. Lumetta, and Rakeshh Kumar. MOPED: Accelerating data communication on future CMPs. IEEE Micro, 31(4):42– 50, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Giaccone:2002:IPS

[GSP02] Paolo Giaccone, Devavrat Shah, and Balaji Prabhakar. An implementable parallel scheduler for input-queued switches. IEEE Micro, 22 (1):19–25, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/m1019abs.htm;
                                         http://dlib.computer.org/
                                        mi/books/mi2002/pdf/m1019.
                                        pdf.

Gueron:2007:WDS

[GSS^{+}07] Shay Gueron, Jean-Pierre Seifert, Geoffrey Strongin, Derek Chiou, Resit Sendag, and Joshua J. Yi. Where does security stand? new vulnerabilities vs. trusted computing. IEEE Micro, 27(6): 25–35, November/December 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gurumurthi:2009:UIP

[GSS09] Sudhanva Gurumurthi, Sriram Sankar, and Mircea R. Stan. Using intradisk parallelism to build energy-efficient storage systems. IEEE Micro, 29(1):50–61, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gomaa:2003:TFR

[GSVP03] Mohamed A. Gomaa, Chad Scarbrough, T. N. Vijaykumar, and Irith Pomeranz. Transient-fault recovery for chip multiprocessors. IEEE Micro, 23(6):76–83, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6076abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/06/m6076.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6076.pdf.

Gustavson:1983:WLT

[GT83] David B. Gustavson and John Theus. Wire-OR logic on transmission-lines. IEEE Micro, 3(3):51–55, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gurumurthi:2022:SIT

[GT22] Sudhanva Gurumurthi and Radu Teodorescu. Special issue on top picks from the 2021 computer architecture conferences. IEEE Micro, 42(4): 6–9, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gonzalez:1997:ATO

[GTF97] R. González, A. Torralba, and L. G. Franquelo. AFAN — tool for optimizing fuzzy controllers. IEEE Micro, 17(5): 50–54, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5050.
pdf;   http://www.computer.
org/micro/mi1997/m5050abs.
htm.

Garzon:2022:AAM

[GTLY22] Esteban Garzón, Adam Teman, Marco Lanuzza, and Leonid Yavits. AIDA: Associative in-memory deep learning accelerator. IEEE Micro, 42(6):67–75, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Goldberg:1998:IDP

[GU98] S. R. Goldberg and S. J. Upadhyaya. Implementing degradable processing arrays. IEEE Micro, 18(1):64–74, January/February 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gunn:2006:CPH

[Gun06] Cary Gunn. CMOS photonics for high-speed interconnects. IEEE Micro, 26(2):58– 66, March/April 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gurumurthi:2009:PAS

[Gur09] Sudhanva Gurumurthi. Prolegomena: Architecting storage for the cloud computing era. IEEE Micro, 29(6): 68–71, November/December 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gustavson:1984:CBT

[Gus84] David B. Gustavson. Computer buses — a tutorial. IEEE Micro, 4(4):7– 22, July/August 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gustavson:1985:MBE

[Gus85] D. B. Gustavson. More on bigendian vs little-endian byte ordering. IEEE Micro, 5 (3):4, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gustavson:1992:SCI

[Gus92] David B. Gustavson. The Scalable Coherent Interface and related standards projects. IEEE Micro, 12 (1):10–22, January/February 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gillingham:1997:SHP

[GV97] Peter Gillingham and Bill Vogley. SLDRAM — highperformance, open-standard memory. IEEE Micro, 17(6): 29–39, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m6029.
pdf;   http://www.computer.
org/micro/mi1997/m6029abs.
htm.

Gomaa:2006:OTF

[GV06] Mohamed A. Gomaa and T. N. Vijaykumar. Opportunistic transient-fault detection. IEEE Micro, 26(1):92– 99, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grant:2021:HI

[GV21] R. E. Grant and M. G. Venkata. Hot Interconnects 27. IEEE Micro, 41(1):6–7, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Guo:2013:NMC

[GXMZ13] Jing Guo, Liyi Xiao, Zhigang Mao, and Qiang Zhao. Novel mixed codes for multiple-cell upsets mitigation in static RAMs. IEEE Micro, 33(6): 66–74, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Genc:2017:FIT

[GZC^{+}17] Hasan Genc, Yazhou Zu, Ting-Wu Chin, Matthew Halpern, and Vijay Janapa Reddi. Flying IoT: Toward low-power vision in the sky. IEEE Micro, 37(6):40–51, November/December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060040-
abs.html.

Gan:2020:UHS

[GZC^{+}20] Y. Gan, Y. Zhang, D. Cheng, A. Shetty, P. Rathi, N. Katarki, A. Bruno, J. Hu, B. Ritchken, B. Jackson, K. Hu, M. Pancholi, Y. He, B. Clancy, C. Colen, F. Wen, C. Leung, S. Wang, L. Zaruvinsky, M. Espinosa, R. Lin, Z. Liu, J. Padilla, and C. Delimitrou. Unveiling the hardware and software implications of microservices in cloud and edge systems. IEEE Micro, 40(3): 10–19, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hollaar:1996:LRD

[HA96] L. Hollaar and A. Asay. Legal recognition of digitalsignatures. IEEE Micro, 16 (3):44–45, May/June 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Holt:2009:SSM

[HAB^{+}09] Jim Holt, Anant Agarwal, Sven Brehmer, Max Domeika, Patrick Griffin, and Frank Schirrmeister. Software standards for the multicore era. IEEE Micro, 29(3):40– 51, May/June 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Haj-Ali:2018:NAM

[HABHW^{+}18] Ameer Haj-Ali, Rotem Ben-Hur, Nimrod Wald, Ronny Ronen, and Shahar Kvatinsky. Not in name alone: A memristive memory processing unit for real in-memory processing. IEEE Micro, 38(5):13– 21, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/05/mmi2018050013-
                                        abs.html.

Hachiya:2001:JUM

[Hac01] Shouichi Hachiya. Java use in mobile information devices: Introducing JTRON. IEEE Micro, 21(4):16–21, July/August 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2001/

pdf/m4016.pdf;   m4016abs.
htm.

Ho:2013:SPI

[HAC^{+}13] Ron Ho, Philip Amberg, Eric Chang, Pranay Koka, Jon Lexau, Guoliang Li, Frankie Y. Liu, Herb Schwetman, Ivan Shubin, Hiren D. Thacker, Xuezhe Zheng, John E. Cunningham, and Ashok V. Krishnamoorthy. Silicon photonic interconnects for large-scale computer systems. IEEE Micro, 33(1):68– 78, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hall:1991:ACM

[Hal91] Douglas V. Hall. Adapting curriculum materials for different course sequences. IEEE Micro, 11(1):34–37, 82–83, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Halliwell:1993:CDA

[Hal93] Chris Halliwell. Camp development: the art of building a market through standards. IEEE Micro, 13(6): 10–18, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hamblen:2000:RPU

[Ham00] James O. Hamblen. Rapid prototyping using field-programmable logic devices. IEEE Micro, 20(3):29–37, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m3029.
                                        pdf;   http://www.computer.
                                        org/micro/mi2000/m3029abs.
                                        htm.

Hanson:1981:IMM

[Han81] Donald F. Hanson. An improved model for a microcomputer component — the 6520 PIA. IEEE Micro, 1(4):17– 25, October/December 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hannum:1984:HOC

[Han84] D. L. Hannum. HP opts for compatibility. IEEE Micro, 4(1):84–85, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hannum:1985:JMD

[Han85] D. L. Hannum. Juki Model 6300 daisywheel printer. IEEE Micro, 5(3):79–82, May/June 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hannum:1987:GPP

[Han87] D. L. Hannum. Graphics packages for the PC and compatibles — 2 of the best. IEEE Micro, 7(1):78–79, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hansen:1996:MMA

[Han96] Craig Hansen. MicroUnity’s MediaProcessor architecture: Extending general-purpose systems for media processing and communications while reducing initial and life cycle system costs. IEEE Micro, 16 (4):34–41, July/August 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hari:2013:RAR

[HANR13] Siva Kumar Sastry Hari, Sarita V. Adve, Helia Naeimi, and Pradeep Ramachandran. Relyzer: Application resiliency analyzer for transient faults. IEEE Micro, 33(3): 58–66, May/June 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Harrouet:2012:DMM

[Har12] Fabrice Harrouet. Designing a multicore and multiprocessor individual-based simulation engine. IEEE Micro, 32 (1):54–65, January/February 2012. CODEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Harrod:2021:MIA

[Har21] Pete Harrod. Memories from IBM 370 to ARM. IEEE Micro, 41(6):153–154, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hastings:1985:SSC

[Has85]  Chuck Hastings.  Second-

sourcing CPUs — emulation, ethics, and electropolitics. IEEE Micro, 5(3):41– 52, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Haskell:1994:PEP

[Has94] B. Haskell. Portable electronics packaging technologies. IEEE Micro, 14(5):72– 78, September/October 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hauptman:1988:GIA

[Hau88a] G. Hauptman. A good idea has arrived. IEEE Micro, 8(4): 3–??, July/August 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hauptman:1988:PAP

[Hau88b] G. Hauptman. Protection against piracy. IEEE Micro, 8 (3):88, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hauptman:1988:JBW

[Hau88c] G. A. Hauptman. Joining Berne is wise — abolishing copyright isn’t. IEEE Micro, 8(6):3, November/December 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Huang:2011:TAA

[HAWC^{+}11] Wei Huang, Malcolm Allen-Ware, John B. Carter, Mircea R. Stan, Kevin Skadron, and Edmund Cheng. Temperatureaware architecture: Lessons and opportunities. IEEE Micro, 31(3):82–86, May/June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Huh:2004:SIC

[HBCS04] Jaehyuk Huh, Doug Burger, Jichuan Chang, and Gurindar S. Sohi. Speculative incoherent cache protocols. IEEE Micro, 24(6):104–109, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6104.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6104.pdf.

Haond:1999:DMC

[HBd^{+}99] Michel Haond, Marie-Thérèse Basso, Walter deCoster, Eric Gerritsen, Jos Guelen, and Christophe Lair. Developing a 0.18-micron CMOS process. IEEE Micro, 19(5):16– 22, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m5016.
                                        pdf.

Hoe:2010:FAS

[HBE^{+}10] James C. Hoe, Doug Burger, Joel Emer, Derek Chiou, Resit Sendag, and Joshua Yi. The future of architectural simulation. IEEE Micro, 30(3): 8–18, May/June 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Heckathorne:1983:AAT

[HC83a] Craig W. Heckathorne and Dudley S. Childress. Applying anticipatory text selection in a writing aid for people with severe motor impairment. IEEE Micro, 3(3):17–23, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Huntsman:1983:MFP

[HC83b] Clayton Huntsman and Duane Cawthron. The MC68881 floating-point coprocessor. IEEE Micro, 3(6):44–54, November/December 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Habekotte:1984:SDR

[HC84] Ernst Habekotte and Stefan Cserveny. A smart digitalreadout circuit for a capacitive microtransducer. IEEE Micro, 4(5):44–54, September/October 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hazendonk:1999:PMT

[HC99] Teus Hazendonk and Giuseppe Coppola. Preparing for multimedia terminals. IEEE Micro, 19(5):44–51, September/October 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1999/
pdf/m5044.pdf;   http://
www.computer.org/micro/
mi1999/m5044abs.htm.

Hill:2002:MWP

[HC02] Jason L. Hill and David E.

Culler. Mica: a wireless platform for deeply embedded networks. IEEE Micro, 22(6): 12–24, November/December 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m6012.
pdf;   http://www.computer.
org/micro/mi2002/m6012abs.
htm.

Hurwitz:2004:EEP

[HcF04] Justin (Gus) Hurwitz and Wu chun Feng. End-to-end performance of 10-gigabit Ethernet on commodity systems. IEEE Micro, 24(1):10– 22, January/February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1010abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/01/m1010.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/01/
m1010.pdf.

Huang:2003:CBP

[HCP^{+}03] Michael C. Huang, Daniel Chaver, Luis Piñuel, Manuel Prieto, and Francisco Tirado. Customizing the branch predictor to reduce complexity and energy consumption. IEEE Micro, 23(5):12– 25, September/October 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/05/m5012abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/05/m5012.
pdf.

Hill:2016:PVO

[HCP^{+}16] Mark D. Hill, Dave Christie, David Patterson, Joshua J. Yi, Derek Chiou, and Resit Sendag. Proprietary versus open instruction sets. IEEE Micro, 36(4):58–68, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040058.
html.

Hennessy:2003:UAS

[HCPS03] John Hennessy, Daniel Citron, David Patterson, and Guri Sohi. The use and abuse of SPEC: An ISCA panel. IEEE Micro, 23(4):73– 77, July/August 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/04/m4073abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/04/m4073.
pdf.

Harris:2007:TMO

[HCU^{+}07] Tim Harris, Adrián Cristal, Osman S. Unsal, Eduard Ayguade, Fabrizio Gagliardi, Burton Smith, and Mateo Valero. Transactional memory: An overview. IEEE Micro, 27(3):8–29, May/June 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hammond:2004:TCC

[HCW^{+}04] Lance Hammond, Brian D.

Carlstrom, Vicky Wong, Michael Chen, Christos Kozyrakis, and Kunle Olukotun. Transactional coherence and consistency: Simplifying parallel hardware and software. IEEE Micro, 24(6):92–103, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6092.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6092.pdf.

Huzaifa:2022:IOT

[HDG^{+}22] Muhammad Huzaifa, Rishi Desai, Samuel Grayson, Xutao Jiang, Ying Jing, Jae Lee, Fang Lu, Yihan Pang, Joseph Ravichandran, Finn Sinclair, Boyuan Tian, Hengzhi Yuan, Jeffrey Zhang, and Sarita V. Adve. ILLIXR: an open testbed to enable extended reality systems research. IEEE Micro, 42(4):97– 106, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hofstee:1998:DG

[HDM^{+}98] H. Peter Hofstee, Sang H. Dhong, David Meltzer, Kevin J. Nowka, Joel A. Silberman, Jeffrey L. Burns, Stephen D. Posluszny, and Osamu Takahashi. Designing for a gigahertz. IEEE Micro, 18 (3):66–74, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3066.
pdf;   http://www.computer.
org/micro/mi1998/m3066abs.
htm.

Hu:1994:CAA

[HDMT94] Xiaobo Hu, Joseph G. D’Ambrosio, Brian T. Murray, and Dah-Lain Tang. Codesign of architectures for automotive powertrain modules. IEEE Micro, 14(4):17–25, July/August 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hoste:2007:MIW

[HE07] Kenneth Hoste and Lieven Eeckhout. Microarchitectureindependent workload characterization. IEEE Micro, 27(3): 63–72, May/June 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Heath:1984:SER

[Hea84] Walter S. Heath. A system executive for real-time microcomputer programs. IEEE Micro, 4(3):20–32, May/June 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Heath:1987:SDR

[Hea87] Walter S. Heath. Software design for real-time multiprocessor VMEbus systems. IEEE Micro, 7(6):71–80, November/December 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hecht:1983:CIS

[Hec83a]  H. Hecht.  Copyrighting

of IEEE standards challenged. IEEE Micro, 3(6): 4–??, November/December 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hecht:1983:PCF

[Hec83b] H. Hecht. Publication creates de-facto standards without proper review. IEEE Micro, 3(3):57–58, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hennessy:1996:RM

[Hen96] J. Hennessy. RISC microprocessors. IEEE Micro, 16 (6):27, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hennessy:2021:YHM

[Hen21a] John L. Hennessy. The 50 year history of the microprocessor as five technology eras. IEEE Micro, 41(6):20–21, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Henning:2021:HMV

[Hen21b] John L. Henning. How many VAXes fit in the palms of your hands? IEEE Micro, 41(6):140–143, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Henry:2021:MM

[Hen21c]  G. Glenn Henry.  From

mainframes to microprocessors. IEEE Micro, 41(6): 89–96, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Herrell:1993:ACA

[Her93] Dennis J. Herrell. Addressing the challenges of advanced packaging and interconnection. IEEE Micro, 13(2):10– 18, March/April 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Herring:2000:MMS

[Her00]  Chris Herring.  Micropro-

cessors, microcontrollers, and systems in the new millennium. IEEE Micro, 20(6): 45–51, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6045.
pdf;   http://www.computer.
org/micro/mi2000/m6045abs.
htm.

Hartmann:1981:VAS

[HF81] Alfred C. Hartmann and Scott Fehr. A VLSI architecture for software structure: the Intel 8086. IEEE Micro, 1(2): 57–69, April/June 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hunter:1984:INA

[HF84] Colin B. Hunter and Erin Farquhar. Introduction to the NS16000 architecture. IEEE Micro, 4(2):26–47, March/ April 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hardavellas:2010:NOC

[HFFA10] Nikos Hardavellas, Michael Ferdman, Babak Falsafi, and Anastasia Ailamaki. Nearoptimal cache block placement with reactive nonuniform cache architectures. IEEE Micro, 30(1):29, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hardavellas:2011:TDS

[HFFA11] Nikos Hardavellas, Michael Ferdman, Babak Falsafi, and Anastasia Ailamaki. Toward dark silicon in servers. IEEE Micro, 31(4):6–15, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hoefler:2012:TPH

[HGPT12] Torsten Hoefler, Patrick Geoffray, Fabrizio Petrini, and Jesper Larsson Traff. Top picks from Hot Interconnects 2011: Petascale network architectures. IEEE Micro, 32(1): 4–7, January/February 2012. CODEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Huang:2017:ACM

[HGS^{+}17] Yipeng Huang, Ning Guo, Mingoo Seok, Yannis Tsividis, and Simha Sethumadhavan. Analog computing in a modern context: A linear algebra accelerator case study. IEEE Micro, 37(3): 30–38, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030030-
abs.html.

Hsu:1999:IDB

[hHH99] Feng hsiung H. Hsu. IBM’s Deep Blue chess grandmaster chips. IEEE Micro, 19 (2):70–81, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2070.
pdf;   http://www.computer.
org/micro/mi1999/m2070abs.
htm.

Hori:2009:ADV

[HHNK09] Yuichi Hori, Yuya Hanai, Jun Nishimura, and Tadahiro Kuroda. Architecture design of versatile recognition processor for Sensornet applications. IEEE Micro, 29(6): 44–57, November/December 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hammond:2000:SHC

[HHS^{+}00] Lance Hammond, Benedict A. Hubbert, Michael Siu, Manohar K. Prabhu, Michael Chen, and Kunle Olukotun. The Stanford Hydra CMP. IEEE Micro, 20 (2):71–84, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m2071.
                                        pdf;   http://www.computer.
                                        org/micro/mi2000/m2071abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Higgins:1985:TPE

[Hig85] W. T. Higgins. Turning a PC into an engineering workstation — 2 approaches. IEEE Micro, 5(5):80–84, September/October 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hill:1987:CAM

[Hil87] G. A. Hill. Change the approach of microstandards. IEEE Micro, 7(1):4, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hill:2019:RRA

[Hil19] M. D. Hill. Reflections and research advice upon receiving the 2019 Eckert–Mauchly Award. IEEE Micro, 39(5): 119–124, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hinnant:1988:AUB

[Hin88] David F. Hinnant. Accurate Unix benchmarking: art, science, or black magic? IEEE Micro, 8(5):64–75, September/October 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Han:2022:MDT

[HIP^{+}22] Donghyeon Han, Dongseok Im, Gwangtae Park, Youngwoo Kim, Seokchan Song, Juhyoung Lee, and Hoi-Jun Yoo. A mobile DNN training processor with automatic bit precision search and finegrained sparsity exploitation. IEEE Micro, 42(2):16–25, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hariharan:1982:MBP

[HK82] R. Hariharan and V. D.

Kulkarni. A microprocessorbased pulse-height analyzer. IEEE Micro, 2(3):86–91, July/September 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hughes:2010:PEI

[HKC10] Christopher Hughes, Changkyu Kim, and Yen-Kuang Chen. Performance and energy implications of many-core caches for throughput computing. IEEE Micro, 30(6):25–35, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Holland:1985:ESS

[HKM^{+}85] Les Holland, Granino Korn, John Matson, Bob Seader, and Phil Wolfe. Engineering support system software. IEEE Micro, 5(5):17– 21, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ho:2016:AAM

[HKS16] Chen-Han Ho, Sung Jin Kim, and Karthikeyan Sankaralingam. Accelerating the accelerator memory interface with access-execute and dataflow. IEEE Micro, 36(3): 31–41, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030031-
abs.html.

Hasegawa:1995:SHC

[HKY^{+}95] Atsushi Hasegawa, Ikuya Kawasaki, Kouji Yamada, Shinichi Yoshioka, Shumpei Kawasaki, and Prasenjit Biswas. SH3 — high code density, low-power. IEEE Micro, 15(6):11–19, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Huang:1986:OSI

[HL86] V. K. L. Huang and P. M.

Lu. Operating-systems — introduction. IEEE Micro, 6(4): 6–7, July/August 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Horel:1999:UID

[HL99] Tim Horel and Gary Lauterbach. UltraSPARC-III: Designing third-generation 64bit performance. IEEE Micro, 19(3):73–85, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3073.
pdf;   http://www.computer.
org/micro/mi1999/m3073abs.
htm.

Hur:2006:AHB

[HL06] Ibrahim Hur and Calvin Lin.

Adaptive history-based memory schedulers for modern processors. IEEE Micro, 26 (1):22–29, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Haworth:1990:EDS

[HLHR90] Guy Haworth, Steve Leunig, Carsten Hammer, and Mike Reeve. The European Declarative System, database, and languages. IEEE Micro, 10(6):20–23, 83–88, November/December 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hu:2020:VSD

[HLIT20] Y. Hu, M. Lokhandwala, T. I, and H. Tseng. Varifocal storage: Dynamic multiresolution data storage. IEEE Micro, 40(3):47–55, May/June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ham:2021:AGD

[HLS^{+}21] Tae Jun Ham, Yejin Lee, Seong Hoon Seo, U. Gyeong Song, Jae W. Lee, David Bruns-Smith, Brendan Sweeney, Krste Asanovic, Young H. Oh, and Lisa Wu Wills. Accelerating genomic data analytics with composable hardware acceleration framework. IEEE Micro, 41(3):42–49, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hauswald:2016:SIF

[HLZ^{+}16] Johann Hauswald, Michael A. Laurenzano, Yunqi Zhang, Cheng Li, Austin Rovinski, Arjun Khurana, Ronald G.

Dreslinski, Trevor Mudge, Vinicius Petrucci, Lingjia Tang, and Jason Mars. Sirius implications for future warehouse-scale computers. IEEE Micro, 36(3):42– 53, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030042-
abs.html.

Hurson:1993:KTS

[HM93] A. R. Hurson and Patrick M.

Miller. A 16-Kbit Theta — search associative memory. IEEE Micro, 13(2):59– 65, March/April 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hidaka:1990:CDA

[HMAF90] Hideto Hidaka, Yoshio Matsuda, Mikio Asakura, and Kazuyasu Fujishima. The cache DRAM architecture: a DRAM with an on-chip cache memory. IEEE Micro, 10 (2):14–25, March/April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hammarlund:2014:HFG

[HMB^{+}14] Per Hammarlund, Alberto J.

Martinez, Atiq A. Bajwa, David L. Hill, Erik Hallnor, Hong Jiang, Martin Dixon, Michael Derr, Mikal Hunsaker, Rajesh Kumar, Randy B. Osborne, Ravi Rajwar, Ronak Singhal, Reynold D’Sa, Robert Chappell, Shiv Kaushik, Srinivas Chennupaty, Stephan Jourdan, Steve Gunther, Tom Piazza, and Ted Burton. Haswell: The fourth-generation Intel core processor. IEEE Micro, 34(2): 6–20, March/April 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Humble:2021:QCH

[HML^{+}21] Travis S. Humble, Alexander McCaskey, Dmitry I. Lyakh, Meenambika Gowrishankar, Albert Frisch, and Thomas Monz. Quantum computers for high-performance computing. IEEE Micro, 41(5):15– 23, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Huck:2000:IIA

[HMR^{+}00] Jerry Huck, Dale Morris, Jonathan Ross, Allan Knies, Hans Mulder, and Rumi Zahir. Introducing the IA-64 architecture. IEEE Micro, 20(5):12– 23, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5012.
pdf;   http://www.computer.
org/micro/mi2000/m5012abs.
htm.

Hill:2019:SMP

[HMR^{+}19] M. D. Hill, J. Masters, P. Ranganathan, P. Turner, and J. L. Hennessy. On the Spectre and Meltdown processor security vulnerabilities. IEEE Micro, 39(2):9–19, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hayes:1986:MBH

[HMS^{+}86] John P. Hayes, Trevor Mudge, Quentin F. Stout, Stephen Colley, and John Palmer. A microprocessor-based hypercube supercomputer. IEEE Micro, 6(5):6–17, September/October 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Homewood:1987:ITT

[HMSS87] Mark Homewood, David May, David Shepherd, and Roger Shepherd. The IMS T800 transputer. IEEE Micro, 7(5): 10–26, September/October 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hilton:2010:ITA

[HNR10] Andrew Hilton, Santosh Nagarakatte, and Amir Roth. iCFP: Tolerating all-level cache misses in in-order processors. IEEE Micro, 30 (1):12–19, January/February 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hagiwara:1999:SDC

[HO99a] Shiro Hagiwara and Ian Oliver. Sega Dreamcast: Creating a unified entertainment world. IEEE Micro, 19(6): 29–35, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m6029.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m6029abs.
                                        htm.

Hangal:1999:PAV

[HO99b] Sudheendra Hangal and Mike O’Connor. Performance analysis and validation of the picoJava processor. IEEE Micro, 19(3):66–72, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m3066.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m3066abs.
                                        htm.

Hoerbst:1992:MEG

[Hoe92] Egon Hoerbst. Microelectronics in Europe — Guest Editors’ introduction. IEEE Micro, 12(4):8–9, July/August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hoefflinger:1993:GEI

[Hoe93] Bernd Hoefflinger. Guest Editor’s introduction: An electronic copilot in your car?

IEEE Micro, 13(1):7–10, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Haring:2012:IBG

[HOF^{+}12] Ruud A. Haring, Martin Ohmacht, Thomas W. Fox, Michael K. Gschwind, David L. Satterfield, Krishnan Sugavanam, Paul W. Coteus, Philip Heidelberger, Matthias A. Blumrich, Robert W. Wisniewski, Alan Gara, George Liang-Tai Chiu, Peter A. Boyle, Norman H. Chist, and Changhoan Kim. The IBM Blue Gene/Q compute chip. IEEE Micro, 32(2):48– 60, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hasper:1999:AME

[HOHCV99] Albert Hasper, Ed Oosterlaken, Frank Huussen, and Tanja Claasen-Vujcic. Advanced manufacturing equipment: a vertical batch furnace for 300-mm wafer processing. IEEE Micro, 19(5):34– 43, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5034.
pdf;   http://www.computer.
org/micro/mi1999/m5034abs.
htm.

Holden:1998:MNG

[Hol98] Happy Holden. Microvias: The next generation of substrates and packages. IEEE Micro, 18(4):10–16, July/ August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1998/
                                        pdf/m4010.pdf;   http://
                                        www.computer.org/micro/
                                        mi1998/m4010abs.htm.

Hootman:1989:NNP

[Hoo89a] J. Hootman. Neural networks — problem-solving tools. IEEE Micro, 9(6): 4–??, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1989:PPF

[Hoo89b] J. Hootman. Past, present, future. IEEE Micro, 9(1):1–2, January/February 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1989:RTC

[Hoo89c] J. Hootman. RISC tutorial comments. IEEE Micro, 9 (3):4, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1990:FEI

[Hoo90a] J. Hootman. The Far-East issue for 1990. IEEE Micro, 10 (2):3, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1990:HC

[Hoo90b] J. Hootman. Hot Chips. 2.

IEEE Micro, 10(3):3–4, May/ June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1990:HMS

[Hoo90c] J. Hootman. How Micro survives. IEEE Micro, 10(5):3– 4, September/October 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1990:LA

[Hoo90d] J. Hootman. Letters and articles. IEEE Micro, 10(4):2– 3, July/August 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1991:RE

[Hoo91] J. Hootman. The role of the EIC. IEEE Micro, 11(1):9–??, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Horst:1995:TRS

[Hor95]  R. W. Horst.  Tnet —

a reliable system area network. IEEE Micro, 15(1):37– 45, January/February 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Heath:1981:HWU

[HP81] J. Robert Heath and Shailesh M. Patel. How to write a universal cross-assembler. IEEE Micro, 1(3):45–66, July/ September 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hanson:1985:ESS

[HP85] Donald F. Hanson and Peggy Cook Power. Electronic scanners with speech output — a communication system for the physically handicapped and mentally retarded. IEEE Micro, 5(2):20–52, March/April 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ha:2023:DCS

[HRC^{+}23] Minho Ha, Junhee Ryu, Jungmin Choi, Kwangjin Ko, Sunwoong Kim, Sungwoo Hyun, Donguk Moon, Byungil Koh, Hokyoon Lee, Myoungseo Kim, Hoshik Kim, and Kyoung Park. Dynamic capacity service for improving CXL pooled memory efficiency. IEEE Micro, 43(2):39– 47, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Han:2024:LPA

[HRK^{+}24] Donghyeon Han, Junha Ryu, Sangyeob Kim, Sangjin Kim, Jongjun Park, and Hoi-Jun Yoo. A low-power artificialintelligence-based 3-D rendering processor with hybrid deep neural network computing. IEEE Micro, 44(1):17–27, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Huang:2011:SDC

[HRSS11] Wei Huang, Karthick Rajamani, Mircea R. Stan, and Kevin Skadron. Scaling with design constraints: Predicting the future of big chips. IEEE Micro, 31(4):16–29, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Harrison:1985:AMC

[HS85] Malcolm C. Harrison and Owen Smith. The Ampos multiprocessor — a computer system for laboratory use. IEEE Micro, 5(1):22–30, January/February 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Herrmann:1992:DAP

[HS92] Frederick P. Herrmann and Charles G. Sodini. A dynamic associative processor for machine vision applications. IEEE Micro, 12(3):

31–41, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hunt:1999:VFM

[HS99] Warren A. Hunt, Jr. and Jun Sawada. Verifying the FM9801 microarchitecture. IEEE Micro, 19 (3):47–55, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3047.
pdf;   http://www.computer.
org/micro/mi1999/m3047abs.
htm.

Hsiao:1991:PSM

[Hsi91] David K. Hsiao. A parallel, scalable, microprocessorbased database computer for performance gains and capacity growth. IEEE Micro, 11(6):44–60, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Heo:2023:IBI

[HSN^{+}23] Tejun Heo, Dan Schatzberg, Andrew Newell, Song Liu, Saravanan Dhakshinamurthy, Iyswarya Narayanan, Josef Bacik, Chris Mason, Chunqiang Tang, and Dimitrios Skarlatos. IOCost: Block input output control for containers in datacenters. IEEE Micro, 43(4):80–87, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hanindhito:2021:IWF

[HSNJ21] Bagus Hanindhito, Karthik Swaminathan, Vijaykrishnan Narayanan, and Lizy Kurian John. Intel wins in four decades, but AMD catches up. IEEE Micro, 41(6):168–171, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Haq:2001:JSS

[HSP^{+}01] Ejaz Haq, Jim Slager, John Pecoraro, John D. Johnson, Mark Santoro, Lee Tavrow, Scott Wakefield, and David Weisner. JAZiO signalswitching technology: a lowcost digital I/O for high-speed applications. IEEE Micro, 21 (1):72–81, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1072.
pdf;   http://www.computer.
org/micro/mi2001/m1072abs.
htm.

Husemann:2018:OSA

[HSR18] Ronaldo Husemann, Altamiro Amadeu Susin, and Valter Roesler. Optimized solution to accelerate in hardware an intra H.264/ SVC video encoder. IEEE Micro, 38(6):8–17, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/06/08536428-abs.
                                        html.

Hsu:1994:DTM

[Hsu94] Peter Yan-Tek Hsu. Designing the TFP microprocessor. IEEE Micro, 14 (2):23–33, March/April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL ftp://

                                        ftp.sgi.com/sgi/doc/TFP/
                                        R8000_Micro_Paper.ps.Z.

Huang:1989:AWD

[HSW^{+}89] Victor K. L. Huang, James W. Seery, William S. Wu, Saul K. Altabet, Michael J. Killian, Simeon Aymeloglu, Thaddeus J. Gabara, Aaron L. Fisher, Inseok S. Hwang, and David W. Thompson. The AT&T WE32200 design challenge. IEEE Micro, 9(2):14– 25, March/April 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hirt:1998:APF

[HSW98] Etienne Hirt, Michael Scheffler, and Jean-Pierre P. Wyss. Area I/O’s potential for future processor systems. IEEE Micro, 18(4):42–49, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4042.
pdf;   http://www.computer.
org/micro/mi1998/m4042abs.
htm.

Hu:2018:SH

[HSX18] Xing Hu, Dylan Stow, and Yuan Xie. Die stacking is happening. IEEE Micro, 38(1):22–28, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010022-
abs.html.

Huang:1989:HPM

[Hua89]  V. K. L. Huang.  High-

performance microprocessors — the RISC dilemma. IEEE Micro, 9(4):13–14, July/ August 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Humphry:1984:FTM

[Hum84] J. A. Humphry. Fault tolerance and micros in the real world. IEEE Micro, 4(6):3– 5, November/December 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hunter:1987:ICA

[Hun87] C. B. Hunter. Introduction to the Clipper architecture. IEEE Micro, 7(4):6– 26, July/August 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hung:1995:DDF

[Hun95]  Donald L. Hung.  Ded-

icated digital fuzzy hardware. IEEE Micro, 15(4):31– 39, July/August 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hurt:1997:MVO

[Hur97] J. R. Hurt. Micro view: OMIs collaborative approach. IEEE Micro, 17(5):80, 79, September/October 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1997/
                                        pdf/m5080.pdf.

Hurt:1998:CFS

[Hur98] J. R. Hurt. Cyrix faces slot-1 challenge. IEEE Micro, 18 (1):88–??, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hodjat:2004:HTP

[HV04] Alireza Hodjat and Ingrid Verbauwhede. High-throughput programmable cryptocoprocessor. IEEE Micro, 24(3): 34–45, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/03/m3034.htm;   http:

//csdl.computer.org/dl/
mags/mi/2004/03/m3034.pdf.

Hoskote:2007:GMI

[HVS^{+}07] Yatin Hoskote, Sriram Vangal, Arvind Singh, Nitin Borkar, and Shekhar Borkar. A 5-GHz mesh interconnect for a teraflops processor. IEEE Micro, 27(5):51–61, September/October 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hill:1991:GEI

[HW91] Mark D. Hill and David A.

Wood. Guest Editors’ introduction: Hot chips II symposium. IEEE Micro, 11(3): 8–9, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hu:2009:GSM

[HWG^{+}09] Weiwu Hu, Jian Wang, Xiang Gao, Yunji Chen, Qi Liu, and Guojie Li. Godson-3: a scalable multicore RISC processor with x86 emulation. IEEE Micro, 29(2):17–29, March/April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hu:1998:NHD

[HY98] Yiming M. Hu and Qing Yang.

A new hierarchical disk architecture. IEEE Micro, 18(6): 64–76, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6064.
pdf;   http://www.computer.
org/micro/mi1998/m6064abs.
htm.

Hyde:2000:TDC

[Hyd00] Daniel C. Hyde. Teaching design in a computer architecture course. IEEE Micro, 20(3):23–28, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3023.
pdf;   http://www.computer.
org/micro/mi2000/m3023abs.
htm.

Hatano:1990:BMP

[HYM^{+}90] Yuji Hatano, Shinichiro Yano, Hiroyuki Mori, Hiroji Yamada, Mikio Hirano, and Ushio Kawabe. A 4-bit, 250-MIPS processor using Josephson-technology. IEEE Micro, 10(2):40–55, March/ April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Horowitz:1998:HSE

[HYS98] M. Horowitz, C. K. K. Yang, and S. Sidiropoulos. High-speed electrical signaling — overview and limitations. IEEE Micro, 18(1):12– 24, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ikeda:2009:GEI

[IA09] Makoto Ikeda and Fumio Arakawa. Guest Editors’ introduction: Cool Chips. IEEE Micro, 29(6):5–6, November/December 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ikeda:2011:CC

[IA11] Makoto Ikeda and Fumio Arakawa. Cool chips. IEEE Micro, 31(6):4–5, November/December 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ikeda:2013:CCG

[IA13] Makoto Ikeda and Fumio Arakawa. Cool chips [guest editors’ introduction]. IEEE Micro, 33(6):4–5, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Ikeda:2022:SIC

[IA22] Makoto Ikeda and Fumio Arakawa. Special issue on cool chips. IEEE Micro, 42(2):6– 7, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Iacobovici:1988:PIH

[Iac88] Sorin Iacobovici. A pipelined interface for high floatingpoint performance with precise exceptions. IEEE Micro, 8(3):77–87, May/June 1988.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ibbett:2000:HDS

[Ibb00] Roland N. Ibbett. HASE DLX simulation model. IEEE Micro, 20(3):57–65, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3057.
pdf;   http://www.computer.
org/micro/mi2000/m3057abs.
htm.

Isci:2005:LTW

[IBM05] Canturk Isci, Alper Buyuktosunoglu, and Margaret Martonosi. Long-term workload phases: Duration predictions and applications to DVFS. IEEE Micro, 25(5):39– 51, September/October 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ishida:2021:SCN

[IBN^{+}21] Koki Ishida, Ilkwon Byun, Ikki Nagaoka, Kosuke Fukumitsu, Masamitsu Tanaka, Satoshi Kawakami, Teruo Tanimoto, Takatsugu Ono, Jangwoo Kim, and Koji Inoue. Superconductor computing for neural networks. IEEE Micro, 41(3):19–26, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Iyer:2021:AMC

[IDI^{+}21] Ravi Iyer, Vivek De, Ramesh Illikkal, David Koufaty, Bhushan Chitlur, Andrew Herdrich, Muhammad Khellah, Fatih Hamzaoglu, and Eric Karl. Advances in microprocessor cache architectures over the last 25 years. IEEE Micro, 41(6):78–88, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Iturbe:2014:RBA

[IEB^{+}14] Xabier Iturbe, Ali Ebrahim, Khaled Benkrid, Chuan Hong, Tughrul Arslan, Jon Perez, Didier Keymeulen, and Marco D. Santambrogio. R3TOSbased autonomous faulttolerant systems. IEEE Micro, 34(6):20–30, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060020-
abs.html.

Ionica:2015:MMA

[IG15] Mircea Horea Ionica and David Gregg. The Movidius Myriad Architecture’s potential for scientific computing. IEEE Micro, 35(1):6–14, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/

                                        mi/2015/01/mmi2015010006-
                                        abs.html.

Ishibashi:1999:SBT

[IGH^{+}99] Kenichi Ishibashi, Tsutomu Goto, Takehisa Hayashi, Tetsuhiko Okada, Akira Yamagiwa, Masabumi Shibata, Kazuhiro Akimoto, Naoki Hamanaka, Toshiro Takahashi, Akio Koyama, and Tatsuhiro Aida. Simultaneous bidirectional transceiver logic. IEEE Micro, 19(1):14– 19, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m1014.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m1014abs.
                                        htm.

Ilitzky:2007:ASC

[IHCE07] David Arditti Ilitzky, Jeffrey D. Hoffman, Anthony Chun, and Brando Perez Esparza. Architecture of the Scalable Communications Core’s Network on Chip. IEEE Micro, 27(5):62– 74, September/October 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Isaak:1998:MVP

[IJ98] Jim Isaak and Lowell Johnson. Micro view: POSIX/ UNIX standards — foundation for 21st-century growth. IEEE Micro, 18(4):88–??, July/August 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m4088.pdf.

Indiveri:1996:SIA

[IKK96] Giacomo Indiveri, Jorg Kramer, and Christof Koch. System implementations of analog VLSI velocity sensors — developing architectures that incorporate new, compact velocity sensors for applications in space exploration, driver assistance, and autonomous navigation. IEEE Micro, 16(5):40– 49, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ikeda:1999:SMV

[IKN^{+}99] Mitsuo Ikeda, Toshio Kondo, Koyo Nitta, Kazuhito Suguri, Takeshi Yoshitome, Toshihiro Minami, Hiroe Iwasaki, Katsuyuki Ochiai, Jiro Naganuma, Makoto Endo, Yutaka Tashiro, Hiroshi Watanabe, Naoki Kobayashi, Tsuneo Okubo, Takeshi Ogura, and Ryota Kasai. SuperEnc: MPEG-2 video encoder chip. IEEE Micro, 19 (4):56–65, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4056.
pdf;   http://www.computer.

                                        org/micro/mi1999/m4056abs.
                                        htm.

Inayoshi:1988:RG

[IKNS88] Hideo Inayoshi, Ikuya Kawasaki, Tadahiko Nishimukai, and Ken Sakamura. Realization of Gmicro/200. IEEE Micro, 8 (2):12–21, March/April 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Iacobovici:1987:VSP

[IN87] S. Iacobovici and C. C. Ng. VLSI and system performance modeling. IEEE Micro, 7 (4):59–72, July/August 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ing:1999:ITM

[Ing99] Davin S. L. Ing. Innovations in a technology museum. IEEE Micro, 19(6): 44–52, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m6044.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m6044abs.
                                        htm.

Iyer:2005:RAN

[INKM05] Ravishankar K. Iyer, Nithin M. Nakka, Zbigniew T. Kalbarczyk, and Subhasish Mitra. Recent advances and new avenues in hardware-level reliability support. IEEE Micro, 25(6):18–29, November/December 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Iyer:2016:VIA

[IO16] Ravi Iyer and Emre Ozer. Visual IoT: Architectural challenges and opportunities; toward a self-learning and energy-neutral IoT. IEEE Micro, 36(6):45–49, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060045-
abs.html.

Ipek:2019:MAD

[Ipe19] E. Ipek. Memristive accelerators for dense and sparse linear algebra: From machine learning to high-performance scientific computing. IEEE Micro, 39(1):58–61, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Im:2023:MDO

[IPL^{+}23] Dongseok Im, Gwangtae Park, Zhiyong Li, Junha Ryu, Sanghoon Kang, Donghyeon Han, Jinsu Lee, Wonhoon Park, Hankyul Kwon, and Hoi-Jun Yoo. A mobile 3-D object recognition processor with deep-learning-based monocular depth estimation. IEEE Micro, 43(3):

74–82, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Isaak:1983:WDB

[Isa83] J. Isaak. What defines a 32bit microprocessor? IEEE Micro, 3(6):3–6, November/December 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Inoue:1991:RRD

[ISH^{+}91] Ushio Inoue, Tetsuji Satoh, Haruo Hayami, Hideaki Takeda, Toshio Nakamura, and Hideki Fukuoka. Rinda: a relational database processor with hardware specialized for searching and sorting. IEEE Micro, 11(6):61–70, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Iyer:2011:CHS

[IST^{+}11] Ravi Iyer, Sadagopan Srinivasan, Omesh Tickoo, Zhen Fang, Ramesh Illikkal, Steven Zhang, Vineet Chadha, Paul M. Stillwell, Jr., and Seung Eun Lee. CogniServe: Heterogeneous server architecture for large-scale recognition. IEEE Micro, 31(3):20–31, May/ June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Iyer:2015:HCG

[IT15] Ravi Iyer and Dean Tullsen.

Heterogeneous computing [Guest Editors’ introduction]. IEEE Micro, 35(4):4–5, July/August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/04/mmi2015040004-
abs.html.

Ingenbleek:1989:IFD

[IWM89] Bernd Ingenbleek, Klaus Woelcken, and Claudia Matthaus. Information flow in digital metal-oxide semiconductor circuits. IEEE Micro, 9(2):69– 79, March/April 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jouppi:1996:GEI

[JA96] Norman P. Jouppi and Hasan S.

Alkhatib. Guest Editors’ introduction: Hot chips and the microprocessor. IEEE Micro, 16(2):6–7, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jacob:2003:CSD

[Jac03] Bruce Jacob. A case for studying DRAM issues at the system level. IEEE Micro, 23 (4):44–56, July/August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/04/m4044abs.htm;

                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/04/m4044.
                                        pdf.

Jaeger:1982:TADa

[Jae82a] R. C. Jaeger. Tutorial: analog data acquisition technology. I. digital-to-analog conversion. IEEE Micro, 2(2): 20–37, April/June 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jaeger:1982:TADb

[Jae82b] R. C. Jaeger. Tutorial: analog data acquisition technology. II. analog-to-digital conversion. IEEE Micro, 2(3): 46–57, July/September 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jaeger:1982:TADc

[Jae82c] R. C. Jaeger. Tutorial: analog data acquisition technology. III. sample-and-holds, instrumentation amplifiers, and analog multiplexers. IEEE Micro, 2(4):20–35, October/December 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jaeger:1983:TAD

[Jae83]  R. C. Jaeger.  Tutorial:

analog data acquisition technology. IV. system design, analysis, and performance. IEEE Micro, 3(1):52–61, January/February 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jaggar:1997:GEI

[Jag97] Dave Jaggar. Guest Editor’s introduction: ARM architecture and systems: ARM licenses three basic processor cores for high-performance, low-cost, and low-power implementations. IEEE Micro, 17(4):9–11, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4009.
pdf.

James:1990:MBE

[Jam90] David V. James. Multiplexed buses — the endian wars continue. IEEE Micro, 10(3): 9–21, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jain:2022:ODD

[JAS^{+}22] Arpan Jain, Nawras Alnaasan, Aamir Shafi, Hari Subramoni, and Dhabaleswar K. Panda. Optimizing distributed DNN training using CPUs and BlueField-2 DPUs. IEEE Micro, 42(2):53–60, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jouppi:1994:DPT

[JBF94] Norman P. Jouppi, Patrick Boyle, and John S. Fitch. Designing, packaging, and testing a 300-MHz, 115 W ECL microprocessor. IEEE Micro, 14(2):50–58, March/April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jaramillo-Botero:1995:PHS

[JBM95] A. Jaramillo-Botero and Y. Miyake. Parallel, highspeed PC fuzzy control. IEEE Micro, 15(6):63, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jackson:1984:PIM

[JC84] Don L. Jackson and Jack Cowan. The proposed IEEE 855 microprocessor operating systems interface standard. IEEE Micro, 4(4):63– 71, July/August 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jin:2008:EBS

[JC08a] Zhanpeng Jin and Allen C. Cheng. Evolutionary benchmark subsetting. IEEE Micro, 28(6):20–36, November/December 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jin:2008:ICP

[JC08b] Zhanpeng Jin and Allen C.

Cheng. ImplantBench: Characterizing and projecting representative benchmarks for emerging bioimplantable computing. IEEE Micro, 28 (4):71–91, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jeffries:1984:PSP

[Jef84] R. Jeffries. PC stat packages (reprinted from the Jeffries Report, Vol 3, 1984). IEEE Micro, 4(6):81–83, November/December 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jelemensky:1989:MM

[JGB^{+}89] Joe Jelemensky, Vernon Goler, Brad Burgess, James Eifert, and Gary Miller. The MC68332 microcontroller. IEEE Micro, 9(4):31–50, July/August 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jimenez:2011:EAA

[JGC^{+}11] Victor Jimenez, Roberto Gioiosa, Francisco J. Cazorla, Mateo Valero, Eren Kursun, Canturk Isci, Alper Buyuktosunoglu, and Pradip Bose. Energy-aware accounting and billing in large-scale computing facilities. IEEE Micro, 31(3):60–71, May/June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

James:1998:SHP

[JGF98] David V. James, David B. Gustavson, and Balint Fleischer. SerialExpress: a highperformance workstation interconnect. IEEE Micro, 18 (3):54–65, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3054.
pdf;   http://www.computer.
org/micro/mi1998/m3054abs.
htm.

Joseph:2020:PAN

[JGM^{+}20] V. Joseph, G. L. Gopalakrishnan, S. Muralidharan, M. Garland, and A. Garg. A programmable approach to neural network compression. IEEE Micro, 40(5):17– 25, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jimenez:2021:TPC

[Jim21] Daniel A. Jimenez. Top picks from the 2020 Computer Architecture Conferences. IEEE Micro, 41(3):6–9, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:2011:RCS

[JJK^{+}11] Daniel R. Johnson, Matthew R.

Johnson, John H. Kelm, William Tuohy, Steven S. Lumetta, and Sanjay J. Patel. Rigel: a 1,024-core single-chip accelerator architecture. IEEE Micro, 31 (4):30–41, July/August 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jia:1996:RFT

[JKN96] Weijia J. Jia, Jorg Kaiser, and Edgar Nett. RMP: Fault-tolerant group communication: Rethinking the token ring approach to decrease overhead and increase reliability. IEEE Micro, 16(2):59– 67, March/April 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jagadish:1989:ESI

[JKP89] N. Jagadish, J. Mohan Kumar, and L. M. Patnaik. An efficient scheme for interprocessor communication using dual-ported RAMs. IEEE Micro, 9(5):10–19, September/October 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jenkins:1987:ASC

[JL87] Robert E. Jenkins and D. Gilbert Lee, Jr. An applicationspecific coprocessor for highspeed cellular logic operations. IEEE Micro, 7(6): 63–70, November/December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jerger:2011:SVL

[JL11] Natalie Enright Jerger and Mikko Lipasti. Systems for very large-scale computing. IEEE Micro, 31(3):4–7, May/ June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jain:2019:PAE

[JLG19] A. Jain, S. Lloyd, and M. Gokhale. Performance assessment of emerging memories through FPGA emulation. IEEE Micro, 39(1):8–16, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Julien:2003:PCM

[JLSM03] Nathalie Julien, Johann Laurent, Eric Senn, and Eric Martin. Power consumption modeling and characterization of the TI C6201. IEEE Micro, 23(5):40–49, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/05/m5040abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/05/m5040.
pdf.

Jia:2020:GSA

[JLWL20] L. Jia, L. Lu, X. Wei, and Y. Liang. Generating systolic array accelerators with reusable blocks. IEEE Micro, 40(4):85–92, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jacob:1998:VMC

[JM98] Bruce Jacob and Trevor Mudge. Virtual memory in contemporary microprocessors. IEEE Micro, 18 (4):60–75, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4060.
pdf;   http://www.computer.
org/micro/mi1998/m4060abs.
htm.

Jiang:2011:CID

[JMZ^{+}11] Xiaowei Jiang, Niti Madan, Li Zhao, Mike Upton, Ravi Iyer, Srihari Makineni, Donald Newell, Yan Solihin, and Rajeev Balasubramonian. CHOP: Integrating DRAM caches for CMP server platforms. IEEE Micro, 31(1):99– 108, January/February 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2021:MTC

[JN21a] Lizy Kurian John and Vijaykrishnan Narayanan. Microprocessor at 50: A time to celebrate and energize for the future. IEEE Micro, 41(6): 10–12, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2021:MIL

[JN21b] Lizy Kurian John and Vijaykrishnan Narayanan. Microprocessor at 50: Industry leaders speak. IEEE Micro, 41(6):13–15, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:1984:FTM

[Joh84] Barry W. Johnson. Faulttolerant microprocessor-based systems. IEEE Micro, 4(6): 6–21, November/December 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:1986:M

[Joh86] B. W. Johnson. Multiprocessing. IEEE Micro, 6(5): 5, September/October 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:1987:SCD

[Joh87] M. Johnson. System considerations in the design of the AM29000. IEEE Micro, 7 (4):28–41, July/August 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:1989:FPD

[Joh89] Matthew Johnson. A fixedpoint DSP for graphics engines. IEEE Micro, 9(4):63– 77, July/August 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:1990:OPC

[Joh90a] M. J. Johnson. One person’s cup of tea is — reply. IEEE Micro, 10(1):93, January/February 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:1990:HCS

[Joh90b] Stephen C. Johnson. Hot chips and soggy software: RISC success springs partially from good system design. take note and eliminate the software bottleneck from your new design. IEEE Micro, 10 (1):23–26, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2019:EHC

[Joh19a] L. John. Emerging hot chips and systems. IEEE Micro, 39(2):4–5, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2019:EIC

[Joh19b] L. John. To the era of intelligent chips and systems.

IEEE Micro, 39(1):4–5, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2019:DCC

[Joh19c] L. K. John. 3-D chips! Chips are getting denser and taller than ever!! IEEE Micro, 39 (6):4–5, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2019:MLA

[Joh19d] L. K. John. Machine learning accelerators and more. IEEE Micro, 39(5):4–5, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2019:SA

[Joh19e] L. K. John. Secure architectures. IEEE Micro, 39(4):4– 5, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2019:TP

[Joh19f] L. K. John. Top picks. IEEE Micro, 39(3):4–5, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2020:AHD

[Joh20a] L. K. John. Agile hardware design. IEEE Micro, 40(4):

4–5, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2020:CCC

[Joh20b] L. K. John. Connectivity!

Connectivity! Connectivity! May you be more connected than ever!! IEEE Micro, 40(1):4–5, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2020:DMC

[Joh20c] L. K. John. Did ML chips heat up the chip design arena? IEEE Micro, 40(2): 4–5, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2020:ETP

[Joh20d] L. K. John. Enjoy these top picks, while you work from home! IEEE Micro, 40(3): 4–5, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2021:CMN

[Joh21a] L. K. John. Connectivity more needed than ever before. IEEE Micro, 41(1):4–5, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2021:ML

[Joh21b] Lizy Kurian John. From the memory lane! IEEE Micro, 41(6):144–147, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2021:MLB

[Joh21c]  Lizy Kurian John.  Mi-

croprocessor at 50: Looking back and looking forward. IEEE Micro, 41(6):5– 9, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2022:AIE

[Joh22a] Lizy Kurian John. Artificial intelligence at the edge: Designs and architectures for pervasive intelligence. IEEE Micro, 42(6):4–5, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2022:ACW

[Joh22b] Lizy Kurian John. Automatic compilation will be key for success of the accelerator revolution! IEEE Micro, 42(5):4– 5, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2022:HCM

[Joh22c] Lizy Kurian John. Hot Chips 33 and more! IEEE Micro, 42 (3):4–5, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2022:SAS

[Joh22d] Lizy Kurian John. Smart agriculture and smart memories. IEEE Micro, 42(1):4–6, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2022:SIC

[Joh22e] Lizy Kurian John. Special issue on cool chips and hot interconnects. IEEE Micro, 42(2):4–5, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2022:TPC

[Joh22f] Lizy Kurian John. Top picks from 2021 computer architecture conferences! IEEE Micro, 42(4):4–5, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2023:ESI

[Joh23a] Lizy Kurian John. Emerging system interconnects enabling more opportunities than ever! IEEE Micro, 43(2): 4–5, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2023:ESC

[Joh23b] Lizy Kurian John. Environmentally sustainable computing. IEEE Micro, 43(1):4–6, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2023:HSP

[Joh23c] Lizy Kurian John. Hardware security and privacy: Threats and opportunities. IEEE Micro, 43(5):4–5, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2023:HCM

[Joh23d] Lizy Kurian John. Hot Chips 34 and more! IEEE Micro, 43 (3):4–6, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2023:TNM

[Joh23e] Lizy Kurian John. TinyML but by no means a tiny feat! IEEE Micro, 43(6):4– 6, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2023:TPC

[Joh23f] Lizy Kurian John. Top picks from computer architecture conferences! IEEE Micro, 43(4):4–5, July/August 2023.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Joshi:1986:HPN

[Jos86]  Sunil P. Joshi.  High-

performance networks — a focus on the fiber distributed data interface (FDDI) standard. IEEE Micro, 6(3):8– 14, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jouppi:1992:HCI

[Jou92] Norman P. Jouppi. Hot Chips-III — introduction. IEEE Micro, 12(2):8–9, March/April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jackson:2017:BMS

[JP17] Scott M. Jackson and JoAnn M.

Paul. Building maze solutions with computational dreaming. IEEE Micro, 37(4):64– 71, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040064-
abs.html.

Jiang:2020:PPF

[JPOB20] S. Jiang, P. Pan, Y. Ou, and C. Batten. PyMTL3: a Python framework for opensource hardware modeling, generation, simulation, and verification. IEEE Micro, 40 (4):58–66, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jaleel:2017:TPC

[JQ17] Aamer Jaleel and Moinuddin Qureshi. Top picks from the 2016 Computer Architecture Conferences. IEEE Micro, 37(3):6–11, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030006.
html.

Jackson:1986:PAN

[JRHM86] William A. Jackson, Paul C. Rogers, Robert J. Hearn, and Jeffrey S. Mattiace. Performance and availability in a network file server. IEEE Micro, 6(4):18–34, July/August 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jerger:2018:AC

[JS18a] Natalie Enright Jerger and Joshua San Miguel. Approximate computing. IEEE Micro, 38(4):8–10, July/August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040008.
html.

John:2018:MBC

[JS18b] Lizy K. John and Earl E. Swartzlander. Memristorbased computing. IEEE Micro, 38(5):5–6, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050005.
html.

Jeffrey:2016:UOP

[JSY^{+}16] Mark C. Jeffrey, Suvinay Subramanian, Cong Yan, Joel Emer, and Daniel Sanchez. Unlocking ordered parallelism with the swarm architecture. IEEE Micro, 36(3):105– 117, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030105-
abs.html.

Jentzsch:2022:RMF

[JUP^{+}22] Felix Jentzsch, Yaman Umuroglu, Alessandro Pappalardo, Michaela Blott, and Marco Platzner. RadioML meets FINN: Enabling future RF applications with FPGA streaming architectures. IEEE Micro, 42(6):125–133, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jouppi:1999:GEI

[JW99] Norman P. Jouppi and John Wawrzynek. Guest Editors’ introduction: Real products, real technology. IEEE Micro, 19(2):10–11, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m2010.
                                        pdf.

Jacobi:2020:HIZ

[JW20] C. Jacobi and C. Webb. History of IBM Z mainframe processors. IEEE Micro, 40(6): 50–58, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jagasivamani:2019:AMI

[JWS^{+}19] M. Jagasivamani, C. Walden, D. Singh, L. Kang, S. Li, M. Asnaashari, S. Dubois, B. Jacob, and D. Yeung. Analyzing the monolithic integration of a ReRAM-based main memory into a CPU’s die. IEEE Micro, 39(6): 64–72, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jouppi:2018:MEF

[JYPP18] Norman Jouppi, Cliff Young, Nishant Patil, and David Patterson. Motivation for and evaluation of the first tensor processing unit. IEEE Micro, 38(3):10–19, May/June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

mi/2018/03/mmi2018030010-
abs.html.

Kandel:1995:FHC

[KAC^{+}95] Abraham Kandel, Giuseppe Ascia, Vincenzo Catania, Biagio Giacalone, Marco Russo, Lorenzo Vita, Andres Jaramillo-Botero, Yoichi Miyake, Hua Harry Li, Nowell Godfrey, Yuandong Ji, Shuwei Guo, Liliane Peters, Krishna Rao Valavala, Mahmoud A. Manzoul, Antonio Ruiz, Julio Gutierrez, and J. A. Felipe Fernandez. Fuzzy hardware challenges. IEEE Micro, 15(6):61–67, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1990:AJ

[Kah90a] D. K. Kahaner. Assignment Japan. IEEE Micro, 10(4): 4–6, July/August 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1990:QI

[Kah90b]  D. K. Kahaner.  Quality

improvement. IEEE Micro, 10(6):48–51, November/December 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1990:SRP

[Kah90c] David K. Kahaner. Software report — the Pax parallel computer. IEEE Micro, 10(5):5–6, 91–93, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1991:GP

[Kah91a] D. K. Kahaner. 6th generation project. IEEE Micro, 11 (2):11–??, March/April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1991:CGT

[Kah91b] D. K. Kahaner. Computer growth in Taiwan. IEEE Micro, 11(3):39–41, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1991:OCA

[Kah91c]  D. K. Kahaner.  Optical

computing activities. IEEE Micro, 11(1):53–56, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1991:SRG

[Kah91d] D. K. Kahaner. Software report: a glimpse of the future. IEEE Micro, 11(5):35– ??, September/October 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1991:SRF

[Kah91e] D. K. Kahaner. Software report: Fuzzy theory, applications. IEEE Micro, 11(4):8– 11, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1992:IT

[Kah92a] D. K. Kahaner. ISA in Taiwan. IEEE Micro, 12(3):69– 71, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1992:IPC

[Kah92b]  D. K. Kahaner.  ISDN

progress, civil vs military technology. IEEE Micro, 12(5):7– ??, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1992:MNC

[Kah92c] D. K. Kahaner. Melco neural chip — holography research. IEEE Micro, 12(4):85– 87, July/August 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1992:RDJ

[Kah92d] D. K. Kahaner. Researchand-development in Japan. IEEE Micro, 12(1):7–??, January/February 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1992:TD

[Kah92e] D. K. Kahaner. Transputers and databases. IEEE Micro, 12(6):88–89, November/December 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1992:SRM

[Kah92f] David Kahaner. Special report — MITI’s real world computing program. IEEE Micro, 12 (4):70–80, July/August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:CJN

[Kah93a] D. K. Kahaner. Cooperation: Japan’s new watchword? (software industry). IEEE Micro, 13(5):90–92, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:MS

[Kah93b] D. K. Kahaner. Malaysia and Singapore. IEEE Micro, 13 (1):75–77, January/February 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:MHT

[Kah93c] D. K. Kahaner. Micromachines — high-tech odds and ends. IEEE Micro, 13(2):79– 82, March/April 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:SRC

[Kah93d] D. K. Kahaner. Software report: completely automated assembly. IEEE Micro, 13(3): 88–92, May/June 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:HRJ

[Kah93e] David K. Kahaner. HDTV research in Japan. IEEE Micro, 13(5):49–53, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:SRI

[Kah93f] David K. Kahaner. Special report: India and China hurdle computing obstacles. IEEE Micro, 13(4):44–48, July/August 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:SRS

[Kah93g] David K. Kahaner. Special report: Supercomputing — the view from Japan. IEEE Micro, 13(1):67–70, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:SRV

[Kah93h] David K. Kahaner. Special report: Virtual reality in Japan. IEEE Micro, 13(2):66– 73, March/April 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahrs:1993:SND

[Kah93i] Mark Kahrs. Short note: Dream chip #1: Timed priority queue. IEEE Micro, 13 (4):49–51, July/August 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaiser:1988:MCS

[Kai88] Jorg Kaiser. MUTABOR, a coprocessor supporting memory management in an objectoriented architecture. IEEE Micro, 8(5):30–46, September/October 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Koc:1996:ACM

[KAK96]  Çetin Kaya Koç, Tolga Acar,

and Burton S. Kaliski, Jr. Analyzing and comparing Montgomery multiplication algorithms — assessing five algorithms that speed up modular exponentiation, the most popular method of encrypting and signing digital data. IEEE Micro, 16(3):26–33, May/June 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2022:YYP

[KAK^{+}22] Joon Kyung Kim, Byung Hoon Ahn, Sean Kinzer, Soroush Ghodrati, Rohan Mahapatra, Brahmendra Yatham, Shu-Ting Wang, Dohee Kim, Parisa Sarikhani, Babak Mahmoudi, Divya Mahajan, Jongse Park, and Hadi Esmaeilzadeh. Yin-Yang: Programming abstractions for cross-domain multi-acceleration. IEEE Micro, 42(5):89–98, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaliski:1993:SES

[Kal93] Burt Kaliski. A survey of encryption standards. IEEE Micro, 13(6):74–81, November/December 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kalapathy:1997:HSI

[Kal97] Paul Kalapathy. Hardwaresoftware interactions on Mpact. IEEE Micro, 17(2):20–26, March/April 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kandel:1995:GEI

[Kan95] Abraham Kandel. Guest Editor’s introduction: The fuzzy boom. IEEE Micro, 15(4): 6–7, July/August 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kongetira:2005:NWM

[KAO05] Poonacha Kongetira, Kathirgamar Aingaran, and Kunle Olukotun. Niagara: a 32-way multithreaded Sparc processor. IEEE Micro, 25(2):21– 29, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2005/02/m2021abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2005/02/m2021.
                                        pdf.

Kartashev:1985:RRS

[Kar85] S. Kartashev. Review referee states concerns — reply. IEEE Micro, 5(5):91– 93, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Karjala:1988:PAP

[Kar88a] D. Karjala. Protection against piracy — comments. IEEE Micro, 8(3):89, May/June 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Karjala:1988:CLF

[Kar88b] D. S. Karjala. Copyright losses to foreign pirates. IEEE Micro, 8(5):2–??, September/October 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Karpuzcu:2021:SIQ

[Kar21] Ulya R. Karpuzcu. Special issue on quantum computing. IEEE Micro, 41(5):6– 7, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Katayama:1997:TSM

[Kat97] Yasunao Katayama. Trends in semiconductor memories. IEEE Micro, 17(6):10–17, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m6010.
pdf;   http://www.computer.
org/micro/mi1997/m6010abs.
htm.

Kunkel:1999:SOO

[KAV99] Steven Kunkel, Bill Armstrong, and Philip Vitale. System optimization for OLTP workloads. IEEE Micro, 19 (3):56–65, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3056.
pdf;   http://www.computer.
org/micro/mi1999/m3056abs.
htm.

Kawamura:1998:CID

[Kaw98] Shoji Kawamura. Capturing images with digital still cameras. IEEE Micro, 18(6): 14–19, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m6014.
                                        pdf;   http://www.computer.
                                        org/micro/mi1998/m6014abs.
                                        htm.

Knaflitz:1991:CAM

[KB91] Marco Knaflitz and Gabriella Balestra. Computer analysis of the myoelectric signal. IEEE Micro, 11(5):12–15, 48– 58, September/October 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kash:2013:SIS

[KB13] Jeffrey A. Kash and Raymond G. Beausoleil. Special issue: Selected research from the First Optical Interconnects Conference. IEEE Micro, 33(1):3–5, January/February 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kozyrakis:2020:HCR

[KB20] C. Kozyrakis and I. Bratt. The hot chips renaissance. IEEE Micro, 40(2):6–7, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Krashinsky:2004:VTA

[KBH^{+}04] Ronny Krashinsky, Christopher Batten, Mark Hampton, Steve Gerding, Brian Pharris, Jared Casper, and Krste Asanovic. The vectorthread architecture. IEEE Micro, 24(6):84–90, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6084.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6084.pdf.

Knauerhase:2008:UOI

[KBH^{+}08] Rob Knauerhase, Paul Brett, Barbara Hohlt, Tong Li, and Scott Hahn. Using OS observations to improve performance in multicore systems. IEEE Micro, 28(3): 54–66, May/June 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2003:NCA

[KBK03] Changkyu Kim, Doug Burger, and Stephen W. Keckler. Nonuniform cache architectures for wire-delay dominated on-chip caches. IEEE Micro, 23(6):99–107, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6099abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/06/m6099.
htm;   http://csdl.computer.

                                        org/dl/mags/mi/2003/06/
                                        m6099.pdf.

Krishnan:2016:EEG

[KBN16] Guhan Krishnan, Dan Bouvier, and Samuel Naffziger. Energy-efficient graphics and multimedia in 28-nm Carrizo accelerated processing unit. IEEE Micro, 36(2):22– 33, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2016/02/mmi2016020022-
                                        abs.html.

Klauer:1995:AP

[KBW95] Bernd Klauer, Andreas Bleck, and Klaus Waldschmidt. The AM^{3} associative processor. IEEE Micro, 15(2):70–78, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kursun:2009:TVC

[KC09] Eren Kursun and Chen-Yong Cher. Temperature variation characterization and thermal management of multicore architectures. IEEE Micro, 29(1):116–126, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaxiras:2018:NSL

[KCAR18] Stefanos Kaxiras, Trevor E. Carlson, Mehdi Alipour, and Alberto Ros. Non-speculative load reordering in total store ordering. IEEE Micro, 38(3): 48–57, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030048-
abs.html.

Kattan:2021:DMCa

[KCHA21a] Hammam Kattan, Sung Woo Chung, Jörg Henkel, and Hussam Amrouch. On-demand mobile CPU cooling with thin-film thermoelectric array. IEEE Micro, 41(4):67– 73, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kattan:2021:DMCb

[KCHA21b] Hammam Kattan, Sung Woo Chung, Jörg Henkel, and Hussam Amrouch. On-demand mobile CPU cooling with thin-film thermoelectric array. IEEE Micro, 41(4):67– 73, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Krishna:2014:SSC

[KCKP14] Tushar Krishna, Chia-Hsin Owen Chen, Woo-Cheol Kwon, and Li-Shiuan Peh. Smart: Singlecycle multihop traversals over a shared network on chip. IEEE Micro, 34(3):43–56, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Kwon:2020:MDC

[KCS^{+}20] H. Kwon, P. Chatarasi, V. Sarkar, T. Krishna, M. Pellauer, and A. Parashar. MAE-STRO: A data-centric approach to understand reuse, performance, and hardware cost of DNN mappings. IEEE Micro, 40(3):20–29, May/ June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2017:HCM

[KCXmWH17] Nam Sung Kim, Deming Chen, Jinjun Xiong, and Wen mei W. Hwu. Heterogeneous computing meets near-memory acceleration and high-level synthesis in the post-Moore era. IEEE Micro, 37(4):10–18, July/ August 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/04/mmi2017040010-
                                        abs.html.

Kanev:2016:PWS

[KDH^{+}16] Svilen Kanev, Juan Pablo Darago, Kim Hazelwood, Parthasarathy Ranganathan, Tipp Moseley, Gu-Yeon Wei, and David Brooks. Profiling a warehouse-scale computer. IEEE Micro, 36(3): 54–59, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

mi/2016/03/mmi2016030054-
abs.html.

Kopetz:1989:DFT

[KDK^{+}89] Hermann Kopetz, Andreas Damm, Christian Koza, Marco Mulazzani, Wolfgang Schwabl, Christoph Senft, and Ralph Zainlinger. Distributed fault-tolerant real-time systems — the Mars approach. IEEE Micro, 9(1):25–40, January/February 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Khailany:2001:IMP

[KDK^{+}01] Brucek Khailany, William J.

Dally, Ujval J. Kapasi, Peter Mattson, Jinyung Namkoong, John D. Owens, Brian Towles, Andrew Chang, and Scott Rixner. Imagine: Media processing with streams. IEEE Micro, 21(2):35–46, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2035.
pdf;   http://www.computer.
org/micro/mi2001/m2035abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Keckler:2011:GFP

[KDK^{+}11] Stephen W. Keckler, William J.

Dally, Brucek Khailany, Michael Garland, and David Glasco. GPUs and the future of parallel computing. IEEE Micro, 31(5):7–17, September/October 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2009:CED

[KDSA09] John Kim, William Dally, Steve Scott, and Dennis Abts. Cost-efficient dragonfly topology for large-scale systems. IEEE Micro, 29(1):33–40, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kabuka:1989:RTI

[KE89] Mansur Kabuka and Rodrigo Escoto. Real-time implementation of the Newton-Euler equations of motion on the NEC mu PD77230 DSP. IEEE Micro, 9(1):66–76, January/February 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kessler:1999:AM

[Kes99] R. E. Kessler. The Alpha-21264 microprocessor. IEEE Micro, 19(2):24–36, March/ April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m2024.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m2024abs.
                                        htm.

Kapadia:2000:PWP

[KFF00] Nirav H. Kapadia, Renato J.

Figueiredo, and José A. B. Fortes. Punch — Web portal for running tools. IEEE Micro, 20(3):38–47, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3038.
pdf;   http://www.computer.
org/micro/mi2000/m3038abs.
htm.

Keller:2005:TBV

[KG05] Jörg Keller and Andreas Grävinghoff. Thread-based virtual duplex systems in embedded environments. IEEE Micro, 25(2):60–69, March/ April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2005/02/m2060abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2005/02/
m2060.pdf.

Kleeberger:2013:CLT

[KGDW^{+}13] Veit B. Kleeberger, Christina Gimmler-Dumont, Christian Weis, Andreas Herkersdorf, Daniel Mueller-Gritschneder, Sani R. Nassif, Ulf Schlichtmann, and Norbert Wehn. A cross-layer technology-based study of how memory errors impact system resilience. IEEE Micro, 33(4):46–55, July/August 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Khazraee:2017:SPC

[KGMT17] Moein Khazraee, Luis Vega Gutierrez, Ikuo Magaki, and Michael Bedford Taylor. Specializing a planet’s computation: ASIC clouds. IEEE Micro, 37(3):62–69, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/03/mmi2017030062-
                                        abs.html.

Kolli:2019:LSM

[KGS^{+}19] A. Kolli, V. Gogte, A. Saidi, S. Diestelhorst, W. Wang, P. M. Chen, S. Narayanasamy, and T. F. Wenisch. Language support for memory persistency. IEEE Micro, 39(3):94– 102, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kumar:2022:PDP

[KGT22] Prabhat Kumar, Govind P. Gupta, and Rakesh Tripathi. PEFL: Deep privacyencoding-based federated learning framework for smart agriculture. IEEE Micro, 42 (1):33–40, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Khalid:2000:VTD

[Kha00] Humayn Khalid. Validating trace-driven microarchitectural simulations. IEEE Micro, 20(6):76–82, November/December 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2000/
pdf/m6076.pdf;   http://
www.computer.org/micro/
mi2000/m6076abs.htm.

Kahaner:1986:MSB

[KHF86] David K. Kahaner, Jeffrey Horlick, and Debra K. Foer. Mathematical software in BA-SIC: RV, generation of uniform and normal random variables. IEEE Micro, 6(3):52– 60, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Knight:1985:ESS

[KHHR85] Don Knight, Mikel J. Harry, Jim Howard, and Juan Rivero. Engineering support systems for engineering managers. IEEE Micro, 5(5):22– 26, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2016:HDS

[KHL^{+}16] Bongjun Kim, Seonyeong Heo, Gyeongmin Lee, Soyeon Park, Hanjun Kim, and Jong Kim. Heterogeneous distributed shared memory for lightweight Internet of Things devices. IEEE Micro, 36(6):16–24, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060016-
abs.html.

Kosuge:2023:FAC

[KHS^{+}23] Atsutake Kosuge, Yao-Chung Hsu, Rei Sumikawa, Mototsugu Hamada, Tadahiro Kuroda, and Tomoe Ishikawa. A 10.7-μJ/frame 88% accuracy CIFAR-10 singlechip neuromorphic fieldprogrammable gate array processor featuring various nonlinear functions of dendrites in the human cerebrum. IEEE Micro, 43(6):19–27, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1985:MSB

[KHW85] David K. Kahaner, Jeffrey Horlick, and Webb L. Wyman. Mathematical software in BASIC-DINT: data integration. IEEE Micro, 5(2):76– 82, March/April 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kidd:2014:PCO

[Kid14] David Kidd. Process and circuit optimization for power reduction using DDC transistors. IEEE Micro, 34(2):54– 62, March/April 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Komuro:2009:QSP

[KII09] Takashi Komuro, Atsushi Iwashita, and Masatoshi Ishikawa. A QVGA-size pixel-parallel image processor for 1,000-fps vision. IEEE Micro, 29(6):58–67, November/December 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kimura:2009:FHM

[KIM^{+}09] Motoki Kimura, Kenichi Iwata, Seiji Mochizuki, Hiroshi Ueda, Masakazu Ehama, and Hiromi Watanabe. A full HD multistandard video codec for mobile applications. IEEE Micro, 29(6):18–27, November/December 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2020:TPC

[Kim20] H. Kim. The 2019 top picks in computer architecture. IEEE Micro, 40(3):6–9, May/June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1983:WDB

[Kir83a] H. Kirrmann. What defines a 32-bit microprocessor? — reply. IEEE Micro, 3(6):

4, November/December 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1983:DFB

[Kir83b] Hubert Kirrmann. Data format and bus compatibility in multiprocessors. IEEE Micro, 3(4):32–47, July/August 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1984:DFB

[Kir84a] H. Kirrmann. Data format and the S-100 bus — reply. IEEE Micro, 4(1):11–14, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1984:MDC

[Kir84b] H. Kirrmann. Microprocessors defy classification — reply. IEEE Micro, 4(3):3–6, May/ June 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1985:RPM

[Kir85a] H. Kirrmann. Report on the Paris Multibus-II meeting and some thoughts about the future of the committee. IEEE Micro, 5(4):82– 89, July/August 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1985:EIT

[Kir85b] Hubert Kirrmann. Events and interrupts in tightly coupled multiprocessors. IEEE Micro, 5(1):53–66, January/February 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1987:FTP

[Kir87] Hubert D. Kirrmann. Fault tolerance in process-control — an overview and examples of European products. IEEE Micro, 7(5):27–50, September/October 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1988:E

[Kir88a] H. Kirrmann. Europe in 1992. IEEE Micro, 8(5): 5, September/October 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1988:EIB

[Kir88b] H. Kirrmann. Europe’s industrial bus scene. IEEE Micro, 8 (6):6–7, November/December 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1989:FTC

[Kir89a] H. Kirrmann. Fault-tolerant computing in Europe. IEEE Micro, 9(2):5–7, March/April 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1989:MSR

[Kir89b] H. Kirrmann. Multiprocessors and supercomputer research in Europe. IEEE Micro, 9(1): 7–8, January/February 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1989:NCN

[Kir89c] H. Kirrmann. Neural computing — the new Gold Rush in informatics. IEEE Micro, 9 (3):7–8, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1989:SSS

[Kir89d] H. Kirrmann. The Smaky story — the Swiss personalcomputer. IEEE Micro, 9 (4):78–79, July/August 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1990:CNW

[Kir90a] H. Kirrmann. COCOM — the next wall to fall. IEEE Micro, 10(1):4–6, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1990:MFL

[Kir90b] H. Kirrmann. Minitel — the French love affair with telematics. IEEE Micro, 10(2):88– 90, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1990:REW

[Kir90c] H. Kirrmann. Reunifying East-West electronics industries. IEEE Micro, 10(6):5– 7, November/December 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1990:TB

[Kir90d] H. Kirrmann. Train buses.

IEEE Micro, 10(5):79–80, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1990:TCS

[Kir90e] H. Kirrmann. Train controlsystems. IEEE Micro, 10 (4):79–80, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1991:EWM

[Kir91a] H. Kirrmann. Europe on the way to the metric system. IEEE Micro, 11(4):36– 39, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1991:LEC

[Kir91b] H. Kirrmann. Light at the end of the chunnel. IEEE Micro, 11(3):4–6, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1991:WCW

[Kir91c] H. Kirrmann. When the computer was still personal. IEEE Micro, 11(1):42–44, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:1992:HES

[Kir92] H. Kirrmann. Hermes, the European space-shuttle. IEEE Micro, 12(2):3–5, March/ April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrman:2001:LEP

[Kir01] Hubert Kirrman. Letter to the editor: On “Preventing abuse of IEEE standards policy”. IEEE Micro, 21(4):5, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4005.
pdf;   m4005abs.htm. See

[Ste01e].

Kim:2019:IRA

[KIR19] Y. Kim, M. Imani, and T. Rosing. Image recognition accelerator design using in-memory processing. IEEE Micro, 39(1):17–23, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kunimatsu:2000:VUA

[KIS^{+}00] Atsushi Kunimatsu, Nobuhiro Ide, Toshinori Sato, Yukio Endo, Hiroaki Murakami, Takayuki Kamei, Masashi Hirano, Fujio Ishihara, Haruyuki Tago, Masaaki Oka, Akio Ohba, Teiji Yutaka, Toyoshi Okada, and Masakazu Suzuoki. Vector unit architecture for emotion synthesis. IEEE Micro, 20(2):40–47, March/ April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2040.
pdf;   http://www.computer.
org/micro/mi2000/m2040abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Kwon:2023:FTT

[KJC^{+}23] Miryeong Kwon, Junhyeok Jang, Hanjin Choi, Sangwon Lee, and Myoungsoo Jung. Failure tolerant training with persistent memory disaggregation over CXL. IEEE Micro, 43(2):66–75, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kelm:2010:TCM

[KJL^{+}10] John H. Kelm, Daniel R.

Johnson, Steven S. Lumetta, Sanjay J. Patel, and Matthew I. Frank. A task-centric memory model for scalable accelerator architectures. IEEE Micro, 30 (1):29–39, January/February 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kannan:2016:EIT

[KJL16] Ajaykumar Kannan, Natalie Enright Jerger, and Gabriel H. Loh. Exploiting interposer technologies to disintegrate and reintegrate multicore processors. IEEE Micro, 36(3):84–93, May/June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/03/mmi2016030084-
                                        abs.html.

Kim:2007:DMP

[KJMP07] Hyesoon Kim, José A. Joao, Onur Mutlu, and Yale N. Patt. Diverge-merge processor: Generalized and energyefficient dynamic predication. IEEE Micro, 27(1):94–104, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2013:ASG

[KJP^{+}13] Youngtaek Kim, Lizy Kurian John, Sanjay Pant, Srilatha Manne, Michael Schulte, W. Lloyd Bircher, and Madhu Saravana Sibi Govindan. Automating stressmark generation for testing processor voltage fluctuations. IEEE Micro, 33(4):66–75, July/August 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kelm:2011:CAH

[KJT^{+}11] John H. Kelm, Daniel R.

Johnson, William Tuohy, Steven S. Lumetta, and Sanjay J. Patel. Cohesion: An adaptive hybrid memory model for accelerators. IEEE Micro, 31(1):42–55, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaxiras:2010:SCS

[KK10] Stefanos Kaxiras and Georgios Keramidas. SARC coherence: Scaling directory cache coherence in performance and power. IEEE Micro, 30(5):54– 65, September/October 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2023:SIE

[KK23] John Kim and Nam Sung Kim. Special issue on emerging system interconnects. IEEE Micro, 43(2):6– 8, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kleinhans:1993:SHS

[KKC93] Uwe Kleinhans, Joerg Kaiser, and Karol Czaja. Spearmints: hardware support for performance measurements in distributed systems. IEEE Micro, 13(5):69–78, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirman:2007:COT

[KKD^{+}07] Nevin Kırman, Meyrem Kırman, Rajeev K. Dokania, José F. Martínez, Alyssa B. Apsel, Matthew A. Watkins, and David H. Albonesi. Onchip optical technology in future bus-based multicore designs. IEEE Micro, 27(1):56– 66, January/February 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2024:CNC

[KKH^{+}24] Sangyeob Kim, Soyeon Kim, Seongyon Hong, Sangjin Kim, Jiwon Choi, Donghyeon Han, and Hoi-Jun Yoo. COOL-NPU: Complementary online learning neural processing unit. IEEE Micro, 44(1):28– 37, January/February 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Krishnaiyer:2000:AOI

[KKL^{+}00] Rakesh Krishnaiyer, Dattatraya Kulkarni, Daniel Lavery, Wei Li, Chu cheow Lim, John Ng, and David Sehr. An advanced optimizer for the IA-64 architecture. IEEE Micro, 20(6): 60–68, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6060.
pdf;   http://www.computer.
org/micro/mi2000/m6060abs.
htm.

Kim:2009:RTO

[KKL^{+}09] Joo-Young Kim, Minsu Kim, Seungjin Lee, Jinwook Oh, Sejong Oh, and Hoi-Jun Yoo. Real-time object recognition with neuro-fuzzy controlled workload-aware task pipelining. IEEE Micro, 29(6): 28–43, November/December 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2022:AXH

[KKL^{+}22] Jin Hyun Kim, Shin-Haeng Kang, Sukhan Lee, Hyeonsu Kim, Yuhwan Ro, Seungwon Lee, David Wang, Jihyun Choi, Jinin So, YeonGon Cho, JoonHo Song, Jeonghyeon Cho, Kyomin Sohn, and Nam Sung Kim. Aquabolt-XL HBM2-PIM, LPDDR5-PIM with in-memory processing, and AXDIMM with acceleration buffer. IEEE Micro, 42(3):20–30, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Krishna:2009:EVC

[KKP^{+}09] Tushar Krishna, Amit Kumar, Li-Shiuan Peh, Jacob Postman, Patrick Chiang, and Mattan Erez. Express virtual channels with capacitively driven global links. IEEE Micro, 29(4):48–61, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2014:ARP

[KKP^{+}14] Gyeonghoon Kim, Donghyun Kim, Seongwook Park, Youchang Kim, Kyuho Lee, Injoon Hong, Kyeongryeol Bong, and Hoi-Jun Yoo. An augmented reality processor with a congestion-aware networkon-chip scheduler. IEEE Micro, 34(6):31–41, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060031-
abs.html.

Koyanagi:1998:FSS

[KKS^{+}98] Mitsumasa Koyanagi, Hiroyuki Kurino, Katsuyuki Sakuma, Kang Wook Lee, Nobuaki Miyakawa, and Hikotaro Itani. Future system-on-silicon LSI chips. IEEE Micro, 18(4):17–22, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4017.
pdf;   http://www.computer.
org/micro/mi1998/m4017abs.
htm.

Kim:2023:SSD

[KKS^{+}23] Kyungsan Kim, Hyunseok Kim, Jinin So, Wonjae Lee, Junhyuk Im, Sungjoo Park, Jeonghyeon Cho, and Hoyoung Song. SMT: Softwaredefined memory tiering for heterogeneous computing systems with CXL memory expander. IEEE Micro, 43 (2):20–29, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kozyrakis:2010:SEI

[KKSV10] Christos Kozyrakis, Aman Kansal, Sriram Sankar, and Kushagra Vaid. Server engineering insights for large-scale online services. IEEE Micro, 30(4):8–19, July/August 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kuo:1991:KES

[KKT^{+}91] Yau-Hwang Kuo, Ling-Yang Kung, Ching-Chung Tzeng, Guang-Huei Jeng, and Wei-Kuo Chia. KMDS: an expert system for integrated hardware/software design of microprocessor-based digital systems. IEEE Micro, 11(4): 32–35, 86–92, July/August 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Karpuzcu:2013:CPV

[KKT13] Ulya R. Karpuzcu, Nam Sung Kim, and Josep Torrellas.

Coping with parametric variation at near-threshold voltages. IEEE Micro, 33(4):6– 14, July/August 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kimura:1988:IVV

[KKY88] Shinya Kimura, Yasuhiko Komoto, and Yoichi Yano. Implementation of the V60/V70 and its FRM function. IEEE Micro, 8(2):22–36, March/ April 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kodi:2005:DHS

[KL05] Avinash Karanth Kodi and Ahmed Louri. Design of a high-speed optical interconnect for scalable sharedmemory multiprocessors. IEEE Micro, 25(1):41–49, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1041.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1041.pdf.

Kodi:2008:OSS

[KL08] Avinash Karanth Kodi and Ahmed Louri. Optisim: a system simulation methodology for optically interconnected HPC systems. IEEE Micro, 28(5):22–36, September/October 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Karlsson:1994:UHI

[KLD^{+}94] Johan Karlsson, Peter Liden, Peter Dahlgren, Rolf Johansson, and Ulf Gunneflo. Using heavy-ion radiation to validate fault-handling mechanisms. IEEE Micro, 14(1): 8–23, January/February 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Klingman:1981:DPM

[Kli81a] Edwin E. Klingman. A design philosophy for microcomputers. IEEE Micro, 1(1):58– 64, January/March 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Klingman:1981:HCM

[Kli81b] Edwin E. Klingman. Hierarchical coding of microcomputers for high-level architecture. IEEE Micro, 1(1):53– 56, January/March 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2015:AAS

[KLM^{+}15] Hyojong Kim, Hongyeol Lim, Dilan Manatunga, Hyesoon Kim, and Gi-Ho Park. Accelerating application startup with nonvolatile memory in Android systems. IEEE Micro, 35(1):15–25, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010015-
abs.html.

Kloker:1986:MDD

[Klo86] Kevin L. Kloker. The Motorola DSP56000 digital signal processor. IEEE Micro, 6(6): 29–48, November/December 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kohn:1989:III

[KM89] Les Kohn and Neal Margulis. Introducing the Intel i860 64bit microprocessor. IEEE Micro, 9(4):15–30, July/August 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Koufaty:2003:HTN

[KM03] David Koufaty and Deborah T. Marr. Hyperthreading technology in the netburst microarchitecture. IEEE Micro, 23(2):56–65, March/April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m2056.
pdf;   http://www.computer.
org/micro/mi2003/m2056abs.
htm.

Krishnan:2005:LCC

[KM05] Venkata Krishnan and David Mayhew. Localized congestion control in advanced switching interconnects. IEEE Micro, 25(1):10–18, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1010.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1010.pdf.

Keltcher:2003:AOP

[KMAC03] Chetana N. Keltcher, Kevin J.

McGrath, Ardsher Ahmed, and Pat Conway. The AMD Opteron processor for multiprocessor servers. IEEE Micro, 23(2):66–76, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m2066.
pdf;   http://www.computer.
org/micro/mi2003/m2066abs.
htm.

Kleveland:2013:IRS

[KMD^{+}13] Bendik Kleveland, Michael John Miller, Ronald B. David, Jay Patel, Rajesh Chopra, Dipak K. Sikdar, Jeff Kumala, Socrates D. Vamvakos, Mike Morrison, Ming Liu, and Jayaprakash Balachandran. An intelligent RAM with serial I/Os. IEEE Micro, 33(6): 56–65, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Karadayi:2003:SMA

[KMG^{+}03] Kerem Karadayi, Vishal Markandey, Jeremiah Golston, Robert J. Gove, and Yongmin Kim. Strategies for mapping algorithms to mediaprocessors for high performance. IEEE Micro, 23 (4):58–70, July/August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/04/m4058abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/04/m4058.
pdf.

Kim:2001:DCD

[KMK01] Donglok Kim, Ravia Managuli, and Yongmin Kim. Data cache and direct memory access in programming mediaprocessors. IEEE Micro, 21(4):33–42, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4033.
pdf;   m4033abs.htm.

Karandikar:2019:FFA

[KMK^{+}19] S. Karandikar, H. Mao, D. Kim, D. Biancolin, A. Amid, D. Lee, N. Pemberton, E. Amaro, C. Schmidt, A. Chopra, Q. Huang, K. Kovacs, B. Nikoli, R. H. Katz, J. Bachrach, and K. Asanovi. FireSim: FPGA-accelerated cycle-exact scale-out system simulation in the public cloud. IEEE Micro, 39(3): 56–65, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kapil:2004:CMP

[KML04] Sanjiv Kapil, Harlan McGhan, and Jesse Lawrendra. A chip multithreaded processor for network-facing workloads. IEEE Micro, 24(2):20– 30, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/02/m2020abs.htm;
http://csdl.computer.org/
dl/mags/mi/2004/02/m2020.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/
m2020.pdf.

Karim:2004:MCA

[KMN^{+}04] Faraydon Karim, Alain Mellan, Anh Nguyen, Utku Aydonat, and Tarek Abdelrahman. A multilevel computing architecture for embedded multimedia applications. IEEE Micro, 24(3): 56–66, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/03/m3056.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/03/m3056.pdf.

Kim:2006:WBE

[KMPS06] Hyesoon Kim, Onur Mutlu, Yale N. Patt, and Jared Stark. Wish branches: Enabling adaptive and aggressive predicated execution. IEEE Micro, 26(1):48–58, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kodi:2014:PIE

[KNB14] Avinash Karanth Kodi, Brian Neel, and William C. Brantley. Photonic interconnects for exascale and datacenter architectures. IEEE Micro, 34(5):18–30, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050018-
abs.html.

Karim:2002:IAN

[KND02] Faraydon Karim, Anh Nguyen, and Sujit Dey. An interconnect architecture for networking systems on chips. IEEE Micro, 22(5):36–45, September/October 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2002/
pdf/m5036.pdf;   http://
www.computer.org/micro/
mi2002/m5036abs.htm.

Knight:1985:EWE

[Kni85] D. O. Knight. The engineering workstation and the engineering support system — present status, future-directions — introduction. IEEE Micro, 5(5):6–9, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaneko:1990:PED

[KNN^{+}90] Katsuyuki Kaneko, Masaitsu Nakajima, Yasuhiro Nakakura, Junji Nishikawa, Ichiro Okabayashi, and Hiroshi Kadota. Processing element design for a parallel computer. IEEE Micro, 10(2):26– 38, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Keshavarzi:2020:FEI

[KNV^{+}20] A. Keshavarzi, K. Ni, W. Van Den Hoek, S. Datta, and A. Raychowdhury. Ferro-Electronics for edge intelligence. IEEE Micro, 40(6): 33–48, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kota:2005:HLS

[KO05] Rajesh Kota and Rich Oehler.

Horus: Large-scale symmetric multiprocessing for Opteron systems. IEEE Micro, 25 (2):30–40, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2030abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2005/02/m2030.
pdf.

Koeman:1986:ASI

[Koe86] Henriecus Koeman. Applicationspecific IC design technologies — a system designers overview. IEEE Micro, 6 (1):42–50, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kumanoya:1995:ADI

[KOI95] Masaki Kumanoya, Toshiyuki Ogawa, and Kazunari Inoue. Advances in DRAM interfaces. IEEE Micro, 15(6): 30–36, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kojima:2023:SBB

[KOKA23] Takuya Kojima, Hayate Okuhara, Masaaki Kondo, and Hideharu Amano. A scalable body bias optimization method toward low-power CGRAs. IEEE Micro, 43 (1):49–57, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Koopman:1988:WP

[Koo88] P. Koopman. WISCs propagate. IEEE Micro, 8(2): 5, March/April 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Koopman:2002:GEI

[Koo02] Philip Koopman. Guest Editor’s introduction: Critical embedded automotive networks. IEEE Micro, 22 (4):14–18, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4014.
pdf;   http://www.computer.
org/micro/mi2002/m4014abs.
htm.

Kumar:1990:DCD

[KP90] Krishna A. Kumar and Brian Petrasko. Designing a custom DSP circuit using VHDL. IEEE Micro, 10(5): 46–53, September/October 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kozyrakis:2003:SVP

[KP03] Christoforos E. Kozyrakis and David A. Patterson. Scalable vector processors for embedded systems. IEEE Micro, 23(6):36–45, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6036abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/06/m6036.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6036.pdf.

Kundu:2007:GEI

[KP07] Partha Kundu and Li-Shiuan Peh. Guest Editors’ introduction: On-chip interconnects for multicores. IEEE Micro, 27(5):3–5, September/October 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/05/mmi2007050003.
pdf.

Krall:2004:XCB

[KPHP04] Andreas Krall, Ivan Pryanishnikov, Ulrich Hirnschrott, and Christian Panis. xDSPcore: a compiler-based configurable digital signal processor. IEEE Micro, 24(4):67– 78, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4067abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/04/m4067.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/04/
m4067.pdf.

Katevenis:2010:ECS

[KPK^{+}10] Manolis Katevenis, Vassilis Papaefstathiou, Stamatis Kavadias, Dionisios Pnevmatikatos, Federico Silla, and Dimitrios Nikolopoulos. Explicit communication and synchronization in SARC. IEEE Micro, 30(5):30–41, September/October 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kumar:2008:TIC

[KPKJ08] Amit Kumar, Li-Shiuan Peh, Partha Kundu, and Niraj K. Jha. Toward ideal on-chip communication using express virtual channels. IEEE Micro, 28(1):80–90, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2011:TCM

[KPMHB11] Yoongu Kim, Michael Papamichael, Onur Mutlu, and Mor Harchol-Balter. Thread cluster memory scheduling. IEEE Micro, 31(1):78–89, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Knechtel:2020:PSC

[KPN^{+}20] J. Knechtel, S. Patnaik, M. Nabeel, M. Ashraf, Y. S. Chauhan, J. Henkel, O. Sinanoglu, and H. Amrouch. Power side-channel attacks in negative capacitance transistor. IEEE Micro, 40(6): 74–84, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kistler:2006:CMC

[KPP06] Michael Kistler, Michael Perrone, and Fabrizio Petrini. Cell multiprocessor communication network: Built for speed. IEEE Micro, 26(3):

10–23, May/June 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kokolis:2022:DDP

[KPR^{+}22] Apostolos Kokolis, Antonis Psistakis, Benjamin Reidys, Jian Huang, and Josep Torrellas. Distributed data persistency. IEEE Micro, 42(4): 107–115, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kornaros:1999:AIS

[KPV^{+}99] Georgios Kornaros, Dionisios Pnevmatikatos, Panagiota Vatsolaki, Georgios Kalokerinos, Chara Xanthaki, Dimitrios Mavroidis, Dimitrios Serpanos, and Manolis Katevenis. ATLAS I: Implementing a single-chip ATM switch with backpressure. IEEE Micro, 19(1):30– 41, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1030.
pdf;   http://www.computer.
org/micro/mi1999/m1030abs.
htm.

Kubiatowicz:2019:HC

[KR19a] J. Kubiatowicz and S. Rusu. Hot Chips 30. IEEE Micro, 39(2):6–8, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kwon:2019:DMS

[KR19b] Y. Kwon and M. Rhu. A disaggregated memory system for deep learning. IEEE Micro, 39(5):82–90, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kramer:1996:ABA

[Kra96] Alan H. Kramer. Arraybased analog computation: Computing billions of regular low-level operations efficiently on mW of power. IEEE Micro, 16(5):20–29, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Khailany:2020:ACD

[KRD^{+}20] B. Khailany, H. Ren, S. Dai, S. Godil, B. Keller, R. Kirby, A. Klinefelter, R. Venkatesan, Y. Zhang, B. Catanzaro, and W. J. Dally. Accelerating chip design with machine learning. IEEE Micro, 40(6): 23–32, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kitahara:1990:GBM

[KS90] Takeshi Kitahara and Taizo Satoh. The Gmicro/300 32-Bit microprocessor. IEEE Micro, 10(3):68–75, May/June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kirrmann:2000:LDF

[KS00] Hubert Kirrmann and Ken Sakamura. Letters: date formats. IEEE Micro, 20(1): 3, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m1003.
                                        pdf.

Kubiatowicz:2007:GEI

[KS07] John Kubiatowicz and Howard Sachs. Guest Editors’ introduction: Hot Chips 18. IEEE Micro, 27(2):7– 9, March/April 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2007/02/m2007.pdf.

Kahng:2011:BC

[KS11] Andrew B. Kahng and Vijayalakshmi Srinivasan. Big chips. IEEE Micro, 31(4):3– 5, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kim:2018:HA

[KS18] Martha Kim and Yakun Sophia Shao. Hardware acceleration. IEEE Micro, 38(6):6– 7, November/December 2018.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08585396-abs.
html.

Kang:2019:EEP

[KSA^{+}19] M. Kang, P. Srivastava, V. Adve, N. S. Kim, and N. R. Shanbhag. An energyefficient programmable mixedsignal accelerator for machine learning algorithms. IEEE Micro, 39(5):64–72, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Krishnan:2021:CNP

[KSB21] V. Krishnan, O. Serres, and M. Blocksome. Configurable Network Protocol Accelerator (COPA). IEEE Micro, 41(1): 8–14, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Koc:2022:CWT

[KSE^{+}22] Fahrettin Koc, Behzad Salami, Oguz Ergin, Osman Unsal, and Adrian Cristal Kestelman. Can we trust undervolting in FPGA-based deep learning designs at harsh conditions? IEEE Micro, 42(3): 57–65, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kondo:1996:TCM

[KSI^{+}96] Toshio Kondo, Kazuhito Suguri, Mitsuo Ikeda, Tetsuya Abe, Hiroaki Matsuda, Tsuneo Okubo, Kenji Ogura, Yutaka Tashiro, Naoki Ono, Toshihiro Minami, Ritsu Kusaba, Takeshi Ikenaga, Nobutaro Shibata, Ryota Kasai, Koji Otsu, Fumiaki Nakagawa, and Yasuhiko Sato. Two-chip MPEG-2 video encoder: Switching to simple profile at main level for a costeffective encoder. IEEE Micro, 16(2):51–58, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Kwon:2018:CCA

[KSK18] Hyoukjun Kwon, Ananda Samajdar, and Tushar Krishna. A communicationcentric approach for designing flexible DNN accelerators. IEEE Micro, 38(6):25–35, November/December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08527533.pdf.

Kim:2017:BLP

[KSLY17] Youchang Kim, Dongjoo Shin, Jinsu Lee, and Hoi-Jun Yoo. BRAIN: A low-power deep search engine for autonomous robots. IEEE Micro, 37(5):11– 19, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050011-
abs.html.

Komori:1989:DDM

[KSM^{+}89] Shinji Komori, Kenji Shima, Souichi Miyata, Toshiya Okamoto, and Hiroaki Terada. The data-driven microprocessor. IEEE Micro, 9(3): 45–59, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kneip:1999:AIM

[KSM99] Johannes Kneip, Bernd Schmale, and Henning Möller. Applying and implementing the MPEG-4 multimedia standard. IEEE Micro, 19(6): 64–74, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6064.
pdf;   http://www.computer.
org/micro/mi1999/m6064abs.
htm.

Kleiman:1999:UNI

[KSR^{+}99] Steven R. Kleiman, Scott Schoenthal, Alan Rowe, Steven H. Rodrigues, and Arputham Benjamin. Using NUMA interconnects for highly available filers. IEEE Micro, 19 (1):42–49, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m1042.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m1042abs.
                                        htm.

Kalla:2010:PIN

[KSSF10] Ron Kalla, Balaram Sinharoy, William J. Starke, and Michael Floyd. Power7: IBM’s next-generation server processor. IEEE Micro, 30(2): 7–15, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kalla:2004:IPC

[KST04] Ron Kalla, Balaram Sinharoy, and Joel M. Tendler. IBM Power5 chip: a dualcore multithreaded processor. IEEE Micro, 24(2):40– 47, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/02/m2040abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2004/02/m2040.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2004/02/
                                        m2040.pdf.

Kamruzzaman:2012:USP

[KST12] Md Kamruzzaman, Steven Swanson, and Dean M. Tullsen. Underclocked software prefetching: More cores, less energy. IEEE Micro, 32 (4):32–41, July/August 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Karageorgos:2021:BSV

[KSV^{+}21] Ioannis Karageorgos, Karthik Sriram, Ján Veselý, Nick Lindsay, Xiayuan Wen, Michael Wu, Marc Powell, David Borton, Rajit Manohar, and Abhishek Bhattacharjee. Balancing specialized versus flexible computation in brain– computer interfaces. IEEE Micro, 41(3):87–94, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaneko:1990:RVS

[KSWM90] Hiroaki Kaneko, Nariko Suzuki, Hiroshi Wabuka, and Koji Maemura. Realizing the V80 and its system support functions. IEEE Micro, 10(2):56– 69, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Keckler:2014:ISC

[KT14] Stephen W. Keckler and Dean Tullsen. 2014 International Symposium on Computer Architecture Influential Paper Award; 2014 Maurice Wilkes Award given to Ravi Rajwar. IEEE Micro, 34(6):95–97, November/December 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060095.
html.

Kornaros:2018:HAS

[KTC18] Georgios Kornaros, Othon Tomoutzoglou, and Marcello Coppola. Hardware-assisted security in electronic control units: Secure automotive communications by utilizing one-time-programmable network on chip and firewalls. IEEE Micro, 38(5):63– 74, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050063-
abs.html.

Kato:2015:OAA

[KTI^{+}15] Shinpei Kato, Eijiro Takeuchi, Yoshio Ishiguro, Yoshiki Ninomiya, Kazuya Takeda, and Tsuyoshi Hamada. An open approach to autonomous vehicles. IEEE Micro, 35(6): 60–68, November/December 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/csdl.computer.org/csdl/
mags/mi/2015/06/mmi2015060060-
abs.html.

Kambadur:2013:PBV

[KTK13] Melanie Kambadur, Kui Tang, and Martha A. Kim. Parallel block vectors: Collection, analysis, and uses. IEEE Micro, 33(3):86–94, May/June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kubo:2024:CLO

[KTY24] Tatsuya Kubo and Shinya Takamaeda-Yamazaki. Cachet: Low-overhead integrity verification on metadata cache in secure nonvolatile memory systems. IEEE Micro, 44 (1):38–48, January/February 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kulkarni:2020:CD

[Kul20] J. P. Kulkarni. Chip Design 2020. IEEE Micro, 40(6):6– 7, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kumar:1997:HPR

[Kum97] Ashok Kumar. The HP PA-8000 RISC CPU. IEEE Micro, 17(2):27–32, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

KurianJohn:2020:CD

[Kur20a] L. Kurian John. Chip Design 2020. IEEE Micro, 40(6):4– 5, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

KurianJohn:2020:MLS

[Kur20b]  L. Kurian John.  Ma-

chine learning for systems, biological computing, and more. IEEE Micro, 40(5):4– 5, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

KurianJohn:2021:CGM

[Kur21a] L. Kurian John. CPUs, GPUs, and more from Hot Chips 32. IEEE Micro, 41(2):4– 5, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

KurianJohn:2021:FCM

[Kur21b] Lizy Kurian John. FPGA computing and more! IEEE Micro, 41(4):4–5, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

KurianJohn:2021:QCM

[Kur21c] Lizy Kurian John. Quantum computing and more! IEEE Micro, 41(5):4–5, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

KurianJohn:2021:TPY

[Kur21d] Lizy Kurian John. Top picks from year 2020. IEEE Micro, 41(3):4–5, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kozyrakis:2009:HCT

[KvdW09] Christos Kozyrakis and Jan-Willem van de Waerdt. Hot Chips turns 20. IEEE Micro, 29(2):4–5, March/April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kinsel:1981:DSG

[KW81] Tracy S. Kinsel and John H.

Wuorinen, Jr. A digital signal generator. IEEE Micro, 1(4):6–15, October/December 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1983:MSB

[KW83] David K. Kahaner and Webb L. Wyman. Mathematical software in Basic: Evaluation of definite integrals. IEEE Micro, 3(5):42– 46, September/October 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kubiatowicz:2002:GEI

[KW02] John Kubiatowicz and Andrew Wolfe. Guest Editors’ introduction: Hot Chips 13. IEEE Micro, 22 (2):6–7, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2006.
pdf;   http://www.computer.
org/micro/mi2002/m2006abs.
htm.

Konig:1995:VII

[KWGG95] Andreas Konig, Peter Windirsch, Michael Gasteier, and Manfred Glesner. Visual inspection in industrial manufacturing — using a specially developed, easily trainable, dedicated VLSI neural network architecture to detect anomalies. IEEE Micro, 15(3):26– 31, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kawasaki:1989:FPV

[KWM89] Shumpei Kawasaki, Mitsuru Watabe, and Shigeki Morinaga. A floating-point VLSI chip for the TRON architecture: an architecture for reliable numerical programming. IEEE Micro, 9(3):26– 44, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kabemoto:1991:ASS

[KY91] Akira Kabemoto and Hiroshi Yoshida. The architecture of the Sure System 2000 communications processor. IEEE Micro, 11(4):28– 31, 73–78, July/August 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaplan:2019:RRP

[KYG19] R. Kaplan, L. Yavits, and R. Ginosar. RASSA: Resistive prealignment accelerator for approximate DNA long read mapping. IEEE Micro, 39 (4):44–54, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kaplan:2017:RCP

[KYGW17] Roman Kaplan, Leonid Yavits, Ran Ginosar, and Uri Weiser. A resistive CAM processing-in-storage architecture for DNA sequence alignment. IEEE Micro, 37(4):20– 28, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040020-
abs.html.

Kirrmann:2001:IIT

[KZ01] Hubert Kirrmann and Pierre A.

Zuber. The IEC/IEEE train communication network. IEEE Micro, 21(2): 81–92, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2081.
pdf;   http://www.computer.
org/micro/mi2001/m2081abs.
htm.

Kozyrakis:2013:SRH

[KZ13] Christos Kozyrakis and Rumi Zahir. Selected research from Hot Chips 24. IEEE Micro, 33(2):6–7, March/April 2013.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ke:2022:NMP

[KZS^{+}22] Liu Ke, Xuan Zhang, Jinin So, Jong-Geon Lee, Shin-Haeng Kang, Sukhan Lee, Songyi Han, YeonGon Cho, Jin Hyun Kim, Yongsuk Kwon, KyungSoo Kim, Jin Jung, Ilkwon Yun, Sung Joo Park, Hyunsun Park, Joonho Song, Jeonghyeon Cho, Kyomin Sohn, Nam Sung Kim, and Hsien-Hsin S. Lee. Nearmemory processing in action: Accelerating personalized recommendation with AxDIMM. IEEE Micro, 42(1):116–127, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lahr:1984:NDP

[Lah84] Roy J. Lahr. The non-death of paper. IEEE Micro, 4(5):18– 25, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Landry:1985:MSM

[Lan85a]  R. Landry.  Micromouse

squeaks at microprocessor forum. IEEE Micro, 5(3): 84, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Landry:1985:WEW

[Lan85b] R. Landry. What is an engineering workstation? IEEE Micro, 5(3):83–85, May/June 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Landry:1987:DLS

[Lan87] Steve L. Landry. ‘Designer’ logic and symbols with logic cell arrays. IEEE Micro, 7 (1):51–59, January/February 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Landolt:1996:ANF

[Lan96]  Olivier Landolt.  Analog

nonlinear function synthesis — designing low-power microsensor, microactuator interfaces. IEEE Micro, 16(5): 50–52, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lukowicz:2001:WML

[LAT^{+}01] Paul Lukowicz, Urs Anliker, Gerhard Tröster, Steven J. Schwartz, and Richard W. DeVaul. The WearARM modular, low-power computing core. IEEE Micro, 21 (3):16–28, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3016.
pdf;   http://www.computer.

                                        org/micro/mi2001/m3016abs.
                                        htm.

Lauterbach:2021:PSW

[Lau21] Gary Lauterbach. The path to successful wafer-scale integration: The Cerebras story. IEEE Micro, 41(6):52–57, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lavagno:2002:GEI

[Lav02] Luciano Lavagno. Guest Editor’s introduction: Systems on a Chip—the next electronic frontier. IEEE Micro, 22(5): 14–15, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m5014.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m5014abs.
                                        htm.

Lazzerini:1989:EVP

[Laz89] Beatrice Lazzerini. Effective VLSI processor architectures for HLL computers: the RISC approach. IEEE Micro, 9 (1):57–65, January/February 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lam:2000:GEI

[LB00] Monica Lam and Forest Baskett. Guest Editors’ introduction: Cutting-edge designs. IEEE Micro, 20(2):

14–15, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2014.

pdf. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Lee:2007:SSR

[LB07] Benjamin C. Lee and David M.

Brooks. Spatial sampling and regression strategies. IEEE Micro, 27(3):74–93, May/ June 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Laes:1999:COS

[LBD^{+}99] Edgard Laes, Livio Baldi, Claus Dahl, Frits R. J. Huisman, and Ludo Deferm. CMOS options for singlechip applications. IEEE Micro, 19(5):23–32, September/October 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1999/
pdf/m5023.pdf;   http://
www.computer.org/micro/
mi1999/m5023abs.htm.

Liu:2022:TME

[LBR^{+}22] Hsin-I Cindy Liu, Marius Brehler, Mahesh Ravishankar, Nicolas Vasilache, Ben Vanik, and Stella Laurenzo. TinyIREE: An ML execution environment for embedded systems from compilation to deployment. IEEE Micro, 42(5):9–16, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2011:IDS

[LBS^{+}11] HyoukJoong Lee, Kevin J. Brown, Arvind K. Sujeeth, Hassan Chafi, Kunle Olukotun, Tirark Rompf, and Martin Odersky. Implementing domain-specific languages for heterogeneous parallel computing. IEEE Micro, 31(5): 42–53, September/October 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:1991:SFV

[LC91] Hua Li and Ching-Ho H. Chen. Simulating a function of visual peripheral processes with an analog VLSI network. IEEE Micro, 11(5):8–11, 44– 47, September/October 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Levy:2009:EMP

[LC09] Markus Levy and Thomas M. Conte. Embedded multicore processors and systems. IEEE Micro, 29(3):7–9, May/June 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2018:AC

[LC18] Hsien-Hsin Sean Lee and Jason Clemons. Automotive computing. IEEE Micro, 38(1):29–30, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010029.
html.

Liu:2022:COA

[LCN^{+}22] Xu Liu, Steven W. Chen, Guilherme V. Nardari, Chao Qu, Fernando Cladera Ojeda, Camillo J. Taylor, and Vijay Kumar. Challenges and opportunities for autonomous micro-UAVs in precision agriculture. IEEE Micro, 42 (1):61–68, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lindtjorn:2011:BTM

[LCP^{+}11] Olav Lindtjorn, Robert Clapp, Oliver Pell, Haohuan Fu, Michael Flynn, and Oskar Mencer. Beyond traditional microprocessors for geoscience high-performance computing applications. IEEE Micro, 31 (2):41–49, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Laes:1992:ADT

[LCS92] Edgard Laes, Herman J.

Casier, and Eric Schutz.

Analog-digital technologies for mixed-signal processing — the driving force to success for the European industry. IEEE Micro, 12(4):34–42, July/August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liang:2008:RSD

[LCWB08] Xiaoyao Liang, Ramon Canal, Gu-Yeon Wei, and David Brooks. Replacing 6T SRAMs with 3T1D DRAMs in the L1 data cache to combat process variability. IEEE Micro, 28 (1):60–68, January/February 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liu:2004:MPC

[LCY^{+}04] Jiuxing Liu, Balasubramanian Chandrasekaran, Weikuan Yu, Jiesheng Wu, Darius Buntinas, Sushmitha Kini, Dhabaleswar K. Panda, and Pete Wyckoff. Microbenchmark performance comparison of high-speed cluster interconnects. IEEE Micro, 24(1):42– 51, January/February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1042abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/01/m1042.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/01/
m1042.pdf.

Li:1987:HND

[LDA87] Kin Fun Li, Nikitas J. Dimopoulos, and J. William Atwood. The HM-Nucleus: distributed kernel design for the Homogeneous Multiprocessor. IEEE Micro, 7(1):14–24, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lucia:2009:AAD

[LDCS09] Brandon Lucia, Joseph Devietti, Luis Ceze, and Karin Strauss. Atom-Aid: Detecting and surviving atomicity violations. IEEE Micro, 29(1):73– 83, January/February 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lefurgy:2013:AGM

[LDF^{+}13] Charles R. Lefurgy, Alan J.

Drake, Michael S. Floyd, Malcolm S. Allen-Ware, Bishop Brock, Jose A. Tierno, John B. Carter, and Robert W. Berry. Active guardband management in Power7+ to save energy and maintain reliability. IEEE Micro, 33(4):35– 45, July/August 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

LaBoda:2017:EDF

[LDL17] Craig LaBoda, Chris Dwyer, and Alvin R. Lebeck. Exploiting dark fluorophore states to implement resonance energy transfer pre-charge logic. IEEE Micro, 37(4):52–62, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040052-
abs.html.

Levy:2018:SER

[LE18] Hank Levy and Susan Eggers. Susan Eggers receives Eckert–Mauchly Award for Outstanding Contributions to Computer Architecture. IEEE Micro, 38(4):71– 75, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040071.
html.

Leahy:1985:EDC

[Lea85] P. Leahy. Electronic data communications privacy. IEEE Micro, 5(2):4–5, March/April 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lea:1988:ACE

[Lea88] R. M. Lea. AsP — a costeffective parallel microcomputer. IEEE Micro, 8(5):10– 29, September/October 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:1990:PDB

[Lee90]  Edward A. Lee.  Pro-

grammable DSPs — a brief overview. IEEE Micro, 10(5): 14–16, September/October 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:1994:TMD

[Lee94] Ruby B. Lee. Trends in microprocessor design. IEEE Micro, 14(2):7–9, March/April 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:1995:AME

[Lee95] Ruby B. Lee. Accelerating multimedia with enhanced microprocessors. IEEE Micro, 15 (2):22–32, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Lee:1996:SPM

[Lee96] Ruby B. Lee. Subword parallelism with MAX-2: Accelerating media processing with a minimal set of instruction extensions supporting efficient subword parallelism. IEEE Micro, 16(4):51– 59, July/August 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2021:SIC

[Lee21] Hsien-Hsin S. Lee. Special issue on commercial products 2021. IEEE Micro, 41(5): 66, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2024:BWF

[Lee24a] Hsien-Hsin S. Lee. Beyond wires: The future of interconnects. IEEE Micro, 44(2):4– 5, March/April 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2024:CCC

[Lee24b] Hsien-Hsin S. Lee. Computing with COOL chips. IEEE Micro, 44(1):4–5, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Leistner:1998:ASS

[Lei98]  Stacy Leistner.  Avoiding

surprises — some thoughts on standards. IEEE Micro, 18(3):25–32, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m3025.
                                        pdf;   http://www.computer.
                                        org/micro/mi1998/m3025abs.
                                        htm.

Levy:2023:SIH

[Lev23] Scott Levy. Special issue on Hot Interconnects 29. IEEE Micro, 43(2):97–98, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liu:2024:PMD

[LG24] Jianqing Liu and Na Gong.

Privacy by memory design: Visions and open problems. IEEE Micro, 44(1):49–58, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:1995:FLB

[LGJ95] H. H. Li, N. Godfrey, and Yuandong D. Ji. A fuzzy logic beam-and-ball controller prototype. IEEE Micro, 15 (6):64, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Loh:2012:SVL

[LH12] Gabriel H. Loh and Mark D.

Hill. Supporting very large DRAM caches with compound-access scheduling and MissMap. IEEE Micro, 32(3):70–78, May/June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Litz:2020:MLS

[LH20] H. Litz and M. Hashemi.

Machine learning for systems. IEEE Micro, 40(5):6– 7, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:2002:HTM

[LHC^{+}02] Jin-Fu Li, Hsin-Jung Huang, Jeng-Bin Chen, Chih-Pin Su, Cheng-Wen Wu, Chuang Cheng, Shao-I Chen, Chi-Yi Hwang, and Hsiao-Ping Lin. A hierarchical test methodology for systems on chip. IEEE Micro, 22(5):69– 81, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5069.
pdf;   http://www.computer.
org/micro/mi2002/m5069abs.
htm.

Li:2012:SPT

[LHC^{+}12] Hengjie Li, Wenting He, Yang Chen, Lieven Eeckhout, Olivier Temam, and Chengyong Wu. SWAP: Parallelization through algorithm substitution. IEEE Micro, 32(4):54– 67, July/August 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Leng:2020:EEV

[LHC^{+}20] Y. Leng, J. Huang, C. Chen, Q. Sun, and Y. Zhu. Energyefficient video processing for virtual reality. IEEE Micro, 40(3):30–36, May/June 2020.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liu:2009:MBS

[LHL09] Hsiang-Ning Liu, Yu-Jen Huang, and Jin-Fu Li. Memory built-in self test in multicore chips with mesh-based networks. IEEE Micro, 29(5): 46–55, September/October 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lentz:1999:SVU

[LHM99] Karen Panetta Lentz, Jamie Heller, and Pier Luca Montessoro. System verification using multilevel concurrent simulation. IEEE Micro, 19 (1):60–67, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1060.
pdf;   http://www.computer.
org/micro/mi1999/m1060abs.
htm.

Lee:1991:VAL

[LHMH91] Kuo Chu Lee, Takako Matoba Hickey, Victor W. Mak, and Gary E. Herman. VLSI accelerators for large database systems. IEEE Micro, 11(6): 8–20, November/December 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Louri:1995:CTP

[LHN95] Ahmed Louri, James A. Hatch, Jr., and Jongwhoa Na. A constant-time parallel sorting algorithm and its optical implementation — using optics to achieve newer and more efficient solutions to old problems. IEEE Micro, 15(3): 60–71, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lie:2023:CAD

[Lie23]  Sean Lie.  Cerebras ar-

chitecture deep dive: First look inside the hardware/ software co-design for deep learning. IEEE Micro, 43(3): 18–30, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lindeburg:1992:EER

[Lin92] M. R. Lindeburg. Electricalengineer registration. IEEE Micro, 12(4):82–84, July/ August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lin:1998:CPC

[Lin98] Jenshan Lin. Chip-package codesign for high-frequency circuits and systems. IEEE Micro, 18(4):24–32, July/ August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m4024.pdf;   http://
www.computer.org/micro/
mi1998/m4024abs.htm.

Lines:2004:AIS

[Lin04] Andrew Lines. Asynchronous interconnect for synchronous SoC design. IEEE Micro, 24(1):32–41, January/ February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1032abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/01/m1032.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/01/
m1032.pdf.

Lindenstruth:2006:EPE

[Lin06] Volker Lindenstruth. An extreme processor for an extreme experiment. IEEE Micro, 26(2):48–57, March/April 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liu:2002:RTC

[Liu02]  Huan Liu.  Routing ta-

ble compaction in ternary CAM. IEEE Micro, 22(1):58– 64, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1058abs.htm;
 http://dlib.computer.org/

                                        mi/books/mi2002/pdf/m1058.
                                        pdf.

Livesay:2023:AFF

[LJM^{+}23] Neal Livesay, Gilbert Jonatan, Evelio Mora, Kaustubh Shivdikar, Rashmi Agrawal, Ajay Joshi, José L. Abellán, John Kim, and David Kaeli. Accelerating finite field arithmetic for homomorphic encryption on GPUs. IEEE Micro, 43(5): 55–63, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lockwood:2002:GEI

[LK02] John Lockwood and Marwan Krunz. Guest Editors’ introduction: Hot Interconnects. IEEE Micro, 22(1): 8–9, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/m1008abs.htm;
                                         http://dlib.computer.org/
                                        mi/books/mi2002/pdf/m1008.
                                        pdf.

Lim:2010:TPU

[LK10] Hyesook Lim and So Yeon Kim. Tuple pruning using Bloom filters for packet classification. IEEE Micro, 30(3): 48–59, May/June 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:2024:CAB

[LKGL24] Yuke Li, Arjun Kashyap, Yanfei Guo, and Xiaoyi Lu. Compression analysis for BlueField-2/-3 data processing units: Lossy and lossless perspectives. IEEE Micro, 44 (2):8–19, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2022:EEC

[LKJ^{+}22] Juhyoung Lee, Jihoon Kim, Wooyoung Jo, Sangyeob Kim, Sangjin Kim, and Hoi-Jun Yoo. ECIM: Exponent computing in memory for an energy-efficient heterogeneous floating-point DNN training processor. IEEE Micro, 42 (1):99–107, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Luo:1992:FGS

[LKM92] Jin Luo, Christof Koch, and Bimal Mathur. Figure-ground segregation using an analog VLSI chip. IEEE Micro, 12(6):46–57, November/December 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lai:2003:PSM

[LL03] Wangyang Lai and Chin-Tau Lea. A programmable state machine architecture for packet processing. IEEE Micro, 23(4):32–42, July/ August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/04/m4032abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/04/m4032.
pdf.

Lee:1990:MLP

[LLC90] K. H. Lee, K. S. Leung, and S. M. Cheang. A microprogrammable list processor for personal computers. IEEE Micro, 10(4):50– 61, July/August 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:2016:AOB

[LLL^{+}16] Sheng Li, Hyeontaek Lim, Victor W. Lee, Jung Ho Ahn, Anuj Kalia, Michael Kaminsky, David G. Andersen, Seongil O, Sukhan Lee, and Pradeep Dubey. Achieving one billion key-value requests per second on a single server. IEEE Micro, 36(3): 94–104, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030094-
abs.html.

Lin:2009:HSC

[LLLL09] Ying-Dar Lin, Po-Ching Lin, Yuan-Cheng Lai, and Tai-Ying Liu. Hardwaresoftware codesign for highspeed signature-based virus scanning. IEEE Micro, 29(5): 56–65, September/October 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lu:2005:ITM

[LLSS05] Zhijian Lu, John Lach, Mircea R. Stan, and Kevin Skadron. Improved thermal management with reliability banking. IEEE Micro, 25(6): 40–49, November/December 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:2008:AVM

[LLT^{+}08] Eric Li, Wenlong Li, Xiaofeng Tong, Jianguo Li, Yurong Chen, Tao Wang, Patricia P. Wang, Wei Hu, Yangzhou Du, Yimin Zhang, and Yen-Kuang Chen. Accelerating video-mining applications using many small, generalpurpose cores. IEEE Micro, 28 (5):8–21, September/October 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lin:2007:SHP

[LLW^{+}07] Yuan Lin, Hyunseok Lee, Mark Woh, Yoav Harel, Scott Mahlke, Trevor Mudge, Chaitali Chakrabarti, and Krisztián Flautner. SODA: a high-performance DSP architecture for software-defined radio. IEEE Micro, 27(1):114– 123, January/February 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:2004:PDE

[LLZ^{+}04] Xiaodong Li, Zhenmin Li, Pin Zhou, Yuanyuan Zhou, Sarita V. Adve, and Sanjeev Kumar. Performancedirected energy management for storage systems. IEEE Micro, 24(6):38–49, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/06/m6038.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/06/m6038.pdf.

Lockwood:2016:IUL

[LM16] John W. Lockwood and Madhu Monga. Implementing ultra-low-latency datacenter services with programmable logic. IEEE Micro, 36(4):18– 26, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/04/mmi2016040018-
                                        abs.html.

Lunney:1983:MBL

[LMC^{+}83] David Lunney, Robert C. Morrison, Margaret M. Cetera, Richard V. Hartness, Raymond T. Mills, Alger D. Salt, and David C. Sowell. A microcomputer-based laboratory aid for visually impaired students. IEEE Micro, 3 (4):19–31, July/August 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liu:2005:EIP

[LMVP05] Jiuxing Liu, Amith Mamidala, Abhinav Vishnu, and Dhabaleswar K. Panda. Evaluating InfiniBand performance with PCI Express. IEEE Micro, 25(1):20–29, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1020.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1020.pdf.

Lange:1994:ONC

[LNK94] E. Lange, Y. Nitta, and K. Kyuma. Optical neural chips. IEEE Micro, 14(6): 29–41, November/December 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lant:2020:TFB

[LNLG20] J. Lant, J. Navaridas, M. Luján, and J. Goodacre. Toward FPGA-based HPC: Advancing interconnect technologies. IEEE Micro, 40(1):25–34, January/February 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lindholm:2008:NTU

[LNOM08] Erik Lindholm, John Nickolls, Stuart Oberman, and John Montrym. NVIDIA Tesla: a unified graphics and computing architecture. IEEE Micro, 28(2):39–55, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Le-Ngoc:1982:PSG

[LNV82] T. Le-Ngoc and L. C. Vroomen. Programming strategies in the game of Pushover. IEEE Micro, 2(3):58–68, July/September 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Le-Ngoc:1989:IPF

[LNV89] Tho Le-Ngoc and Minh Tue Vo. Implementation and performance of the Fast Hartley Transform. IEEE Micro, 9(5): 20–27, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lockwood:2003:GEI

[Loc03] John W. Lockwood. Guest Editor’s introduction: Hot Interconnects 10—thinking beyond the Internet. IEEE Micro, 23(1):8–9, January/ February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2003/

pdf/m1008.pdf;   http://
www.computer.org/micro/
mi2003/m1008abs.htm.

Louri:1991:TDO

[Lou91]  Ahmed Louri.  Three-

dimensional optical architecture and data-parallel algorithms for massively parallel computing. IEEE Micro, 11(2):24–27, 65–82, March/ April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liu:1989:HMM

[LP89] An-Chi C. Liu and Ranjani Parthasarathi. Hardware monitoring of a multiprocessor system. IEEE Micro, 9(5):44– 51, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:2021:FAQ

[LP21] He Li and Yaru Pang. FPGAaccelerated quantum computing emulation and quantum key distillation. IEEE Micro, 41(4):49–57, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:2012:AEC

[LPC12] Feng Li, Antoniu Pop, and Albert Cohen. Automatic extraction of coarse-grained data-flow threads from imperative programs. IEEE Micro, 32(4):19–31, July/August 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lakhotia:2022:AAN

[LPKP22] Kartik Lakhotia, Fabrizio Petrini, Rajgopal Kannan, and Viktor Prasanna. Accelerating Allreduce with innetwork reduction on Intel PI-UMA. IEEE Micro, 42(2):44– 52, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lazzerini:1986:PDA

[LPL86] Beatrice Lazzerini, Cosimo Antonio Prete, and Lanfranco Lopriore. A programmable debugging aid for real-time software development. IEEE Micro, 6(3):34–42, May/June 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lustig:2015:VCM

[LPM15] Daniel Lustig, Michael Pellauer, and Margaret Martonosi. Verifying correct microarchitectural enforcement of memory consistency models. IEEE Micro, 35(3):72– 82, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/03/mmi2015030072-
                                        abs.html.

Lim:2009:SDW

[LRC^{+}09] Kevin Lim, Parthasarathy Ranganathan, Jichuan Chang, Chandrakant Patel, Trevor Mudge, and Steven K. Reinhardt. Server designs for warehouse-computing environments. IEEE Micro, 29 (1):41–49, January/February 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:1996:GEI

[LS96] Ruby B. Lee and Michael D.

Smith. Guest Editors’ introduction: Media processing: a new design target. IEEE Micro, 16(4):6–9, July/August 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Litch:1998:SPC

[LS98a] Tim Litch and Jeff Slaton.

StrongARMing portable communications. IEEE Micro, 18 (2):48–55, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2048.
pdf;   http://www.computer.
org/micro/mi1998/m2048abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24– 26, 1997.

Lorch:1998:AME

[LS98b] Jacob R. Lorch and Alan Jay Smith. Apple Macintosh’s energy consumption. IEEE Micro, 18(6):54–63, November/December 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m6054.pdf;   http://
www.computer.org/micro/
mi1998/m6054abs.htm.

Lenjani:2022:SMD

[LS22] Marzieh Lenjani and Kevin Skadron. Supporting moderate data dependency, position dependency, and divergence in PIM-based accelerators. IEEE Micro, 42(1):108– 115, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Levy:2024:SIH

[LS24] Scott Levy and Whit Schonbein. Special issue on Hot Interconnects 30. IEEE Micro, 44(2):6–7, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lustig:2017:TMM

[LSBM17] Daniel Lustig, Geet Sethi, Abhishek Bhattacharjee, and Margaret Martonosi. Transistency models: Memory ordering at the hardware–OS interface. IEEE Micro, 37(3): 88–97, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.

computer.org/csdl/mags/
mi/2017/03/mmi2017030088-
abs.html.

Langguth:2015:SHC

[LSL^{+}15] Johannes Langguth, Mohammed Sourouri, Glenn Terje Lines, Scott B. Baden, and Xing Cai. Scalable heterogeneous CPU–GPU computations for unstructured tetrahedral meshes. IEEE Micro, 35(4):6–15, July/August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040006-
abs.html.

Lee:2001:EPI

[LSY01] Ruby B. Lee, Zhijie Shi, and Xiao Yang. Efficient permutation instructions for fast software cryptography. IEEE Micro, 21(6):56–69, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6056abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m6056.
pdf.

Loucks:1982:VPB

[LSZ82] Wayne M. Loucks, Martin Snelgrove, and Safwat G. Zaky. A vector processor based on one-bit microprocessors. IEEE Micro, 2(1):53– 62, January/March 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Li:1997:GEI

[LTL97] Qiang N. Li, Chuck Thacker, and Kai Li. Guest Editors’ introduction: The hottest interconnects. IEEE Micro, 17 (1):7, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lu:2007:ADA

[LTQZ07] Shan Lu, Joseph Tucek, Feng Qin, and Yuanyuan Zhou. AVIO: Detecting atomicity violations via accessinterleaving invariants. IEEE Micro, 27(1):26–35, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lunscher:1985:SSZ

[Lun85] W. Lunscher. Semaphore strategy for Z80. IEEE Micro, 5(3):4, May/June 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Luu:1990:CCR

[Luu90a] J. Luu. Comments on ‘A comparison of RISC architectures’ by R. S. Piepho and W. S. Wu. IEEE Micro, 10(2):5, March/ April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Luu:1990:TL

[Luu90b] J. Luu. On 2nd thought . . ..

IEEE Micro, 10(2):5, March/ April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:1994:VRC

[LW94] Isaac Yi-Yuan Y. Lee and Sheng-De D. Wang. A versatile ring-connected hypercube. IEEE Micro, 14(3): 60–67, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liang:2009:RVT

[LWB09] Xiaoyao Liang, Gu-Yeon Wei, and David Brooks. Revival: a variation-tolerant architecture using voltage interpolation and variable latency. IEEE Micro, 29(1):127–138, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2016:AAB

[LWC^{+}16] Yunsup Lee, Andrew Waterman, Henry Cook, Brian Zimmer, Ben Keller, Alberto Puggelli, Jaehwa Kwak, Ruzica Jevtic, Stevo Bailey, Milovan Blagojevic, Pi-Feng Chiu, Rimas Avizienis, Brian Richards, Jonathan Bachrach, David Patterson, Elad Alon, Bora Nikolic, and Krste Asanovic. An agile approach to building RISC-V microprocessors. IEEE Micro, 36(2):

8–20, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020008-
abs.html.

Lazzaro:1994:STS

[LWK94] John Lazzaro, John Wawrzynek, and Alan Kramer. Systems technologies for silicon auditory models. IEEE Micro, 14(3):7–15, May/June 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Liu:2016:NSC

[LWML16] Fangfei Liu, Hao Wu, Kenneth Mai, and Ruby B. Lee. Newcache: Secure cache architecture thwarting cache side-channel attacks. IEEE Micro, 36(5):8–16, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050008-
abs.html.

Loh:2010:PSM

[LX10] Gabriel H. Loh and Yuan Xie. Prolegomena: 3D stacked microprocessor: Are we there yet? IEEE Micro, 30(3): 60–64, May/June 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Loh:2007:PDS

[LXB07] Gabriel H. Loh, Yuan Xie, and Bryan Black. Processor design in 3D die-stacking technologies. IEEE Micro, 27(3): 31–48, May/June 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Luo:2004:NNP

[LYBZ04] Yan Luo, Jun Yang, Laxmi N.

Bhuyan, and Li Zhao. NeP-Sim: a network processor simulator with a power evaluation framework. IEEE Micro, 24(5):34–44, September/October 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/05/m5034.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/05/m5034.pdf.

Lyles:2004:GEI

[Lyl04] J. Bryan Lyles. Guest Editor’s introduction: Hot interconnects 11- solving network bottlenecks. IEEE Micro, 24(1): 8–9, January/February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1008.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/01/m1008.
htm.

Liu:2018:SAD

[LYP^{+}18] Zhenhong Liu, Amir Yazdanbakhsh, Taejoon Park, Hadi Esmaeilzadeh, and Nam Sung Kim. SiMul: An algorithmdriven approximate multiplier design for machine learning. IEEE Micro, 38(4):50– 59, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040050-
abs.html.

Leon:2018:WTE

[LZX^{+}18] Vasileios Leon, Georgios Zervakis, Sotirios Xydis, Dimitrios Soudris, and Kiamal Pekmestzi. Walking through the energy-error Pareto frontier of approximate multipliers. IEEE Micro, 38(4):40– 49, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040040-
abs.html.

Lee:2010:PCT

[LZY^{+}10] Benjamin C. Lee, Ping Zhou, Jun Yang, Youtao Zhang, Bo Zhao, Engin Ipek, Onur Mutlu, and Doug Burger. Phase-change technology and the future of main memory. IEEE Micro, 30(1):143, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Marshall:1983:FDD

[MA83] Trevor G. Marshall and John A. Attikiouzel. Floppy disk data transfer techniques. IEEE Micro, 3(6): 17–23, November/December 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Marchand:1994:DPR

[MA94] P. J. Marchand and P. Ambs.

Developing a parallel-readout optical-disk system. IEEE Micro, 14(6):20–27, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Maas:2020:TMS

[Maa20]  M. Maas.  A taxonomy

of ML for systems problems. IEEE Micro, 40(5):8– 16, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Macgregor:1984:MCC

[Mac84] D. Macgregor. The MC68020 — corrections and comparisons — reply. IEEE Micro, 4(5):3–6, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

MacKay:1993:AIE

[Mac93] Colin A. MacKay. Amalgams for improved electronics interconnection. IEEE Micro, 13 (2):46–58, March/April 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

MacKernan:1998:ALM

[Mac98] Colm MacKernan. Avoiding the legal mire. IEEE Micro, 18(3):34–42, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3034.
pdf;   http://www.computer.
org/micro/mi1998/m3034abs.
htm.

Maenner:1987:FIB

[Mae87] Reinhard Maenner. A fast integer binary logarithm of large arguments. IEEE Micro, 7(6): 41–45, November/December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Majithia:1987:NGM

[Maj87] K. Majithia. The new generation of microprocessors. IEEE Micro, 7(4):4–5, July/August 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moshovos:2018:VBD

[MAJ^{+}18] Andreas Moshovos, Jorge Albericio, Patrick Judd, Alberto Delmas Lascorz, Sayeh Sharify, Tayler Hetherington, Tor Aamodt, and Natalie Enright Jerger. Valuebased deep-learning acceleration. IEEE Micro, 38(1):41– 55, January/February 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010041-
abs.html.

Maas:2019:HAT

[MAK19] M. Maas, K. Asanovic, and J. Kubiatowicz. A hardware accelerator for tracing garbage collection. IEEE Micro, 39(3): 38–46, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Minkenberg:2006:DCS

[MAM^{+}06] Cyriel Minkenberg, François Abel, Peter Müller, Raj Krishnamurthy, Mitchell Gusat, Peter Dill, Ilias Iliadis, Ronald Luijten, B. Roe Hemenway, Richard Grzybowski, and Enrico Schiattarella. Designing a crossbar scheduler for HPC applications. IEEE Micro, 26(3):58–71, May/June 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mange:1986:C

[Man86a] D. A. Mange. Correction.

IEEE Micro, 6(3):78, May/ June 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mange:1986:HLLa

[Man86b] Daniel A. Mange. A highlevel-language programmable controller. I. A controller for structured microprogramming. IEEE Micro, 6(1):25– 41, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mange:1986:HLLb

[Man86c] Daniel A. Mange. A highlevel-language programmable controller. II microcompilation of the high-level language Micropascal. IEEE Micro, 6 (2):47–63, March/April 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mann:1992:UAM

[Man92]  Daniel Mann.  Unix and

the Am29000 microprocessor. IEEE Micro, 12(1):23–31, January/February 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mansur:2009:NNF

[Man09] Dan Mansur. A new 40nm FPGA and ASIC common platform. IEEE Micro, 29 (2):46–53, March/April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Marshall:1984:C

[Mar84] T. G. Marshall. Correction. IEEE Micro, 4(1):78, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Marx:1985:ESW

[Mar85]  Esther R. Marx.  EDIF:

the standard for workstation intercommunication. IEEE Micro, 5(5):68–75, September/October 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Marr:1986:NVM

[Mar86] J. D. Marr. Von Neumann versus microcoding. IEEE Micro, 6(3):5, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Markoff:1996:MIS

[Mar96] John Markoff. The microprocessor’s impact on society: At 25 years old, has the microprocessor fulfilled its early promise? what does it offer for the future? IEEE Micro, 16(6):54–59, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mar:1998:LMS

[Mar98] L. Mar. Letters: Microsoft supporter speaks out. IEEE Micro, 18(3):2–3, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3002.
pdf.

Martonosi:2014:ISC

[Mar14] Margaret Martonosi. 2013 International Symposium on Computer Architecture Influential Paper Award. IEEE Micro, 34(1):91–92, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Martonosi:2017:MWA

[Mar17] Margaret Martonosi. 2016 Maurice Wilkes Award Given to Timothy Sherwood. IEEE Micro, 37(2):104–105, March/ April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

https://www.computer.org/
csdl/mags/mi/2017/02/mmi2017020104.
html.

Martin:2021:BIW

[Mar21]  Bob Martin.  8 bits in

an IoT world: Legacy chips simplify advanced architecture interfaces. IEEE Micro, 41(6):109–111, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mashey:1993:HCC

[Mas93] J. R. Mashey. Hot and cool chips — Guest Editors’ introduction. IEEE Micro, 13(3): 9–10, May/June 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Maeda:2005:RTS

[MAS^{+}05] Seiji Maeda, Shigehiro Asano, Tomofumi Shimada, Koichi Awazu, and Haruyuki Tago. A real-time software platform for the cell processor. IEEE Micro, 25(5):20– 29, September/October 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mysore:2007:II

[MAS^{+}07] Shashidhar Mysore, Banit Agrawal, Navin Srivastava, Sheng-Chih Lin, Kaustav Banerjee, and Timothy Sherwood. 3D integration for introspection. IEEE Micro, 27 (1):77–83, January/February 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mashey:2021:IIC

[Mas21] John R. Mashey. Interactions, impacts, and coincidences of the first golden age of computer architecture. IEEE Micro, 41(6):131–139, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mathias:1983:CSP

[Mat83]  C. M. Mathias.  Com-

puter software protection law — technical expertise needed. IEEE Micro, 3(4):

5–6, July/August 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Morton:1985:ICT

[MAT85] Steven G. Morton, Enrique Abreu, and Fred Tse. ITT cap — toward a personal supercomputer. IEEE Micro, 5(6): 37–49, November/December 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1987:PTP

[Mat87] Richard Mateosian. Producing technical papers. IEEE Micro, 7(6):4, 89, November/December 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1988:ME

[Mat88] Richard Mateosian. Macworld Expo. IEEE Micro, 8(2):8– ??, March/April 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1989:MIW

[Mat89a] Richard Mateosian. Macintosh issues welcomed — reply. IEEE Micro, 9(5): 7, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1989:NME

[Mat89b]  Richard Mateosian.  A

new Macintosh environment. IEEE Micro, 9(4):11–12, July/August 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1990:IS

[Mat90a] Richard Mateosian. Impressive software. IEEE Micro, 10 (4):81–82, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1990:PJL

[Mat90b] Richard Mateosian. The problems of Japanese-language computing. IEEE Micro, 10(1):7–8, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1990:WCE

[Mat90c] Richard Mateosian. A writing course for engineers. IEEE Micro, 10(5):76–77, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1991:CC

[Mat91a] Richard Mateosian. Computers and creativity. IEEE Micro, 11(5):4–6, September/October 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1991:MHS

[Mat91b] Richard Mateosian. Mathematica help stack. IEEE Micro, 11(4):43, July/August 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1991:MT

[Mat91c] Richard Mateosian. Mousetrak. IEEE Micro, 11(4):43– 44, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1992:HSM

[Mat92a]  Richard Mateosian.  How

I spent my Christmas vacation. IEEE Micro, 12(1):65– 68, January/February 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1992:ID

[Mat92b]  Richard Mateosian.  In-

stant definitions. IEEE Micro, 12(5):72–73, September/October 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1992:PM

[Mat92c] Richard Mateosian. PC miscellany. IEEE Micro, 12(6): 86–87, November/December 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1993:F

[Mat93a] Richard Mateosian. Framemaker 3.0. IEEE Micro, 13(3):84– 85, May/June 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1993:MWM

[Mat93b] Richard Mateosian. Microsoft Word 5.1 for the Macintosh and Word for Windows 2.0. IEEE Micro, 13(1):79–80, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1993:MT

[Mat93c] Richard Mateosian. MKS Toolkit 4.1 for DOS. IEEE Micro, 13(1):80, January/ February 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1993:M

[Mat93d]  Richard Mateosian.  MS-

DOS 6. IEEE Micro, 13(3): 85, May/June 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1993:PW

[Mat93e]  Richard Mateosian.  Pro-

gramming Windows. IEEE Micro, 13(5):4–??, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1993:SRW

[Mat93f] Richard Mateosian. Speed Reader Windows version. IEEE Micro, 13(1):80, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1994:PP

[Mat94]  Richard Mateosian.  The

PowerPC in perspective. IEEE Micro, 14(5):7, September/October 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1995:B

[Mat95a] Richard Mateosian. Backlog. IEEE Micro, 15(5):2– ??, September/October 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1995:MRF

[Mat95b] Richard Mateosian. Micro review: Fireside reading: How is technology shaping our lives? IEEE Micro, 15(6): 74–75, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1995:MRP

[Mat95c] Richard Mateosian. Micro review: From the profound to the mundane: digital mantras and MacInTax. IEEE Micro, 15(2):79–80, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1995:MRU

[Mat95d] Richard Mateosian. Micro review: Upgrading to Windows 95 won’t be painless, but you can do it. also, how should we address Internet privacy and publishing concerns? IEEE Micro, 15(5):2– 3, 84–85, September/October 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1995:PT

[Mat95e]  Richard Mateosian.  Past

themes. IEEE Micro, 15(3): 72–74, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1996:MRB

[Mat96a]  Richard Mateosian.  Mi-

cro review: Books, books, books. IEEE Micro, 16(4):77– 79, July/August 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Brief reviews of several Java books.

Mateosian:1996:MRR

[Mat96b] Richard Mateosian. Micro review: The Road Ahead outlines Gates’ vision of the coming information highway and the critical issues it presents. IEEE Micro, 16(1):5–6, 72, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1996:MRH

[Mat96c] Richard Mateosian. Micro review: Heavy-duty publishing: the thud factor. IEEE Micro, 16(6):77, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1996:MRS

[Mat96d] Richard Mateosian. Micro review: Software that makes your life easier. IEEE Micro, 16(5):8–9, 75, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1996:MRM

[Mat96e] Richard Mateosian. Micro review: The microprocessor at 25: It’s invisible. IEEE Micro, 16(2):75–76, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1996:MRU

[Mat96f]  Richard Mateosian.  Mi-

cro review: Unlike C++, which seeks to extend C and maintain a large degree of backward compatibility, Java starts from C and rips out its most troublesome and errorprone features. IEEE Micro, 16(3):3–5, May/June 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1997:MND

[Mat97a] R. M. Mateosian. Micro news: DARPA aids Tera MTA. IEEE Micro, 17(5):5– 6, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5005.
pdf.

Mateosian:1997:MNV

[Mat97b] R. M. Mateosian. Micro news: Visiting Hot Chips IX. IEEE Micro, 17(5):5, 72, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5005.
pdf.

Mateosian:1997:MRA

[Mat97c] Richard Mateosian. Micro review: Adobe Framemaker-5.5 for Windows, Macintosh, and UNIX systems. IEEE Micro, 17(6):86–88, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m6086.
pdf.

Mateosian:1997:MRM

[Mat97d] Richard Mateosian. Micro review: MKS-Toolkit-6.1 for Windows-95 and Windows-NT. IEEE Micro, 17(6):87, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1997/pdf/m6086.
                                        pdf.

Mateosian:1997:RS

[Mat97e] Richard Mateosian. Real sips. IEEE Micro, 17(3):3–5, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1998:LF

[Mat98a] Richard Mateosian. Looking forward. IEEE Micro, 18(1): 9–??, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1998:MRV

[Mat98b] Richard Mateosian. Micro review: a vision for distributed objects; interface design analysis. IEEE Micro, 18(2):9–10, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m2009.
                                        pdf.

Mateosian:1998:MRW

[Mat98c]  Richard Mateosian.  Mi-

cro review: Working on the Web. IEEE Micro, 18(5):2– 3, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5002.
pdf.

Mateosian:1998:MRY

[Mat98d] Richard Mateosian. Micro review: Year 2000, Windows 98; help! IEEE Micro, 18(3):5–79, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3005.
pdf.

Mateosian:1998:R

[Mat98e] Richard Mateosian. Robohelp-5.5. IEEE Micro, 18(3):78– 79, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1999:MRCb

[Mat99a] Richard Mateosian. Micro review: Changes: Extreme Programming Explained; The Cathedral and The Bazaar. IEEE Micro, 19(6):84–85, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6084.
pdf.

Mateosian:1999:MRCa

[Mat99b]  Richard Mateosian.  Mi-

cro review: Creating documents. IEEE Micro, 19 (4):4, 85, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4004.
pdf.

Mateosian:1999:MRH

[Mat99c]  Richard Mateosian.  Mi-

cro review: Happy new year: MacWorld Expo; handbook of programming languages. IEEE Micro, 19(1): 4–5, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1004.
pdf.

Mateosian:1999:MRPa

[Mat99d]  Richard Mateosian.  Mi-

cro review: Pot pourri: Jini. IEEE Micro, 19(5):10– 11, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5010.
pdf.

Mateosian:1999:MRPb

[Mat99e] Richard Mateosian. Micro review: Pot pourri: Winwriters. IEEE Micro, 19(5): 11, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1999:MRW

[Mat99f] Richard Mateosian. Micro review: Words of wisdom. IEEE Micro, 19(2):5, 85, March/ April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2005.
pdf.

Mateosian:2000:MRD

[Mat00a] Richard Mateosian. Micro review: Doing it right! IEEE Micro, 20(3):4–5, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3004.
pdf.

Mateosian:2000:MRH

[Mat00b]  Richard Mateosian.  Mi-

cro review: Happy New Year. IEEE Micro, 20(1):77– 78, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1077.
pdf.

Mateosian:2000:MRI

[Mat00c]  Richard Mateosian.  Mi-

cro review: Interaction design. IEEE Micro, 20(5):5– 6, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m5005.
                                        pdf.

Mateosian:2000:MRS

[Mat00d] Richard Mateosian. Micro review: Summer cleanup. IEEE Micro, 20(4):7–9, 85, July/August 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2000/
                                        pdf/m4007.pdf.

Mateosian:2000:MRW

[Mat00e]  Richard Mateosian.  Mi-

cro review: Windows 2000. IEEE Micro, 20(2):12–13, 96, March/April 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2000/
                                        pdf/m2012.pdf.

Mateosian:2001:MRC

[Mat01a]  Richard Mateosian.  Mi-

cro review: Confronting the context of technology’s advances. IEEE Micro, 21 (1):90–92, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1090.
pdf.

Mateosian:2001:MRH

[Mat01b]  Richard Mateosian.  Mi-

cro review: Holiday reading. IEEE Micro, 21(6): 78–79, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6078abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m6078.
pdf.

Mateosian:2001:MRMa

[Mat01c]  Richard Mateosian.  Mi-

cro review: Making it work. IEEE Micro, 21(4): 70–71, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4070.
pdf;   m4070abs.htm.

Mateosian:2001:MRMb

[Mat01d] Richard Mateosian. Micro review: Managing development. IEEE Micro, 21(5):86– 87, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5086abs.htm;
 http://dlib.computer.org/

                                        mi/books/mi2001/pdf/m5086.
                                        pdf.

Mateosian:2001:MRPa

[Mat01e]  Richard Mateosian.  Mi-

cro review: Pervasive technologies. IEEE Micro, 21 (2):94–95, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m2094.
                                        pdf.

Mateosian:2001:MRPb

[Mat01f] Richard Mateosian. Micro review: Project tools. IEEE Micro, 21(3):78–79, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m3078.
                                        pdf.

Mateosian:2002:MRE

[Mat02a]  Richard Mateosian.  Mi-

cro review: Enterprise computing. IEEE Micro, 22 (3):71–72, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m3071.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m3071abs.
                                        htm.

Mateosian:2002:MRL

[Mat02b]  Richard Mateosian.  Mi-

cro review: Learning proven lessons. IEEE Micro, 22 (1):94–95, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1094abs.htm;
 http://dlib.computer.org/
mi/books/mi2002/pdf/m1094.
pdf.

Mateosian:2002:MRPb

[Mat02c] Richard Mateosian. Micro review: Personal effectiveness. IEEE Micro, 22(5):94– 96, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5094.
pdf;   http://www.computer.
org/micro/mi2002/m5094abs.
htm.

Mateosian:2002:MRPa

[Mat02d]  Richard Mateosian.  Mi-

cro review: Programming books. IEEE Micro, 22 (4):10–11, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4010.
pdf;   http://www.computer.
org/micro/mi2002/m4010abs.
htm.

Mateosian:2003:MRE

[Mat03a] Richard Mateosian. Micro review: Evolution. IEEE Micro, 23(3):70–72, May/June 2003.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2003/03/m3070.pdf.

Mateosian:2003:MRL

[Mat03b] Richard Mateosian. Micro review: Leadership annoyances. IEEE Micro, 23(1):82– 83, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m1082.
                                        pdf;   http://www.computer.
                                        org/micro/mi2003/m1082abs.
                                        htm.

Mateosian:2003:MRMa

[Mat03c] Richard Mateosian. Micro review: Managing software projects. IEEE Micro, 23 (4):11–13, July/August 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2003/04/m4011.pdf.

Mateosian:2003:MRN

[Mat03d] Richard Mateosian. Micro review: Nuts and bolts. IEEE Micro, 23(2):7, 77, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m2007.
                                        pdf.

Mateosian:2003:MRMb

[Mat03e] Richard Mateosian. Micro review: So many books, so little time. IEEE Micro, 23(5):6–7, 79–80, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2003/05/m5006.pdf.

Mateosian:2003:MRW

[Mat03f] Richard Mateosian. Micro review: Where we’re going. IEEE Micro, 23(6):6– 7, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2003/06/m6006.htm;   http:
//csdl.computer.org/dl/
mags/mi/2003/06/m6006.pdf.

Mateosian:2004:MRA

[Mat04a] Richard Mateosian. Micro review: Attacking complexity. IEEE Micro, 24(4):88, 87, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/04/m4088.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/04/m4088.pdf.

Mateosian:2004:MRB

[Mat04b] Richard Mateosian. Micro review: Back to the future. IEEE Micro, 24(3):

70–71, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/03/m3070.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/03/m3070.pdf.

Mateosian:2004:MRM

[Mat04c] Richard Mateosian. Micro review: More on old themes. IEEE Micro, 24(6):133–134, November/December 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/06/m6133.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/06/m6133.
htm.

Mateosian:2004:MRSb

[Mat04d]  Richard Mateosian.  Mi-

cro review: Seek and show. IEEE Micro, 24(5):88, 86– 87, September/October 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/05/m5088.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/05/m5088.pdf.

Mateosian:2004:MRSa

[Mat04e] Richard Mateosian. Micro review: Single sourcing Mount Fuji. IEEE Micro, 24(1):74– 75, January/February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/01/m1074.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/01/m1074.pdf.

Mateosian:2005:MRD

[Mat05a] Richard Mateosian. Micro review: Dealing with globalization. IEEE Micro, 25(3): 13–15, May/June 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2005:MRG

[Mat05b] Richard Mateosian. Micro review: Going through the database. IEEE Micro, 25 (4):80, 79, July/August 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2005:MRTb

[Mat05c] Richard Mateosian. Micro review: Thinking about history and design. IEEE Micro, 25(2):6–7, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2006.pdf;
http://csdl.computer.org/
comp/mags/mi/2005/02/m2006abs.
htm.

Mateosian:2005:MRTa

[Mat05d] Richard Mateosian. Micro review: Too much information. IEEE Micro, 25(1):98– 99, January/February 2005.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2005/01/m1098.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/mi/2005/01/m1098.
                                        htm.

Mateosian:2005:MRY

[Mat05e] Richard Mateosian. Micro review: Year-end cleanup. IEEE Micro, 25(6):82–84, November/December 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2006:MRMa

[Mat06a] Richard Mateosian. Micro review: More on old topics. IEEE Micro, 26(3):86– 87, May/June 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2006/03/m3086.pdf.

Mateosian:2006:MRO

[Mat06b]  Richard Mateosian.  Mi-

cro review: Old and new. IEEE Micro, 26(4):83–85, July/August 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2006/04/m4083.pdf.

Mateosian:2006:MRMb

[Mat06c] Richard Mateosian. Micro review: So many books. IEEE Micro, 26(5):82–83, September/October 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/05/m5082.pdf.

Mateosian:2006:MRF

[Mat06d] Richard Mateosian. Micro review: The future will soon be here. IEEE Micro, 26(1):141– 142, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/01/m1141.pdf.

Mateosian:2007:MRA

[Mat07a] Richard Mateosian. Micro review: Advice for investigators. IEEE Micro, 27(6):60–61, November/December 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/06/mmi2007060060.
pdf.

Mateosian:2007:MRE

[Mat07b] Richard Mateosian. Micro review: Economics. IEEE Micro, 27(1):128–130, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/01/m1128.pdf.

Mateosian:2007:MRL

[Mat07c] Richard Mateosian. Micro review: Looking back. IEEE Micro, 27(2):83–85, March/ April 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2007:MRT

[Mat07d] Richard Mateosian. Micro review: Thinking about technology. IEEE Micro, 27(4):59– 61, July/August 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2007/04/mmi2007040059.
                                        pdf.

Mateosian:2008:MRS

[Mat08a] Richard Mateosian. Micro review: Software development patterns. IEEE Micro, 28(5): 72, 71, September/October 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2008:MRP

[Mat08b] Richard Mateosian. Micro review: The paradigms, they are a-changin’. IEEE Micro, 28(2):72, 70–71, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://

                                        bell.computer.org/dlcomments/
                                        ;   http://csdl.computer.
                                        org/comp/mags/mi/2008/02/
                                        mmi2008020072.pdf.

Mateosian:2009:MRL

[Mat09a] Richard Mateosian. Micro review: Life and work. IEEE Micro, 29(5):66–68, September/October 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2009:MRN

[Mat09b] Richard Mateosian. Micro review: No more wishful thinking. IEEE Micro, 29(2):68– 71, March/April 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2009:MRSb

[Mat09c]  Richard Mateosian.  Mi-

cro review: Software architects. IEEE Micro, 29(3): 62–64, May/June 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2009:MRSa

[Mat09d] Richard Mateosian. Micro review: System Green. IEEE Micro, 29(1):144–147, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2009:MRT

[Mat09e] Richard Mateosian. Micro review: Twitter. IEEE Micro, 29(4):87–88, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2010:BG

[Mat10a]  Richard Mateosian.  Be-

ing geek. IEEE Micro, 30 (6):78, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2010:MRD

[Mat10b] Richard Mateosian. Micro review: Designing for discovery. IEEE Micro, 30(2):90– 92, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2010:MRM

[Mat10c] Richard Mateosian. Micro review: Miscellany. IEEE Micro, 30(4):86–88, July/August 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2010:MRT

[Mat10d] Richard Mateosian. Micro review: Technical writing. IEEE Micro, 30(1):144–147, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2011:MRE

[Mat11a] Richard Mateosian. Micro review: Effective communication. IEEE Micro, 31(5):76– 78, September/October 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2011:T

[Mat11b] Richard Mateosian. Technology. IEEE Micro, 31(2):100– 102, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2012:MRF

[Mat12a] Richard Mateosian. Micro review: Forewords by celebrities. IEEE Micro, 32(5):76– 77, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2012:MRM

[Mat12b] Richard Mateosian. Micro review: Miscellany. IEEE Micro, 32(2):61–63, March/April 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2013:MRE

[Mat13a] Richard Mateosian. Micro review: Ethics of big data. IEEE Micro, 33(2):60–61, March/ April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2013:MRU

[Mat13b] Richard Mateosian. Micro review: Unconscious meaning [review of A User’s Guide to Thought and Meaning; jackendoff, r.; 2012]. IEEE Micro, 33(3):116–118, May/June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:2013:TDT

[Mat13c] Richard Mateosian. Technical design [two books reviewed]. IEEE Micro, 33(6): 76–78, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Mateosian:2014:HBR

[Mat14] Richard Mateosian. How to [3 books reviewed]. IEEE Micro, 34(4):53–55, July/ August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040053.
html.

Mateosian:2015:FW

[Mat15a]  Richard Mateosian.  The

future of work. IEEE Micro, 35(1):54–56, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010054.
html.

Mateosian:2015:NT

[Mat15b]  Richard Mateosian.  New

tools. IEEE Micro, 35(6): 69–71, November/December 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/csdl.computer.org/csdl/
mags/mi/2015/06/mmi2015060069.
html.

Mateosian:2015:WW

[Mat15c] Richard Mateosian. Writing well. IEEE Micro, 35(3):147– 149, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030147.
html.

Mateosian:2017:RF

[Mat17]  Richard Mateosian.  Re-

sistance is futile. IEEE Micro, 37(1):74–76, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010074.
html.

Maruyama:2018:SXF

[MAT^{+}18] Takumi Maruyama, Yasunobu Akizuki, Takekazu Tabata, Kenichi Kitamura, Noriko Takagi, Hiroyuki Ishii, Shingo Watanabe, and Fumihiro Tawa. SPARC64 XII: Fujitsu’s latest 12core processor for missioncritical servers. IEEE Micro, 38(5):75–84, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050075-
abs.html.

Mateosian:2019:WM

[Mat19]  R. Mateosian.  What I

missed. IEEE Micro, 39(5): 114–116, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mattioli:2021:AGC

[Mat21a] Michael Mattioli. The Apollo guidance computer. IEEE Micro, 41(6):179–182, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mattioli:2021:PTP

[Mat21b] Michael Mattioli. PCs take a page from Xbox with Pluton. IEEE Micro, 41(5): 125–128, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mattioli:2021:RMA

[Mat21c] Michael Mattioli. Rome to Milan, AMD continues its tour of Italy. IEEE Micro, 41(4):78– 83, July/August 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mattioli:2022:MF

[Mat22] Michael Mattioli. Meet the FaM1ly. IEEE Micro, 42(3): 78–84, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

May:2012:XAX

[May12] David May. The XMOS architecture and XS1 chips. IEEE Micro, 32(6):28–37, November/December 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Maynard:2021:MCJ

[May21] Ann Marie G. Maynard. My computer journey. IEEE Micro, 41(6):160, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McKeown:1999:GEI

[MB99] Nick McKeown and Chase Bailey. Guest Editors’ introduction: The increasingly important interconnect. IEEE Micro, 19(1):12– 13, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1012.
pdf.

McNairy:2005:MDC

[MB05] Cameron McNairy and Rohit Bhatia. Montecito: a dualcore, dual-thread Itanium processor. IEEE Micro, 25(2):10– 20, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2010abs.htm;

 http://csdl.computer.org/
dl/mags/mi/2005/02/m2010.
pdf.

Mutlu:2015:IMT

[MB15] Onur Mutlu and Rich Belgard. Introducing the MICRO Test of Time Awards: Concept, process, 2014 winners, and the future. IEEE Micro, 35(2):85– 87, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020085.
html.

Mignolet:2009:MPA

[MBA^{+}09] Jean-Yves Mignolet, Rogier Baert, Thomas J. Ashby, Prabhat Avasare, Hye-On Jang, and Jae Cheol Son. MPA: Parallelizing an application onto a multicore platform made easy. IEEE Micro, 29(3):31–39, May/June 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mutlu:2016:CBM

[MBG^{+}16] Onur Mutlu, Rich Belgard, Thomas R. Gross, Norman R. Jouppi, John L. Hennessy, Steven Przybylski, Chris Rowen, Yale N. Patt, Wen-Mei W. Hwu, Stephen W. Melvin, Michael C. Shebanow, Tse-Yu Yeh, and Andy Wolfe. Common bonds: MIPS, HPS, two-level branch prediction, and compressed code RISC processor. IEEE Micro, 36 (4):70–85, July/August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040070-
abs.html.

Minnich:1995:MIN

[MBH95] R. Minnich, D. Burns, and F. Hady. The memoryintegrated network interface. IEEE Micro, 15(1):11–19, January/February 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Muralimanohar:2008:AEI

[MBJ08] Naveen Muralimanohar, Rajeev Balasubramonian, and Norman P. Jouppi. Architecting efficient interconnects for large caches with CACTI 6.0. IEEE Micro, 28(1):69–79, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Muller:1992:ASP

[MBK^{+}92] Urs A. Muller, Bernhard Baumle, Peter Kohler, Anton Gunzinger, and Walter Guggenbuhl. Achieving supercomputer performance for neural net simulation with an array of digital signal processors. IEEE Micro, 12(5):55– 65, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mukherjee:2002:ANA

[MBL^{+}02] Shubhendu S. Mukherjee, Peter Bannon, Steven Lang, Aaron Spink, and David Webb. The Alpha 21364 network architecture. IEEE Micro, 22(1):26–35, January/February 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2002/
                                        m1026abs.htm;   http://
                                        dlib.computer.org/mi/books/
                                        mi2002/pdf/m1026.pdf.

Mellichamp:1985:RTC

[MBP^{+}85] Duncan Mellichamp, Dave Bedworth, Odd Pettersen, Peter Rony, Lew Bezanson, Walter Higgins, and Granino Korn. Real-time computing and the engineering support system. IEEE Micro, 5(5):27– 35, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moreau:1992:ETL

[MBS92] Jean Pierre Moreau, Joseph Borel, and Davoud Samani. European trends in library development. IEEE Micro, 12 (4):43–53, July/August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Meixner:2008:ALC

[MBS08] Albert Meixner, Michael E. Bauer, and Daniel J. Sorin.

Argus: Low-cost, comprehensive error detection in simple cores. IEEE Micro, 28(1):52– 59, January/February 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Maier:2002:TTA

[MBSP02] Reinhard Maier, Günther Bauer, Georg Stöger, and Stefan Poledna. Timetriggered architecture: a consistent computing platform. IEEE Micro, 22(4): 36–45, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4036.
pdf;   http://www.computer.
org/micro/mi2002/m4036abs.
htm.

Mutlu:2016:MTT

[MBTS16] Onur Mutlu, Rich Belgard, Nick Tredennick, and Michael Schlansker. The 2014 MI-CRO test of time award winners: From 1978 to 1992. IEEE Micro, 36(1):60–c3, January/February 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/01/mmi2016010060.
html.

McCallum:1987:SIM

[MC87] John C. McCallum and Tat-Seng Chua. A synthetic instruction mix for evaluating microprocessor performance. IEEE Micro, 7(3): 63–80, May/June 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1990:SGE

[MC90] C. Miller and R. Crawford. Silicon Glen — the European challenge. IEEE Micro, 10 (3):7–8, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moors:1992:CCA

[MC92] Tim Moors and Antonio Cantoni. Cascading contentaddressable memories. IEEE Micro, 12(3):56–66, May/ June 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Milligan:1995:PIU

[MC95] Michael K. Milligan and Harvey G. Cragon. Processor implementations using queues. IEEE Micro, 15 (4):58–66, July/August 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moreno:1994:ASN

[MCC^{+}94] Juan M. Moreno, Francisco Castillo, Joan Cabestany, Jordi Madrenas, and Andrzej Napieralski. An analog systolic neural processing architecture. IEEE Micro, 14(3):

51–59, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McDonald:2007:TMH

[MCC^{+}07] Austen McDonald, Brian D.

Carlstrom, JaeWoong Chung, Chi Cao Minh, Hassan Chafi, Christos Kozyrakis, and Kunle Olukotun. Transactional memory: The hardwaresoftware interface. IEEE Micro, 27(1):67–76, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McDermott:2021:MMO

[McD21]  Mark McDermott.  Mo-

torola MC68332: One of the first true SoCs. IEEE Micro, 41(6):105–106, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Merritt:1985:DAE

[MCF^{+}85] Mike Merritt, Bob Cavett, Lou Fields, Dick Kaplan, Tom Lazear, Don Miller, Brice Carnahan, Dave Bray, Harry Lundgren, Mark Turnquist, and Gary Whitehouse. Desires and aspirations of the engineering support system user. IEEE Micro, 5(5):10– 17, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McGovern:1982:RCC

[McG82] M. J. McGovern. Is ROMed code copyrightable after all? IEEE Micro, 2(2):3–6, April/ June 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Murray:1994:PSV

[MCH^{+}94] Alan F. Murray, Stephen Churcher, Alister Hamilton, Andrew J. Holmes, Geoff B. Jackson, H. Martin Reekie, and Robin J. Woodburn. Pulse stream VLSI neural networks. IEEE Micro, 14(3): 29–39, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McIntosh:1985:WN

[McI85] I. McIntosh. What’s new? IEEE Micro, 5(6):88, November/December 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McKerrow:1983:MSE

[McK83] P. J. McKerrow. Microcomputers, slotcars, and education. IEEE Micro, 3(1):62–65, January/February 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McLauchlan:1987:ECI

[McL87] Derek J. McLauchlan. European cooperation in the information technology industry. IEEE Micro, 7(5):6– 9, September/October 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McLellan:1993:AAA

[McL93] E. McLellan. The Alpha AXP architecture and 21064 processor. IEEE Micro, 13(3): 36–47, May/June 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mudge:2016:IFT

[MCM^{+}16] Trevor Mudge, Frederic T.

Chong, Igor L. Markov, Resit Sendag, Joshua J. Yi, and Derek Chiou. Impact of future technologies on architecture. IEEE Micro, 36(4):48– 56, July/August 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/04/mmi2016040048.
html.

McMahan:2018:AA

[MCN^{+}18] Joseph McMahan, Michael Christensen, Lawton Nichols, Jared Roesch, Sung-Yee Guo, Ben Hardekopf, and Timothy Sherwood. An architecture for analysis. IEEE Micro, 38(3): 107–115, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

mi/2018/03/mmi2018030107-
abs.html.

Manne:2017:IYB

[MCR17] Srilatha Manne, Bryan Chin, and Steven K. Reinhardt. If you build it, will they come? IEEE Micro, 37(6):6–12, November/December 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060006-
abs.html.

Morari:2014:SSG

[MCV^{+}14] Alessandro Morari, Vito Giovanni Castellana, Oreste Villa, Antonino Tumeo, Jesse Weaver, David Haglin, Sutanay Choudhury, and John Feo. Scaling semantic graph databases in size and performance. IEEE Micro, 34 (4):16–26, July/August 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040016-
abs.html.

Moreau:2019:HSB

[MCV^{+}19] T. Moreau, T. Chen, L. Vega, J. Roesch, E. Yan, L. Zheng, J. Fromm, Z. Jiang, L. Ceze, C. Guestrin, and A. Krishnamurthy. A hardware software blueprint for flexible deep learning specialization. IEEE Micro, 39(5):8– 16, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Morris:1988:FPD

[MD88] L. R. Morris and S. A. Dyer.

Floating-point digital signalprocessing chips — the end of the supercomputer era. IEEE Micro, 8(6):86, November/December 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Medina:2020:HLP

[MD20] E. Medina and E. Dagan. Habana Labs purpose-built AI inference and training processor architectures: Scaling AI training systems using standard Ethernet with Gaudi processor. IEEE Micro, 40 (2):17–24, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Toong:1981:ACC

[mDTG81] Hoo min D. Toong and Amar Gupta. An architectural comparison of contemporary 16bit microprocessors. IEEE Micro, 1(2):26–37, April/June 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Madisetti:1995:VPE

[ME95] Vijay K. Madisetti and Thomas W. Egolf. Virtual prototyping of embedded microcontroller-based DSP systems. IEEE Micro, 15 (5):9–21, September/October 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mead:1996:SMT

[Mea96] C. A. Mead. Scaling of MOS technology. IEEE Micro, 16 (6):48, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Murray:2020:SVO

[MEB^{+}20] K. E. Murray, M. A. Elgammal, V. Betz, T. Ansell, K. Rothman, and A. Comodi. SymbiFlow and VPR: An open-source design flow for commercial and novel FP-GAs. IEEE Micro, 40(4):49– 57, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Meindl:2003:IOG

[Mei03] James D. Meindl. Interconnect opportunities for gigascale integration. IEEE Micro, 23(3):28–35, May/June 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/03/m3028abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/03/m3028.
                                        pdf.

Melamed:1987:PAU

[Mel87]  Anna S. Melamed.  Per-

formance analysis of Unixbased network file systems. IEEE Micro, 7(1):25–38, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). CSZ108.

Melear:1989:DRF

[Mel89] Charles Melear. The design of the 88000-RISC family. IEEE Micro, 9(2):26– 38, March/April 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Meyer:2004:NPA

[Mey04] Matthias Meyer. A novel processor architecture with exact tag-free pointers. IEEE Micro, 24(3):46–55, May/June 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/03/m3046.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/03/m3046.pdf.

Marsh:1985:MSQ

[MF85] Jackie Marsh and James J.

Farrell III. Motorola’s Silver Quill program. IEEE Micro, 5(3):53–57, May/June 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Molinero-Fernandez:2002:TSE

[MFM02] Pablo Molinero-Fernández and Nick McKeown. TCP switching: Exposing circuits to IP. IEEE Micro, 22 (1):82–89, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1082abs.htm;
 http://dlib.computer.org/
mi/books/mi2002/pdf/m1082.
pdf.

McKeown:2017:PMP

[MFN^{+}17] Michael McKeown, Yaosheng Fu, Tri Nguyen, Yanqi Zhou, Jonathan Balkind, Alexey Lavrov, Mohammad Shahrad, Samuel Payne, and David Wentzlaff. Piton: A manycore processor for multitenant clouds. IEEE Micro, 37(2):70– 80, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020070-
abs.html.

Miller:1988:HTT

[MG88] C. Miller and B. Gross. Hypercard — tool or toy. IEEE Micro, 8(3):92–93, May/June 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

MehdiOwrangO:1989:LDT

[MG89] M. Mehdi Owrang O. and W. Gamini Gunaratna. A logical design tool for relational databases. IEEE Micro, 9(3): 76–83, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Misoczki:2019:TPS

[MGG^{+}19] R. Misoczki, S. Gulley, V. Gopal, M. G. Dixon, H. Vrsalovic, and W. K. Feghali. Toward postquantum security for embedded cores. IEEE Micro, 39(4):17–26, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mattioli:2021:NSF

[MGP21] Michael Mattioli, Tom Garrison, and Baiju V. Patel. The next security frontier: Taking the mystery out of the supply chain. IEEE Micro, 41 (3):103–107, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Marty:2008:VH

[MH08] Michael R. Marty and Mark D.

Hill. Virtual hierarchies. IEEE Micro, 28(1):99–109, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mudge:2010:COE

[MH10] Trevor Mudge and Urs Holzle.

Challenges and opportunities for extremely energy-efficient processors. IEEE Micro, 30 (4):20–24, July/August 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hwu:2018:AAT

[mHP18] Wen mei Hwu and Sanjay Patel. Accelerator architectures: A ten-year retrospective. IEEE Micro, 38(6):56–62, November/December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08585394-abs.
html.

Manzhosov:2023:RRC

[MHP^{+}23] Evgeny Manzhosov, Adam Hastings, Meghna Pancholi, Ryan Piersma, Mohamed Tarek Ibn Ziad, and Simha Sethumadhavan. Revisiting residue codes for modern memories. IEEE Micro, 43 (4):53–61, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Masa:1994:HSA

[MHW94] Peter Masa, Klaas Hoen, and Hans Wallinga. A High-Speed analog neural processor. IEEE Micro, 14(3):40– 50, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Martin:2003:TCN

[MHW03] Milo M. K. Martin, Mark D. Hill, and David A. Wood. Token coherence: a new framework for shared-memory multiprocessors. IEEE Micro, 23(6):108–116, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6108abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/06/m6108.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6108.pdf.

Martinez:2009:DMR

[Mİ09] José F. Martínez and Engin İpek. Dynamic multicore resource management: a machine learning approach. IEEE Micro, 29(5):8–17, September/October 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1986:YN

[Mil86] A. R. Miller. 1986 — the year of networking. IEEE Micro, 6 (3):6–7, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1987:NNI

[Mil87] C. Miller. Neurocomputing — a new information-processing paradigm. IEEE Micro, 7(6): 6, 91, November/December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1988:AB

[Mil88a] C. Miller. After the breakthrough. IEEE Micro, 8 (2):82–83, March/April 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1988:CRW

[Mil88b]  C. Miller.  CISC, RISC,

WISC — what’s in a name. IEEE Micro, 8(1):6–7, January/February 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1988:RP

[Mil88c] C. Miller. The reason processor. IEEE Micro, 8(5):76– ??, September/October 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1988:WR

[Mil88d] C. Miller. Why RISC? IEEE Micro, 8(2):84–85, March/ April 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1989:QL

[Mil89] C. Miller. The quantum leap. IEEE Micro, 9(1):95–96, January/February 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Milenkovic:1990:MMM

[Mil90] Milan Milenkovic. Microprocessor memory management units. IEEE Micro, 10(2):70– 85, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McKeown:1997:TTP

[MIM^{+}97] Nick McKeown, Martin Izzard, Adisak Mekkittikul, William Ellersick, and Mark Horowitz. Tiny Tera: a packet switch core: Using new scheduling algorithms to build a 1-terabits packet switch with a central hub no larger than a can of soda. IEEE Micro, 17 (1):26–33, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mineta:1984:FP

[Min84] N. Mineta. Factors of production. IEEE Micro, 4(3): 4–6, May/June 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Misunas:1993:GEI

[Mis93] David Misunas. Guest Editor’s introduction: Advanced packaging and interconnection technology. IEEE Micro, 13(2):7–9, March/April 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moreira:2010:CFT

[MK10] Jose E. Moreira and John Karidis. The case for fullthrottle computing: An alternative datacenter design strategy. IEEE Micro, 30(4):25– 28, July/August 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Misra:2022:SIA

[MK22] Sudip Misra and Neeraj Kumar. Special issue on artificial intelligence, edge, and internet of things for smart agriculture. IEEE Micro, 42(1): 6–7, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mezzetti:2018:HIP

[MKAC18] Enrico Mezzetti, Leonidas Kosmidis, Jaume Abella, and Francisco J. Cazorla. Highintegrity performance monitoring units in automotive chips for reliable timing V&V. IEEE Micro, 38(1):56–65, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010056-
abs.html.

Mahdiani:2020:DNP

[MKG^{+}20] H. Mahdiani, A. Khadem, A. Ghanbari, M. Modarressi, F. Fattahi-Bayat, and M. Daneshtalab. Δ NN: Power-efficient neural network acceleration using differential weights. IEEE Micro, 40 (1):67–74, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Manatunga:2015:SCS

[MKM15] Dilan Manatunga, Hyesoon Kim, and Saibal Mukhopadhyay. SP-CNN: A scalable and programmable CNNbased accelerator. IEEE Micro, 35(5):42–50, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050042-
abs.html.

Maejima:1983:VCS

[MKNK83] Hideo Maejima, Koyo Katsura, Hideo Nakamura, and Toshimasa Kihara. The VLSI control structure of a CMOS microcomputer. IEEE Micro, 3(6):9–16, November/December 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miyata:1988:TBM

[MKOK88] Misao Miyata, Hidechika Kishigami, Kosei Okamoto, and Shigeo Kamiya. The TX1 32-bit microprocessor — performance analysis, and debugging support. IEEE Micro, 8 (2):37–46, March/April 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mutlu:2006:ERE

[MKP06] Onur Mutlu, Hyesoon Kim, and Yale N. Patt. Efficient runahead execution: Powerefficient memory latency tolerance. IEEE Micro, 26(1):10– 20, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Milne:1997:MDA

[MKRC97] George Milne, Ashis Khan, Simon Rayne, and Juha Christensen. Microcontroller design advantages for portable computing: Using the ARM7100 microcontroller to enable two OEM applications to reach the market quickly. IEEE Micro, 17(4):49–55, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4049.
pdf.

Miura:2013:SHM

[MKT^{+}13] Noriyuki Miura, Yusuke Koizumi, Yasuhiro Take, Hiroki Matsutani, Tadahiro Kuroda, Hideharu Amano, Ryuichi Sakamoto, Mitaro Namiki, Kimiyoshi Usami, Masaaki Kondo, and Hiroshi Nakamura. A scalable 3D heterogeneous multicore with an inductive ThruChip interface. IEEE Micro, 33(6): 6–15, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Madhusudan:2005:HAS

[ML05] Bharath Madhusudan and John W. Lockwood. A hardware-accelerated system for real-time worm detection. IEEE Micro, 25(1):60– 69, January/February 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1060.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1060.pdf.

Mattioli:2021:HPW

[ML21] M. Mattioli and A. Lahtiranta. Hidden potential within video game consoles. IEEE Micro, 41(2):72–77, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ma:2015:NPA

[MLL^{+}15] Kaisheng Ma, Xueqing Li, Shuangchen Li, Yongpan Liu, John Jack Sampson, Yuan Xie, and Vijaykrishnan Narayanan. Nonvolatile processor architecture exploration for energy-harvesting applications. IEEE Micro, 35(5):32–40, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050032-
abs.html.

Ma:2018:IIA

[MLL^{+}18] Kaisheng Ma, Jinyang Li, Xueqing Li, Yongpan Liu, Yuan Xie, Mahmut Kandemir, Jack Sampson, and Vijaykrishnan Narayanan. IAA: Incidental approximate architectures for extremely energyconstrained energy harvesting scenarios using IoT nonvolatile processors. IEEE Micro, 38(4):11–19, July/ August 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040011-
abs.html.

Murali:2020:ANI

[MLM^{+}20] P. Murali, N. M. Linke, M. Martonosi, A. J. Abhari, N. H. Nguyen, and C. H. Alderete. Architecting noisy intermediate-scale quantum computers: A real-system study. IEEE Micro, 40(3): 73–80, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ma:2016:NPA

[MLS^{+}16] Kaisheng Ma, Xueqing Li, Karthik Swaminathan, Yang Zheng, Shuangchen Li, Yongpan Liu, Yuan Xie, John Jack Sampson, and Vijaykrishnan Narayanan. Nonvolatile processor architectures: Efficient, reliable progress with unstable power. IEEE Micro, 36(3): 72–83, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

mi/2016/03/mmi2016030072-
abs.html.

MacGregor:1983:VMM

[MM83] Douglas MacGregor and David S. Mothersole. Virtual memory and the MC68010. IEEE Micro, 3(3):24–39, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McNeley:1987:ECI

[MM87] Kevin J. McNeley and Veljko M.

Milutinovic. Emulating a complex instruction set computer with a reduced instruction set computer. IEEE Micro, 7 (1):60–72, January/February 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moore:1996:IFC

[MM96] Jason Moore and Mahmoud A. Manzoul. An interactive fuzzy CAD tool: Mapping fuzzy controllers onto VLSI systolic arrays. IEEE Micro, 16(2):68–74, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Montrym:2005:G

[MM05] John Montrym and Henry Moreton. The GeForce 6800. IEEE Micro, 25(2):41–51, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.

computer.org/comp/mags/
mi/2005/02/m2041abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2005/02/m2041.
pdf.

Mutlu:2009:PAB

[MM09] Onur Mutlu and Thomas Moscibroda. Parallelismaware batch scheduling: Enabling high-performance and fair shared memory controllers. IEEE Micro, 29 (1):22–32, January/February 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mazumder:2023:RTH

[MM23] Arnab Neelim Mazumder and Tinoosh Mohsenin. Reg-TuneV2: a hardware-aware and multiobjective regressionbased fine-tuning approach for deep neural networks on embedded platforms. IEEE Micro, 43(6):74–83, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mogul:2008:UAS

[MMB^{+}08] Jeffrey C. Mogul, Jayaram Mudigonda, Nathan Binkert, Parthasarathy Ranganathan, and Vanish Talwar. Using asymmetric single-ISA CMPs to save energy on operating systems. IEEE Micro, 28(3): 26–41, May/June 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miguelez:2012:RMS

[MMB12] Ismael Gomez Miguelez, Vuk Marojevic, and Antoni Gelonch Bosch. Resource management for software-defined radio clouds. IEEE Micro, 32 (1):44–53, January/February 2012. CODEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Misra:2022:OIC

[MMC^{+}22] Pulkit A. Misra, Ioannis Manousakis, Esha Choukse, Majid Jalili, Íñigo Goiri, Ashish Raniwala, Brijesh Warrier, Husam Alissa, Bharath Ramakrishnan, Phillip Tuma, Christian Belady, Marcus Fontoura, and Ricardo Bianchini. Overclocking in immersion-cooled datacenters. IEEE Micro, 42(4):10– 17, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mutlu:2018:IMS

[MMCH18] Onur Mutlu, Scott Mahlke, Tom Conte, and Wen-Mei Hwu. Iterative modulo scheduling. IEEE Micro, 38(1):115–117, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/

mi/2018/01/mmi2018010115.
html.

Maglione-Mathey:2020:PLI

[MMESG^{+}20] G. Maglione-Mathey, J. Escudero-Sahuquillo, P. J. Garcia, F. J. Quiles, and J. Duato. Path2SL: Leveraging InfiniBand resources to reduce head-of-line blocking in fat trees. IEEE Micro, 40(1):8– 14, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Maglione-Mathey:2022:RII

[MMESGQ22] German Maglione-Mathey, Jesus Escudero-Sahuquillo, Pedro Javier Garcia, and Francisco J. Quiles. Reducing the impact of interjob interference in Dragonfly networks using virtual partitions. IEEE Micro, 42(3):50–56, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McCormack:1999:INB

[MMG^{+}99] Joel McCormack, Robert Mc-Namara, Christopher Gianos, Norman P. Jouppi, Todd Dutton, John Zurawski, Larry Seiler, and Ken Correll. Implementing Neon: a 256-bit graphics accelerator. IEEE Micro, 19(2):58–69, March/ April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2058.

pdf;   http://www.computer.
org/micro/mi1999/m2058abs.
htm.

MacGregor:1984:MM

[MMM84] Doug MacGregor, Dave Mothersole, and Bill Moyer. The Motorola MC68020. IEEE Micro, 4(4):101–118, July/ August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Markovic:2015:KUM

[MNU^{+}15] Nikola Markovic, Daniel Nemirovsky, Osman Unsal, Mateo Valero, and Adrian Cristal. Kernel-to-user-mode transition-aware hardware scheduling. IEEE Micro, 35 (4):37–47, July/August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040037-
abs.html.

Monden:1987:III

[Mon87]  Hiroshi Monden.  Intro-

duction to ITRON — the industry-oriented operating system. IEEE Micro, 7(2):45– 52, March/April 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Montague:1997:JEJ

[Mon97] Bruce R. Montague. JN:

OS for an embedded Java Network Computer: Supporting Java’s Virtual Machine on a single-chip embedded PC attached to the Internet. IEEE Micro, 17 (3):54–60, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3054.
pdf.

Moore:2003:PTM

[Moo03]  Charles Moore.  Parting

thoughts: Managing the transition from complexity to elegance: Knowing when you have a problem. IEEE Micro, 23(5):88, 86–87, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2003/05/m5088.pdf.

Moore:2004:MTC

[Moo04a] Charles Moore. Managing the transition from complexity to elegance: Design convergence. IEEE Micro, 24(1):80, 79, January/February 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/01/m1080.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/01/m1080.pdf.

Moore:2004:GIR

[Moo04b] Chuck Moore. Getting it right. IEEE Micro, 24(2):80, 79, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/02/m2080.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/02/m2080.pdf.

Morris:1984:PDD

[Mor84] L. R. Morris. Price, Derek, Desolla and the Antikythera mechanism — an appreciation 1922–1983 — obituary. IEEE Micro, 4(1):15–21, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Morris:1986:DSP

[Mor86a] L. R. Morris. Digital signalprocessing microprocessors — forward to the past. IEEE Micro, 6(6):6–8, November/December 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Morris:1986:GFS

[Mor86b] L. R. Morris. Good FFT software stretches processor performance. IEEE Micro, 6(2): 4–5, March/April 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Morris:1988:PBD

[Mor88] L. Robert Morris. A PCbased digital speech spectrograph. IEEE Micro, 8(6):

68–85, November/December 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moussouris:1996:M

[Mou96] J. Moussouris. Mediaprocessors. IEEE Micro, 16 (6):51, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

MacGregor:1985:PAM

[MR85] Doug MacGregor and Jon Rubinstein. A performance analysis of MC68020-based systems. IEEE Micro, 5(6): 50–70, November/December 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mattson:2020:MIS

[MRC^{+}20] P. Mattson, V. J. Reddi, C. Cheng, C. Coleman, G. Diamos, D. Kanter, P. Micikevicius, D. Patterson, G. Schmuelling, H. Tang, G. Wei, and C. Wu. MLPerf: An industry standard benchmark suite for machine learning performance. IEEE Micro, 40(2):8–16, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mahajan:2015:AAA

[MRJ^{+}15] Divya Mahajan, Kartik Ramkrishnan, Rudra Jariwala, Amir Yazdanbakhsh, Jongse Park, Bradley Thwaites, Anandhavel Nagendrakumar, Abbas Rahimi, Hadi Esmaeilzadeh, and Kia Bazargan. Axilog: Abstractions for approximate hardware design and reuse. IEEE Micro, 35(5):16– 30, September/October 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050016-
abs.html.

Moore:2003:GEI

[MRLB03] Charles Moore, Kevin W.

Rudd, Ruby B. Lee, and Pradip Bose. Guest Editors’ introduction: Micro’s top picks from microarchitecture conferences. IEEE Micro, 23(6):8–10, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6008.pdf;
http://csdl.computer.org/
dl/mags/mi/2003/06/m6008.
htm.

Meredith:2011:PIN

[MRSV11] Jeremy S. Meredith, Philip C.

Roth, Kyle L. Spafford, and Jeffrey S. Vetter. Performance implications of nonuniform device topologies in scalable heterogeneous architectures. IEEE Micro, 31(5):66– 75, September/October 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Muller-Schloer:1983:MBC

[MS83] C. Müller-Schloer. A microprocessorbased cryptoprocessor. IEEE Micro, 3(5):5–15, September/October 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McGill:1984:FTC

[MS84] William F. McGill and Steven E. Smith. Fault tolerance in continuous processcontrol. IEEE Micro, 4(6): 22–33, November/December 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Micheletti:1987:LCD

[MS87] Giancarlo Micheletti and Claudio Salati. A lowcost distributed architecture for telecommunication systems. IEEE Micro, 7(5):70– 82, September/October 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McNairy:2003:IPM

[MS03] Cameron McNairy and Don Soltis. Itanium 2 processor microarchitecture. IEEE Micro, 23(2):44–55, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m2044.

pdf;   http://www.computer.
org/micro/mi2003/m2044abs.
htm.

Martin:2016:TPC

[MS16] Milo Martin and Daniel Sorin.

Top picks from the 2015 computer architecture conferences. IEEE Micro, 36(3): 6–9, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030006-
abs.html.

Magklis:2003:DFV

[MSA^{+}03] Grigorios Magklis, Greg Semeraro, David H. Albonesi, Steven G. Dropsho, Sandhya Dwarkadas, and Michael L. Scott. Dynamic frequency and voltage scaling for a multipleclock-domain microprocessor. IEEE Micro, 23(6):62–68, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6062abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/06/m6062.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6062.pdf.

Maenner:1987:HPS

[MSB87] Reinhard Maenner, Richard L.

Shoemaker, and Peter H. Bartels. The Heidelberg Polyp System. IEEE Micro, 7(1): 5–13, January/February 1987.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miyamura:2017:NBF

[MSB^{+}17] Makoto Miyamura, Toshitsugu Sakamoto, Xu Bai, Yukihide Tsuji, Ayuka Morioka, Ryusuke Nebashi, Munehiro Tada, Naoki Banno, Koichiro Okamoto, Noriyuki Iguchi, Hiromitsu Hada, Tadahiko Sugibayashi, Yuya Nagamatsu, Soichi Ookubo, Takuma Shirai, Fumihito Sugai, and Masayuki Inaba. NanoBridge-based FPGA in high-temperature environments. IEEE Micro, 37(5):32– 42, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050032-
abs.html.

Mahmoud:2019:AIS

[MSP^{+}19] M. Mahmoud, D. M. Stuart, Z. Poulos, A. D. Lascorz, P. Judd, S. Sharify, M. Nikoli , K. Siu, I. E. Vivancos, J. Albericio, and A. Moshovos. Accelerating image-sensor-based deep learning applications. IEEE Micro, 39(5):26–35, September/October 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Madhavan:2015:RLA

[MSS15] Advait Madhavan, Timothy Sherwood, and Dmitri Strukov. Race logic: Abusing hardware race conditions to perform useful computation. IEEE Micro, 35(3):48– 57, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030048-
abs.html.

Mange:1985:BDB

[MST^{+}85] D. Mange, E. Sanchez, A. Thayse, Paul J. A. Zsombor-Murray, L. J. Vroomen, R. Hudson, T. Le-Ngoc, and Darryl J. Stewart. Binarydecision-based programmable controllers. IEEE Micro, 5(3): 58–72, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mutlu:2003:REE

[MSWP03] Onur Mutlu, Jared Stark, Chris Wilkerson, and Yale N. Patt. Runahead execution: An effective alternative to large instruction windows. IEEE Micro, 23(6):20–25, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6020abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/06/m6020.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2003/06/
                                        m6020.pdf.

Meng:2022:TRR

[MSY^{+}22] Jian Meng, Wonbo Shim, Li Yang, Injune Yeo, Deliang Fan, Shimeng Yu, and Jae sun Seo. Temperatureresilient RRAM-based inmemory computing for DNN inference. IEEE Micro, 42 (1):89–98, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Martinez:2003:SSP

[MT03] José F. Martínez and Josep Torrellas. Speculative synchronization: Programmability and performance for parallel codes. IEEE Micro, 23(6):126–134, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/06/m6126abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/06/m6126.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2003/06/
                                        m6126.pdf.

Marculescu:2005:EAU

[MT05] Diana Marculescu and Emil Talpes. Energy awareness and uncertainty in microarchitecture-level design. IEEE Micro, 25(5):64– 76, September/October 2005.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mars:2012:IUM

[MTS^{+}12] Jason Mars, Lingjia Tang, Kevin Skadron, Mary Lou Soffa, and Robert Hundt. Increasing utilization in modern warehouse-scale computers using bubble-up. IEEE Micro, 32(3):88–99, May/ June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mudge:2010:GEI

[Mud10] Trevor Mudge. Guest Editor’s introduction: Top picks from the computer architecture conferences of 2009. IEEE Micro, 30(1):8–11, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mudge:2015:TWE

[Mud15] Trevor Mudge. Thoughts on winning the 2014 Eckert– Mauchly Award. IEEE Micro, 35(3):144–146, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/03/mmi2015030144.
html.

Murray:1989:PAV

[Mur89] Alan F. Murray. Pulse arithmetic in VLSI neural networks. IEEE Micro, 9(6):

64–74, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Murari:2003:INC

[Mur03]  Bruno Murari.  Integrat-

ing nonelectronic components into electronic microsystems. IEEE Micro, 23(3): 36–44, May/June 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/03/m3036abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/03/m3036.
                                        pdf.

Murmann:2006:DAA

[Mur06] Boris Murmann. Digitally assisted analog circuits. IEEE Micro, 26(2):38–47, March/ April 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Moudgill:1996:PI

[MV96] Mayan Moudgill and Stamatis Vassiliadis. Precise interrupts. IEEE Micro, 16(1):58– 67, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mirhosseini:2019:QFA

[MW19] A. Mirhosseini and T. F. Wenisch. The queuingfirst approach for tail management of interactive services. IEEE Micro, 39(4):55– 64, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mukherjee:2003:MAV

[MWE^{+}03] Shubhendu S. Mukherjee, Christopher T. Weaver, Joel Emer, Steven K. Reinhardt, and Todd Austin. Measuring architectural vulnerability factors. IEEE Micro, 23(6):70–75, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6070abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/06/m6070.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6070.pdf.

Moudgill:1999:EPM

[MWM99] Mayan Moudgill, John-David D.

Wellman, and Jaime H. Moreno. Environment for PowerPC microarchitecture exploration. IEEE Micro, 19(3):15–25, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3015.
pdf;   http://www.computer.
org/micro/mi1999/m3015abs.
htm.

Mirapuri:1992:MRP

[MWV92] Sunil Mirapuri, Michael Woodacre, and Nader Vasseghi. The MIPS R4000 processor. IEEE Micro, 12(2):10–22, March/ April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips III, Stanford University, 1992.

Miki:1995:FIA

[MY95] Tsutomu Miki and Takeshi Yamakawa. Fuzzy inference on an analog fuzzy chip. IEEE Micro, 15(4):8– 18, July/August 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1981:CRS

[Mye81] W. Myers. Control robotic systems with distributed microprocessors. IEEE Micro, 1 (1):112, January/March 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1982:CS

[Mye82a] W. Myers. Compcon Spring 82. IEEE Micro, 2(2):81– 83, April/June 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1982:LNS

[Mye82b] W. Myers. Local-network standard on track. IEEE Micro, 2(4):74–76, October/December 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1982:TLN

[Mye82c] W. Myers. Toward a localnetwork standard. IEEE Micro, 2(3):28–45, July/ September 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1982:PCA

[Mye82d] Ware Myers. Personal computers aid the handicapped. IEEE Micro, 2(1):26–40, January/March 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1983:DGM

[Mye83a]  W. Myers.  Deaf to get

microprocessor-based autocuer. IEEE Micro, 3(4):84–85, July/August 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1983:LTR

[Mye83b] W. Myers. Language translator runs on as little as 64K. IEEE Micro, 3(6): 64–68, November/December 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1983:SCB

[Mye83c] W. Myers. Single-chip 32bit microprocessors arriving. IEEE Micro, 3(6): 65–??, November/December 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1984:AIE

[Mye84a] W. Myers. Active interconnections enable parallel microprocessors to manage vast relational database. IEEE Micro, 4(1):76–79, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1984:ABM

[Mye84b] W. Myers. Additional 32bit microprocessors reported. IEEE Micro, 4(1):75–81, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1984:CMC

[Mye84c]  W. Myers.  Concurrent

microprocessors are costeffective for scientific problems. IEEE Micro, 4(1):79–??, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1984:SOF

[Mye84d] W. Myers. Standards are our friends, Gordon Bell says. IEEE Micro, 4(2):71, March/ April 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1985:APC

[Mye85a] W. Myers. The AT&T Personal Computer-6300. IEEE Micro, 5(1):68–71, January/ February 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1985:PWV

[Mye85b] W. Myers. PFS — write version-B. IEEE Micro, 5 (2):92–94, March/April 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1989:DCP

[Mye89a] W. Myers. Design choices power the next wave. IEEE Micro, 9(3):96–??, May/June 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1989:RSC

[Mye89b] W. Myers. The road to the supersmart card. IEEE Micro, 9 (4):96–??, July/August 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1990:USN

[Mye90] W. Myers. Users said — no sacrifices. IEEE Micro, 10(1): 9–??, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1991:HFI

[Mye91a] W. Myers. HDTV faces intertwined challenges. IEEE Micro, 11(2):8–9, March/April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1991:NC

[Mye91b] W. Myers. News from Compcon 91. IEEE Micro, 11(3): 47–49, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1991:DY

[Mye91c] Ware Myers. The drive to the year 2000. IEEE Micro, 11(1):10–13, 68–74, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1992:DTP

[Mye92a] W. Myers. From desk-top to palmtop. IEEE Micro, 12(3): 7–??, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1992:LCD

[Mye92b] W. Myers. The limits of chip density. IEEE Micro, 12 (1):75–77, January/February 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1992:SGR

[Mye92c] W. Myers. Second-generation RISCs. IEEE Micro, 12(2):86– 87, March/April 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1993:GMF

[Mye93a] W. Myers. Get to market faster with FPGAs. IEEE Micro, 13(3):73–74, May/June 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1993:IPV

[Mye93b] W. Myers. ICs per vehicle increasing rapidly. IEEE Micro, 13(1):4–6, January/February 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Myers:1993:MVD

[Mye93c]  W. Myers.  Market vol-

ume drives neural-net technology. IEEE Micro, 13(2):3– 5, March/April 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Maruyama:2010:SVN

[MYK^{+}10] Takumi Maruyama, Toshio Yoshida, Ryuji Kan, Iwao Yamazaki, Shuji Yamamura, Noriyuki Takahashi, Mikio Hondou, and Hiroshi Okano. Sparc64 VIIIfx: a newgeneration octocore processor for petascale computing. IEEE Micro, 30(2):30– 40, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Noakes:1984:NPT

[NA84] P. D. Noakes and R. Aish.

A new peripheral for threedimensional computer input. IEEE Micro, 4(5):26– 35, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nagendra:2020:AUM

[NAA^{+}20] N. P. Nagendra, G. Ayers, D. I. August, H. K. Cho, S. Kanev, C. Kozyrakis, T. Krishnamurthy, H. Litz, T. Moseley, and P. Ranganathan. AsmDB: Understanding and mitigating frontend stalls in warehouse-scale computers. IEEE Micro, 40(3):56–63, May/June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Naithani:2022:VRI

[NAJE22] Ajeya Naithani, Sam Ainsworth, Timothy M. Jones, and Lieven Eeckhout. Vector runahead for indirect memory accesses. IEEE Micro, 42(4):116–123, July/August 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nakamura:1999:GEI

[Nak99] Tadao Nakamura. Guest Editor’s introduction: Introducing cool chips. IEEE Micro, 19(4):9–10, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi1999/pdf/m4009.
pdf.

Nakamura:2000:CCI

[Nak00] Tadao Nakamura. Cool Chips III. IEEE Micro, 20(6): 83–84, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6083.
pdf;   http://www.computer.
org/micro/mi2000/m6083abs.
htm.

Narayanan:2019:GVF

[Nar19] V. Narayanan. Going vertical: The future of electronics. IEEE Micro, 39(6):6– 7, November/December 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nepal:2006:MRP

[NBM^{+}06] Kundan Nepal, R. Iris Bahar, Joseph Mundy, William R. Patterson, and Alexander Zaslavsky. MRF reinforcer: a probabilistic element for space redundancy in nanoscale circuits. IEEE Micro, 26(5):19– 27, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nag:2018:NGT

[NBS^{+}18] Anirban Nag, Rajeev Balasubramonian, Vivek Srikumar, Ross Walker, Ali Shafiee, John Paul Strachan, and Naveen Muralimanohar. Newton: Gravitating towards the physical limits of crossbar acceleration. IEEE Micro, 38(5):41–49, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050041-
abs.html.

Neri:1986:MMB

[NC86] Giovanni Neri and Tullio Salmon Cinotti. The Modiac multiprocessor — a 286-based design. IEEE Micro, 6(1):7–15, January/ February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Normoyle:1998:UIE

[NCT^{+}98] Kevin B. Normoyle, Michael A. Csoppenszky, Allan Tzeng, Timothy P. Johnson, Christopher D. Furman, and Jamshid Mostoufi. UltraSPARC-IIi: Expanding the boundaries of a system on a chip. IEEE Micro, 18(2):14–24, March/ April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2014.
pdf;   http://www.computer.
org/micro/mi1998/m2014abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24– 26, 1997.

Nickolls:2010:GCE

[ND10] John Nickolls and William J.

Dally. The GPU computing era. IEEE Micro, 30(2):56– 69, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Neves:2022:CVE

[NDR^{+}22] Nuno Neves, Joao Mario Domingos, Nuno Roma, Pedro Tomás, and Gabriel Falcao. Compiling for vector extensions with streambased specialization. IEEE Micro, 42(5):49–58, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nelson:1984:C

[Nel84]  J. Nelson.  Correction.

IEEE Micro, 4(1):78, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nelson:1981:MBC

[NF81] Victor P. Nelson and Hugh L.

Fellows, Jr. A microcomputerbased controller for an amusement park ride. IEEE Micro, 1(3):13–22, July/September 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Neve:2003:STF

[NFQ03] Amaury Nève, Denis Flandre, and Jean-Jacques Quisquater.

SOI technology for future high-performance smart cards. IEEE Micro, 23(3):58–67, May/June 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/03/m3058abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/03/m3058.
pdf.

Naused:1987:BMG

[NG87] Barbara A. Naused and Barry K. Gilbert. A 32-bit, 200-MHz GaAs RISC for highthroughput signal-processing environments. IEEE Micro, 7(6):8–20, November/ December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nowatzki:2016:HNE

[NGS16] Tony Nowatzki, Vinay Gangadhar, and Karthikeyan Sankaralingam. A heterogeneous Von Neumann/ explicit dataflow processor. IEEE Micro, 36(3):20– 30, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030020-
abs.html.

Nowatzki:2017:DSG

[NGSW17] Tony Nowatzki, Vinay Gangadhar, Karthikeyan Sankaralingam, and Greg Wright.

Domain specialization is generally unnecessary for accelerators. IEEE Micro, 37(3): 40–50, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030040-
abs.html.

Noyce:1981:HMD

[NH81] Robert N. Noyce and Marcian E. Hoff, Jr. A history of microprocessor development at Intel. IEEE Micro, 1(1): 8–11, 13–21, January/March 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Navardi:2023:MTM

[NHMM23] Mozhgan Navardi, Edward Humes, Tejaswini Manjunath, and Tinoosh Mohsenin. MetaE2RL: Toward meta-reasoning for energyefficient multigoal reinforcement learning with squeezededge you only look once. IEEE Micro, 43(6):29–39, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nahas:2022:AIE

[NHY^{+}22] Hassan Nahas, Sean Huver, Billy Y. S. Yiu, Chris M. Kallweit, Adrian J. Y. Chee, and Alfred C. H. Yu. Artificialintelligence-enhanced ultrasound flow imaging at the edge. IEEE Micro, 42(6): 96–106, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Najjar:2014:RC

[NI14] Walid A. Najjar and Paolo Ienne. Reconfigurable computing. IEEE Micro, 34(1): 4–6, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Nicoud:1984:AMP

[Nic84] J. D. Nicoud. Advances in microcomputer peripherals — introduction. IEEE Micro, 4(5):7–8, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nicoud:1988:VRS

[Nic88] Jean-Daniel D. Nicoud. Video RAMs — structure and applications. IEEE Micro, 8(1): 8–27, January/February 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nicoud:1991:DTM

[Nic91] Jean-Daniel D. Nicoud. Dedicated tools for microprocessor education. IEEE Micro, 11(1):14–17, 62–68, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nickolls:2003:CLP

[NIJ^{+}03] John Nickolls, L. J. Madar III, Scott Johnson, Viresh Rustagi, Ken Unger, and Mustafiz Choudhury. Calisto: a low-power singlechip multiprocessor communications platform. IEEE Micro, 23(2):29–43, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m2029.
pdf;   http://www.computer.
org/micro/mi2003/m2029abs.
htm.

Najafi:2017:OTB

[NJZL^{+}17] M. Hassan Najafi, Shiva Jamali-Zavareh, David J. Lilja, Marc D. Riedel, Kia Bazargan, and Ramesh Harjani. An overview of time-based computing with stochastic constructs. IEEE Micro, 37(6):62–71, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060062-
abs.html.

Nadehara:1995:LPM

[NKDN95] Kouhei Nadehara, Ichiro Kuroda, Masayuki Daito, and Takashi Nakayama. Lowpower multimedia RISC. IEEE Micro, 15(6):20–29, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nojiri:2009:DPT

[NKI^{+}09] Tohru Nojiri, Yuki Kondo, Naohiko Irie, Masayuki Ito, Hajime Sasaki, and Hideo Maejima. Domain partitioning technology for embedded multicore processors. IEEE Micro, 29(6):7–17, November/December 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nelson:1983:MKM

[NKPC83] Peter J. Nelson, Larry Korba, Gordon Park, and David Crabtree. The MOD keyboard (modified keyboard). IEEE Micro, 3(4):7–17, July/ August 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nossal:2002:MBS

[NL02] Roman Nossal and Roland Lang. Model-based system development: An approach to building X-by-wire applications. IEEE Micro, 22 (4):56–63, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m4056.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m4056abs.
                                        htm.

Nguyen:2019:NDR

[NLM^{+}19] K. Nguyen, K. Lyu, X. Meng, V. Sridharan, and X. Jian. Nonblocking DRAM refresh. IEEE Micro, 39(3):103–109, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Naccache:1996:CSC

[NM96] David Naccache and David M’Raïhi. Cryptographic smart cards — comparing the existing cryptographydedicated microprocessors and describing possible directions for their evolution. IEEE Micro, 16(3):14, 16–24, May/ June 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nakamura:1999:PLD

[NM99] Ichiya Nakamura and Hideki Mori. Play and learning in the digital future. IEEE Micro, 19(6):36–42, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6036.
pdf;   http://www.computer.
org/micro/mi1999/m6036abs.
htm.

Nguyen:2022:ROC

[NM22] Thien Nguyen and Alexander McCaskey. Retargetable optimizing compilers for quantum accelerators via a multilevel intermediate representation. IEEE Micro, 42(5):17– 33, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nugier:2024:ACM

[NM24] Cyrius Nugier and Vincent Migliore. Acceleration of a classic McEliece postquantum cryptosystem with cache processing. IEEE Micro, 44 (1):59–68, January/February 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nesbit:2008:MRM

[NMC^{+}08] Kyle J. Nesbit, Miquel Moreto, Francisco J. Cazorla, Alex Ramirez, Mateo Valero, and James E. Smith. Multicore resource management. IEEE Micro, 28(3): 6–16, May/June 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nazari:2023:AAA

[NMF^{+}23] Najmeh Nazari, Hosein Mohammadi Makrani, Chongzhou Fang, Behnam Omidi, Setareh Rafatirad, Hossein Sayadi, Khaled N. Khasawneh, and Houman Homayoun. Adversarial attacks against machine learning-based resource provisioning systems. IEEE Micro, 43(5):35–44, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nowatzki:2015:ASC

[NMHS15] Tony Nowatzki, Jaikrishnan Menon, Chen-Han Ho, and Karthikeyan Sankaralingam. Architectural simulators considered harmful. IEEE Micro, 35(6):4–12, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060004-
abs.html.

Nemirovsky:2015:RHC

[NMU^{+}15] Daniel Nemirovsky, Nikola Markovic, Osman Unsal, Mateo Valero, and Adrian Cristal. Reimagining heterogeneous computing: A functional instruction-set architecture computing model. IEEE Micro, 35(5):6–14, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050006-
abs.html.

Nagarakatte:2013:HEC

[NMZ13] Santosh Nagarakatte, Milo M. K. Martin, and Steve Zdancewic. Hardware-enforced comprehensive memory safety. IEEE Micro, 33(3):38–47, May/June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nagle:1981:DFI

[NN81a] H. Troy Nagle, Jr. and Victor P. Nelson. Digital filter implementation on 16-bit microcomputers. IEEE Micro, 1(1): 23–41, January/March 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nelson:1981:DFP

[NN81b] V. P. Nelson and H. T. Nagle, Jr. Digital filtering performance comparison of 16-bit microcomputers. IEEE Micro, 1(1):32–41, January/March 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Naffziger:2014:HC

[NN14] Samuel Naffziger and Donald Newell. Hot Chips 25. IEEE Micro, 34(2):4–5, March/ April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Neusser:1993:NLV

[NNS^{+}93] Stefan Neusser, Jos Nijhuis, Lambert Spaanenburg, Bernd Hoefflinger, Uwe Franke, and Hans Fritz. Neurocontrol for lateral vehicle guidance. IEEE Micro, 13(1):57–66, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Noh:2019:TEF

[Noh19] S. H. Noh. Has the time for EMT finally come? IEEE Micro, 39(1):67–68, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Noyce:1985:AIN

[Noy85] R. N. Noyce. Action on imports needed now. IEEE Micro, 5(5):85, September/October 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Narayanasamy:2006:BRA

[NPC06] Satish Narayanasamy, Gilles Pokam, and Brad Calder. BugNet: Recording applicationlevel execution for deterministic replay debugging. IEEE Micro, 26(1):100–109, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Norrie:2021:DPG

[NPY^{+}21] T. Norrie, N. Patil, D. H.

Yoon, G. Kurian, S. Li, J. Laudon, C. Young, N. Jouppi, and D. Patterson. The design process for Google’s training chips: TPUv2 and TPUv3. IEEE Micro, 41 (2):56–63, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Neelakantam:2008:HAE

[NRS^{+}08] Naveen Neelakantam, Ravi Rajwar, Suresh Srinivas, Uma Srinivasan, and Craig Zilles. Hardware atomicity: An effective abstraction for reliable software speculation. IEEE Micro, 28(1):21–31, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Neto:2006:UBB

[NRV^{+}06] Egas Henes Neto, Ivandro Ribeiro, Michele Vieira, Gilson Wirth, and Fernanda Lima Kastensmidt. Using bulk built-in current sensors to detect soft errors. IEEE Micro, 26(5):10– 18, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Newman:1981:MDM

[NS81] Mike Newman and David Wayne Smith. The MC6809 in DMA mode on the IEEE-488 bus. IEEE Micro, 1(4):56–66, 68– 75, October/December 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nesbit:2005:DCP

[NS05] Kyle J. Nesbit and James E. Smith. Data cache prefetching using a global history buffer. IEEE Micro, 25(1):90– 97, January/February 2005.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1090.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1090.pdf.

Naffziger:2015:HC

[NS15] Samuel Naffziger and Guri Sohi. Hot Chips 26. IEEE Micro, 35(2):4–5, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/02/mmi2015020004-
abs.html.

Nakamura:1993:FIF

[NSN^{+}93] Kazuo Nakamura, Narumi Sakashita, Yasuhiko Nitta, Ken’ichi Shimomura, and Takeshi Tokuda. Fuzzy inference and a fuzzy inference processor. IEEE Micro, 13(5):37– 48, September/October 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nanomura:1997:MDI

[NST97a] Yasuhiro Nanomura, Toru Shimizu, and Osamu Tomisawa. M32R/D—integrating DRAM and microprocessor. IEEE Micro, 17(6):40–48, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m6040.

                                        pdf;   http://www.computer.
                                        org/micro/mi1997/m6040abs.
                                        htm.

Nunomura:1997:MDI

[NST97b] Y. Nunomura, T. Shimizu, and O. Tomisawa. M32R/D— integrating DRAM and microprocessor. IEEE Micro, 17(6): 40–48, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nicoud:1989:TTI

[NT89] Jean-Daniel D. Nicoud and Andrew Martin Tyrrell. The transputer-T414 instruction set. IEEE Micro, 9(3):60– 75, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nunez-Yanez:2022:FAD

[NY22] Jose Nunez-Yanez. Fused architecture for dense and sparse matrix processing in Tensor-Flow Lite. IEEE Micro, 42(6): 55–66, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Orlando:1981:OMF

[OA81] Richard V. Orlando and Thomas L. Anderson. An overview of the 9900 microprocessor family. IEEE Micro, 1(3):38–44, July/September 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oehler:1991:IRS

[OB91] Richard R. Oehler and Michael W.

Blasgen. IBM RISC System/ 6000: architecture and performance. IEEE Micro, 11(3): 14–17, 56–62, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Owens:2007:RCC

[ODH^{+}07] John D. Owens, William J.

Dally, Ron Ho, D. N. (Jay) Jayasimha, Stephen W. Keckler, and Li-Shiuan Peh. Research challenges for onchip interconnection networks. IEEE Micro, 27(5): 96–108, September/October 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Olmedilla:2021:DLI

[OESGG^{+}21] C. Olmedilla, J. Escudero-Sahuquillo, P. J. Garcia-Garcia, F. Alfaro-Cortés, J. L. Sánchez, F. J. Quiles, W. Sun, X. Yu, Y. Xu, and J. Duato. DVL-Lossy: Isolating congesting flows to optimize packet dropping in lossy data-center networks. IEEE Micro, 41 (1):37–44, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ozaki:1988:SFT

[OFG88] Brenda M. Ozaki, Eduardo B.

Fernandez, and Ehud Gudes. Software fault tolerance in architectures with hierarchical protection levels. IEEE Micro, 8(4):30–43, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oleksenko:2023:RTB

[OFKS23] Oleksii Oleksenko, Christof Fetzer, Boris Köpf, and Mark Silberstein. Revizor: Testing black-box CPUs against speculation contracts. IEEE Micro, 43(4):37–44, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oberman:1999:ATA

[OFW99] Stuart Oberman, Greg Favor, and Fred Weber. AMD 3DNow! technology: Architecture and implementations. IEEE Micro, 19(2): 37–48, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2037.
pdf;   http://www.computer.
org/micro/mi1999/m2037abs.
htm.

OConnor:2001:IAP

[OG01] Mike O’Connor and Christopher A. Gomez. The iFlow address processor. IEEE Micro, 21(2):16–23, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2016.

pdf;   http://www.computer.
org/micro/mi2001/m2016abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Oliveira:2024:CNC

[OG24] Rafael Oliveira and Ada Gavrilovska. Comprex: Innetwork compression for accelerating IoT analytics at scale. IEEE Micro, 44(2):20– 30, March/April 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oliveira:2022:ANN

[OGLG^{+}22] Geraldo F. Oliveira, Juan Gómez-Luna, Saugata Ghose, Amirali Boroumand, and Onur Mutlu. Accelerating neural network inference with Processing-in-DRAM: From the edge to the cloud. IEEE Micro, 42(6):25–38, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Olukotun:1994:SHC

[OHLR94] Kunle A. Olukotun, Rachid Helaihel, Jeremy Levitt, and Ricardo Ramirez. A softwarehardware cosynthesis approach to digital system simulation. IEEE Micro, 14(4):48– 58, July/August 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oh:2012:LPR

[OKH^{+}12] Jinwook Oh, Gyeonghoon Kim, Injoon Hong, Junyoung Park, Seungjin Lee, Joo-Young Kim, Jeong-Ho Woo, and Hoi-Jun Yoo. Low-power, real-time object-recognition processors for mobile vision systems. IEEE Micro, 32(6): 38–50, November/December 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Otani:2011:PMC

[OKN^{+}11] Sugako Otani, Hiroyuki Kondo, Itaru Nonomura, Toshihiro Hanawa, Shin’ichi Miura, and Taisuke Boku. Peach: a multicore communication system on chip with PCI express. IEEE Micro, 31(6): 39–50, November/December 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

OGrady:1985:PSM

[OL85] E. Pearse O’Grady and Raul Lozano. A performance study of mutual exclusion/ synchronization mechanisms in an IEEE 796 bus multiprocessor. IEEE Micro, 5(4):32– 47, July/August 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ollivier:2023:SAP

[OLT^{+}23] Sébastien Ollivier, Sheng Li, Yue Tang, Stephen Cahoon, Ryan Caginalp, Chayanika Chaudhuri, Peipei Zhou, Xulong Tang, Jingtong Hu, and Alex K. Jones. Sustainable AI processing at the edge. IEEE Micro, 43(1):19– 28, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ogras:2007:CPR

[OML^{+}07] Umit Y. Ogras, Radu Marculescu, Hyung Gyu Lee, Puru Choudhary, Diana Marculescu, Michael Kaufman, and Peter Nelson. Challenges and promising results in NoC prototyping using FP-GAs. IEEE Micro, 27(5):86– 95, September/October 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ophir:2013:SPM

[OMMB13] Noam Ophir, Christopher Mineo, David Mountain, and Keren Bergman. Silicon photonic microring links for highbandwidth-density, low-power chip I/O. IEEE Micro, 33 (1):54–67, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ondrusch:1996:TAP

[Ond96] S. Ondrusch. Toward available personal security tokens. IEEE Micro, 16(3):15, May/ June 1996. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic).

Oswald:2023:HAS

[ONS^{+}23] Nicolai Oswald, Vijay Nagarajan, Daniel J. Sorin, Vasilis Gavrielatos, Theo X. Olausson, and Reece Carr. HeteroGen: Automatic synthesis of heterogeneous cache coherence protocols. IEEE Micro, 43(4):62–70, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oka:1999:DPE

[OS99] Masaaki Oka and Masakazu Suzuoki. Designing and programming the Emotion Engine. IEEE Micro, 19(6): 20–28, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m6020.
                                        pdf.

Owen:2008:NAA

[OS08] Jonathan Owen and Maurice Steinman. Northbridge architecture of AMD’s Griffin microprocessor family. IEEE Micro, 28(2):10–18, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oshio:2024:CSN

[OSS^{+}24] Reon Oshio, Takuya Sugahara, Atsushi Sawada, Mutsumi Kimura, Renyuan Zhang, and Yasuhiko Nakashima. A compressed spiking neural network onto a memcapacitive in-memory computing array. IEEE Micro, 44(1):8–16, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

OConnor:1997:PJV

[OT97] J. Michael O’Connor and Marc Tremblay. Picojava-I — the Java Virtual Machine in hardware. IEEE Micro, 17 (2):45–53, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Okada:1982:RPP

[OTM82] Y. Okada, H. Tajima, and R. Mori. A reconfigurable parallel processor with microprogram control. IEEE Micro, 2 (4):48–60, October/December 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Omnes:1990:GEI

[OVT90] Jean-François Omnes, Thierry Van der Pyl, and Philip Treleaven. Guest Editors’ introduction: Parallel computing in Europe. IEEE Micro, 10(6):8–10, November/December 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Opris:2001:FAF

[OW01] Ion E. Opris and Seiichiro Watanabe. A fast analog front-end processor for digital imaging systems. IEEE Micro, 21(2):48–54, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2048.
pdf;   http://www.computer.
org/micro/mi2001/m2048abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Ohkubo:1987:CCK

[OWK87] Toshikazu Ohkubo, Tetsuo Wasano, and Ichizo Kogiku. Configuration of the Ctron kernel. IEEE Micro, 7(2):33– 44, March/April 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ozdal:2017:GAA

[OYK^{+}17] Muhammet Mustafa Ozdal, Serif Yesil, Taemin Kim, Andrey Ayupov, John Greth, Steven Burns, and Ozcan Ozturk. Graph analytics accelerators for cognitive systems. IEEE Micro, 37(1):42– 51, January/February 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010042-
abs.html.

Ozaki:2011:CMA

[OYS^{+}11] Nobuaki Ozaki, Yoshihiro Yasuda, Yoshiki Saito, Daisuke Ikebuchi, Masayuki Kimura, Hideharu Amano, Hiroshi Nakamura, Kimiyoshi Usami, Mitaro Namiki, and Masaaki Kondo. Cool mega-arrays: Ultralow-power reconfigurable accelerator chips. IEEE Micro, 31(6):6–18, November/December 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ollivier:2022:PRB

[OZT^{+}22] Sébastien Ollivier, Xinyi Zhang, Yue Tang, Chayanika Choudhuri, Jingtong Hu, and Alex K. Jones. Pod-racing: bulk-bitwise to floating-point compute in racetrack memory for machine learning at the edge. IEEE Micro, 42(6): 9–16, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Patterson:1997:CIR

[PAC^{+}97] David Patterson, Thomas Anderson, Neal Cardwell, Richard Fromm, Kimberly Keeton, Christoforos Kozyrakis, Randi Thomas, and Katherine Yelick. A case for intelligent RAM. IEEE Micro, 17 (2):34–44, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Polig:2014:GTA

[PAC^{+}14] Raphael Polig, Kubilay Atasu, Laura Chiticariu, Christoph Hagleitner, H. Peter Hofstee, Frederick R. Reiss, Huaiyu Zhu, and Eva Sitaridi. Giving text analytics a boost. IEEE Micro, 34(4):6–14, July/August 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/04/mmi2014040006-
                                        abs.html.

Palamara:1982:RCR

[Pal82] P. Palamara. And is ROMed code really different from source code. IEEE Micro, 2 (2):6, April/June 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Palmquist:1993:ICC

[Pal93] Ulf Palmquist. Intelligent cruise control and roadside information. IEEE Micro, 13 (1):20–28, January/February 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pullini:2007:BNN

[PAM^{+}07] Antonio Pullini, Federico Angiolini, Srinivasan Murali, David Atienza, Giovanni De Micheli, and Luca Benini. Bringing NoCs to 65 nm. IEEE Micro, 27(5):75– 85, September/October 2007.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Papazoglou:1989:EDS

[Pap89] Mike P. Papazoglou. An extensible DBMS for small and medium systems. IEEE Micro, 9(2):52–68, March/April 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Papworth:1996:TPP

[Pap96] David B. Papworth. Tuning the Pentium Pro microarchitecture: Refining a design from the initial goals, performance simulations, trade-offs, and dies to the final product. IEEE Micro, 16(2):8– 15, March/April 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Parhami:2000:LRM

[Par00] Behrooz Parhami. Letters:

Reinventing the mousetrap. IEEE Micro, 20(2):4, March/ April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2004.
pdf.

Paterson:1984:DFB

[Pat84] T. Paterson. Data format and the S-100 bus. IEEE Micro, 4(1):11–14, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Paterson:1990:TL

[Pat90] T. Paterson. On 2nd thought . . .. IEEE Micro, 10(2): 5, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Paulino:2021:BTF

[PBFC21] Nuno Paulino, João Bispo, João C. Ferreira, and João M. P. Cardoso. A binary translation framework for automated hardware generation. IEEE Micro, 41 (4):15–23, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Perez:2006:SMF

[PBT06] Daniel Gracia Pérez, Hugues Berry, and Olivier Temam. A sampling method focusing on practicality. IEEE Micro, 26(6):14–28, November/December 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:1993:ICG

[PC93] D. Price and J. H. Crawford. Intel chief gauges chip trends. IEEE Micro, 13 (6):104, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Provan:2001:MBF

[PC01] Gregory Provan and Yi-Liang Chen. Model-based fault-tolerant control reconfiguration for general network topologies. IEEE Micro, 21(5):64–76, September/October 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
m5064abs.htm;   http://
dlib.computer.org/mi/books/
mi2001/pdf/m5064.pdf.

Putnam:2015:RFA

[PCC^{+}15] Andrew Putnam, Adrian M.

Caulfield, Eric S. Chung, Derek Chiou, Kypros Constantinides, John Demme, Hadi Esmaeilzadeh, Jeremy Fowers, Gopi Prashanth Gopal, Jan Gray, Michael Haselman, Scott Hauck, Stephen Heil, Amir Hormati, Joo-Young Kim, Sitaram Lanka, James Larus, Eric Peterson, Simon Pope, Aaron Smith, Jason Thong, Phillip Yi Xiao, and Doug Burger. A reconfigurable fabric for accelerating large-scale datacenter services. IEEE Micro, 35(3): 10–22, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/

                                        mi/2015/03/mmi2015030010-
                                        abs.html.

Pistol:2010:AIN

[PCDL10] Constantin Pistol, Wutichai Chongchitmate, Christopher Dwyer, and Alvin R. Lebeck. Architectural implications of nanoscale-integrated sensing and computing. IEEE Micro, 30(1):110–120, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Petrini:2002:QNH

[PcFH^{+}02] Fabrizio Petrini, Wu chun Feng, Adolfy Hoisie, Salvador Coll, and Eitan Frachtenberg. The Quadrics network: Highperformance clustering technology. IEEE Micro, 22 (1):46–57, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/m1046abs.htm;
                                         http://dlib.computer.org/
                                        mi/books/mi2002/pdf/m1046.
                                        pdf.

Poovey:2009:BCE

[PCLGO09] Jason A. Poovey, Thomas M. Conte, Markus Levy, and Shay Gal-On. A benchmark characterization of the EEMBC Benchmark Suite. IEEE Micro, 29(5):18– 29, September/October 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pelley:2015:MPS

[PCW15] Steven Pelley, Peter M. Chen, and Thomas F. Wenisch. Memory persistency: Semantics for byte-addressable nonvolatile memory technologies. IEEE Micro, 35(3):125– 131, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030125-
abs.html.

Peh:2001:DMR

[PD01] Li-Shiuan Peh and William J.

Dally. A delay model for router microarchitectures. IEEE Micro, 21(1):26– 34, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1026.
pdf;   http://www.computer.
org/micro/mi2001/m1026abs.
htm.

Pistol:2008:NOC

[PDL08] Constantin Pistol, Chris Dwyer, and Alvin R. Lebeck. Nanoscale optical computing using resonance energy transfer logic. IEEE Micro, 28(6): 7–18, November/December 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pinckney:2013:LPB

[PDS^{+}13] Nathaniel Pinckney, Ronald G.

Dreslinski, Korey Sewell, David Fick, Trevor Mudge, Dennis Sylvester, and David Blaauw. Limits of parallelism and boosting in dim silicon. IEEE Micro, 33(5):30– 37, September/October 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Poulton:1998:TCR

[PDT98] J. Poulton, W. J. Dally, and S. Tell. A tracking clock recovery receiver for 4-Gbps signaling. IEEE Micro, 18(1):25–27, January/February 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pease:1995:TFL

[Pea95] R. A. Pease. 3rd thoughts on fuzzy-logic. IEEE Micro, 15 (4):78–80, July/August 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Peels:1987:DDS

[Pee87] Arno J. H. M. Peels. Designing digital systems — SSI and MSI vs. LSI and VLSI. IEEE Micro, 7(2):66–80, March/ April 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pennello:1990:CCR

[Pen90] Thomas J. Pennello. Compiler challenges with RISCs. IEEE Micro, 10(1):37–43, January/February 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pentland:1999:WCN

[Pen99] Alex (Sandy) Pentland. Wearable computing NEW!: Wearable computers. IEEE Micro, 19(6):9–11, November/December 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1999/
pdf/m6009.pdf.

Pentland:2001:GEI

[Pen01] Alex (Sandy) Pentland. Guest Editor’s introduction: Wearable information devices. IEEE Micro, 21(3):12–15, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
pdf/m3012.pdf.

Perez:1983:BWC

[Per83] Aram Perez. Byte-wise CRC calculations. IEEE Micro, 3(3):40–50, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Petajan:1992:DVC

[Pet92] Eric Petajan. Digital video coding techniques for US highdefinition TV. IEEE Micro, 12(5):13–21, September/October 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pal:2019:OMG

[PEZ^{+}19] S. Pal, E. Ebrahimi, A. Zulfiqar, Y. Fu, V. Zhang, S. Migacz, D. Nellans, and P. Gupta. Optimizing multi-GPU parallelization strategies for deep learning training. IEEE Micro, 39(5): 91–101, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pfaffenberger:1994:IPE

[Pfa94] B. Pfaffenberger. Internet in plain English. IEEE Micro, 14(6):6, November/December 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Perkowski:2002:LHUa

[PFC^{+}02a] Marek Perkowski, David Foote, Qihong Chen, Anas Al-Rabadi, and Lech Jozwiak. Learning hardware using multiple-valued logic, Part 1: Introduction and approach. IEEE Micro, 22 (3):41–51, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m3041.
pdf;   http://www.computer.
org/micro/mi2002/m3041abs.
htm.

Perkowski:2002:LHUb

[PFC^{+}02b] Marek Perkowski, David Foote, Qihong Chen, Anas Al-Rabadi, and Lech Jozwiak. Learning hardware using multiple-valued logic, Part 2: Cube calculus and architecture. IEEE Micro, 22 (3):52–61, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m3052.
pdf;   http://www.computer.
org/micro/mi2002/m3052abs.
htm.

Prete:1997:CTT

[PGL97] Cosimo Antonio Prete, Marco Graziano, and Francesco Lazzarini. The ChARM tool for tuning embedded systems: Selecting and tuning system configurations to meet cost, performance, and power consumption requirements. IEEE Micro, 17 (4):67–76, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4067.
pdf.

Petrisko:2020:BAO

[PGW^{+}20] D. Petrisko, F. Gilani, M. Wyse, D. C. Jung, S. Davidson, P. Gao, C. Zhao, Z. Azad, S. Canakci, B. Veluri, T. Guarino, A. Joshi, M. Oskin, and M. B. Taylor. Black-Parrot: An agile open-source RISC-V multicore for accelerator SoCs. IEEE Micro, 40 (4):93–102, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

PeytonJones:1991:FIS

[PH91] Simon L. Peyton Jones and Mark S. Hardie. A Futurebus interface from off-the-shelf parts. IEEE Micro, 11(1):38– 41, 84–93, January/February 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pichai:2015:ATT

[PHB15] Bharath Pichai, Lisa Hsu, and Abhishek Bhattacharjee. Address translation for throughput-oriented accelerators. IEEE Micro, 35(3):102– 113, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/03/mmi2015030102-
                                        abs.html.

Pedrycz:1995:RFN

[PHC95] Witold Pedrycz, C. Hart Poskar, and Peter J. Czezowski. A reconfigurable fuzzy neural network with in-situ learning. IEEE Micro, 15(4):19– 30, July/August 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Phillips:1985:ZM

[Phi85] David Phillips. The Z80000 microprocessor. IEEE Micro, 5(6):23–36, November/December 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Piroumian:1997:ISJ

[Pir97] Vartan Piroumian. Internationalization support in Java: Making software globalization easy. IEEE Micro, 17 (3):20–29, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3020.
pdf.

Pittman:1991:ISR

[Pit91] T. Pittman. IEEE and the Software Rental Act. IEEE Micro, 11(3):2, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pittman:1995:MVR

[Pit95] Tom Pittman. Micro view:

The RISC penalty. IEEE Micro, 15(6):5, 76–80, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pittman:1996:RPD

[Pit96a] T. Pittman. RISC proponents in denial. IEEE Micro, 16 (2):2, March/April 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pittman:1996:RVC

[Pit96b] T. Pittman. RISC versus CISC — reply. IEEE Micro, 16(1):2, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pollard:1991:AEM

[PJ91] L. Howard Pollard and Ramiro Jordan. An Advanced Educational Microprocessor System. IEEE Micro, 11(1):22–25, 78–79, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pugsley:2014:CIN

[PJB^{+}14] Seth H. Pugsley, Jeffrey Jestes, Rajeev Balasubramonian, Vijayalakshmi Srinivasan, Alper Buyuktosunoglu, Al Davis, and Feifei Li. Comparing implementations of near-data computing with inmemory MapReduce workloads. IEEE Micro, 34(4):44– 52, July/August 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/04/mmi2014040044-
abs.html.

Purkiser:1988:IFE

[PK88] Clif Purkiser and Jim Kardach. The Intel 376 family for embedded processor applications. IEEE Micro, 8(3): 10–26, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Papazian:2015:IBS

[PKB^{+}15] Irma Esmer Papazian, Sailesh Kottapalli, Jeff Baxter, Jeff Chamberlain, Geetha Vedaraman, and Brian Morris. Ivy Bridge server: A converged design. IEEE Micro, 35(2):16– 25, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020016-
abs.html.

Palframan:2013:RHP

[PKL13] David J. Palframan, Nam Sung Kim, and Mikko H. Lipasti. Resilient high-performance processors with spare RIBs. IEEE Micro, 33(4):26–34, July/August 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Passas:2015:CIO

[PKP15] Giorgos Passas, Manolis Katevenis, and Dionisios Pnevmatikatos. The combined input-output queued crossbar architecture for highradix on-chip switches. IEEE Micro, 35(6):38–47, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060038-
abs.html.

Patel:1992:DAS

[PKR92] Parimal A. Patel, Hemal N. Kothari, and James F. Robb. Development of an ASIC set for signal processing. IEEE Micro, 12(6):24–33, November/December 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Petrini:2006:GEI

[PLB06] Fabrizio Petrini, Olav Lysne, and Ron Brightwell. Guest Editors’ introduction: Highperformance interconnects. IEEE Micro, 26(3):7–9, May/ June 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2006/03/m3007.pdf.

Petracca:2009:PNS

[PLBC09] Michele Petracca, Benjamin G. Lee, Keren Bergman, and Luca P. Carloni. Photonic NoCs: System-level design exploration. IEEE Micro, 29 (4):74–85, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pannuto:2016:MSI

[PLK^{+}16] Pat Pannuto, Yoonmyung Lee, Ye-Sheng Kuo, ZhiYoong Foo, Benjamin Kempke, Gyouho Kim, Ronald G. Dreslinski, David Blaauw, and Prabal Dutta. MBus: A system integration bus for the modular microscale computing class. IEEE Micro, 36(3): 60–70, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/03/mmi2016030060-
abs.html.

Patt:2011:TP

[PM11] Yale N. Patt and Onur Mutlu.

Top picks. IEEE Micro, 31(1): 6–10, January/February 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pangracious:2015:DOH

[PMM15] Vinod Pangracious, Zied Marrakchi, and Habib Mehrez. Design and optimization of a horizontally partitioned, highspeed, 3D tree-based FPGA. IEEE Micro, 35(6):48–59, November/December 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060048-
abs.html.

Ponzina:2022:HSC

[PMR^{+}22] Flavio Ponzina, Simone Machetti, Marco Rios, Benoît Walter Denkinger, Alexandre Levisse, Giovanni Ansaloni, Miguel Peón-Quirós, and David Atienza. A hardware/software co-design vision for deep learning at the edge. IEEE Micro, 42(6): 48–54, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pereira:2023:DTM

[PMS23] Eduardo S. Pereira, Leonardo S. Marcondes, and Josemar M. Silva. On-device tiny machine learning for anomaly detection based on the extreme values theory. IEEE Micro, 43(6): 58–65, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Patel:2008:AA

[PmWH08] Sanjay Patel and Wen mei W. Hwu. Accelerator architectures. IEEE Micro, 28(4): 4–12, July/August 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2008/04/mmi2008040004.
pdf.

Papaefstathiou:2004:GEI

[PNDG04] Ioannis Papaefstathiou, Nikos A. Nikolaou, Bharat Doshi, and Eric Grosse. Guest Editors’ introduction: Network processors for future highend systems and applications. IEEE Micro, 24(5):7– 9, September/October 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/05/m5007.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/05/m5007.
htm.

Petrov:2004:TBC

[PO04] Peter Petrov and Alex Orailoglu.

Transforming binary code for low-power embedded processors. IEEE Micro, 24(3):21– 33, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/03/m3021.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/03/m3021.pdf.

Powell:1994:DFT

[Pow94] David Powell. Distributed fault-tolerance — lessons from Delta-4. IEEE Micro, 14 (1):36–47, January/February 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Patterson:1982:ARH

[PP82] D. A. Patterson and R. S.

Piepho. Assessing RISCs in high level language support. IEEE Micro, 2(4):9–19, October/December 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Privat:1992:PHR

[PP92] Gilles Privat and Eric Petajan. Processing hardware for real-time video coding — introduction. IEEE Micro, 12 (5):9–12, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Parashar:2014:ESP

[PPA^{+}14] Angshuman Parashar, Michael Pellauer, Michael Adler, Bushra Ahsan, Neal Crago, Daniel Lustig, Vladimir Pavlov, Antonia Zhai, Mohit Gambhir, Aamer Jaleel, Randy Allmon, Rachid Rayess, Stephen Maresh, and Joel Emer. Efficient spatial processing element control via triggered instructions. IEEE Micro, 34 (3):120–137, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Pan:2003:AFA

[PPBS03] Rong Pan, Balaji Prabhakar, Lee Breslau, and Scott Shenker. Approximate fair allocation of link bandwidth. IEEE Micro, 23 (1):36–43, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m1036.
                                        pdf;   http://www.computer.
                                        org/micro/mi2003/m1036abs.
                                        htm.

Papaefstathiou:2004:PHN

[PPO^{+}04] Ioannis Papaefstathiou, Stylianos Perissakis, Theofanis G. Orphanoudakis, Nikos A. Nikolaou, George Kornaros, Nicholas A. Zervos, George Konstantoulakis, Dionisios N. Pnevmatikatos, and Kyriakos Vlachos. PRO3: a hybrid NPU architecture. IEEE Micro, 24(5):20–33, September/October 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/05/m5020.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/05/m5020.pdf.

Psounis:2001:AFD

[PPP01] Konstantinos Psounis, Rong Pan, and Balaji Prabhakar. Approximate fair dropping for variable-length packets. IEEE Micro, 21(1):48– 56, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1048.
pdf;   http://www.computer.
org/micro/mi2001/m1048abs.
htm.

Prete:1991:RCM

[Pre91] Cosimo A. Prete. RST cache memory design for a tightly coupled multiprocessor system. IEEE Micro, 11(2):16– 19, 40–52, March/April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Polfliet:2011:AFS

[PRE11] Stijn Polfliet, Frederick Ryckbosch, and Lieven Eeckhout. Automated full-system power characterization. IEEE Micro, 31(3):46–59, May/June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Preethichandra:2021:ZMS

[Pre21] D. M. G. Preethichandra. Z80: The 1970s microprocessor still alive. IEEE Micro, 41(6):156–157, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Prital:1986:VSB

[Pri86] S. Prital. The VME subsystem bus. IEEE Micro, 6 (2):66–71, March/April 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:1989:BT

[Pri89] Walter J. Price. A benchmark tutorial. IEEE Micro, 9(5): 28–43, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Priem:1990:DGG

[Pri90] Curtis R. Priem. Developing the GX graphics accelerator architecture. IEEE Micro, 10 (1):44–54, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:1993:BPP

[Pri93a] D. Price. Berkeley’s Patterson predicts exciting times. IEEE Micro, 13(4):4–??, July/August 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:1993:CC

[Pri93b] D. Price. Cows and computers. IEEE Micro, 13(6): 95–97, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:1994:MVC

[Pri94a] D. Price. Micro view: Clipper: soon a de facto standard? IEEE Micro, 14(4):80, 79, July/August 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:1994:SSM

[Pri94b] D. Price. A surprising source of material — organic compounds — may lead to devices one-hundredth the size of current silicon chips. IEEE Micro, 14(3):80–??, May/June 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:1995:PFF

[Pri95] Dick Price. Pentium FDIV flaw — lessons learned. IEEE Micro, 15(2):88, 86– 87, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://

ieeexplore.ieee.org/iel1/
40/8521/00372360.pdf.

Papamichalis:1988:TFP

[PS88] Panos Papamichalis and Ray Simar, Jr. The TMS320C30 floating-point digital signal processor. IEEE Micro, 8(6): 13–29, November/December 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Panigrahy:2003:SSU

[PS03] Rina Panigrahy and Samar Sharma. Sorting and searching using ternary CAMs. IEEE Micro, 23(1):44–53, January/February 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2003/
pdf/m1044.pdf;   http://
www.computer.org/micro/
mi2003/m1044abs.htm.

Perais:2015:ETP

[PS15] Arthur Perais and Andre Seznec. EOLE: Toward a practical implementation of value prediction. IEEE Micro, 35(3):114–124, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//www.computer.org/csdl/
mags/mi/2015/03/mmi2015030114-
abs.html.

Patterson:2020:CP

[PS20] D. A. Patterson and Y. S.

Shao. Commercial products. IEEE Micro, 40(6):49, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pellegrini:2020:ANN

[PSB^{+}20] A. Pellegrini, N. Stephens, M. Bruce, Y. Ishii, J. Pusdesris, A. Raja, C. Abernathy, J. Koppanalil, T. Ringe, A. Tummala, J. Jalal, M. Werkheiser, and A. Kona. The Arm Neoverse N1 platform: Building blocks for the next-gen cloud-to-edge infrastructure SoC. IEEE Micro, 40(2):53– 62, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Parmar:2023:EMO

[PSL^{+}23] Vivek Parmar, Syed Shakib Sarwar, Ziyun Li, Hsien-Hsin S. Lee, Barbara De Salvo, and Manan Suri. Exploring memory-oriented design optimization of edge AI hardware for extended reality applications. IEEE Micro, 43(6): 40–49, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Porter:2014:HSD

[PSP14] George Porter, Alex C. Snoeren, and George Papen. High-speed datacenter interconnects [Guest Editors’ introduction]. IEEE Micro, 34(5):6–7, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/05/mmi2014050006-
                                        abs.html.

Popescu:1991:MA

[PSS^{+}91] Val Popescu, Merle Schultz, John Spracklen, Gary Gibson, Bruce Lightner, and David Isaman. The Metaflow architecture. IEEE Micro, 11(3): 10–13, 63–73, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Peterson:1991:IML

[PSW91] Craig Peterson, James Sutton, and Paul Wiley. Iwarp — a 100-Mops, LIW microprocessor for multicomputers. IEEE Micro, 11(3):26–29, 81– 87, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pflanz:1998:GRE

[PV98] Matthias Pflanz and Heinrich Theodor Vierhaus. Generating reliable embedded processors. IEEE Micro, 18(5): 33–41, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5033.
pdf;   http://www.computer.
org/micro/mi1998/m5033abs.
htm.

Pflanz:2001:OCR

[PV01] Matthias Pflanz and Heinrich Theodor Vierhaus. Online check and recovery techniques for dependable embedded processors. IEEE Micro, 21(5): 24–40, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5024abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m5024.

pdf. See corrections [Ano01a].

Peltenburg:2020:TOS

[PVB^{+}20] J. Peltenburg, J. Van Straten, M. Brobbel, Z. Al-Ars, and H. P. Hofstee. Tydi: An open specification for complex data structures over hardware streams. IEEE Micro, 40(4): 120–130, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Papa:2011:PSC

[PVS^{+}11] David Papa, Natarajan Viswanathan, Cliff Sze, Zhuo Li, Gi-Joon Nam, Charles Alpert, and Igor L. Markov. Physical synthesis with clock-network optimization for large systems on chips. IEEE Micro, 31(4):51– 62, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Putic:2017:HTM

[PVS17] Mateja Putic, A. J. Varshneya, and Mircea R. Stan. Hierarchical temporal memory on the automata processor. IEEE Micro, 37(1):52–59, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2017/01/mmi2017010052-
                                        abs.html.

Potter:1994:RDC

[PVYU94] T. Potter, M. Vaden, J. Young, and N. Ullah. Resolution of data and control-flow dependencies in the PowerPC 601. IEEE Micro, 14(5):18– 29, September/October 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Piepho:1989:CRA

[PW89] Richard S. Piepho and William S. Wu. A comparison of RISC architectures. IEEE Micro, 9 (4):51–62, July/August 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Peleg:1996:MTE

[PW96] Alex Peleg and Uri Weiser. MMX technology extension to the Intel architecture — improving multimedia and communications application performance by 1.5 to 2 times. IEEE Micro, 16(4):42– 50, July/August 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Perlmutter:1987:A

[PY87] D. Perlmutter and A. K. W.

Yuen. The 80387 and its applications. IEEE Micro, 7(4):42– 57, July/August 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pentapati:2019:LMP

[PZB^{+}19] S. Pentapati, L. Zhu, L. Bamberg, D. E. Shim, A. García-Ortiz, and S. K. Lim. A logic-on-memory processorsystem design with monolithic 3-D technology. IEEE Micro, 39(6):38–45, November/December 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Prabhakar:2018:PRA

[PZK^{+}18] Raghu Prabhakar, Yaqi Zhang, David Koeplinger, Matt Feldman, Tian Zhao, Stefan Hadjis, Ardavan Pedram, Christos Kozyrakis, and Kunle Olukotun. Plasticine: A reconfigurable accelerator for parallel patterns. IEEE Micro, 38(3): 20–31, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030020-
abs.html.

Park:2006:MPA

[PZL06] Yong-Joon Park, Zhao Zhang, and Gyungho Lee. Microarchitectural protection against stack-based buffer overflow attacks. IEEE Micro, 26(4):62– 71, July/August 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Qureshi:2008:SDC

[QJP^{+}08] Moinuddin K. Qureshi, Aamer Jaleel, Yale N. Patt, Simon C. Steely, Jr., and Joel Emer. Set-dueling-controlled adaptive insertion for highperformance caching. IEEE Micro, 28(1):91–98, January/February 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Qian:2015:VTD

[QLLG15] Chenxiong Qian, Xiapu Luo, Yu Le, and Guofei Gu. Vul-Hunter: Toward discovering vulnerabilities in Android applications. IEEE Micro, 35(1):44–53, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010044-
abs.html.

Qureshi:2021:QCD

[QT21] Moinuddin Qureshi and Swamit Tannu. Quantum computing and the design of the ultimate accelerator. IEEE Micro, 41 (5):8–14, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Quach:2000:HAR

[Qua00] Nhon Quach. High-availability and reliability in the Itanium processor. IEEE Micro, 20(5): 61–69, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5061.
pdf;   http://www.computer.
org/micro/mi2000/m5061abs.
htm.

Qureshi:2019:NMC

[Qur19] M. Qureshi. With new memories come new challenges. IEEE Micro, 39(1):52–53, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rodrigo:2021:DFS

[RAA^{+}21] Santiago Rodrigo, Sergi Abadal, Eduard Alarcón, Medina Bandic, Hans van Someren, and Carmen G. Almudéver. On double full-stack communicationenabled architectures for multicore quantum computers. IEEE Micro, 41(5):48– 56, September/October 2021.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Raghavendra:1984:FTR

[Rag84] C. S. Raghavendra. Fault tolerance in regular network architectures. IEEE Micro, 4(6): 44–53, November/December 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Riera:2019:CCG

[RAG19] M. Riera, J. Arnau, and A. González. CGPA: Coarsegrained pruning of activations for energy-efficient RNN inference. IEEE Micro, 39(5):36– 45, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Raghunathan:2021:HMF

[Rag21] K. Raghu Raghunathan. History of microcontrollers: First 50 years. IEEE Micro, 41(6): 97–104, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rajbenbach:1994:EBC

[Raj94] H. Rajbenbach. The exotic becomes the credible. IEEE Micro, 14(6):8–9, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Randall:1997:TMP

[Ran97] Martin Randall. Talisman — multimedia for the PC. IEEE Micro, 17(2):11–19, March/ April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Riesebos:2021:UGB

[RBB21] Leon Riesebos, Brad Bondurant, and Kenneth R. Brown. Universal graph-based scheduling for quantum systems. IEEE Micro, 41(5):57– 65, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rupley:2019:SMP

[RBGZ19] J. Rupley, B. Burgess, B. Grayson, and G. D. Zuraski. Samsung M3 processor. IEEE Micro, 39 (2):37–44, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rohr:2011:MGD

[RBKL11] David Rohr, Matthias Bach, Matthias Kretz, and Volker Lindenstruth. Multi-GPU DGEMM and high performance Linpack on highly energy-efficient clusters. IEEE Micro, 31(5):18–27, September/October 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ranganathan:2012:RDD

[RC12] Parthasarathy Ranganathan and Jichuan Chang. (Re)Designing data-centric data centers. IEEE Micro, 32(1):66–70, January/February 2012. CO-DEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rogenmoser:2013:RTV

[RC13] Robert Rogenmoser and Lawrence T. Clark. Reducing transistor variability for high performance low power chips. IEEE Micro, 33(2):18–26, March/ April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reis:2007:AIL

[RCA07] George A. Reis, Jonathan Chang, and David I. August. Automatic instructionlevel software-only recovery. IEEE Micro, 27(1):36–47, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reed:2000:ANA

[RCBL00] Benjamin C. Reed, Edward G. Chron, Randal C. Burns, and Darrell D. E. Long. Authenticating network-attached storage. IEEE Micro, 20 (1):49–57, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi2000/pdf/m1049.
pdf;   http://www.computer.
org/micro/mi2000/m1049abs.
htm.

Radhakrishnan:2007:BNC

[RCC07] Sivakumar Radhakrishnan, Sundaram Chinthamani, and Kai Cheng. The Blackford Northbridge chipset for the Intel 5000. IEEE Micro, 27 (2):22–33, March/April 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reda:2012:APC

[RCC12] Sherief Reda, Ryan Cochran, and Ayse K. Coskun. Adaptive power capping for servers with multithreaded workloads. IEEE Micro, 32(5):64– 75, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ramirez:2010:SA

[RCJ^{+}10] Alex Ramirez, Felipe Cabarcas, Ben Juurlink, Mauricio Alvarez Mesa, Friman Sanchez, Arnaldo Azevedo, Cor Meenderinck, Catalin Ciobanu, Sebastian Isaza, and Georgi Gaydadjiev. The SARC architecture. IEEE Micro, 30(5):16–29, September/October 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reddi:2021:VBM

[RCK^{+}21] Vijay Janapa Reddi, Christine Cheng, David Kanter, Peter Mattson, Guenther Schmuelling, and Carole-Jean Wu. The vision behind MLPerf: Understanding AI inference performance. IEEE Micro, 41(3):10–18, May/ June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rajagopal:2004:DSE

[RCR04] Sridhar Rajagopal, Joseph R. Cavallaro, and Scott Rixner. Design space exploration for real-time embedded stream processors. IEEE Micro, 24 (4):54–66, July/August 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2004/04/m4054abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2004/04/m4054.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2004/04/
                                        m4054.pdf.

Rounce:1990:AWE

[RD90] Peter A. Rounce and Jose Delgado. Architectures within the Esprit Span Project. IEEE Micro, 10(6):24–27, 88–97, November/December 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rettberg:1998:GEI

[RDC98] R. Rettberg, W. J. Dally, and D. E. Culler. Guest Editors’ introduction — the bleeding edge. IEEE Micro, 18(1):10– 11, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ruhl:2013:IPW

[RDJ^{+}13] Gregory Ruhl, Saurabh Dighe, Shailendra Jain, Surhud Khare, and Sriram R. Vangal. IA-32 processor with a widevoltage-operating range in 32nm CMOS. IEEE Micro, 33 (2):28–36, March/April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Renau:2011:HC

[RE11] Jose Renau and Will Eatherton. Hot Chips 22. IEEE Micro, 31(2):4–5, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ready:1986:VRT

[Rea86] James F. Ready. VRTX: a real-time operating system for embedded microprocessor applications. IEEE Micro, 6(4): 8–17, July/August 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reddi:2013:RAM

[Red13] Vijay Janapa Reddi. Reliabilityaware microarchitecture design [guest editor’s introduction]. IEEE Micro, 33(4):4– 5, July/August 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Regula:1992:PSS

[Reg92] Jack Regula. The proposed SSBLT standard doubles the VME64 transfer rate. IEEE Micro, 12(2):64–71, March/ April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Raghavan:2013:UDS

[RES^{+}13] Arun Raghavan, Laurel Emurian, Lei Shao, Marios Papaefthymiou, Kevin P. Pipe, Thomas F. Wenisch, and Milo M. K. Martin. Utilizing dark silicon to save energy with computational sprinting. IEEE Micro, 33(5):20–28, September/October 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Riedel:1986:GFS

[RFGM86] N. K. Riedel, C. Fisher, N. B. Goldstein, and D. A. McAninch. Good FFT software stretches processor performance — reply. IEEE Micro, 6(2):5, March/April 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Radhakrishnan:1985:CDC

[RG85] Thiruvengadam Radhakrishnan and Clifford P. Grossner. Cuenet — a distributed computing facility. IEEE Micro, 5(1):42–52, January/February 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ramabadran:1988:TCC

[RG88] Tenkasi V. Ramabadran and Sunil S. Gaitonde. A tutorial on CRC computations. IEEE Micro, 8(4):62– 75, July/August 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rajwar:2003:TET

[RG03] Ravi Rajwar and James Goodman. Transactional execution: Toward reliable, highperformance multithreading. IEEE Micro, 23(6):117–125, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6117abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/06/m6117.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6117.pdf.

Ronen:2007:GEI

[RG07] Ronny Ronen and Antonio González. Guest Editors’ introduction: Micro’s top picks from the Microarchitecture Conferences. IEEE Micro, 27(1):8–11, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2007/01/m1008.pdf.

Ruiz:1995:FCO

[RGF95] A. Ruiz, J. Gutierrez, and J. A. Felipe Fernandez. A fuzzy controller with an optimized defuzzification algorithm. IEEE Micro, 15 (6):67, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). See correction [RGF96].

Ruiz:1996:CFC

[RGF96] A. Ruiz, J. Gutierrez, and J. A. F. Fernandez. Correction: A fuzzy controller with an optimized defuzzification algorithm (vol. 15, p. 67, 1995). IEEE Micro, 16(1):2, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). See [RGF95].

Reddi:2010:PVD

[RGH^{+}10] Vijay Janapa Reddi, Meeta Gupta, Glenn Holloway, Michael D. Smith, Gu-Yeon Wei, and David Brooks. Predicting voltage droops using recurring program and microarchitectural event activity. IEEE Micro, 30(1):110, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ramrakhyani:2019:SPI

[RGK19] A. Ramrakhyani, P. V. Gratz, and T. Krishna. Synchronized Progress in Interconnection Networks (SPIN): A new theory for deadlock freedom. IEEE Micro, 39(3):110– 117, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ruping:1995:CSO

[RGR95] Stefan Ruping, Karl Goser, and Ulrich Ruckert. A chip for self-organizing feature maps. IEEE Micro, 15(3): 57–59, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Russell:1991:CVM

[RH91] David W. Russell and Kirtley B. Haden. A configurable, virtual microprocessor system for instructional use in real-time, real-world studies. IEEE Micro, 11(1):26–29, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Richardson:2003:IMS

[RHH^{+}03] Nick Richardson, Lun Bin Huang, Razak Hossain, Julian Lewis, Tommy Zounes, and Naresh Soni. The iCore 520-MHz synthesizable CPU core. IEEE Micro, 23(3):46– 57, May/June 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/03/m3046abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/03/m3046.
pdf.

Ritchie:1997:SPJ

[Rit97] Stuart Ritchie. Systems programming in Java: Reducing complexity and shortening debugging time by using Java at the system level. IEEE Micro, 17(3):30–35, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3030.
pdf.

Rony:1991:MB

[RJ91] P. R. Rony and R. C. Jaeger. A magazine is born. IEEE Micro, 11(1):7–8, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rossetto:1989:AVS

[RJHK89] Olivier Rossetto, Christian Jutten, Jeanny Herault, and Ingo Kreuzer. Analog VLSI synaptic matrices as building blocks for neural networks. IEEE Micro, 9(6):

56–63, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rowen:1988:MRF

[RJR88] Chris Rowen, Mark Johnson, and Paul Ries. The MIPS R3010 floating-point coprocessor. IEEE Micro, 8(3):53– 62, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reddi:2016:IT

[RK16] Vijay Janapa Reddi and Hyesoon Kim. On the Internet of Things. IEEE Micro, 36(6):5– 7, November/December 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060005-
abs.html.

Roy:2020:HTN

[RKA^{+}20] S. Roy, A. Kaushik, R. Agrawal, J. Gergen, W. Rouwet, and J. Arends. A high-throughput network processor architecture for latency-critical applications. IEEE Micro, 40 (1):50–56, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reddi:2011:VNP

[RKK^{+}11] Vijay Janapa Reddi, Svilen Kanev, Wonyoung Kim, Simone Campanoni, Michael D.

Smith, Gu-Yeon Wei, and David Brooks. Voltage noise in production processors. IEEE Micro, 31(1):20– 28, January/February 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Raghavan:2013:DRC

[RLC^{+}13] Arun Raghavan, Yixin Luo, Anuj Chandawalla, Marios Papaefthymiou, Kevin P. Pipe, Thomas F. Wenisch, and Milo M. K. Martin. Designing for responsiveness with computational sprinting. IEEE Micro, 33(3):8– 15, May/June 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reichel:1994:UOS

[RLG94] F. Reichel, W. Loeffler, and E. Gaertner. Using optical space-frequency analysis for real-time pattern recognition. IEEE Micro, 14(6): 49–60, November/December 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Romanescu:2011:ATA

[RLS11] Bogdan F. Romanescu, Alvin R. Lebeck, and Daniel J. Sorin. Address translation aware memory consistency. IEEE Micro, 31(1):109–118, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rose:1985:FTM

[RLV85] Jonathan Rose, Wayne Loucks, and Zvonko Vranesic. Fermtor — a tunable multiprocessor architecture. IEEE Micro, 5 (4):5–17, July/August 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reddi:2023:SIT

[RM23] Vijay Janapa Reddi and Boris Murmann. Special issue on TinyML. IEEE Micro, 43(6): 7–10, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ryan:1981:ILN

[RMBK81] Robert Ryan, George D. Marshall, Robert Beach, and Steven R. Kerman. Intel Local Network Architecture. IEEE Micro, 1(4):26–41, October/December 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rusu:2004:IPH

[RMC04] Stefan Rusu, Harry Muljono, and Brian Cherkauer. Itanium 2 processor 6M: Higher frequency and larger L3 cache. IEEE Micro, 24(2):10–18, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.

computer.org/comp/mags/
mi/2004/02/m2010abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/02/m2010.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/
m2010.pdf.

Riedel:1985:SPI

[RMFG85] Neal K. Riedel, David A. McAninch, Cameron Fisher, and Nahum B. Goldstein. A signal processing implementation for an IBM-PC-based workstation. IEEE Micro, 5(5):52–67, September/October 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Regnier:2004:EEI

[RMM^{+}04] Greg Regnier, Dave Minturn, Gary McAlpine, Vikram A. Saletore, and Annie Foong. ETA: Experience with an Intel Xeon processor as a packet processing engine. IEEE Micro, 24(1):24–31, January/February 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1024abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/01/m1024.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/01/
m1024.pdf.

Rotem:2012:PMA

[RNA^{+}12] Efraim Rotem, Alon Naveh, Avinash Ananthakrishnan, Eliezer Weissmann, and Doron Rajwan. Power-management architecture of the Intel microarchitecture code-named Sandy Bridge. IEEE Micro, 32 (2):20–27, March/April 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ravichandran:2023:PAA

[RNLY23] Joseph Ravichandran, Weon Taek Na, Jay Lang, and Mengjia Yan. PACMAN: Attacking ARM pointer authentication with speculative execution. IEEE Micro, 43(4):11– 18, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Roca:2016:EBI

[RNN^{+}16] Damian Roca, Daniel Nemirovsky, Mario Nemirovsky, Rodolfo Milito, and Mateo Valero. Emergent behaviors in the Internet of Things: The ultimate ultralarge-scale system. IEEE Micro, 36(6):36–44, November/December 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/06/mmi2016060036-
abs.html.

Rogers:2013:CCT

[ROA13] Timothy G. Rogers, Mike O’Connor, and Tor M.

Aamodt. Cache-conscious thread scheduling for massively multithreaded processors. IEEE Micro, 33(3):78– 85, May/June 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Roberts:1991:RPE

[Rob91] Charles E. Roberts. A RISC processor for embedded applications within an ASIC. IEEE Micro, 11(5):20–23, 68– 72, September/October 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Robinson:1992:PAM

[Rob92] Ian N. Robinson. Patternaddressable memory. IEEE Micro, 12(3):20–30, May/ June 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Robinson:1997:MSAa

[Rob97a]  G. S. Robinson.  Micro

standards: Active projects list. IEEE Micro, 17(2):70– 75, March/April 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Robinson:1997:MSAb

[Rob97b] G. S. Robinson. Micro standards: ANSI’s role in developing standards. IEEE Micro, 17(6):84–85, November/December 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1997/
pdf/m6084.pdf.

Robinson:1997:MSH

[Rob97c] G. S. Robinson. Micro standards: How does the IEEE do it? IEEE Micro, 17 (4):5–6, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4005.
pdf.

Robinson:1997:MSY

[Rob97d] G. S. Robinson. Micro standards: So, you want fasttrack standards. IEEE Micro, 17(3):71–72, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3071.
pdf.

Robinson:1997:MST

[Rob97e] G. S. Robinson. Micro standards: Truth and spurious urban myths. IEEE Micro, 17 (5):9–10, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5009.
pdf.

Robinson:1998:JP

[Rob98a] G. S. Robinson. Java and the PAS process. IEEE Micro, 18(1):4–5, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Robinson:1998:MSE

[Rob98b]  G. S. Robinson.  Micro

standards: Exploiting the acronym craze. IEEE Micro, 18(6):8, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6008.
pdf.

Robinson:1998:MSSb

[Rob98c] G. S. Robinson. Micro standards: Standards give you control. IEEE Micro, 18(5): 6, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5006.
pdf.

Robinson:1998:GEI

[Rob98d] Gary S. Robinson. Guest Editor’s introduction: Standards and the market. IEEE Micro, 18(3):16–17, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3016.
pdf.

Robinson:1998:MSSa

[Rob98e] Gary S. Robinson. Micro standards: Starting an international standard. IEEE Micro, 18(4):2–3, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4002.
pdf.

Robinson:1999:MSH

[Rob99a]  G. S. Robinson.  Micro

standards: Healthcare needs standards. IEEE Micro, 19(3):5, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3005.
pdf.

Robinson:1999:MSD

[Rob99b] Gary S. Robinson. Micro standards: a different look at standards. IEEE Micro, 19 (6):4–5, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6004.
pdf.

Robinson:1999:MSI

[Rob99c] Gary S. Robinson. Micro standards: IEEE Registration Authority Committee. IEEE Micro, 19(2):3–4, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2003.
pdf.

Robinson:1999:MSS

[Rob99d] Gary S. Robinson. Micro standards: Skiing and standards don’t mix! IEEE Micro, 19 (5):3–4, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5003.
pdf.

Robinson:1999:MSL

[Rob99e] Gary S. Robinson. Micro standards: The little committee that grew. IEEE Micro, 19 (1):80–81, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1080.
pdf.

Robinson:1999:MSW

[Rob99f]  Gary S. Robinson.  Mi-

cro standards: When is two too many? IEEE Micro, 19(4):3, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4003.
pdf.

Robinson:2000:MSF

[Rob00a] Gary S. Robinson. Micro standards: Formal SDOs: They’re still alive! IEEE Micro, 20(5):84, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5084.
pdf;   http://www.computer.
org/micro/mi2000/m5084abs.
htm.

Robinson:2000:MSJ

[Rob00b] Gary S. Robinson. Micro standards: Join a standards group and see the world. IEEE Micro, 20(3):85–86, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3085.
pdf.

Robinson:2000:MSM

[Rob00c] Gary S. Robinson. Micro standards: Making standards simple. IEEE Micro, 20 (6):8–9, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6008.
pdf.

Robinson:2000:MSS

[Rob00d]  Gary S. Robinson.  Mi-

cro standards: Standards intellectual property licensing. IEEE Micro, 20(1):6– 8, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1006.
pdf.

Robinson:2000:MSG

[Rob00e] Gary S. Robinson. Micro standards: The good, the bad, and the ugly. IEEE Micro, 20(2):3–4, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2003.
pdf.

Robinson:2001:MSC

[Rob01a] Gary S. Robinson. Micro standards: Centipedes, SDOs, and consortia. IEEE Micro, 21(3):4–5, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3004.
pdf.

Robinson:2001:MSI

[Rob01b] Gary S. Robinson. Micro standards: International standards: Why do them? IEEE Micro, 21(2):4, 96, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2004.
pdf.

Robinson:2001:MSWb

[Rob01c] Gary S. Robinson. Micro standards: Why standards are here to stay. IEEE Micro, 21(4):6–7, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4006.
pdf;   m4006abs.htm.

Robinson:2001:MSWa

[Rob01d] Gary S. Robinson. Micro standards: Working in informal groups. IEEE Micro, 21(1): 13–13, 92, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1013.
pdf.

Roesgen:1986:ADM

[Roe86]  John P. Roesgen.  The

ADSP-2100 DSP microprocessor. IEEE Micro, 6(6): 49–59, November/December 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ryckbosch:2010:FAV

[RPE10] Frederick Ryckbosch, Stijn Polfliet, and Lieven Eeckhout. Fast, accurate, and validated full-system software simulation of x86 hardware. IEEE Micro, 30(6):46–56, November/December 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Raman:2000:ISS

[RPK00] Srinivas K. Raman, Vladimir Pentkovski, and Jagannath Keshava. Implementing streaming SIMD extensions on the Pentium III processor. IEEE Micro, 20(4): 47–57, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4047.
pdf;   http://www.computer.
org/micro/mi2000/m4047abs.
htm.

Rossi:2017:EEN

[RPL^{+}17] Davide Rossi, Antonio Pullini, Igor Loi, Michael Gautschi, Frank Kagan Gurkaynak, Adam Teman, Jeremy Constantin, Andreas Burg, Ivan Miro-Panades, Edith Beigne, Fabien Clermidy, Philippe Flatresse, and Luca Benini. Energy-efficient near-threshold parallel computing: The PULPv2 cluster. IEEE Micro, 37(5):20–31, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050020-
abs.html.

Ramadan:2008:MTT

[RRP^{+}08] Hany E. Ramadan, Christopher J. Rossbach, Donald E.

Porter, Owen S. Hofmann, Aditya Bhandari, and Emmett Witchel. MetaTM/ TxLinux: Transactional memory for an operating system. IEEE Micro, 28(1):42– 51, January/February 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rumsey:1990:AMM

[RS90] Michael Rumsey and John Sackett. An ASIC methodology for mixed analog-digital simulation. IEEE Micro, 10 (4):34–40, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Russell:1993:SRW

[RS93] Gordon Russell and Paul Shaw. Shifting register windows. IEEE Micro, 13(4):28– 35, July/August 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Renau:2006:EET

[RSC^{+}06] Jose Renau, Karin Strauss, Luis Ceze, Wei Liu, Smruti R. Sarangi, James Tuck, and Josep Torrellas. Energyefficient thread-level speculation. IEEE Micro, 26(1):80– 91, January/February 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ranganathan:2022:WSV

[RSC^{+}22] Parthasarathy Ranganathan, Daniel Stodolsky, Jeff Calow, Jeremy Dorfman, Marisabel Guevara, Clinton Wills Smullen IV, and Aki Kuusela. Warehousescale video acceleration. IEEE Micro, 42(4):18–26, July/ August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Richardson:2001:FTA

[RSE01] Paul Richardson, Larry Sieh, and Ali M. Elkateeb. Faulttolerant adaptive scheduling for embedded real-time systems. IEEE Micro, 21(5):41– 51, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/m5041abs.htm;
                                         http://dlib.computer.org/
                                        mi/books/mi2001/pdf/m5041.
                                        pdf.

Reick:2008:FTD

[RSS^{+}08] Kevin Reick, Pia N. Sanda, Scott Swaney, Jeffrey W. Kellington, Michael Mack, Michael Floyd, and Daniel Henderson. Fault-tolerant design of the IBM Power6 microprocessor. IEEE Micro, 28 (2):30–38, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reinemo:2010:EHP

[RSW10] Sven-Arne Reinemo, Tor Skeie, and Manoj K. Wadekar. Ethernet for high-performance data centers: On the new IEEE datacenter bridging standards. IEEE Micro, 30 (4):42–51, July/August 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rap:1986:MPI

[RT86] Michael D. Rap and R. Scott Tetrick. Microstandards — P1296: the interprocessor communication standard. IEEE Micro, 6(3):72– 77, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ruetz:1992:MIP

[RT92] Peter A. Ruetz and Po Tong.

A 160 Mpixel/s IDCT processor for HDTV. IEEE Micro, 12(5):28–32, September/October 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rusci:2023:DCT

[RT23] Manuele Rusci and Tinne Tuytelaars. On-device customization of tiny deep learning models for keyword spotting with few examples. IEEE Micro, 43(6):50–57, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rashid:2005:PEE

[RTHA05] M. Wasiur Rashid, Edwin J. Tan, Michael C. Huang, and David H. Albonesi. Powerefficient error tolerance in chip multiprocessors. IEEE Micro, 25(6):60–70, November/December 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rangan:2020:CLU

[RTJ20] R. Rangan, N. Turakhia, and A. Joly. Countering load-to-use stalls in the NVIDIA Turing GPU. IEEE Micro, 40(6): 59–66, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). See corrections [RTJ21].

Rangan:2021:CCL

[RTJ21] R. Rangan, N. Turakhia, and A. Joly. Corrections to Countering Load-to-Use Stalls in the NVIDIA Turing GPU. IEEE Micro, 41(1):83, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). See [RTJ20].

Ren:2010:GWP

[RTM^{+}10] Gang Ren, Eric Tune, Tipp Moseley, Yixin Shi, Silvius Rus, and Robert Hundt. Google-wide profiling: a continuous profiling infrastructure for data centers. IEEE Micro, 30(4):65–79, July/August 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ruckert:2002:UAA

[Rüc02] Ulrich Rückert. ULSI architectures for artificial neural networks. IEEE Micro, 22(3):10–19, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m3010.
pdf;   http://www.computer.
org/micro/mi2002/m3010abs.
htm.

Raina:2021:BPH

[RY21] P. Raina and C. Young. Best papers from Hot Chips 32. IEEE Micro, 41(2):6, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ryan:1988:IAO

[Rya88] David P. Ryan. Intel’s 80960:

an architecture optimized for embedded control. IEEE Micro, 8(3):63–76, May/June 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reddi:2018:TBD

[RYK18] Vijay Janapa Reddi, Hongil Yoon, and Allan Knies. Two billion devices and counting.

IEEE Micro, 38(1):6–21, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010006-
abs.html.

Rotem:2022:IAL

[RYR^{+}22] Efraim Rotem, Adi Yoaz, Lihu Rappoport, Stephen J. Robinson, Julius Yuli Mandelblat, Arik Gihon, Eliezer Weissmann, Rajshree Chabukswar, Vadim Basin, Russell Fenger, Monica Gupta, and Ahmad Yasin. Intel Alder Lake CPU architectures. IEEE Micro, 42(3):13–19, May/June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ryshpan:1984:MCC

[Rys84] J. Ryshpan. The MC68020 — corrections and comparisons. IEEE Micro, 4(5):3– 6, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sharangpani:2000:IPM

[SA00] Harsh Sharangpani and Ken Arora. Itanium processor microarchitecture. IEEE Micro, 20(5):24–43, September/October 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2000/

pdf/m5024.pdf;   http://
www.computer.org/micro/
mi2000/m5024abs.htm.

Savage:1999:DII

[SAA^{+}99] Stefan Savage, Thomas Anderson, Amit Aggarwal, David Becker, Neal Cardwell, Andy Collins, Eric Hoffman, John Snell, Amin Vahdat, Geoff Voelker, and John Zahorjan. Detour — informed Internet routing and transport. IEEE Micro, 19(1):50– 59, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1050.
pdf;   http://www.computer.
org/micro/mi1999/m1050abs.

htm. See correction [Sav99a].

Srinivasan:2005:LRT

[SABR05] Jayanth Srinivasan, Sarita V.

Adve, Pradip Bose, and Jude A. Rivers. Lifetime reliability: Toward an architectural solution. IEEE Micro, 25(3):70–80, May/June 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shalev:2020:COT

[SABS20] L. Shalev, H. Ayoub, N. Bshara, and E. Sabbag. A cloudoptimized transport protocol for elastic and scalable HPC. IEEE Micro, 40(6): 67–73, November/December 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slegel:1999:IGM

[SAC^{+}99] Timothy J. Slegel, Robert M. Averill III, Mark A. Check, Bruce C. Giamei, Barry W. Krumm, Christopher A. Krygowski, Wen H. Li, John S. Liptay, John D. MacDougall, Thomas J. McPherson, Jennifer A. Navarro, Eric M. Schwarz, Kevin Shum, and Charles F. Webb. IBM’s S/390 G5 microprocessor design. IEEE Micro, 19(2): 12–23, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1999/pdf/m2012.
                                        pdf;   http://www.computer.
                                        org/micro/mi1999/m2012abs.
                                        htm.

Singh:2021:FBN

[SAC^{+}21] Gagandeep Singh, Mohammed Alser, Damla Senol Cali, Dionysios Diamantopoulos, Juan Gómez-Luna, Henk Corporaal, and Onur Mutlu. FPGA-based near-memory acceleration of modern dataintensive applications. IEEE Micro, 41(4):39–48, July/ August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1987:LFT

[Sak87a] K. Sakamura. Looking into the future with Tron. IEEE Micro, 7(2):4–6, March/April 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1987:ATV

[Sak87b] Ken Sakamura. Architecture of the Tron VLSI CPU. IEEE Micro, 7(2):17–31, March/ April 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1987:BBO

[Sak87c] Ken Sakamura. Btron — the business-oriented operating system. IEEE Micro, 7 (2):53–65, March/April 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1987:TP

[Sak87d] Ken Sakamura. The Tron Project. IEEE Micro, 7(2):8– 14, March/April 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1988:RT

[Sak88] K. Sakamura. Recent trends.

IEEE Micro, 8(2):10–11, March/April 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1989:SFE

[Sak89] K. Sakamura. Special Far-East issue — computer projects in Japan — introduction. IEEE Micro, 9(3): 12–13, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1990:TIH

[Sak90a] K. Sakamura. The Tron intelligent house. IEEE Micro, 10(2):6–7, March/April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1990:GEI

[Sak90b] Ken Sakamura. Guest Editor’s introduction: The current Japanese computer scene. IEEE Micro, 10(2):12–13, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1991:GEI

[Sak91] Ken Sakamura. Guest Editor’s introduction: Presenting the Far East Special Issue for 1991. IEEE Micro, 11(4):12– 15, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1993:TWF

[Sak93] Ken Sakamura. Toward a world filled with computers. IEEE Micro, 13(5):6– 11, September/October 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1995:GEI

[Sak95] Ken Sakamura. Guest Editor’s introduction: Microelectronics in Japan. IEEE Micro, 15(6):8–10, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1997:GEI

[Sak97] Ken Sakamura. Guest Editors’ introduction: DRAM technology. IEEE Micro, 17 (6):8–9, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m6008.
pdf.

Sakamura:1999:ECMb

[Sak99a] Ken Sakamura. Editor-in-Chief message: Selecting a processor’s architecture must be based on more than its technical impact. IEEE Micro, 19(2):2, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2002.
pdf.

Sakamura:1999:ECMa

[Sak99b] Ken Sakamura. Editor-in-Chief message: Welcoming the 21st century. IEEE Micro, 19(1):2–3, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1002.
pdf.

Sakamura:1999:ECMe

[Sak99c] Ken Sakamura. Editor-in-Chief’s message: Advancing microelectronics technology. IEEE Micro, 19(5): 2, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5002.
pdf.

Sakamura:1999:ECMd

[Sak99d] Ken Sakamura. Editor-in-Chief’s message: Designing low-power CPUs. IEEE Micro, 19(4):2, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4002.
pdf.

Sakamura:1999:ECMc

[Sak99e] Ken Sakamura. Editor-in-Chief’s message: Relying on market forces. IEEE Micro, 19(3):2, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m3002.
pdf.

Sakamura:1999:GEI

[Sak99f] Ken Sakamura. Guest editor’s introduction: Entertainment and edutainment. IEEE Micro, 19(6):15–19, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6015.
pdf.

Sakamura:2000:ECMc

[Sak00a] Ken Sakamura. Editor-in-Chief’s message: Developing new computer architectures. IEEE Micro, 20(3):2, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3002.
pdf.

Sakamura:2000:ECMb

[Sak00b] Ken Sakamura. Editor-in-Chief’s message: New applications and demands. IEEE Micro, 20(2):2, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2002.
pdf.

Sakamura:2000:ECMa

[Sak00c]  Ken Sakamura.  Editor-

in-Chief’s message: Performance in the new millennium. IEEE Micro, 20(1):

2, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1002.
pdf.

Sakamura:2000:EMCa

[Sak00d] Ken Sakamura. EIC message: 21st-century microprocessors. IEEE Micro, 20 (4):10–11, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4010.
pdf;   http://www.computer.
org/micro/mi2000/m4010abs.
htm.

Sakamura:2000:EMCb

[Sak00e] Ken Sakamura. EIC message: Connecting will be commonplace. IEEE Micro, 20(5):2, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5002.
pdf.

Sakamura:2000:GEI

[Sak00f] Ken Sakamura. Guest Editor’s introduction: Stepping into the future. IEEE Micro, 20(6): 10–11, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi2000/pdf/m6010.
pdf.

Sakamura:2001:EMJ

[Sak01a] Ken Sakamura. EIC message: a Java-enabled evolution. IEEE Micro, 21(4): 2–3, 5, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4002.
pdf;   m4002abs.htm.

Sakamura:2001:EMN

[Sak01b] Ken Sakamura. EIC message: a new challenge. IEEE Micro, 21(2):2, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2002.
pdf.

Sakamura:2001:EMB

[Sak01c] Ken Sakamura. EIC message: Beginnings and endings. IEEE Micro, 21(1): 4–5, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1004.
pdf.

Sakamura:2001:EMSa

[Sak01d] Ken Sakamura. EIC message: In search of perfect computing. IEEE Micro, 21(3):2, May/June 2001.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3002.
pdf.

Sakamura:2001:EMSb

[Sak01e]  Ken Sakamura.  EIC’s

message: Surviving the slump, building the future. IEEE Micro, 21(5): 5, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5005abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m5005.
pdf.

Sakamura:2001:GEI

[Sak01f] Ken Sakamura. Guest editor’s introduction: Radio frequency identification and noncontact smart cards. IEEE Micro, 21 (6):4–6, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6004abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m6004.
pdf.

Sakamura:2002:EMNb

[Sak02a] Ken Sakamura. EIC’s message: a new definition for high-performance computing. IEEE Micro, 22(2):2, March/ April 2002. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2002.
pdf;   http://www.computer.
org/micro/mi2002/m2002abs.
htm.

Sakamura:2002:EMFb

[Sak02b]  Ken Sakamura.  EIC’s

message: Farewell message. IEEE Micro, 22(6): 2, November/December 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m6002.
pdf.

Sakamura:2002:EMFa

[Sak02c] Ken Sakamura. EIC’s message: Future SOC possibilities. IEEE Micro, 22(5): 7, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5007.
pdf;   http://www.computer.
org/micro/mi2002/m5007abs.
htm.

Sakamura:2002:EML

[Sak02d] Ken Sakamura. EIC’s message: Large market on the horizon. IEEE Micro, 22(4):4, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4004.

                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m4004abs.
                                        htm.

Sakamura:2002:EMNa

[Sak02e]  Ken Sakamura.  EIC’s

message: The next challenge. IEEE Micro, 22(1): 2, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/m1002abs.htm;
                                         http://dlib.computer.org/
                                        mi/books/mi2002/pdf/m1002.
                                        pdf.

Sakamura:2002:EMT

[Sak02f] Ken Sakamura. EIC’s message: The test of time. IEEE Micro, 22(3):2, May/June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m3002.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m3002abs.
                                        htm.

Sakamura:2002:GEI

[Sak02g] Ken Sakamura. Guest Editor’s introduction: Making computers invisible. IEEE Micro, 22(6):7–11, November/December 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2002/
                                        pdf/m6007.pdf.

Samueli:2000:BR

[Sam00] Henry Samueli. The broadband revolution. IEEE Micro, 20(2):16–26, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2016.
pdf;   http://www.computer.
org/micro/mi2000/m2016abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Sanders:1997:JCB

[San97a] B. Sanders. Javaone-97 — catching the buzz. IEEE Micro, 17(3):5–6, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3005.
pdf.

Sanz:1997:UTF

[San97b] Alfredo Sanz. A unified tool for fuzzy/neural network systems. IEEE Micro, 17(2):65– 69, March/April 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suzuki:1998:VAE

[SANK98] Kazumasa Suzuki, Tomohisa Arai, Kouhei Nadehara, and Ichiro Kuroda. V830R/ AV: An embedded multimedia superscalar RISC processor. IEEE Micro, 18 (2):36–47, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2036.
pdf;   http://www.computer.
org/micro/mi1998/m2036abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24– 26, 1997.

SEP:2010:PPM

[SAR10] SARC European Project. Parallel programming models for heterogeneous multicore architectures. IEEE Micro, 30(5):42–53, September/October 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Savage:1999:CDI

[Sav99a] S. Savage. Corrigendum: “Detour: Informed Internet Routing and Transport” (vol. 19, p. 55, 1999). IEEE Micro, 19 (2):9, March/April 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). See [SAA^{+}99].

Savell:1999:EDA

[Sav99b]  Thomas C. Savell.  The

EMU10K1 digital audio processor. IEEE Micro, 19 (2):49–57, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi1999/pdf/m2049.
pdf;   http://www.computer.
org/micro/mi1999/m2049abs.
htm.

Shacham:2010:RDD

[SAW^{+}10] Ofer Shacham, Omid Azizi, Megan Wachs, Stephen Richardson, and Mark Horowitz. Rethinking digital design: Why design must change. IEEE Micro, 30(6):9–24, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schmitter:1984:BFT

[SB84] Ernst J. Schmitter and Peter Baues. The Basic Fault-Tolerant System. IEEE Micro, 4(1):66–74, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Steinberg:2000:EAI

[SB00] Dan Steinberg and Yitzhak Birk. An empirical analysis of the IEEE-1394 serial bus protocol. IEEE Micro, 20 (1):58–65, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1058.
pdf;   http://www.computer.
org/micro/mi2000/m1058abs.
htm.

Shacham:2007:BUL

[SB07] Assaf Shacham and Keren Bergman. Building ultralowlatency interconnection networks using photonic integration. IEEE Micro, 27(4):6– 20, July/August 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Snelgrove:2023:SPT

[SB23] Martin Snelgrove and Robert Beachler. speedAI240: a 2petaflop, 30-Teraflops/W atmemory inference acceleration device with 1456 RISC-V cores. IEEE Micro, 43(3): 58–63, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stephens:2017:ASV

[SBB^{+}17] Nigel Stephens, Stuart Biles, Matthias Boettcher, Jacob Eapen, Mbou Eyole, Giacomo Gabrielli, Matt Horsnell, Grigorios Magklis, Alejandro Martinez, Nathanael Premillieu, Alastair Reid, Alejandro Rico, and Paul Walker. The ARM scalable vector extension. IEEE Micro, 37(2):26– 39, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020026-
abs.html.

Sarma:2001:RFI

[SBE01] Sanjay Sarma, David Brock, and Daniel Engels. Radio frequency identification and the electronic product code. IEEE Micro, 21(6): 50–54, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6050abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m6050.
pdf.

Sureshbabu:1997:DHR

[SBG97] R. M. Suresh babu, B. B.

Biswas, and G. Govindarajan. Developing highly reliable software. IEEE Micro, 17(5): 56–63, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1997/pdf/m5056.
pdf;   http://www.computer.
org/micro/mi1997/m5056abs.
htm.

Skadron:2007:LPD

[SBG^{+}07] Kevin Skadron, Pradip Bose, Kanad Ghose, Resit Sendag, Joshua J. Yi, and Derek Chiou. Low-power design and temperature management. IEEE Micro, 27(6): 46–57, November/December 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shum:2013:IZT

[SBJ13] C. Kevin Shum, Fadi Busaba, and Christian Jacobi. IBM zEC12: The third-generation high-frequency mainframe microprocessor. IEEE Micro, 33 (2):38–47, March/April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schmidt:1991:DSC

[SC91] Ulrich Schmidt and Knut Caesar. Datawave — a single-chip multiprocessor for video applications. IEEE Micro, 11(3): 22–25, 88–94, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sharma:2024:PPF

[SC24] Debendra Das Sharma and Swadesh Choudhary. Pipelined and partitionable forward error correction and cyclic redundancy check circuitry implementation for PCI Express 6.0 and Compute Express Link 3.0. IEEE Micro, 44 (2):50–59, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Scannell:1998:WD

[Sca98] E. Scannell. Windows-98 debut. IEEE Micro, 18(2): 3, March/April 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sharma:2012:ULE

[SCA^{+}12] Vibhu Sharma, Stefan Cosemans, Maryam Ashouie, Jos Huisken, Francky Catthoor, and Wim Dehaene. Ultra low-energy SRAM design for smart ubiquitous sensors. IEEE Micro, 32(5):10– 24, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stasiak:2005:CPL

[SCC^{+}05] Daniel Stasiak, Rajat Chaudhry, Dennis Cox, Stephen Posluszny, Jim Warnock, Steve Weitzel, Dieter Wendel, and Michael Wang. Cell processor lowpower design methodology. IEEE Micro, 25(6):71–78, November/December 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Segars:1995:ECP

[SCG95] Simon Segars, Keith Clarke, and Liam Goudge. Embedded control problems, Thumb, and the ARM7TDMI. IEEE Micro, 15(5):22–30, September/October 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schulthess:1984:RHL

[Sch84] Peter U. Schulthess. A reduced high-level-language instruction set. IEEE Micro, 4(3):55–67, May/June 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schachner:1991:OA

[Sch91a] J. M. Schachner. On an open architecture. IEEE Micro, 11(6):2, November/December 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schultz:1991:PHH

[Sch91b] Thomas W. Schultz. Peripheral hardware and a hands on multitasking lab. IEEE Micro, 11(1):30–33, 80–82, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schachner:1996:RVC

[Sch96] J. Schachner. RISC versus CISC. IEEE Micro, 16(1):2, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sugawara:2023:DMA

[SCH^{+}23] Yutaka Sugawara, Dong Chen, Ruud A. Haring, Abdullah Kayi, Eugene Ratzlaff, Robert M. Senger, Krishnan Sugavanam, Ralph Bellofatto, Ben J. Nathanson, and Craig Stunkel. Data movement accelerator engines on a prototype Power10 processor. IEEE Micro, 43(1):67–75, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Scott:1996:GC

[Sco96] Steve Scott. The GigaRing channel. IEEE Micro, 16 (1):27–34, January/February 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Scott:2014:MWA

[Sco14]  Steve Scott.  2013 Mau-

rice Wilkes Award Given to Parthasarathy (Partha) Ranganathan. IEEE Micro, 34 (1):90–91, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Seiler:2009:LMC

[SCS^{+}09] Larry Seiler, Doug Carmean, Eric Sprangle, Tom Forsyth, Pradeep Dubey, Stephen Junkins, Adam Lake, Robert Cavin, Roger Espasa, Ed Grochowski, Toni Juan, Michael Abrash, Jeremy Sugerman, and Pat Hanrahan. Larrabee: a many-core x86 architecture for visual computing. IEEE Micro, 29(1):10–21, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sood:1993:DM

[SCSR93] A. K. Sood, J. C. Carson, M. F. Suer, and R. Raphael. Developing 3D memories. IEEE Micro, 13 (6):6–7, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Samaras:2001:IIP

[SCV01] William A. Samaras, Naveen Cherukuri, and Srinivas Venkataraman. The IA-64 Itanium processor cartridge. IEEE Micro, 21(1):82–89, January/February 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2001/
                                        pdf/m1082.pdf;   http://
                                        www.computer.org/micro/
                                        mi2001/m1082abs.htm.

So:2011:MUI

[SCYY11] Hayden K.-H. So, Junying Chen, Billy Y. S. Yiu, and Alfred C. H. Yu. Medical ultrasound imaging: To GPU or not to GPU? IEEE Micro, 31(5):54–65, September/October 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sriraman:2021:UAO

[SD21] Akshitha Sriraman and Abhishek Dhanotia. Understanding acceleration opportunities at hyperscale. IEEE Micro, 41(3):34–41, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sethumadhavan:2004:SHM

[SDB^{+}04] Simha Sethumadhavan, Rajagopalan Desikan, Doug Burger, Charles R. Moore, and Stephen W. Keckler.

Scalable hardware memory disambiguation for high-ILP processors. IEEE Micro, 24(6):118–127, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6118.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6118.pdf.

Song:1994:PRM

[SDC94] S. P. Song, M. Denman, and J. Chang. The PowerPC-604 RISC microprocessor. IEEE Micro, 14(5):8–17, September/October 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Scherer:2023:TCT

[SDF^{+}23] Moritz Scherer, Alfio Di Mauro, Tim Fischer, Georg Rutishauser, and Luca Benini. TCN-CUTIE: A 1,036-TOp/s/W, 2.72-μJ/Inference, 12.2-mW all-digital ternary accelerator in 22-nm FDX technology. IEEE Micro, 43(1):42–48, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Skarlatos:2021:BFA

[SDG^{+}21] Dimitrios Skarlatos, Umur Darbaz, Bhargava Gopireddy, Nam Sung Kim, and Josep Torrellas. BabelFish: Fusing address translations for containers. IEEE Micro, 41(3): 57–62, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Segars:1997:APC

[Seg97] Simon Segars. ARM7TDMI power consumption: Reducing power in CPUs for portable and handheld products through efficient circuitand system-level design. IEEE Micro, 17(4):12–19, July/ August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://

                                        pascal.computer.org/mi/
                                        books/mi1997/pdf/m4012.
                                        pdf.

Sell:2018:XOX

[Sel18]  John Sell.  The Xbox

One X Scorpio Engine. IEEE Micro, 38(2):53–60, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/02/mmi2018020053-
                                        abs.html.

Sen:1986:DCS

[Sen86] C. Sen. Design case-study. IEEE Micro, 6(4):4, 79, July/August 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stone:1995:SP

[SF95] Janice M. Stone and Robert P.

Fitzgerald. Storage in the PowerPC. IEEE Micro, 15 (2):50–58, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shreejith:2018:SNI

[SF18] Shanker Shreejith and Suhaib A.

Fahmy. Smart network interfaces for advanced automotive applications. IEEE Micro, 38 (2):72–80, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/02/mmi2018020072-
abs.html.

Satyanarayanan:2022:ASW

[SFG^{+}22] Mahadev Satyanarayanan, Ziqiang Feng, Shilpa George, Jan Harkes, Roger Iyengar, Haithem Turki, and Padmanabhan Pillai. Accelerating silent witness storage. IEEE Micro, 42(6):39–47, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Singh:2023:XAL

[SFP^{+}23] Sarabjeet Singh, Xiong Fan, Ananth Krishna Prasad, Lin Jia, Anirban Nag, Rajeev Balasubramonian, Mahdi Nazm Bojnordi, and Elaine Shi. XCRYPT: Accelerating latticebased cryptography with memristor crossbar arrays. IEEE Micro, 43(5):45–54, September/October 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sanchez:2000:ADL

[SG00] Jesús Sánchez and Antonio González. Analyzing data locality in numeric applications. IEEE Micro, 20 (4):58–66, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4058.
pdf;   http://www.computer.
org/micro/mi2000/m4058abs.
htm.

Samadzadeh:2001:HSC

[SG01a] Mansur H. Samadzadeh and Loai E. Garalnabi. Hardware/software cost analysis of interrupt processing strategies. IEEE Micro, 21 (3):69–76, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3069.
pdf;   http://www.computer.
org/micro/mi2001/m3069abs.
htm.

Shah:2001:FUA

[SG01b] Devavrat Shah and Pankaj Gupta. Fast updating algorithms for TCAMs. IEEE Micro, 21(1):36–47, January/February 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
pdf/m1036.pdf;   http://
www.computer.org/micro/
mi2001/m1036abs.htm.

Spainhower:1994:IMF

[SGC94] Lisa Spainhower, Thomas A.

Gregg, and Ram Chillarege. IBM’s ES/9000 Model 982S fault-tolerant design for consolidation. IEEE Micro, 14 (1):48–59, January/February 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sodani:2016:KLS

[SGC^{+}16] Avinash Sodani, Roger Gramunt, Jesus Corbal, Ho-Seop Kim, Krishna Vinod, Sundaram Chinthamani, Steven Hutsell, Rajat Agarwal, and Yen-Chen Liu. Knights Landing: Second-generation Intel Xeon Phi product. IEEE Micro, 36(2):34–46, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2016/02/mmi2016020034-
abs.html.

Shah:2012:STD

[SGG^{+}12] Manish Shah, Robert Golla, Gregory Grohoski, Paul Jordan, Jama Barreh, Jeff Brooks, Mark Greenberg, Gideon Levinsky, Mark Luttrell, Christopher Olson, Zeid Samoail, Matt Smittle, and Tom Ziaja. Sparc T4: a dynamically threaded server-on-a-chip. IEEE Micro, 32(2):8– 19, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolens:2004:FBS

[SGK^{+}04] Jared C. Smolens, Brian T. Gold, Jangwoo Kim, Babak Falsafi, James C. Hoe, and Andreas G. Nowatzyk. Fingerprinting: Bounding softerror-detection latency and bandwidth. IEEE Micro, 24(6):22–29, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/06/m6022.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/06/m6022.pdf.

Seger:1993:VAS

[SGL93] Ulrich Seger, Heinz-Gerd G. Graf, and Marc E. Landgraf. Vision assistance in scenes with extreme contrast. IEEE Micro, 13(1):50– 56, January/February 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shah:2002:ERA

[SGP02] Devavrat Shah, Paolo Giaccone, and Balaji Prabhakar. Efficient randomized algorithms for inputqueued switch scheduling. IEEE Micro, 22(1):10–18, January/February 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2002/
m1010abs.htm;   http://
dlib.computer.org/mi/books/
mi2002/pdf/m1010.pdf.

Shapiro:1982:EDC

[Sha82] M. D. Shapiro. Errors detected in Caps programmers card. IEEE Micro, 2(4):6, October/December 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shaffer:1996:PC

[Sha96] R. A. Shaffer. The personalcomputer. IEEE Micro, 16 (6):23, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sharma:2022:LLL

[Sha22] Debendra Das Sharma. A low-latency and low-power approach for coherency and memory protocols on PCI Express 6.0 PHY at 64.0 GT/s with PAM-4 signaling. IEEE Micro, 42(2):37–43, March/ April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sharma:2023:CEL

[Sha23a] Debendra Das Sharma. Compute Express Link (CXL):

Enabling heterogeneous datacentric computing with heterogeneous memory hierarchy. IEEE Micro, 43(2):99– 109, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sharma:2023:NCS

[Sha23b] Debendra Das Sharma. Novel composable and scaleout architectures using compute express link. IEEE Micro, 43(2): 9–19, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sharma:2023:SPI

[Sha23c] Debendra Das Sharma. System on a package innovations with universal chiplet interconnect express (UCIe) interconnect. IEEE Micro, 43 (2):76–85, March/April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sturton:2019:FAS

[SHKS19] C. Sturton, M. Hicks, S. T. King, and J. M. Smith. FinalFilter: Asserting security properties of a processor at runtime. IEEE Micro, 39 (4):35–42, July/August 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shladover:1993:RDN

[Shl93] Steven E. Shladover. Research and development needs for advanced vehicle control systems. IEEE Micro, 13 (1):11–19, January/February 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shouse:1985:FCB

[Sho85]  D. V. Shouse.  ‘On the

fly’ CRC-16 byte-wise calculation for 8088-based computers. IEEE Micro, 5(2):67– 75, March/April 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1985:MHL

[SHS85] M. F. Smith, Yigal Hoffner, and Mark A. Sealey. Mapping high-level syntax and structure into assembly language. IEEE Micro, 5(4):67– 81, July/August 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakai:2008:MPM

[SHTE08] Junji Sakai, Inoue Hiroaki, Sunao Torii, and Masato Edahiro. Multitasking parallel method for high-end embedded appliances. IEEE Micro, 28(5):54–62, September/October 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sibigtroth:1984:MMD

[Sib84] James M. Sibigtroth. Motorola’s MC68HC11: Definition and design of a VLSI microprocessor. IEEE Micro, 4(1):54–65, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schulte:2015:AEC

[SIL^{+}15] Michael J. Schulte, Mike Ignatowski, Gabriel H. Loh, Bradford M. Beckmann, William C. Brantley, Sudhanva Gurumurthi, Nuwan Jayasena, Indrani Paul, Steven K. Reinhardt, and Gregory Rodgers. Achieving exascale capabilities through heterogeneous computing. IEEE Micro, 35 (4):26–36, July/August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040026-
abs.html.

Sima:1997:SII

[Sim97] Dezsõ Sima. Superscalar instruction issue. IEEE Micro, 17(5):28–39, September/October 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1997/
pdf/m5028.pdf;   http://
www.computer.org/micro/
mi1997/m5028abs.htm.

Sima:2000:DSR

[Sim00] Dezsö Sima. The design space of register renaming techniques. IEEE Micro, 20(5):70– 83, September/October 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m5070.
pdf;   http://www.computer.
org/micro/mi2000/m5070abs.
htm.

Shah:2002:MSC

[SIPM02] Devavrat Shah, Sundar Iyer, Balaji Prabhakar, and Nick McKeown. Maintaining statistics counters in router line cards. IEEE Micro, 22(1):76– 81, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1076abs.htm;
 http://dlib.computer.org/
mi/books/mi2002/pdf/m1076.
pdf.

StAmant:2009:MSA

[SJB09] Renée St. Amant, Daniel A.

Jiménez, and Doug Burger. Mixed-signal approximate computation: a neural predictor case study. IEEE Micro, 29(1):104–115, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shen:2019:AEE

[SJFM19] Y. Shen, T. Ji, M. Ferdman, and P. Milder. Argus: An end-to-end framework for accelerating CNNs on FP-GAs. IEEE Micro, 39(5):17– 25, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shah:2023:EMB

[SJL23] Ishan Shah, Akanksha Jain, and Calvin Lin. Effective mimicry of Bélády’s MIN policy. IEEE Micro, 43(4):45– 52, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schiele:2001:SAC

[SJO01] Bernt Schiele, Tony Jebara, and Nuria Oliver. Sensoryaugmented computing: Wearing the museum’s guide. IEEE Micro, 21(3):44–52, May/ June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m3044.
                                        pdf;   http://www.computer.
                                        org/micro/mi2001/m3044abs.
                                        htm.

Sohie:1988:DSP

[SK88] Guy R. L. Sohie and Kevin L. Kloker. A digital signal processor with IEEE floatingpoint arithmetic. IEEE Micro, 8(6):49–67, November/December 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sibai:1997:TMR

[SK97] Fadi N. Sibai and Sunil D.

Kulkarni. A time-multiplexed reconfigurable neuroprocessor — combining analog and digital technology in a neuroprocessor using time-multiplexed pulse stream communication. IEEE Micro, 17(1):58–65, January/February 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:2001:EWA

[SK01] Ken Sakamura and Noboru Koshizuka. The eTRON widearea distributed-system architecture for E-commerce. IEEE Micro, 21(6):7–12, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6007abs.htm;
 http://dlib.computer.org/
mi/books/mi2001/pdf/m6007.
pdf.

Sakamura:2002:EOR

[SK02] Ken Sakamura and Noboru Koshizuka. T-engine: The open, real-time embeddedsystems platform. IEEE Micro, 22(6):48–57, November/December 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2002/
                                        pdf/m6048.pdf;   http://
                                        www.computer.org/micro/
                                        mi2002/m6048abs.htm.

Sanchez:2012:SEF

[SK12] Daniel Sanchez and Christos Kozyrakis. Scalable and efficient fine-grained cache partitioning with Vantage. IEEE Micro, 32(3):26–37, May/ June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slijepcevic:2014:TVF

[SKA^{+}14a] Mladen Slijepcevic, Leonidas Kosmidis, Jaume Abella, Eduardo Quinones, and Francisco J. Cazorla. Timing verification of fault-tolerant chips for safety-critical applications in harsh environments. IEEE Micro, 34(6):8–19, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/06/mmi2014060008-
                                        abs.html.

Sung:2014:DEH

[SKA14b] Hyojin Sung, Rakesh Komuravelli, and Sarita V. Adve. DeNovoND: Efficient hardware for disciplined nondeterminism. IEEE Micro, 34(3): 138–148, May/June 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Suresh:2023:NAN

[SKC^{+}23] Kaushik Kandadi Suresh, Kawthar Shafie Khorassani, Chen Chun Chen, Bharath Ramesh, Mustafa Abduljabbar, Aamir Shafi, Hari Subramoni, and Dhabaleswar K. Panda. Network-assisted noncontiguous transfers for GPUaware MPI libraries. IEEE Micro, 43(2):131–139, March/ April 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stuecheli:2011:CDL

[SKJ^{+}11] Jeffrey Stuecheli, Dimitris Kaseridis, Lizy K. John, David Daly, and Hillery C. Hunter. Coordinating DRAM and last-level-cache policies with the virtual write queue. IEEE Micro, 31(1):90–98, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Simar:1992:FPP

[SKL^{+}92] Ray Simar, Jr., Peter Koeppen, Jerald Leach, Steve Marshall, Dave Francis, Greg Mekras, Jeffrey Rosenstrauch, and Scott Anderson. Floatingpoint processors join forces in parallel processing architectures. IEEE Micro, 12(4):60– 69, July/August 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Semenov:1997:DAP

[SKLY97] Alex Semenov, Albert M. Koelmans, Lee Lloyd, and Alexandre Yakovlev. Designing an asynchronous processor using Petri nets. IEEE Micro, 17(2):54–64, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sourdis:2016:RCM

[SKM^{+}16] Ioannis Sourdis, Danish Anis Khan, Alirad Malek, Stavros Tzilis, Georgios Smaragdos, and Christos Strydis. Resilient chip multiprocessors with mixed-grained reconfigurability. IEEE Micro, 36(1):35–45, January/February 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/01/mmi2016010035-
                                        abs.html.

Sunaga:2023:AGB

[SKM23] Kazuki Sunaga, Masaaki Kondo, and Hiroki Matsutani. Addressing the gap between training data and deployed environment by on-device learning. IEEE Micro, 43(6): 66–73, November/December 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Skordalakis:1983:MA

[Sko83]  E. Skordalakis.  Meta-

assemblers. IEEE Micro, 3(2):

6–16, March/April 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1989:OBS

[SKO89] Ken Sakamura, Yoshiaki Kushiki, and Kazuhiro Oda. An overview of the BTRON/ 286 specification. IEEE Micro, 9(3):14–25, May/June 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Swaminathan:2013:SSD

[SKS^{+}13] Karthik Swaminathan, Emre Kultursay, Vinay Saripalli, Vijaykrishnan Narayanan, Mahmut T. Kandemir, and Suman Datta. Steep-slope devices: From dark to dim silicon. IEEE Micro, 33(5):50– 59, September/October 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sato:2022:CDS

[SKTO22] Mitsuhisa Sato, Yuetsu Kodama, Miwako Tsuji, and Tesuya Odajima. Co-design and system for the supercomputer Fugaku. IEEE Micro, 42 (2):26–34, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sriram:2023:HHS

[SKW^{+}23] Karthik Sriram, Ioannis Karageorgos, Xiayuan Wen, Ján Veselý, Nick Lindsay, Michael Wu, Lenny Khazan, Raghavendra Pradyumna Pothukuchi, Rajit Manohar, and Abhishek Bhattacharjee. HALO: a hardware software co-designed processor for brain computer interfaces. IEEE Micro, 43(3): 64–72, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1984:AAS

[SL84a] M. F. Smith and B. E. Luff. Automatic assembler source translation from the Z80 to the MC6809. IEEE Micro, 4(2):3–9, March/April 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stigall:1984:MCM

[SL84b] Paul D. Stigall and Brian E. Lenharth. A microprocessorcontrolled message display system. IEEE Micro, 4(2):10– 25, March/April 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Soderquist:1997:DSR

[SL97] Peter Soderquist and Miriam Leeser. Division and square root: Choosing the right implementation: Exploring the major design choices for microprocessor implementations of floating-point division and square root. IEEE Micro, 17 (4):56–66, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4056.
pdf.

Schuehler:2003:TST

[SL03] David V. Schuehler and John W. Lockwood. TCP splitter: a TCP/IP flow monitor in reconfigurable hardware. IEEE Micro, 23(1):54– 59, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1054.
pdf;   http://www.computer.
org/micro/mi2003/m1054abs.
htm.

Slater:1989:VSB

[Sla89] M. Slater. VLIW, superscalar, and 64 bit or not. IEEE Micro, 9(6):104–??, November/December 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1990:AVI

[Sla90a]  M. Slater.  AMD v. In-

tel. IEEE Micro, 10(6): 104–??, November/December 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1990:FPS

[Sla90b] M. Slater. Failings of the patent system. IEEE Micro, 10(4):96–??, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1990:PCA

[Sla90c] M. Slater. Protecting computer architectures. IEEE Micro, 10(5):96–??, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1990:VF

[Sla90d] M. Slater. The view from 10,000 feet. IEEE Micro, 10 (1):96–??, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1990:WR

[Sla90e] M. Slater. What is RISC? IEEE Micro, 10(3):96–??, May/June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1990:WNF

[Sla90f] M. Slater. Who needs faster processors? IEEE Micro, 10 (2):96–??, March/April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1991:EM

[Sla91a] M. Slater. The end of the 386monopoly. IEEE Micro, 11 (2):88–??, March/April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1991:EA

[Sla91b] M. Slater. Evolving architectures. IEEE Micro, 11(1):96– ??, January/February 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slater:1996:MTM

[Sla96] Michael Slater. The microprocessor today — MicroDesign resources’ founder outlines technology and business issues in today’s microprocessor industry. IEEE Micro, 16(6):32–44, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suh:2004:ICCa

[SLB04a] Taeweon Suh, Hsien-Hsin S.

Lee, and Douglas M. Blough. Integrating cache coherence protocols for heterogeneous multiprocessor systems, part 1. IEEE Micro, 24(4):33– 41, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4033abs.htm;
http://csdl.computer.org/
dl/mags/mi/2004/04/m4033.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/04/
m4033.pdf.

Suh:2004:ICCb

[SLB04b] Taeweon Suh, Hsien-Hsin S. Lee, and Douglas M. Blough. Integrating cache coherence protocols for heterogeneous multiprocessor systems, part 2. IEEE Micro, 24(5):70– 78, September/October 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/05/m5070.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/05/m5070.pdf.

Schares:2014:TOO

[SLC^{+}14] Laurent Schares, Benjamin G. Lee, Fabio Checconi, Russell Budd, Alexander Rylyakov, Nicolas Dupuis, Fabrizio Petrini, Clint L. Schow, Pablo Fuentes, Oliver Mattes, and Cyriel Minkenberg. A throughput-optimized optical network for dataintensive computing. IEEE Micro, 34(5):52–63, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050052-
abs.html.

Shin:2018:DEE

[SLL^{+}18] Dongjoo Shin, Jinmook Lee, Jinsu Lee, Juhyoung Lee, and Hoi-Jun Yoo. DNPU: An energy-efficient deep-learning processor with heterogeneous multi-core architecture. IEEE Micro, 38(5):85–93, September/October 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050085-
abs.html.

Skellern:1997:HSW

[SLM^{+}97] David J. Skellern, L. H. Charles Lee, Tom McDermott, Neil H. E. Weste, John Dalton, Jeffrey Graham, Tan F. Wong, Andrew F. Myles, Terence M. P. Percival, and Philip J. Ryan. A high-speed wireless LAN — using millimeter-wave frequencies, which can accommodate link speeds of hundreds of mbits per second, for indoor communications systems. IEEE Micro, 17(1):40– 47, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sim:2014:CSR

[SLSO14] Jaewoong Sim, Gabriel H.

Loh, Vilas Sridharan, and Mike O’Connor. A configurable and strong RAS solution for die-stacked DRAM caches. IEEE Micro, 34(3):80– 90, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Snyder:2023:RCC

[SLZ23] John Snyder, Alvin R. Lebeck, and Danyang Zhuo. RDMA congestion control: It is only for the compliant. IEEE Micro, 43(1):76–82, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stelzer:1985:MBC

[SM85] Eric H. Stelzer and Randy H. Moss. A microcomputer-based control system for a threejoint robot arm. IEEE Micro, 5(3):22–40, May/June 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suga:2000:IFE

[SM00] Atsuhiro Suga and Kunihiko Matsunami. Introducing the FR500 embedded microprocessor. IEEE Micro, 20 (4):21–27, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m4021.
                                        pdf;   http://www.computer.
                                        org/micro/mi2000/m4021abs.
                                        htm.

Suresh:2016:CSA

[SMAS16] Chandra K. H. Suresh, Bodhisatwa Mazumdar, Sk Subidh Ali, and Ozgur Sinanoglu. A comparative security analysis of current and emerging technologies. IEEE Micro, 36(5):50–61, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.

computer.org/csdl/mags/
mi/2016/05/mmi2016050050-
abs.html.

Sterling:1987:EIP

[SMCT87] Thomas L. Sterling, Albert J.

Musciano, Ellery Y. Chan, and Douglas A. Thomae. Effective implementation of a parallel language on a multiprocessor. IEEE Micro, 7(6): 46–62, November/December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sachs:1991:DIT

[SMHB91] Howard G. Sachs, Harlan McGhan, Lee F. Hanson, and Nathan A. Brookwood. Design and implementation trade-offs in the Clipper C400 architecture. IEEE Micro, 11 (3):18–21, 74–80, May/June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1982:NOI

[Smi82] T. Smith. The near-optimal instruction set. IEEE Micro, 2 (3):5–7, July/September 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1985:WNR

[Smi85] M. F. Smith. What’s new?

— reply. IEEE Micro, 5 (6):88, November/December 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1986:ASPa

[Smi86a] G. Smith. Assembler study praised. IEEE Micro, 6(1): 4–6, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1986:ASPb

[Smi86b] M. F. Smith. Assembler study praised — reply. IEEE Micro, 6(1):5, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1992:HRD

[Smi92] Michael R. Smith. How RISCy is DSP? IEEE Micro, 12(6): 10–23, November/December 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1996:MEM

[Smi96a] H. Smith. Marketing the early microprocessors. IEEE Micro, 16(6):18, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:1996:MDM

[Smi96b] P. Smith. Microprocessor design in the mid-1990s. IEEE Micro, 16(6):29, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smith:2017:RAS

[Smi17] James E. Smith. Research agenda: Spacetime computation and the neocortex. IEEE Micro, 37(1):8–14, January/February 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010008-
abs.html.

Suleman:2011:DMM

[SMJ^{+}11] M. Aater Suleman, Onur Mutlu, Jose A. Joao, Khubaib, and Yale Patt. Data marshaling for multicore systems. IEEE Micro, 31(1):56– 64, January/February 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schuehler:2004:AHB

[SML04] David V. Schuehler, James Moscola, and John W. Lockwood. Architecture for a hardware-based, TCP/IP content-processing system. IEEE Micro, 24(1):62–69, January/February 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1062abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/01/m1062.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/01/
m1062.pdf.

Shaikh:2022:AIB

[SMM^{+}22] Faisal Karim Shaikh, Mohsin Ali Memon, Naeem Ahmed Mahoto, Sherali Zeadally, and Jamel Nebhen. Artificial intelligence best practices in smart agriculture. IEEE Micro, 42 (1):17–24, January/February 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1986:PPW

[Smo86a]  M. Smolin.  Publish and

or perish (or, who wants to use a trial-use standard?). IEEE Micro, 6(6):80, November/December 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1986:M

[Smo86b] M. D. Smolin. Microstandards. IEEE Micro, 6(5):63– 64, September/October 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1987:FMS

[Smo87a] M. Smolin. Future micro standards projects. IEEE Micro, 7(6):88–89, November/December 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1987:M

[Smo87b] M. Smolin. Microstandards. IEEE Micro, 7(1):76–77, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1987:MMI

[Smo87c] M. Smolin. MicroStandards — more on IEEE standards generation. IEEE Micro, 7(5):92– 94, September/October 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1987:RTW

[Smo87d] M. Smolin. Rebuttal — them windmills got teeth. IEEE Micro, 7(2):90–91, March/April 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1988:ES

[Smo88a] M. Smolin. Excitement in standards. IEEE Micro, 8(2): 81, March/April 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1988:IBB

[Smo88b] M. Smolin. IEEE 32-bit, backplane bus comparisons. IEEE Micro, 8(3):5–7, May/June 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Smolin:1988:WHS

[Smo88c] M. Smolin. What’s happening with study-groups? IEEE Micro, 8(4):78–81, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suleman:2010:ACS

[SMQP10] M. Aater Suleman, Onur Mutlu, Moinuddin K. Qureshi, and Yale N. Patt. Accelerating critical section execution with asymmetric multicore architectures. IEEE Micro, 30 (1):60–70, January/February 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sha:2007:NSL

[SMR07] Tingting Sha, Milo M. K. Martin, and Amir Roth. NoSQ: Store-load communication without a store queue. IEEE Micro, 27(1):106–113, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stanley-Marbell:2018:PCA

[SMR18] Phillip Stanley-Marbell and Martin Rinard. Perceivedcolor approximation transforms for programs that draw. IEEE Micro, 38(4):20–29, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040020-
abs.html.

Stanley-Marbell:2020:WHP

[SMR20] P. Stanley-Marbell and M. Rinard. Warp: A hardware platform for efficient multimodal sensing with adaptive approximation. IEEE Micro, 40 (1):57–66, January/February 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Szafaryn:2013:EOM

[SMS13] Lukasz G. Szafaryn, Brett H.

Meyer, and Kevin Skadron. Evaluating overheads of multibit soft-error protection in the processor core. IEEE Micro, 33(4):56–65, July/August 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sukhwani:2014:DAR

[SMT^{+}14] Bharat Sukhwani, Hong Min, Mathew Thoennes, Parijat Dube, Bernard Brezzo, Sameh Asaad, and Donna Eng Dillenberger. Database analytics: A reconfigurable-computing approach. IEEE Micro, 34 (1):19–29, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Sarangi:2007:PPD

[SNC^{+}07] Smruti Sarangi, Satish Narayanasamy, Bruce Carneal, Abhishek Tiwari, Brad Calder, and Josep Torrellas. Patching processor design errors with programmable hardware. IEEE Micro, 27(1):12–25, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sankaralingam:2003:EIT

[SNL^{+}03] Karthikeyan Sankaralingam, Ramadass Nagarajan, Haiming Liu, Changkyu Kim, Jaehyuk Huh, Doug Burger, Stephen W. Keckler, and Charles Moore. Exploiting ILP, TLP, and DLP with the polymorphous TRIPS architecture. IEEE Micro, 23(6):46–51, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6046abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/06/m6046.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6046.pdf.

Singh:2013:SFA

[SNM^{+}13] Abhayendra Singh, Satish Narayanasamy, Daniel Marino, Todd Millstein, and Madanlal Musuvathi. A safety-first approach to memory models. IEEE Micro, 33(3):96– 104, May/June 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shiri:2022:ELG

[SNM^{+}22] Aidin Shiri, Mozhgan Navardi, Tejaswini Manjunath, Nicholas R. Waytowich, and Tinoosh Mohsenin. Efficient languageguided reinforcement learning for resource-constrained autonomous systems. IEEE Micro, 42(6):107–114, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sinanoglu:2002:ECA

[SO02] Ozgur Sinanoglu and Alex Orailoglu. Efficient construction of aliasing-free compaction circuitry. IEEE Micro, 22(5):82–92, September/October 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2002/
pdf/m5082.pdf;   http://
www.computer.org/micro/
mi2002/m5082abs.htm.

Sell:2014:XOS

[SO14] John Sell and Patrick O’Connor.

The Xbox One system on a chip and Kinect sensor. IEEE Micro, 34(2):44–53, March/ April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Solihin:2019:PMA

[Sol19] Y. Solihin. Persistent memory: Abstractions, abstractions, and abstractions. IEEE Micro, 39(1):65–66, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sood:1993:ETR

[Soo93] A. K. Sood. On the edge — thermal resistance (ICs). IEEE Micro, 13(4):52–58, July/August 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sosnowski:1994:TFT

[Sos94] Janusz Sosnowski. Transient fault tolerance in digital systems. IEEE Micro, 14(1):24– 35, January/February 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Schmalzel:1992:GEI

[SP92] John L. Schmalzel and Parimal A. Patel. Guest Editors’ introduction: Transforming the world of digital signal processing. IEEE Micro, 12 (6):8–9, November/December 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shenck:2001:ESS

[SP01] Nathan S. Shenck and Joseph A. Paradiso. Energy scavenging with shoe-mounted piezoelectrics. IEEE Micro, 21 (3):30–42, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m3030.
                                        pdf;   http://www.computer.
                                        org/micro/mi2001/m3030abs.
                                        htm.

Selvaggi:2009:BMP

[SP09] Richard Selvaggi and Larry Pearlstein. Broadcom mediaDSP: a platform for building programmable multicore video processors. IEEE Micro, 29(2):30–45, March/April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sherwood:2003:DEP

[SPH^{+}03] Timothy Sherwood, Erez Perelman, Greg Hamerly, Suleyman Sair, and Brad Calder. Discovering and exploiting program phases. IEEE Micro, 23(6):84–93, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6084abs.htm;
http://csdl.computer.org/
dl/mags/mi/2003/06/m6084.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6084.pdf.

Shang:2006:TAC

[SPKJ06] Li Shang, Li-Shiuan Peh, Amit Kumar, and Niraj K. Jha. Temperature-aware on-chip networks. IEEE Micro, 26(1):130–139, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sprunt:2002:BPM

[Spr02a] Brinkley Sprunt. The basics of performance-monitoring hardware. IEEE Micro, 22(4): 64–71, July/August 2002.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4064.
pdf;   http://www.computer.
org/micro/mi2002/m4064abs.
htm.

Sprunt:2002:PPM

[Spr02b] Brinkley Sprunt. Pentium 4 performance-monitoring features. IEEE Micro, 22(4): 72–82, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4072.
pdf;   http://www.computer.
org/micro/mi2002/m4072abs.
htm.

Shah:2004:NCP

[SPRK04] Niraj Shah, William Plishker, Kaushik Ravindran, and Kurt Keutzer. NP-Click: a productive software development approach for network processors. IEEE Micro, 24(5):45– 54, September/October 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/05/m5045.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/05/m5045.pdf.

Storer:1992:APM

[SPT^{+}92] Richard Storer, Mike R. Pout, Andrew R. Thomson, Erik L. Dagless, Andrew W. G. Duller, A. Paul Marriott, and Peter J. Hicks. An associative processing module for a heterogeneous vision architecture. IEEE Micro, 12(3):42– 55, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Srinivasan:2004:CFP

[SRA^{+}04] Srikanth T. Srinivasan, Ravi Rajwar, Haitham Akkary, Amit Gandhi, and Michael Upton. Continual flow pipelines: Achieving resourceefficient latency tolerance. IEEE Micro, 24(6):62–73, November/December 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6062.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6062.pdf.

Seaborn:1991:SGI

[SRJ^{+}91] True Seaborn, Peter R. Rony, Richard C. Jaeger, James J. Farrell, and Joe Hootman. Sometimes a good idea beats a good plan (but not often). IEEE Micro, 11(1):5–6, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sha:1991:SFR

[SRL91] Lui Sha, Ragunathan Rajkumar, and John P. Lehoczky. Special feature: Real-time computing with IEEE Futurebus+. IEEE Micro, 11(3): 30–??, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sinclair:2023:FYI

[SRU^{+}23] Matthew D. Sinclair, Parthasarathy Ranganathan, Gaurang Upasani, Adrian Sampson, David Patterson, Rutwik Jain, Nidhi Parthasarathy, and Shaan Shah. Fifty years of the International Symposium on Computer Architecture: a datadriven retrospective. IEEE Micro, 43(6):109–124, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shao:2015:AAA

[SRWB15] Yakun Sophia Shao, Brandon Reagen, Gu-Yeon Wei, and David Brooks. The Aladdin approach to accelerator design and modeling. IEEE Micro, 35(3):58–70, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030058-
abs.html.

Stern:1982:CWS

[SS82] Richard H. Stern and Jeffrey L. Squires. Can we stop software theft? IEEE Micro, 2(1):13–25, January/March 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sterbenz:2005:GEI

[SS05] James P. G. Sterbenz and Dimitrios Stiliadis. Guest Editors’ introduction: Hot Interconnects 12. IEEE Micro, 25(1):8–9, January/ February 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/01/m1008.pdf;
http://csdl.computer.org/
dl/mags/mi/2005/01/m1008.
htm.

Sell:2006:GEI

[SS06] John Sell and Alan Jay Smith.

Guest Editors’ introduction: Hot Chips 17. IEEE Micro, 26(2):8–9, March/April 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suresh:2016:AAE

[SS16] Amoghavarsha Suresh and John Sartori. Automated algorithmic error resilience based on outlier detection. IEEE Micro, 36(1):46–59, January/February 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/01/mmi2016010046-
abs.html.

Scherer:2022:SIH

[SS22] Alisa Scherer and Guri Sohi.

Special issue on Hot Chips 33. IEEE Micro, 42(3):6, May/ June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Seznec:2016:PMB

[SSA16] Andre Seznec, Joshua San Miguel, and Jorge Albericio. Practical multidimensional branch prediction. IEEE Micro, 36(3):10– 19, May/June 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2016/03/mmi2016030010-
                                        abs.html.

Skinner:1995:ONN

[SSB95] Steven R. Skinner, James E. Steck, and Elizabeth C. Behrman. Optical neural network using Kerr-type nonlinear materials. IEEE Micro, 15(3):52–54, May/June 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suggs:2020:AZP

[SSB20] D. Suggs, M. Subramony, and D. Bouvier. The AMD Zen 2 processor. IEEE Micro, 40 (2):45–52, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sullivan:2022:CMS

[SSC^{+}22] Michael B. Sullivan, Nirmal R. Saxena, Mike O Connor, Donghyuk Lee, Paul Racunas, Saurabh Hukerikar, Timothy Tsai, Siva Kumar Sastry Hari, and Stephen W. Keckler. Characterizing and mitigating soft errors in GPU DRAM. IEEE Micro, 42 (4):69–77, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Singh:2014:CCG

[SSF^{+}14] Inderpreet Singh, Arrvindh Shriraman, Wilson W. L. Fung, Mike O’Connor, and Tor M. Aamodt. Cache coherence for GPU architectures. IEEE Micro, 34(3):69– 79, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Sakamura:1988:ITS

[SSH88] Ken Sakamura, Ryoichi Sano, and Kazuhiko Honma. Introducing Tobus: the system bus in the TRON architecture. IEEE Micro, 8(2):47– 59, March/April 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Skadron:2003:TAC

[SSH^{+}03] Kevin Skadron, Mircea R.

Stan, Wei Huang, Sivakumar Velusamy, Karthik Sankaranarayanan, and David Tarjan. Temperature-aware computer systems: Opportunities and challenges. IEEE Micro, 23(6):52–61, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/06/m6052abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/06/m6052.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2003/06/
                                        m6052.pdf.

Sudhakar:2023:DCW

[SSK23] Soumya Sudhakar, Vivienne Sze, and Sertac Karaman. Data centers on wheels: Emissions from computing onboard autonomous vehicles. IEEE Micro, 43(1):29–39, January/February 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Silverman:1982:MMM

[SSL82] G. Silverman, A. Stundel, and J. Lehman. The modular multiprocessor — a model for laboratory instrument design. IEEE Micro, 2(2):51– 62, April/June 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sabne:2015:UPH

[SSLV15] Amit Sabne, Putt Sakdhnagool, Seyong Lee, and Jeffrey S. Vetter. Understanding portability of a high-level programming model on contemporary heterogeneous architectures. IEEE Micro, 35 (4):48–58, July/August 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/04/mmi2015040048-
abs.html.

Sanchez:1987:GHP

[SSMI87] Eduardo Sanchez, Patrick Sommer, Jacques Menu, and Christian Iseli. A general heap processor. IEEE Micro, 7(6): 29–40, November/December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sugumar:2021:MTN

[SSR21] R. Sugumar, M. Shah, and R. Ramirez. Marvell ThunderX3: Next-generation Armbased server processor. IEEE Micro, 41(2):15–21, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sumbul:2024:FDR

[SsSMB24] H. Ekin Sumbul, Jae sun Seo, Daniel H. Morris, and Edith Beigne. A fully digital and row-pipelined computein-memory neural network accelerator with system-on-chip-level benchmarking for augmented/virtual reality applications. IEEE Micro, 44 (2):61–70, March/April 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sase:1997:MLA

[SSY97] Ichiro Sase, Nobuyuki Shimizu, and Takashi Yoshikawa. Multimedia LSI accelerator with embedded DRAM. IEEE Micro, 17(6):49–54, November/December 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1997/
                                        pdf/m6049.pdf;   http://
                                        www.computer.org/micro/
                                        mi1997/m6049abs.htm.

Sethumadhavan:2019:SA

[ST19] S. Sethumadhavan and M. Tiwari. Secure architectures. IEEE Micro, 39(4):6– 7, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Simar:2021:HVW

[ST21] Ray Simar and Reid Tatge. How VLIWs were adopted as digital signal processors. IEEE Micro, 41(6):121–128, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Starner:2001:CWCa

[Sta01a] Thad Starner. The challenges of wearable computing: Part 1. IEEE Micro, 21(4):44–52, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m4044.
                                        pdf;   m4044abs.htm.

Starner:2001:CWCb

[Sta01b] Thad Starner. The challenges of wearable computing: Part 2. IEEE Micro, 21(4):54–67, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4054.
pdf;   m4054abs.htm.

Stern:1983:MLP

[Ste83a] R. H. Stern. Micro law: a proposal for a system of computer software protection. IEEE Micro, 3(3):61–65, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1983:ML

[Ste83b] Richard H. Stern. Micro law. IEEE Micro, 3(4):88– 92, July/August 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1983:MLCa

[Ste83c] Richard H. Stern. Micro law:

Can software be tied to hardware? Part I. IEEE Micro, 3(1):23–33, January/February 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1983:MLCb

[Ste83d] Richard H. Stern. Micro law:

Can software be tied to hardware? Part II. IEEE Micro, 3 (2):17–25, March/April 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1984:MLCa

[Ste84a] R. H. Stern. Micro law: 9th Circuit overturns Data General decision, rules that software-hardware tie-in is illegal per se. IEEE Micro, 4(5): 66–67, September/October 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1984:MLCb

[Ste84b] R. H. Stern. Micro law: Copyprotection-defeating programs — should Congress act? IEEE Micro, 4(6):84–85, November/December 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1984:MLM

[Ste84c] R. H. Stern. Micro law: More on the copyright controversy — who owns software. IEEE Micro, 4(2):69– 70, March/April 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1984:MLW

[Ste84d] R. H. Stern. Micro law: What is the function of an operating system? IEEE Micro, 4(1):80–81, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1984:PWG

[Ste84e] R. G. Stewart. P854 working group completes radixindependent floating-point draft. IEEE Micro, 4(1):82– 83, January/February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1985:ARQ

[Ste85a] R. H. Stern. Answers to readers questions. IEEE Micro, 5(6):86–88, November/December 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1985:MLF

[Ste85b] R. H. Stern. Micro law:

Further chip rights developments. IEEE Micro, 5(4):90– 92, July/August 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1985:MLM

[Ste85c] R. H. Stern. Micro law: Is microcode hardware or software? IEEE Micro, 5(2):89– 91, March/April 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1985:MLP

[Ste85d] R. H. Stern. Micro law:

Proprietary rights in cell libraries. IEEE Micro, 5(3): 73–78, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1985:MLS

[Ste85e] R. H. Stern. Micro law: Source code difference no protection against infringement suit. IEEE Micro, 5(2):88– 91, March/April 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1985:LCD

[Ste85f] D. J. Stewart. Logical choice and the discriminating designer. IEEE Micro, 5(3): 68–70, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1985:LYL

[Ste85g] R. G. Stewart. 8 long years … but success at last. IEEE Micro, 5(3):80–82, May/June 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1985:SEH

[Ste85h] R. G. Stewart. In search of excellence in high technology companies. IEEE Micro, 5 (2):86–87, March/April 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1986:MLF

[Ste86a] R. H. Stern. Micro law:

Field-programmable logic devices — are they hardware or software — can their programmed configurations be protected against copying. IEEE Micro, 6(5):61– 78, September/October 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1986:MLP

[Ste86b] R. H. Stern. Micro law:

Protecting semicustom chips. IEEE Micro, 6(1):72–76, January/February 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1986:MLRa

[Ste86c] R. H. Stern. Micro law:

Reverse engineering of chips.

1. the legislative background. IEEE Micro, 6(3):79–83, May/June 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1986:MLRb

[Ste86d] R. H. Stern. Micro law: Reverse engineering of chips. 2. A case example. IEEE Micro, 6(4):59–75, July/August 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1986:MLS

[Ste86e] R. H. Stern. Micro law: Software copyright developments. IEEE Micro, 6(6): 74–79, November/December 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1986:MLL

[Ste86f] R. H. Stern. Micro law: The look, feel, taste, and smell of software. IEEE Micro, 6 (2):64–65, March/April 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1986:BWR

[Ste86g] R. G. Stewart. Bus wars — response. IEEE Micro, 6(6): 5–??, November/December 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1986:MPP

[Ste86h] R. G. Stewart. Microstandards — promises, promises, promises. IEEE Micro, 6 (4):66–68, July/August 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1987:MLL

[Ste87a] R. H. Stern. Micro law: Legal mythology, or micromyths. IEEE Micro, 7(1):73–75, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1987:MLMb

[Ste87b] R. H. Stern. Micro law: Manufacturers disclaimers of liability. IEEE Micro, 7(6): 86–87, November/December 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1987:MLMa

[Ste87c] R. H. Stern. Micro law: Microcode revisited — further implications of the NEC vs Intel case — microcode, instruction sets, and compatibility. IEEE Micro, 7(2):81– 92, March/April 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1987:MLSa

[Ste87d] R. H. Stern. Micro law:

Software copyright developments. IEEE Micro, 7(3): 81–89, May/June 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1987:MLSb

[Ste87e] R. H. Stern. Micro law:

Software models of hardware. IEEE Micro, 7(5):85– 89, September/October 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1988:MLC

[Ste88a] R. H. Stern. Micro law: Copywriting screen displays. IEEE Micro, 8(4):4–5, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1988:MLE

[Ste88b] R. H. Stern. Micro law: On emulators and the procreation of porcupines. IEEE Micro, 8(5):6–7, September/October 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1988:MLP

[Ste88c] R. H. Stern. Micro law: Protection against piracy — reply. IEEE Micro, 8(3):88– 89, May/June 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1988:MLR

[Ste88d] R. H. Stern. Micro law: Reflections on Dirty Harry and dBASE. IEEE Micro, 8(1):70– 72, January/February 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1988:MLB

[Ste88e] R. H. Stern. Micro law: The Berne Convention — a bad idea whose time has come. IEEE Micro, 8(2):6–7, March/ April 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1989:MLAd

[Ste89a] R. H. Stern. Micro law:

Appropriate and inappropriate legal protection of user interfaces and screen displays. IV. screen display protection and policy concerns. IEEE Micro, 9(6):84–89, November/December 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1989:MLF

[Ste89b] R. H. Stern. Micro law: Framing prints, giving the Mona Lisa a moustache, speeding up video games, and marketing add-on software — a comment on the Mirage case. IEEE Micro, 9(2):8–??, March/April 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1989:MLAa

[Ste89c] Richard H. Stern. Micro law: Appropriate and inappropriate legal protection of user interfaces and screen displays. I. IEEE Micro, 9(3): 84–88, May/June 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1989:MLAb

[Ste89d] Richard H. Stern. Micro law:

Appropriate and inappropriate legal protection of user interfaces and screen displays. II. technical aspects of screen design raising legal policy issues. IEEE Micro, 9(4):7– 10, 92–94, July/August 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1989:MLAc

[Ste89e] Richard H. Stern. Micro law: Appropriate and inappropriate legal protection of user interfaces and screen displays. III. copyright law, the courts, and the copyright office. IEEE Micro, 9(5):8– 9, 75–79, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1989:MLP

[Ste89f] Richard H. Stern. Micro law: Protecting hardware against competition by copyrighting it as a compilation of data. IEEE Micro, 9(1):2–5, January/February 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1990:MLM

[Ste90a] R. H. Stern. Micro law: More on software patents. IEEE Micro, 10(4):7–9, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1990:MLPa

[Ste90b] R. H. Stern. Micro law: Professional ethics and the law. IEEE Micro, 10(3):83– 84, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1990:MLPb

[Ste90c] R. H. Stern. Micro law:

The paperback case. IEEE Micro, 10(5):7–10, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1990:MLPc

[Ste90d] R. H. Stern. Micro law:

The paperback case. 2. A nonliteral analysis. IEEE Micro, 10(6):39–41, November/December 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1990:MLA

[Ste90e] Richard H. Stern. Micro law:

Appropriate and inappropriate legal protection of user interfaces and screen displays. V. how different forms of copyright protection interact with policy. IEEE Micro, 10(1):79– 84, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1990:MLS

[Ste90f] Richard H. Stern. Micro law:

Software patents. IEEE Micro, 10(2):8–11, March/April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1990:FAHa

[Ste90g] Bob Stewart, editor. First Annual Hot Chips Symposium, Part 1, volume 10(1) of IEEE Micro. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1990:FAHb

[Ste90h] Bob Stewart, editor. First Annual Hot Chips Symposium, Part 2, volume 10(3) of IEEE Micro. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLCb

[Ste91a] R. H. Stern. Micro law: (C) — greater-than-software legal HLP. IEEE Micro, 11(3): 42–46, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLCa

[Ste91b] R. H. Stern. Micro law: Computer software rentals restricted in United States. IEEE Micro, 11(1):52, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLD

[Ste91c] R. H. Stern. Micro law:

Database system copyrights. IEEE Micro, 11(6):78–80, November/December 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLFb

[Ste91d] R. H. Stern. Micro law:

Fraud on the Copyright Office. IEEE Micro, 11(5):33– 34, September/October 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLI

[Ste91e] R. H. Stern. Micro law: IEEE and the Software Rental Act — reply. IEEE Micro, 11(3): 3–??, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLPa

[Ste91f] R. H. Stern. Micro law:

The paperback case. 3. misconceptions about functionality. IEEE Micro, 11(1):48–51, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLPb

[Ste91g] R. H. Stern. Micro law: The paperback case. 4. what’s really going on. IEEE Micro, 11 (2):30–33, March/April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1991:MLFa

[Ste91h] Richard H. Stern. Micro law: The first chip-layout copying case. IEEE Micro, 11(4):3–6, 94, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1992:MLC

[Ste92a] R. H. Stern. Micro law: Current developments in United States computer and electronics-industry tie-in law. IEEE Micro, 12(4):3–??, July/August 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1992:MLE

[Ste92b] R. H. Stern. Micro law: Engineers can be disqualified, too. IEEE Micro, 12(1):4–??, January/February 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1992:MLG

[Ste92c] R. H. Stern. Micro law: Game genie — copyrights and addons. IEEE Micro, 12(2):74– 79, March/April 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1992:MLN

[Ste92d] R. H. Stern. Micro law: No accolades for accolade court. IEEE Micro, 12(3):3–6, May/ June 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1992:MLP

[Ste92e] R. H. Stern. Micro law:

Penalties for reverse engineering. IEEE Micro, 12(5):2– 4, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1992:MLU

[Ste92f] R. H. Stern. Micro law: Unobserved demise of exhaustion doctrine. IEEE Micro, 12 (6):5–7, November/December 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1993:MLH

[Ste93a] R. H. Stern. Micro law:

Hewlett–Packard’s license finesses Intel’s patent in foundry deal. IEEE Micro, 13(4):64– 67, July/August 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1993:MLPb

[Ste93b] R. H. Stern. Micro law:

Patents and internationaltrade issues. IEEE Micro, 13(6):89–92, November/December 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1993:MLB

[Ste93c] Richard H. Stern. Micro law: a back door to protecting look and feel? IEEE Micro, 13 (2):74–76, March/April 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1993:MLGa

[Ste93d] Richard H. Stern. Micro law: a guardedly cheerful note — for a change (computer program copyright). IEEE Micro, 13(1):73–75, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1993:MLGb

[Ste93e] Richard H. Stern. Micro law: Glitches left in software copyright system. IEEE Micro, 13(3):75–77, May/June 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1993:MLPa

[Ste93f] Richard H. Stern. Micro law: Protecting industrial-property rights. IEEE Micro, 13(5): 2–3, 100, September/October 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1993:MLR

[Ste93g] Richard H. Stern. Micro law:

Reverse engineering for future compatibility. IEEE Micro, 13 (4):6–7, 76–79, July/August 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1994:MLSc

[Ste94a] R. H. Stern. Micro law: a short look at time bombs. IEEE Micro, 14(6):6–7, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1994:MLD

[Ste94b] Richard H. Stern. Micro law:

Disassembling object code — a misdeed. IEEE Micro, 14 (1):2–4, 78, January/February 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1994:MLSb

[Ste94c] Richard H. Stern. Micro law: Setting standards on the information superhighway. IEEE Micro, 14(4):4– 5, July/August 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1994:MLSa

[Ste94d] Richard H. Stern. Micro law: Sweat equity investments. IEEE Micro, 14(2):3– 4, 80, March/April 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1994:MLT

[Ste94e] Richard H. Stern. Micro law: Tilting at Gates’s windmill (Microsoft licensing). IEEE Micro, 14(5):5–6, 79– 80, September/October 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1994:MLU

[Ste94f] Richard H. Stern. Micro law: US intellectual property law can protect almost any technical advance in neural network design. IEEE Micro, 14 (3):5, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1995:MLF

[Ste95a] Richard H. Stern. Micro law: Fuzziness versus all or nothing. IEEE Micro, 15(3):7, 77–78, May/June 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1995:MLH

[Ste95b] Richard H. Stern. Micro law: Hauling manufacturers into the ITC (US International Trade Commission). IEEE Micro, 15(1):6–7, 80, January/February 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1995:MLM

[Ste95c]  Richard H. Stern.  Mi-

cro law: Microsoft and vaporware: “Sticking it to Phillippe”. IEEE Micro, 15 (2):6–7, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1995:MLP

[Ste95d] Richard H. Stern. Micro law:

The PTO on software patents. IEEE Micro, 15(4):2–3, 77– 78, July/August 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1995:MLW

[Ste95e] Richard H. Stern. Micro law: Winsocking the competition. IEEE Micro, 15(6):6– 7, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1996:MLA

[Ste96a] Richard H. Stern. Micro law:

Anti-knockoff article protection law. IEEE Micro, 16(4): 4–5, July/August 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1996:MLN

[Ste96b] Richard H. Stern. Micro law:

Net access — divvying up the pie. IEEE Micro, 16(3): 6–9, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1996:MLPc

[Ste96c] Richard H. Stern. Micro law: Patenting computerized methods of doing business. IEEE Micro, 16(6):4–6, 75, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1996:MLPa

[Ste96d] Richard H. Stern. Micro law: Patenting software, revisited. IEEE Micro, 16(2): 5, 77–79, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1996:MLPb

[Ste96e] Richard H. Stern. Micro law: Patents on selling via the Net — really? IEEE Micro, 16(5):6–7, 74–75, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1996:MLS

[Ste96f] Richard H. Stern. Micro law: Should a BB or net access provider be liable for copyright infringement when a user posts infringing material on a user newsgroup of forum? IEEE Micro, 16(1):7–9, 70–72, January/February 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1997:MLA

[Ste97a] R. H. Stern. Micro law: AOL:

essential for sending junk email? IEEE Micro, 17(2): 7–8, March/April 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1997:MLB

[Ste97b] R. H. Stern. Micro law:

Being right isn’t everything. IEEE Micro, 17 (4):77–78, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m4077.
pdf.

Stern:1997:MLC

[Ste97c] R. H. Stern. Micro law: Content pirates, beware! IEEE Micro, 17(3):7–9, 76–77, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://pascal.computer.
org/mi/books/mi1997/pdf/
m3007.pdf.

Stern:1997:MLE

[Ste97d] R. H. Stern. Micro law:

Episode 2, part 1: The browser wars. IEEE Micro, 17 (6):2–5, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1997/pdf/m6002.
                                        pdf.

Stern:1997:MLT

[Ste97e] R. H. Stern. Micro law: The tyranny of paradigms. IEEE Micro, 17(5):3–4, September/October 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi1997/
                                        pdf/m5003.pdf.

Stern:1997:MLS

[Ste97f]  Richard H. Stern.  Mi-

cro law: Shrink-wrap license restrictions — preempted? IEEE Micro, 17(1):75–78, January/February 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1998:MLM

[Ste98a] R. H. Stern. Micro law: Microsoft wins readers’ support — the author responds. IEEE Micro, 18(3):3, May/ June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1998:MLS

[Ste98b] R. H. Stern. Micro law: Should signals be patented? IEEE Micro, 18(2):6–8, March/ April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2006.
pdf.

Stern:1998:MLG

[Ste98c] R. H. Stern. Micro law:

The gnat versus the Borg. IEEE Micro, 18(1):6–8, January/February 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:1998:MLY

[Ste98d] R. H. Stern. Micro law: Y2K product liability. IEEE Micro, 18(5):7, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5007.
pdf.

Stern:1998:MLI

[Ste98e] Richard H. Stern. Micro law: Inviting participants in standard setting. IEEE Micro, 18(3):6–7, 79, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m3006.
pdf.

Stern:1998:MLR

[Ste98f] Richard H. Stern. Micro law: Restraints on technology advances. IEEE Micro, 18 (4):4–6, 78–81, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4004.
pdf.

Stern:1999:MLLa

[Ste99a] Richard H. Stern. Micro law: Licensing IP embodied in standards. IEEE Micro, 19 (4):7–8, 82–83, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4007.
pdf.

Stern:1999:MLLb

[Ste99b]  Richard H. Stern.  Mi-

cro law: Licensing IP embodied in standards, Part 2: ANSI position; the Web Standards Project position. IEEE Micro, 19(5):7–9, 81– 83, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m5007.
pdf.

Stern:1999:MLWb

[Ste99c] Richard H. Stern. Micro law: Web concerns. IEEE Micro, 19(2):6–7, March/April 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m2006.
pdf.

Stern:1999:MLWc

[Ste99d] Richard H. Stern. Micro law: When compliance with a standard gets too expensive. IEEE Micro, 19(6):12– 14, 86, November/December 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m6012.
pdf.

Stern:1999:MLWa

[Ste99e] Richard H. Stern. Micro law:

When elephants dance, mice watch out! IEEE Micro, 19 (1):6–7, 82, January/February 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m1006.
pdf.

Stern:2000:MLIc

[Ste00a] Richard H. Stern. Micro law: IP-related refusals to deal: Part 21/2: a postscript. IEEE Micro, 20(3):6–7, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3006.
pdf.

Stern:2000:MLIa

[Ste00b] Richard H. Stern. Micro law: IP-related refusals to deal: Part 1: Updating the Intel-Intergraph controversy. IEEE Micro, 20(1):9– 12, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1009.
pdf.

Stern:2000:MLIb

[Ste00c] Richard H. Stern. Micro law: IP-related refusals to deal: Part 2: Pretext and misconduct as standards. IEEE Micro, 20(2):8–11, 96, March/ April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m2008.
pdf.

Stern:2000:MLN

[Ste00d]  Richard H. Stern.  Mi-

cro law: Napster: a walking copyright infringement? IEEE Micro, 20(6):4–5, 95, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6004.
pdf.

Stern:2001:MLAa

[Ste01a] Richard H. Stern. Micro law: Amazon’s one-click patent loses its teeth. IEEE Micro, 21(2):7–10, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2007.
pdf.

Stern:2001:MLAb

[Ste01b]  Richard H. Stern.  Mi-

cro law: Another update on standardization skullduggery. IEEE Micro, 21(5):8– 10, September/October 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m5008abs.htm;
http://dlib.computer.org/
mi/books/mi2001/pdf/m5008.
pdf.

Stern:2001:MLMb

[Ste01c] Richard H. Stern. Micro law: Is mousetrapping unfair? IEEE Micro, 21(6): 72–77, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6072abs.htm;
http://dlib.computer.org/
mi/books/mi2001/pdf/m6072.
pdf.

Stern:2001:MLMa

[Ste01d] Richard H. Stern. Micro law:

More standardization skullduggery. IEEE Micro, 21 (4):12–15, 69, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/

books/mi2001/pdf/m4012.
pdf;   m4012abs.htm.

Stern:2001:MLP

[Ste01e] Richard H. Stern. Micro law: Preventing abuse of IEEE standards policy. IEEE Micro, 21(3):8–11, May/June 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m3008.

pdf. See letter [Kir01].

Stern:2001:MLW

[Ste01f] Richard H. Stern. Micro law: Who invented hyperlinks? IEEE Micro, 21 (1):8–10, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1008.
pdf.

Stern:2002:MLC

[Ste02a] Richard H. Stern. Micro law: Challenging search engines under copyright law: Part 1. IEEE Micro, 22(3):6–7, May/ June 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m3006.
pdf;   http://www.computer.
org/micro/mi2002/m3006abs.
htm.

Stern:2002:MLF

[Ste02b] Richard H. Stern. Micro law: FTC piles onto Rambus’ standardization skullduggery. IEEE Micro, 22 (4):6–7, 86–87, July/August 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m4006.
pdf;   http://www.computer.
org/micro/mi2002/m4006abs.
htm.

Stern:2002:MLG

[Ste02c] Richard H. Stern. Micro law: Is gatoring unfair or illegal? IEEE Micro, 22(1): 6–7, 92–93, January/February 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/m1006abs.htm;
http://dlib.computer.org/
mi/books/mi2002/pdf/m1006.
pdf.

Stern:2002:MLS

[Ste02d]  Richard H. Stern.  Mi-

cro law: Standardization and competitive advantage. IEEE Micro, 22(6):4–5, 73, November/December 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m6004.
pdf.

Stern:2003:MLU

[Ste03a] Richard H. Stern. Micro law: Unresolved legal questions about patents and standard setting. IEEE Micro, 23(5):5, 72–74, September/October 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2003/05/m5005.pdf.

Stern:2003:MLW

[Ste03b]  Richard H. Stern.  Mi-

cro law: Weird turn of events in continuing Rambus saga. IEEE Micro, 23 (1):76–80, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1076.
pdf;   http://www.computer.
org/micro/mi2003/m1076abs.
htm.

Stern:2004:MLCa

[Ste04a] Richard H. Stern. Micro law: Challenging search engines and pop-ups under copyright law: Part 2. IEEE Micro, 24(1):7, 71–73, January/February 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/01/m1007.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/01/m1007.
htm.

Stern:2004:MLCb

[Ste04b]  Richard H. Stern.  Mi-

cro law: Challenging search engines and pop-ups under copyright law: Part 3. IEEE Micro, 24(2):6, 70– 72, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/02/m2006.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/02/m2006.pdf.

Stern:2004:MLCc

[Ste04c] Richard H. Stern. Micro law:

Collecting patent infringement damages on unpatented products. IEEE Micro, 24(3): 6–7, May/June 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/03/m3006.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/03/m3006.
htm.

Stern:2004:MLF

[Ste04d] Richard H. Stern. Micro law: FTC turns back challenge on patent coverage. IEEE Micro, 24(4):7, 85– 86, July/August 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/04/m4007.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/04/m4007.
htm.

Stern:2004:MLV

[Ste04e] Richard H. Stern. Micro law: Vicarious liability for infringement. IEEE Micro, 24(5):6, 79–82, September/October 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/05/m5006.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/05/m5006.pdf.

Stern:2005:MLT

[Ste05a] Richard Stern. Micro law: Transnational electronic systems and patent infringement. IEEE Micro, 25(6): 85–88, November/December 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2005:MLF

[Ste05b] Richard H. Stern. Micro law: FTC cracks down on spyware and PC hijacking, but not true lies. IEEE Micro, 25(1):6–7, 100–101, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2005/01/m1006.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2005/01/m1006.pdf.

Stern:2005:MLS

[Ste05c] Richard H. Stern. Micro law: Standardization skullduggery update: UMTS standard. IEEE Micro, 25(4):73– 76, July/August 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2005:MLA

[Ste05d] Richard H. Stern. Micro law: The antitrust ghost in the standard-setting machine. IEEE Micro, 25(3):7–9, May/ June 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2006:MLC

[Ste06a] Richard Stern. Micro law:

Court dismisses “Copyright champion’s” source code copyright suit. IEEE Micro, 26(3): 88–90, May/June 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2006:MLN

[Ste06b] Richard Stern. Micro law:

New Jersey federal court holds Qualcomm’s unFRANDly acts no antitrust violation. IEEE Micro, 26(5):9, 84– 85, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2007:MLA

[Ste07a] Richard Stern. Micro law:

Antitrust division gives IEEE Standard setters the okay to ask patentees how RAND they are. IEEE Micro, 27(3):106– 109, May/June 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2007:MLC

[Ste07b] Richard Stern. Micro law: Coming down the home stretch in the Rambus standardization skullduggery saga: To levy or not to levy royalties. IEEE Micro, 27(2):80– 82, March/April 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2007:MLF

[Ste07c] Richard Stern. Micro law: Federal Appeals Court sees potential antitrust violation in standardization skullduggery. IEEE Micro, 27(5): 109–110, September/October 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2007:MLS

[Ste07d] Richard Stern. Micro law: Supreme Court to hear semiconductor chip patent “Exhaustion” case. IEEE Micro, 27(6):11–13, November/December 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2007:MLW

[Ste07e] Richard Stern. Micro law: West Coast Federal Appeals Court upholds Chip Protection Act violation finding. IEEE Micro, 27(1):124–126, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2008:MLA

[Ste08a] Richard Stern. Micro law:

AAI asks FTC to investigate RAND issues concerning Digital TV Standard. IEEE Micro, 28(3):72, 70– 71, May/June 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Stern:2008:MLU

[Ste08b] Richard Stern. Micro law:

An update on “Exhaustion”— Supreme Court decides Quanta case. IEEE Micro, 28(6): 57–54, November/December 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2008:MLF

[Ste08c] Richard Stern. Micro law:

FTC sues N-Data for violating standards commitment to IEEE. IEEE Micro, 28(2):66– 69, March/April 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2008:MLWa

[Ste08d] Richard Stern. Micro law:

What kinds of computersoftware-related advances (if any) are eligible for patents? Part I. IEEE Micro, 28(4):

96, 91–95, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/
                                        ;   http://www.adobe.com/
                                        products/acrobat/readstep2.
                                        html.

Stern:2008:MLWb

[Ste08e] Richard Stern. Micro law: What kinds of computersoftware-related advances (if any) are eligible for patents? Part II: The “Useful arts” requirement. IEEE Micro, 28(5):64–70, September/October 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2008:MN

[Ste08f]  Richard Stern.  Micro

news. IEEE Micro, 28(6):64, November/December 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2009:MLA

[Ste09a] Richard Stern. Micro law: All Bilski briefs filed and case set for oral argument. IEEE Micro, 29(6):72, November/December 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2009:MLI

[Ste09b] Richard Stern. Micro law: IEEE-USA tells Congress that patent reform is essential to economic recovery. IEEE Micro, 29(2):64–65, March/April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2009:MLE

[Ste09c] Richard H. Stern. Micro law:

An end to the Rambus skullduggery saga. IEEE Micro, 29(4):86, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2009:MLO

[Ste09d] Richard H. Stern. Micro law: One of the last updates on Rambus standardization skullduggery. IEEE Micro, 29(1):139–143, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2011:SSR

[Ste11] Richard H. Stern. Standardization skullduggery revisited. IEEE Micro, 31 (2):96–99, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2012:MLS

[Ste12] Richard H. Stern. Micro law:

Standardization skullduggery never ends: Apple v. Motorola. IEEE Micro, 32(2): 3–5, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2013:MTC

[Ste13]  Richard H. Stern.  Mi-

crosoft tells court that without FRAND, standard-setting would be “Blatant antitrust violation”. IEEE Micro, 33 (4):76–77, July/August 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stern:2014:AVCa

[Ste14a] Richard H. Stern. Alice v. CLS Bank: Are US businessmethod and software patents doomed? Part 1. IEEE Micro, 34(5):64–69, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050064-
abs.html.

Stern:2014:AVCb

[Ste14b] Richard H. Stern. Alice v. CLS Bank: Are US businessmethod and software patents doomed? Part 2. IEEE Micro, 34(6):98–c3, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060098-
abs.html.

Stern:2015:FCS

[Ste15a] Richard H. Stern. Federal Circuit speaks out on determining RAND royalties for standards. IEEE Micro, 35(1):57– 61, January/February 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010057-
abs.html.

Stern:2015:JDA

[Ste15b] Richard H. Stern. Justice Department agrees IEEE’s new RAND policy isn’t price fixing. IEEE Micro, 35(2):78– 84, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020078-
abs.html.

Stenstrom:2016:MWA

[Ste16] Per Stenstrom. 2015 Maurice Wilkes Award given to Christos Kozyrakis. IEEE Micro, 36(3):128–129, May/June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2016/03/mmi2016030128.
html.

Stern:2017:FASa

[Ste17a]  Richard H. Stern.  FTC

and Apple sue Qualcomm for cell phone standardization skullduggery: Part 1.

IEEE Micro, 37(2):81–89, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020081-
abs.html.

Stern:2017:FASb

[Ste17b] Richard H. Stern. FTC and Apple sue Qualcomm for cell phone standardization skullduggery, part 2: Apple’s claims. IEEE Micro, 37(4):72– 81, July/August 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040072-
abs.html.

Stern:2017:FAS

[Ste17c]  Richard H. Stern.  FTC

and Apple sue Qualcomm for cell phone standardization skullduggery, part 3: Determining SEP reasonable royalty. IEEE Micro, 37(5):61– 69, September/October 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050061-
abs.html.

Stern:2018:FAS

[Ste18]  Richard H. Stern.  FTC

and Apple sue Qualcomm for cell phone standardization skullduggery, part 4. IEEE Micro, 38(1):100–114, January/February 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/01/mmi2018010100-
abs.html.

Steck:2021:MAM

[Ste21] Randy Steck. The middleaged microprocessor. IEEE Micro, 41(6):22–28, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stitt:2011:FPG

[Sti11]  Greg Stitt.  Are field-

programmable gate arrays ready for the mainstream? IEEE Micro, 31(6):58–63, November/December 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stiles:2019:HSB

[Sti19] D. Stiles. The hardware security behind Azure Sphere. IEEE Micro, 39(2):20–28, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamura:1988:AMB

[STK88] Ken Sakamura, Kanehisa Tsurumi, and Hiro Kato. Applying the Mu-Btron bus to a music LAN. IEEE Micro, 8 (2):60–66, March/April 1988.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sadasivam:2017:IPP

[STKS17] Satish Kumar Sadasivam, Brian W. Thompto, Ron Kalla, and William J. Starke. IBM Power9 processor architecture. IEEE Micro, 37 (2):40–51, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020040-
abs.html.

Sauer:1992:EAE

[STL92] Anton Sauer, Jean Pierre Tual, and Robin La Fontaine. European activities for EDA standardization. IEEE Micro, 12(4):54–59, July/August 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shimada:2002:USI

[STM02] Shigeru Shimada, Masaaki Tanizaki, and Kishiko Maruyama. Ubiquitous spatial-information services using cell phones. IEEE Micro, 22(6):25–34, November/December 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m6025.
pdf;   http://www.computer.
org/micro/mi2002/m6025abs.
htm.

Stockton:1986:M

[Sto86] John F. Stockton. Megacells. IEEE Micro, 6(1):59–67, January/February 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stock:1990:LCC

[Sto90] S. J. Stock. Low-cost CAD drawings. IEEE Micro, 10 (4):77–78, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stockton:1994:PES

[Sto94] John F. Stockton. Portable electronic storage systems. IEEE Micro, 14(1):69–76, January/February 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Strumpen:1998:PFT

[Str98] Volker Strumpen. Portable and fault-tolerant software systems. IEEE Micro, 18(5): 22–32, September/October 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m5022.
pdf;   http://www.computer.
org/micro/mi1998/m5022abs.
htm.

Stadler:2001:DVS

[STR^{+}01] Manfred Stadler, Markus Thalmann, Thomas Röwer, Hubert Kaeslin, Norbert Felber, and Wolfgang Fichnter. Design and verification of a stack processor virtual component. IEEE Micro, 21(2):69–80, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2069.
pdf;   http://www.computer.
org/micro/mi2001/m2069abs.
htm.

Srinivasan:2013:HSD

[STR^{+}13] Sudharsanan Srinivasan, Yongbo Tang, Graham Read, Nadir Hossain, Di Liang, Stephen J. Sweeney, and John E. Bowers. Hybrid silicon devices for energy-efficient optical transmitters. IEEE Micro, 33 (1):22–31, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stormon:1992:GPC

[STS^{+}92] Charles D. Stormon, Nikos B. Troullinos, Edward M. Saleh, Abhijeet V. Chavan, Mark R. Brule, and John V. Oldfield. A general-purpose CMOS associative processor IC and system. IEEE Micro, 12(6): 68–78, November/December 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Starke:2021:IPP

[STSM21] W. J. Starke, B. W. Thompto, J. A. Stuecheli, and J. E. Moreira. IBM’s POWER10 processor. IEEE Micro, 41(2):7– 14, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sakamoto:2015:STB

[STT^{+}15] Toshitsugu Sakamoto, Yukihide Tsuji, Munehiro Tada, Hideki Makiyama, Takumi Hasegawa, Yoshiki Yamamoto, Shinobu Okanishi, Keiichi Maekawa, Naoki Banno, Makoto Miyamura, Koichiro Okamoto, Noriyuki Iguchi, Hidekazu Oda, Shiro Kamohara, Yasushi Yamagata, Nobuyuki Sugii, Hiromitsu Hada, and Yasuhiro Ogasahara. A silicon-on-thin-buried-oxide CMOS microcontroller with embedded atom-switch ROM. IEEE Micro, 35(6):13–23, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060013-
abs.html.

Surmann:1995:FRB

[SU95] Hartmut Surmann and Ansgar P. Ungering. Fuzzy-rule-based systems on general purpose processors. IEEE Micro, 15(4):40–48, July/August 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suito:2012:DRM

[SUF^{+}12] Kazutoshi Suito, Rikuhei Ueda, Kei Fujii, Takuma Kogo, Hiroki Matsutani, and Nobuyuki Yamasaki. The dependable responsive multithreaded processor for distributed real-time systems. IEEE Micro, 32(6):52–61, November/December 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sangiovanni-Vincentelli:2003:ESD

[SV03] Alberto Sangiovanni-Vincentelli. Electronic-system design in the automobile industry. IEEE Micro, 23(3):8–18, May/June 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/03/m3008abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/03/m3008.
                                        pdf.

Swaminathan:2021:HSC

[SV21] Karthik Swaminathan and Augusto Vega. Hardware specialization: From cell to heterogeneous microprocessors everywhere. IEEE Micro, 41(6):112–120, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Safa:2022:NNS

[SVA^{+}22] Ali Safa, Jonah Van Assche, Mark Daniel Alea, Francky Catthoor, and Georges G. E. Gielen. Neuromorphic nearsensor computing: From event-based sensing to edge learning. IEEE Micro, 42(6): 88–95, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Steffen:2001:TQC

[SVC01] Matthias Steffen, Lieven M. K. Vandersypen, and Isaac L. Chuang. Toward quantum computation: a five-qubit quantum processor. IEEE Micro, 21(2): 24–34, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2024.
pdf;   http://www.computer.
org/micro/mi2001/m2024abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Sangiovanni-Vincentelli:2003:GEI

[SVL03] Alberto Sangiovanni-Vincentelli and Luciano Lavagno. Guest Editors’ introduction: Trends and directions in microelectronics. IEEE Micro, 23(3): 6–7, May/June 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.

                                        computer.org/dl/mags/mi/
                                        2003/03/m3006.pdf.

Sardashti:2014:DCC

[SW14] Somayeh Sardashti and David A. Wood. Decoupled compressed cache: Exploiting spatial locality for energy optimization. IEEE Micro, 34(3):91– 99, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Swanson:2019:RFS

[Swa19] S. Swanson. Redesigning file systems for nonvolatile main memory. IEEE Micro, 39 (1):62–64, January/February 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Salapura:2006:EWP

[SWG06] Valentina Salapura, Robert Walkup, and Alan Gara. Exploiting workload parallelism for performance and power optimization in Blue Gene. IEEE Micro, 26(5):67– 81, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Saggese:2005:ESS

[SWK^{+}05] Giacinto P. Saggese, Nicholas J. Wang, Zbigniew T. Kalbarczyk, Sanjay J. Patel, and Ravishankar K. Iyer. An experimental study of soft errors in microprocessors. IEEE Micro, 25(6):30–39, November/December 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sibai:1990:PUM

[SWL90] F. N. Sibai, K. L. Watson, and Mi Lu. A parallel unification machine. IEEE Micro, 10 (4):21–33, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Seong:2011:SRP

[SWL11] Nak Hee Seong, Dong Hyuk Woo, and Hsien-Hsin S. Lee. Security refresh: Protecting phase-change memory against malicious wear out. IEEE Micro, 31(1):119–127, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sanamrad:1987:HSA

[SWM87] M. A. Sanamrad, K. Wada, and H. Matsumoto. A hardware syntactic analysis processor. IEEE Micro, 7(4):73– 80, July/August 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stephenson:2020:PDM

[SWM^{+}20] A. Stephenson, M. Willsey, J. McBride, S. Newman, B. Nguyen, C. Takahashi, K. Strauss, and L. Ceze. PurpleDrop: a digital microfluidicsbased platform for hybrid molecular–electronics applications. IEEE Micro, 40(5):76– 86, September/October 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sherwood:2006:GEI

[SY06] Timothy Sherwood and Joshua J. Yi. Guest Editors’ introduction: Computer architecture simulation and modeling. IEEE Micro, 26(4):5– 7, July/August 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2006/04/m4005.pdf.

Skarlatos:2020:MEM

[SYG^{+}20] D. Skarlatos, M. Yan, B. Gopireddy, R. Sprabery, J. Torrellas, and C. W. Fletcher. MicroScope: Enabling microarchitectural replay attacks. IEEE Micro, 40(3):91–98, May/June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shimamoto:2011:ACT

[SYKM11] Hiroshi Shimamoto, Takayuki Yamashita, Misao Kubota, and Hirotaka Maruyama. Advanced camera technologies for broadcasting. IEEE Micro, 31(6):51–57, November/December 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sampson:2014:SMA

[SYW^{+}14] Richard Sampson, Ming Yang, Siyuan Wei, Chaitali Chakrabarti, and Thomas F. Wenisch. Sonic Millip3De: An architecture for handheld 3D ultrasound. IEEE Micro, 34 (3):100–108, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Suzuki:2011:HTL

[SYY^{+}11] Tomoya Suzuki, Hideki Yamada, Toshiyuki Yamagishi, Daisuke Takeda, Koji Horisaki, Toshio Fujisawa, Yasuo Unekawa, Tom Vander Aa, and Liesbet Van der Perre. High-throughput, lowpower software-defined radio using reconfigurable processors. IEEE Micro, 31(6): 19–28, November/December 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stigall:1982:PSB

[SZH82] P. D. Stigall, R. E. Ziemer, and L. Hudec. A performance study of 16-bit microcomputer-implemented FFT algorithms. IEEE Micro, 2(4):61–66, October/December 1982. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stigall:1981:PSM

[SZP81] Paul D. Stigall, Rodger E.

Ziemer, and Van T. Pham. A performance study of a microcomputer-implemented FSK receiver. IEEE Micro, 1(1):43–51, January/March 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shang:2001:CDI

[SZZ01] Zhao Shang, Zhichun Zhu, and Xiaodong Zhang. Cached DRAM for ILP processor memory access latency reduction. IEEE Micro, 21 (4):22–32, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4022.
pdf;   m4022abs.htm.

Tiwari:2016:ASS

[TA16] Mohit Tiwari and Todd Austin. On architectural support for systems security. IEEE Micro, 36(5):6– 7, September/October 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050006-
abs.html.

Tabak:1984:DAL

[Tab84] Daniel Tabak. Dynamic architecture and LSI modular computer systems — a review of “Designing and Programming Modern Computers and Systems, Volume I — LSI Modular Computer Systems”. IEEE Micro, 4(2):48– 66, March/April 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tabak:1991:IP

[Tab91] D. Tabak. Improving the product. IEEE Micro, 11(2): 2, March/April 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tanaka:2021:DDL

[TAI^{+}21] K. Tanaka, Y. Arikawa, T. Ito, K. Morita, N. Nemoto, K. Terada, J. Teramoto, and T. Sakamoto. Distributed deep learning with GPU-FPGA heterogeneous computing. IEEE Micro, 41(1):15–22, January/February 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Talia:1993:MRS

[Tal93] D. Talia. Message-routing systems for transputer-based multicomputers. IEEE Micro, 13(3):62–72, May/June 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tuck:2009:SSE

[TATC09] James Tuck, Wonsun Ahn, Josep Torrellas, and Luis Ceze. SoftSig: Softwareexposed hardware signatures for code analysis and optimization. IEEE Micro, 29 (1):84–95, January/February 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Taub:1984:ACA

[Tau84] D. M. Taub. Arbitration and control acquisition in the proposed IEEE 896 Futurebus. IEEE Micro, 4(4):28– 41, July/August 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Taub:1986:BW

[Tau86]  M. Taub.  Bus wars.

IEEE Micro, 6(6):5, November/December 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Taub:1987:ICA

[Tau87] D. Matthew Taub. Improved control acquisition scheme for the IEEE 896 Futurebus. IEEE Micro, 7(3):52– 62, May/June 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Taylor:2013:LND

[Tay13] Michael B. Taylor. A landscape of the new dark silicon design regime. IEEE Micro, 33(5):8–19, September/October 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Turakhia:2019:DGC

[TBD19] Y. Turakhia, G. Bejerano, and W. J. Dally. Darwin: A genomics coprocessor. IEEE Micro, 39(3):29–37, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Trichina:2001:SCH

[TBDL01] Elena Trichina, Marco Bucci, Domenico De Seta, and Raimondo Luzzi. Supplemental cryptographic hardware for smart cards. IEEE Micro, 21(6):26–35, November/December 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2001/
m6026abs.htm;   http://
dlib.computer.org/mi/books/
mi2001/pdf/m6026.pdf.

Temam:2015:ACD

[TC15] Olivier Temam and Luis Ceze.

Alternative computing designs and technologies. IEEE Micro, 35(5):4–5, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050004.
html.

Tremblay:2000:MAS

[TCC^{+}00] Marc Tremblay, Jeffrey Chan, Shailender Chaudhry, Andrew W. Conigliaro, and Shing Sheung Tse. The MAJC architecture: a synthesis of parallelism and scalability. IEEE Micro, 20(6): 12–25, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m6012.
pdf;   http://www.computer.
org/micro/mi2000/m6012abs.
htm.

Takahashi:2005:PCD

[TCD^{+}05] Osamu Takahashi, Scott Cottier, Sang H. Dhong, Brian Flachs, and Joel Silberman. Power-conscious design of the CELL processor’s synergistic processor element. IEEE Micro, 25(5):10–18, September/October 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Torralba:1996:FLB

[TCF96] Antonio J. Torralba, Jorge Chavez, and L. G. Franquelo. Fuzzy-logic-based analog design tools combining fuzzy logic with conventional approaches to automate difficult analog design tasks. IEEE Micro, 16(4):60–68, July/August 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Teachey:1982:SRX

[Tea82] R. D. Teachey. Square-root-X comparison — new results discovered. IEEE Micro, 2(4):

5–6, October/December 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tarafdar:2018:GFS

[TES^{+}18] Naif Tarafdar, Nariman Eskandari, Varun Sharma, Charles Lo, and Paul Chow. Galapagos: A full stack approach to FPGA integration in the cloud. IEEE Micro, 38(6):18–24, November/December 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08541106-abs.
html.

Tang:2020:OOS

[TGC^{+}20] X. Tang, E. Giacomin, B. Chauviere, A. Alacchi, and P. Gaillardon. OpenFPGA: An open-source framework for agile prototyping customizable FPGAs. IEEE Micro, 40 (4):41–48, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tsoukarellas:1995:STR

[TGE95] Manthos A. Tsoukarellas, Vasilis C. Gerogiannis, and Kostis D. Economides. Systematically testing a realtime operating system. IEEE Micro, 15(5):50–60, September/October 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Thakkar:1988:BMS

[TGF88] Shreekant Thakkar, Paul Gifford, and Garay Fielland. The Balance multiprocessor system. IEEE Micro, 8(1):57–69, January/February 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tziantzioulis:2018:TAF

[THC18] Georgios Tziantzioulis, Nikos Hardavellas, and Simone Campanoni. Temporal approximate function memoization. IEEE Micro, 38(4):60– 70, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        mi/2018/04/mmi2018040060-
                                        abs.html.

Thomborson:1992:VSD

[Tho92]  Clark Thomborson.  The

V.42bis standard for datacompressing modems. IEEE Micro, 12(5):41–53, September/October 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Talati:2019:CCO

[THP^{+}19] N. Talati, H. Ha, B. Perach, R. Ronen, and S. Kvatinsky. CONCEPT: A columnoriented memory controller for efficient memory and PIM operations in RRAM. IEEE Micro, 39(1):33–43, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Talla:2004:APD

[THT^{+}04] Deepu Talla, Ching-Yu Hung, Raj Talluri, Frank Brill, David Smith, David Brier, Bruce Xiong, and Derek Huynh. Anatomy of a portable digital mediaprocessor. IEEE Micro, 24(2):32– 39, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/02/m2032abs.htm;
http://csdl.computer.org/
dl/mags/mi/2004/02/m2032.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/
m2032.pdf.

Tanaka:2013:HBO

[TIT^{+}13] Kazuhiro Tanaka, Satoshi Ide, Yukito Tsunoda, Takashi Shiraishi, Takatoshi Yagisawa, Tadashi Ikeuchi, Tsuyoshi Yamamoto, and Tomohiro Ishihara. High-bandwidth optical interconnect technologies for next-generation server systems. IEEE Micro, 33(1): 6–13, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tsiokanos:2021:EEH

[TK21] I. Tsiokanos and G. Karakonstantis. ExHero: Execution history-aware error-rate estimation in pipelined designs. IEEE Micro, 41(1):61– 68, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tamura:2014:ESC

[TKI^{+}14] Hikaru Tamura, Kiyoshi Kato, Takahiko Ishizu, Wataru Uesugi, Atsuo Isobe, Naoaki Tsutsui, Yasutaka Suzuki, Yutaka Okazaki, Yukio Maehashi, Jun Koyama, Yoshitaka Yamamoto, Shunpei Yamazaki, Masahiro Fujita, James Myers, and Pekka Korpinen. Embedded SRAM and Cortex-M0 core using a 60-nm crystalline oxide semiconductor. IEEE Micro, 34(6):42–53, November/December 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060042-
abs.html.

Taylor:2002:RMC

[TKM^{+}02] Michael Bedford Taylor, Jason Kim, Jason Miller, David Wentzlaff, Fae Ghodrat, Ben Greenwald, Henry Hoffman, Paul Johnson, Jae-Wook Lee, Walter Lee, Albert Ma, Arvind Saraf, Mark Seneski, Nathan Shnidman, Volker Strumpen, Matt Frank, Saman Amarasinghe, and Anant Agarwal. The Raw microprocessor: a computational fabric for software circuits and general-purpose programs. IEEE Micro, 22 (2):25–35, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2025.
pdf;   http://www.computer.
org/micro/mi2002/m2025abs.
htm.

Tsoutsouras:2022:LMT

[TKS^{+}22] Vasileios Tsoutsouras, Orestis Kaparounakis, Chatura Samarakoon, Bilgesu Bilgin, James Meech, Jan Heck, and Phillip Stanley-Marbell. The Laplace microarchitecture for tracking data uncertainty. IEEE Micro, 42(4):78–86, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Trippel:2019:SVA

[TLM19] C. Trippel, D. Lustig, and M. Martonosi. Security verification via automatic hardware-aware exploit synthesis: The CheckMate approach. IEEE Micro, 39(3): 84–93, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tiwari:2010:GLI

[TLW^{+}10] Mohit Tiwari, Xun Li, Hassan M. G. Wassel, Bita Mazloom, Shashidhar Mysore, Frederic T. Chong, and Timothy Sherwood. Gate-level information-flow tracking for secure architectures. IEEE Micro, 30(1):92–100, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tan:2004:OBC

[TLYL04] Zhangxi Tan, Chuang Lin, Hao Yin, and Bo Li. Optimization and benchmark of cryptographic algorithms on network processors. IEEE Micro, 24(5):55–69, September/October 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/05/m5055.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/05/m5055.pdf.

Tomikawa:1981:IJP

[TM81] Takehiko Tomikawa and Kinji Matsumoto. Ink jet printing of Japanese Kanji characters. IEEE Micro, 1(2):39–42, 44–46, April/June 1981. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Taylor:1982:BSA

[TM82] G. D. Taylor and S. J. Mc-Cormick. Best starting approximations. IEEE Micro, 2(2):64–65, April/June 1982.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tomasevic:1994:HACb

[TM94a] M. Tomasevic and V. Milutinovic. Hardware approaches to cache coherence in sharedmemory multiprocessors, part 2. IEEE Micro, 14(6):61–66, November/December 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tomasevic:1994:HACa

[TM94b] Milo Tomašević and Veljko Milutinović. Hardware approaches to cache coherence in shared-memory multiprocessors, part 1. IEEE Micro, 14(5):52–59, September/October 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Thottethodi:2014:TPC

[TM14] Mithuna S. Thottethodi and Shubu Mukherjee. Top picks from the 2013 computer architecture conferences. IEEE Micro, 34(3):4–7, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Tokusashi:2017:MNC

[TM17] Yuta Tokusashi and Hiroki Matsutani. Multilevel NoSQL cache combining in-NIC and in-kernel approaches. IEEE Micro, 37(5):44–51, September/October 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/05/mmi2017050044-
abs.html.

Tomesh:2021:QC

[TM21] Teague Tomesh and Margaret Martonosi. Quantum codesign. IEEE Micro, 41(5):33– 40, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tunali:2018:DTL

[TMA18] Onur Tunali, M. Ceylan Morgul, and Mustafa Altun. Defect-tolerant logic synthesis for memristor crossbars with performance evaluation. IEEE Micro, 38(5):22– 31, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050022-

abs.html. See erratum [Ano18d].

Taghizadeh:1994:DDO

[TMBT94] M. R. Taghizadeh, J. M. Miller, P. Blair, and F. A. P. Tooley. Developing diffractive optics for optical computing. IEEE Micro, 14(6): 10–19, November/December 1994. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tan:2013:OIH

[TMJ13] Michael R. T. Tan, Moray McLaren, and Norman P. Jouppi. Optical interconnects for high-performance computing systems. IEEE Micro, 33 (1):14–21, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Trippel:2018:FSM

[TML^{+}18] Caroline Trippel, Yatin A.

Manerkar, Daniel Lustig, Michael Pellauer, and Margaret Martonosi. Full-stack memory model verification with TriCheck. IEEE Micro, 38(3):58–68, May/June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030058-
abs.html.

Teodorescu:2006:SPE

[TNT06] Radu Teodorescu, Jun Nakano, and Josep Torrellas. SWICH: a prototype for efficient cachelevel checkpointing and rollback. IEEE Micro, 26(5):28– 40, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tremblay:1996:UFI

[TO96] Marc Tremblay and J. Michael O’Connor. UltraSparc I: a four-issue processor supporting multimedia: Combining on-chip multimedia instructions with a highperformance, four-issue architecture. IEEE Micro, 16 (2):42–50, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Tremblay:1996:VSN

[TONH96] Marc Tremblay, J. Michael O’Connor, Venkatesh Narayanan, and Liang He. VIS speeds new media processing — enhancing conventional RISC instruction sets to significantly accelerate media-processing algorithms. IEEE Micro, 16 (4):10–20, July/August 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Torrellas:2006:GEI

[Tor06]  Josep Torrellas.  Guest

editor’s introduction: Micro’s top picks from Microarchitecture Conferences. IEEE Micro, 26(1):8–9, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2006/01/m1008.pdf.

Torrellas:2012:ISC

[Tor12] Josep Torrellas. 2012 International Symposium on Computer Architecture Influential Paper Award. IEEE Micro, 32 (5):4–5, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tabatabaei:2010:SMO

[TP10] Sassan Tabatabaei and Aaron Partridge. Silicon MEMS oscillators for high-speed digital systems. IEEE Micro, 30 (2):80–89, March/April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Treleaven:1989:VAN

[TPV89] Philip Treleaven, Marco Pacheco, and Marley Vellasco. VLSI architectures for neural networks. IEEE Micro, 9(6): 8–27, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Troster:1998:GEI

[Trö98] Gerhard Tröster. Guest Editor’s introduction: New route in system integration: Chippackage codesign. IEEE Micro, 18(4):7–9, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m4007.
pdf.

Tan:2009:HSO

[TRY^{+}09] Michael R. T. Tan, Paul Rosenberg, Jong-Souk Yeo, Moray McLaren, Sagi Mathai, Terry Morris, Huei Pei Kuo, Joseph Straznicky, Norman P. Jouppi, and Shih-Yuan Wang. A high-speed optical multidrop bus for computer interconnections. IEEE Micro, 29 (4):62–73, July/August 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Takada:1991:IMA

[TS91] Hiroaki Takada and Ken Sakamura. Itron-MP — an adaptive real-time kernel specification for shared-memory multiprocessor systems. IEEE Micro, 11(4):24–27, 78–85, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Takada:1995:ISS

[TS95] Hiroaki Takada and Ken Sakamura. μITRON for small-scale embedded systems. IEEE Micro, 15(6):46–54, November/December 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tan:2006:ABS

[TS06] Lin Tan and Timothy Sherwood. Architectures for bitsplit string scanning in intrusion detection. IEEE Micro, 26(1):110–117, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Taylor:2013:DSG

[TS13] Michael B. Taylor and Steven Swanson. Dark silicon [Guest Editors’ introduction]. IEEE Micro, 33(5):6–7, September/October 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tredennick:2014:PRS

[TS14] Nick Tredennick and Brion Shimamoto. Prospects for reconfigurable systems. IEEE Micro, 34(1):72–78, January/ February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Temucin:2022:ADL

[TSA^{+}22] Yiltan Hassan Temuçin, Amir Hossein Sojoodi, Pedram Alizadeh, Benjamin Kitor, and Ahmad Afsahi. Accelerating deep learning using interconnect-aware UCX communication for MPI collectives. IEEE Micro, 42(2):68– 76, March/April 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tsai:2021:LST

[TSFS21] Po-An Tsai, Andres Sanchez, Christopher W. Fletcher, and Daniel Sanchez. Leaking secrets through compressed caches. IEEE Micro, 41(3): 27–33, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Thomas:2023:EBE

[TSMS23] Kyle Thomas, Muhammad Santriaji, David Mohaisen, and Yan Solihin. Exploration of bitflip’s effect on deep neural network accuracy in plaintext and ciphertext. IEEE Micro, 43(5):24– 34, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tandler:2002:RMT

[TSP02] Peter Tandler, Norbert Streitz, and Thorsten Prante. Roomware-moving toward ubiquitous computers. IEEE Micro, 22(6):36–47, November/December 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2002/
pdf/m6036.pdf;   http://
www.computer.org/micro/
mi2002/m6036abs.htm.

Tang:2018:MSA

[TSS18] Adrian Tang, Simha Sethumadhavan, and Salvatore Stolfo. Motivating securityaware energy management. IEEE Micro, 38(3):98–106, May/June 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030098-
abs.html.

Talpes:2020:CST

[TSV^{+}20] E. Talpes, D. D. Sarma, G. Venkataramanan, P. Bannon, B. McGee, B. Floering, A. Jalote, C. Hsiong, S. Arora, A. Gorti, and G. S. Sachdev. Compute solution for Tesla’s full self-driving computer. IEEE Micro, 40 (2):25–35, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tremaine:2001:PIM

[TSW^{+}01] R. Brett Tremaine, T. Basil Smith, Mike Wazlowski, David Har, Kwok-Ken Mak, and Sujith Arramreddy. Pinnacle: IBM MXT in a memory controller chip. IEEE Micro, 21(2):56–68, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m2056.
pdf;   http://www.computer.
org/micro/mi2001/m2056abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Talpes:2023:MDT

[TSW^{+}23] Emil Talpes, Debjit Das Sarma, Doug Williams, Sahil Arora, Thomas Kunjan, Benjamin Floering, Ankit Jalote, Christopher Hsiong, Chandrasekhar Poorna, Vaidehi Samant, John Sicilia, Anantha Kumar Nivarti, Raghuvir Ramachandran, Tim Fischer, Ben Herzberg, Bill McGee, Ganesh Venkataramanan, and Pete Banon. The microarchitecture of DOJO, Tesla’s exascale computer. IEEE Micro, 43(3):31–39, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tseng:2012:ERC

[TT12] Hung-Wei Tseng and Dean M. Tullsen. Eliminating redundant computation and exposing parallelism through datatriggered threads. IEEE Micro, 32(3):38–47, May/June 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tombs:1996:PFL

[TTF96] Jon Tombs, Antonio J. Torralha, and Leopoldo G. Franguelo. A PWM fuzzy logic controller: Using pulse width modulation to improve fuzzy digital controller design. IEEE Micro, 16(5):68– 71, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tual:1999:MGA

[Tua99] Jean-Pierre P. Tual. MASSC: a generic architecture for multiapplication smart cards. IEEE Micro, 19(5):52–61, September/October 1999. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1999/
pdf/m5052.pdf;   http://
www.computer.org/micro/
mi1999/m5052abs.htm.

Takaragi:2001:USI

[TUI^{+}01] Kazuo Takaragi, Mitsuo Usami, Ryo Imura, Rei Itsuki, and Tsuneo Satoh. An ultra small individual recognition security chip. IEEE Micro, 21(6):43–49, November/December 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/m6043abs.htm;
http://dlib.computer.org/
mi/books/mi2001/pdf/m6043.
pdf.

Taram:2019:CSD

[TVT19] M. Taram, A. Venkat, and D. M. Tullsen. Contextsensitive decoding: Ondemand microcode customization for security and energy management. IEEE Micro, 39(3):75–83, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tzimpragos:2021:TCS

[TVV^{+}21] Georgios Tzimpragos, Jennifer Volk, Dilip Vasudevan, Nestan Tsiskaridze, George Michelogiannakis, Advait Madhavan, John Shalf, and Timothy Sherwood. Temporal computing with superconductors. IEEE Micro, 41 (3):71–79, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tredennick:2000:MVP

[TW00] Nick Tredennick and Steven J. Wallach. Micro view: Predicting the future. IEEE Micro, 20(4):88, 87, July/August 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m4088.
pdf.

Takata:1999:DMM

[TWN^{+}99] Hidehiro Takata, Tetsuya Watanabe, Tetsuo Nakajima, Takashi Takagaki, Hisakazu Sato, Atsushi Mohri, Akira Yamada, Toshiki Kanamoto, Yoshio Matsuda, Shuhei Iwade, and Yasutaka Horiba. The D30V/MPEG multimedia processor. IEEE Micro, 19(4):38–47, July/August 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1999/pdf/m4038.
pdf;   http://www.computer.
org/micro/mi1999/m4038abs.
htm.

Tabachnick:1981:SCS

[TZMVLN81] Ritchie L. Tabachnick, Paul J. A. Zsombor-Murray, Louis J. Vroomen, and Tho Le-Ngoc. Sequence controllers with standard hardware and custom firmware. IEEE Micro, 1(2):9–25, April/June 1981. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Uchiyama:1993:GSM

[UAN^{+}93] Kunio Uchiyama, Fumio Arakawa, Susumu Narita, Hirokazu Aoki, Ikuya Kawasaki, Shigezumi Matsui, Mitsuyoshi Yamamoto, Norio Nakagawa, and Ikuo Kudo. The Gmicro/ 500 superscalar microprocessor with branch buffers. IEEE Micro, 13(5):12–22, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Uchiyama:2005:GEI

[UB05] Kunio Uchiyama and Pradip Bose. Guest Editors’ introduction: Energy-efficient design. IEEE Micro, 25(5):6– 9, September/October 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/05/m5006.pdf.

Undy:1994:LCG

[UBH^{+}94] Steve Undy, Mick Bass, Dave Hollenbeck, Wayne Kever, and Larry Thayer. A lowcost graphics and multimedia workstation chip set. IEEE Micro, 14(2):10–22, March/ April 1994. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic).

Unger:1982:OSZ

[UBL^{+}82] B. Unger, D. Bidulock, G. Lomow, P. Belanger, C. Hankins, and N. Jain. An OASIS simulation of the ZNET microcomputer network. IEEE Micro, 2(3):70–84, July/September 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ungerer:2010:MME

[UCS^{+}10] Theo Ungerer, Francisco Cazorla, Pascal Sainrat, Guillem Bernat, Zlatko Petrov, Christine Rochange, Eduardo Quinones, Mike Gerdes, Marco Paolieri, Julian Wolf, Hugues Casse, Sascha Uhrig, Irakli Guliashvili, Michael Houston, Floria Kluge, Stefan Metzlaff, and Jorg Mische. Merasa: Multicore execution of hard real-time applications supporting analyzability. IEEE Micro, 30(5):66– 75, September/October 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Uht:2000:BRC

[ULS^{+}00] Augustus K. Uht, Jien-Chung Lo, Ying Sun, James C. Daly, and James Kowalski. Building real computer systems. IEEE Micro, 20(3):48–56, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3048.
pdf;   http://www.computer.
org/micro/mi2000/m3048abs.
htm.

Updegrove:1993:FFO

[Upd93] Andrew Updegrove. Forming, funding, and operating standard-setting consortia. IEEE Micro, 13(6): 52–61, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Urquhart:1997:GEI

[Urq97] Ken Urquhart. Guest editor’s introduction: Java’s open future. IEEE Micro, 17(3):10–13, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3010.
pdf.

UlMustafa:2023:PMS

[US23] Naveed Ul Mustafa and Yan Solihin. Persistent memory security threats to interprocess isolation. IEEE Micro, 43(5): 16–23, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Usselmann:1991:OAR

[Uss91]  R. Usselmann.  On an

open architecture — reply. IEEE Micro, 11(6):2, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Unsal:2006:IPV

[UTB^{+}06] Osman S. Unsal, James W. Tschanz, Keith Bowman, Vivek De, Xavier Vera, Antonio González, and Oguz Ergin. Impact of parameter variations on circuits and microarchitecture. IEEE Micro, 26(6):30–39, November/December 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vachss:1987:CMF

[Vac87] Raymond Vachss. The Cordic magnification function. IEEE Micro, 7(5):83–84, September/October 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vinnakota:2021:ODS

[VAD^{+}21] B. Vinnakota, I. Agarwal, K. Drucker, D. Jani, G. Miller, M. Mittal, and R. Wang. The Open Domain-Specific Architecture. IEEE Micro, 41 (1):30–36, January/February 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vahdat:2010:SND

[VAFF^{+}10] Amin Vahdat, Mohammad Al-Fares, Nathan Farrington, Radhika Niranjan Mysore, George Porter, and Sivasankar Radhakrishnan. Scale-out networking in the data center. IEEE Micro, 30(4):29– 41, July/August 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

VanSingel:2021:RAS

[Van21] Steven Van Singel. The Renesas automotive story in the history of the microprocessor. IEEE Micro, 41(6):107–108, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

VonEicken:1995:LLC

[VBB95] T. Von Eicken, A. Basu, and V. Buch. Low-latency communication over ATM networks using active messages. IEEE Micro, 15(1):46–53, January/February 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vega:2014:SSH

[VBB14] Augusto Vega, Alper Buyuktosunoglu, and Pradip Bose. Special series on harsh chips [Guest Editors’ introduction]. IEEE Micro, 34(6):6– 7, November/December 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060006.
html.

Vasiljevic:2021:CSS

[VBC^{+}21] J. Vasiljevic, L. Bajic, D. Capalija, S. Sokorac, D. Ignjatovic, L. Bajic, M. Trajkovic, I. Hamer, I. Matosevic, A. Cejkov, U. Aydonat, T. Zhou, S. Z. Gilani, A. Paiva, J. Chu, D. Maksimovic, S. A. Chin, Z. Moudallal, A. Rakhmati, S. Nijjar, A. Bhullar, B. Drazic, C. Lee, J. Sun, K. M. Kwong, J. Connolly, M. Dooley, H. Farooq, J. Y. T. Chen, M. Walker, K. Dabiri, K. Mabee, R. S. Lal, N. Rajatheva, R. Retnamma, S. Karodi, D. Rosen, E. Munoz, A. Lewycky, A. Knezevic, R. Kim, A. Rui, A. Drouillard, and D. Thompson. Compute substrate for Software 2.0. IEEE Micro, 41 (2):50–55, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

vonHanxleden:1998:CAS

[vBK98] Reinhard von Hanxleden, Ali Botorabi, and Slawomir Kupczyk. A codesign approach for safety-critical automotive applications. IEEE Micro, 18(5):66–79, September/October 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1998/
pdf/m5066.pdf;   http://
www.computer.org/micro/
mi1998/m5066abs.htm.

Vuduc:2011:PWG

[VC11] Richard Vuduc and Kent Czechowski. Prolegomena: What GPU computing means for high-end systems. IEEE Micro, 31(4):74– 78, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Venkataramani:2016:DHC

[VCD16] Guru Venkataramani, Jie Chen, and Milos Doroslovacki. Detecting hardware covert timing channels. IEEE Micro, 36(5):17–27, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050017-
abs.html.

VanBiesbrouck:2006:ESS

[VCE06] Michael Van Biesbrouck, Brad Calder, and Lieven Eeckhout. Efficient sampling startup for SimPoint. IEEE Micro, 26 (4):32–42, July/August 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Valamehr:2013:IRM

[VCK^{+}13] Jonathan Kaveh Valamehr, Melissa Chase, Seny Kamara, Andrew Putnam, Daniel Shumow, Vinod Vaikuntanathan, and Timothy Sherwood. Inspection-resistant memory architectures. IEEE Micro, 33(3):48–56, May/ June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Venkataramani:2019:DCE

[VCS^{+}19] S. Venkataramani, J. Choi, V. Srinivasan, W. Wang, J. Zhang, M. Schaal, M. J. Serrano, K. Ishizaki, H. Inoue, E. Ogawa, M. Ohara, L. Chang, and K. Gopalakrishnan. DeepTools: Compiler and execution runtime extensions for RaPiD AI accelerator. IEEE Micro, 39(5): 102–111, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vetter:2017:APM

[VDC17] Jeffrey S. Vetter, Erik P. DeBenedictis, and Thomas M. Conte. Architectures for the post-Moore era. IEEE Micro, 37(4):6–8, July/August 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/04/mmi2017040006.
html.

vanderHoeven:1990:MHL

[vdDD90] A. J. van der Hoeven, A. A. J. de Lange, E. F. Deprettere, and P. M. Dewilde. A model for the high-level description and simulation of VLSI networks. IEEE Micro, 10(4):41– 48, July/August 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

VanErtvelde:2010:WRG

[VE10] Luk Van Ertvelde and Lieven Eeckhout. Workload reduction and generation techniques. IEEE Micro, 30(6): 57–65, November/December 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Veidenbaum:2004:GEI

[Vei04] Alex Veidenbaum. Guest Editor’s introduction: Applicationspecific processors. IEEE Micro, 24(3):8–9, May/June 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/csdl.computer.org/comp/
mags/mi/2004/03/m3008.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/03/m3008.
htm.

Venkataramani:2023:SIS

[Ven23] Guru Venkataramani. Special issue on security and privacypreserving execution environments. IEEE Micro, 43(5):6– 7, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vickers:1993:DAS

[Vic93] Richard Vickers. The development of ATM standards and technology: a retrospective. IEEE Micro, 13(6): 62–73, November/December 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Verleysen:1989:AVI

[VJ89] Michel Verleysen and Paul G. A. Jespers. An analog VLSI implementation of Hopfield’s neural network. IEEE Micro, 9(6):46–55, November/December 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Volos:2017:FCS

[VJFG17] Stavros Volos, Djordje Jevdjic, Babak Falsafi, and Boris Grot. Fat caches for scale-out servers. IEEE Micro, 37(2): 90–103, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/02/mmi2017020090-
abs.html.

Varma:2000:GEI

[VL00] Anujan Varma and Mark Laubach. Guest Editors’ introduction: Solving interconnection problems. IEEE Micro, 20(1):15–17, January/February 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2000/
pdf/m1015.pdf.

Vlahos:1988:GMD

[VM88] Harry Vlahos and Veljko Milutinovic. GaAs microprocessors and digital systems: an overview of R&D efforts. IEEE Micro, 8(1):28–56, January/February 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Valavala:1995:FPS

[VM95] K. Rao Valavala and M. A.

Manzoul. A fuzzy partitioning system. IEEE Micro, 15 (6):66, November/December 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

VanBulck:2019:BVM

[VMW^{+}19] J. Van Bulck, M. Minkin, O. Weisse, D. Genkin, B. Kasikci, F. Piessens, M. Silberstein, T. F. Wenisch, Y. Yarom, and R. Strackx. Breaking virtual memory protection and the SGX ecosystem with foreshadow. IEEE Micro, 39(3): 66–74, May/June 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vittoz:1996:GEI

[VN96] Eric A. Vittoz and Jean-Daniel D. Nicoud. Guest Editors’ introduction: Biologylnspired circuits. IEEE Micro, 16(5):10–11, September/October 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Valero:2010:MVE

[VN10] Mateo Valero and Nacho Navarro. Multicore: The view from Europe. IEEE Micro, 30 (5):2–4, September/October 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vassoler:2014:TDI

[VPRS14] Gilmar L. Vassoler, Marcia H. M. Paiva, Moises R. N. Ribeiro, and Marcelo E. V. Segatto. Twin datacenter interconnection topology. IEEE Micro, 34(5):8–17, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/05/mmi2014050008-
                                        abs.html.

Verdu:2012:PEC

[VPV12] Javier Verdu, Alex Pajuelo, and Mateo Valero. The problem of evaluating CPU– GPU systems with 3D visualization applications. IEEE Micro, 32(6):17–27, November/December 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vega:2020:LTH

[VRMC20] L. Vega, J. Roesch, J. McMahan, and L. Ceze. Last-Layer: Toward hardware and software continuous integration. IEEE Micro, 40(4):103– 111, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

VanderAuweraer:1987:FIA

[VS87] H. Van der Auweraer and R. Snoeys. FFT implementation alternatives in advanced measurement systems. IEEE Micro, 7(1):39–49, January/February 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Verleysen:1994:APA

[VTVM94] Michel Verleysen, Philippe Thissen, Jean-Luc Voz, and Jordi Madrenas. An analog processor architecture for a neural network classifier. IEEE Micro, 14(3):16– 28, May/June 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vidal-Verdu:1995:UBB

[VVRV95] Fernando Vidal-Verdu and Angel Rodriguez-Vazquez. Using building blocks to design analog neuro-fuzzy controllers. IEEE Micro, 15 (4):49–57, July/August 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

vanTilborg:1983:OSM

[vW83] Andre M. van Tilborg and Larry D. Wittie. Operating systems for the Micronet network computer. IEEE Micro, 3(2):38–47, March/April 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

vanderLinden:1985:IDE

[vW85] Frits van der Linden and Ian Wilson. An interactive debugging environment. IEEE Micro, 5(4):18–31, July/August 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Viredaz:2003:PEH

[VW03] Marc A. Viredaz and Deborah A. Wallach. Power evaluation of a handheld computer. IEEE Micro, 23(1):66– 74, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1066.
pdf;   http://www.computer.
org/micro/mi2003/m1066abs.
htm.

Vassiliadis:2003:MPP

[VWC03] Stamatis Vassiliadis, Stephan Wong, and Sorin Cotofana. Microcode processing: Positioning and directions. IEEE Micro, 23(4):21– 30, July/August 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/04/m4021abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2003/04/m4021.
pdf.

Watkins:2011:RRA

[WA11] Matthew A. Watkins and David H. Albonesi. ReMAP: a reconfigurable architecture for chip multiprocessors. IEEE Micro, 31(1):65–77, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wong:2013:SEP

[WA13] Daniel Wong and Murali Annavaram. Scaling the energy proportionality wall with KnightShift. IEEE Micro, 33(3):28–37, May/June 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wade:2020:TCT

[WAA^{+}20] M. Wade, E. Anderson, S. Ardalan, P. Bhargava, S. Buchbinder, M. L.Davenport, J. Fini, H. Lu, C. Li, R. Meade, C. Ramamurthy, M. Rust, F. Sedgwick, V. Stojanovic, D. Van Orden, C. Zhang, C. Sun, S. Y. Shumarayev, C. O’Keeffe, T. T. Hoang, D. Kehlet, R. V. Mahajan, M. T. Guzy, A. Chan, and T. Tran. TeraPHY: A chiplet technology for lowpower, high-bandwidth inpackage optical I/O. IEEE Micro, 40(2):63–71, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wesolowski:2021:DSA

[WAA^{+}21] Lukasz Wesolowski, Bilge Acun, Valentin Andrei, Adnan Aziz, Gisle Dankel, Christopher Gregg, Xiaoqiao Meng, Cyril Meurillon, Denis Sheahan, Lei Tian, Janet Yang, Peifeng Yu, and Kim Hazelwood. Datacenter-scale analysis and optimization of GPU machine learning workloads. IEEE Micro, 41(5): 101–112, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Walsh:1997:RSC

[Wal97] Dave Walsh. Reducing system cost with software modems: Accelerating product development, simplifying international approval, and ensuring upgradability. IEEE Micro, 17(4):37–43, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

                                        /pascal.computer.org/mi/
                                        books/mi1997/pdf/m4037.
                                        pdf.

Warren:1989:EB

[War89a] C. Warren. Establishing the basics. IEEE Micro, 9(6): 92–94, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1989:US

[War89b] C. Warren. Update on standards. IEEE Micro, 9(6):

95–97, November/December 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1990:BM

[War90a] C. Warren. Backplane measurements. IEEE Micro, 10 (4):75–77, July/August 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1990:BYA

[War90b] C. Warren. A busy year ahead. IEEE Micro, 10(1):85– 86, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1990:KUS

[War90c]  C. Warren.  Keeping up

with Uncle Sam. IEEE Micro, 10(5):72–73, September/October 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1990:RTM

[War90d] C. Warren. Realizing a transmission model. IEEE Micro, 10(3):76–79, May/June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1990:SCI

[War90e] C. Warren. The scalable coherent interface. IEEE Micro, 10(3):80–82, May/June 1990.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1990:SC

[War90f] C. Warren. A standard to consider. IEEE Micro, 10(6): 42–45, November/December 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1990:WWI

[War90g] C. Warren. Wire-to-wire interaction. IEEE Micro, 10 (1):87–88, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1991:BR

[War91a] C. Warren. A bountiful return. IEEE Micro, 11(4):40– 41, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1991:CI

[War91b] C. Warren. Computing interconnections. IEEE Micro, 11(6):84–86, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1991:ISP

[War91c] C. Warren. IEEE standard P1754 — an open microprocessor architecture. IEEE Micro, 11(5):30–33, September/October 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1991:SOB

[War91d] C. Warren. Sbus — an open bus architecture. IEEE Micro, 11(6):80–83, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1991:SCK

[War91e] C. Warren. Sorry, Captain Kirk. IEEE Micro, 11(2):34– 35, March/April 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1991:TSH

[War91f] C. Warren. There’s a standard hiding out there. IEEE Micro, 11(1):45–??, January/February 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1991:EES

[War91g] Carl Warren. On the edge — evaluating shielded twistedpair cable. IEEE Micro, 11 (1):46–47, January/February 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1992:FS

[War92a] C. Warren. Firmware standards. IEEE Micro, 12(1):73– 75, January/February 1992.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1992:POI

[War92b] C. Warren. PCMCIA — the other interface. IEEE Micro, 12(2):72–73, March/April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1992:PF

[War92c] C. Warren. A plan for the future. IEEE Micro, 12(3): 76–77, May/June 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1992:UAT

[War92d]  C. Warren.  Understand-

ing the acronym Tower-of-Babel. IEEE Micro, 12(1):69– 72, January/February 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wenisch:2012:EAC

[WB12] Thomas F. Wenisch and Alper Buyuktosunoglu. Energyaware computing. IEEE Micro, 32(5):6–8, September/October 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:1995:DMP

[WBC^{+}95] Karl Wang, Chris Bryant, Mike Carlson, Tom Elmer, Adrian Harris, Michael Garcia, C. S. Hui, C. K. Leung, Brian Reynolds, Raymond Tang, Laura Weber, Jim Wenzel, Glen Wilson, and Mike Becker. Designing the MPC105 PCI bridge/memory controller. IEEE Micro, 15 (2):44–49, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Welsh:1998:MMU

[WBHv98] Matt Welsh, Anindya Basu, Xun Wilson Huang, and Thorsten von Eicken. Memory management for user-level network interfaces. IEEE Micro, 18(2):77–82, March/ April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m2077.
pdf;   http://www.computer.
org/micro/mi1998/m2077abs.
htm.

Wu:2014:HPB

[WBKR14] Lisa Wu, Raymond J. Barker, Martha A. Kim, and Kenneth A. Ross. Hardware partitioning for big data analytics. IEEE Micro, 34(3):109– 119, May/June 2014. CODEN IEMIDZ. ISSN 0272-1732.

Washwell:1994:OCS

[WCH94] E. R. Washwell, G. O. Cheen, and C. H. Huang. Optical correlators for space applications: prospects and problems. IEEE Micro, 14(6):42–47, November/December 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2004:HTV

[WCW^{+}04] Perry H. Wang, Jamison D. Collins, Hong Wang, Dongkeun Kim, Bill Greene, Kai-Ming Chan, Aamir B. Yunus, Terry Sych, Stephen F. Moore, and John P. Shen. Helper threads via virtual multithreading. IEEE Micro, 24(6):74–82, November/December 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/mi/
                                        2004/06/m6074.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/mi/2004/06/m6074.pdf.

Wawrzynek:2003:GEI

[WD03] John Wawrzynek and Keith Diefendorff. Guest Editors’ introduction: Hot Chips 14— innovation in the face of uncertain economics. IEEE Micro, 23(2):8–11, March/April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m2008.
                                        pdf.

Whatmough:2020:CAR

[WDK^{+}20] P. N. Whatmough, M. Donato, G. G. Ko, S. K. Lee, D. Brooks, and G. Wei.

CHIPKIT: An agile, reusable open-source framework for rapid test chip development. IEEE Micro, 40(4):32–40, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:1993:IPI

[WE93] Chia-Jiu J. Wang and Frank Emnett. Implementing precise interruptions in pipelined RISC processors. IEEE Micro, 13(4):36–43, July/August 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Weatherford:1997:MVI

[Wea97a] M. Weatherford. Micro view:

IBM bets on JavaBeans. IEEE Micro, 17(3):80–??, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3080.
pdf.

Weatherford:1997:MVM

[Wea97b]  M. Weatherford.  Micro

view: MOSIS eyes the future. IEEE Micro, 17(2): 80–??, 79, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Webb:2008:IZN

[Web08] Charles F. Webb. IBM z10:

The next-generation mainframe microprocessor. IEEE Micro, 28(2):19–29, March/ April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Webb:2021:MAM

[Web21] Charles Webb. Microprocessor advances and the mainframe legacy. IEEE Micro, 41(6):68–70, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Weiser:2017:IEM

[Wei17] Uri Weiser. Insights from the 2016 Eckert–Mauchly Award recipient. IEEE Micro, 37(3): 126–128, May/June 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/03/mmi2017030126-
abs.html.

Weaver:2004:RSE

[WEMR04] Christopher T. Weaver, Joel Emer, Shubhendu S. Mukherjee, and Steven K. Reinhardt. Reducing the soft-error rate of a high-performance microprocessor. IEEE Micro, 24(6):30–37, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6030.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6030.pdf.

Wenisch:2018:TPC

[Wen18] Thomas F. Wenisch. Top picks from the 2017 computer architecture conferences. IEEE Micro, 38(3): 5–9, May/June 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/03/mmi2018030005.
html.

West:1989:MIW

[Wes89] L. F. West. Macintosh issues welcomed. IEEE Micro, 9(5): 7, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wettersten:1986:ATP

[Wet86] P. B. Wettersten. America’s trade-policy. IEEE Micro, 6(1):4–6, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wenisch:2010:MAC

[WFA^{+}10] Thomas F. Wenisch, Michael Ferdman, Anastasia Ailamaki, Babak Falsafi, and Andreas Moshovos. Making addresscorrelated prefetching practical. IEEE Micro, 30(1):50–59, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2021:HPM

[WFW^{+}21] Junbin Wang, Shaoxia Fang, Xi Wang, Jiangsha Ma, Taobo Wang, and Yi Shan. High-performance mixed-low-precision CNN inference accelerator on FPGA. IEEE Micro, 41(4):31–38, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warren:1992:TME

[WG92] C. Warren and J. Gafford. Tools to make the engineer’s life easier. IEEE Micro, 12(5): 66–68, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wilcke:1997:GEI

[WG97] Winfried Wilcke and Robert Garner. Guest Editors’ introduction: Celebrating chips and architectures. IEEE Micro, 17(2):9–10, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wilkerson:2009:TCC

[WGA^{+}09] Chris Wilkerson, Hongliang Gao, Alaa R. Alameldeen, Zeshan Chishti, Muhammad Khellah, and Shih-Lien Lu. Trading off cache capacity for low-voltage operation. IEEE Micro, 29(1):96–103, January/February 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wentzlaff:2007:CIA

[WGH^{+}07] David Wentzlaff, Patrick Griffin, Henry Hoffmann, Liewei Bao, Bruce Edwards, Carl Ramey, Matthew Mattina, Chyi-Chang Miao, John F. Brown III, and Anant Agarwal. On-chip interconnection architecture of the Tile Processor. IEEE Micro, 27(5):15–31, September/October 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wolinski:2002:PCF

[WGM02] Christophe Wolinski, Maya Gokhale, and Kevin McCabe. A polymorphous computing fabric. IEEE Micro, 22(5):56– 68, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5056.
pdf;   http://www.computer.
org/micro/mi2002/m5056abs.
htm.

Wassel:2014:NCP

[WGO^{+}14] Hassan M. G. Wassel, Ying Gao, Jason K. Oberg, Ted Huffmire, Ryan Kastner, Frederic T. Chong, and Timothy Sherwood. Networks on chip with provable security properties. IEEE Micro, 34(3):57–68, May/June 2014. CODEN IEMIDZ. ISSN 0272-

Wang:2009:MSH

[WH09] Feng Wang and Mounir Hamdi. Memory subsystems in high-end routers. IEEE Micro, 29(3):52–63, May/June 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Walker:1989:CNN

[WHA89] Mark Walker, Paul Hasler, and Lex Akers. A CMOS neural network for pattern association. IEEE Micro, 9(5):68– 74, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wharton:1997:LSR

[Wha97] J. Wharton. Letters: Setting the record straight. IEEE Micro, 17(2):2, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2018:HHP

[WHCK18] Xiebing Wang, Kai Huang, Long Chen, and Alois Knoll. h^{2}ECU: A high-performance and heterogeneous electronic control unit for automated driving. IEEE Micro, 38(5): 53–62, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/mi/2018/05/mmi2018050053-
                                        abs.html.

Wu:2023:BFP

[WHJ^{+}23] Guiming Wu, Qianwen He, Jiali Jiang, Zhenxiang Zhang, Yunfeng Shi, Xin Long, Linquan Jiang, Shuangchen Li, Yuan Xie, Changzheng Wei, Yuan Zhao, Ying Yan, Hui Zhang, and Yinchao Zou. E-Booster: a fieldprogrammable gate arraybased accelerator for secure tree boosting using additively homomorphic encryption. IEEE Micro, 43(5):88– 96, September/October 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

White:1993:HDPa

[WHKM93a] Steven W. White, Phil D.

Hester, Jack W. Kemp, and G. Jeanette McWilliams. How does processor MHz relate to end-user performance? part 1: Pipelines and functional units. IEEE Micro, 13(4):8– 16, July/August 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

White:1993:HDPb

[WHKM93b] Steven W. White, Phil D.

Hester, Jack W. Kemp, and G. Jeanette McWilliams. How does processor MHz relate to end-user performance? part 2: Memory subsystem and instruction set. IEEE Micro, 13(5):79–89, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wyglinski:2013:SAS

[WHP^{+}13] Alexander M. Wyglinski, Xinming Huang, Taskin Padir, Lifeng Lai, Thomas R. Eisenbarth, and Krishna Venkatasubramanian. Security of autonomous systems employing embedded computing and sensors. IEEE Micro, 33(1):80– 86, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Williams:1984:DRI

[Wil84] Theodore J. Williams. The development of reliability in industrial control systems. IEEE Micro, 4(6): 66–80, November/December 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Williams:1986:WNN

[Wil86] J. Williams. When is a von Neumann not a von Neumann? IEEE Micro, 6(1): 5–6, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wilson:1995:DSH

[Wil95a] J. Wilson. Data-security hits home. IEEE Micro, 15(5):86– 88, September/October 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wilson:1995:MVA

[Wil95b] Janet Wilson. Micro view:

Anyone who doesn’t design computer networks for security is providing irresponsible engineering. IEEE Micro, 15(5):88, 86–87, September/October 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wilson:1996:OV

[Wil96] J. Wilson. An outsider’s view. IEEE Micro, 16(5): 80, September/October 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wilson:1997:MVV

[Wil97] J. Wilson. Micro view: VSI alliance: Selfish interests the key? IEEE Micro, 17(1):80, January/February 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wilkes:2003:L

[Wil03] Maurice V. Wilkes. Letters.

IEEE Micro, 23(6):5, November/December 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2003/06/m6005.pdf;
http://csdl.computer.org/
dl/mags/mi/2003/06/m6005.
htm.

Wu:2005:FCT

[WJM^{+}05] Qiang Wu, Philo Juang, Margaret Martonosi, Li-Shiuan Peh, and Douglas W. Clark. Formal control techniques for power-performance management. IEEE Micro, 25(5):52– 62, September/October 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Williams:1988:OSF

[WJR88] Ronald D. Williams, Barry W. Johnson, and Thomas E. Roberts. An operating system for a fault-tolerant multiprocessor controller. IEEE Micro, 8(4):18–29, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2013:ITP

[WK13] Hao Wang and Nam Sung Kim. Improving throughput of power-constrained many-core processors based on unreliable devices. IEEE Micro, 33 (4):16–24, July/August 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2014:SOC

[WKK^{+}14] Xiaofei Wang, John Keane, Tony Tae-Hyoung Kim, Pulkit Jain, Qianying Tang, and Chris H. Kim. Silicon odometers: Compact in situ aging sensors for robust system design. IEEE Micro, 34(6):74–85, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060074-
abs.html.

Wittenbrink:2011:FGG

[WKP11] Craig M. Wittenbrink, Emmett Kilgariff, and Arjun Prabhu. Fermi GF100 GPU architecture. IEEE Micro, 31 (2):50–59, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

White:1992:ERT

[WL92] Lee White and Hareton K. N.

Leung. On the edge — regression testability. IEEE Micro, 12(2):81–84, March/April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2015:NRE

[WLD15] Siyang Wang, Alvin R.

Lebeck, and Chris Dwyer. Nanoscale resonance energy transfer-based devices for probabilistic computing. IEEE Micro, 35(5):72–84, September/October 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/05/mmi2015050072-
abs.html.

Woo:2008:PIB

[WLF^{+}08] Dong Hyuk Woo, Hsien-Hsin S. Lee, Joshua B. Fryman, Allan D. Knies, and Marsha Eng. POD: a 3Dintegrated broad-purpose acceleration layer. IEEE Micro, 28(4):28–40, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Weng:2022:USA

[WLKN22] Jian Weng, Sihao Liu, Dylan Kupsh, and Tony Nowatzki. Unifying spatial accelerator compilation with idiomatic and modular transformations. IEEE Micro, 42(5):59– 69, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wu:2015:QDP

[WLP^{+}15] Lisa Wu, Andrea Lottarini, Timothy K. Paine, Martha A. Kim, and Kenneth A. Ross. The Q100 database processing unit. IEEE Micro, 35(3): 34–46, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2015/03/mmi2015030034-
                                        abs.html.

Wang:2021:CMN

[WLY^{+}21a] Zixuan Wang, Xiao Liu, Jian Yang, Theodore Michailidis, Steven Swanson, and Jishen Zhao. Characterizing and modeling nonvolatile memory systems. IEEE Micro, 41(3): 63–70, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wu:2021:UUC

[WLY^{+}21b] Di Wu, Jingjie Li, Ruokai Yin, Hsuan Hsiao, Younghyun Kim, and Joshua San Miguel. uGEMM: Unary computing for GEMM applications. IEEE Micro, 41(3): 50–56, May/June 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wainwright:1985:MBM

[WM85] Richard E. Wainwright and Randy H. Moss. A microcomputerbased model robot system with pulse-width modulation control. IEEE Micro, 5(1): 7–21, January/February 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wu:2006:DCD

[WMC^{+}06] Qiang Wu, Margaret Martonosi, Douglas W. Clark, Vijay Janapa Reddi, Dan Connors, Youfeng Wu, Jin Lee, and David Brooks. Dynamiccompiler-driven control for microprocessor energy and performance. IEEE Micro, 26(1):119–129, January/February 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2010:TNP

[WMH^{+}10] Peng Wang, Dan Meng, Jizhong Han, Jianfeng Zhan, Bibo Tu, Xiaofeng Shi, and Le Wan. Transformer: a new paradigm for building dataparallel programming models. IEEE Micro, 30(4):55– 64, July/August 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Watts:2009:VPB

[WMSH09] Lloyd Watts, Dana Massie, Allen Sansano, and Jim Huey. Voice processors based on the human hearing system. IEEE Micro, 29(2):54–63, March/ April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Weider:1992:CTT

[WN92] Armin W. Weider and Franz Neppl. CMOS technology trends and economics. IEEE Micro, 12(4):10–19, July/August 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Weiler:1994:SGL

[WN94] B. Weiler and E. Nett. Speed-Log. A generic log service supporting efficient node-crash recovery. IEEE Micro, 14(5):60– 71, September/October 1994.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Watson:2016:FPD

[WNW^{+}16] Robert N. M. Watson, Robert M.

Norton, Jonathan Woodruff, Simon W. Moore, Peter G. Neumann, Jonathan Anderson, David Chisnall, Brooks Davis, Ben Laurie, Michael Roe, Nirav H. Dave, Khilan Gudka, Alexandre Joannou, A. Theodore Markettos, Ed Maste, Steven J. Murdoch, Colin Rothwell, Stacey D. Son, and Munraj Vadera. Fast protection-domain crossing in the CHERI capabilitysystem architecture. IEEE Micro, 36(5):38–49, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050038-
abs.html.

Welling:2001:CBA

[WOM01] Girish Welling, Maximilian Ott, and Saurabh Mathur. A cluster-based active router architecture. IEEE Micro, 21 (1):16–25, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1016.
pdf;   http://www.computer.
org/micro/mi2001/m1016abs.
htm.

Wong:2003:MTC

[Won03] Alfred K. Wong. Microlithography: Trends, challenges, solutions, and their impact on design. IEEE Micro, 23 (2):12–21, March/April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m2012.
                                        pdf;   http://www.computer.
                                        org/micro/mi2003/m2012abs.
                                        htm.

Wang:2023:HTP

[WPH^{+}23] Yingchen Wang, Riccardo Paccagnella, Elizabeth Tang He, Hovav Shacham, Christopher W. Fletcher, and David Kohlbrenner. Hertzbleed: Turning power side-channel attacks into remote timing attacks on x86. IEEE Micro, 43 (4):19–27, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2003:PMR

[WPM03] Hang-Sheng Wang, Li-Shiuan Peh, and Sharad Malik. A power model for routers: Modeling Alpha 21364 and InfiniBand routers. IEEE Micro, 23(1):26–35, January/February 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2003/
                                        pdf/m1026.pdf;   http://

www.computer.org/micro/
mi2003/m1026abs.htm.

Wawrzynek:2007:RRA

[WPO^{+}07] John Wawrzynek, David Patterson, Mark Oskin, Shih-Lien Lu, Christoforos Kozyrakis, James C. Hoe, Derek Chiou, and Krste Asanović. RAMP: Research accelerator for multiple processors. IEEE Micro, 27(2):46–57, March/April 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2020:LPL

[WPSR20] S. Wang, R. T. Possignolo, H. B. Skinner, and J. Renau. LiveHD: A productive live hardware development flow. IEEE Micro, 40 (4):67–75, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2014:RCU

[WRA^{+}14] Dan Wang, Aravindkumar Rajendiran, Sundaram Ananthanarayanan, Hiren Patel, Mahesh V. Tripunitara, and Siddharth Garg. Reliable computing with ultra-reduced instruction set coprocessors. IEEE Micro, 34(6):86–94, November/December 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/06/mmi2014060086-
abs.html.

Whitby-Strevens:1990:TPP

[WS90] Colin Whitby-Strevens. Transputers — past, present, and future. IEEE Micro, 10 (6):16–19, 76–82, November/December 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2013:IPW

[WS13] Liang Wang and Kevin Skadron. Implications of the power wall: Dim cores and reconfigurable logic. IEEE Micro, 33(5):40–48, September/October 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Woh:2010:AAA

[WSM^{+}10] Mark Woh, Sangwon Seo, Scott Mahlke, Trevor Mudge, Chaitali Chakrabarti, and Krisztián Flautner. AnySP: Anytime anywhere anyway signal processing. IEEE Micro, 30(1):81–91, January/February 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wu:2005:PET

[WSZS05] Weidong Wu, Jian Shi, Ling Zuo, and Bingxin Shi. Powerefficient TCAMS for bursty access patterns. IEEE Micro, 25(4):64–72, July/August 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wong:1998:GEI

[WT98] H. S.-Philip P. Wong and Albert J. P. Theuwissen. Guest Editors’ introduction: Digital imaging. IEEE Micro, 18(6): 12–13, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6012.
pdf.

Wu:2023:SIE

[Wu23] Carole-Jean Wu. Special issue on environmentally sustainable computing. IEEE Micro, 43(1):7–8, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Woudsma:1992:CAD

[Wv92] Rob Woudsma and Jef L.

van Meerbergen. Consumer applications: a driving force for high-level synthesis of signal-processing architectures. IEEE Micro, 12(4):20– 33, July/August 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wenisch:2006:SSS

[WWF^{+}06] Thomas F. Wenisch, Roland E.

Wunderlich, Michael Ferdman, Anastassia Ailamaki, Babak Falsafi, and James C. Hoe. SimFlex: Statistical sampling of computer system simulation. IEEE Micro, 26 (4):18–31, July/August 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wollrath:1997:JCD

[WWR97] Ann Wollrath, Jim Waldo, and Roger Riggs. Java-centric distributed computing: Providing a homogeneous view of a heterogeneous group of machines. IEEE Micro, 17 (3):44–53, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/pascal.computer.org/mi/
books/mi1997/pdf/m3044.
pdf.

Wen:2008:CMS

[WWZ^{+}08] Mei Wen, Nan Wu, Chunyuan Zhang, Qianming Yang, Jun Ren, Yi He, Wei Wu, Jun Chai, Maolin Guan, and Changqing Xun. On-chip memory system optimization design for the FT64 scientific stream accelerator. IEEE Micro, 28(4):51–70, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wang:2020:TNL

[WZL20] K. Wang, H. Zheng, and A. Louri. TSA-NoC: Learningbased threat detection and mitigation for secure networkon-chip architecture. IEEE Micro, 40(5):56–63, September/October 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xu:2007:HMR

[XBH07] Min Xu, Rastislav Bodík, and Mark D. Hill. A hardware memory race recorder for deterministic replay. IEEE Micro, 27(1):48–55, January/February 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xia:2021:KFN

[XCZ^{+}21] Jing Xia, Chuanning Cheng, Xiping Zhou, Yuxing Hu, and Peter Chun. Kunpeng 920: The first 7-nm chiplet-based 64-core ARM SoC for cloud services. IEEE Micro, 41(5): 67–75, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xie:2012:TIM

[XLW^{+}12] Min Xie, Yutong Lu, Kefei Wang, Lu Liu, Hongjia Cao, and Xuejun Yang. Tianhe-1A interconnect and messagepassing services. IEEE Micro, 32(1):8–20, January/February 2012. CODEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xin:2023:PUA

[XLX^{+}23] Yao Xin, Wenjun Li, Gaogang Xie, Yang Xu, and Yi Wang. A parallel and updatable architecture for FPGA-based packet classification with large-scale rule sets. IEEE Micro, 43(2):110– 119, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xie:2019:NSR

[XPZ^{+}19] M. Xie, C. Pan, Y. Zhang, J. Hu, Y. Liu, and C. Xue. A novel STT-RAM-based hybrid cache for intermittently powered processors in IoT devices. IEEE Micro, 39(1):24– 32, January/February 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xu:2009:DDR

[XWZ09] Jiang Xu, Wayne Wolf, and Wei Zhang. Double-data-rate, wave-pipelined interconnect for asynchronous NoCs. IEEE Micro, 29(3):20–30, May/ June 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xia:2002:DCS

[XYCS02] Fei Xia, Alex V. Yakovlev, Ian G. Clark, and Delong Shang. Data communication in systems with heterogeneous timing. IEEE Micro, 22(6): 58–69, November/December 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m6058.
pdf;   http://www.computer.

org/micro/mi2002/m6058abs.
htm.

Xu:2023:TDH

[XYT^{+}23] Yinan Xu, Zihao Yu, Dan Tang, Ye Cai, Dandan Huan, Wei He, Ninghui Sun, and Yungang Bao. Toward developing high-performance RISC-V processors using agile methodology. IEEE Micro, 43 (4):98–106, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Xie:2019:EMT

[XZ19] Y. Xie and J. Zhao. Emerging memory technologies. IEEE Micro, 39(1):6–7, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yesil:2018:TDP

[YAK18] Serif Yesil, Ismail Akturk, and Ulya R. Karpuzcu. Toward dynamic precision scaling. IEEE Micro, 38(4):30– 39, July/August 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/04/mmi2018040030-
abs.html.

Yao:1985:CAP

[Yao85]  Neng Fred Yao.  The

computer-aided programming system — a friendly programming environment. IEEE Micro, 5(2):9–19, March/April 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yazdanbakhsh:2015:CCF

[YBNS15] Amir Yazdanbakhsh, Raghuraman Balasubramanian, Tony Nowatzki, and Karthikeyan Sankaralingam. Comprehensive circuit failure prediction for logic and SRAM using virtual aging. IEEE Micro, 35(6):24–36, November/December 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
mi/2015/06/mmi2015060024-
abs.html.

Yang:2017:HDS

[YBS17] Kaiyuan Yang, David Blaauw, and Dennis Sylvester. Hardware designs for security in ultra-low-power IoT systems: An overview and survey. IEEE Micro, 37(6):72–89, November/December 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/06/mmi2017060072-
abs.html.

Yuzuguler:2019:ANN

[YCD^{+}19] A. C. Yüzügüler, F. Celik, M. Drumond, B. Falsafi, and P. Frossard. Analog neural networks with deep-submicrometer nonlinear synapses. IEEE Micro, 39(5): 55–63, September/October 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yoon:2011:VEF

[YE11] Doe Hyun Yoon and Mattan Erez. Virtualized ECC: Flexible reliability in main memory. IEEE Micro, 31(1):11–19, January/February 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yeager:1996:MRS

[Yea96] Kenneth C. Yeager. The MIPS R10000 superscalar microprocessor: Emphasizing concurrency and latencyhiding techniques to efficiently run large, real-world applications. IEEE Micro, 16(2):28– 41, March/April 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Yeh:2007:LPH

[Yeh07]  Tse-Yu Yeh.  Low-power,

high-performance architecture of the PWRficient processor family. IEEE Micro, 27(2):69– 78, March/April 2007. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yao:2019:LCM

[YFDV19] F. Yao, H. Fang, M. Doroslovački, and G. Venkataramani. Leveraging cache management hardware for practical defense against cache timing channel attacks. IEEE Micro, 39(4): 8–16, July/August 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yu:2020:CFS

[YHHF20] J. Yu, L. Hsiung, M. E. Hajj, and C. W. Fletcher. Creating foundations for secure microarchitectures with data-oblivious ISA extensions. IEEE Micro, 40(3):99– 107, May/June 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yoshida:2015:SXF

[YHT^{+}15] Toshio Yoshida, Mikio Hondou, Takekazu Tabata, Ryuji Kan, Naohiro Kiyota, Hiroyuki Kojima, Koji Hosoe, and Hiroshi Okano. Sparc64 XIfx: Fujitsu’s next-generation processor for high-performance computing. IEEE Micro, 35 (2):6–14, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/02/mmi2015020006-
abs.html.

Yi:2021:AHP

[Yi21a] Joshua J. Yi. Analysis of historical patenting behavior and patent characteristics of computer architecture companies. IEEE Micro, 41(5):

114–123, September/October 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2021:MPT

[Yi21b] Joshua J. Yi. Microarchitecture patents over time and interesting early microarchitecture patents. IEEE Micro, 41(6):172–178, November/December 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2021:RPL

[Yi21c] Joshua J. Yi. Recent patents for leading computer architecture companies. IEEE Micro, 41(4):74–77, July/August 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2022:AHPa

[Yi22a] Joshua J. Yi. Analysis of historical patenting behavior and patent characteristics of computer architecture companies. Part II: Prosecution time and effective patent term length. IEEE Micro, 42(1):128–136, January/February 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2022:AHPb

[Yi22b]  Joshua J. Yi.  Analysis

of historical patenting behavior and patent characteristics of computer architecture companies. Part III: Claims. IEEE Micro, 42(4): 124–132, July/August 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2022:AHPc

[Yi22c]  Joshua J. Yi.  Analysis

of historical patenting behavior and patent characteristics of computer architecture companies. Part IV: Claims. IEEE Micro, 42(5): 119–127, September/October 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2022:AHPd

[Yi22d] Joshua J. Yi. Analysis of historical patenting behavior and patent characteristics of computer architecture companies. Part V: References. IEEE Micro, 42(6):135–140, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2022:RPIb

[Yi22e] Joshua J. Yi. Review of patents issued to computer architecture companies in 2021 — Part II. IEEE Micro, 42(3): 67–77, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2022:RPIa

[Yi22f] Joshua J. Yi. Review of patents issued to computer architecture companies in 2021 [micro law]. IEEE Micro, 42 (2):77–84, March/April 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2023:AHPa

[Yi23a] Joshua J. Yi. Analysis of historical patenting behavior and patent characteristics of computer architecture companies — Part VI: Relationship between prosecution time and claims. IEEE Micro, 43(4): 119–123, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2023:AHPb

[Yi23b]  Joshua J. Yi.  Analysis

of historical patenting behavior and patent characteristics of computer architecture companies. Part VII: Relationship between prosecution time and claims. IEEE Micro, 43(6):103–108, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2023:DAR

[Yi23c] Joshua J. Yi. Does academic research drive industrial innovation in computer architecture? analyzing citations to academic papers in patents. IEEE Micro, 43 (1):83–88, January/February 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2024:AHPb

[Yi24a] Joshua J. Yi. Analysis of historical patenting behavior and patent characteristics of computer architecture companies. Part IX: Patent families. IEEE Micro, 44(2):72– 77, March/April 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yi:2024:AHPa

[Yi24b] Joshua J. Yi. Analysis of historical patenting behavior and patent characteristics of computer architecture companies part VIII: Patent families. IEEE Micro, 44(1):70–74, January/February 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yavits:2018:EFA

[YKG18] Leonid Yavits, Roman Kaplan, and Ran Ginosar. Enabling full associativity with memristive address decoder. IEEE Micro, 38(5):32– 40, September/October 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/05/mmi2018050032-
abs.html.

Yin:2019:MIR

[YKH^{+}19] S. Yin, Y. Kim, X. Han, H. Barnaby, S. Yu, Y. Luo, W. He, X. Sun, J. Kim, and J. Seo. Monolithically integrated RRAM- and CMOS-based in-memory computing optimizations for efficient deep learning. IEEE Micro, 39(6):54–63, November/December 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yu:2005:EMP

[YKL05] Fang Yu, Randy H. Katz, and T. V. Lakshman. Efficient multimatch packet classification and lookup with TCAM. IEEE Micro, 25(1):50–59, January/February 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1050.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1050.pdf.

Yoshida:2013:SXF

[YMA^{+}13] Toshio Yoshida, Takumi Maruyama, Yasunobu Akizuki, Ryuji Kan, Naohiro Kiyota, Kiyoshi Ikenishi, Shigeki Itou, Tomoyuki Watahiki, and Hiroshi Okano. Sparc64 X: Fujitsu’s newgeneration 16-core processor for Unix servers. IEEE Micro, 33(6):16–24, November/ December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Yoon:2012:FPP

[YMC^{+}12] Doe Hyun Yoon, Naveen Muralimanohar, Jichuan Chang, Parthasarathy Ranganathan, Norman P. Jouppi, and Mattan Erez. Free-p: a practical end-to-end nonvolatile memory protection mechanism. IEEE Micro, 32(3): 79–87, May/June 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yao:2014:FST

[YNS^{+}14] Jun Yao, Yasuhiko Nakashima, Mitsutoshi Saito, Yohei Hazama, and Ryosuke Yamanaka. A flexible, self-tuning, faulttolerant functional unit array processor. IEEE Micro, 34(6):54–63, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        mi/2014/06/mmi2014060054-
                                        abs.html.

Young:2021:AAM

[You21] Cliff Young. Atari’s AN-TIC: My favorite microprocessor. IEEE Micro, 41 (6):161, November/December 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yang:2022:TDS

[YRC^{+}22] Lita Yang, Robert M. Radway, Yu-Hsin Chen, Tony F. Wu, Huichu Liu, Elnaz Ansari, Vikas Chandra, Subhasish Mitra, and Edith Beigné. Three-dimensional stacked neural network accelerator architectures for AR/ VR applications. IEEE Micro, 42(6):116–124, November/December 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yoshida:1991:GBM

[YSMH91] Toyohiko Yoshida, Toru Shimizu, Shigeo Mizugaki, and Junichi Hinata. The Gmicro/100 32-bit microprocessor. IEEE Micro, 11(4):20–23, 62– 72, July/August 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yee:2001:GEI

[YT01] James Yee and Manu Thapar. Guest Editors’ introduction: Red hot chili papers. IEEE Micro, 21(1):14– 15, January/February 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m1014.
pdf;   http://www.computer.
org/micro/mi2001/m1014abs.
htm.

Yu:1998:PSO

[YTR^{+}98] B. Y. Yu, P. Toliver, R. J.

Runser, K. L. Deng, D. Y. Zhou, I. Glesk, and P. R. Prucnal. Packet-switched optical networks. IEEE Micro, 18(1):28–38, January/February 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yu:1996:FMI

[Yu96] Albert Y. C. Yu. The future of microprocessors — Intel’s head of microprocessor products looks 10 years ahead to 2006. IEEE Micro, 16(6): 46–53, November/December 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yun:2001:TMS

[Yun01] Kenneth Y. Yun. A terabit multiservice switch. IEEE Micro, 21(1):58–70, January/February 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/mi/books/mi2001/
                                        pdf/m1058.pdf;   http://
                                        www.computer.org/micro/
                                        mi2001/m1058abs.htm.

Yates:1988:FTM

[YW88] Steven W. Yates and Ronald D. Williams. A fault-tolerant multiprocessor controller for magnetic bearings. IEEE Micro, 8(4):6–17, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yang:1994:MFT

[YW94] Chu-Sing S. Yang and Shun-Yue Y. Wu. Modular faulttolerant Boolean N-cubes. IEEE Micro, 14(4):68–77, July/August 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Yeung:1998:DWS

[YYH98] Minerva M. Yeung, Boon-Lock L. Yeo, and Matthew Holliman. Digital watermarks — shedding light on the invisible. IEEE Micro, 18(6): 32–41, November/December 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi1998/pdf/m6032.
pdf;   http://www.computer.
org/micro/mi1998/m6032abs.
htm.

Yu:2020:STT

[YYK^{+}20] J. Yu, M. Yan, A. Khyzha, A. Morrison, J. Torrellas, and C. W. Fletcher. Speculative Taint Tracking (STT): A comprehensive protection for speculatively accessed data. IEEE Micro, 40(3):81–90, May/ June 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

You:2023:EET

[YZW^{+}23] Haoran You, Yang Zhao, Cheng Wan, Zhongzhi Yu, Yonggan Fu, Jiayi Yuan, Shang Wu, Shunyao Zhang, Yongan Zhang, Chaojian Li, Vivek Boominathan, Ashok Veeraraghavan, Ziyun Li, and Yingyan Celine Lin. Eye-CoD: Eye tracking system acceleration via FlatCam-based algorithm and hardware codesign. IEEE Micro, 43(4):88– 97, July/August 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zilberman:2014:NST

[ZACM14] Noa Zilberman, Yury Audzevich, G. Adam Covington, and Andrew W. Moore. NetFPGA SUME: Toward 100 Gbps as research commodity. IEEE Micro, 34(5):32– 41, September/October 2014. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050032-
abs.html.

Zhang:2015:PDF

[ZBES15] Meng Zhang, Jesse D. Bingham, John Erickson, and Daniel J. Sorin. PVCoherence: Designing flat coherence protocols for scalable verification. IEEE Micro, 35(3): 84–91, May/June 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030084-
abs.html.

Zhang:2000:SAP

[ZBH^{+}00] L. Louis Zhang, Brent Beacham, Massoud Reza Hashemi, Paul Chow, and Alberto Leon-Garcia. A scheduler ASIC for a programmable packet switch. IEEE Micro, 20 (1):42–48, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m1042.
pdf;   http://www.computer.
org/micro/mi2000/m1042abs.
htm.

Zang:2014:DNS

[ZCW^{+}14] Dawei Zang, Zheng Cao, Zhan Wang, Xiaoli Liu, Lin Wang, and Ninghui Sun. Decentralized NIC-switching architecture using SR-IOV PCI Express network device. IEEE Micro, 34(5):42–50, September/October 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2014/05/mmi2014050042-
abs.html.

Zahir:2013:MSI

[ZES13] Rumi Zahir, Mark Ewert, and Hari Seshadri. The Medfield smartphone: Intel architecture in a handheld form factor. IEEE Micro, 33(6): 38–46, November/December 2013. CODEN IEMIDZ. ISSN 0272-1732.

Zhang:2023:NET

[ZFW^{+}23] Shunyao Zhang, Yonggan Fu, Shang Wu, Jyotikrishna Dass, Haoran You, and Yingyan Lin. NetDistiller: Empowering tiny deep learning via in situ distillation. IEEE Micro, 43(6):84–92, November/December 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zuquete:1996:TAC

[ZG96] Andre Zuquete and Paulo Guedes. Transparent authentication and confidentiality for stream sockets: Ensuring private network communications for Unix and Windows systems. IEEE Micro, 16(3): 34–41, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:1991:C

[Zha91a] X. D. Zhang. Correction. IEEE Micro, 11(3):6, May/ June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:1991:SEI

[Zha91b] Xiaodong D. Zhang. System effects of interprocessor communication latency in multicomputers. IEEE Micro, 11(2):12–15, 52–55, March/ April 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zu:2017:TSP

[ZHPR17] Yazhou Zu, Wei Huang, Indrani Paul, and Vijay Janapa Reddi. Ti-states: Power management in active timing margin processors. IEEE Micro, 37(3):106–114, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https:

//www.computer.org/csdl/
mags/mi/2017/03/mmi2017030106-
abs.html.

Zhu:2015:RCE

[ZHR15] Yuhao Zhu, Matthew Halpern, and Vijay Janapa Reddi. The role of the CPU in energyefficient mobile Web browsing. IEEE Micro, 35(1):26–33, January/February 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/01/mmi2015010026-
abs.html.

Zuo:2019:WDH

[ZHZ^{+}19] P. Zuo, Y. Hua, M. Zhao, W. Zhou, and Y. Guo. Write deduplication and hash mode encryption for secure nonvolatile main memory. IEEE Micro, 39(1):44–51, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhao:2007:ELS

[ZIM^{+}07] Li Zhao, Ravi Iyer, Jaideep Moses, Ramesh Illikkal, Srihari Makineni, and Don Newell. Exploring largescale CMP architectures using ManySim. IEEE Micro, 27 (4):21–33, July/August 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:2023:OCL

[ZKP^{+}23] Yuxuan Zhang, Tanvir Ahmed Khan, Gilles Pokam, Baris Kasikci, Heiner Litz, and Joseph Devietti. Online code layout optimizations via OCOLOS. IEEE Micro, 43 (4):71–79, July/August 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zahedi:2015:SIF

[ZL15] Seyed Majid Zahedi and Benjamin C. Lee. Sharing incentives and fair division for multiprocessors. IEEE Micro, 35 (3):92–100, May/June 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
mi/2015/03/mmi2015030092-
abs.html.

Zhang:2016:MAV

[ZL16] Tianwei Zhang and Ruby B. Lee. Monitoring and attestation of virtual machine security health in cloud computing. IEEE Micro, 36(5):28– 37, September/October 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2016/05/mmi2016050028-
abs.html.

Zhao:2006:NPB

[ZLBI06] Li Zhao, Yan Luo, Laxmi N.

Bhuyan, and Ravi Iyer. A network processor-based, content-aware switch. IEEE Micro, 26(3):72–84, May/ June 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zortman:2013:BER

[ZLTW13] William A. Zortman, Anthony L. Lentine, Douglas C. Trotter, and Michael R. Watts. Bit-error-rate monitoring for active wavelength control of resonant modulators. IEEE Micro, 33(1):42– 52, January/February 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zsombor-Murray:1983:BDBb

[ZMVH^{+}83a] Paul J. A. Zsombor-Murray, Louis J. Vroomen, Robert D. Hudson, Peter H. Holck, and Tho Le-Ngoc. Binarydecision-based programmable controllers. II. IEEE Micro, 3(5):16–26, September/October 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zsombor-Murray:1983:BDBc

[ZMVH^{+}83b] Paul J. A. Zsombor-Murray, Louis J. Vroomen, Robert D. Hudson, Peter H. Holck, and Tho Le-Ngoc. Binarydecision-based programmable controllers. III. IEEE Micro, 3(6):24–39, November/December 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zsombor-Murray:1983:BDBa

[ZMVH^{+}83c] Paul J. A. Zsombor-Murray, Louis J. Vroomen, Robert D. Hudson, Tho Le-Ngoc, and Peter H. Holck. Binarydecision-based programmable controllers. Part I. IEEE Micro, 3(4):67–78, 81–83, July/August 1983. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zanoni:1993:IRS

[ZP93] E. Zanoni and P. Pavan. Improving the reliability and safety of automotive electronics. IEEE Micro, 13(1):30–48, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhou:2004:ISG

[ZQL^{+}04] Pin Zhou, Feng Qin, Wei Liu, Yuanyuan Zhou, and Josep Torrellas. iWatcher: Simple, general architectural support for software debugging. IEEE Micro, 24(6):50–56, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6050.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6050.pdf.

Zhu:2017:CCS

[ZRA^{+}17] Yuhao Zhu, Vijay Janapa Reddi, Robert Adolf, Saketh Rama, Brandon Reagen, Gu-Yeon Wei, and David Brooks. Cognitive computing safety: The new horizon for reliability/the design and evolution of deep learning workloads. IEEE Micro, 37(1):15– 21, January/February 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2017/01/mmi2017010015-
abs.html.

Zhou:2020:SSG

[ZRA^{+}20] Y. Zhou, S. Roy, A. Abdolrashidi, D. L. Wong, P. Ma, Q. Xu, A. Mirhoseini, and J. Laudon. A single-shot generalized device placement for large dataflow graphs. IEEE Micro, 40(5):26–36, September/October 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhou:2022:MHC

[ZRB^{+}22] Chuteng Zhou, Fernando García Redondo, Julian Büchel, Irem Boybat, Xavier Timoneda Comas, S. R. Nandakumar, Shidhartha Das, Abu Sebastian, Manuel Le Gallo, and Paul N. Whatmough. ML-HW codesign of noise-robust TinyML models and always-on analog compute-in-memory edge accelerator. IEEE Micro, 42(6): 76–87, November/December 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zaruba:2021:MCR

[ZSB21] F. Zaruba, F. Schuiki, and L. Benini. Manticore: A 4096core RISC-V chiplet architecture for ultraefficient floatingpoint computing. IEEE Micro, 41(2):36–42, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zschau:1984:NPM

[Zsc84]  E. Zschau.  A national

policy to maintain United States technological leadership. IEEE Micro, 4(1):12–14, January/February 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:2019:RAC

[ZSS^{+}19] Z. Zhang, X. Si, S. Srinivasa, A. K. Ramanathan, and M. Chang. Recent advances in compute-in-memory support for SRAM using monolithic 3-D integration. IEEE Micro, 39(6):28–37, November/December 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:2018:DPA

[ZUNN18] Renyuan Zhang, Noriyuki Uetake, Takashi Nakada, and Yasuhiko Nakashima. Design of programmable analog calculation unit by implementing support vector regression for approximate computing. IEEE Micro, 38(6):73–82, November/December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
mi/2018/06/08486976-abs.
html.

Zsombormurray:1985:LCDa

[ZV85] P. J. Zsombormurray and L. J.

Vroomen. Logical choice and the discriminating designer. IEEE Micro, 5(3): 63–68, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zsombormurray:1985:LCDb

[ZVH85] P. J. Zsombormurray, L. J.

Vroomen, and R. Hudson. Logical choice and the discriminating designer — reply. IEEE Micro, 5(3):71– 72, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zsombormurray:1985:CBD

[ZVHL85] P. J. Zsombormurray, L. J.

Vroomen, R. Hudson, and T. Lengoc. Comments on binary-decision-based programmable controllers — reply. IEEE Micro, 5(3):60– 63, May/June 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:2019:CBB

[ZWB19] S. Zhang, A. Wright, and T. Bourgeat. Composable building blocks to open up processor design. IEEE Micro, 39(3):47–55, May/June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhu:2002:AMP

[ZZ02] Zhichun Zhu and Xiaodong Zhang. Access-mode predictions for low-power cache design. IEEE Micro, 22 (2):58–71, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m2058.
pdf;   http://www.computer.
org/micro/mi2002/m2058abs.
htm.

Zhu:2005:LAA

[ZZ05] Zhichun Zhu and Xiaodong Zhang. Look-ahead architecture adaptation to reduce processor power consumption. IEEE Micro, 25(4):10– 19, July/August 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:2023:PNW

[ZZNT^{+}23] Shiqing Zhang, Ziyue Zhang, Mahmood Naderan-Tahan, Hossein SeyyedAghaei, Xin Wang, He Li, Senbiao Qin, Didier Colle, Guy Torfs, Mario Pickavet, Johan Bauwelinck, Günther Roelkens, and Lieven Eeckhout. Photonic networkon-wafer for multichiplet GPUs. IEEE Micro, 43(2):86– 95, March/April 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zhang:1997:TFH

[ZZY97] Chenxi X. Zhang, Xiaodong D.

Zhang, and Yong Yan. Two fast and high-associativity cache schemes. IEEE Micro, 17(5):40–49, September/October 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi1997/
pdf/m5040.pdf;   http://
www.computer.org/micro/
mi1997/m5040abs.htm.